A Bibliography of Publications about the GNU (Gnu is Not Unix) System

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

17 November 2023 Version 3.259

Title word cross-reference

#3 [Bon11]. #35 [Rog09a]. #37 [Rog09b]. #52 [Bri09a]. #54 [Bri09b]. #56 [Och09]. #81 [Rog11]. #95 [Och12].

(zI − H)x = b [HKY^{+}21]. 1 + 118 [APHV19]. 2 [ZK21]. $22.96 [Cas02]. $24.95 [Ano99a]. 3 [Ano01j, CZS^{+}21, MGYC18, SDeaK^{+}09, Wen02]. $34.95 [Ano00a, Ano00b]. $39.95 [Ano97a]. $39.99 [Kuc06]. 4 [DO16]. $49.99 [Fox08]. 5 [MGPB20]. $95.00 [Aji17]. ^{x}3 [KGW^{+}21]. hp [CMC^{+}15]. i [HWM^{+}15]. Kω [HKY^{+}21]. μ [TACA15]. N [HPT17, PHT17].

*BSD [Den99].

-Compiler [PKP02, PKP05, PKP05]. -D [DO16, SDeaK^{+}09, Wen02]. -diff [TACA15]. -electron [HPT17, PHT17]. -programming [KORP95]. -v3 [Car04].

//sagemath.org [Den13]. /GNOME [Wri00]. /Octave [MBR21]. /Should [HR11].

0 [Fox08]. 0-13-234971-X [Fox08]. 0-596-00287-4 [Cas02]. ’00s [Zad02]. 01 [DWP^{+}14]. 02 [KY16]. 03 [Zha16]. 06 [BVLF14].

1 [Kuc06]. 1-4 [Ano06]. 1-59059-503-3 [Kuc06]. 1-8 [AFS82]. 1.0 [KPK^{+}17, RHR^{+}21, Weh03]. 1.1 [Ano01a]. 10 [Uni77]. 10.0 [Bau06b]. 10x [Ano04b]. 11 [Dig80a, Dig80b, PH82]. 118 [PMG^{+}09]. 11th [ACM93a, ACM94, Kap92]. 12th [Bun94]. 131-3 [AM18]. 141 [PKG^{+}10]. 158 [GNR^{+}09]. 166MHz [Ano96c]. 18 [BJJ14]. 1970 [MSLH71]. 1983 [Neu84]. 19A [DWP^{+}14]. 1st [FFvdH01]. 1UAXe [Ano00j].

2 [Ano15a, BB91, CCG^{+}02, Dye03, EKJ^{+}03, Goe07, Kro00, Mau05, NRG^{+}99, War04]. 2.0 [Ano04b, Coc01b, DCS05, EHP14]. 2.2 [Bra04]. 2.3.x [LS04]. 2.6 [BS05]. 2.x [LSM^{+}00]. 20 [Ano97c]. 2000 [Ste00b]. 2000/3000 [ZKCS91]. 2005 [MSK05]. 2009 [Hea09]. 2010 [Gre11b]. 2012 [Rui13, SP12]. 2013 [BK14]. 2014 [HMW15]. 2017 [BBdD17]. 20th [IEE90]. 213 [AZ17a]. 21st [DMP^{+}02, IEE05, Jef08, Har05]. 22-24 [IEE93]. 23rd [DMP^{+}02]. 242 [CFCA13a]. 24th [BBdD17]. 253 [DKMB14]. 2597-31 [MSS95]. 26-28 [IEE90]. 264DP [Ano98]. 28-nm [CCA^{+}19]. 2B [Gom99]. 2e [Str94]. 2nd [HDR04].

3 [Ahm08b, Kuc06, Wes00]. 3-D [Wes00]. 3.0 [GHM^{+}05, HMO^{+}18, MS00, Váz16]. 3.1 [Bau06a, HHV05]. 3.8 [EaoGOBHW14]. 32 [TG99]. 33rd [ACM95]. 390 [Pen03]. 393 [WKA^{+}08]. 3D [DDJ99]. 3DLDF [Fin22a, Fin22b].

4 [G^{+}06]. 4.0 [Rob11]. 42 [GKL^{+}14]. 465 [TBPS15]. 4G [CTP^{+}22]. 4th [EKR91, HY14, HDR04, PT91, USE00a, FFHL05].

5 [Ano97d, Bra97, Kro00]. 5-8 [ACM93a]. 500 [Pra03]. 50th [SHB^{+}20]. 511-Core [DXT^{+}18]. 56th [Bon93]. 570 [PBJ^{+}12]. 5G [LLEL^{+}23]. 5th [AK95, FFH^{+}05, USE01a].

6 [Ano04b]. 6.0 [Ano01a, Gra99]. 64-bit [Ano00h, GHL^{+}04, Jae03]. 6th [Lio96].

7 [Kro99a]. 746 [Sta78a]. 75 [KMF^{+}07]. 7th [Yuk94].

8 [Ano03d, Ano11]. 816 [PPG^{+}11]. 852 [GB06]. ’87 [BSK87]. 871 [SC08]. ’88 [SMNF88].

9 [Ano95f, Ano95g, Sta12]. ’90 [Mio90, Ano96e, Ano97d, Bra97]. ’91 [PT91].

’92 [Ano94a, IEE92d, VW92, Vor92, VRS^{+}99a, VRS^{+}99b]. 92-07 [VW92]. 92-19 [AY93]. ’93 [ACM93a, ACM93b, Ano94a, Bao93, Bon93, Lei93a, SS93]. ’94 [BGG^{+}94]. ’95 [AK95, IEE95b, Lev95a, Bro03, FK99, Smy97, VGD^{+}97]. 97 [Laz98]. 978 [Ano11, Ano15a]. 978-0-12-802916-9 [SD16]. 978-1-4422-4992-9 [Aji17]. 978-1-59327-649-2 [Ano15a]. 978-1-59749-627-8 [Ano11]. 99 [PKP05]. 9A [BVLF14]. 9X [BG95, KTP95].

Abaco [Ano01j]. Abandoned [KCAS23]. Abandonment [LMPT22, KC22]. Abbotsbrook [Ano00k]. abgeschwächte [NO03]. ABI [Tro04]. AbiSource [Knu99a]. AbiWord [Knu99a]. Absoft [Ano96b]. Absorption [HW17a]. Abstraction [CSD^{+}05, BR95]. AC [CD95]. Academia [Rob20]. academic [AKF21]. academy [MTBS09]. Accelerated [Ano96c, Ano97c, SAC^{+}15, TL17]. Accelerated-X [Ano97c]. accelerates [Ped05]. Accelerator [DXT^{+}18, GCE^{+}21, KY16, LGW^{+}22, PGW^{+}20, CCA^{+}13]. Accelerators [dlCKK15, HXS20]. Acceptable [Sha10]. Access [Sta04a, Bow05, CD95, PDG^{+}87, Sta96b, WK93]. accessibility [Aji17, Fri16, HBC^{+}05]. Accessing [Tan11b]. Accidental [Ray99b, Ray99c, Ray01b]. Accomplishments [MN04, SZAB98]. Accurate [XXCL19, HR94, SC16]. ACFlow [Hua23]. Achievement [Coc01a]. Achieving [Abe07]. Acid [Lew99b]. ACLs [BS98]. ACM [Abr81, Bar00b, Bar00a, DGBH93, FP95, FMA02]. ACM/SIGAPP [DGBH93]. acne [DSB^{+}16]. acoustic [HKvH16]. acquisition [Mei92, MGFRG12]. across [CHE^{+}10, Fri97, PAB^{+}17]. Act [MSZ02]. Action [Jan08, MMD12, NR03, RCP^{+}12]. activation [DARJ23]. ActiveX [Kro99b]. activism [Mau05]. Activities [Ave06, BY14, LL14, LC12b, WFW^{+}20, AKHG16, SSA08, VGP^{+}19, ZE00]. activity [CF09, RCGB^{+}22]. actor [KF17]. actor-oriented [KF17]. Acumen [Kro99b]. ad [SH11]. ADA [ACM94, ACM93b, Ano87, GB94, Smy97, ACM94, BG95, BOM97, BCHR12, FK99, Kan12, KTP95, Kle21, MGM^{+}02, MSK05, Och09, RAH^{+}01, RTH15, Rui13, SP12, Smy97, VGD^{+}97]. Ada83 [Fel93]. ADA94 [CGS94]. Ada95 [Gar09, Due97]. Ada95/C [Gar09]. Ada’97 [ACM97]. Ada9x [Fel93]. AdaCore [Bro19]. adapted [WPAV14]. Adapting [Man92, YM93]. Adaptive [Joh18, CV22, SA15]. Add [Bar01]. Adding [SZAB97, Ano03d, CLL05]. Address [CDsJ^{+}00, WCG22]. Addressing [Sha04, ZWH21]. Adds [Ano00j, Sur04]. Admin [Plo97]. Administration [Ano00e, Ano01j, G^{+}00, G^{+}02, GA04b, Har94, Kre03, USE94, USE98b, Ron05b, TB05]. Administrations [SC02]. Administrator [Ano90c, Mag04, DRP01]. Administrators [FT09, SHN97]. Admitted [PLO^{+}23, YXS^{+}19, HSX^{+}18, ZFD21]. Adobe [Ano02b]. ADOC [KG20]. Adopt [MSC19, MFS15, VVM08, Ano00g]. Adopter [RNR17]. Adopters [Goo14]. Adopting [ACC^{+}12, GHM^{+}05, SF15]. Adoption [ACHC11, DD17, Muw09, WW01, AW07, BGL^{+}22, CM06, KKA^{+}19, eLAA^{+}23, NYB10, NDDH^{+}21, PdSCJM22, RH21, SG12]. Adopts [GGB17]. Advance [RLTD23]. Advanced [Ano88b, MYU89, Ron05a, Wes00, Bar00c]. Advancement [Sca19, Spi06]. Advances [DDJ98a, DDJ98b]. Advancing [FVD^{+}12]. Advection [RAW^{+}16, WFV14]. advection-diffusion [WFV14]. Advertisers [Ano95a]. aeroacoustics [MVAXP22]. Aerodynamics [Jen01]. aerosol [WNS^{+}21, WSK^{+}22]. aerospace [ZLF^{+}22]. aerothermodynamics [DFU20]. affect [CH11, PSL21, STB23]. Affecting [SS04, KC22, LRD^{+}19, MP12]. affiliations [ESM19]. affordable [MPE^{+}11]. Africa [May06]. aftaleret [MG05]. After [Bon11, MSZ02]. Against [Hoh01, SG92, ZRNA20, JKS02, Rau04]. Age [Fer03, Hef97, PSP^{+}22, Rus14, San03, FN21, GJLT11, Zic01]. Agencies [PBH01]. agenda [GRJS01]. Agent [EKJ^{+}03, SCFR06, LQR17, NZPWR22]. Agent-based [SCFR06, NZPWR22]. Agere [Ano02b]. Aggregation [TGC^{+}21, ZJS^{+}20]. Agile [ABC^{+}14, BC20a, FQYS23, GEI^{+}11, PGW^{+}20, TGC^{+}20, WDK^{+}20, ATM22, KHA^{+}03, Mac18, You08]. Agreement [Bar01]. AI [Lou96, PSP^{+}22]. aid [Lal91, Deo90]. AIDA [Ano87]. AIDA-87 [Ano87]. aide [Rus88]. aided [Deo90]. aids [Ing92]. air [ACW04, MGPB20]. Aircraft [SKSM19, CBRSH22]. AJIS [Och09]. al [Ano04c, Bur04a]. Alabama [IEE92d]. Alberta [ACM88]. AlDraTex [Ber22]. Algebra [Coo95b, DMP^{+}02, WR71, Coo95a, Joy08, Joy09b, JČMG11, LR08, MZE13, Pag07]. Algebraic [Lev95b, Lev95a, SAC^{+}15, GJMPAM^{+}14, Kli90]. Algorithm [BdP13, Bar00c, Bar01, Joh18, LLWM23, LGW^{+}22, ZRNA20, AFZ17, AFZ18, ATCZ19, CLS95, CSEP14, DD10, FLA^{+}16, LZ11a, LZ11b, YZC22, GB06, SC08]. Algorithms [QR92, Sha95, Val93, Ban16, Ban17, PC13b]. Alias [Gup03]. aliasing [ZC01]. Aligning [BMZ14]. Alignment [vWHvW09]. alkali [SPAW17]. alle [Mol01]. Allegiance [CH10]. aller [DF00]. Allocation [YLL^{+}07, FG92, HC07]. allocator [Mat03]. Almost [BH17]. Almquist [Ano00e]. Alon [Aji17]. Alone [DDJ98a, ESM19, DDJ98b]. Alpha [Ano98, Ano00i]. Already [CGK^{+}02]. Alternative [Ian02, Liu06, MS12, PK10, SS06, Ste08]. alternatives [Pot06]. Altmetrics [ZW17]. am [Bud10, FBY^{+}17]. Amalgamation [IAS16]. Amant [Waa09]. AMASS [dlVRB21]. Amber [Cha13]. Ambidexterity [FN21]. Ambidextrous [O’S02]. AMD [Ano01c, SuS01, Zad02]. AMD-PCs [SuS01]. AMD64 [Hub03]. Americans [Sta96b]. AMGKQ [Joh18]. Amherst [IEE92a]. AML [Esp96]. among [Col09a]. AMS [Joy09a]. Amsterdam [Ahm08b]. Amtec [Ano96c]. Anaheim [USE90]. Analyse [Rau04]. Analyses [SJW22, BMS^{+}22, GFD^{+}24]. Analysing [HYA20, PSSH16]. Analysis [Ano18, Bak20, Bar01, BFC02, Gol06, Gup03, KY16, KS11, KJRD16, MTM^{+}19, Mar22, MRGP20, Mor08, NS01, Omb20, PSR16, PMG^{+}09, PKG^{+}10, PPG^{+}11, PBJ^{+}12, SSP17, SSP18, SDD06, WMK^{+}17, AAA^{+}12, Amb15, AG22, AKMS23, BOL14, Ban16, Ban17, BJM^{+}22, BSK^{+}15, Bow05, BDP^{+}14, CWZ06, DP09, Feixx, FM10, Gal01, GF17, Gv14, GV16, GJS^{+}02, HLL^{+}95, HFO^{+}12, HBB^{+}12, Höp04, IC22, Kam21, Koc09, KFYI13, MG12, Mas05, MPE^{+}11, MRS07, MOT^{+}18, NDDH^{+}21, Och12, PKGA22, PKH07, QB21, Raj13, RZWW23, Rau04, RAMB18, RP08, SBM^{+}10, SMS16, SGM^{+}08, SM08, SAOB02, THG23, UBR^{+}17, Váz16, WLD^{+}17, YLXZ16, Yes12, ZLL04, dA15]. Analyst [Wil71]. Analytic [Wut12, Hua23]. analytical [FBY^{+}17]. Analyzer [LO89, Pax88]. Analyzing [CZ22, APHV19, HPT17, PHT17, SAHP15]. Anarchism [Mog99]. Anarchy [Bar00a]. Anatomy [Lor95, Ros02a, Yan92]. Ancient [SCDS15]. Andrew [Ano00c, Teo13]. Android [CMTA19, CCK21, QB21, SL01]. Andy [Ano97a]. Angle [LHZ12]. Animal [GNR^{+}09]. animation [MWG^{+}90, MWG^{+}91]. anisotropic [YSVM^{+}16]. anisotropy [RZWW23]. Anniversary [SHB^{+}20]. annotation [GKP^{+}14]. Announcements [Ano96e, Ano01a]. Annual [ACM93b, ACM94, ACM95, Ano87, Ano88b, Den99, IEE92c, IEE05, MS91, USE99, USE00a, USE01a, USE01b, USE02b, USE02c, ACM93a, Ano94b, Bon93, IEE95b, Jef08, USE98a]. anomaly [HC07]. Anonymity [CGK^{+}02, CSD^{+}05]. Anonymization [TRB22]. Anonymous [CdR99, BP14]. Anpassen [Gün02]. ANSI [PKP02, PKP05, PKP05]. answers [Phi12]. Ant [HL02, XWZ^{+}23, ZK05]. Antarctica [Ana99]. anthropological [Zei03]. anthropomorphic [GV16]. Anti [Ano00k, Mau05]. Anti-spam [Mau05]. Anti-Virus [Ano00k]. Antique [Jes03a]. Antonio [IEE92c, IEE94b]. Anup [GAS^{+}01]. Anwenderhandbuch [Ano01c, Ron05a, Ano01c, Ron01a, Ron01b]. Anwendung [PKP05, G^{+}00, G^{+}02, GA04b]. Anwendungen [PKP02, PKP05]. Aonix [Kro99b, Kro99a]. AP [BSP11]. AP-1-mediated [BSP11]. Apache [BGL^{+}20, CJ17, DFT21, DGC^{+}07, Fie99, FvH03, LW03, MFH02, Sai01, THG23, Ude97]. API [Ano00i, CCK21, Zag14]. Aplio [Ano00j]. App [DDJ99, CCK21]. Appearance [LR11]. Appendix [TRM16]. Appgen [Ano00k]. applet [GM02]. Application [AJ05, Ano00c, Ano01i, Ano01j, CWB^{+}04, FFH^{+}05, GBG^{+}16, GB00, HMP^{+}15, LGW18, PPG^{+}11, VGdlP01, XXAD21, ACW04, AHG94, BH11, CFCA13a, CFCA13b, EKR91, FRBRF19, Fri97, HM10, Jon05a, Mac18, PSSH16, Thi99, ZLF^{+}22, DFCPSF15]. Applications [Ano00d, Ano00k, Ano00l, Ano02b, BK91, BJJ14, CKB^{+}05, DFP23, FHH11, Gag02, GD12, GM05, IEE92c, IEE94b, IEE95b, IEE05, JJ00, MRGP20, MS12, Per00, PK10, She01, SPDQ22, ACM94, AH19, APK14a, APK14b, BSW95, BSW^{+}14, Cha11, DP09, Hen92, Hin87, LGS^{+}17, LS04, McL92, Pet05, RHR^{+}21, Sai13, Yad07, Yeo05, ZSW14, TDBEE11]. Applied [Ano01a, DGBH93, ZK21]. Applikationen [GGK99]. Applix [Knu99b]. Applixware [GGK99]. Applying [SDD06]. appreciate [Fei23]. Apprentice [RW87, Wat85a, Wat87, Wat85b]. Approach [BSA22, DKMT11, GGB17, MH07, Mud97, TPK21a, Fra13, HKP02, TPSZ19, VB19, WFV14, ZD05]. Approaches [Egy01, ASAAM^{+}19]. Approval [CWB^{+}04]. approximation [MSB09]. approximations [PC13a]. apps [CMTA19]. Apress [Kuc06]. April [Ano04c, IEE92d, Uni01, Yuk94]. AQSES [PSS^{+}07]. Arbeiten [GGK99, Ste00a]. arbitrary [LH22, MRH23]. arbitrary-precision [MRH23]. Arc [LGW18, SPAW17]. Arc-Routing [LGW18]. Archimedean [GHH20]. Architectural [KGM^{+}16, BLG^{+}17]. Architecture [AWD^{+}18, BE06, CTP^{+}22, Hub03, KORP95, LHCH93, Lat03, SJV^{+}05, XXCL19, ACB^{+}16, CLL05, Nor23]. Architectures [DXT^{+}18, Val93, BSW95, CSP09, QR92, WFV14]. Archive [EKJ^{+}03]. Archiving [Ano01i]. Arden [LHCH93]. ARDI [Ano96c]. Arduino [KSH14]. area [BVT06, LLS11]. Arena [MS12]. Aren’t [BHP^{+}01]. ArgoUML [MAF22]. ArgoUML-SPL [MAF22]. Aristotle [HLL^{+}95, SMS16]. ARITH [BBdD17]. Arithmetic [BBdD17, Cse99, GSR^{+}04, LZ16, LZ17, Abb12, Bee17, Fär05]. Arizona [IEE05]. ARM [Jor04, Kuk98, DVC^{+}07, Jor04, TTB09]. ARM-based [Jor04]. ARM-basierender [Jor04]. Arms [Boy08]. arrays [Che95]. Art [BlG12, DGBH93, MS08]. Article [Gom99]. Articles [Ano03b, Ano03c]. Artificial [AT92, Ano87, Ano88b, BPG94, IEE94b, AK95, ODP15]. Artist [Log99, Ham07]. artists [Ham07]. Asbestos [Lew99a]. ASCII [Mud97]. Asia [IEE94a, CH06a, GAS^{+}01]. Asia-Pacific [IEE94a]. Asian [CM06]. ASIS [Bon93, RSZ96, RSKF96]. Asked [And03]. Aspect [CC05, CSP09, HOST05]. Aspect-oriented [CC05, CSP09, HOST05]. AspectJ [CC05]. Aspects [BCB^{+}17, Jak03, HK03]. Aspekte [HK03, Jak03]. Assays [DKMB14]. Assembler [Mah13]. assembly [Win95]. Assessing [BJJ14, CH06b, HK09, LRP21]. Assessment [LMM02, LMPT22, GSW08, JD19, NDDH^{+}21]. asset [FN21, GGH10]. Assignment [Sha03, Liu08, SBS20]. assignments [GSW08]. assistance [LC12a]. assisted [LF90, LSF94, LFA92, SBM^{+}10]. Association [Ano90c, Ano00j, Sie04, The04]. Assurance [CKB^{+}05, KY16, KMF^{+}07, dlVRB21, GSW08, ZE03]. Asterisk [Sch04, VSM06]. ASTERIX [AAA^{+}12]. AsterixDB [AAA^{+}14]. asteroid [GVOM09]. astonishing [BH17]. Astronomy [Bar01, MP00]. Astrophysical [Owe01]. asynchronous [BG95]. AT75C310 [Ano00j]. ATARI [TSM88]. Atlanta [Dig82, USE00a]. atomistic [AZ17a, AZ17b]. atoms [SPAW17]. Attack [Ahm08a, Sta06, YDZ19]. Attacks [JKS02]. Audacity [JP09b]. Audio [TF21, FG16, MMR95]. Augmentative [PK10]. August [HY14, IEE95b, Lei93a, SS93, USE00b, Ano96e]. ausgereifte [Bra04]. Austria [Jef08]. Austrian [WP04]. authentication [Coc01a]. author [MG05]. Authoring [Ano01i, CBB06, WRSG92]. authors [Mol01]. authorship [APHV19]. auto [KTTK17]. auto-scaling [KTTK17]. Autoconf [VETT00, Cal10, Fri97]. autogeneration [SC08]. Automake [VETT00, Cal10]. Automata [AWD^{+}18]. Automata-Processing [AWD^{+}18]. Automated [BY92, Bun94, GV16, Kap92, LMW12, RDKT12, TPK21a, TBPS15, Vor92, BCHR12, KG20, PSS^{+}07, THG23, BY91]. Automatic [DKMB14, Kro99a, UNF^{+}08, KGMI06, MSM10]. automatically [Wen90]. automatically-generated [Wen90]. Automating [KKT17, Mof02, PSP^{+}22, YXS^{+}19, SD16]. Automation [Bax01, IEE89, MPG^{+}16, SL01, WMLM22]. automaton [Mak03]. autonomic [BJWZ08]. Autotools [Cal10, Zad02]. Autovectorization [Nai04]. Autumn [Ano89]. av [Esp96]. Availability [Bon11, Reh01a]. Available [Ano04b, GM02, Kop05, ODP15, PH16]. Avanti [Ano03a]. Avenue [Ano00j]. Avilon [Kro99b]. Avionics [PG02]. avoiding [Sta96b]. AVP [Ano00i]. Award [Bar00b]. Awards [GAS^{+}01]. Aware [ZZZ22]. Away [Bro19, Den99]. Awk [Anoxx, Lie92, Rob11, AKW88, Rob96, Rob97]. Axiom [Dal02, Joy08, Pag07].

B [Ano00l, Waa09, Sca05]. Baan [Kro99b]. Baby [Ano08c]. Back [Ano05a, EJS^{+}01]. Backend [Liu06]. backers [ZWH21]. Background [Per02]. backs [Ano01f]. Backup [Ano00i]. Bad [BHP^{+}01, Ros02b, Wil99]. bAIcis} [ZRNA20]. Bakar [BCHR12]. balance [Gal01, LGA20]. Balancing [KHA^{+}03, Kro99b, TPK^{+}21b, ZSW14]. band [MFB23]. Bandwagon [SSC^{+}00]. Bandwidth [CKB^{+}05, ZPH^{+}15]. Banff [ACM88]. barriers [NYB10]. Base [DFLS05, Sta92c, Sta97c, WCA^{+}14, Sta88d, Yeo05]. Based [AMS03, Ano01j, Ber96, BKP05, CTP^{+}22, CPJ^{+}98, EXA^{+}05, GB00, HMKC12, KMF^{+}07, KG01, Li18, LLWM23, LL14, LO89, MFS15, Mor96, PK10, SJV^{+}05, SSP17, SSP18, WCHRM21, Zha16, ZZZ22, AVA^{+}16, Ano06, AG22, Ban16, Ban17, CCA^{+}13, Car89, CV13, CM06, CJ17, CJ19, Cla90, CDR^{+}15, CWZ06, DC23, DFU20, DBLF16, EKUR10, FRBRF19, FMT^{+}08, GLMS18, GJMPAM^{+}14, GM84, GDK21, GF17, GSW08, HLL^{+}95, HWL^{+}23, HNH03, HZS^{+}16, HMX21a, HMX21b, Iwa02, JP09a, Jor04, KN93, KSH14, KKA^{+}19, KRR23, KFYI13, Mak03, Mam01, McH92, Mił10, MM04, Noj01, NZPWR22, PPC^{+}15, PH16, PM21, PC13a, PKP05, RW89, RTH15, SBS20, SMO^{+}13, SA15, SGD05a, SGD05b, SCFR06, SM08, SS23, SWTC23, TMM^{+}13, WN15, WGG^{+}19, Wat94, WG00, Wil14, YKSH20, YMCF23, YKK23, Zag14, ZRZ^{+}21, ZDM10, ASAAM^{+}19]. based [MSS95]. Bash [DLT^{+}23, Ram94a, Ram94c, Ram94b]. Basic [Ano01i, Ano01j, Tro96a]. Basics [EJS^{+}01]. basierender [Jor04]. Basis [GA04b, PKP02, PKP05, CYOS19, Ano01j]. Bastard [RAH^{+}01]. battery [CCK21]. Bayesian [BSA14, ZRNA20]. Bayh [MSZ02]. Bayonne [Sug02a, Sug02b]. Bazaar [Ray99b, Ray99c, FFvdH01, Ray01b]. BCI [SB08]. BDMS [AAA^{+}14]. Be [HR11, Ano00h, CK08, CPG^{+}04, Ell12, JDB09, Sta96b, Coc01b, Kro00]. BEA [Ano04a]. Beachtung [Stö04]. Beats [Ano00h, Bar01]. Became [Boy07]. Beck [Oms03, Oms03]. Becomes [Wol03a]. Bedingungen [SG05]. Beer [Gre18]. Befehle [Gün02]. Before [CRW^{+}04]. Beginning [Pec08, Ron05a, Wri00]. Beginnings [Mur09]. Begrifflichkeit [Geh96]. Behavior [LR11, JCNS^{+}22]. behaviors [jFFR16]. Behaviour [TTB09]. behaviours [HMP^{+}15]. BEHEMOTH [Bar00c]. behind [Cas19]. Beijing [Bao93]. Being [SvGH15, Fei23]. Beispiel [Bud10]. Beispiele [PKP05]. Beitrag [Oms03]. Bekämpfung [Rau04]. Belly [DB05]. Ben [Ano00e]. Bench [DFP23]. Benchmark [CH91, DFP23, ILG10, LGW18, Bes04]. benchmarks [WCS20, ZK21]. Benefit [Ebe09]. Benefits [CPJ^{+}98, CK10, Ros05]. benutzergerechter [FG85]. BeOS [Kro00]. beräkningar [Jön05c]. Berkeley [Gil88, Yad07]. Berlin [EHP94]. Bessel [VRS^{+}99a, VRS^{+}99b, SG99, VRS^{+}95]. Best [Ano00j, CFL23, Sid03, UMV15, Xia08]. Beta [Ano98, LSM^{+}99, Ano01a, LMOS93]. Bethesda [MSLH71]. Betriebssystem [CK06a, CK06b, CK06c, CK06d, CK06e, CK06f, CK06g, CK06h, SuS01]. Betriebssysteme [Bud10]. Better [Fie88, PM00, RMAM19, Ano08b, Li91, NXC13, Sta98b]. Between [CFM08, BD03b, CFMRL11, DRM21, Gal01, HPM^{+}08, Lam09, LC12b, MRS07, San08, SC02, ZFD21]. Beyond [Ano91, BRH10, CFL23, CSD^{+}05, DC00, Jin18, RLTD23, Van22, Wol98, GL14, MSR10]. BFD [Cha91, Tay99]. biases [STB23]. Bible [PR96, HHV05, Hun01, WY94, W^{+}95, vGS10]. bibliography [HM89, Lie92]. Bibliometric [NGJ03]. BibTEX-Mode [Che87a]. Big [AAA^{+}12, BW00, Fra13, Ing92, Val91]. Bigot [CPG^{+}04]. Bikeshed [Kam14a]. Bildbearbeitung [DF00]. Bildbearbeitungsprogramm [GGK99]. Bilder [DF00]. Bill [Ano00a, Ano00b]. Binaries [ASWD18]. Binary [Lew99a, Lew99b, XXCL19, BCR^{+}08, FHL^{+}07, Cha91]. Binary128 [LZ16, LZ17]. Binary64 [LZ16, LZ17]. Binding [Ano01j, Ano02b, Coo95b, Bad07, Coo95a]. Bindings [LFN^{+}11]. BinPo [MFB23]. Biogeography [FVD^{+}12]. Bioinformatics [CKB^{+}05, SHK^{+}03, DD08, KTTK17]. Biological [DKMB14]. Biology [Car01, KTH^{+}22, WLD^{+}17]. biomass [XAPK14]. biomedical [AJLM18, DP09, KTTK17, MVF20]. Biometry [MMD12]. biomolecular [LHZ12]. Biopolis [Ano06]. BioSig [SB08]. BIOTC [XAPK14]. bioweapons [JH16]. bipartite [PPR19]. Bipolar [WSK^{+}22]. bird [Rob11]. Birds [Dew07]. Birds-of-a-feather [Dew07]. Birmingham [IEE92d]. Bison [DS99, Vol89, DS88, HSC89, DS90, DS99, DS00, DS02]. bit [Ano96b, Ano00h, GHL^{+}04, Jae03, STS92]. Bit-Mapped [STS92]. Bitcoin [Cap12]. Bitcoins [Hol15]. Bits [Rus09, Eub05]. Bittco [Kuk98]. Black [Ano08a]. BlackParrot [PGW^{+}20]. Blame [CWB^{+}04]. BLAST [Ano96b]. Blender [JP09b]. Blind [WRDP17, Man92]. Bloch [RJ21, SDL^{+}16]. Blockchain [PLO^{+}23, TNM17, MQN19]. Blockchain-Technologie [TNM17]. blocking [VGSN18]. Blocks [Ano00j]. blog [PM13]. Blossom [SH19]. blueprint [Mon03]. Board [Bar01]. bodies [SNC^{+}06]. Body [BY14, JWC18]. Body-Worn [BY14]. Boltzmann [ASC^{+}21, FBY^{+}17, KKA^{+}21, ZCG17]. Bonas [QR92]. Book [Aji17, Ang01, Ano97b, Ano97a, Ano99a, Ano00c, Ano00a, Ano00b, Ano00e, Ano00d, Ano11, Ano15a, Bar16a, Bra92, Cas02, Cha13, Chi97, Cho09, CDsJ^{+}00, Cro00, Gil06, Jen97, LMW12, Men12, PKP05, SD16, Teo13, TG15, Waa09, LD13, PKP02, Yad07, Ano15a, Cha13]. books [Sta01a]. Bookshelf [GF99, GS00]. Bootstrapping [Tay19]. BORIS [FG16]. borne [Eds16]. borrow [Sib17]. borrowing [Har05]. Boston [IEE92b, USE01b]. both [KHA^{+}03, YLXZ16, YZC22]. Bots [HBGS19, WWSG21]. Boundary [WP04, KD23, MVF20]. bounded [Rog09a, Rog09b]. Bounds [Wut12]. bounty [ZWH21]. Bourne [Ano00k]. Boussinesq [TL17]. Boussinesq-type [TL17]. Box [Ano00i, Hae02, RAH^{+}01]. Boy [RAH^{+}01]. Branch [Sim00, PdSCJM22]. branches [GK92]. Brave [TG99]. Brazil [EKJ^{+}03, Hal02]. Brazilian [Bro04]. BRB [KY16]. Breaking [BKHT21]. breast [WHJ15]. Breed [Vol89]. Breeze [Ano02b]. Bremen [EKR91]. Bridge [Ano00i, Ano01i]. Brief [Ano15b, AD04, GB21, Kri90, San01]. Briefs [Gla99, PM00]. Brighton [AT92]. bring [VMKB05]. Bringing [San03, Fri06, Sea04]. British [MG94, SM89b]. Broadband [GEMN07]. Broken [Ahm08a]. Broker [YMLT14]. Brooks [Bar00b]. BROOM [CCA^{+}19]. Browser [Hau01, HBC^{+}05, Yac88]. browsing [RM92]. brukergrensesnitt [Esp96]. BSD [DF00, Lin02a, Cor05, Guy00, Lin02a]. BSDCon [USE02a]. BSDL [Jak03]. BSDs [Luc99a]. BSE [RAMB18]. bubbly [LMHL20]. Buch [GGK99]. Bucks [Ano00k]. Buffer [CPG^{+}04, Wag03, Ano99b, Rog09a, Rog09b, ZLL04]. Bug [CPJ^{+}98, TLL^{+}14, ACB18, DRM21, FM10, JCNS^{+}22, SBS20]. bug-assignment [SBS20]. bug-fixing [ACB18]. Bug-Free [CPJ^{+}98]. Buggy [CPJ^{+}98]. Bugs [DLT^{+}23, SIK^{+}13, TZH22, VGSN18, ZKDP22, ZFY^{+}19, ZRGJ21]. Bugzilla [ZK05]. build [APK14a, APK14b, Kop05]. builder [Mei92]. Building [AAB^{+}05a, AMS03, Ano00d, BdSI15, Bur04b, CC04, KTTK17, Mac18, Nej12, Per00, PK10, Raf23, SJV^{+}05, Sch03, Ste00b, TG99, WHJ15, WKB14, Woo01, Yeo05, DPH16, EHP14, PSSH16]. Builds [Mof02, dCdCM14, Mac18]. built [JDB09]. Builtins [RMAM19]. bulletin [Fre87]. Burdens [CGB^{+}05]. Büropakete [GGK99]. Business [Ano00k, Ano01j, Bar01, Bro19, Hec99, LRP11, Rie21, Hay05, WBY^{+}08]. businesses [BGL^{+}22]. Buttons [STS92]. Buy [CDsJ^{+}00]. Bye [Coc01b].

C [Ano90d, Ano96b, DF00, FD92, Lea88, Lea92, Lea93, LMOS93, LSMO96, LSM^{+}01, Loo15, Oms03, PKP02, PKP05, SGD00, Tie88, Tie93, Ahm08b, Ano88c, Ano93a, Ano01j, BM06, BR95, Ber96, CR92b, CWM^{+}20, Che95, Col02, Dav91, DP09, DC00, Edd96, Eig03, FHH11, FG92, Gar09, Gil88, GK92, HR19, He95, Ho95, KOI94, Kle21, KORP95, Lea94, Lla06, LSM^{+}99, LSM^{+}00, LS04, MLA^{+}19, Mei92, MRGP20, MVF20, MSS95, Mit94, NZPWR22, OK94, Oma89, PKH07, PQM11, PH82, Pin02, PKP02, PKP05, Sai13, Sal88, SC08, Ste95, Ste99, Ste00b, Ste01, Tie90, USE88, Vol96, Wal93, Win95, Yac88, Yes12, YSVM^{+}16]. C# [BHP^{+}01]. C-Tree [Ano96b]. C. [Oms03]. C/C [FD92, PKP02, PKP05, Eig03, SC08]. C99 [She07]. CA [Ano97a, Ano99a, Ano00a, Ano00b, Cas02, Lei93a, USE94, USE01a, USE02a]. Cache [RLVdS21, Cha92, Fär05, GGSRMPM20, GYW^{+}23]. cache-clustering [GGSRMPM20]. cache-partitioning [GGSRMPM20]. CACSD [Ano96d]. Cactus [HL02]. CAD [Ano04b, RP08]. CAD-oriented [RP08]. CADE [Bun94, Kap92]. CADE-11 [Kap92]. CADE-12 [Bun94]. CAGE [RHW^{+}21]. Calc [Gil93, Ano96a, BMR^{+}23]. calculate [Yap11].

Calculates [Coc01b]. Calculating [ALA20, GFZ16, SC16, SPAW17, YKSH20]. calculation [VRS^{+}95, VRS^{+}99a, VRS^{+}99b, XOTI22, BMR^{+}23]. calculations [Dan11, KPK^{+}17, MSB09, DARJ23, RAMB18]. calculator [MBTB21]. Caldera [Ano98, Ano00e]. Calendar [ACM05]. Calendaring [Sif00]. Calibration [BVLF14]. California [ACM92, Ano94b, USE90, USE99, USE02b, USE02c]. Call [Ano03b, Ano03c, KG01, PMBM^{+}15, Ano04a, GKM82, GKM04, Hub04b]. calls [Och09]. Caltech [Bar00a]. camera [GTMR23]. Can [Ebe09, EMdL^{+}07, Ell12, JH16, Sta01a, Ste08, Wea03, BR03]. Canada [ACM88, Ano00i, Lev95b, Lev95a, MG94, Ass95, HDR03, HDR04]. CAOVerif [ABF^{+}14]. capabilities [Bri09a, Bri09b, FN21, KGT22]. Capability [MMD^{+}22]. Capacity [Ano01j]. Capital [Got05, CH11]. Capitalism [CMJ^{+}04]. capture [Far23, JCNS^{+}22]. Capturing [Ale92, Pal87, PDG^{+}88]. CAPWAP [BCI^{+}09]. car [MSR10]. Carbopolis [EXA^{+}05]. Card [Chi93, Dig80b]. cardiac [BSW^{+}14]. Cardiff [Ano01i]. Cards [SSC93]. care [EKUR10]. Career [Rie15]. Carlo [Adk11, HWM^{+}15, Hua17, Hua23, MMY^{+}19, SMRM^{+}17]. cas [Cor00]. Cascade [YYL^{+}15]. CASCON [BGG^{+}94]. Case [AtHR11, Gau07, Goo14, GYW^{+}23, KGM^{+}16, LMZP19, O’D07, Rie10, TTB09, VGdlP01, WKS^{+}14, Zad02, vdLLM09, BSW^{+}14, BGL^{+}20, DTB05, DDHS03, Emb06, FvH03, Gal01, Gau03, Ger03, GPPT16, GGH05, HZ14, JK11, KFYI13, LLEL^{+}23, MFH02, MLWR18, NDDH^{+}21, PAB^{+}17, Pya06, SSS22, Shi12, VSGM14, YAS91, YA11, ZWU22, vKSL03]. cases [MAF22]. Cassandra [SMRM^{+}17]. Cat [IEE94c, MS91, MSNS91, Pow14]. CATCH [LF90]. Catches [Bar01]. categorization [KGMI06, NRRS20]. Cathedral [Ray99b, Ray99c, USE02a, Ray01b]. Cause [MTM^{+}19]. Cause-and-Effect [MTM^{+}19]. Causes [MTM^{+}19, CMTA19]. cautionary [Bur04a]. Cayenne [Ano98]. CC [Sta88e, Sta92b, Sta00b]. CCA [Com84, CCA84]. CCG [Mor91]. CD [Ano96b, Ano97a, Ano00a, Ano00b, Ano00i, Ano01d, Ano01c, DF00, PKP02, PKP05, SuS01, PKP05, Rod00]. CD-ROM [Ano00a, Ano00b, Rod00]. CD-ROMs [Ano01d]. CDK [SHK^{+}03]. CDs [Gün02]. CE [Ano95a, SSC^{+}00]. ceiling [MSM^{+}03]. Celebrate [CSD^{+}05]. Celebrating [SHB^{+}20]. Celeris [TL17]. Celerity [DXT^{+}18]. Cell [GEMN07, BDAW15, BCP^{+}16, DBP^{+}18]. cells [NGCI^{+}12]. Cellular [MPE^{+}11]. censorship [Shi12]. Center [ACM00, Bao93, Cha98, GB00, KG01, WBB^{+}74]. centered [For12]. Centers [LLWM23]. Central [Har94]. centralizada [VD01]. centralized [VD01]. Centre [Rit88]. Centric [BE06]. Century [ST10, Sal94, Sch09, Har05]. CEO [Ano04a]. CERN [VW92]. CERNLIB [FP94]. certain [DB02]. Certification [BCPS10, BS14, DBBA10, Mag04, Mir07, Mag00, Mag01a, Mag01b, Mag01c, MS02, Sai01, Sai02, dlVRB21]. Certified [Ano00c, Mag04, DRP01]. Certifying [SK04]. CFD [AH19, CDR^{+}15, Gro01, KDM17, XAPK14]. Cfengine [BS98]. cflow [Gol06]. Château [QR92]. Chain [Har20, Sch03, You08]. Chains [Mar22]. Challenge [XMGM21, XMGM22]. Challenges [Her20, Kle21, PWA^{+}19, SDD06, TZ22, Ten93, VH04, WCG22, WWSG21, WG06, dCdCM14, BAR16b, BGL^{+}22, Cap13, Dei10, Eds16, PSDG18, WBGM02]. Change [BAP00, CSY^{+}04, CDsJ^{+}00, Fly87b, YXS^{+}19, Ban16, Ban17, GGT05, Joh94a, KT05, KL07, NS05, TG99, ZK21]. Change-Level [YXS^{+}19]. change-prone [KL07]. Changed [Lus04, BH17]. Changes [BKHT21, DDJ98a, RAH^{+}01, DDJ98b, AKMS23, CJ19, CCK21]. Changing [Edw98, Law09, Par03, Rie15, Sal08]. channel [GYW^{+}23]. Chaospy [FL15]. Chapter [AFS81]. Chapters [AFS82]. Characteristics [MSZ02, SBDR22, ABC18, TLL^{+}14]. Characterization [HKA^{+}19, AJ05, CH91, ZKDP22]. Characterizing [CJ17, VGSN18, KL07, ZSW14]. charging [WSK^{+}22]. Charles [Ano00e]. Chassis [Ano00i]. Chaste [BSW^{+}14]. Chauhan [TG15]. CHEBINT [PC13a]. Chebyshev [PC13a]. Checker [Kro99b, LQR17]. Checking [CCG^{+}02, BCHR12, DDHS03, Eig03, SP12, WK93]. checkpoint [LFA92]. checkpointing [LF90, LSF94]. checks [Due97]. chemical [ABNÅ05, RHW^{+}21]. Chemistry [BH07, Boy07, Boy13, LB00, Lip07, MSLH71, Par03, SW13, WBB^{+}74, HHG^{+}21, PM21, WPAV14, HHG^{+}21, SHK^{+}03]. Chemists [HW17b]. Chemo [SHK^{+}03]. Chemo- [SHK^{+}03]. CheMPS2 [WPAV14]. Chicken [Les03]. Chinese [CLM^{+}08, EKJ^{+}03]. Chip [FQYS23, KRB^{+}22, Sta06, WDK^{+}20, Ano96c, Don04]. CHIPKIT [WDK^{+}20]. Chips [DXT^{+}18]. Chisnall [Fox08]. Choice [PMD13, CF09]. Choosing [Eng10, Spi11]. Chord [CSP^{+}03]. Chosen [JKS02]. Chosen-Ciphertext [JKS02]. Chris [Ano99a]. Christopher [Pri19]. ciberespacio [Les01]. CIO [CGB^{+}05]. Ciphertext [JKS02]. Circuit [PGC21]. Circuits [SFWD12]. City [CDsJ^{+}00, SG05]. claim [Sie99]. Claiming [Mog03b]. Claims [GB00, Mog03a, Stö04]. Clara [Ano94b]. Clarification [GWT^{+}01]. Clarifying [GF99]. Clarity [Hol23]. Clark [Kim01a]. Class [BHP^{+}01]. Classes [Mor92, Ban16, Ban17, How98, KL07, SK12]. Classic [MZH22]. classical [Aki16, GLT08]. classification [KKA^{+}19]. Classifiers [BSA14]. Classifying [Li91]. Classroom [Alf05, Pow14, RT05]. cleaning [MCQF21]. ClearMail [Ano04b]. Cliche [Wat94]. Cliche-based [Wat94]. cliches [Car89]. Client [Ano96b, Ano00k, MSM10]. Climate [HR11]. Clone [MOMM11, Joh94a, WCS20]. clones [RC10]. CLOS [SO91]. close [Hac98]. Closed [Ano04a, Ano04c, Eri01, KKN^{+}21, Dwa04, For07, PSE04]. Closed-Source [Ano04c, PSE04]. Cloud [FVD^{+}12, Gan17, IAS16, LLWM23, MSC19, MB16, PPC^{+}15, VOM12, YMLT14, KTTK17, KG20, PPR19, ZFY^{+}19]. Clouds [BdSI15]. cluster [Kop05, Kop05]. Clustering [Kre00, Cre07, GGSRMPM20]. Clutter [BNST99]. CMMI [vWHvW09]. CMMI-DEV [vWHvW09]. CMOS [CCA^{+}19]. Cnest [KC92]. CNNParted [KSS^{+}23]. CNNs [LGW^{+}22]. co [Ano02a, KS02, ZVvDD11, USE88]. co-evolution [ZVvDD11]. co-operation [KS02]. co-ordination [Ano02a, KS02]. coagulating [Mor91]. coagulation [WNS^{+}21]. coaxial [YMCF23]. Cobol [EJS^{+}01]. coco [Ano01a]. CODE [BY91, Bak20, BHP^{+}01, BCB^{+}17, Bro19, CJ19, Col09a, Cou17, DLT^{+}23, Gup03, Han00, Kro00, Les99, LPFD21, Lio96, Lók04, Mir07, SAOB02, SHS^{+}93, STB23, Van22, WFF18, XXCL19, ALGE12, AHG94, Añe11, Ano99c, Ano02a, Ano04a, APHV19, BOL14, BY92, CAWK22, CZS^{+}21, Cas19, CLS95, CMTA19, DGJH19, DIK^{+}23, DBP^{+}18, DFU20, Eub05, GCK^{+}17, HMP^{+}15, HC07, KKA^{+}21, MFB23, MLA^{+}19, Moo01b, Moo01a, Mor91, MSR10, NMS14, Nor23, Pit16, RDZ20, RVLS14, SSAO04, SC16, SHW^{+}21, Smi17, SH11, Tai13, WCS20, Wen90, XAPK14, XTY^{+}22, Yac88, Zag14, ZVvDD11, ZLL04, Gho07, MZE13]. code-clone [WCS20]. Code-Issue-Introducing [CJ19]. codebase [Big13]. Codec [Haf01]. codecs [GLCMC17]. CodeFutures [Ano04b]. Codemesh [Ano01i]. Codes [UNF^{+}08, SG99]. CodeWarrior [Kro99c]. código [Les01]. Coding [Ano95e, Fow00, FG16, IC22]. coefficients [MBTB21]. Cognitive [CWB^{+}04]. coherence [CKS16]. Coherent [KRB^{+}22]. cohesive [QC18]. ColdFusion [Ano02b]. CoLiS [BJM^{+}22]. Collaboration [FFHL05, GGB17, SBDR22, MG12, NS05, Sin10a, Wii91a]. Collaborations [Fre23, PRRL12]. Collaborative [ACC^{+}12, Gho07, Hef97, PMBM^{+}15, PFL^{+}12, Rav00, CV13, DBP^{+}18, Wii91b]. Collation [Noj01]. Colleagues [PLS^{+}91]. Collect [CSD^{+}05]. collecting [ZWU22]. Collection [AHB^{+}09, CR92b, Sta99, Sta00c, BOL14, Sta03b, Ahm08b, BBM^{+}21, EGH^{+}05]. collections [WMK^{+}17]. CollectionSpace [For12]. collectives [ZWU22]. College [Bar01]. Collider [ZC95]. Collision [TBPS15]. collocations [WFF18]. colloidal [BMT^{+}20]. Colorado [USE00b]. colorimetric [Amb15, BSK^{+}15]. coloring [Mat03]. colour [MM04]. colour-map [MM04]. Columbia [MG94]. Columbus [Bon93]. COM [Ano00j]. Combination [BSA14]. combining [Kan12, SHW^{+}21]. Come [Bar00a, Kam11]. comes [Fie90b]. Comm [AZ17a]. Command [Tan11a, Tan11b, Cra90]. Command-Line [Tan11a]. Commands [Coh82, Li91]. Commentary [Lio96, O’D07, Max01]. Comments [PLS^{+}91, Wil71, MOT^{+}18]. commerce [Sca05]. COMMERCIAL [BY91, BBD^{+}96a, BB02, FL16, Gre80, MD04, MEB^{+}20, VOM12, ALGE12, ACKT20, BY92, GGH05, HBR19, PC13b, Sie99]. Commercializing [Kar03]. Commodification [vdLLM09]. commodity [Kop05]. Common [KRR23, RAH^{+}01, Phi12, Gad88, Hen92, LH03]. Commons [VH04, Mah03, Mor11]. Commun [VRS^{+}99a, VRS^{+}99b]. CommunesPlone [VJ23]. Communication [Ano00k, AD04, Bao93, CK06b, CK06c, KRB^{+}22, PK10, PBJ^{+}12, jFFR16, Koc09, Lin02a]. Communications [Ano96b, Ano01i]. Communism [CMJ^{+}04]. Communities [ASS^{+}23, APCs22, BKR^{+}20, BSFR22, CH06b, GBICMR13, GB21, Lev23, TZ22, ESM19, Far23, jFFR16, GL14, PM13, RCO20, SSR02, WGS07]. Community [BGL^{+}21, DFT21, PRRL12, San08, Sca04, TPK21a, Ter00, Wil99, XMGM21, vKSL03, AAB^{+}05a, Ano02a, CSEP14, Emb06, ES23, jFFR16, LH03, Mah03, MLMFN^{+}15, TPSZ19, VSGM14, XMGM22, Pel89]. community-oriented [Emb06]. como [RÓ01]. Compact [Ano03a, CRB^{+}18, PPR19]. compaction [DVC^{+}07]. Companies [EKJ^{+}03, GBICMR13]. Company [BGL^{+}21, Ano01a]. Compaq [Ano01a]. Comparative [BSK^{+}15, Amb15]. Comparing [Ahm08b, BH11, DGC^{+}07, KT05, PC13b, vGPB10]. Comparison [SO91, Sin10b, VOM12, ALVV17, ACKT20, Aye97, GJS^{+}02, JLL23, Lam09, She07, YSC^{+}06, ZFD21]. Compatibility [Egy01, KKT17]. Compatible [HWZ01, AAB^{+}05b, DS88, DS99, DS00, DS02, Guy00, Woo01, DS90]. CompatibleOne [YMLT14]. Competition [Gau07, HK03, Cor00]. competitive [Dan11]. compilation [Big13, GJS^{+}02]. Compile [Bot03]. Compiler [Ahm08b, AS97, Ano01i, BBM^{+}21, Col02, EGH^{+}05, FKM^{+}11, Gil88, LSF94, LFA92, Mir07, PKP02, PKP05, SZAB97, SZAB98, SZAB99, Sta99, Sta00c, YLL^{+}07, Ano01a, BB91, CGS94, DuB02, FG92, FMT^{+}08, GHL^{+}04, GK92, He95, Ho95, Kir12, LF90, MSK05, MRS07, She07, Smy97, Sta03b, TG99, CZ99, ZC01]. Compiler-assisted [LSF94, LFA92, LF90]. compilers [ALGE12, Bee17, Gou04, Sal88, Win95]. Compiling [DC00]. complaints [Raj13]. Complete [Ano98, Gri02, LD13]. completely [JP09a]. Complex [MTM^{+}19, Neh04, Neh07, BJWZ08, BG12, CZS^{+}21, YMCF23]. Compliance [Cou20, GD12, Omb20, PZ20, Sch19]. compliant [AM18, VSdCCR23]. complicated [PH16]. Component [Kro99b, Kro99a, PMM17, PMM18, PK10, BSC^{+}21, CKB11, HWL^{+}23, PPC^{+}15, SNC^{+}06]. Component-Based [PK10, HWL^{+}23, PPC^{+}15]. Components [ACC^{+}12, CYL^{+}23, LMPT22, RW87, Spi11, Spi19, BGL^{+}22, CLM^{+}08, LMZT22, PSSH16, XTG^{+}11]. Composability [HS15]. COMPOSER [CRB^{+}18]. Composition [Omb20]. Comprehensive [BDP^{+}14, HKA^{+}19, ZRNA20, Fra13, Gar00, SAHP15, VBG^{+}10, WM01, You08]. compressible [BSC^{+}21, HWL^{+}23, SPLD20]. Compression [KW94, SC00, BGM99, Fow00, SGD05a, SGD05b]. Compromise [Ahm08a]. COMPSAC [IEE95b]. Comput [AZ17a, CFCA13a, VRS^{+}99a, VRS^{+}99b]. Computation [Jef08, Lev95b, Lev95a, Mio90, PQM11, WBB^{+}74, GJMPAM^{+}14, GVOM09, KGT22, KD23, MBTB21, Sai13, SDL^{+}16]. Computational [Ano01a, BH07, Boy07, Boy13, HW17b, LB00, Lip07, MSLH71, MCGA22, NMS14, SKSM19, TDBEE11, WKA^{+}08, BCP^{+}16, BSW^{+}14, CFCA13a, CFCA13b, MDRN18, SHB^{+}20, WNS^{+}21, WSK^{+}22]. computations [ABNÅ05, Eat97, Eat00, Eat02, Eat05, EBH08, Jön05c, RBM^{+}23]. compute [KSV16, MFB23, VB19]. compute-and-forward [KSV16]. Computer [AFS81, AFS82, Ano88b, Ano00j, Ano00k, Bao93, Bar01, BKP05, BBdD17, Coc01b, Cse99, DMP^{+}02, Est06, IEE92c, IEE94c, IEE95b, IEE05, Kuk98, SM89b, Ten93, And01, BSK87, Cra90, Deo90, DDA^{+}07, DD08, Dre94, EKR91, HETD09, Hol05, JP09a, Joy08, Joy09b, JČMG11, LR08, Lla06, MZE13, MP00, NN00, Neu84, Pag07, SBM^{+}10, SA15, SS93, YAS91, CPJ^{+}98]. computer-assisted [SBM^{+}10]. Computer-Implemented [Est06].

Computers [IEE94c, Par03, SNF04, SM89b, Kro00]. Computersysteme [FG85]. Computing [ACM00, AY93, And03, Ano97d, Ano00j, Bar01, Bra97, BMB^{+}18, CSD^{+}05, Cse99, DGBH93, FVD^{+}12, Gan17, HE17, Hom00, IEE90, IEE92b, IEE93, IEE95a, MSC19, Ten93, VW92, ZZZ22, Zim10, AH19, BPG94, Bik96, Bor09, CZS^{+}21, Dan11, EHP94, HTU96, HXS20, JP09a, Kaw92, MTD^{+}09, MM10, PCAJ^{+}23, SS05a, SZ05, Spi21, TACA15, Wol02, YM93]. Con [DDJ99]. concept [Bow05]. Concepts [TG15, Geh96, Kra05, Sai02]. conceptual [KG20, Tai13]. Concern [HOST05]. CONCERTO [Maz15, SA15]. concurrency [Dan11]. Concurrent [KK94, MWB89, MR94]. condensed [GCK^{+}17]. Conditions [SG05, WP04, MVF20]. Conduct [LPFD21]. Conducting [LGA20]. Conference [ACM89, ACM92, ACM93a, ACM95, ACM97, AT92, Ano87, Ano88b, Ano88c, Ano89, Ano90a, Ano90b, Ano90c, Ano90d, Ano91, Ano93c, Ano94a, Ano94b, Ano06, Bao93, BSW95, BK14, BSK87, Bun94, EHP94, FMA02, Fur90, HY14, IEE89, IEE92c, IEE94a, IEE94c, IEE94b, IEE95a, IEE95b, IEE05, Kap92, MS91, MSNS91, MG94, SS93, ACM05, SC00, Uni01, USE88, USE90, USE94, USE98b, USE99, USE00a, USE01a, USE01b, USE02b, USE02c, VW92, Vor92, Yuk94, AK95, Lei93a, MSLH71, SM89b, USE98a, Den99, Mar01, Xia08]. Confession [Gla08]. Configurable [RLVdS21]. configuraciones [VD01]. Configuration [Bla06, BCI^{+}09, Bur95, HPT17, Jan01, Mag00, Mag01a]. configurations [VD01]. configured [LLEL^{+}23]. Configuring [Reh01a]. conflict [ABC18, FFHL05]. Conflicts [GMBv20]. Conformal [GNR^{+}09]. Congress [Ano14, FvdHJ10]. conjoint [BH11]. Connect [Ano00i]. Connection [Uni01]. cons [Ano04d, Gre80]. consciousness [Ale92]. conservation [KTH^{+}22]. Considerations [Amb15, BGL^{+}22]. considering [YZC22]. Consoles [GAS^{+}01]. Consortium [Bar01, Sta06, Hoh01]. Constitutional [DPL^{+}91]. constraint [GB06]. Construction [KK17, BJWZ08, Fin22a, HMR93, Muw09]. Consultants [Ano95a]. Consumer [GB00]. Contained [Ano03a]. containers [Dav91]. Contemporary [BCB^{+}17, RCP^{+}12, IDSM23]. Content [MSW09, BCvE^{+}05, Cap13, Gal01, GRJS01, Shi12]. Contest [Bar00b, Bar00a]. Context [Men10, AH19]. Contexts [LBF^{+}22]. Continental [Bao93]. Continuation [GM05, Hua23]. continuing [DCS05]. Continuous [PZ20, HWM^{+}15, Hua17, KG20, SS23, WMLM22]. continuous-time [HWM^{+}15, Hua17]. Contract [Jon02, BCHR12, MG05, Oms03]. Contrast [Jes03b]. contribute [Ano04a, Spi21]. Contribution [BE06, AKF21, HK09, JZ09, Oms03]. Contributions [BGL^{+}21, EGH^{+}05, FCTP21, KCAS23, BS14, GSW08, LMWM18]. Contributor [KCAS23, Jon01]. Contributor-Abandoned [KCAS23]. contributor-run [Jon01]. Contributors [TGW^{+}22, WCG22, HNH03]. Control [Bao93, BVLF14, NRG^{+}99, RCB^{+}14, SSC^{+}00, Sim00, TV99, Ano96d, BG95, FFHL05, Gal01, Gar09, MGFRG12, RHW^{+}21, RCGB^{+}22]. Controlled [CW15a, CW15b, Kir12]. Controller [Ben78, AM18].

Controlling [Rie11, KSH14]. convection [ZK21]. convection-coupled [ZK21]. Convention [ACM00, Bao93, JH16]. conventions [CNSR23]. Conversation [Flo94]. Conversations [LPFD21]. conversion [QLC^{+}12]. Converters [Ano95c]. Converting [Kro00, Rad92]. Convolution [LGW^{+}22]. Convolutional [KSS^{+}23]. Cookbook [Ano00c, Fin80a, Fin80b, G^{+}06, Phi12]. ’COOL [MR94, CPG^{+}04, Eub05]. cooperation [MG05]. Cooperative [Dan11, Bro04]. Coordinating [HBGS19]. coordination [Koc09, SMS16]. Coping [ZC01]. Copley [USE01b]. Coprocessor [Gut00, AV04]. Copulas [GHH20]. Copyfuture [Kenxx]. Copyleft [Hef97, Kenxx, O’S02, Gom99]. Copyright [Gil05, Kenxx, Mog99, O’S02, San03, Sto09, Dre94, Geh96, NO03, Oms03, Stö04, Sur01b, Zic01, Zic01]. Copyrights [Vai01, Gil04]. copywrongs [Vai01]. CORBA [Ang01, BES^{+}01, Pud04]. Core [AML^{+}10, DXT^{+}18, Hub04a, MMD^{+}22, Max01, ACB^{+}16, Sai02, SGNB08]. Corp [Ano98, Ano00j, Kuk98]. Corporate [Fer03, GGH10, YLG05]. Corporation [Ano00i, Ano00j, Ano00k, Ano00i]. Corporations [San08]. corpus [MNS19]. correct [FHL^{+}07]. Correction [BHP^{+}01, BTL^{+}11, YLXZ16, YZC22]. Corrections [Ano95a]. CORRECTNESS [BY91, MD22, BY92]. Corrector [Dig75a, Dig75b]. correlates [PdSCJM22]. correlation [MRS07]. Corresponding [LRP11]. Corrigendum [AZ17a, CFCA13a]. corrigés [Rod00]. cortex [Gv14]. COS [Ano96c]. COS/Print [Ano96c]. Cosmology [Coc01b]. Cost [CTP^{+}22, Gal60, RDKT12, Smy97]. CoStLy [Neh04, Neh07]. Costs [Kam14a, Kam14b]. could [Gal04, Sta96b]. Counterpoint [HR11]. Countries [FCTP21, CF07a, CH06a]. Country [Men10, YA11]. coupled [MZE13, MVAXP22, ZK21]. coupling [KGW^{+}21, ZLF^{+}22]. Course [AFS81, AFS82, Col05]. courses [HPM^{+}08, PBOP07]. courseware [BHMB03]. Court [Maj03, Bea04, Höp04]. CPC [SHB^{+}20]. CPP [SW15]. CPPPO [MGR16]. CppyABM [NZPWR22]. CPRM [GWT^{+}01]. CPS [PSSH16, Maz15]. CPU2006 [ALGE12]. CPUs [Ano00f]. CQF [Kop20]. Craft [Fin91]. Crafting [JP09b]. Craig [DPL^{+}91]. Cram [Ano00c]. Create [AG95, SKB23, Ste08]. created [Raj13]. Creating [HPM^{+}08, STS92, Och09, SSR02, Yad07]. Creation [Cha01b, Coc01b, CK06b, CK06c, Far23, Sin08]. Creative [Jam09, VH04, Ham07, Mor11]. creativity [Vai01]. Creator [Coc01a]. creators [Smi17]. Credit [EKJ^{+}03, Smi17]. Crediting [AKF21]. Crimes [Mar01]. criminal [Rau04]. criminal-political [Rau04]. criminological [Rau04]. Crisis [Les03]. Crispin [Neu84]. criteria [BH11, CWZ06]. Critic [Lew99a, Lew99b]. Critical [BdP13, Fit04, NK04, RDKT12, Sca19, Gar09, GEI^{+}11, JLL23, eLAA^{+}23, Pya06, dCdCM14]. Critical-Line [BdP13]. critique [WB07]. Cross [Gui00, HW17a, Mit94, Mit95, NRG^{+}99, Sch03, VOK^{+}22, XXCL19, CFW17, DuB02, GJS^{+}02, He95, Ho95, May17, NRRS20, RA16]. Cross-Architecture [XXCL19]. cross-compiler [DuB02, He95, Ho95]. Cross-Development [Mit94]. cross-language [May17]. Cross-layer [VOK^{+}22]. cross-machine [CFW17]. Cross-OS [XXCL19]. Cross-Platform [Gui00, NRG^{+}99, RA16]. Crowder [Ano00e]. Cruise [CTP^{+}22]. Crusade [Wil02, Cas02]. Crypto [Ahm08a]. Cryptographic [Gut00, ABF^{+}14, AV04]. cryptography [Gen99, McA08]. crystal [AZ17a, AZ17b, AFZ17, AFZ18, ATCZ19, FLA^{+}16, LZ11a, LZ11b, LZ12, Wen02]. crystallography [GFZ16, TV13]. CS [AFS81, AFS82, CWB^{+}04, EMdL^{+}07]. Cscope [KC92]. CSiBE [Bes04]. CT [Zha16]. CTG [PH16]. cubature [CDSV10, CDSV11]. CUDA [WGG^{+}19]. CUDA-based [WGG^{+}19]. CUG285 [HSC89]. CUG333 [Wit90]. CUG334 [Kot91]. CUG359 [FD92]. CUG368 [Ohl92]. CUG372 [Mor92]. CUG392 [Arc94]. CUIRRE [ZSW14]. Cults [Gla00]. Cultural [NS05, Rus09]. culture [DDA^{+}07, DD10]. Cultures [Gla00]. Cumming [Ano15a]. Curl [Ano01i]. Curley [Ano00e]. currency [Hol15]. Current [PKG^{+}10]. Curricula [CWB^{+}04]. Curricular [BCB07, CBB06]. Curriculum [BMB^{+}18, DDA^{+}07]. custom [PSSH16]. Customizable [Sta79, Sta84, TGC^{+}20, Sta80a, Sta81e, Sta81a, Sta81d, Sta81c, Sta81b]. Customization [FQYS23]. Customized [GKL^{+}14]. customizer [Rus88]. customizing [Rus88]. CVS [CdR99, Sim00, ZK05]. Cyber [PSSH16, AM18, dlVRB21, ZKDP22]. cyber-physical [dlVRB21]. Cyberpiracy [Rau04]. Cyberpiraterie [Rau04]. Cybersecurity [Lev23, VSN22]. cyberspace [Les99]. Cyclic [Due97]. Cygnus [Laz99]. Cygwin [Rac00]. Cyrix [Ano96c, Ano01c]. cytometry [MPE^{+}11]. cytoplasmic [SNC^{+}06].

D [BVLF14, Kuk98, Ano01j, CZS^{+}21, DO16, MGYC18, SDeaK^{+}09, Wen02, Wes00, ZK21]. daemon [Sal08]. Daily [BY14]. Dalek [OCH90a, OCH90b]. Dallas [ACM00, IEE95b]. damage [CBRSH22]. dance [DB05]. Danger [Cha01b]. Daniel [Ano00d]. Danish [Fre23, MG05]. dans [Séd02]. dansk [MG05]. DARPA [Coc01a, CGB^{+}05]. Data [AHB^{+}09, Ano96b, Ano00k, Ano01j, Ano02b, Ano04b, Bak20, CSD^{+}05, Dig82, Ell12, FY18, FMA02, Gil05, IAS16, Jan08, LLWM23, McC02b, Mor08, Noj01, QC18, SSC^{+}00, SC02, SSH22, Sta92c, Sta97c, SDD06, SC00, TRB22, WKB14, ALVV17, BGM99, CD95, DP09, EKUR10, Fra13, HFO^{+}12, HBB^{+}12, Hua23, Koc09, Lal91, LH14, Mas05, Mei92, MNS19, MGYC18, MGFRG12, MCQF21, MGR16, ODP15, QLC^{+}12, SCR05, Sta88d, Wii91b, YZC22, AAA^{+}12, Dat85, Kro00, Lin02a]. Data-Binding [Ano01j, Ano02b]. Database [Ano00e, Ano00d, Ano01i, Ano04b, ASAAM^{+}19, ABNÅ05, Bon02, CYOS19, EKUR10, MSZ^{+}01, Mon03, Phi93, Qui00, XTG^{+}11, Yad07]. Databases [CC03, Ell12, Mar01, Noj01, Pau04, Tan11b, Bon02, EKUR10, MQN19, PC13b]. Datacomm [Ano96c]. Dataless [CGB^{+}05]. Dataset [CGZ17]. datasets [CHE^{+}10, SKB23]. Dateien [DF00]. Datenbank [GGK99]. Datenverschlüsselung [Lin02a]. d’autore [Mol01, Zic01]. David [Fox08]. Day [Pow00, McC02b]. Days [Boy00]. DB [Yad07]. DBURLs [Tan11b]. DC [IEE89, IEE95a]. DCC [SC00]. DDD [MS08]. DDF [LQ17].

DDPredictor [HLS^{+}13a, HLS^{+}13b]. Dead [Hoh01, MI07]. Deadlock [GWT^{+}01]. Deal [Jon02]. Dealing [HR19]. Death [Mog99]. deb [VD01]. Debacle [Ian02]. Debain [Dum05]. Debate [CDsJ^{+}00, McL05, PLS^{+}91, DPL^{+}91]. Debian [Ano96c, Ano98, Ano00a, Ano00e, Ano01c, CK06a, CK06b, CK06e, CK06g, DF00, Her04, Kuk98, Ahm08a, Ano99b, Ano01b, Ano01d, Ano01c, Ano05b, Bau06a, BJM^{+}22, Bel00, BGO02, Bra04, CK06a, CK06b, Den99, EJS^{+}01, Epp00, G^{+}00, G^{+}02, GA04b, Gan04, GO99, HHV05, Hun01, Kra05, Mac02, Mac99, McC99a, Mur94, Pre16a, Pre05, Pre08, Pre16b, Ron01a, Ron01b, Ron05b, Ron05a, SS05b, Sta01b, VD01, WN15]. Debian-GNU-Linux-Powerpack [Gan04]. Debit [EKJ^{+}03]. Debsources [CGZ17]. Debt [PLO^{+}23, YXS^{+}19, HSX^{+}18, TKSC20, ZFD21]. debts [OMA^{+}22]. Debug [Per02]. Debugger [MZG14, OCH90a, OCH90b, Per02, Sta89a, SPS^{+}00, SPS^{+}02, But94, But95, Sta88a, Sta89b, SPG92, SP93, SP95, Sta96a, Sta98a]. Debugging [LL14, SP93, SP95, Sta96a, Sta98a, SPS^{+}00, SPS^{+}02, Zac01, MS08, Mit95]. Debuts [Ano02b]. DEC [AFS81, AFS82, Dig75a, PH82]. DEC-20 [AFS81, AFS82]. decade [Pes93, Sch09]. decades [CGZ17]. December [IEE92c, IEE94a, IEE05]. decentralised [PWA^{+}19]. Decentralization [TZ22]. Deception [CGK^{+}02, Sta02a]. Decimal [Bee17]. Decimal-arithmetic [Bee17]. decision [Höp04, SSS22, Wen90]. decision-making [SSS22]. Declarative [Dvo04]. Decomposition [BSA22]. Deconstructing [SBDR22]. Deconvolution [GF17, SDeaK^{+}09]. Decreasing [WM19]. Decryption [Bar00b]. DECsystem [Uni77]. DECsystem-10 [Uni77]. DECUS [Dig82]. deduction [Bun94, Kap92]. deductive [ABF^{+}14]. Deduplication [Gal10]. deep [ASAAM^{+}19, MLZ^{+}23, PNK^{+}23]. DeepFlame} [MLZ^{+}23]. DeepOtolith [PSP^{+}22]. DeepPlayer [PNK^{+}23]. Defamation [Ros02c]. Defect [KT04, Raj13]. Defects [UMV15, PdSCJM22]. Defend [CSD^{+}05]. Defense [CWB^{+}04, Bol02, Sca19]. Define [CSD^{+}05]. defined [SSS^{+}14]. Defining [Bar22, RT12]. definite [Yad07]. Definition [Per05, Sta96c, Lal91]. definitions [THG20]. Definitive [Fox08, WvH04, vHW03, vHW06]. Deformable [GKL^{+}14]. deformation [GBG^{+}16]. del [Les01]. delivering [Hen92]. delivery [SA15]. Delta3D [DMJ05]. demo [AAA^{+}12]. demystified [Sut02]. Deneb [YKK23]. Denning [CSP^{+}03, PLS^{+}91]. density [GCK^{+}17, HPT17, RAMB18, SHW^{+}21, SAHP15, THG23, WPAV14]. DensToolKit [SAHP15]. Denver [USE88, USE00b]. Department [Bol02, Sca19, BHP^{+}01]. dependability [LG02]. Dependable [EHP94]. dependence [CH06a, HMR93]. dependencies [PSL21]. Dependency [Gus20]. Dependent [HW17a, YSVM^{+}16, YSMA^{+}17]. Deploying [Maz15]. Depth [CWB^{+}04, SJW22]. Derivative [Maj03, Väl04, SAHP15]. description [Pyr84, SLC88]. descriptions [SC88a]. descriptor [Yap11, Cha91]. descriptors [GJMPAM^{+}14, Yap11]. Design [AML^{+}10, BGM99, Bar00b, Bax01, CFM08, CMJ^{+}04, DXT^{+}18, GCE^{+}21, IEE94c, Kro00, LOW91, MPG^{+}16, Mat03, Mio90, MEB^{+}20, Nov04, SKSM19, SFWD12, TMM^{+}13, Wal93, Bor88, FK99, For12, KP93, KSD^{+}12, PDG^{+}87, Pal87, PDG^{+}88, Váz16, Wii91a, Yan92, dA15]. designed [Mud97]. designers [Ham07]. Designing [Bar00a, DFCPSF15, Maz15, CG17, FL15]. Designs [Ano00j]. Desktop [Bra04, Dye03, EKJ^{+}03, LGW18, RB92, dILM98, PS^{+}09]. Desktop- [Bra04]. detailed [JD19]. Details [CRW^{+}04]. Detecting [GWT^{+}01, XTY^{+}22, ODP15]. Detection [Cha01a, CYL^{+}23, Kro99a, TBPS15, WM19, Wen00, XGF^{+}23, YDZ19, ASAAM^{+}19, AG22, HYA20, Joh94a, KOI94, OK94, SKB23, WCS20, Wan21, YLXZ16, YZC22]. detector [NRRS20]. Determinants [RH21]. Determination [YXS^{+}19, VRS^{+}95, VRS^{+}99a, VRS^{+}99b]. Determine [MOMM11]. Determining [Pow00, MD17]. deTestSet [MCS12]. detonation [SWTC23]. detonationFoam [SWTC23]. deutsche [Oms03]. deutsches [Hüp01]. DEV [vWHvW09]. Develop [RMAM19, MSS95]. developed [Fie90b]. Developer [Ano96c, Ano98, Bar01, Fei23, KJRD16, RM99, RAH^{+}01, Rie15, SBDR22, WFW^{+}20, DRM21, DWJG02, KC22, Sin10a, War04, ZVvDD11]. Developers [Ano95a, Col09a, FT09, Gau07, HDR03, HDR04, McL05, PLZ^{+}22, SV19, AMWH19, HNH03, JLH^{+}17, CCK21]. Developing [AHB^{+}09, Ano03b, Ano03c, BGR89, FK04, Iwa02, MH07, Men10, Mił10, Pet05, Tra95, CF07a, CH06a, YA11]. Development [ABC^{+}14, Ano00k, Ano01i, Ano01j, Ber96, BNSW15, BSA22, BH07, BCB^{+}17, CFM08, CPJ^{+}98, Coh03, Fox08, FCTP21, GAS^{+}01, Got07, Gui00, HOL^{+}07, HR11, HM19, Jin18, Kro99b, Kro99a, LHCH93, Lal91, LLWM23, LC12b, Mit94, NR03, NK04, O’Rxx, PG02, DARJ23, Rav00, RCP^{+}12, RCGB^{+}22, Rob20, RE04, Rui02, SBDR22, Sca04, SFF^{+}06, Sch03, SCSC04, SS04, Sta02c, SHK^{+}03, SF15, SuS01, Thi22, TDBEE11, WKS^{+}14, WDK^{+}20, WL01, XXAD21, Zha16, ZZZ22, vdLLM09, AW07, ATM22, And11, ASAB02, AHM^{+}07, BM06, BD03b, BAR16b, But94, Car89, CM06, CLM^{+}08, CH11, CF09, CSP09, CHA06, CWHW12, DB02, DTB05, EMdL^{+}07, Emb06, For12, Gar09, GHL^{+}04, Ger03, HOST05, HLL^{+}95, HWL^{+}23, IC22, Jør01, KMG^{+}07, LG02, LW03, LFB^{+}21, LRD^{+}19, Lov06, MG12, MFH02, MSR10, NGJ03, O’R99]. development [Pea16, Ped05, PH16, PT91, Pya06, Qui00, Rob05, SSAO04, Sca05, SA15, SNC^{+}06, SAOB02, SDD05, SG06, Thi99, Twi04, VMKB05, Yac88, You08, Zei03, ZE03, vKvH03, FN21, Kro99b]. developments [Ros05, Cse99]. Device [Bud10]. Device-driver [Bud10]. devices [BM22, KPK^{+}17]. DevSecOps [KG20]. d’exploitation [Cor00]. dfemtoolz [MVF20]. DFINT [QC18]. DG [YKK23]. DGDBM [Fra95]. DHTML [BHP^{+}01]. Diagramming [Kro99a]. diagrams [Ber22]. Dialing [DDJ98a, DDJ98b]. Dialogue [Lew99a]. DiBona [Ano99a]. DICOM [PBJ^{+}12]. DICOM-RT [PBJ^{+}12]. Dictionary [Ray91, Neu84, SGD05a, SGD05b]. dictionary-based [SGD05a, SGD05b]. Diego [USE94]. Dies [Coc01a]. diff [TACA15]. Differences [HW17a, LC12b]. Different [FCTP21, Tan11b, DC23, Mud97]. Differential [ALA20, LMW12, YLHW21, MZE13]. Differentiation [UNF^{+}08].

Difficulties [BBM^{+}21]. diffraction [HMYH22]. Diffusion [FRAK15, SS23, WFV14, WSK^{+}22]. Digest [IEE90, IEE92a, IEE92b, IEE93]. digit [PPR19]. Digital [CTP^{+}22, Gho07, Hef97, IAS16, Par03, QC18, Rus14, Thi22, Wat01, JP09b, Jon01, PPR19, Sin08, Zic01, PBJ^{+}12]. digitale [Zic01]. Dilemma [CGB^{+}05, Sil13]. Dimension [Ano96c, CGK^{+}02, JWC18]. Dimensional [PBJ^{+}12, BM22, CKS16, EHP14, MFB23]. dimensionality [SP12]. Dimensions [Sta97b]. Dinoflagellate [HW17a]. Dirac [MN21]. directed [Kli90, SMS04]. Directing [SM00a, SM00b, SM02, SM89a]. Directories [Ano95a]. Directory [Ano95a]. dired [STS92]. diritto [Mol01, Zic01]. disaster [CdSV07, LQ17]. disaster-tolerant [LQ17]. Disastrous [PFL^{+}12]. Discipline [Par03]. DISCO [Mio90]. discontinuous [FRAK15, HWL^{+}23, JRA^{+}18, RAW^{+}16, RHR^{+}21, ZJS^{+}20]. discourse [Kit94]. Discovering [TPSZ19]. discovery [Aki16, ALVV17, BH17]. Discrete [PSR16, DPH16, NMS14]. discussed [J^{+}05]. Disentangle [MTM^{+}19]. disk [Ano88a, Wil14]. disk-based [Wil14]. disks [TACA15]. disparate [Hin87]. Display [Ano97c, BVLF14, Sta79, Sta84, MYU89, Sta80a, Sta81e, Sta81a, Sta81d, Sta81c, Sta81b]. dispositions [Mol01]. disposizioni [Mol01]. Dispute [Sta04a]. Distance [Far06, MTBS09, WN15]. distortion [GV16]. Distributed [AY93, CD95, Cha01a, GF17, IEE92a, BM02, BD03b, CKH91, Fra95, KMG^{+}93, Kaw92, LQ17, MQN19, RAMB18, TTL06, VGD^{+}97]. Distributing [Bar01]. Distribution [Ano00k, Ano15c, MSSvK08, McC02b, Väl04, LH14, BVT06]. distributions [AKHG16]. Distributor [Rie21]. Disturb [WWSG21]. Disturbing [CKB^{+}05]. Diversity [IHSR19, AG22]. Divide [Wat01]. Division [GS12, GM94, GLT08]. DivX [Zic01]. DIY [Sea04]. DIY-IT [Sea04]. DJGPP [FP94]. DL [WBB01]. DLBFoam [TPK^{+}21b]. DMFTwDFT [SHW^{+}21]. Do [Gau07, LMZP19, OG07, PLZ^{+}22, SV19, WCG22, AKMS23, CWHW12, Fei23, CCK21, PM13, Sea04, STB23]. do-it-yourself [Sea04]. Document [Ano01i, Fur90, HKA^{+}19, JV01]. Documentation [CS95, CS96, CS99, CPJ^{+}98, KGM^{+}16, CS91, CS93, MGYC18, SC88b]. Documenting [Sta79, Sta84, Sta80a, Sta81e, Sta81a, Sta81d, Sta81c, Sta81b]. documents [Bec93, Iwa02, Rod00]. Does [WLC01]. Doing [Lev23]. Doku [Ano01d]. Dole [MSZ02]. DOLFIN.COM [Ano00k]. DOM [CP01]. Domain [Bar00c, Geh96, NO03, HKvH16, KD23]. domains [CZS^{+}21, PAB^{+}17]. dominant [JZ09]. Donald [Neu84]. donations [ZWU22]. DoNOF [PM21]. Don’t [ATM22, CWB^{+}04, CSD^{+}05, CGB^{+}05, DDJ99, EKJ^{+}03, WWSG21]. Door [Wea03]. Doors [Eri01, Jon01]. DOS-Programme [Str94]. Dose [PBJ^{+}12, Zha16]. Dose-Volume [PBJ^{+}12]. dot [KPK^{+}17]. Double [Wut12]. DoubleVision [Ano00i]. Down [NS01, Sav23, Ano03e]. downgrader [McH92]. download [Ano88a]. Dragonfly [AMR18]. DRAM [SJW22]. DRAMSys4.0 [SJW22]. draw [EMdL^{+}07]. drawbacks [Mud97]. Drawing [EGK^{+}02]. Dress [Ahm08a]. Drink [Sta06]. drinks [Ano03e].

Driven [Fav23, LFB^{+}21, MT94, Phi93]. Driver [Ano96b, Ano00k, Bud10]. Drivers [Ano97c, BMZ14]. Drives [Ebe07]. DRM [YKK23]. DRM-DG [YKK23]. drone [ASAAM^{+}19]. drones [Ros14]. Drops [EKJ^{+}03]. Drosophila [SNC^{+}06]. Drucken [Gün02]. Drug [Ano14, EKUR10]. Drupal [JP09b, WKS^{+}14]. DSP [Lea94]. DSPs [Ros02a]. Dual [Ano98, CH10, Sil13, Gar09]. Dual-use [Sil13]. DualSPHysics [CDR^{+}15]. dugksFoam [ZCG17]. Dumb [BHP^{+}01]. dummies [Bel00]. DuMu [KGW^{+}21]. duplicate [LZ12]. DVD [SuS01, Ken02]. Dynamic [LLWM23, LMZP19, MSM^{+}03, NS01, FN21, HFO^{+}12, KGT22, MLMFN^{+}15, Och12, TPK^{+}21b]. Dynamical [AG22, Dan11, Gv14, SHW^{+}21]. Dynamically [MZG14, Don04, HH88]. Dynamics [BSA22, BCB^{+}17, KCAS23, TDBEE11, Aki16, APK14a, APK14b, CFCA13a, CFCA13b, HPT17, JNN12, PHT17, WNS^{+}21, WSK^{+}22, Zag14, ZJS^{+}20].

E-Address [CDsJ^{+}00]. e-books [Sta01a]. e-government [NDDH^{+}21]. e-Voting [CKB^{+}05, PL05, ADF^{+}21]. Earlier [Sta12]. Early [Par03, VOM12, KTP95, SNC^{+}06]. Earthquake [PFL^{+}12]. Easy [CX23, Kre00, Yes12]. Easysoft [Ano00i]. echo [PSS^{+}07]. Eclipse [Bea21, DGC^{+}07, GP05, GHM^{+}05, MS08, Wol03a, ZK05]. ECOGEN [SPLD20]. Ecologists [Kri03]. Ecology [PMBM^{+}15, Eds16, WLD^{+}17]. Econometric [BD03a]. Econometrics [Edd00, Edd96]. Economic [Rie07, Rie10, Waa09, Ano01g, GLT08]. economics [AW07, Gla04]. economies [Zei03]. economists [YL08]. Economy [Gho07, May06]. Ecosystem [AWD^{+}18, KS11, Li18, Ron15, KHMA12, LGA20, VSGM14, dlVRB21]. Ecosystems [BKHT21, MTM^{+}19, TH04, CFL23, DRM21, FN21]. EDCC [EHP94]. EDCC-1 [EHP94]. edge [BCHR12, Kan12, Tor99]. Edgewatch [San98]. Edit [von88]. Edited [Ano99a, SD16]. Editieren [Str94]. Editing [Bk94, Bee91a, Bee91b, Fin91, KB90, San78b, SBA92, Sta78b, BK91, Bec93, FK90, Goe07, Kli90, PH82, Vie97, vdHGG^{+}13]. Edition [Ano00c, Ano00e, Kro99c, Ano00j, Ano11, Cha13, CS91, Eat00, G^{+}01, PKP02, PKP05, Teo13, Lio96]. Editor [Ano95a, Dig75a, Dig75b, DM97, Joh92, KK94, Kro00, McC99c, RAH^{+}01, Sta79, Sta84, von88, Cic78, Dat85, Gos83, Ham90, HK95, Man92, Sch91a, SLC88, CAC09, Sta80a, Sta81e, Sta81a, Sta81d, Sta81c, Sta81b, Tho90a, Tho90b, Uni85a, Uni85e, Uni85c, Uni85d, Uni85f, Ano99d, Bur04a, Ebe09, Gal60, Xia08]. Editorial [Ano01e, Ano01g, Eri99, Eri00, Eri01]. Editorials [SFF^{+}06]. Editors [Coh82, Far92, Par91, Val93, Fin80a, Fin80b, GM84, KP93, Ude89, Wat94, AM03, AM04, SS04]. EDRIXS [WFDK19]. Eds [Waa09]. Education [EXA^{+}05, HE17, HM19, FHH11, HETD09, LFB^{+}21, MTBS09, MWG08, MTD^{+}09, SSS^{+}14, YAS91]. education-friendly [MWG08]. Educational [Bro01, BB08, RP08, RT05]. Edwin [Bir93]. Effect [JK11, LO89, MTM^{+}19, AH19, Sin10a]. Effective [CFCA13a, CFCA13b, LBF^{+}22, Rob96, Rob97, SPDQ22]. Effectiveness [SK12]. effectivity [Höp04]. Effects [DKMB14, KGM^{+}16, AW07, Ham07, Jes03a, Koc09]. Efficiency [HW17a, Koc09]. Efficient [ASWD18, HMR93, Joh18, PKH07, TGC^{+}21, WK93, YLL^{+}07, Fär05, Hen92, IHBS14, KTP95, Kir12, KSS^{+}23, MBR21, MVF20, MGR16, MRN20]. Efficiently [MGM^{+}02]. Effort [CIC13, CFM08, GS12, Joh99, KS02, QB21, SS02, Asu05, KFYI13, RCGB^{+}22, Yu06]. eGovernment [CH06a]. EHTS [Wii91a]. Eiffel [CZ99, ZC01, CZ99, SO91]. Eigensystem [GBDM77, SBD^{+}76]. Eighteenth [Uni01]. Eighth [IEE92c, USE94]. Eilmer [GDJG23]. Einführung [CK06a, CK06b, CK06c, CK06d, CK06e, CK06f, CK06g, CK06h, PKP05]. einrichten [Ste00a]. Einsatz [SG05]. Einsteiger [Ron05a]. EISPACK [GBDM77, SBD^{+}76]. EJB [Kro99a]. elastic [HMX21a, HMX21b, RZWW23]. elastography [WHJ15]. Electromagnetic [LFN^{+}11, ORI^{+}10, SDL^{+}16, WGG16, WGG^{+}19]. electroMicroTransport [GDK21]. electromigrative [DSK19, GDK21]. electron [HMYH22, HPT17, MFB23, PHT17, SAHP15]. Electronic [Ano00i, Bax01, Fur90, Par03, Bea04, CSV^{+}07, DPL^{+}91, MSB09, Rud10, Sca05]. electronic-structure [MSB09]. electrophysiology [BSW^{+}14]. electrospinning [LPC^{+}15]. Electrotechnical [Yuk94]. Elektronische [Ano01b, Ano01d, Ano01c]. elektronisches [Rud10]. Element [HMX21a, HMX21b, LMW12, BSW^{+}14, CMC^{+}15, HMP^{+}15, KD23, MVF20, NMS14]. elements [ZK21]. ELF [SG99]. ELF90 [Ano96e]. Eliminate [Bar01]. Eliminating [GK92]. Elite [WFW^{+}20]. Elk [TF21]. ellipse [Fin22a]. Ellis [Coc01a]. ELS [YZC22]. Elsevier [TG15]. Elusive [Gla03a]. Emacs [Ano97c, Bla89a, Bla89b, Bor88, CR91, CR92a, CEL^{+}05, Che86, Che87b, Chi97, Fin80b, HNT93, Lew88, LLG90, LLG93, LLG94, LLSt99, LLSt00, Mit84, Rit88, RW89, Rod00, Rus88, Sta87, Str94, Uni85b, STS92, AFS81, AFS82, Buc82a, Buc82b, Com84, CCA84, Cic78, Dat85, Gos81, Gos83, KX86, Man92, McH92, PDG^{+}87, PDG^{+}88, She87, Smi90, Sta79, Sta80b, Sta80a, Sta81e, Sta81a, Sta81d, Sta81c, Sta81b, Sta84, Tra95, Uni85e, Uni85c, Uni85d, Uni85f, Ano96a, Aye97, Aye01, Bec93, Ber96, Bir93, Bra92, CRR96, Cam99, Cam00, Cha97, Chaxx, Cha01c, Cha04, Che87a, Chi93, Coo91, Cur02, Esp96, Fin80a, Fin91, Gli97, Gos84, Gre80, HH88, KK94, Lan89, Laz98, LLSt99, LLSt00, Li91, LO89, McA85, MS20, MC91, NS01, Pal87, P^{+}99, Pel89]. Emacs [Pyr84, Rei93, RB92, Rod00, RHS^{+}04, SBA92, Sta86a, Sta86b, Sta86c, Sta88b, Sta88c, Sta92a, Sta93a, Sta93b, Sta94, Sta95, Sta97a, Sta02b, Uni85a, Uni86, Vie97, Wel94a, Wel94b, Yac88, von88, Tho92]. Emacs-Based [LO89]. EMACS-Intr [AFS81]. Emacs-type [KK94]. Embedded [Ano94b, Ano01j, BES^{+}01, Hau01, ILG10, Jor04, Kro99b, LMPT22, Sut02, TTB09, BM06, KW94, KSS^{+}23, LLS11, Twi04, Yad07, Hau01, Jor04]. embedding [GCK^{+}17]. embrace [Sam06]. Embracing [Ano03e, CO12]. Emergency [CTP^{+}22]. emission [HFO^{+}12]. Emotional [Sta04a]. emotions [MOT^{+}18]. Empirical [AW07, Ano01e, Ano04c, Ban16, Ban17, CFM08, CWM^{+}20, CASA22, FM10, GFS05, KJRD16, MG12, MSSvK08, MRS07, RSBP23, ZXB^{+}23, AKHG16, AMC16, ACKT20, AKMS23, CLM^{+}08, FG92, HZS^{+}16, HM10, JK12, LGS^{+}17, LMZT22, CCK21, NXC13, OMA^{+}22, PKB17, PSE04, RNR17, RC10, SSA08, Tai13, THG20, TKSC20, VGP^{+}19, ZKDP22, ZFY^{+}19, ZD05, ZRGJ21, FMFZ19]. employing [LC12a]. empowered [MLZ^{+}23]. Empowering [MNS19]. Empress [Ano97c]. Emulating [Smi90]. Emulation [XXCL19]. Emulator [MZG14]. EMUstack [SDL^{+}16]. Enabled [FKM^{+}11, KG01, Bla06]. Enabling [Ano02b, BGG^{+}15, FQYS23]. Encoder [Kro00]. Encoding [CPJ^{+}98, Mud97, Lin02a]. encouraging [GSW08]. encrypting [Pow14]. Encryption [Ano86, Bar00c, PM00]. End [Ano00k, TGW^{+}22, Tho92, Tro05]. ended [YMCF23]. EndNote [Lie92]. Endorses [Sta04a]. Energy [BPG94, BMR^{+}23, VW92, GFD^{+}24, HMYH22, MBTB21, PCAJ^{+}23, ZPH^{+}15]. energy-and-bandwidth [ZPH^{+}15]. Enforceability [Rav00]. Enforcing [Mog01a, Mog01b]. engage [HETD09]. engagement [BAR16b]. Engaging [Nej12, Rob20, MdL09]. Engine [BM12, DMJ05, GEMN07, HKA^{+}19, ACM05, Wen02, AJLM18, Bur95, CP01, CBRSH22, Ert94, WGG^{+}19]. Engineer [Ano00c, GAS^{+}01]. Engineering [Ano96c, Ano00j, Ano01e, Ano04c, Bao93, Bea21, BPG94, Cha98, CC03, Fit11, IEE94a, IAS16, MCGA22, Tot06, ACM93b, Ano01g, CV13, FHH11, FFvdH01, FFHL05, FFH^{+}05, Ano01a, FMFZ19]. Engineers [CSD^{+}05]. England [IEE90]. Enhanced [Ano00i, McC05, SZAB98, SZAB99, EHHH06, HBC^{+}05]. Enhancement [HNT93, WRDP17, LA10, SSS22]. Enhancing [vWHvW09]. enjeux [Cor00]. Enlightenment [Fer03]. Enough [Bar00a, Wes03]. Ensure [CKB^{+}05]. Entangled [Bar00c, JWC18]. Enterprise [BVLF14, G^{+}06, Kop05, Kro99a, McC04, Sor06, SVAGB20, BH11, Dei10, Men12, RNR17, WG05]. Entertainment [GAS^{+}01]. enthalpy [ZK21]. Entire [Mof02, Mer03]. entrepreneur [Gil04]. Entrepreneurial [PRRL12, Far23]. Entropy [SSP17, SSP18, VB19]. Entrust [Kro99a]. Entur [RLTD23]. Entwicklung [O’Rxx]. Entwurf [Lei93b]. envelope [Fin22a]. envelopment [Koc09]. Environment [Ano01i, Ano01j, III01, Dye03, GEMN07, Har94, KP84, Kro99b, LHCH93, San78a, Ste00a, WKA^{+}08, dILM98, Ano96d, AHM^{+}07, Big13, Gad88, Gar09, GM84, GHL^{+}04, HH88, KORP95, KC92, KTTK17, McA85, PSS^{+}07, RT05, YM93, HOST05]. Environmental [EXA^{+}05, SSM^{+}07, DFCPSF15]. Environments [BY14, BSS84, CW15a, CW15b, VOM12, KK17, Oma89, Sim05]. Eof [Fri06, Bro04, Bro05, Coh03, DiB04, Lei04, O’S04, Ped05]. EP90 [Fur90]. Epilogue [TRM16]. epiretinal [CKS16]. Episodic [BKR^{+}20, BSFR22]. Epoch [OSM94a, OSM94b]. Epsilon [Joh92]. EPTs [GS02]. eQE [GCK^{+}17]. eqtools [CFW17]. equal [Raj13]. equals [Mur20]. Equation [eLAA^{+}23, HXS20, MBR21, ORS^{+}14, SM08, YSVM^{+}16, YSMA^{+}17, ZCG17]. equation-free [MBR21]. Equations [ALA20, LMW12, HWL^{+}23, MZE13, RJ21]. equilibria [CFW17, MZE13]. equipment [MGFRG12]. Equivalent [Coc01b]. Era [BGG^{+}00]. erfolgreiches [GGK99]. erklärt [J^{+}05]. ERMOCTAVE [MSC19]. ERP [BH11]. ERPs [SS06]. Erratum [VRS^{+}99a, VRS^{+}99b]. Error [Gla03a, Gla08, Kro99a, TV99, Wut12, SK12]. Error-Detection [Kro99a]. Error-Free [Gla03a, Gla08, TV99]. erstellen [Gün02]. ES-Kit [CKH91]. Escaping [May06]. eServer [AV04]. Essays [Fio03, Gay02]. Essential [Ano00k, ZK05, Har00, Bar16a]. established [Kle21]. Establishment [DSB^{+}16]. estimating [YZC22, YMCF23]. Estimation [WM05, Asu05, CIC13, QB21, RCGB^{+}22]. Estudio [RÓ01]. Etc [Pet06]. Etc/rant [Pet06]. EtherPage [Kuk98]. Ethical [GW09]. Ethics [EE01, KHA^{+}03, SHS^{+}93, Fei23, Mag01c]. EU [PSSH16]. Euler [HWL^{+}23]. Eulerian [LMHL20]. EUODHILOS [OSM94a, OSM94b]. EUODHILOS-II [OSM94a, OSM94b]. EuroFORTH [Ano94a]. Europe [PT91, Dre94]. European [Ano90a, EHP94, ACW04, DSB^{+}16, MTBS09]. EurOpen [Ano92]. EurOpen/USENIX [Ano92]. EUUG [Ano89]. Evacuation [CTP^{+}22]. Evaluate [VOK^{+}22, ALGE12, SG99]. Evaluates [Maj03]. Evaluating [CHE^{+}10, DDJ99, KGM^{+}16, RT12, GM02, GYW^{+}23, VSdCCR23]. Evaluation [ALVV17, BTL^{+}11, CWZ06, DKK22, NMX19, TRB22, ZRNA20, AMOS19, BH11, BC20b, Fug03, GGSRMPM20, GLCMC17, HK95, KMG^{+}07, PDG^{+}87, Pal87, PDG^{+}88, SCH^{+}91b, TPSZ19, YT22, ZW17]. Evangelizing [Coc01a]. even [SSAO04]. event [DPH16, FG16, Mei92]. event-logging [FG16]. Events [PFL^{+}12, ACM05, CBRSH22, Wii91b]. ever [Höp04]. everyone [JH16]. everything [LD13]. Everywhere [Ana99]. Evidence [GS12, Goo14, CV13]. evidence-based [CV13]. Evolution [MS20, PPRB07, Sca06, SDD06, Wal93, BLG^{+}17, CMTA19, DCS05, FRBRF19, GPPT16, JK11, Koc07, LGS^{+}17, NXC13, SCR05, SCFR06, WGS07, ZVvDD11, Zic01]. evolutionary [AFZ17, AFZ18, ATCZ19, FLA^{+}16, JCNS^{+}22, LZ11a, LZ11b, O’S03, WLD^{+}17]. evoluzione [Zic01]. evolve [AKMS23]. evolved [GL14]. evolving [Kle21]. Exam [AFS81, Ano00c, GM02, DRP01]. Examination [Mag04, AMC16, MSR10, RNR17]. Example [Wil99, Bud10, PSSH16, RHR^{+}21, THG23]. Examples [PKP05]. exascale [DFU20]. exascale-oriented [DFU20]. Exception [LZWH22, LRP21, Tie90]. Exchange [Boy13, PKGA22]. excited [RHW^{+}21]. Executable [GB20]. Executables [BHP^{+}01, DVC^{+}07]. executing [KTTK17]. Execution [AWD^{+}18, GKM82, GKM04]. Executor [Ano96c]. Exempt [Ano15c]. exercices [Rod00]. ExoLab [Ano01j]. Expanding [HM19]. expenses [ZWU22]. Experience [Bro19, III01, MSZ02, Raj23, San78a, SK04, Mei92, Pes93, RVLS14, RT05]. Experiences [AAB^{+}05b, GS02, LBF^{+}22, SS02, Tot06]. experiment [KTP95, MWB89, MHP94, WFF18]. Experimental [TRB22, Yan90, Amb15, BC20b, KSH14]. experimented [Fei23]. experiments [Adk11, CG17, Fei23, NMG11, WN15]. Expert [BPG94, Cla90, Rus88]. expertext [WRSG92]. Expertise [Col09a, CH11].

Experts [AS03, Sta04a]. Explained [Bal19, J^{+}05]. Explaining [Col09b, YYL^{+}15]. explicit [Car89, DDHS03]. explicit-state [DDHS03]. exploitable [ZLL04]. exploitation [Ron05b]. exploiting [Nob08]. exploration [VGP^{+}19]. Explorations [BGG^{+}15]. explorative [MSR10]. Exploratory [BNSW15, ZDM10, CNSR23, DD17, GW10, MOT^{+}18]. explore [Phi12, RCO20]. Explorer [Nas04]. Explorer‘s [Dum05]. Exploring [Goo14, HM19, KC22, Koc09, LMZT22, TPK21a, Blo04]. Expo [Ano06]. Exponential [Wut12]. Expose [O’S02]. Express [dlCKK15]. Expressing [LR11]. expression [MSZ^{+}01]. Extend [Ano03d, Mud97]. Extending [Che95]. Extensible [Sta79, Sta84, Ude89, CLL05, CFW17, HH88, Ham90, Sta80a, Sta81e, Sta81a, Sta81d, Sta81c, Sta81b, Tho90a, Tho90b]. Extension [GBDM77, BR95, Hin87, KX86, Sta78a]. Extensions [Gli97, Bec93, Rit88, RW89, Chi97]. Extensive [LMZP19]. external [THG23]. extra [Ano03d]. Extracting [CJ19]. Extraction [YDZ19, Kit94, KKA^{+}19, MAF22]. extreme [HXS20, HL02]. extreme-scale [HXS20]. Extremely [Gla03a, RBM^{+}23]. Eye [Jes03c, GW10, Hol05]. Eyelet [Ano00j].

F [UNF^{+}08]. f2c [Ano95c]. F77 [Ano95c]. F90 [Ano95c]. f95 [Ano01a]. FaaS [ES23]. Fabric [DXT^{+}18]. Face [AS03, Edw98, KS03, Law09, Sta96b]. Face-off [KS03]. faces [Cas19]. facilitate [HBR19]. Facilitating [Rav00]. Facilitator [Boy13]. Facility [Tra95, Yan92]. facto [Egy01]. Factor [Ano02b]. Factoring [Lók04]. Factors [MP12, Sca19, WCG22, CH11, KC22, LMZT22, LRD^{+}19]. Factory [DFP23, PSSH16, Ano01i, Kro99b]. Facts [ACHC11]. failure [CBRSH22]. Failures [Bar01, YYL^{+}15]. Faircloth [Ano11]. Fall [KOI94]. Fall-in [KOI94]. falsification [BA15]. Fame [Bar00c]. family [AVA^{+}16, GM84]. Fast [Bar00a, DXT^{+}18, Wen90, Yes12, CDSV10, CDSV11, Pax88, Pax95, PC13a, TPK^{+}21b, TG99, XAPK14, XOTI22, CKB11, IKW23]. fast-update [XOTI22]. Faster [Bar00c]. Fault [IEE90, IEE92a, IEE92b, IEE93, WM19, YYL^{+}15, GFS05, KTP95, LH14, Wan21, Yad07, YLXZ16, YZC22]. Fault-Tolerant [IEE90, IEE92a, IEE92b, IEE93, Yad07]. faults [OMA^{+}22]. FCS [QLC^{+}12]. FCSTrans [QLC^{+}12]. FDTD [ORI^{+}10, WGG^{+}19]. Fear [Mog03c, Ros01d]. fears [Ray01a]. Feasibility [SC02]. feather [Dew07]. feature [HZ14, KKA^{+}19, PdSCJM22]. Features [EKJ^{+}03, Bau06b, Bau06a, Fri97, GRJS01, HMO^{+}18, Phi12]. February [CS95, Lea92, Sta92b, USE02a]. Federal [BSK87]. Fedora [Men12, McC04]. feedback [RA16]. Feel [PM00]. Feinberg [Ano00c]. Felix [Ano00l]. FEM [HMX21a, HMX21b]. Femtoseconds [Bar00a]. Fenglin [FQYS23]. Fenglin-I [FQYS23]. FEniCS [Kam21, LMW12]. Fermi [MN21]. fermionic [XOTI22]. Fernandez [Ano00l]. Festschrift [Oms03, Oms03]. FESTUNG [FRAK15, JRA^{+}18, RAW^{+}16, RHR^{+}21]. Fewer [Bar00a]. FF [GNR^{+}09, KMF^{+}07, PMG^{+}09]. ffLink [dlCKK15]. FFT [DO16, Fär05].

FIASCO [Haf01]. Field [Bar00a, EW01, SHW^{+}21, Bow05, Gv14, PKH07]. field-sensitive [PKH07]. fields [RHW^{+}21, SAHP15]. Fifteen [Gri16]. Fifth [FP95, MSNS91, SS93, SM89b]. Fight [Rau04, Sta00a]. Fighting [DDHS03, Mak04]. Figlet [Ano10]. filament [GF17]. File [Cha91, PPR19, BGM99, LQ17, QLC^{+}12, RAMB18, STS92]. Files [Kro00, Fra95, MRS07]. Filesystem [BS98]. FileZerver [Ano00j]. filing [Bea04]. filling [Fel93]. Film [KMF^{+}07, Row02, Jes03a]. Film-Based [KMF^{+}07]. filopodia [UBR^{+}17]. Filopodyan [UBR^{+}17]. filosofia [Cor05]. filtering [CV22, MGR16, Shi12]. Filters [NRG^{+}99]. Final [AFS81, Ano01c]. financial [Hic04]. Find [Wea03]. Finder [LOW91, LO92]. findOSSLicense [KC21]. fine [MT94]. fine-grain [MT94]. fingerprints [Yap11]. Finite [HMX21a, HMX21b, LMW12, Zag14, BSW^{+}14, Mak03, MVF20, ZK21]. finite-element [BSW^{+}14]. Finkel [Neu84]. Finnish [Ano00j]. firewall [Pot06]. Firm [CFMRL11, JZ09, WG06]. Firms [CH10, MD04, Rie20]. Firmware [CYL^{+}23, AHM^{+}07, Fra19, GHL^{+}04, Sut02]. First [CCSW10, IEE94a, EHP94, Höp04, KD23, MBTB21, Sch09, SG99]. first- [KD23]. first-ever [Höp04]. Fish [DDJ99, PSP^{+}22]. fisheries [HBB^{+}12]. fishing [BMS^{+}22]. Fitting [Bur04a]. Five [ACHC11, FRBRF19, MRS07]. five-year [FRBRF19]. Fixing [Jes03b, ACB18, FM10]. Fjords [Coc01b]. Flame [DGJH19, DD10]. Flash [Ano03d]. flavors [Fri97]. Flaws [SV03]. fleurons [Wil13]. Flex [Nic93, Pax95, Pax88]. Flexible [SAC^{+}15, Bor88, ORI^{+}10, SC16, YMCF23]. float [Abb12]. Floating [FL16, He95, Ho95, FHL^{+}07]. Floating-Point [FL16, FHL^{+}07]. Florida [SS93]. FLOSS [CIC13, GS12, HBR19, O’S03, YA11]. Flow [CSP^{+}03, Gol06, MEB^{+}20, ACW04, BZB17, BSC^{+}21, GDJG23, HSF^{+}15, JD19, KGW^{+}21, MLZ^{+}23, TPK^{+}21b, YKK23]. flows [CAWK22, CFCA13a, CFCA13b, FTZ^{+}23, LMHL20, SPLD20]. FLR [KMG^{+}07]. Fluid [TDBEE11, ZAC^{+}23, CFCA13a, CFCA13b, DIK^{+}23, Kam21, LH22, DARJ23, WNS^{+}21, WSK^{+}22, Zag14]. Fluid-Structure [ZAC^{+}23]. FLUNED [DARJ23]. fluorescence [BDAW15, GF17, SDeaK^{+}09]. flux [MBTB21, WFV14]. fluxes [ORS^{+}14]. fly [BGM99, RBM^{+}23]. foaming [ASC^{+}21]. foams [KDM17]. focus [KGW^{+}21]. Foe [Wel94a]. Fog [ZZZ22]. folding [ZJS^{+}20]. folks [GMPS14]. Follow [CKB^{+}05, GAS^{+}01]. font [Wil13]. fonts [Ano03d, Wil13, Wil13]. Fookes [Kro00]. Foot [Wea03]. force [YKSH20, Kro00]. forces [Ano00f]. forciert [Ano00f]. forecasting [TKSC20]. Forensic [IAS16, QC18]. Foresighting [YA11]. ForeSys [Ano96e]. Forge [Kuk98]. fork [GL14, JLH^{+}17]. form [HKY^{+}21, ZK21]. formal [BCPS10, GB20, PT91]. Formalizing [RW87]. formally [MRH23]. Format [CS95, CS96, CS99, CS91, CS93, SC88b]. formats [CF07b]. FORML [Ano94a]. formulation [jFFR16, JRA^{+}18]. Förster [BMR^{+}23]. Fortgeschritten [Ron05a]. Forth [Ano94a, Ert94]. Fortran [Ano96b, Ano01a, AS97, AG95, Ano96e, Ano97d, Ano01a, Bad07, Bra97, Bro03, SZAB97, SZAB98, SZAB99, UNF^{+}08, YSVM^{+}16, YSMA^{+}17, Zag14].

FORTRAN77 [But95]. FortranPlus [Ano01a]. Fortune [Pra03]. Forum [CGK^{+}02, CSP^{+}03, CWB^{+}04, CMJ^{+}04, CRW^{+}04, CKB^{+}05, CSD^{+}05, CGB^{+}05, San78b, Sta78b, Sta02a, Sta03c, SHS^{+}93]. forward [KSV16]. Foss [MdL09, ASWD18, Bol02, BMB^{+}18]. FOSS4G [MS12, BK14]. FOSSES [AMOS19]. FOSSIL [ASWD18]. Fotos [DF00]. Fought [Kos21]. foundation [BYV08, Ano01i, Bro05, CJ17, CRW^{+}04]. Foundations [Rie10, You08]. Four [Van22, Cre07]. Fourier [JP09a, Wut12]. Fourth [Ano88b, Ano90c, MS91]. FPGA [CCA^{+}13, LGW^{+}22, ZKCS91]. FPGA-based [CCA^{+}13]. FPGAs [FL16, MEB^{+}20, TGC^{+}20]. Fractal [Haf01]. fractional [CZS^{+}21]. Fragen [Sie99, Stö04]. fragility [CMTA19]. FrameMaker [Ano02b]. Frames [Men10]. Framework [AMOS19, Ano01j, BMF^{+}16, CBB06, HMKC12, HS15, JPOB20, MSC19, MMD12, MK12, Mor08, PMBM^{+}15, PK10, PGC21, SAC^{+}15, SJW22, Sto09, TGC^{+}20, WDK^{+}20, BC20b, CMC^{+}15, CBRSH22, DP09, HWL^{+}23, HPT17, Hub04b, JNN12, KMG^{+}07, KSH14, KH05, KGT22, KSS^{+}23, MVF20, MGFRG12, PCAJ^{+}23, PHT17, SSR02, WFV14, ZAC^{+}23, ZLF^{+}22]. frameworks [FRBRF19, IDSM23, May17, YWA07, YT22]. France [Bun94, IEE93, QR92, DMP^{+}02]. Francisco [ACM92, Lei93a, USE02a]. fray [Sch09]. Fred [Bar00b]. Free [Ano86, Ano96e, Ano97d, Ano99c, Ano99d, Ano01a, Lin02b, Ano10, BBD^{+}96a, Bal19, Bar00a, BKR^{+}20, BSFR22, BFC02, Bon11, Bra97, Bro96, CPJ^{+}98, Coc01a, CPG^{+}04, Col09a, CK10, CGK^{+}02, CGB^{+}05, Cur02, DDJ98a, Del01, DM15a, EXA^{+}05, Ell12, Far91, Fav23, FFHL07, Fie88, Fie90a, Fie90b, Fio03, Gal10, Gay02, Gla03a, Gla08, GBICMR13, GB21, GM05, Gre18, GW09, mH00, Hal02, Hug95, Jam09, Kam14a, Kam14b, Keh94, KS11, Kro00, Lam09, Luc99b, May06, Mee12, MTBS09, Mic04, MN04, Mog01c, Nej12, O’S04, Omb20, Ous99, Pom04, Pri19, Rad89, Rus09, Sca04, SFF^{+}06, SB08, Shi12, DDJ98b, Sta96c, Sta98b, Sta02a, Sta04b, TV99, Tay00, Tro96d, Tro96a, Tro96c, Tro96b, Tro97, Ude97, Whe03, Wil02, dA15, Bab02, BVT06, BMR^{+}23, BAE14]. free [Bro04, CGZ17, Cas02, Col09b, Cor05, CHA06, CWZ06, Cus04, DFCPSF15, DB05, DM15b, Eds16, Fog06, jFFR16, FG16, GJMPAM^{+}14, GFZ16, GW10, Hea09, KJ03, Kop05, MG12, MBR21, MTD^{+}09, O’S03, ORI^{+}10, PH16, PWA^{+}19, RCGB^{+}22, Ros05, Sal08, Sca05, SSA08, Sta09, Ste08, SG06, WPAV14, XTG^{+}11, YA11, Zag14, Zic01, CPG^{+}04, Ano01i, BES^{+}01, Bol02, Bro05, CWHW12, Jak03, Mog99, Mog01a, Mog01b, Mol01, MS12, NR03, Séd02, Sie99, Sie04, St.04, Stö04, Cas02]. free-software [GJMPAM^{+}14, ORI^{+}10]. Free/ [CWHW12]. Free/Libre [BKR^{+}20]. Free/Libre/Open [BSFR22, O’S03, YA11]. Free/Open [Lam09, SFF^{+}06, CWZ06, RCGB^{+}22, Sca05, SSA08, SG06, NR03]. FreeBSD [Ano00i, And01, Coc01a, DTB05, GS12, Jør01, YSC^{+}06]. Freedom [Cha01b, EKJ^{+}03, Wil02, Cas02, Mar05, Sta01a, YL08, Jak04]. Freely [GM02, ODP15]. Freemont [Ano00j]. Freenet [Bar00a]. Freenets [Hug95]. FREENIX [USE01b, USE02b, USE98a]. FreeRTOS [GPPT16]. Freetype [Ano10]. Freeware [Edw98, Geh96, Ude97, Fri97]. freie [Lin02b, Jak03, Sie04, Stö04]. Freier [Sie99]. Freiheit [Jak04]. French [Ron05b, Séd02]. Frenzy [GWT^{+}01]. Frequency [PSR16, Blo04, KD23]. Frequently [And03]. FRET [BMR^{+}23]. FRET-Calc [BMR^{+}23]. freundlicher [Oms03]. Friedman [Aji17, Ano00c]. Friend [Wel94a]. friendly [MWG08, Oms03, Sch90b]. front [Tho92, Tro05]. Frontmatter [ACM05]. Fronts’ [Sta03a]. FS [Whe03]. FSF [DDJ99]. FTCS [IEE92b, IEE93]. FTCS-22 [IEE92b]. FTCS-23 [IEE93]. FTM [MHP94]. fuel [BCP^{+}16, HMP^{+}15, NGCI^{+}12]. Fujitsu [Ano01a, YM93]. Full [RSKF96, Dan11, LSF94, Phi12]. fully [FBY^{+}17]. Fun [GAS^{+}01, Pri19, Ros14]. Function [Ano15c, Cod75, Wut12, CYOS19, RC10, Sch90a]. Functional [ACM92, Cou13, HW17a, SDD06, GCK^{+}17, HPT17, PM21, RAMB18, SHW^{+}21]. functionality [Ano03d, SRGCPB^{+}09]. Functions [ASWD18, CYOS19, LN92, MBR21, Mer03, Neh04, Neh07, PHT17, SG99, VRS^{+}95, VRS^{+}99a, VRS^{+}99b]. Funds [Cha98, Coc01a]. FUNPACK [Cod75]. further [Ano02a]. fusion [GTMR23]. Future [LLdI00, MCGA22, San98, SZAB98, Sta04a, CK08, PWA^{+}19, Tay19, WCS20]. Fuzz [MZH22].

G [Ano00c, KY16, Men12, Gou04, Sid03, Ano04b, MGPB20]. G-air-simulator [MGPB20]. g77 [Ano95c]. Gaelyne [Ano00d]. Gaining [WG00]. Gaithersburg [Fur90]. Galerkin [FRAK15, HWL^{+}23, JRA^{+}18, RAW^{+}16, RHR^{+}21]. Game [BM12, DMJ05, Sca04]. Games [CSP^{+}03, GAS^{+}01, SuS01]. Gap [San08, Joy09b]. GAR [Mof02]. Garage [Pra03]. Garanties [Séd02]. Garbage [CR92b]. Garcia [Ano00l]. gas [FTZ^{+}23]. gaseous [SWTC23]. Gasson [Ano00d]. GATE [GLMS18]. GATE-based [GLMS18]. Gateway [Ano00j, AML^{+}10]. Gauss [Joh18]. GaussDal [ABNÅ05]. Gawk [Gor96, Anoxx, Clo89, Clo91, GAW87a, GAW87b, Lou96, Wit90]. Gawky [SNC^{+}06]. GCC [Sta00c, Ave06, BHP^{+}01, Bot03, Bro03, DC23, Dvo03, Edd96, Ede04, FMT^{+}08, FKM^{+}11, Gou04, Gri02, GJS^{+}02, Gui00, Hag04, HC07, Hub03, Hub04b, HDR03, HDR04, Jon05b, KSK09, Lat03, LC12a, Liu06, Lók04, Mak03, Mak04, Mat03, MS00, MRS07, Nai04, Nov03, Pen03, Pun04, RMAM19, RVLS14, Sha03, Sha04, Sid04, Sta99, Sta03b, SW15, Ste01, Tro05, Wag03, WvH04, Wei03, WRDP17, Xia08, ZRGJ21, vHW03, vHW06, Bee17, Bow05]. GCC-NMF [WRDP17]. GCC185 [BSP11]. Gcj [Tro04]. Gcom [Ano00j]. GCompris [CK06h, CK06g]. GCTP [Sif00]. GDB [MS08, MZG14, Mit95, Sta88a, Sta89a, Sta89b, SPG92, SP93, SP95, Sta96a, Sta98a, SPS^{+}00, SPS^{+}02, But95]. GDPR [VSdCCR23]. GDPR-compliant [VSdCCR23]. ge [Gün02]. ge-packt [Gün02]. GEANT [FP94]. gear [BMS^{+}22]. Geek [Cas19, Ros01a]. gekonnte [Ste00a]. Gelato [Ano06]. Geld [Cap12]. Gem [Bri09a, Bri09b, Och09, Och12, Rog09a, Rog09b, Rog11]. gems [Ude97]. Gen3 [dlCKK15]. Gender [IHSR19, STB23]. Gene [MPG^{+}16, MSZ^{+}01]. GeNeDA [MPG^{+}16]. Genehmigung [Oms03]. General [Dat85, Hüp01, MG05, Neg15, O’S03, TGC^{+}21, USE02c, Väl04, EKUR10]. générale [CF98]. generated [ALGE12, Wen90]. Generating [LGW18, AZ17a, AZ17b, KN93]. Generation [GAS^{+}01, JPOB20, Lat03, MZG14, BD03a, CKGW22, JP09a, KSK09, Wen90]. Generator [DS88, DS99, DS00, DS02, Mey18, Nic93, AHG94, BCR^{+}08, CLS95, DS90, HMP^{+}15, Mor91, Pax88, Pax95, Sch90a]. generators [Sib17]. Generic [Ano96c, CWM^{+}20, Dav91, Rob05, Mer03]. genesis [TV13]. GeneX [MSZ^{+}01]. GeNUAdmin [Har94]. Geoffrey [Neu84]. geographic [Eds16]. Geometric [WCA^{+}14, LH22]. geometries [SS23]. Geometry [MK12, SKSM19, CZS^{+}21]. GeoPDEs [Váz16]. GeoRef [Lie92]. Georgia [Dig82, USE00a]. Geospatial [BK14, MS12, PFL^{+}12]. Geostatistics [Pil09]. Gerätetreiber [Bud10]. Gerätetreiber-Reengineering [Bud10]. German [Ano00f, Lin02b, Bra04, Bud10, Cap12, CK06a, CK06b, CK06c, EW01, Fal03, Feixx, GA04b, Geh96, HK03, Hüp01, J^{+}05, Jak03, Jak04, Jor04, Lin02a, Mag04, Man03, NO03, O’Rxx, Oms03, PKP05, Rau04, Ron05a, Rud10, SG05, Sie99, Sie04, Spe01, Spi03, Stö04, Sur01a, Sur01b, WP04]. Germany [BPG94, BSK87, EHP94, EKR91, Ano01f]. Gers [QR92]. Geschichte [Geh96]. Gesichtspunkte [Stö04]. gestalten [DF00]. Gestaltung [FG85]. gestión [VD01]. Get [ATM22, Coc01a, Fie89, Sid03, Sta06, Ste93, Far91]. Gets [AS03, CWB^{+}04, Sta06, Gen99]. Getting [McA19, Nob08, PM00, Woe94a, ACM93a, Ano08b]. gewährleistungsrechtliche [Stö04]. GFX [Row02]. GG [PKG^{+}10, WKA^{+}08]. Ghosh [GAS^{+}01]. Gift [Zei03]. GIMP [DF00, GGK99, Neu00, Bun00, Bur04b, Bus99, Goe07, Ham07, Har00, JP09b, Jes03a, Jes03b, Jes03c, KK99, LD13, Pec08, Row02, vGS10, DF00, Log99]. GIMP- [DF00]. GIMPLE [Mer03]. GIS [FVD^{+}12, Lit14, MN04, Pil09, SNF04]. Git [BBM^{+}21]. GitHub [Far23, GMBv20, JLH^{+}17, KKN^{+}21, LMZP19, MNS19, WBRH23, ZWU22, LPFD21]. Give [CSP^{+}03, Smi17]. Given [Den99]. Giving [Bro19]. Glasgow [Ano14]. Glibc [Gar00]. Glisterings [Wil13]. Global [Ahm08a, BB08, BK14, De’15, FVD^{+}12, Uni01, Ano99c, Ger03, Lla06, MG12, Ano98]. Glow [CK10]. Glue [Car89]. GMP [BMZ02]. GNARL [GB94]. GNAT [CFCA13a, Fly87a, Fly87b, Ano95d, BOM97, Big13, Bri09a, Bri09b, CFCA13b, CGS94, CDG97, Dew07, GS02, Gre14, JD19, KTP95, Kir12, MB98, MGM^{+}02, MSM^{+}03, Mir03, MSK05, Mir07, MD22, Och09, Och12, PG02, RTH15, Rog09a, Rog09b, Rog11, Rui13, RSZ96, RSKF96, Sch10, Smy97, VGdlP01, dlPRGB99, Shi03]. GNAT-AJIS [Och09]. GNAT/ORK [VGdlP01]. GNATProve [Kan12, HMW15]. GNATS [Plo97]. GNATTest [Kan12]. GNL [NN16a, NN16b]. Gnome [LR11, Cro00, Dye03, EKJ^{+}03, GM84, Ger03, GWT^{+}01, UCLxx, Ben78, KS02, LLdI00, Lov06, Pet05, Pin02, SG99, Ste00a, War04, Wri00, dILM98, CP01, She01, VSGM14]. GNOME/GTK [Cro00]. Gnomes [Ano05a]. GNU [Ano97b, Ano97a, Ano98, Ano00e, Ano01b, Ano01d, Ano01c, Lin02b, Ano05b, Ano15a, Bra04, Bud10, Chi97, DF00, FRAK15, G^{+}00, G^{+}02, GA04b, Gan04, Gün02, HHV05, Hüp01, Jor04, Kuk98, Mac02, Mag04, Per02, PKP02, PKP05, Ron01a, Ron01b, Ron05a, SW15, Ste00a, SuS01, Ygg93, Ygg94, Ano97c, Fis69, Sal08, Ahm08b, AS97, AAB^{+}04, ACW04, Ana99, And11, Ano93a, Ano93b, Ano95b, Ano95c, Ano95e, Ano95f, Ano95g, Anoxx, Lin02b, ATHW92, Arc94, Avi98, Aye01, BGM99, Bad07, BD03a, Bak20, BM06, Bec93, Bee01, Bel00, BBM^{+}21, Bla89a, Bla89b, Blo04, BGO02, Bra92, Bud10, Bur95, BS98, But95, Cal10, CR91, CR92a, CRR96, Cam99, Cam00, CEL^{+}05, CS91, CS93, CS95, CS96, CS99, Che86, Che87a, Che87b, Chi93, CF98]. GNU [Coc01a, Col02, CZ99, CGS94, Cor00, Cor05, Cou17, DRP01, Deo90, DC00, Don04, DDHS03, DuB02, Dum05, Eat97, Eat00, Eat02, Eat05, EBH08, EaoGOBHW14, EGH^{+}05, Esp96, Fre87, FY18, Fär05, Fie89, Fin22a, Fin22b, FG92, FD92, GIM07, G^{+}01, GDT^{+}02, GDT^{+}05, Gal09, GA04b, Gar00, GB94, GHL^{+}04, Gil88, Gli97, GO99, Gol06, Gou04, GK92, GSR^{+}04, GS00, GvdHPR14, HWZxx, HWZ01, HH88, Ham99, HNT93, He95, Hei16, Ho95, Hol05, Hom00, Hun01, Hüp01, Ing92, Jan01, JRA^{+}18, Jen97, Jon05a, Jön05c, Jor04, Ken02, KSV16, Kre00, Kro99c, Lan89, Lea88, Lea92, Lea93, LZ16, LZ17, Leó98, Lew97, Lew88, LLSt99, LLSt00, LMOS93, LSMO96, LSM^{+}99, LSM^{+}00, LSM^{+}01, LS04, Loo15, LO97, Mac99, Mag00, Mag01a, Mag01b, Mag01c, MS02, Mag04, Mah13]. GNU [Man00, MG05, McC99a, Mec05, MSS95, Mit94, Mit95, MC91, Mur94, NR03, Neg15, Nor02, O’S03, OSM94a, OSM94b, OCH90a, OCH90b, Pel89, Lás05, Pes93, PKP05, PSR16, PR96, Rac00, Ram94a, Ram94c, Ram94b, RAW^{+}16, RHR^{+}21, Rit88, RW89, Rob94a, Rob94b, Rob94c, Rob94d, Rob95a, Rob95b, Rob95c, Rob96, Rob11, Rod00, RÓ01, Ron15, Ron05b, Ron05a, Rui02, Sai01, Sai02, SS05b, SBA92, SZAB97, SZAB98, SZAB99, Sha05, SPDQ22, Smy97, Sta85, Sta86a, Sta86b, Sta86c, Sta87, Sta88b, Sta88c, Sta88a, SC88b, Sta88e, Sta89a, Sta89b, SM89a, Sta92a, SPG92, Sta92b, SP93, Sta93a, Sta93b, Sta94, SP95, Sta95, Sta96a, Sta97a, Sta98a, Sta99, SPS^{+}00, SM00a, SM00b, Sta00b, SPS^{+}02, Sta02b, SM02, Sta03b, SMS04, SNF04, Sug02a, Sug02b, SGD00]. GNU [Tan11a, Tan11b, Tho92, Tie88, Tie93, TG99, Twi04, UZ97, VD01, Väl04, VETT00, VWM98, Vie97, Vol89, Vol96, WB07, Wal99, WY94, W^{+}95, Wes00, Wil14, Woe94a, Woe94b, Wol04, Wol03b, Yac88, YL08, YSMA^{+}17, Zac01, Zad02, ZC01, ZPH^{+}15, Zim10, vdH04, vdHGG^{+}13, Aye97, Ber96, Sec95, Sta00c, Zic01, Ano00a]. GNU-ADA94 [CGS94]. GNU-Aid [Deo90]. GNU-Emacs [Lan89, Rod00, Rod00]. GNU-Linux-Zertifizierung [Mag04]. GNU-Softwaretools [Jor04]. GNU/Cfengine [BS98]. GNU/Linux [Ano98, Ano00e, Ano01c, Lin02b, Ano05b, G^{+}02, GA04b, Gün02, HHV05, Kuk98, Ron05a, Lás05, RÓ01, ACW04, Ana99, BGM99, Bel00, BGO02, Cor00, Cor05, Don04, Dum05, GA04b, Gar00, GO99, Hom00, Hun01, Jon05a, Ken02, Mac99, Mag04, Man00, McC99a, Mur94, NR03, Nor02, Ron05b, Ron05a, SS05b, SNF04, VD01, YL08, Zac01, Lin02b, Cor05, RÓ01, Zic01, Ano00a, Wil13]. GNU/Linux-Anwenderhandbuch [Ano01c]. GNU/Linux-Systeme [Gün02]. GNUC [HS15]. GNUlib [Ohl92]. GnuPG [JKS02]. Gnuplot [Phi12, Kot90, Kot91, Ste93, WKC^{+}90, AG95, Jan08, Vau96, Rac06]. GNUPro [Ano00j, Laz99]. GNUstep [Mar03]. Go [AS03, BlG12, Dew07, NN20, Ano98]. GO-Global [Ano98]. goals [RSZ96]. Goes [DDJ98a, DDJ98b, GMPS14]. Going [Bon02, Rui13]. Gold [GB00, TGW^{+}22, Cap12]. Golgi [BSP11]. Good [Joh02, McL05, KS03]. Goodfellow [Neu84]. Gopher [Mam01]. GOTO [Nag18]. Governance [APCs22, CFM08, DFT21, IC23, McA19, HBR19]. Government [Coc03, GGB17, SV03, Wea03, NDDH^{+}21, Sim05]. Governmental [CPJ^{+}98]. Governments [AS03]. gperf [Sch90a]. GPGPU [BGG^{+}15]. GPL [J^{+}05, Jak03, MG05, Gal04, Guy00, Höp04, J^{+}05, Jak03, KS03, Mog01a, Mog01b, Mog03b, Pow00]. gprMax [WGG16, WGG^{+}19]. Gprof [GKM82, GKM04, GJLT11]. GPU [DFU20, SAC^{+}15, TL17, WGG^{+}19]. GPU-accelerated [TL17]. GPUs [ZSW14]. Graduate [Kos21]. Grafik [PKP02, PKP05, Str94]. Grafiken [DF00]. Graham [Ano15a]. Graham-Cumming [Ano15a]. grain [MT94]. grammars [EKR91]. Grand [Bao93, GB00, Ten93]. Graner [Cha13]. Grant [Cha98, LLEL^{+}23]. Graph [EKR91, EGK^{+}02, GKM82, GKM04, Hub04b, Mat03, Sch91a]. graphic [Tho90a, Tho90b]. Graphical [CGK^{+}02, BBNP93, Fin22a]. Graphics [Ham99, Kro99b, Kuk98, NRG^{+}99, PKP05, Wal99, Wen02, Wes00, Bus99, Lin00, Sta78a]. Graphing [Vau96, Phi12]. Graphite [Ano10]. GraphOn [Ano98]. Graphs [Jan08, HMR93, Sha05]. Graphviz [EGK^{+}02]. GRASS [MN04, SNF04]. Great [Bus99, RAH^{+}01, Ano01i]. greater [SSAO04]. Greek [KKA^{+}19]. Green [CPJ^{+}98, CYOS19]. Greenstone [WBB01]. GRef2End [Lie92]. GRETL [BD03a, Adk11]. Grid [BE06, Bor09, BE06, HOL^{+}07]. grids [LH22]. Griffin [Ano00l]. Groff [Ano10, RAH^{+}01, Rob94b]. Grokking [Bun00]. Gross [YSVM^{+}16, YSMA^{+}17]. Ground [WGG16]. Group [BSA22, Sif00, SM89b, XWZ^{+}23, KP93, WPAV14, Ano01j]. Groups [Men10]. groupware [Mar11]. growing [KHMA12]. Grown [CPG^{+}04]. growth [GLT08, WNS^{+}21]. Grundlagen [DF00, G^{+}00, G^{+}02, GA04b, Str94]. GSAS [TV13]. GSAS-II [TV13]. GSE [HM19]. GSL [G^{+}01, Rap94]. GTK [DF00, Ste00a, Cro00, Wri00]. GTK-Versionen [DF00]. guarantee [Stö04]. Guarantees [Séd02]. Guarding [Mah03]. Guest [Ano01g, SFF^{+}06, Ebe09, SS04, Xia08]. GUI [Ano00j, CMTA19, GF17, Kro99a]. Guide [AFS81, AFS82, ABB^{+}92, ABB^{+}95, ABB^{+}99, Anoxx, DMBS79, Fox08, GBDM77, GO99, UCLxx, Log99, McC02a, Nas04, NN20, RM99, Rob96, RB92, SBD^{+}76, WvH04, vHW03, vHW06, BGO02, Buc82b, Cal10, Coo91, Dig74, Dig80a, Dig80b, DRP01, Gil04, Ham07, KX86, Lea92, LD13, Men12, Neu84, NN21, PH82, She87, SLC88, SPG92, Uni85e, Uni85c, Uni86, War04, WM01, Yad07]. Guided [GIA^{+}06, Mir03]. Guidelines [STG19]. Guile [Lor95]. Guilty [MOMM11]. Guix [Cou13, CW15a, CW15b, Cou17]. gut [Cam00]. Gutenberg [DDJ99]. Guy [Neu84]. Guys [Pra03]. gVOF [LH22]. GW [MSB09, RAMB18]. GXP [TMM^{+}13].

H [Oms03, BMB^{+}18]. H/FOSS [BMB^{+}18]. H8 [Ano93a]. H8/ [Ano93a]. Hacker [Ray91, Neu84]. hackers [Cio01]. Hacking [Mar01, DPL^{+}91, TG15]. Hacks [ORC06]. HACopula [GHH20]. Hadoop [CZ22]. Haftung [EW01, Geh96]. haftungs [Stö04]. haftungs- [Stö04]. haftungsrechtliche [Sie99]. Haftungsrisiko [Spe01]. Hagenberg [Jef08]. Haifa [FG92]. Halcyon [Boy00]. Half [ST10]. Hall [Bar00c, Fox08]. Halt [CSP^{+}03]. Hamm [Ano98]. Handbook [Bra04, Hig93, Kor11, Ron05a, Waa09]. Handbuch [Bra04]. Handcrafted [SH19]. Handheld [SNF04]. Handling [KT04, LRP21, Tie90]. Handook [RAH^{+}01]. handout [Dig82]. handover [LRD^{+}19]. Happy [EKJ^{+}03]. Hard [Mog03a]. hardback [Aji17]. Hardware [BGG^{+}15, BC20a, GNGS17, HT21, Her20, HW17a, JPOB20, PKP05, BMF^{+}19, Kop05, Pea16, RH21, Tay19, VMKB05]. hardwarenahe [PKP02, PKP05]. Harmful [Nag18]. harmonic [VB19]. Harmony [ATM22]. harnessing [GGT05]. Harper [Neu84]. HART [PMG^{+}09, PKG^{+}10, PPG^{+}11, PBJ^{+}12]. Harvey [Ano00c]. hash [Sch90a, Wil14]. hashing [NMX19]. hats [Ano08a]. Hawaii [MS91, MSNS91]. hazard [Mak03]. HBE [Ano00k]. HC6811C [He95, Ho95]. HCI [Rit88, SS93]. HDG [JRA^{+}18]. Health [CH06b, GGL21, MSLH71, Thi22, EKUR10, XFS^{+}22]. Heart [Coc03]. Heartbleed [Kam14a, Kam14b]. heat [Gal04, MBTB21]. heat-transport [MBTB21]. Heavy [IKW23]. Held [Ano15c, BSK87, MSLH71]. Helmke [Cha13, Teo13]. Help [Sil13, Tra95, JH16, KN93, PDG^{+}87, Pal87, PDG^{+}88]. helpful [Sta96b]. Helps [EKJ^{+}03]. HEPLike [BC20b]. Here [Bar00a, Far91]. Hertin [Oms03, Oms03]. Heterogeneous [DWP^{+}14, Har94, AAB^{+}04, FvH03]. HEVC [GLCMC17]. Hewlett [Ano00f]. Hi [GTMR23, Kan12]. Hi-Lite [Kan12]. Hi-ROS [GTMR23]. Hibernate [WACBL03]. hidradenitis [DSB^{+}16]. Hierarchical [GHH20, JEB^{+}23]. hierarchy [Rog09a, Rog09b]. High [ACM00, Ano94c, Ano16, BPG94, Ede04, KRB^{+}22, LGW^{+}22, PG02, Reh01a, RLVdS21, Ten93, VW92, dlCKK15, BSC^{+}21, CCA^{+}13, CKGW22, DFU20, Eat97, Eat00, Eat02, Eat05, EBH08, GFD^{+}24, HMYH22, HWL^{+}23, HYA20, KT05, KGT22, MVS15, PCAJ^{+}23, Smy97, YKK23, Zag14, DBLF16]. High-Availability [Reh01a]. high-change [KT05]. high-energy [HMYH22]. High-Integrity [PG02]. High-Level [Ede04, CCA^{+}13, Eat97, Eat00, Eat02, Eat05, EBH08, MVS15]. high-level/high-performance [MVS15]. high-order [BSC^{+}21, DFU20, HWL^{+}23, YKK23, Zag14, DBLF16]. High-Performance [KRB^{+}22, RLVdS21, dlCKK15, MVS15, YKK23]. high-precision [KGT22]. high-quality [Smy97]. high-speed [CKGW22, HYA20]. higher [Car89]. highest [KT05]. Highly [MGR16, BCHR12, BB91, KORP95, Kop05]. highly-portable [BB91]. Hill [USE02a]. Hilton [ACM89, ACM93b]. hinder [Sil13]. HIPPI [Ano00k]. Histogram [PBJ^{+}12, PMG^{+}09, PKG^{+}10, PPG^{+}11, PBJ^{+}12]. Historians [Ens04]. Historical [CK08]. History [Ano15b, Boy13, CK07, Geh96, GB21, Pom04, Pri19, Rus14, CFGS05, Mos12, Sch10]. Hit [Fie90a, Ano08a]. HIV [Ano14]. hoc [SH11]. HODG [HWL^{+}23]. Hogs [DDJ99]. Hold [CSP^{+}03, CSD^{+}05, PM00, ST10]. Hold-Up [ST10]. Holes [BNSW15]. HoloGen [CKGW22]. hologram [CKGW22]. holograms [JP09a]. home [PH16]. home-monitoring [PH16]. Homework [GM02]. Homework/ [GM02]. Homo [BH17]. homophily [KF17]. honeypots [IDSM23]. Hong [Uni01]. Honors [DDJ99]. HONPAS [QSX^{+}15]. Hood [Cha98]. Hopes [Bar01]. HORATIO [McL92]. horizons [Bab02]. Horizontal [HZ14]. HOS [DBLF16]. HOS-ocean [DBLF16]. host [AHM^{+}07]. Hosted [GMBv20]. Hot [GB00, BCI^{+}09]. hot-spots [BCI^{+}09]. Hotel [Bao93, USE01b, USE02a]. Hours [P^{+}99]. House [Zha16, Han00]. HP [Ano00f, Ano00f, Dol91]. HP-ST [Dol91]. HPC [BDP^{+}14, CW15a, CW15b, Cre07]. HRSTS [Har77]. HTML [Kro99b]. HTR [DFU20]. http [Den13, EKJ^{+}03]. Hudson [Teo13, Teo13]. Hues [Row02]. Huge [BHP^{+}01]. Human [BSK87, CGK^{+}02, FCTP21, SS93, SM89b, BH17]. Human-Computer [SM89b, BSK87, SS93]. Humanitarian [HE17, Nej12, EMdL^{+}07]. humanity [MTD^{+}09]. humans [Cas19]. Hungarian [Lás05]. Hungary [Cse99]. hunters [ZWH21]. Hurd [Ano01b, Bud10, Bud10, Epp00, WB07]. Hybrid [DO16, JEB^{+}23, LQ17, Sch91a, SSR02, HPT17]. hybrid-open [SSR02]. Hybridized [JRA^{+}18]. hydrodynamic [KD23]. Hydrodynamics [CDR^{+}15, Owe01, ZRZ^{+}21]. hydrological [MLMFN^{+}15]. Hype [Gla99]. Hyperbola [Par03]. hypercubes [PC13a]. hyperparameter [XFS^{+}22]. hypersonic [DFU20, GDJG23]. Hypertext [Con87, Rad92, Pes93, RM92, Wii91a]. HyperTransport [SGNB08]. Hypervisor [Fox08, SJV^{+}05].

i-protocol [DDHS03]. i386 [Ano01b, Ano01c]. IA [Ano00j, TG99]. IA-32 [TG99]. IA-64 [Ano00j]. IaaS [BdSI15]. IBM [Bee86, Rad89, AJ05, Ano86, Ano00g, Ano04a, AV04, AHM^{+}07, CFGS05, Coc03, G^{+}06, Kro99a, Pen03, Sam06, SCH^{+}91b, Uni85a, YLG05]. IBM-PC [Ano86, Uni85a]. IBM-PC/MS-DOS [Uni85a]. IBPM [HMKC12]. ICANN [AT92]. ICANN-92 [AT92]. Icarus [WB02]. ICE [Ano06]. ICMS [FvdHJ10, HY14]. iCompression [Kro00]. Icon [Mit84]. ICs [Bar00c, Bar00a]. IDA [SC02]. IDE [Ano01i, Ano04b, Avi98, Cur02, Sur04]. idea [Ano04a]. Ideas [Wal93, Eub05, BHP^{+}01]. Identification [BNSW15, ASAAM^{+}19]. identify [Ban16, Ban17, SK12]. Identifying [ASWD18, HSX^{+}18, KL07, LZ12, MGYC18, LSM09]. Identity [Par03]. Ideology [Rus14]. IDEs [Sor01]. IDL [Kro00]. IDS [WJM22]. IEC [AM18]. IEEE [Ano04c, BBdD17, Bao93, Rob20]. if [Sta96b]. IFIP [BSK87]. Ignore [CGB^{+}05]. IGSTK [GIA^{+}06]. II [Ano01d, HPT17, MS02, Mog01b, OSM94a, OSM94b, QR92, RAW^{+}16, Rob95c, Ros01c, Sai01, Sai02, TV13]. III [Ano93d, BPG94, JRA^{+}18]. ILS [Ano03a]. im [EW01, GGK99, Sie04, Ste01].

Image [GKL^{+}14, GIA^{+}06, Haf01, Leó98, PBJ^{+}12, RDKT12, BK91, GIM07, Goe07, GRJS01, HFO^{+}12, KORP95, MM04, SS23, YA05]. image-based [SS23]. Image-Guided [GIA^{+}06]. ImageJ [MM04]. Images [PSP^{+}22, BTL^{+}11, Bk94, Bur04b, CKS16, GF17, KB90]. imaginary [CYOS19]. imaginary-time [CYOS19]. Imaging [BJJ14, Kro99b, CG17, SDeaK^{+}09]. Immersive [Ano04b, Coc01a, SSM^{+}07]. immersogeometric [Kam21]. Impact [BB02, Car01, CF09, Gil05, SBDR22, SvGH15, ZW17, DC23]. impacts [YLHW21]. Impatient [AL92]. Implantación [VD01]. Implantation [VD01]. Implement [VOK^{+}22]. implementación [RÓ01]. Implementation [AML^{+}10, BdP13, BGG^{+}15, HAC^{+}23, JKS02, JD19, Joh18, KGT22, KSV16, KG01, Mam01, Men10, Mio90, Nov04, Pud04, RSKF96, SP12, TBPS15, Wut12, BGM99, BCI^{+}09, CFCA13a, CFCA13b, Fär05, GS02, HK95, He95, Ho95, HM10, Kir12, LLEL^{+}23, Lla06, Mat03, MSM^{+}03, MSK05, Neh07, PM21, RÓ01, RSZ96, TMM^{+}13, Tie90, Váz16, VB19, WPAV14, XOTI22, ZC01, Ang01]. implementations [ABF^{+}14, DO16]. Implemented [Est06]. Implementierung [Lei93b]. Implementing [AS97, BG95, But95, CH06a, DFLS05, KTP95, Sim12, VSdCCR23, Wel95]. implements [MZE13]. Implications [GM05, McG01, Thi22, dCdCM14, BA15, Gom99, Tro04]. importance [Añe11, BH11]. Important [Boy07, CK06b, Mud97]. imposition [MVF20]. Impossible [CSD^{+}05]. Impressions [CCSW10, BCG^{+}14]. imprisonment [Sta96b]. Improve [WLC01]. Improved [OCH90a, OCH90b, LH14, Qui00]. Improvement [Bes03, PBJ^{+}12]. Improvements [BOM97, PMG^{+}09, WKA^{+}08]. improves [WMLM22]. Improving [Ave06, BBM^{+}21, DD08, HBC^{+}05, LLWM23, SRGCPB^{+}09, BR95, PYM^{+}06, ZDM10]. impurity [HWM^{+}15, Hua17]. IMRT [KMF^{+}07]. IMSI [Ano04b]. In-depth [SJW22]. In-House [Zha16]. included [Ano97a, Ano00a, Ano00b]. including [GVOM09, HL02, PH16]. Incompatibility [XGF^{+}23]. Incomplete [MRGP20]. Inconsistencies [WBRH23]. inconsistency [WMK^{+}17]. incorporating [YLXZ16]. Increased [HJ07]. Increasing [Tay00]. incremental [Jør01]. incrementalism [GGT05]. Incubators [DGC^{+}07]. indecent [Sta96b]. Indent [Arc94]. Independence [NRG^{+}99]. Independent [MV05, Lin02b, BSP11]. Index [Ano95a]. indexing [PPR19]. India [Ano03e, BAR16b, GAS^{+}01]. Indian [Bab02, Nor02]. indicators [XFS^{+}22]. Indirectly [Yu06]. induced [AH19]. Industrial [BCB^{+}17, PMBM^{+}15, vdLLM09, AM18, HZS^{+}16, MRS07, PWA^{+}19, ZVvDD11]. Industrials [Ros05]. Industry [Boy07, Car01, Ebe08, Ebe09, Edw98, GGB17, HBR19, Kam11, Rob20, CLM^{+}08, EMD03, KS03, LLEL^{+}23, ZFD21]. Inel [SuS01]. Inel- [SuS01]. inelastic [WFDK19]. INET [Lei93a]. Infancy [LRBM23]. Infection [Ano14, Ros01d]. Inference [MRGP20, HFO^{+}12, KSS^{+}23, PNK^{+}23]. InfiniBand [HMKC12]. influence [KF17, Sin10a, YLHW21]. Info [Pes93]. informatics [KSD^{+}12]. Information [AMS03, Ano96e, CF07a, CSP^{+}03, CKB^{+}05, CHA06, FK04, Goo14, Has05, Pel89, San03, SCSC04, ACM05, Wat01, YDZ19, Aji17, BBNP93, Eds16, Fri16, HK95, KMG^{+}93, Kit94, KRR23, Sai13, Sea04, Ano95a]. infrared [LA10, SC16]. Infrastructure [Coc01b, FK04, Nov03, WKB14, BOL14, SGM^{+}08]. infrastructures [PWA^{+}19]. ing [Jon01]. inherited [Big13]. inhibitors [AMC16]. Initial [MCS12, JJ91, YLHW21]. Initiates [HW17b]. Initiative [Coc01a, WG00, ASAAM^{+}19, FMFZ19, MTBS09]. initiatives [Man92]. initio [RAMB18, WPAV14, CSV^{+}07]. Inner [LR11, SF15]. Innovation [Ebe07, GNGS17, PRRL12, BAR16b, Far23, FvH03, MSR09, MLWR18, WG06, vKSL03, LMWM18]. Innovations [Boy08, Gil05, Rie19]. Innovative [Ano96c]. innovativeness [Lam09]. Input [GF11]. ins [DF00]. insertion [LFA92]. Insider [EJS^{+}01]. insight [KTF15]. insightful [SDL^{+}16]. Insights [BCB^{+}17, MSR09]. inspection [ZDM10]. inspiration [Dvo04]. Inspired [MPG^{+}16, MSR09]. InstallAnywhere [Ano04b]. Installation [Ano90c, G^{+}00, G^{+}02, GA04b, GO99, Str94, Jan01, Mag01a, Ron05b, SLC88, Mag00]. installed [XOTI22]. Installing [EXA^{+}05, Gre11a, Gre11b]. Instances [LGW18]. Institute [ABC^{+}14]. Institutes [MSLH71]. Institutional [BMB^{+}18, JDB09, MSZ02]. Institutions [Bro01, ST10]. Instruction [SCH^{+}91b, Cha92, Deo90, Mak03, UZ97]. Instrumenting [MZG14]. Instrumentino [KSH14]. instruments [KSH14]. integer [MRH23]. integers [GM94]. integral [MN21]. Integrands [Joh18]. integrate [NZPWR22]. Integrated [BGG^{+}94, Li18, PG02, PGC21, HM89, MSZ^{+}01, Ano02b]. Integrates [Mor08]. Integrating [APK14a, APK14b, GP05, Hin87]. Integration [CPJ^{+}98, GCE^{+}21, Kro00, AKHG16, BSW95, CH11, FRBRF19, LMZT22, PC13a, THG20, WMLM22, vGPB10]. integration-oriented [vGPB10]. Integrity [PG02]. Intel [Ano01c, Ano00f, Ano00h, BHP^{+}01, BKP05, YSMA^{+}17]. Intellectual [Lin08, Mar05, Vai01, WBGM02]. intellectuelle [Cor00]. Intelligence [Ano87, Ano88b, BPG94, IEE94b, SSH22, TG15, AK95, BA15, GLMS18, KTH^{+}22, KRR23, QC18, VSN22, SD16]. Intelligent [Deo90]. Intels [Ano00f]. intensities [BDAW15]. intensity [Dan11]. intensive [BGL^{+}22]. inter [Hub04b, ESM19]. Inter-organizational [ESM19]. inter-procedural [Hub04b]. Interact [GBICMR13, BSK87]. Interacting [WWSG21]. Interaction [BSK87, SM89b, ZAC^{+}23, HPT17, Kam21, SS93]. Interactions [PMM17, PMM18, BFI^{+}21, CKB11, DRM21, YKSH20]. Interactive [BSS84, Coo95b, DKMT11, RAH^{+}01, San78a, WKC^{+}90, Coo95a, Eat97, Eat00, Eat02, Eat05, EBH08, FHH11, KK94, Rac06, TL17]. InterBase [Ano98]. Interchange [SC02]. interconnection [AVA^{+}16]. Interest [GW09]. INterface [LH03, ACM88, AG95, Ano96b, Lor95, SMNF88, dlCKK15, BJWZ08, Li91, MGPB20, Pel89, Sch90b]. Interfaces [KMG^{+}93, BBNP93, MSK05]. Interfacing [Pil09]. Intermediate [Gup03, CYOS19]. intermediate-representation [CYOS19]. Internals [BBD^{+}96b, Tay99]. International [AT92, Ano91, Ano00k, Bao93, BPG94, Bun94, Cse99, FMA02, Fur90, HY14, IEE89, IEE90, IEE92b, IEE93, IEE94c, IEE95a, IEE95b, Jef08, Kap92, Lei93a, Lev95b, Lev95a, MS91, MSNS91, Mio90, MG94, QR92, SS93, Uni01, VW92, Vor92, ACM93a, EKR91, FvdHJ10, PT91, Ano01j, Ano14]. Internationalizing [dM99, Rei93]. Internet [Ano96c, CK06b, DF00, Bar00a, BGG^{+}00, Boy13, CK06b, CK06c, CRW^{+}04, EMD03, HNH03, Ian02, Li18, MD18, MSW09, PM00, Shi12, SH11, TF21, Zic01]. Internet-based [HNH03]. Internetkommunikation [CK06b, CK06c]. Internetprogramme [CK06b]. Internetprogrammen [CK06c]. InterNetwork [Ano01j]. Interoperability [BES^{+}01, MMD12, BGL^{+}20]. interpolation [CDSV10, CDSV11, PC13a]. Interpretation [AFS81, AFS82, FY18, Tra95]. interpreter [HC07, Rap94]. Interval [AS97, SZAB97, SZAB98, SZAB99, GB06, Hei16, Neh04]. Interval-Enhanced [SZAB99]. Interview [Li94, Ray99a]. Intonational [Fyk97]. Intr [AFS81]. intramolecular [VB19]. Intranet [Ano96c]. Intranet/Web [Ano96c]. Intrinsics [AS97]. Intro [AFS82]. Introduce [HOL^{+}07, MAMC05]. Introduces [Ano01j]. Introducing [BMS^{+}22, CJ19, HMP^{+}15, KNS18, Kle21]. Introduction [BV87, Cha97, Chaxx, Cha01c, Cha04, CK06a, CK06b, CK06c, Ebe09, Gor96, Kri03, Mar03, SF05, SC88a, SS04, Xia08, Cic78, Dig75a, Fin22b, Gou04, NS05, Wan21]. Introductory [MMP^{+}22, PKP05, GM84, MdL09]. Intrusion [Cha01a, Wen00, AG22, HYA20]. Intuitive [WLD^{+}17]. Invalidity [Mog03b]. invariant [GM94]. Inventions [Est06]. Inventors [Bar00c]. inversa [DSB^{+}16]. inversion [MN21]. Investigating [AMWH19, RB05a, Ano01g, IKW23]. Investigation [DC23, IAS16, PLO^{+}23, Sha10, YWA07, Ano02a, FG92, HM10, Koc07, KKA^{+}19, MLA^{+}19, Tai13, THG23, WHJ15]. investigations [BA15]. investment [Pea16, WG06]. InvFD [MN21]. InvisibleWeb [Ano96c]. involvement [CFGS05, CFMRL11]. IOb [RLVdS21]. IOb-Cache [RLVdS21]. IONA [Kro99a]. IoT [CTP^{+}22, CYL^{+}23, PWA^{+}19, ZZZ22]. IP [Str94, Ano01h, AML^{+}10, BGG^{+}00, KG01, PC13b]. IP-PBX [AML^{+}10]. IP-PBX/VoIP [AML^{+}10]. IPv6 [Bla06]. IPv6-enabled [Bla06]. i}QIST [Hua17]. irbasis [CYOS19]. Iron [BW00]. irregular [ZSW14]. ISBN [Aji17, Ano11, Ano15a, Cas02, Fox08, Kuc06, SD16]. isogeometric [Váz16, ZLF^{+}22]. ISP [Ave06]. ISPH [MRN20]. ISSAC [Jef08, Lev95a]. ISSAC’95 [Lev95b]. Issue [CJ19, DBBA10, MOT^{+}18, ZWH21, vKvH03]. Issues [Kenxx, Kim01b, KP93, LG02, Wii91a, ZXB^{+}23, Ano01g, Ano02a, HYA20, Man92, MV05, TB05]. Italian [Cor05, Fav23, Lam09, Mol01, Zic01]. Itanium [Ano00f, Ano06, Ano00f, Ano00h, Ave06, Liu06]. Itanium-based [Ano06]. iterative [CHE^{+}10, MRN20]. ITHACA [PK10]. IUC18 [Uni01]. IV [BSW95]. Ivan [Ano00l]. IVR [Ano00k].

J [CFCA13a, GKL^{+}14, GNR^{+}09, Pri19]. J2ME [Vir05]. J2SE [Ano00j].

JafSoft [Kro00]. James [Ano04c, Kim01a]. January [Hea09]. Japan [FvdHJ10, IEE94a, SS02]. Japanese [HK95]. Jason [Ano00l]. Java [Ano96c, Ano00k, Kuc06, Kuk98, Sur04, ABC18, Ano01i, Ano01j, Ano02b, Ano03a, Ano03d, Ano04a, Ano04d, Avi98, Bar01, BHP^{+}01, CJ17, CJ19, CDG97, DDJ99, Egy01, EXA^{+}05, Eub05, GD12, GMBv20, GM02, HL02, Kro99b, Liu08, Mam01, MB98, MM04, Nas04, NRG^{+}99, Och09, Pet05, Pra03, Shi03, Sta04b, SHK^{+}03, Sur04, THG20, Tro97, Vir05, WACBL03, YLL^{+}07]. Java-Based [Ano01j, CJ17, CJ19]. JavaScript [Ano00c, CASA22]. Javier [Ano00l]. jedermann [Str94]. Jeremy [Ano11]. Jesus [Bar16a]. JETSPIN [LPC^{+}15]. Jikes [AAB^{+}05a]. Jim [Coc01a]. John [Ano00l, Ano15a]. joining [vKSL03]. Joins [Bar00c]. Joint [ZPH^{+}15]. Jones [Ano00l]. Jorgenson [Ano00l]. JOSS [KNS18]. journal [Ano01a, FMFZ19, Rud10, KNS18]. Journey [Fav23, RLTD23, ES23]. Journeyman [DuB02]. Journyx [Ano98]. JPEG [CKS16]. JTruss [RP08]. Judgement [Pow00]. Julia [HMX21a, HMX21b]. Julia-based [HMX21a, HMX21b]. July [ACM93b, Bun94, BBdD17, IEE92a, IEE92b, Jef08, Lev95b, Lev95a, Ass95]. Junction [Kro00]. June [ACM92, ACM93b, ACM94, Bun94, FMA02, HDR04, IEE90, IEE93, Kap92, LMOS93, QR92, USE90, USE98a, USE99, USE01b, USE02b, USE02c, Abr81]. JUnit [HL02, WACBL03, ZK05]. JupyterLab [HHG^{+}21]. Juristische [HK03]. juSFEM [HMX21a, HMX21b]. Just [ATM22, CMJ^{+}04, YLL^{+}07, HZ14]. Just-in-Time [YLL^{+}07, HZ14]. Justifying [CGK^{+}02]. Jx [YKSH20].

K42 [AAB^{+}05b]. Kanji [MYU89]. Kaspersky [Ano00i]. KBEmacs [Wat85b, Wat85a, Wat87]. KDE [Ano98, SuS01, EJS^{+}01, GWT^{+}01, Thi99]. keep [Ude97]. KEK [AY93]. Kerberos [Coc01a]. Kernel [BBD^{+}96b, Mar01, Max01, Ros02a, SuS01, Wel95, Wes00, AAB^{+}05b, CIC13, HNH03]. Kernels [dlPRGB99, YSC^{+}06]. Key [Bar00a, Men10, O’Rxx]. Keynote [Tho90a, Tho90b]. Keys [Noj01]. KGI [Bud10, Bud10]. Kiasan [BCHR12]. Kickstarting [GNGS17]. Killer [DDJ99]. kind [Avi98, SG99]. Kinder [CK06g, CK06h]. kinematic [IHBS14]. Kinetic [HFO^{+}12]. Kingdom [AT92]. Kit [Ano96b, Ano96c, Ano97c, Ano00k, Ano01j, CKH91, Kro00, Rui02, SHK^{+}03, Weh03, Ahm08b]. KNOPPIX [Ano01d]. Know [Pit16, CWHW12, NN20, Woe94a]. knowing [TG99]. Knowledge [CH10, DFLS05, WCA^{+}14, ALVV17, AD04, HMP^{+}15, KK17, RCO20, Sin08, SSA08]. Kobe [FvdHJ10]. Kode [NN16b, NN21]. kommentiert [J^{+}05]. kommerziellen [Sie99]. Kommunikation [Lin02a]. kompromisslos [Bra04]. Kong [Uni01]. Konkurrenz [HK03]. Konqueror [Hau01]. Konqueror/Embedded [Hau01]. Konvertieren [Gün02]. Korea [HY14]. Krause [Ano00c]. Kremlin [GJLT11]. kriminalpolitische [Rau04]. kriminologische [Rau04]. Kroll [Ano00l]. Kronrod [Joh18]. Krylov [HKY^{+}21]. Kubernetes [DKK22, RSBP23]. Kubuntu [CK06c, CK06d, CK06f, CK06h, CK06c, CK06d]. Kudos [DDJ98a, DDJ98b].

Kumar [TG15]. kurz [Cam00]. KYZO [Ano00k].

L [Neu84]. Lab [Ano00i, Coh03]. labor [GLT08]. Laboratory [Ano94a, PH82, CFW01, GBG^{+}16, Rud10, Jen01]. Laborjournal [Rud10]. LaGrande [And03]. Lahey [Ano01a]. Lahey/Fujitsu [Ano01a]. LALR [Mey18]. LAMP [LW03]. LAN [Ano02b, RÓ01]. Land [Ber96]. Landeshauptstadt [SG05]. Language [AKW88, Ano01j, DLT^{+}23, NRG^{+}99, Nor02, Tro96d, Wal93, BR95, Eat97, Eat00, Eat02, Eat05, EBH08, Hin87, MR94, May17, NRRS20, Nob08, SC88a, SLC88, Tho90a, Tho90b, Win95]. Languages [Ano94c, CPJ^{+}98, Kim01a, Cra90, Mud97, Val91]. Lanham [Aji17]. LAPACK [ABB^{+}92, ABB^{+}95, ABB^{+}99]. Laplace [Wut12]. Laplacian [CZS^{+}21]. Laptop [Ano97c]. Laptops [Ano00l]. Large [Ano90c, CYL^{+}23, KGM^{+}16, KCAS23, KT04, KS11, PSR16, ZXB^{+}23, ASAAM^{+}19, ACW04, BOL14, BLG^{+}17, Big13, CJ19, DRM21, Koc07, KT05, KL07, LSJ^{+}06, MNS19, QSX^{+}15, RB05a, RBM^{+}23, RAMB18, SMS16, TTL06, VBG^{+}10, VGP^{+}19, WFF18, WMK^{+}17, XTY^{+}22, vGPB10]. Large-Scale [ZXB^{+}23, CYL^{+}23, BOL14, BLG^{+}17, CJ19, Koc07, KT05, KL07, LSJ^{+}06, MNS19, RB05a, RAMB18, SMS16, VGP^{+}19, WFF18, XTY^{+}22]. Larry [DDJ99]. laser [AH19]. laser-induced [AH19]. last [Cra89, Cre07, Lea92, LMOS93, Sta92b]. Latency [CKB^{+}05]. Later [WB02]. lateral [SKB23]. LaTeX [Str94]. LaTTe [YLL^{+}07]. lattice [KKA^{+}21, ASC^{+}21]. Launched [Bar00b]. Law [CPJ^{+}98, Doh01, Gil05, Hüp01, Ros01a, Ros02b, WP04, Mar05, MG05, Oms03, Bur04a, MSSvK08, NGJ03]. laws [Ano01h, Kam11, Les99]. layer [VOK^{+}22]. Layton [SD16]. Lazenby [Ano00d]. LBfoam [ASC^{+}21]. LBsoft [BMT^{+}20]. LCA [Mag04]. LCP [Mag04]. LCP/LCA [Mag04]. Lead [Fre23]. leadership [Far23, Fie99]. Leading [BCHR12, Far23, Kan12]. Leading-edge [BCHR12, Kan12]. Leads [Bar00a]. learn [NN20]. Learned [BBM^{+}21, NK04, XWZ^{+}23, Kle21]. Learning [Ano00e, BY14, BKP05, CR91, CR92a, CRR96, CEL^{+}05, Cio01, FKM^{+}11, GF11, Mac02, McC99a, McC99b, McC02a, McC03, McC04, MSZ02, Nej12, PMD13, Raf23, ZRNA20, ASAAM^{+}19, CV13, Cha11, CP04, FMT^{+}08, HBZ09, MLZ^{+}23, PNK^{+}23, SKB23, XTY^{+}22, ZDM10, Ano00a, Ano00b]. Learning-Based [BKP05]. Least [eLAA^{+}23]. led [Pya06]. Left [EKJ^{+}03]. Legacy [BHP^{+}01, Joh94b, Kle21, RVLS14]. Legal [Col09a, CGB^{+}05, Gil05, Jak03, Kenxx, McG01, Pom04, SCB04, Sto09, WP04, Feixx, O’S03, Sie99, Sie04, HK03, Man03, SG05, Spi03]. Legality [Pow00]. Legalize [CSP^{+}03, Sta03c]. Legally [Sam06]. legislative [Mol01]. LegUp [CCA^{+}13]. Leif [SC88a, SLC88]. LENA [LLEL^{+}23]. length [Che95]. lengths [GF17]. LEO [CSD^{+}05, CGB^{+}05]. Lern [CK06g, CK06h]. Lern- [CK06g, CK06h]. Lernprogramme [CK06e, CK06f]. Lernprogrammen [CK06e, CK06f]. Less [Coh02]. L’essentiel [Rod00]. Lessons [BBM^{+}21, Bor88, Ens04, Fit11, NK04, O’R99, RCP^{+}12, XWZ^{+}23, Kle21, AD04]. Let [DDJ98a, DDJ98b, STG19, NN20]. Letter [Ano99d, EKJ^{+}03, The04]. Letters [BES^{+}01, BHP^{+}01, Bur04a, CPJ^{+}98, CPG^{+}04, CDsJ^{+}00, DM97, DuB02, EJS^{+}01, EKJ^{+}03, Gal60, GWT^{+}01, KHA^{+}03, NRG^{+}99, RAH^{+}01, SSC^{+}00, ACM05, CAC09, Sta01b]. Level [Ano94c, BOM97, BGG^{+}15, Ede04, PMBM^{+}15, Sta89a, SPS^{+}00, SPS^{+}02, YXS^{+}19, But94, CCA^{+}13, Eat97, Eat00, Eat02, Eat05, EBH08, EKUR10, HC07, MBR21, Mag01a, Mag01b, Mag01c, MS02, Sai01, Sai02, Sin10a, Sta88a, Sta89b, SPG92, SP93, SP95, Sta96a, Sta98a, SCH^{+}91b, Mag00]. level/high [MVS15]. Levels [FCTP21, Luc99a, Sha10]. Leveraging [PCAJ^{+}23]. Lexical [Nic93, Pax88]. leyes [Les01]. LF95 [Ano01a]. LGPL [Jak03, Jak03]. Liabilities [EW01]. liability [Geh96, Kam11, Spe01]. LIB [Cha91]. libC [Gar00]. libdbh [Wil14]. libero [Cor05, Mol01]. libg [Lea88]. Libkrylov [RBM^{+}23]. libkww [Wut12]. Libra [Ano00k, Aki16]. Libranet [Ano00k]. Libraries [Ano00l, BHP^{+}01, Gro01, Has05, LS04, Eub05, Jon01, LRP21, McL92, Pya06, Sch09]. Library [Ano96c, Ano01j, ALA20, BKP05, Coo95b, FL16, GDT^{+}02, GDT^{+}05, GSR^{+}04, KY16, KSD^{+}12, Lea88, LSM^{+}99, LSM^{+}00, MD22, PQM11, SB08, SKSM19, Sta92c, Sta97c, SHK^{+}03, Woo01, Aji17, Aki16, BFI^{+}21, Cha91, Coo95a, CKB11, DSM^{+}19, Fär05, FHL^{+}07, Fow00, Fri16, G^{+}01, Gal09, HKY^{+}21, Lea92, Lea93, LS04, MRH23, MCQF21, MGR16, Neh07, NZPWR22, NGCI^{+}12, Pya06, RBM^{+}23, Sai13, SPAW17, Sta88d, Yad07, Yes12, YMCF23, ZSW14, ZRZ^{+}21, AAB^{+}04, Bad07, Bee01, GB94, LMOS93, LSMO96, LSM^{+}01, Loo15, Neh04, PQM11, WACBL03]. library-led [Pya06]. Libre [BKR^{+}20, BSFR22, CWHW12, CF98, O’S03, YA11]. LibreOffice [GL14]. libres [Cor00, Séd02]. libstdc [Car04]. Libtool [VETT00, Cal10]. licence [Mor11, O’S03]. License [Cou20, Eng10, GD12, Hüp01, Maj03, Omb20, PMD13, PZ20, Ros02c, Sch19, SH11, Sta12, WBRH23, XGF^{+}23, CF98, CF09, Gom99, KTF15, KKT17, WMK^{+}17, Guy00, Hüp01, Neg15, Sal20, Sie99, Väl04]. Licenses [Bal19, Hol23, Jak03, KC21, LRP11, MG05, Opexx, Rav00, SSH22, O’S04]. Licensing [Hef97, Kenxx, Lee99, Mal02, MOMM11, St.04, Sto09, VH04, AMWH19, Gil04, Mar05, PKGA22]. LiDetector [XGF^{+}23]. Life [Cas19, EKJ^{+}03, mH00, Str02, ZXB^{+}23]. Lifecycle [Kro99b]. Lifecycles [BE06]. Lifetime [Coc01a, Ano21]. Lifting [PMBM^{+}15]. Light [MD22, SAC^{+}15, Mol01, IKW23]. Lightweight [dlCKK15]. Like [Raf23, Ano96d, Dew07, Ude97]. Likelihood [WM05, BC20b]. Liley [Gv14]. Lilypond [Sha05]. limitations [HKP02]. limiting [RAW^{+}16]. Line [Ano00l, BdP13, Tan11a, Tan11b, Tra95, GM02, KN93, MAF22]. Lineage [CRW^{+}04]. Linear [Bar01, KY16, WR71, FBY^{+}17, QSX^{+}15]. lines [vGPB10]. Link [Ano98, Ano00i, DVC^{+}07]. Link-time [DVC^{+}07]. linking [GRJS01, May17]. LinkScan [Ano00i]. LINPACK [DMBS79]. Linus [CPJ^{+}98, Flo94, Li94]. LINUX [Bra04, GGK99, Str94, Ano95a, Ano96b, Ano96c, Ano97c, Ano98, Ano00e, Ano00d, Ano00f, Ano00i, Ano00j, Ano00k, Ano00l, Ano01a, Ano01b, Ano01d, Ano01c, Lin02b, Ano05b, Ano06, Bud10, CK06a, CK06b, CK06d, Cor05, DF00, G^{+}00, G^{+}02, GA04b, Gan04, G^{+}06, Gra99, Gün02, HHV05, Jor04, Kuk98, Lin02a, Mag04, RÓ01, Ron05a, SuS01, USE00a, USE01a, Lás05, RÓ01, Str94, ACW04, AJ05, Ana99, Ano99b, Ano00c, Ano00f, Ano00h, Ano01c, Ano01j, Ano03d, AAB^{+}05b, AS03, BGM99, BBD^{+}96b, Bel00, BS05, BGO02, BK02, Bud10, CIC13, CM06, Cha11, CK06a, CK06b, Cor00, Cor05, DRP01, III01, Den99, Don04, Dum05, EKJ^{+}03, FT09, GA04b, Gar00, G^{+}06, Gla99, GO99, Gün02, Hau01, HNH03, Hom00, Hun01, Jae03, Jan01]. Linux [JJ00, Jon05a, Ken02, Kop05, Kre00, Kro99b, Kro99c, Kro99d, Laz99, LW03, Lin02a, Lin00, Mac02, Mac99, Mag00, Mag01a, Mag01b, Mag01c, MS02, Mag04, MSSvK08, Man00, dM99, Max01, McC99a, McC99b, McC02a, McC03, McC04, McC05, MZG14, Moo01b, Moo01a, Mur94, NR03, NN16a, NN16b, Noj01, Nor02, PR96, Ray99b, Ray99c, Ray99a, Ray01b, RAH^{+}01, Ron01a, Ron01b, Ron05b, Ron05a, RT05, Sai01, Sai02, San98, SS05b, Sea02, Sea04, SK04, Sor06, Sor01, SNF04, SGD00, TTB09, Tor99, Tro96b, TB05, Ude97, VD01, VMKB05, WY94, Wel95, W^{+}95, Wen00, WG00, YL08, Yeo05, Ygg93, Ygg94, YSC^{+}06, Zac01, Zic01, Ano00a, Ano00b, Men12, Wil13]. Linux-Anwenderhandbuch [Ano01c, Ano01c, Ron01a, Ron01b]. LINUX-Applikationen [GGK99]. Linux-based [CM06, WG00]. Linux-Betriebssystem [SuS01]. LINUX-Büropakete [GGK99]. Linux-compatible [AAB^{+}05b]. Linux-KGI-Treibers [Bud10]. Linux-Systeme [Gün02]. Linux/FreeBSD [Ano00i]. Linux/GNU [Jan01]. Linux/GNU/X [Ygg93, Ygg94]. Linux/Unix [Sor01, Ano03d]. Linux4.TV [Hae02]. LinuxCAD [Kuk98]. Linuxzeitung [Lin02b]. Lions’ [Lio96]. Liquid [Ano96c]. LISA [USE94, USE98b]. Lisp [Bir93, Lew88, LLG90, LLG93, ACM92, Gre80, Cha97, Chaxx, Cha01c, Cha04, Gad88, Hen92, Iwa02, LLG94, LLSt99, LLSt00, MS20, NS01, Rap94, Sta78b, San78a, San78b]. Lisp-based [Iwa02]. List [TTB09, VWM98, Uni77]. listings [Win95]. Lite [Kan12]. literacy [YAS91]. literate [KC92]. Literature [TWS^{+}22, AJLM18, YT22]. Little [Ano00k, Val91, Les03]. Littlefield [Aji17]. Live [FK90, Gre11b, FG16, Gre11a, Pre16a, Pre05, Pre08, Pre16b]. livelock [DDHS03]. Lives [CGB^{+}05]. living [MSR09]. Lizenz [Sie99]. Lizenz- [Sie99]. LJ [EKJ^{+}03, RAH^{+}01]. LLC [Ano00k, Ano98]. Llunatic [GMPS14]. LLVM [ZRGJ21]. LMAKE [Lei93b]. Load [Ano01j, EJS^{+}01, Kro99b, TPK^{+}21b, ZSW14]. Loadable [Wel95]. Loader [BHP^{+}01]. loads [KD23]. Locale [Noj01]. Locales [Noj01]. locality [Fär05]. Localization [YYL^{+}15]. Locking [Sav23]. logbook [HBB^{+}12]. Logging [CJ19, LZWH22, CJ17, FG16]. Logic [Vor92]. Logical [CRW^{+}04, LO89]. logiciel [CF98]. logiciels [Cor00, Séd02]. logistics [dA15]. Logs [CSY^{+}04, SKB23]. London [BBdD17]. Long [Far06, Mas05]. Long-Distance [Far06]. long-timescale [Mas05]. longer [Ant16]. Longhorn [And03]. Longitudinal [JEB^{+}23, Mas05]. Look [CSD^{+}05, Fit04, Mog03a, Whe03, Gla03b, Gla04, Oma89, Sal88]. Loop [Cha92, Dvo03, Ede04]. Loosa [RAH^{+}01]. Lost [CGB^{+}05]. Lösungen [Bra04]. Lotka [Bur04a, NGJ03]. LOTOS [Lal91, Sch90b]. Lotus [Ano01i]. Louis [ACM97]. Loukides [Ano97a]. Lout [Ano10]. Low [BOM97, BGG^{+}15, CCA^{+}19, CTP^{+}22, RDKT12, Smy97]. Low-Cost [CTP^{+}22, RDKT12, Smy97]. Low-Level [BOM97, BGG^{+}15]. Low-Voltage [CCA^{+}19]. Loyal [Gla99, Gla00]. LPAR [Vor92]. Ltd [Ano00i, Ano00k]. Ltd. [Ano00j, Ano00k]. LTE [ACB^{+}16, CTP^{+}22]. LTE/4G [CTP^{+}22]. Lunch [Gre18, Mic04, Fie90b]. Lyon [DMP^{+}02].

M [Ano00l, Fio03, Gay02, Neu84]. MAC [SJV^{+}05, Sta06]. MAC-Based [SJV^{+}05]. Mach [Ano93d, Kup93, Mor96]. Mach-Based [Mor96]. MACHINE [BY91, BY14, BSW95, FKM^{+}11, Raf23, BY92, CFW17, CK06a, CK06c, FMT^{+}08, HBZ09, SKB23, AAB^{+}05a, CDG97, JCNS^{+}22, MB98, Shi03]. Machines [LLWM23]. Macintosh [Ano88a]. Macmillan [Ano00k]. macro [Sin10a]. macro-level [Sin10a]. Macromedia [Ano02b]. Macs [STS92]. MACSYMA [Cla90, Mos12]. maddog [EKJ^{+}03]. Made [CX23, Kre00]. Madison [FMA02]. Madness [CPG^{+}04]. Magma [Kop20, Ste08]. Magnetic [Bar00a, BM22, CFW17, YKSH20]. Mailing [TTB09, VWM98]. Mailman [VWM98]. mainframe [YM93]. Mainstream [RSAT19]. maintain [Big13]. Maintainability [YSC^{+}06, FRBRF19, SSAO04]. maintainer [BJM^{+}22]. Maintaining [BGL^{+}20, Luc99a]. Maintenance [MG94, Wei03, BSP11, Car89, DFCPSF15, KH05, KFYI13, Yu06]. maintenance-free [DFCPSF15]. Major [FL16]. Make [Ano15a, Bar00a, BKHT21, CGK^{+}02, DDJ99, EKJ^{+}03, Mec05, OT91, Sha05, SM00a, SM00b, SM02, SM89a, SMS04, Lei93b, Bak20, TMM^{+}13]. Make-Programmes [Lei93b]. Makes [Fri97]. Making [Ber22, Bro96, FFvdH01, GGL21, O’D07, O’S02, SSS22]. Malware [ASWD18]. Manage [Bak20, DM15a, DM15b]. managed [Mah03]. Management [AtHR11, Ano96c, Ano02b, BAP00, BVLF14, Bro01, Cou13, FMA02, Kro99a, Kro00, MSC19, MSW09, SPDQ22, Sim00, Sta02c, VOK^{+}22, WKB14, Zha16, ACW04, ABNÅ05, AAA^{+}12, AD04, BJWZ08, BCvE^{+}05, BCI^{+}09, CP04, Col05, CdSV07, Cus04, Dig82, EKUR10, HMP^{+}15, KMG^{+}07, Mon03, MSR09, VD01, ZFY^{+}19]. Manager [Ano00i, STS92, VWM98]. Managers [ATM22]. Managing [BSFR22, BG12, BS98, CCSW10, GGH10, Gus20, Har99, Har20, Mec05, NN20, OT91, Plo97, San01, NN21, WG05]. Manifesto [Sta85]. Manifests [RSBP23]. Manipulation [Fur90, HOST05, Leó98, Abr81, GIM07]. ManPy [DPH16]. Manual [AFS81, Ano96a, DS00, DS02, DF00, GAW87a, GAW87b, GDT^{+}02, GDT^{+}05, The03, LO89, LO92, LSM^{+}99, LSM^{+}00, LSM^{+}01, Mey18, MC91, Sta80b, Sta89a, Sta92c, Sta97a, Sta97c, Sta02b, Com84, CCA84, Clo89, Clo91, Dig75b, Dat85, EBH08, EaoGOBHW14, GIM07, G^{+}01, Gal09, Gil93, Gos84, Ham88, Har77, KK99, Lew88, LMOS93, LSMO96, Loo15, PS^{+}09, SLC88, Sta80a, Sta81d, Sta81c, Sta81b, Sta86a, Sta86b, Sta86c, Sta87, Sta88b, Sta88c, Sta88a, Sta89b, Sta92a, Sta93a, Sta93b, Sta94, Sta95, Sta03b, SW15, TSM88, Uni85d, Uni85f]. manuel [Rod00]. manufacturing [DPH16, VMKB05]. Many [GA04a, JWC18, MMY^{+}19]. many-body [JWC18]. many-variable [MMY^{+}19]. Manycore [BMF^{+}16]. map [MM04]. Maple [Ste08]. Mapped [STS92]. Mapping [YLL^{+}07, CV13]. MapQuant [LSJ^{+}06]. MapReduce [CZ22]. Maps [PFL^{+}12, GJMPAM^{+}14]. March [EKR91, IEE94b, SC00]. Margins [Tay00]. Marine [FVD^{+}12]. Marjorie [Ano00e]. Mark [Ano99a, Ano00l, III01, Men12, Neu84]. Markenrecht [Fal03]. Market [All02a, All02b, BMZ14, CWB^{+}04, GB00, GAS^{+}01, Rav00]. Marketplace [Pau04, Wal93]. Markov [Mar22]. Markup [Kim01a]. Marriott [USE01b]. MaRTE [RTH15]. Maryland [Fur90, MSLH71]. masked [YZC22]. Mass [Rav00, LGA20, NGCI^{+}12]. Mass-Market [Rav00]. mass-transfer [NGCI^{+}12]. Massachusetts [IEE92a, IEE92b, USE01b]. massively [ORS^{+}14]. MasterConsole [Kuk98]. Mastering [WM01, HL02]. Matching [LRP11, CLS95, Joh94a, OK94, ZLF^{+}22]. Material [BCB07, CBB06, HMP^{+}15, MVF20]. materials [WZS^{+}18]. Math [DDJ98a, DDJ98b]. Mathematica [Ano95a, Ste08]. Mathematical [Hig93, HY14, JS07, FvdHJ10, Joy09a, KJ03, FvdHJ10]. Mathematics [Den13, BAE14]. MATLAB [Bar16a, FRAK15, Ano96d, AMR18, GHH20, JRA^{+}18, Joh18, PC13a, PSR16, RAW^{+}16, RHR^{+}21, Ano97d, Bra97, CDSV10, CDSV11, MBR21, Ste08, TACA15, Váz16]. MATLAB-like [Ano96d]. MATLAB/GNU [FRAK15, JRA^{+}18, PSR16, RAW^{+}16, RHR^{+}21]. MATLAB/Octave [AMR18, PC13a, CDSV10, CDSV11]. Matrix [Edd96, GBDM77, JWC18, SBD^{+}76, RBM^{+}23, Sai13, SDL^{+}16, WPAV14]. Matter [MI07]. Matters [Mog01a, Mog01b, Mog01c]. Matthew [Cha13, Teo13]. Mature [Bra04, MS12]. maturity [KKA^{+}19, WMLM22]. Maverik [Wes00]. Max [Ano00k]. Maxed [CGK^{+}02]. Maxima [LR08]. maximize [Fra13]. Maximum [WM05]. Maxspeed [Ano00j]. Maxwell [RJ21]. May [AK95, Bar00a, Dig82, DMP^{+}02, HDR03, IEE95a, MSLH71, PM00]. MBDyn [ZAC^{+}23]. mbsolve [RJ21]. McCarty [Ano00a, Ano00b]. McLean [ACM93b, ACM94]. MCMAS [LQR17]. MD [Aji17]. MDD [Ano01j]. MDK [Rui02]. Me [CRW^{+}04, Lit14, STG19, Ste01, WWSG21]. Mean [Bar00c, SHW^{+}21, Gv14]. mean-field [Gv14]. Meaning [Maj03]. Meanings [GA04a]. means [Zic01]. meant [TG99]. Measure [RT12, TTL06]. Measurement [Ano02b, BDP^{+}14, GF17, KT05, Lla06, SA15]. Measurements [Bes03]. measures [CHA06]. Measuring [APHV19, BSA22, WGS07]. mechanical [HMP^{+}15]. Mechanics [BCB07]. Mechanism [MSSvK08, ZZZ22, KTP95, RCO20]. Mechanisms [BE06, May17, PDG^{+}87]. Mechanistic [ORS^{+}14]. med [Jön05c, Lan89, MG05]. Media [Ano04b, HSF^{+}15, JP09b, KGW^{+}21, Kus05, SS23]. Median [NRG^{+}99]. mediated [BSP11]. Medical [Mam01, BTL^{+}11, YA05]. Medicine [PBJ^{+}12]. Mediterranean [Yuk94]. Medium [DWP^{+}14, KT04]. Meep [LFN^{+}11, ORI^{+}10]. meeting [Bon93, Jef08]. Meets [BW00, Got05, RCB^{+}14, Fär05, HBZ09]. Melding [Wes03]. Melodic [Fyk97]. membrane [CKS16]. Memcached [RVLS14]. memory [WK93]. Menu [RW89, Pal87, PDG^{+}88]. Menu-based [RW89]. Merge [GMBv20, ABC18]. Merged [Coc03]. merges [PdSCJM22]. Merging [Kit94]. Merits [Wat01]. Merlin [Ano00k]. Message [Kro99a, MT94]. message-driven [MT94]. Message-Queueing [Kro99a]. Messages [dM99, KN93]. messaging [RA16]. Meta [Tro96c, WLD^{+}17]. meta-analysis [WLD^{+}17]. metabolites [LSM09]. MetaCard [Ano97c, Kuk98]. MetaCreation [Kro99b]. Metadata [VOK^{+}22]. metaheuristic [DC23]. metaheuristic-based [DC23]. MetaSys [VOK^{+}22]. Metcalf [Ano96e]. Method [GD12, HMX21a, HMX21b, LMW12, CFCA13a, CFCA13b, DBLF16, FRAK15, HKvH16, HKY^{+}21, JRA^{+}18, KN93, MMY^{+}19, MVAXP22, NMS14, ORI^{+}10, PPR19, RAW^{+}16, SDL^{+}16, WNS^{+}21, YKK23, ZDM10, ASC^{+}21, CKB11]. Methoden [FG85]. Methodologies [DXT^{+}18]. Methodology [Maz15, McC99c, Aki16, BJWZ08, HPT17, LC12a, Rob05]. Methods [Ano01a, BNST99, DKMT11, KCAS23, TDBEE11, WM05, BTL^{+}11, BCPS10, FL15, GEI^{+}11, HWL^{+}23, LH22, NS05, PM21, PT91, RHR^{+}21, You08]. Metrics [MOMM11, Sha10, FRBRF19, GYW^{+}23, GFS05, JK11, JK12, KFYI13, LMZT22, SK12]. Metro [Ano98, Ano00i]. Metro-X [Ano00i]. Metrowerks [Kro99b]. mezzi [Zic01]. MFC [BSC^{+}21]. mfront [HMP^{+}15]. MIAOW [BGG^{+}15]. Michael [Ano96e]. MicMac [GBG^{+}16]. MICO [Ang01, Pud04]. Micro [KFYI13, Vol96, Wes00]. Micro-C [Vol96]. Micro-Kernel [Wes00]. Microblogging [WKS^{+}14]. Microelectronics [MPG^{+}16]. Microkernel [Bud10]. Microkernel-Betriebssysteme [Bud10]. micromolecular [XTG^{+}11]. micronC [Ano93a]. microprocessor [BY92, BY91]. microprocessors [But94, UZ97]. Microscopy [Ano16, BTL^{+}11, SDeaK^{+}09]. Microservice [DFP23, AKMS23]. microservices [AKMS23]. Microsoft [PKP02, PKP05, Ano02b, BCB^{+}17, Gal04, PKP05, Ray01a]. Microsystems [Ano00j]. Microtest [Ano00j]. Microway [Ano98]. MIDAS [GJMPAM^{+}14]. Middleground [WG00]. Middleware [WG00, AJ05]. MigraTEC [Ano01j]. Migration [Ano01j, BDAW15, KKA^{+}19]. Mike [Ano97a, Ste99]. MILEPOST [FMT^{+}08, FKM^{+}11]. million [Cre07]. Mimic [EKJ^{+}03]. Mind [AM03, PSL21]. minds [NS05]. Mineralogy [PH82]. Mini [Gra99]. Mini-Review [Gra99]. mining [ALVV17, HSX^{+}18, ZVvDD11]. MINIX [Ahm08b, TSM88]. Misconception [RAH^{+}01]. Misconfigurations [RSBP23]. Mismatches [ACC^{+}12]. miss [RC10]. misses [Sta09]. Mission [eLAA^{+}23, NK04]. Mission-Critical [NK04, eLAA^{+}23]. Mistakes [Bar00a, Gla08]. misunderstanding [MLA^{+}19]. Misunderstood [DM97]. mix [QC18, Rui02]. Mixed [KCAS23, ZK21]. Mixed-Methods [KCAS23]. Mixing [Wil13]. Mixture [AL07]. MIZ [LN92]. MIZ-PR [LN92]. MKSAWK [Mor87]. MLAnalysis [GFD^{+}24]. MLISP [Uni85c]. MMBase [BCvE^{+}05]. MNCaRT [AWD^{+}18]. mnemonics [Cra90]. MO [ACM97]. Mobile [Ano01i, Ano01j, SNF04, Cha11, Kus05, LGS^{+}17, MLWR18]. Mode [Che87a, Sha04, Che86, Che87b, Chi93, MC91, SDL^{+}16, Vie97]. Model [AL07, Ano01j, CCG^{+}02, CPJ^{+}98, CK10, HR11, LFB^{+}21, LLWM23, LL14, NR03, Rie21, Ste00a, UMV15, VJ23, WM19, ZZZ22, AH19, ASAB02, BCP^{+}16, CFCA13a, CFCA13b, DDHS03, Gv14, JD19, KN93, KGW^{+}21, KG20, eLAA^{+}23, LQ17, LQR17, LGA20, MLMFN^{+}15, Pya06, SK12, SM08, Tai13, TPK^{+}21b, Wan21, ZE03, ZD05, ZCG17]. Model-Based [CPJ^{+}98]. Model-driven [LFB^{+}21]. Modeling [BSA22, DDJ99, JPOB20, KC21, KG20, Owe01, PMBM^{+}15, WCHRM21, YLXZ16, DSM^{+}19, KF17, KSD^{+}12, LA10, LGA20, MGPB20, NZPWR22, NGCI^{+}12, YMCF23, eLAA^{+}23]. Modellbahnsteuerungssoftware [Ano01c]. Modeller [Jön05c]. Modelling [TTB09, HMP^{+}15, LH14, MVAXP22, ORS^{+}14, WNS^{+}21, WSK^{+}22]. Models [HR11, Jin18, Jön05c, LRP11, Raf23, RGCS14, Asu05, DPH16, GBG^{+}16, SS23, YT22]. moderating [SG06]. Modern [DKMT11, Far92, Fin91, LMZP19, TV13, THG20]. Modes [CWB^{+}04, Sin08, jFFR16]. Modification [Ano94a]. Modul [Per02]. Modula [Ano91, BB91]. Modula-2 [Ano91, BB91]. Modular [CFW17, PG02, UNF^{+}08, AJLM18, KSH14, RBM^{+}23, Zag14]. Modularity [NR03]. module [Hub04b]. Modules [Wel95, KT05]. Modulo [Hag04]. MoJo [Ano00j]. molecular [APK14a, APK14b, GJMPAM^{+}14, RAMB18, SMRM^{+}17, SC16, VBG^{+}10, Yap11, Yes12, ZJS^{+}20, KSD^{+}12]. Monatsblatt [Lin02b]. mondo [Cor05]. Mondrian [SRGCPB^{+}09]. Money [CKB^{+}05, Kam14a, Kam14b]. Monitor [BVLF14]. Monitoring [DFCPSF15, HMKC12, McC02b, Zha16, GBG^{+}16, JH16, Kis90, PH16]. Mono [Lov06]. monoclonal [Ewe18]. monokinetic [LMHL20]. MontaVista [DuB02]. Monte [Adk11, HWM^{+}15, Hua17, Hua23, MMY^{+}19, SMRM^{+}17]. Monterey [USE99, USE02b, USE02c]. monthly [Lin02b]. Montreal [Lev95a, Lev95b]. Moodle [Col05, CP04]. MooseFS [LQ17]. most [CK06b, Fie90a, WM01]. Mothballed [Bar00c]. Mother [Ano00j]. Motif [Ano98, AL07]. Motifs [AL07]. motility [SBM^{+}10, SMO^{+}13]. motion [Hol05]. Motivating [LMWM18]. Motivation [HNH03, Rie07]. Motorola [He95, Ho95]. Mouse [Mor92]. Move [Pau04, Ano93a, Hic04]. Movement [GF99, Ano01h, MdL09, SKB23]. Movies [EKJ^{+}03]. Moving [KKN^{+}21, Par91, ACM93a]. Mozilla [EKJ^{+}03, MFH02]. MP [GSR^{+}04]. mp3 [Zic01, EMD03]. MPEG [Chi01, Kro00]. MPEG-2 [Kro00]. MPFR [FHL^{+}07, LZ16, LZ17, Zim10]. MPI [AVA^{+}16, Coo95a, Coo95b]. MPI-based [AVA^{+}16]. MR [GV16]. Mr. [Lew97]. mRNA [SNC^{+}06]. MRS [PSS^{+}07]. MS [Ano93a, HWZxx, HWZ01, HSC89, Ohl92]. MS-DOS [Ano93a, HWZxx, HWZ01, HSC89, Ohl92, Uni85a]. MS-Windows [HWZxx, HWZ01]. MSL [KSD^{+}12]. MSP [YM93]. MST [EKJ^{+}03]. MUDABlue [KGMI06]. Mudflap [Eig03]. Mule [HNT93]. Multi [AWD^{+}18, AML^{+}10, BMB^{+}18, PBJ^{+}12, SSP17, SSP18, ACB^{+}16, AMR18, BSC^{+}21, DBP^{+}18, DFU20, GDJG23, GTMR23, LQR17, Nor23, WB07, YLXZ16, YKK23, ZRZ^{+}21]. multi-agent [LQR17]. Multi-Architecture [AWD^{+}18, Nor23]. multi-camera [GTMR23]. multi-component [BSC^{+}21]. Multi-Core [AML^{+}10]. Multi-Dimensional [PBJ^{+}12]. multi-GPU [DFU20]. Multi-Institutional [BMB^{+}18]. multi-operator [ACB^{+}16]. multi-phase [BSC^{+}21]. multi-physical [YKK23]. multi-physics [GDJG23, ZRZ^{+}21]. multi-platform [AMR18]. multi-purpose [DBP^{+}18]. multi-release [YLXZ16]. multi-resolution [ZRZ^{+}21]. multi-scale [BSC^{+}21]. multi-server [WB07]. Multi-Version [SSP17, SSP18]. Multicode [Mud97, Mud97]. multicomputers [MT94]. Multiconference [Ten93]. Multicore [PGW^{+}20, GJLT11, Nob08]. Multics [Gre80]. Multideterminantal [PHT17]. multilaterally [SS05a]. multilinear [GJMPAM^{+}14]. Multilingual [Mud97, HNT93]. multimanipulator [MWB89]. Multimedia [CO12, IEE95a, SuS01, ACM93a, Kus05]. Multimodal [MMD12]. multiphase [HSF^{+}15, SPLD20]. multiphysics [SPLD20]. Multiplatform [And11]. Multiple [ALA20, GSR^{+}04, Fär05, FHL^{+}07, KP93, TACA15]. multiple-precision [FHL^{+}07]. multiplication [GM94]. Multipole [CKB11]. multiprecision [Sai13, SC08]. multiprocess [VGD^{+}97]. Multiprocessor [ILG10, RTH15]. multiscale [FTZ^{+}23]. multisystem [DRM21]. multiuser [Wii91a]. Multivariate [Joh18, PC13a]. München [SG05]. Munich [SG05]. municipalities [KKA^{+}19]. muon [BFI^{+}21]. muon-nuclear [BFI^{+}21]. Music [CDsJ^{+}00, CSP^{+}03, Les03, Sta03c, EMD03, Tho90a, Tho90b]. Musical [TF21]. Musings [Ray99b, Ray99c, Ray01b]. Must [CSD^{+}05, Sta00a]. Musterlösungen [PKP02]. Musterlöungen [PKP05]. mVMC [MMY^{+}19, XOTI22]. mVMC-Open-source [MMY^{+}19]. MX [Ano02b]. MX4 [Kuk98]. My [mH00, LR11, Lus04, Sta06]. MyMolDB [XTG^{+}11]. MySQL [LW03, Mon03]. Mysty [Ste01]. myth [Sch11].

N. [Ano96e, Ano01a]. nach [WP04]. NAGWare [Ano01a]. naledi [BH17]. Named [GAS^{+}01]. Names [Coc01a, RAH^{+}01]. Naming [Ros01a, Ros01b, Ros01c]. Nancy [Bun94]. nanobodies [Ewe18]. Nanoelectromechanical [DDJ98a, DDJ98b]. Nanoengineering [Bar00a]. nanofiber [LPC^{+}15]. Nanoseconds [Bar00a]. National [Cha98, WBB^{+}74, MSLH71]. natural [PM21]. natural-orbital-functional-based [PM21]. Nature [GMBv20, SBDR22]. Naval [LMM02]. Navier [HWL^{+}23, MVS15]. Navigated [RLTD23]. navigating [Hol15]. Near [RC10, MZE13]. Near-miss [RC10]. nearly [LD13]. Nebulous [Mog03a]. Need [Coc01a, SS06, TGS22, Asu05]. needs [FvH03, Ous99]. Neidorf [DPL^{+}91]. Nektar [CMC^{+}15]. nell’era [Zic01]. Nelson [Ano00l]. NEMOH [KD23]. Net [FSB^{+}01, Kro99b, Kro99a]. NetBeans [Ano98, Sur04]. NetBSD [YSC^{+}06]. Netherlands [PT91]. Netscape [Cha98, SSC^{+}00]. NetWare [Ano98]. Network [AY93, Ano01j, Cha98, EKJ^{+}03, Hom00, Kis90, Kre03, RCB^{+}14, Ste00a, ZRNA20, AVA^{+}16, Ano99c, AG22, BP14, CLL05, Coc01a, KSV16, LS04, Sta96b, BVT06, KSS^{+}23]. networked [BBNP93]. Networking [ACM00, FQYS23, MSR10, SuS01, Lei93a, Mag01b, NN00]. Networks [MPG^{+}16, Mar22, Rus14, ACB^{+}16, AT92, HYA20, Sin10a, DFCPSF15]. Netzwerkspiele [Str94]. neuem [Ano01d]. neuen [PKP02]. Neural [KSS^{+}23, AT92, Ale92]. Newby [Bur04a]. Newcastle [IEE90]. Newcomers [STG19]. newmat [Edd96]. NewMedia [Ano00j]. News [Aki16, APK14b, Ano97d, Ano03e, Ant16, AS03, Bar00b, Bar00c, Bar00a, Bar01, Bra97, Cha98, Coc01a, Coc01b, Coc03, DDJ98a, DDJ99, GJMPAM^{+}14, Gla99, GB00, GAS^{+}01, HLS^{+}13b, Law09, Mar01, Mau05, McL05, PM00, QSX^{+}15, Sav23, SMRM^{+}17, DDJ98b, Sta04a, Sta06, SV03, Yes12]. newsletter [Fre87, Uni85b]. Next [GB00, GAS^{+}01, Lat03]. Next-Generation [GAS^{+}01]. NGSCB [And03]. NHS [Thi22]. NIC [RHW^{+}21]. NIC-CAGE [RHW^{+}21]. nicht [Sur01a]. night [Cre07]. nineteenth [IEE95b]. Ninth [USE00b]. nixes [Ano04a]. nm [CCA^{+}19]. NMF [WRDP17]. NMR [LSM09]. No [Ano15a, CPJ^{+}98, DuB02, Mic04, Sie04, Sta06, Ant16]. No.91TH0350 [MS91]. No.91TH0350-9 [MS91]. No.91TH0394 [MSNS91]. No.91TH0394-7 [MSNS91]. No.94CH35712 [IEE94c]. Nodal [WNS^{+}21]. noise [IHBS14]. noise-reduction [IHBS14]. Nolan [Ano00k]. Non [Goo14, KRB^{+}22, NN20, IC22, NN21, ZLF^{+}22]. Non-Adopters [Goo14]. non-coding [IC22]. Non-Coherent [KRB^{+}22]. non-matching [ZLF^{+}22]. Non-psychopath [NN20, NN21]. nonequilibrium [FTZ^{+}23]. Nonfree [Kos21]. Nonlinear [BFC02, CFCA13a, CFCA13b, DBLF16, Wan21]. Nonsense [Fal03]. Noordwijkerhout [PT91]. normalization [MZE13]. Norway [AK95]. Note [Mam01]. Notebook [Dum05]. Notes [AFS81, AFS82]. Nothing [SSC^{+}00]. Nottingham [SM89b]. Novel [MEB^{+}20, ZRNA20, ASAB02, GV16, KDM17, WZS^{+}18]. November [ACM89, ACM00, Ano14, IEE92c, USE01a, ACM97]. novice [Pec08]. Noweb [Ano10]. npm [CASA22]. NR [LLEL^{+}23]. ns [LLEL^{+}23]. ns-3 [LLEL^{+}23]. NSA [McC05]. NSF [Cha98]. Nuclear [BPG94, BFI^{+}21]. nucleation [WNS^{+}21]. Number [Fär05, Thi22, Sib17, VRS^{+}95, VRS^{+}99a, VRS^{+}99b]. Numbers [Sta04a, Whe03, MRS07]. Numerical [Ano01j, Lea94, TDBEE11, Wil71, XXAD21, Eat97, Eat00, Eat02, Eat05, EBH08, MN21, Nob08, NGCI^{+}12, VRS^{+}95, VRS^{+}99a, VRS^{+}99b]. Numerics [Cse99, Ano01j]. NuSMV [CCG^{+}02]. Nutan [TG15]. Nutch [CC04]. Nutzung [Sie99]. NWChem [VBG^{+}10]. NY [Kap92].

Oakland [USE01a]. OASIS [BBNP93, MVS15]. Oberammergau [BPG94]. Oberfläche [Ste00a]. Object [Edd96, LOW91, LO92, Mir07, Sha10, Ste00a, ZC95, CKH91, GFS05, Hin87, JK12, MR94, Zag14, Kro99b]. Object-Oriented [Mir07, Sha10, Edd96, ZC95, CKH91, GFS05, MR94, Zag14]. objective [Oma89, Sal88]. Objects [SSC^{+}00, BG95]. ObjectSpace [Ano96c, Ano96c].

ObjectTeam [Ano98]. objektorientiert [Ste00a]. OBOSS [VGdlP01]. Observations [KKN^{+}21, SDD05, Bur04a, FG16]. observatory [VSN22]. Obstetric [PH16]. ocean [DBLF16, DBLF16]. Oceanographic [LMM02]. Ockman [Ano99a]. OCL [MNS19]. Octave [Eat00, EBH08, FRAK15, YDZ19, MN21, Ano96d, AMR18, CDSV10, CDSV11, Eat97, Eat00, Eat02, Eat05, Edd00, FY18, GHH20, Hei16, JRA^{+}18, Joh18, Jön05c, LGW^{+}22, MBR21, MAMC05, Mar22, PC13a, PSR16, RAW^{+}16, RHR^{+}21, Váz16, Bar16a]. Octave/MATLAB [Joh18]. Octaves [Fyk97]. OctCNN [LGW^{+}22]. October [ACM88, ACM93a, Ano06, Bao93, BSW95, BPG94, Bon93, CS96, EHP94, PT91, USE88, USE00a]. ODBC [Ano96b, Ano00i]. ODBC-ODBC [Ano00i]. OECP [YMCF23]. off [CFMRL11, KS03, Zag14]. Offended [RAH^{+}01]. Offene [HK03]. offered [Lam09]. Offers [Ano01j, Ano04b, Avi98]. Office [BH11, GGK99, GvdHPR14, LMM02, XWZ^{+}23]. official [Cha13, PS^{+}09, War04]. Offline [Ano96c]. Offshoring [O’D07]. Ogg [RAH^{+}01]. OGSA [BE06]. Oh [Sta06]. Ohio [Bon93]. OK [CSD^{+}05]. OLAP [SRGCPB^{+}09]. Old [SS06, Rob11]. Omnicore [Ano01j]. On- [Tra95]. On-Chip [KRB^{+}22]. On-Line [Ano00l, GM02, KN93]. on-the-fly [BGM99, RBM^{+}23]. Onboarding [FGBM14, STG19]. One [Ano00j, Bes04, MZH22, Bou05, JWC18]. online [ACM93a, PDG^{+}87, Pal87, PDG^{+}88]. Only [HW17b]. Ontario [ACM93a, HDR03, HDR04, Ass95]. Ontology [GLMS18]. OO [CGB^{+}05]. OP31.05 [PH16]. OPAL [Mei92]. Open [Abe07, ASS^{+}23, AtHR11, Adl00, ABC^{+}14, APCs22, ATM22, Alf05, All02a, All02b, AM03, AM04, AMS03, And08, Ang01, AWD^{+}18, ACC^{+}12, AHB^{+}09, Ano96c, Ano97c, Ano00e, Ano00f, Ano00h, Ano00k, Ano00l, Ano01e, Ano01g, Ano01h, Ano01i, Ano01j, Lin02b, Ano02b, Ano03a, Ano03b, Ano03c, Ano04b, Ano04c, Ano08b, Ano08c, Ano11, Ano15c, Ano16, Ano18, Ant16, AML^{+}10, AS03, ALA20, AD04, AHM^{+}07, ACHC11, BdP13, BGG^{+}15, BHMB03, BRH10, BMF^{+}16, Bal19, BJWZ08, BC20a, Bar00c, Bar01, Bar22, BKR^{+}20, BSFR22, BdSI15, BDAW15, BY14, BlG12, BMZ14, Bax01, BCP^{+}16, Bea21, BK14, BCB07, Bel22, BYV08, BAP00, BM12, BVLF14, BNSW15, Bla06, BSA22, BKHT21, BNST99, Bol02, Bon11, BGG^{+}00, Bor09, BJJ14, BCB^{+}17, BSA14, Bou05, Boy08, BKP05, BP14, Bro01]. Open [BB02, BK02, BGL^{+}21, CC04, CFM08, Car01, CCA^{+}19, CF07b, CO12, Cha01a, CH10, Cha11, Cha07, Cha98, CP04, CTP^{+}22, CSY^{+}04, CWM^{+}20, CZ22, CX23, CYL^{+}23, Chi01, CYOS19, CCSW10, CC03, CBB06, CPJ^{+}98, Coc01b, Coc03, CPG^{+}04, Coh02, Col09a, CK10, CSEP14, Cou20, Cow03, CWB^{+}04, CMJ^{+}04, CRW^{+}04, CKB^{+}05, CGB^{+}05, CDR^{+}15, CH06b, CB12, Cur99, CdSV07, Dal02, DSK19, DBBA10, DMJ05, DXT^{+}18, III01, DB02, De’15, DKK22, Del01, Den13, DWP^{+}14, DKMB14, DFP23, DOS99, DKMT11, Doh01, DM15a, DFT21, DBLF16, DGC^{+}07, DMP^{+}02, Dwa04, Ebe07, Ebe08, Ebe09, EW01, Egy01, EE01, EJS^{+}01, EGK^{+}02, Ell12, Eng10, Ens04, Ens05, Eri99, Eri00, Eri01, Est06, Ewe18, FGBM14, Fal03, FL16, Far06, FSB^{+}01, Fav23, FFvdH01, FFHL05, FFH^{+}05, FFHL07, Fer03, Fit04]. Open [Fit11, For07, Fox08, FT09, Fra19, Fre23, FQYS23, Fug03, FVD^{+}12, FCTP21, GP12, Gag02, GKL^{+}14, Gal10, Gan17, Gar09, GIA^{+}06, GLT08, GF11, Gau07, GDK21, GD12, GKP^{+}14, GMBv20, Gil06, GCE^{+}21, Gla99, Gla00, GGL21, GB21, Goo14, GAS^{+}01, Got05, Got07, Gra01, GM05, GVOM09, GGB17, GNR^{+}09, Gv14, Gri16, GW09, Gro01, GEMN07, GPPT16, GTMR23, GNGS17, Gus20, Gut00, HK03, HKA^{+}19, Hae02, Haf01, HMO^{+}18, HMYH22, Han00, HT21, HKP02, Har99, Har20, HCH^{+}20, HOL^{+}07, Hau01, Hay05, HLS^{+}13a, HLS^{+}13b, Hec99, HR11, Her20, HBB^{+}12, HE17, HM19, HW17a, HMKC12, Hoh01, Hol23, HBZ09, HKY^{+}21, Hub04a, HBGS19, IHBS14, IAS16, IKW23, ILG10, IC23, JLL23, Jen01, JPOB20, Jin18, JEB^{+}23, Joh99, Joh02, JJ00, Jon01, Jon02, JS07, Joy08, Joy09b]. Open [JČMG11, Kam21, KC21, Kar03, KNS18, KTH^{+}22, KGM^{+}16, Kenxx, KY16, KCAS23, KMF^{+}07, Kim01b, Knu99a, Knu99b, KKN^{+}21, KH05, KHA^{+}03, KKA^{+}21, Kre03, KS11, KJRD16, Kri03, Kro99a, Kuc06, KG01, KD23, KRB^{+}22, Kus05, Lam09, Law02, Law09, LW03, Lei04, LSJ^{+}06, LMM02, Lev23, Lew99a, Lew99b, Li18, LPFD21, LMPT22, LZWH22, LLWM23, LL14, LRBM23, Lin08, LRP11, Lin02a, Lit14, LC12b, Luc99a, LGW18, LBF^{+}22, MPG^{+}16, MSSvK08, MMD^{+}22, Mam01, MMP^{+}22, MD17, MD18, Man03, Mar11, MTM^{+}19, Mar01, Mar05, MH07, May06, Maz15, McA19, McC99c, Mee12, MSW09, Men10, MFS15, MCGA22, MMD12, MMY^{+}19, MN04, MK12, Mog01c, MOMM11, MAF22, MS12, Mor08, MLWR18, Mur20, MEB^{+}20, MB16, Nas04, Nej12, NN20, NGJ03, NO03, Noj01, O’D07, Opexx, OG07]. Open [Omb20, Owe01, PSSH16, PMM17, PMM18, PLZ^{+}22, PMBM^{+}15, PM00, Pau04, Ped05, PBOP07, PQM11, PMD13, Per00, PH16, PPRB07, PGW^{+}20, PZ20, PLO^{+}23, PK10, PRRL12, PGC21, PSP^{+}22, PFL^{+}12, PBH01, Pot06, Pri19, PMG^{+}09, PKG^{+}10, PPG^{+}11, PBJ^{+}12, QC18, Qui00, Raf23, RSBP23, RB05a, RT12, Raj23, Ray98, Ray99b, Ray99a, Reh01b, RDKT12, Rie07, Rie10, Rie11, Rie15, Rie19, Rie20, Rie21, RCP^{+}12, RGCS14, RSAT19, Rob20, RLVdS21, Ros02a, Ros01a, Ros14, RCB^{+}14, RLTD23, Rus14, SD16, SJV^{+}05, SSAO04, SBDR22, San98, San08, SS02, Sav23, SFF^{+}06, Sca06, Sca19, SB08, SC02, Sch09, Sch19, SvGH15, ST10, Sea99, Sea02, SG05, SS06, SK04, Sha10, SSM^{+}07, SSH22, SAC^{+}15, Sie04, Sif00, SKSM19, SV19, SCB04, Sim12, SSP17, SSP18, SCDS15, SL01, SSS^{+}14]. Open [SFWD12, Sor06, Sor01, Spe01, Spi03, Spi06, Spi11, Spi19, Sta04a, ACM05, Sta98b, SVAGB20, Sta06, Sta02c, SHK^{+}03, SJW22, STG19, Ste99, SDD06, SF15, SV03, Sto99, Str02, SH19, TPK21a, TZ22, TGC^{+}20, TBPS15, TRM16, TNM17, Ter00, Thi22, TRB22, TTB09, Tot06, TGC^{+}21, TGW^{+}22, TGS22, TWS^{+}22, TG15, TF21, The04, UMV15, UNF^{+}08, Van22, Veg06, VVM08, VOK^{+}22, VJ23, VOM12, Waa09, WCHRM21, WCS20, WACBL03, WW01, WKS^{+}14, WCA^{+}14, WM19, WFW^{+}20, Wan21, WCG22, WGG16, WGG^{+}19, Wat01, Wea03, Wen00, Wen02, WWSG21, WG00, WDK^{+}20, Whe03, WB02, WM05, WKB14, Wil99, WLC01, WBB01, WBGM02, Wol03a, WBRH23, WNS^{+}21, WSK^{+}22, Woo01, WG05, WL01, WKA^{+}08, Wut12, XWZ^{+}23, XMGM21, XXAD21, XGF^{+}23, YLL^{+}07, YMLT14, YA05, Zha16, ZRNA20, ZZZ22, ZXB^{+}23, ZD05, ZK21, dCdCM14]. Open [dBLMT11, dlCKK15, vdLLM09, vWHvW09, ABC18, AKHG16, AMOS19, ADF^{+}21, AH19, Aji17, AW07, ACB18, Aki16, ASAAM^{+}19, ACB^{+}16, ALGE12, APK14a, APK14b, ABF^{+}14, AJLM18, AMWH19, AMC16, AAB^{+}05a, ABNÅ05, AAA^{+}12, AAA^{+}14, ALVV17, AM18, ACKT20, Amb15, And01, AVA^{+}16, Ano99a, Ano99c, Ano00g, Ano01f, Ano02a, Ano03e, Ano04a, Ano04d, Ano19, Ano21, ASAB02, AKF21, AAB^{+}05b, AG22, AKMS23, Asu05, ASC^{+}21, APHV19, AZ17a, AZ17b, AFZ17, AFZ18, ATCZ19, BTL^{+}11, BM02, BOL14, BCR^{+}08, BMF^{+}19, Ban16, Ban17, BB08, BCPS10, BA15, BD03b, Bea04, BCvE^{+}05, BLG^{+}17, BH11, BBE^{+}20, Ber22, BSW^{+}14, BCI^{+}09, BAR16b, BC20b, BSK^{+}15, BM22, BMT^{+}20, BFI^{+}21, Bon02, BCG^{+}14, BAE14, BZB17, BMS^{+}22, BG12, BDP^{+}14, BSC^{+}21, BGL^{+}20, BGL^{+}22, CF07a, CK08, CGZ17, CCA^{+}13]. open [CMC^{+}15, CAWK22, Cap12, CFGS05, Cap13, CZS^{+}21, CFL23, Cas19, CV13, CV22, CRB^{+}18, CM06, CNSR23, CLL05, CLM^{+}08, CJ17, CJ19, CFW17, CG17, CKS16, CH11, CKGW22, CBRSH22, Cio01, CFW01, Col05, Col09b, Com99, CH06a, CMTA19, CSP09, CSV^{+}07, CHA06, CWHW12, CWZ06, CKB11, Cus04, DPH16, Dan11, DGJH19, DRM21, DSB^{+}16, DIK^{+}23, Dei10, DFCPSF15, DP09, DWJG02, DBP^{+}18, DD17, DSM^{+}19, DFU20, DTB05, DDA^{+}07, DD08, DM15b, DD10, DO16, EMD03, ESM19, Eds16, EKUR10, EHHH06, EMdL^{+}07, Emb06, EHP14, ES23, Eub05, FLA^{+}16, FN21, Far23, FL15, Fei23, FBY^{+}17, FTZ^{+}23, FMFZ19, FHH11, Fog06, For12, jFFR16, Fow00, FM10, Fra13, FRBRF19, FG16, Fri06, Fri16, GBG^{+}16, Gal01, GLMS18, GGSRMPM20, GLCMC17, GEI^{+}11, Gau03, GMPS14, GHL^{+}04, Gen99]. open [GCK^{+}17, GGT05, Ger03, GDJG23, Gla03b, Gla04, Goe07, GF17, GRJS01, GSW08, GV16, GFZ16, GYW^{+}23, GFD^{+}24, GGH05, GGH10, GW10, GFS05, HK09, HAC^{+}23, HBC^{+}05, Har05, HBR19, HXS20, HFO^{+}12, HWL^{+}23, HZ14, HMP^{+}15, HPM^{+}08, HPT17, Hic04, HL02, HETD09, HJ07, Hol15, HSF^{+}15, HKvH16, HYA20, HWM^{+}15, Hua17, HSX^{+}18, Hua23, HZS^{+}16, HMX21a, HMX21b, HM10, IDSM23, IC22, JP09a, JCNS^{+}22, JP09b, JH16, JNN12, JK11, JK12, JDB09, Jør01, Joy09a, JZ09, KTF15, KKT17, KDM17, KC22, KMG^{+}07, KHMA12, KPK^{+}17, Kle21, KS02, Koc07, Koc09, KGW^{+}21, KSH14, KKA^{+}19, KK17, Kor11, KT05, KL07, KGT22, KSS^{+}23, KTTK17, KRR23, KFYI13, KSD^{+}12, KG20, eLAA^{+}23, LA10, LLEL^{+}23, LPC^{+}15, LG02, LFB^{+}21, LSM09, LR08, LGS^{+}17, LQ17, LMHL20, LMZT22, LRP21, Liu08, LHZ12, LRD^{+}19, Lla06, LQR17]. open [LZ11a, LZ11b, LZ12, LH22, LH03, LH14, LGA20, MG12, Mac18, MV05, MLZ^{+}23, MBTB21, CCK21, MSB09, MLMFN^{+}15, MFB23, MGPB20, Mas05, May17, MCS12, McA08, McC05, McC02b, MLA^{+}19, MTBS09, MDRN18, MWG08, MNS19, MZE13, MVF20, Mił10, MGYC18, MSM10, MPE^{+}11, MFH02, Mon03, MM04, Moo01b, MVAXP22, MdL09, MTD^{+}09, MVS15, MGFRG12, MSR09, MSR10, MCQF21, MOT^{+}18, MRN20, Muw09, MQN19, NRRS20, NYB10, NXC13, NN00, NMG11, Neu99, NS05, NN21, NDDH^{+}21, NT06, NMS14, Nob08, Nor23, NZPWR22, NGCI^{+}12, NMX19, O’R99, O’S03, OMA^{+}22, ORS^{+}14, ODP15, PCAJ^{+}23, PM13, Pag07, PKGA22, PKB17, PAB^{+}17, PSE04, Pay02, Pea16, PL05, PPC^{+}15, Pet06, Phi12, DARJ23, PSDG18, PM21, Pit16, PS^{+}09, PWA^{+}19, PNK^{+}23, PHT17, PPR19, PC13b]. open [PYM^{+}06, PSS^{+}07, Pow14, PSL21, PdSCJM22, Pya06, QB21, QLC^{+}12, QSX^{+}15, Raj13, RZWW23, RBM^{+}23, RCO20, Ray01b, Ray01a, RHW^{+}21, RH21, RJ21, Rob05, RCGB^{+}22, RAMB18, RP08, RA16, RNR17, RC10, Rud10, RT05, SS05a, SZ05, SBS20, Sal08, Sam06, SBM^{+}10, SMO^{+}13, San01, Sca05, SA15, SPLD20, Sch11, SGM^{+}08, Sch04, SHB^{+}20, Sea04, SRGCPB^{+}09, SMRM^{+}17, SSR02, SSS22, SIK^{+}13, SC16, SPAW17, Sil13, Sim05, Sin08, Sin10b, Sin10a, SK12, SHW^{+}21, SGNB08, SKB23, SCR05, SCFR06, SM08, SAHP15, SSA08, SG12, Spi21, CAC09, Sta09, SAOB02, SS23, Ste08, SDD05, SG06, SDL^{+}16, STB23, SDeaK^{+}09, SWTC23, TZH22, Tai13, TPSZ19, TLL^{+}14, TL17, Tay19, TPK^{+}21b, TTL06, TACA15, TG99, TV13, THG20, THG23, TKSC20, UBR^{+}17, VGSN18, VBG^{+}10, VGP^{+}19, VSdCCR23, Vir05, VSN22]. open [VB19, WJM22, WLD^{+}17, WFF18, WGS07, WHJ15, WFDK19, WMLM22, WBY^{+}08, Web04, Wes03, WG06, WFV14, Wol02, WPAV14, WMK^{+}17, WZS^{+}18, XTG^{+}11, XFS^{+}22, XMGM22, XAPK14, XTY^{+}22, XOTI22, Yad07, YM93, YLG05, YWA07, YLXZ16, YLHW21, YZC22, Yap11, Yes12, YA11, YT22, YKSH20, YMCF23, YKK23, You08, Yu06, YSC^{+}06, ZVvDD11, ZFD21, ZKDP22, ZAC^{+}23, Zei03, ZK05, ZSW14, ZRZ^{+}21, ZE00, ZE03, ZDM10, ZW17, ZLF^{+}22, ZFY^{+}19, ZJS^{+}20, ZWH21, ZWU22, ZCG17, Zic01, ZLL04, dA15, dlVRB21, vGPB10, vKSL03, vKvH03, Ano00d, Bar00b, BS14, BES^{+}01, BW00, BR03, BE06, Cap12, CFMRL11, Coh03, CF09, CDsJ^{+}00, CRW^{+}04, DOS99, DiB04, DCS05, DFLS05, DH01, EW01, Eri00, Feixx, FFH^{+}05, FK04, GA04a, GL14, Gom99, GF99, HK03, Hac98, HHG^{+}21, Has05, HNH03, JV01, JWC18, KGMI06, KT04, Kro99a]. Open [LFN^{+}11, Lee99, LMWM18, Lin02a, LLS11, Lus04, MD04, Maj03, Mal02, MSZ^{+}01, Man03, Mau05, McG01, McL05, MP12, Moo01a, MM10, NR03, NRG^{+}99, NO03, NK04, O’Rxx, Oms03, Per05, Pra03, Pud04, Ray99c, Rob20, Ros01b, Ros01c, Ros00, Rud10, RE04, SSC^{+}00, Sca04, SMS16, SGM^{+}08, SG05, SCSC04, Shi12, Sie04, Spe01, Spi03, SS04, St.04, Ste99, Sur01a, Sur01b, TH04, VSM06, Wal01, WP04, Zag14, Cho09]. open-ended [YMCF23]. Open-Right [Sur01a]. Open-Source [AWD^{+}18, Ano00l, Ano01j, Ano04c, Ano08c, ALA20, BdP13, BGG^{+}15, BC20a, BNSW15, BSA22, BNST99, Bol02, BJJ14, BSA14, Boy08, Bro01, BB02, CCA^{+}19, Cha07, CTP^{+}22, CSY^{+}04, CX23, CYL^{+}23, CC03, CBB06, Coc03, Cow03, DXT^{+}18, DKK22, DKMB14, DFP23, Doh01, FL16, FSB^{+}01, Fre23, FQYS23, GKL^{+}14, GCE^{+}21, Gut00, Haf01, Hau01, Hec99, ILG10, Jen01, JPOB20, JJ00, Jon02, KY16, KCAS23, KMF^{+}07, Kri03, KRB^{+}22, LZWH22, LLWM23, LGW18, MPG^{+}16, MMD^{+}22, MMP^{+}22, McC99c, MCGA22, Mor08, MEB^{+}20, MB16, Nej12, NO03, Owe01, PLZ^{+}22, Per00, PGW^{+}20, PLO^{+}23, PGC21, PSP^{+}22, PMG^{+}09, PKG^{+}10, Ray99a, RDKT12, RGCS14, RLVdS21, Ros02a, Ros01a, SJV^{+}05, Sea02, SG05, Sha10, SSM^{+}07, Sif00, SL01, Sta02c, SHK^{+}03, SJW22, SH19, TPK21a, TGC^{+}20, TRB22, UNF^{+}08, VOM12, WCHRM21]. Open-Source [Wat01, Wen00, Wen02, WG00, WDK^{+}20, WB02, WM05, Wil99, Wol03a, WKA^{+}08, Wut12, XXAD21, YLL^{+}07, Zha16, ZRNA20, ZZZ22, dlCKK15, ATM22, And08, Ano00h, BCP^{+}16, Bor09, Bou05, Car01, CP04, CYOS19, CDR^{+}15, DSK19, DBLF16, Ewe18, FVD^{+}12, GDK21, GVOM09, Gv14, GTMR23, HMO^{+}18, HMYH22, HLS^{+}13a, HLS^{+}13b, HBB^{+}12, HBZ09, HKY^{+}21, IHBS14, JEB^{+}23, Kam21, KTH^{+}22, KD23, Kus05, LSJ^{+}06, Mam01, Mar11, MAF22, NN20, Ped05, Ros14, TGC^{+}21, VOK^{+}22, WCS20, WBB01, WNS^{+}21, WSK^{+}22, Woo01, XMGM21, dCdCM14, ABC18, AMOS19, AH19, Aki16, ALGE12, APK14a, APK14b, ABF^{+}14, AAB^{+}05a, ACKT20, Amb15, And01, AVA^{+}16, Ano99c, Ano00f, Ano04a, Ano21, AAB^{+}05b, AKMS23, ASC^{+}21, AZ17a, AZ17b, AFZ17, AFZ18, ATCZ19, BOL14, BCR^{+}08, BCPS10, Bea04, BCvE^{+}05, BLG^{+}17, BH11, BBE^{+}20, BSK^{+}15]. open-source [BM22, BMT^{+}20, BFI^{+}21, Bon02, BDP^{+}14, BSC^{+}21, CCA^{+}13, CMC^{+}15, Cap12, CFGS05, CZS^{+}21, CV22, CRB^{+}18, CNSR23, CLL05, CFW17, CG17, CKGW22, CBRSH22, Cio01, Com99, CMTA19, CSP09, CSV^{+}07, CWHW12, DPH16, Dan11, DGJH19, DIK^{+}23, DBP^{+}18, DD17, DSM^{+}19, DFU20, DO16, EMdL^{+}07, Emb06, ES23, Eub05, FLA^{+}16, Fei23, FBY^{+}17, FTZ^{+}23, FHH11, For12, jFFR16, Fow00, FG16, GBG^{+}16, GLMS18, GGSRMPM20, GLCMC17, GCK^{+}17, GGT05, GDJG23, GF17, GFZ16, GYW^{+}23, GFD^{+}24, HAC^{+}23, HBC^{+}05, HXS20, HFO^{+}12, HWL^{+}23, HMP^{+}15, HPT17, HSF^{+}15, HYA20, HZS^{+}16, HMX21a, HMX21b, IDSM23, JCNS^{+}22, JNN12, JZ09, KDM17, KMG^{+}07, KPK^{+}17, KGW^{+}21, KSH14, KT05, KL07, KGT22, KSD^{+}12, KG20, eLAA^{+}23, LA10, LLEL^{+}23, LPC^{+}15, LGS^{+}17, LQ17, LMHL20, LRP21, Liu08, LHZ12, Lla06, LQR17, LZ11a, LZ11b, LZ12]. open-source [LH22, LH03, MLZ^{+}23, MBTB21, MLMFN^{+}15, MFB23, MGPB20, McA08, McC02b, MLA^{+}19, MDRN18, MWG08, MNS19, MZE13, MVF20, Mił10, MGYC18, MSM10, MPE^{+}11, MVAXP22, MGFRG12, MCQF21, NXC13, NMG11, Neu99, NS05, NDDH^{+}21, Nob08, Nor23, NZPWR22, NGCI^{+}12, O’R99, OMA^{+}22, PSE04, PPC^{+}15, DARJ23, PM21, PNK^{+}23, PHT17, PC13b, PYM^{+}06, Pow14, PSL21, PdSCJM22, QSX^{+}15, RZWW23, RBM^{+}23, RHW^{+}21, RH21, RJ21, RAMB18, RP08, RA16, RT05, SBS20, SBM^{+}10, SMO^{+}13, SA15, SPLD20, Sch04, SHB^{+}20, SRGCPB^{+}09, SC16, SPAW17, Sim05, Sin10a, SHW^{+}21, SGNB08, SKB23, SAHP15, SS23, SDeaK^{+}09, SWTC23, Tai13, TPSZ19, TPK^{+}21b, TTL06, TACA15, TG99, TV13, THG20, TKSC20, UBR^{+}17, VBG^{+}10, VSdCCR23, VSN22, VB19, WJM22, WLD^{+}17, WFF18, WG06, Wol02, WPAV14, WZS^{+}18, XAPK14]. open-source [XOTI22, YLG05, YWA07, Yes12, YKSH20, YMCF23, YKK23, Yu06, YSC^{+}06, ZFD21, ZKDP22, ZAC^{+}23, ZSW14, ZRZ^{+}21, ZLF^{+}22, dA15, HK03, McG01, NO03, SG05, Spi03, Sur01a, Sur01b, WP04]. Open-Source- [TNM17]. Open-Source-Based [HMKC12]. Open-Source-Geld [Cap12].

Open-Source-Software [HK03, Rud10, Sur01b]. Open-Source-Software-Entwicklung [O’Rxx]. Open-Source-Systems [Ano00f]. open-sourcing [PSDG18]. Open-standard [AHM^{+}07]. OpenAD [UNF^{+}08]. OpenAD/F [UNF^{+}08]. OpenBSD [YSC^{+}06]. OpenCAPWAP [BCI^{+}09]. OpenCert [BS14]. OpenChain [Cou20]. openclean [MCQF21]. OpenDesk.com [Ano00k]. OpenEC [Sca05]. OpenEC/B [Sca05]. opened [SIK^{+}13]. Opener [Bon11]. OpenFFT [DO16]. OpenFlipper [MK12]. OpenFlock [Sif00]. OpenFOAM [SWTC23, TPK^{+}21b, ZCG17]. OpenFPGA [TGC^{+}20]. OpenGeo [WCA^{+}14]. OpenGL [Ano96c, FHH11]. Opening [CdR99, GB00, JWC18, NS05]. OpenLB [KKA^{+}21]. OpenMEE [WLD^{+}17]. OpenMP [Nob08, SPDQ22, YSVM^{+}16, YSMA^{+}17]. Openness [HW17b]. Openoffice [Ano01c]. OpenOffice.org [Ano02b, Weh03]. OpenOrb [GVOM09]. OpenPath [Ano01i]. OpenPiton [BMF^{+}16, BMF^{+}19]. OpenPLC [AM18]. openPSTD [HKvH16]. Opens [Kro00, Van22]. OpenSolaris [FT09]. opensource [MGR16, Sur04, CCG^{+}02]. OpenSources [Ano99a]. OpenSSL [WN15]. OpenStack [IHSR19, ZFY^{+}19]. OPERA [CLL05]. Operating [Maz15, MHP94, Reh01b, TF21, And01, AAB^{+}05b, Bud10, CK06a, CK06b, CK06c, CFW01, Don04, GPPT16, Mac99, PH82, RÓ01, WB07, YM93, Ygg93, Ygg94, YSC^{+}06]. operating-system [AAB^{+}05b]. Operation [CCA^{+}19, KS02]. Operations [AS97, LH03, MMR95]. operativo [RÓ01]. operator [ACB^{+}16, FRAK15, RAW^{+}16]. ophavs [MG05]. ophavs- [MG05]. Opinion [Hic04, Sec95, BHMB03]. Opportunities [CGB^{+}05, Kim01b, SDD06, Cap13, HKP02, MV05]. Opportunity [Law02]. Opposition [Gla99, Gla00]. optical [CKS16, MSB09]. optical-properties [MSB09]. Optimal [Sha03, RHW^{+}21, Ste00a]. Optimization [BdP13, CX23, Jon05b, Kro99b, LH03, Nov03, CHE^{+}10, DVC^{+}07, Hub04b, CCK21, XFS^{+}22, ZRGJ21]. Optimizations [CZ99, Ede04, VOK^{+}22]. Optimized [LZ16, LZ17, XOTI22]. Optimizer [Dvo03]. optimizers [KSK09]. Optimizing [Bes03, EM93]. optional [Ant16]. Oracle [Ano00e, DH01]. Oram [Ano97a]. ORB [McC02b]. orbit [GVOM09]. orbital [PM21]. Orbits [ALA20]. orchestration [Far23, PPC^{+}15]. Order [CCA^{+}19, Mam01, BSC^{+}21, Car89, DFU20, HWL^{+}23, JD19, KD23, YKK23, Zag14, DBLF16]. orders [SG99]. ordinary [MZE13]. ordination [Ano02a, KS02]. O’Reilly [Ano97a, Ano99a, Ano00a, Ano00b, Cas02, Mar01]. Organization [MH07, Gal01]. Organizational [GF11, Goo14, JEB^{+}23, LBF^{+}22, NYB10, SG12, ESM19]. Organizations [SBDR22, SF15, BCG^{+}14, Sil13]. organized [jFFR16]. Orientation [RDKT12]. Oriented [Ano02b, Coh82, Kro99b, Mir07, Sha10, CKH91, CC05, CSP09, DFU20, Edd96, Emb06, GFS05, HOST05, Hin87, JK12, KF17, MR94, RP08, Sch91a, SLC88, Zag14, ZC95, vGPB10]. Origin [Sal20]. Original [HR19, VJ23]. Origins [Sch10]. ORK [VGdlP01]. Orlando [SS93]. Orleans [USE98a]. OS/2 [DF00]. oscillations [Dan11]. OSes [Den99, MI07]. OSGeo [BK14]. OSGi [GHM^{+}05]. OSINT [QC18]. OSMOSIS [LA10].

OSPREY [HMO^{+}18, Liu06]. OSS [KF17, Muw09, ACM05, Whe03]. OSS/FS [Whe03]. OSSARA [LMPT22]. Österreichischem [WP04]. Other [Mar01, Bou05, EKUR10, JP09b, Les99]. Otolith [PSP^{+}22]. otras [Les01]. Ottawa [HDR03, HDR04]. our [BH17]. Out-of-Order [CCA^{+}19]. outcomes [STB23]. output [She07]. overcomes [Mud97]. Overdog [CPJ^{+}98]. Overflow [Wag03, Ano99b]. overflows [ZLL04]. Overhauling [Zad02]. Overrun [CPG^{+}04]. Overview [JLL23, Mur94, RDKT12, RHR^{+}21, WL01]. Ownership [Gho07]. oxide [BCP^{+}16, NGCI^{+}12]. Oxymoron [Law02].

P [Ano00i]. P-STAT [Ano00i]. PA [Ano00f, BSW95]. PA-RISC [Ano00f]. PA-Risc-CPUs [Ano00f]. Paced [Ell12]. PACHI [BlG12]. Pacific [IEE94a, GAS^{+}01]. Pack [G^{+}02]. Package [Ano95a, BJJ14, CX23, Cod75, Cou13, MCS12, Ano99b, ASC^{+}21, BM22, CG17, CSV^{+}07, DO16, EHHH06, EHP14, HAC^{+}23, Hei16, HMX21a, HMX21b, KPK^{+}17, Kop20, LH22, MSB09, Mar22, MZE13, MM04, ORI^{+}10, RHW^{+}21, Rog09a, Rog09b, SMRM^{+}17, SAHP15, SDeaK^{+}09, TV13, VRS^{+}95, VRS^{+}99a, VRS^{+}99b, WZS^{+}18]. Packages [Ano97d, Bra97, CASA22, IKW23, BJM^{+}22, Bra04, FP94, SHW^{+}21, VD01]. Packaging [Kro99b]. Packard [Ano00f, Ano00f]. packet [MSM10]. packt [Gün02]. Padasip [CV22]. PaDEL [HLS^{+}13a, HLS^{+}13b, Yap11]. PaDEL-DDPredictor [HLS^{+}13a, HLS^{+}13b]. PaDEL-descriptor [Yap11]. Padua [CDSV10, CDSV11]. Padua2DM [CDSV10, CDSV11]. page [Har05]. pages [Ano97a, Ano99a, Ano00a, Ano00b]. Pain [EKJ^{+}03]. Paint [Sta04a, Wol98]. Pajamas [Lew99a]. Palladium [And03]. Palm [RM99]. Pan [BV87]. Panda [TG15]. Pandemic [WCHRM21]. Panel [BGG^{+}00]. panoramic [Bur04b]. paper [Bur04a, GDK21, MDRN18]. paper-based [GDK21]. Papers [BK14, Cse99, Xia08, ACM92, IEE90, IEE92a, IEE92b, IEE93, Kri90, MCQF21, Ros00]. paquetes [VD01]. para [RÓ01, VD01]. Parable [All02b]. Parade [Fie90a]. paradigm [VGD^{+}97]. paradox [WG06]. parahybae [SMO^{+}13]. Parallel [Ano00j, ALA20, Coo95b, FP95, Gro01, IEE92a, TDBEE11, ALGE12, AMR18, BMT^{+}20, BZB17, CZS^{+}21, Coo95a, CDR^{+}15, CKB11, DO16, EM93, GGSRMPM20, GJMPAM^{+}14, HMX21a, HMX21b, KORP95, MGR16, Nor23, ORS^{+}14, QR92, SGM^{+}08, Zag14, Kro99a, Tan11a, TDBEE11]. parallelen [Lei93b]. parallelisation [BSW^{+}14]. Parallelism [BBM^{+}21]. parallelizability [AMC16]. Parallelization [AAB^{+}04, GLCMC17]. ParallelKnoppix [Cre07]. Parameters [RGCS14, SMO^{+}13]. Parametric [SKSM19]. Paranoid [Bau06b, Bau06a]. ParaSoft [Kro99a, Ano04b]. Pareto [LH14]. Parliament [Sta04a]. ParMiBench [ILG10]. Parser [DS88, DS99, DS00, DS02, Mey18, DS90]. Parsing [GWT^{+}01]. Part [Bri09a, Bri09b, PH82, Rob95b, Rob95c, ATM22, Ano95f, Ano95g, Ram94c, Ram94b, Rog09a, Rog09b, FRAK15, JRA^{+}18, RAW^{+}16, Ros01b, Ros01c]. parted [SuS01]. Partial [eLAA^{+}23]. Participation [KS11, KJRD16, TWS^{+}22, STB23, YLHW21]. Particle [CDR^{+}15, CKB11, DBP^{+}18, FTZ^{+}23, VB19, ZRZ^{+}21]. particle-in-cell [DBP^{+}18]. partitioning [GGSRMPM20, KSS^{+}23]. Partnerships [San08]. Pascal [KP81]. Passport [Sta04a]. Password [Ano01j, NMX19]. passwords [Ano03e]. past [WCS20]. PASTIS [MWG^{+}90, MWG^{+}91]. PASTIS-program [MWG^{+}90]. Patch [Luc99a]. patches [PKB17]. Patent [Est06, Gil05, MSZ02, ACM05]. Patenting [GM05]. Patents [DDJ99, McL05, MSZ02, SG92, Gil04, MMB^{+}90]. path [Com99]. Pathogen [WM05]. patient [CKS16, EKUR10]. patient-level [EKUR10]. pattern [CLS95, OK94]. Patterns [GS00, MLA^{+}19, NGJ03, SDD05, TPSZ19]. Paul [Ano00e, Ano00k, Oms03, SD16, Teo13, Oms03]. Paulson [Ano04c]. PAW [FP94]. Pay [Bro96]. PB [FBY^{+}17]. PB-AM [FBY^{+}17]. PBBCache [GGSRMPM20]. PBE [Rus88]. PBX [Sch04]. PBX/VoIP [AML^{+}10]. PC [Str94, Ano86, Ano04b, Bee86, JJ91, MSS95]. PC-based [MSS95]. PC/MS [Uni85a]. PCI [dlCKK15]. PCIXCC [AV04]. PCMSolver [DSM^{+}19]. PCs [PKP02, SuS01, FP94, Rad89]. PD [HLS^{+}13a, HLS^{+}13b]. PD-PK-T [HLS^{+}13a, HLS^{+}13b]. PDEs [XXAD21]. PDF [DF00]. PDF-Dateien [DF00]. PDFBox [BGL^{+}20]. PDP [Dig80a, Dig80b]. PDP-11 [Dig80a, Dig80b]. Pearls [Ano00c, Zha16]. Pedagogy [CMJ^{+}04, DD08]. Peer [BCG^{+}14, CWB^{+}04, RCP^{+}12, RGCS14, Sta02c, Sin08]. Peering [BNST99]. PegaSoft [Ano00i]. Penetrating [WGG16]. Penetration [CB12, Ano11]. penguin [Bau06b, Bau06a, Sal08]. Penguins [Ana99]. Pennsylvania [ACM89]. Pens [DDJ99]. Penultimate [RAH^{+}01]. People [Jam09, SM89b, GTMR23]. Perceive [WCG22, CCK21]. perceptions [GW10, LLS11]. PERCIVAL [MMD^{+}22]. perfect [Cra89, Sch90a]. PerfectBACKUP [Ano00k]. Perform [LMZP19]. Performance [ACM00, Ahm08b, BOM97, Car04, CZ22, DKK22, Fyk97, HMKC12, KRB^{+}22, LA10, PRRL12, RAMB18, RLVdS21, Ten93, ZRNA20, ZXB^{+}23, dlCKK15, AJ05, BBE^{+}20, CH11, GLCMC17, GM02, HYA20, Lla06, MVS15, MM10, NMS14, PYM^{+}06, SGM^{+}08, SRGCPB^{+}09, SG06, TB05, YKK23]. Perhaps [JH16]. periodic [HAC^{+}23]. Peripheral [KJRD16]. Periphery [BKR^{+}20]. Perl [DF00, Ham99, Sha95, GWT^{+}01, LW03, MGFRG12, Ude97, Wal99]. Perl5 [Bea94]. permission [Oms03]. permittivity [YMCF23]. Persistence [Ano04b]. Personal [CSD^{+}05, DDJ98a, Kuk98, DDJ98b, And01, MP00, Mos12]. persons [Man92]. Perspective [Bes04, BMB^{+}18, Kuk98, Bab02, CH11, GLT08, LMWM18, Muw09, Tay19, YA11, ZW17]. Perspectives [FFHL07, MN04, Rie07, Waa09]. Pete [Gal10]. Peter [Ano00l]. Pfaffians [XOTI22]. PGI [Ano00i]. PGP [JKS02]. Phänomenologie [Rau04]. phantom [GV16]. phantoms [CG17]. Pharmaceutical [Boy07]. phase [BSC^{+}21, CAWK22, GCK^{+}17, GFZ16, ZK21]. phase-change [ZK21]. Phaser [MAF22]. PhasicFlow [Nor23]. Phenomenology [Rau04]. Philadelphia [BSW95]. philosophies [Cor05]. philosophy [Gan95].

Phonetics [Ano96b]. Photo [Jes03b, RHW^{+}21]. photo-excited [RHW^{+}21]. Photochemical [HW17a]. photoelectric [WSK^{+}22]. Photogenics [Ano00k]. photogrammetric [GBG^{+}16]. photographers [Goe07, Ham07]. Photonic [PGC21]. Photons [Bar00c]. photos [Pow14]. photovoltaic [AH19]. PHP [LW03, WN15]. PHP3 [Ano00d]. PHP4 [Per00]. Phynx [EKUR10]. Phys [AZ17a, CFCA13a, VRS^{+}99a, VRS^{+}99b]. Physical [CKB^{+}05, Gre14, PSSH16, ZKDP22, YKK23, dlVRB21]. Physics [BCB07, MMP^{+}22, VW92, BPG94, GDJG23, GFD^{+}24, PCAJ^{+}23, SHB^{+}20, ZRZ^{+}21]. Pick [Kuk98]. Picking [CPJ^{+}98]. Picks [PM00]. pigs [Lei04]. Pining [Coc01b]. Pinnacle [TBPS15]. Pioneers [DDJ98a, DDJ98b, SW13]. pipeline [Fel93, Mak03, SS23, UBR^{+}17]. pipelined [AHG94, Kri90]. piracy [EMD03]. Pitaevskii [YSVM^{+}16, YSMA^{+}17]. Pitfall [OK94, KOI94]. Pitfalls [Zha16]. PittPack [HXS20]. Pittsburgh [ACM89]. PizzaBox [Ano00k]. PK [HLS^{+}13a, HLS^{+}13b]. PKI [Ahm08a]. PL8 [GHL^{+}04]. Place [USE01b]. Placement [LLWM23, ZZZ22]. Plan [Sta04a, Ano95f, Ano95g, Rob95b, Rob95c, Sta12]. Planes [RCB^{+}14]. Planning [CPJ^{+}98, DWP^{+}14, GNR^{+}09, PFL^{+}12, TBPS15, Tra95]. Plans [Bes04, Coc03, PBJ^{+}12, SZAB98]. plant [MLMFN^{+}15]. plasma [DIK^{+}23, DBP^{+}18]. plate [IHBS14]. Platform [Gui00, KRB^{+}22, Li18, NRG^{+}99, PSP^{+}22, WCHRM21, Wol03a, ABF^{+}14, AMR18, BMF^{+}19, BJM^{+}22, BS05, CRB^{+}18, Don04, GHM^{+}05, HMP^{+}15, MLZ^{+}23, MDRN18, Pen03, PPC^{+}15, RTH15, RA16, SA15, WHJ15, Wes03, Wol02, YL08, vdHGG^{+}13]. Platforms [Ano06, DKK22, HW17a, VOM12, MV05, NMX19, SS05a]. Play [Ste01]. Player [Ken02]. plays [BSP11]. PLEIADES [HMP^{+}15]. Plots [BFC02, Ste93]. Plotting [WKC^{+}90, Rac06]. Plug [DF00]. Plug-ins [DF00]. plugin [MM04, PNK^{+}23]. PlugSys [Ano00k]. pluralistic [O’S03]. Plus [Ano96b, Ano04d, Ano96c, Ano96e]. Pocket [Cam99, Neu00, Uni77]. POINCARÉ [MZE13]. Point [FL16, HR11, FHL^{+}07, He95, Ho95, Sta09]. Point/Counterpoint [HR11]. pointer [Eig03, PKH07]. points [CDSV10, CDSV11, ODP15]. Poised [GAS^{+}01]. Poisson [FBY^{+}17, HXS20]. Policies [BKHT21, CF07a, GGSRMPM20]. Policy [Cho09]. Polish [Vir05]. Political [May06, Rau04]. PolyAWK [Cor87]. Polychromy [SCDS15]. polymorphic [LN92]. polymorphism [BR95]. POLYS [EHP14]. polysaccharides [EHP14]. polyurethane [KDM17]. Ponders [Coc03]. Pooled [WM05]. Pooling [SC02, VJ23]. Poor [EKJ^{+}03, Jon02]. Pops [Cha98]. Popular [Ano04b, Col05, Fie90a]. popularity [CFMRL11]. population [GLMS18]. porous [HSF^{+}15, KGW^{+}21, SS23]. port [Bea94]. Portability [BOM97, Keh94, ATHW92, BBE^{+}20, Fri97]. Portable [Fri97, BB91, Ert94, Rac06, Rap94, VRS^{+}95, VRS^{+}99a, VRS^{+}99b]. portal [YLG05, YWA07]. Portfolio [BdP13]. Porting [Gil88, Hub03, Jae03, JJ91, MB98, Pen03, SSC^{+}00, Shi03, Sta88e, Sta99, Sta00b, Sta00c, VGdlP01, ACW04, Ano00f, Sta92b]. Portland [Abr81]. Portterung [Ano00f]. posed [MD17]. Posit [MMD^{+}22]. Position [The04]. positron [HMYH22, HFO^{+}12]. POSIX [Ano00l, Coc03, GB20, dlPRGB99]. Possibilities [Bes03]. Poster [GV16]. Poster-05 [GV16]. PostgreSQL [GGK99, Per00]. postings [Sta96b]. Potato [Ano01b]. Pots [TGW^{+}22]. pour [Cor00, Rod00]. Power [Bao93, DDJ99, G^{+}02, Tan11a]. Powerful [CSP^{+}03, HMO^{+}18, WM01]. Powerpack [Gan04]. PowerPC [Ano01j, But94]. PowerRAC [Ano00i]. PowerWindows [Ano00k]. pp [Aji17, Cas02]. pp. [Ano04c]. ppohDEM [NMS14]. PPoPP [FP95]. PR [LN92]. Practical [EM93, Gla03b, Gla04, She07, VOK^{+}22, Mei92, Men12, PSSH16, RAMB18]. Practice [BCB^{+}17, DGBH93, FP95, KP99, Par03, EKUR10, Fin80a, Fin80b, MCQF21]. Practices [BKHT21, BB02, LZWH22, Sca04, SF15, TZ22, WCG22, CFL23, CJ17, GGH05, LRP21, Sai02, ZFD21, vGPB10]. Practitioner [LLS11, Cal10]. pragmatic [Sam06]. Praise [CGK^{+}02, CSP^{+}03]. Praxis [Cla90]. Pre [Boy13, Ano01d]. Pre-Internet [Boy13]. preCICE [ZAC^{+}23]. precise [Sta96b]. Precision [FL16, GSR^{+}04, Fär05, FHL^{+}07, KGT22, MRH23]. precompiler [SC08]. Predict [UMV15]. Predicting [ACB18, Ano21, BSA14, XFS^{+}22, RHW^{+}21, VGSN18, Yu06]. prediction [AFZ17, AFZ18, ATCZ19, FLA^{+}16, GFS05, HLS^{+}13a, HLS^{+}13b, KRR23, LZ11a, LZ11b]. preeminent [Phi12]. Preface [Boy00, RB05b]. prefix [PPR19]. pregnancies [PH16]. preliminary [WHJ15]. Prentice [Fox08]. Preparation [Mag04]. Preprocess [MGM^{+}02]. Preprocessor [HZS^{+}16, Iwa02]. Preprocessor-based [HZS^{+}16]. present [WCS20]. presented [ACM92, Cse99]. Press [Ano15a]. Presses [Ray98]. pressure [Mak04]. pressures [MLMFN^{+}15]. Prevailed [Hoh01]. prevails [Höp04]. Prevalence [WM05, AMC16]. previewer [KK94]. Prices [Pra03]. Primary [LBF^{+}22]. Prime [McC99c, Fär05]. Primer [Kenxx, RE04, Aji17, Buc82a, Fri16]. Princeton [Ano01i]. Principles [CRW^{+}04, FP95, MBTB21]. Print [Ano96c]. printer [MGYC18]. printing [CKS16]. Priorities [SPDQ22, MSM^{+}03]. Prioritization [LMZP19]. prison [Sta96b]. Privacy [Coc01b, CSD^{+}05, HKP02, Mag01c]. Private [CK10, Joh02, Ano99c]. Prize [Bar01]. Pro [FT09, Vir05]. proactive [RCO20]. probe [YMCF23]. probits [Cre07]. Problem [jFFR16, Kam11, MCS12, ST10, Tie93, DC23, FK99, KK17, Tie88]. Problem-formulation [jFFR16]. problem-solving [jFFR16, KK17]. Problems [Gro01, Mal02, MTM^{+}19, Sta12, XXAD21, Hay05, HMX21a, HMX21b, KGW^{+}21, Man92, Phi12, RSZ96, TACA15, WFV14]. Problemsets [AFS81]. procedural [Hub04b]. Proceedings [ACM88, ACM92, ACM95, ACM97, AY93, Abr81, Ano87, Ano88b, Ano88c, Ano89, Ano90b, Ano90c, Ano92, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Bao93, BGG^{+}94, DGBH93, FMA02, HDR03, HDR04, IEE89, IEE92c, IEE92d, IEE94c, IEE94b, IEE95a, IEE95b, Lev95a, MS91, MSNS91, Mio90, Ten93, USE90, USE94, Ass95, USE98a, USE98b, USE00a, USE00b, USE01a, USE01b, USE02a, USE02b, USE02c, VW92, Vor92, ACM89, ACM93a, AT92, BPG94, Bon93, BSK87, Bun94, EHP94, EKR91, Fur90, IEE94a, Kap92, Lev95b, MG94, PT91, QR92, SS93, SC00, SM89b, USE88, Yuk94, Ano06, HY14, IEE05, Jef08, ACM93b, FvdHJ10]. Process [BCB^{+}17, GS00, Kro99b, MZG14, TV99, ASAB02, FM10, JCNS^{+}22, KKT17, KH05, KFYI13, LRD^{+}19, ZWH21]. Process-Oriented [Kro99b]. Processes [SFF^{+}06, MAF22, NT06, PYM^{+}06, SSS22, YLXZ16, YZC22, ZVvDD11]. Processing [AWD^{+}18, FY18, Gre80, MK12, TGC^{+}21, AH19, HM89, HFO^{+}12, HBB^{+}12, How98, Kit94, MM04, Phi93, SNC^{+}06, YA05]. Processor [Ano00i, CCA^{+}19, Knu99a, AHG94, CCA^{+}13, Cra89, GYW^{+}23, JV01, KORP95]. processor/accelerator [CCA^{+}13]. Processors [IEE94c, Ano00h, Cha92, Kri90]. Procurement [RLTD23]. Producing [Fog06]. Product [JWC18, Kro99a, RE04, MAF22, WMLM22, vGPB10]. Production [CK10, GF11, ZVvDD11]. Productivity [Ano02b, CRW^{+}04, GS12, Kuk98, SMS16]. Products [Ano95a, Ano96b, Ano96c, Ano97c, Ano97d, Ano98, Ano00i, Ano00j, Ano00k, Ano01i, Ano01j, Ano02b, Ano04b, Bra97, Kro00, Kuk98, CWZ06, HBR19, Hic04, KT05, KL07, PSE04, Ano01i, Ano04c]. Profession [San03]. Professional [Ano00i, Mag04, PKP05, Spi06, Ste93, SuS01, Pec08]. Professional/Administrator [Mag04]. Professionalism [CSP^{+}03]. professionelle [PKP05]. professionellen [DF00]. professionelles [SuS01]. profiler [GKM82, GKM04]. profiling [SCH^{+}91b]. Profit [Pri19, Ous99, Ros14]. Profitable [Wal01]. Program [ATM22, BlG12, BMZ14, Boy13, SM00a, SM00b, SM02, TBPS15, WKC^{+}90, XWZ^{+}23, AZ17a, AZ17b, Bra04, Car89, Dig75b, Dan11, Dre94, FK99, GIM07, Gar09, GFD^{+}24, HMR93, HLL^{+}95, Kle21, KFYI13, MWG^{+}90, MWG^{+}91, PH82, SM89a, SMS04, Tai13, Wat94, ZJS^{+}20, Leó98]. Programmable [OCH90a, OCH90b]. Programmbeispiele [PKP02]. Programme [Str94, SuS01]. programmed [Sch91a]. Programmer [GF99, GS00, Joh92, RW87, Wat85a, Wat87, Wei03, Dig75b, Gla03b, Gla04, Wat85b]. Programmers [Coc01a, Kro00, Wel94b, Ude89]. Programmes [Lei93b]. Programmieren [Ste00a, Jor04]. Programmierung [Jor04, PKP02, PKP05]. Programming [ACM92, AKW88, Ano00c, Ano00d, Ano00k, Bar00a, BM06, BSS84, Cha97, Chaxx, Cha01c, Cha04, CWM^{+}20, CGB^{+}05, Cro00, FP95, Ham99, Jen97, Jor04, KP84, KP99, LO97, Mar03, Pin02, PKP05, RM99, Rob96, Rob97, San78a, SBA92, Ste95, Ste99, Ste00b, Ste01, Tro96a, Vor92, Wal99, Wri00, dlPRGB99, Ano04d, Cio01, CC05, Edd96, Fra95, GM84, GSW08, HH88, HL02, Jon05a, KORP95, KC92, Lin00, Liu08, Man92, MWB89, MAMC05, MT94, NGJ03, PBOP07, WACBL03, ZK05, Ano97b, Ano97a]. Programmming [Dig80b]. Programmpakete [Bra04]. PROGRAMS [BY91, AFS81, AFS82, AG95, CR92b, dM99, MRGP20, BY92, CK06b, CK06c, CZ99, EM93, Fie90b, KOI94, KW94, OK94, SC08, YSVM^{+}16, YSMA^{+}17]. Progress [Ano00j, Ano04c, EKJ^{+}03, Sch91a]. PROGRESS-editor [Sch91a].

Project [All02a, All02b, Ano95b, Bar01, CZ22, DGC^{+}07, Fie89, GS12, GGL21, Hae02, KGM^{+}16, Kro00, Nej12, NN20, Noj01, Owe01, RT12, ACB18, AAB^{+}05a, BGL^{+}20, Car04, CGS94, CWZ06, DTB05, Eub05, Fie99, Fog06, jFFR16, GL14, Gau03, Ger03, Jør01, KS02, MSR10, NN21, NDDH^{+}21, Pel89, Sin10a, SG06, You08, Coc01a, DDJ99, Kan12, Liu06, Maz15]. Projects [Bak20, BMB^{+}18, BGL^{+}21, CFM08, CCSW10, Eri00, FGBM14, GMBv20, HBGS19, JEB^{+}23, KCAS23, KKN^{+}21, KT04, KS11, KJRD16, LPFD21, LZWH22, LMZP19, Mec05, MFS15, OT91, PLZ^{+}22, PLO^{+}23, Rie11, RGCS14, SV19, Sta02c, STG19, TTB09, WFW^{+}20, WWSG21, ZXB^{+}23, ABC18, Ano21, ASAB02, Asu05, APHV19, CIC13, Cap13, CFMRL11, CJ17, CFW01, CSP09, Fei23, FM10, HPM^{+}08, HNH03, HSX^{+}18, KC22, KF17, Koc07, Koc09, MG12, Mah03, MLA^{+}19, MWG08, MRS07, MSR09, NT06, PSSH16, PKB17, PBOP07, PSDG18, PSL21, PdSCJM22, QB21, SBS20, San01, SMS16, SSA08, Spi21, SDD05, STB23, THG20, THG23, VGSN18, VGP^{+}19, WMLM22, WMK^{+}17, XFS^{+}22, YLHW21, ZWH21, ZWU22, vGPB10]. Prolog [DC00]. Promises [Her20, MSR09]. Promote [SV19]. Promotes [Bar00b]. Promoting [LH03, CH06a]. prone [Ban16, Ban17, KL07, SK12]. pronunciation [HK95]. pronunciation-information [HK95]. Proof [BMZ02]. proofreading [Mił10]. PROOFS [BY91, BY92]. propagation [HKvH16, WGG16]. Properties [Bar00a, MSB09, SPAW17, Sin10a]. Property [Lin08, Mar05, Vai01]. Proposals [Coc01b, SSS22]. Proposed [Bar01]. Proprietary [Egy01, SCSC04, BH11, Bou05, DD17, KHMA12, Lam09, PSDG18, Sin10b, Wes03]. propriété [Cor00]. pros [Ano04d]. Prose [Gre80]. prospects [KKA^{+}19]. protect [Mah03]. protected [BG95, Dre94]. Protecting [III01]. Protection [Est06, Wag03, Dre94, NO03, Zic01]. protein [HMO^{+}18, LSJ^{+}06, ZJS^{+}20]. proteins [EHHH06]. Protention [Fyk97]. ProteoAnnotator [GKP^{+}14]. proteogenomics [GKP^{+}14]. Protest [Col09a]. protezione [Zic01]. Protocol [CWB^{+}04, Kro00, DDHS03]. protocols [CLL05]. Proton [DWP^{+}14]. Prototype [LO89, RSKF96, Mor91, Yan90]. Prototyping [Sha95, TGC^{+}20, BBNP93, GGSRMPM20, SSS^{+}14]. ProtoVIew [Kro99b, Kro99a]. protrusions [BDAW15]. provenance [RDZ20]. prover [LN92]. Providers [GB00]. Providing [BS05]. Proving [MD22]. Provision [Joh02]. Proxy [Ano96c]. Prüfungsvorbereitung [Mag04]. pryzm [Ano97c]. PSCs [DDJ98a, DDJ98b]. pseudo [WN15]. pseudospectral [HKvH16]. PSI [GKP^{+}14]. PSI3 [CSV^{+}07]. psychopath [NN20, NN21]. psychosocial [Ano01g]. PTCLab [GFZ16]. Pteros [Yes12]. PTR [Fox08]. Public [Bar00c, Fav23, Fre23, Geh96, Hol23, Hüp01, Joh02, LRBM23, MG05, Neg15, NO03, O’S03, Rav00, RLTD23, ST10, Väl04, VJ23, Ano01a, Bon02, RDZ20]. Public-Domain-Software [NO03]. Publications [PMBM^{+}15, TG15]. publique [CF98]. Publish [KNS18]. publisher [Oms03]. Publishers [Aji17, Neu84]. Publishing [Ano00i, Ano00j, Ano00k, Fur90, DPL^{+}91, PS^{+}09]. PUFoam [KDM17]. Pull [APCs22, KCAS23, Ano21, AKF21]. Pune [KJ03]. purpose [DBP^{+}18, LPC^{+}15, TV13]. Put [CWB^{+}04, EKJ^{+}03, OG07, Gal04]. Putting [Jør01]. PyFR [WFV14]. PYMES [RÓ01]. PyMTL3 [JPOB20]. PyOECP [YMCF23]. pyrolysis [XAPK14]. PySCF [HAC^{+}23]. Python [DF00, RAH^{+}01, CV22, CX23, CFW17, How98, JPOB20, JNN12, KPK^{+}17, KSH14, LFN^{+}11, MMP^{+}22, NZPWR22, SSS22, Tro96d, Tro96a, Tro96c, Tro97, You08]. PyVCI [SC16].

QBMM [LMHL20]. QccPack [Fow00]. QCPE [Boy00]. QEHeat [MBTB21]. Qemu [CK06a, CK06b, CK06c, CK06d, CK06e, CK06f, CK06g, CK06h, MZG14, CK06a, CK06b, CK06c, CK06d, CK06e, CK06f, CK06g, CK06h]. QI [BJJ14]. QIST [HWM^{+}15]. QmeQ [KPK^{+}17]. QoS [ZZZ22]. QoS-Aware [ZZZ22]. QT [Ste01]. Quadrature [Joh18]. Qualitative [BKR^{+}20, MOT^{+}18, SCR05]. Quality [Abe07, ASS^{+}23, BVLF14, CFM08, CKB^{+}05, DM15a, GS12, Kam14a, Kam14b, KY16, KMF^{+}07, PMBM^{+}15, Sta02c, YT22, ZE03, Ano02a, CFMRL11, DM15b, GSW08, PYM^{+}06, RB05a, Smy97, SAOB02, THG23, WMLM22, ZE00]. quantification [BDAW15, FL15, LSJ^{+}06]. Quantify [HW17a]. Quantifying [DKMB14, LSM09]. quantitation [PSS^{+}07]. Quantitative [BJJ14, Sha10, WMLM22, MOT^{+}18, SDeaK^{+}09]. quantization [Fow00]. Quantum [BCB07, Par03, SW13, Aki16, ABNÅ05, HHG^{+}21, HWM^{+}15, Hua17, Hua23, JWC18, JNN12, KPK^{+}17, OMA^{+}22, PM21, Sai13, WPAV14, Boy13]. quarter [Sal94]. quasi [O’S03, VB19]. quasi-legal [O’S03]. quasistatic [WHJ15]. QUATTRO [BJJ14]. QuBiLS [GJMPAM^{+}14]. quelen [SBM^{+}10]. Quelltext [DF00]. quels [Cor00]. query [Phi93]. Questioning [Mog03a]. Questions [And03, CPJ^{+}98, Man03, PKGA22, Sie99, Spi03, Stö04, Val91]. Queue [Kre00]. Queueing [Kro99a, LL14, Mar22]. Quick [Ano00i, Kri03, Ste93, Fie90a, Phi12]. QuickStart [Ano97c]. Quincy [Ste00b]. Quire [MMD^{+}22]. quite [Hac98]. Quiz [AFS81, GM02]. Quiz/Exam [GM02]. quo [WBGM02]. QuTiP [JNN12].

R [Aji17, Ano00d, Neu84, Fri16, HBZ09, MCS12, SA15]. R-based [SA15]. r0a [Ano05b]. r10 [AFZ17]. r11 [AFZ18]. r12 [ATCZ19]. r7 [LZ11b]. r9 [FLA^{+}16]. RackMount [Ano00j]. RackMount-1UAXe [Ano00j]. Radar [WGG16]. Radiation [DKMB14, PMG^{+}09, PKG^{+}10, PPG^{+}11, PBJ^{+}12]. Radio [FSB^{+}01, Sea02, Blo04, SSS^{+}14, ZPH^{+}15, KSV16, Ron15]. Radiotherapy [GNR^{+}09, WKA^{+}08]. Raging [GGT05]. Rahmenbedingungen [WP04]. Rainbow [TGW^{+}22]. rainy [McC02b]. Ralph [Ano00c]. ran [Cre07]. random [Sib17, WN15]. RandSpg [AZ17a, AZ17b]. range [Phi12]. ranging [GVOM09]. ranked [MG12]. Ranking [FRBRF19]. rant [Pet06]. Raphael [Neu84]. Rapid [But94, HOL^{+}07, Lov06, WDK^{+}20, GGSRMPM20, SSS^{+}14, Yan90]. rarefied [CAWK22]. rarefiedMultiphaseFoam [CAWK22]. Raritan [Kuk98, Kuk98]. RAS [Ave06]. Rate [LL14, WM19]. Rate-Based [LL14]. Raum [Sie04]. Rave [Ano00j]. Ravenscar [VGdlP01]. ray [CG17, WFDK19]. Raytheon [Ano01j]. re [SIK^{+}13, Sin08]. re-opened [SIK^{+}13]. re-use [Sin08]. reacting [MLZ^{+}23, TPK^{+}21b]. reaction [SS23]. Reactors [Ano05a]. Read [Sta97b]. Reading [CSD^{+}05, PSP^{+}22]. Ready [EKJ^{+}03, McC99c, Ano00h]. Real [Ano01i, Cur99, Reh01b, RAH^{+}01, SSC^{+}00, ZXB^{+}23, dlPRGB99, GPPT16, GTMR23, HAC^{+}23, MD17, Rui13, SG99, Sta80a, Sta81d, Sta81c, Sta81b, TL17, VGD^{+}97]. Real-Life [ZXB^{+}23]. Real-Time [Ano01i, Reh01b, dlPRGB99, Cur99, GPPT16, GTMR23, HAC^{+}23, Rui13, Sta80a, Sta81d, Sta81c, Sta81b, TL17, VGD^{+}97]. réalisation [Rod00]. Realisering [Esp96]. Realisierung [Per02]. Reality [Ano96c, Coc01a, CGK^{+}02, CKB^{+}05, Sta02a]. Really [CPG^{+}04, Guy00]. RealPaver [GB06]. realtime [MDRN18]. reaping [WG05]. Reasoning [Vor92]. Rebel [Moo01a, Moo01b]. rebooting [GJLT11]. Rebutting [The04]. recenti [Mol01]. Recht [Hüp01, WP04]. Rechtliche [Jak03, WP04, Feixx, SG05]. Rechtsfragen [Man03, Spi03]. Rechtsfreier [Sie04]. Rechtssicherheit [Sie04]. recipes [Phi12]. Recipients [Coc01a]. Recognition [GAS^{+}01]. recognizer [Mak03]. Recognizing [BY14]. recombinant [Ewe18]. recommender [SBS20]. Recommending [KC21]. Recompilation [SM00a, SM00b, SM02, SM89a, SMS04]. Reconfigurable [dlCKK15, Don04]. Reconnaissance [TG15]. Reconsidered [SHS^{+}93]. Reconsidering [Nag18]. Reconstruction [SAC^{+}15, SCDS15, WFV14]. reconstructions [IHBS14]. Record [BHP^{+}01, EJS^{+}01, Sta01b]. Recorded [Les03]. Records [Eri01]. Recovery [LOW91, MMR95]. Recurrence [BFC02]. recursive [LN92]. Red [Ano96b, Ano00c, Ano00b, Ano00j, GB00, Jes03c, Men12, Ano95a, Ano03d, Ano08a, G^{+}06, Gra99, Kro99b, Kro99c, McC99b, McC02a, McC03, McC04, Reh01a]. Red-Eye [Jes03c]. Red-Hot [GB00]. REDBACK [IHBS14]. redesign [HMO^{+}18]. Reduce [Lev23, Hea09]. reduced [JD19]. reduced-order [JD19]. reduction [CFCA13a, CFCA13b, IHBS14, MZE13]. redundancy [Due97, Joh94b]. redundant [Gar09]. reengineering [Bud10, Bud10]. Refactoring [Bow05, SK12, VGP^{+}19]. Refactorings [PLZ^{+}22]. Reference [Cam99, Chi93, GDT^{+}02, GDT^{+}05, The03, Kro00, LLSt99, LLSt00, LSM^{+}99, LSM^{+}00, LSM^{+}01, Neu00, SSC93, Buc82b, Dig75b, EaoGOBHW14, G^{+}01, Gal09, Gar00, Gri02, Har00, LLG90, LLG93, LLG94, LMOS93, LSMO96, Loo15, She87, Sta80a, Sta81d, Sta81c, Sta81b, Sta03b, SW15, Uni77, Uni85d, Uni85f]. References [Ano95a, KHA^{+}03]. reflection [HMYH22]. Reflections [Cus04, Luc99b, CK08, Zei03]. regard [Stö04]. Region [Bao93, GAS^{+}01]. Register [Pun04, YLL^{+}07, FG92, HC07, Mak04, Mat03]. Registration [GKL^{+}14]. Registry [DSB^{+}16]. Regulation [Thi22]. Regulatory [MPG^{+}16]. Rehearsal [Ahm08a]. Reimplementing [VGD^{+}97]. Reinvent [CRW^{+}04]. Reject [PLZ^{+}22]. related [CCK21, ZE00]. relation [DRM21]. Relationship [CFM08, BD03b]. Relationships [GS12]. relative [BH11, THG23]. relay [KSV16]. Release [Ano96c, Bar01, CNSR23, Ano01a, Ano01b, Pyr84, YLXZ16]. Released [Bar00c]. Releases [Ano01i, Ano01j, Ano04b, Bar00c, Den99, IKW23, MFS15]. Relevance [MZH22]. Relevant [Wol04, Wol03b]. Reliability [SSP17, SSP18, UMV15, WM19, Wan21, YLXZ16, YZC22, ZD05]. Reliable [Cse99, Maz15, Zim10, Bou05]. Remains [Gla03a]. Rematerialization [Pun04]. remedies [NYB10]. Remote [Ano01j, Joh99, Eds16, Ano96b]. Removing [Jes03c]. Rendering [MK12]. rendszerek [Lás05]. renormalization [WPAV14]. Renovating [MQN19]. reopened [TZH22]. repeat [DRM21]. replicated [Fra95]. replication [ACB18, CJ17, DTB05]. Report [IHSR19, KJ03, Sto99, WW01, Ano95d, FFHL05, GS02, MSLH71, MSM^{+}03, MOT^{+}18, RVLS14]. reporting [MSM10]. reports [JCNS^{+}22, Raj13]. repositories [AKMS23, JDB09, KGMI06, TKSC20, XTY^{+}22]. Repository [Bar00b, CFL23, CdR99, Mas05, MWG08, ZVvDD11]. representation [CYOS19, Mer03]. reproduce [AVA^{+}16]. Reproducibility [Bar22]. Reproducible [CW15a, CW15b, Sto09, LGA20]. Republic [BSK87]. Republish [CSD^{+}05]. Reputation [CK10, PC13b]. Request [APCs22]. Requests [KCAS23, Ano21, AKF21, HZ14]. required [SNC^{+}06]. Requirements [ACC^{+}12, BNSW15, HBR19, HZ14]. Research [AAB^{+}05a, AWD^{+}18, Ano00j, BMF^{+}16, Bar22, BM12, FVD^{+}12, Gra01, HCH^{+}20, KHA^{+}03, Mog03c, San08, SB08, SZAB98, SCB04, Sto09, Waa09, WKB14, WKA^{+}08, AM18, AKF21, BMF^{+}19, BPG94, BYV08, Bon02, EKUR10, FMT^{+}08, GRJS01, Har05, Joy09a, LH03, MNS19, SSS^{+}14, Kro00, LH03]. researcher [Fei23]. Researchers [Rob20, CFL23]. Residual [UMV15, MGYC18]. Resilient [ASWD18, CCA^{+}19]. Resolution [ACC^{+}12, Ano16, Wat01, DRM21, ZRZ^{+}21]. Resolve [Wat01]. Resolving [CGB^{+}05]. Resonance [BMR^{+}23]. resonant [WFDK19]. Resource [BE06, GB00, VJ23, BDP^{+}14, MS02, PPC^{+}15, Weh03]. Resource-Centric [BE06]. Resources [Ano95a, Ano96e, Cas19, Zha16, BJWZ08, BB08, FHH11, KTTK17, Qui00]. Respect [AS03, NN20, HKP02]. respecting [YL08]. Respond [Gau07]. responsabilités [Séd02]. Response [BNST99, BJJ14, PFL^{+}12, Sta03a, BM22, EMD03]. responses [MLMFN^{+}15]. responsibilities [Séd02]. Ressource [Ano01b, Ano01d, Ano01c]. REST [HHG^{+}21]. Restore [Ano00i]. result [Ano99c]. results [MHP94, PH16]. Retargetable [GJS^{+}02]. Retargeting [Col02, LC12a]. retention [RCO20]. rethinking [GJLT11]. Retrieval [Has05, AJLM18, PPR19]. retrospective [BTL^{+}11, Bor88]. Return [Pea16]. retuschieren [DF00]. Reusability [PAB^{+}17, Tai13]. Reusable [RW87, WDK^{+}20]. Reuse [MD04, SH11, vGPB10]. Reusing [BB02]. Revealed [HW17b]. Reversible [SFWD12]. Review [Aji17, Ang01, Ano97b, Ano97a, Ano99a, Ano00a, Ano00b, Ano11, Ano15a, Bar16a, BCB^{+}17, Bra92, Cas02, Cha13, Chi97, Cho09, Cro00, Fox08, Gil06, Gra99, Jen97, Ken02, Kuc06, Men12, Neu84, RCP^{+}12, RGCS14, SD16, Teo13, TG15, Waa09, Weh03, Edd96, EKUR10, Ens05, Fie90a, FRBRF19, Joh92, Pri19, Rac06, RH21, Rus09, Sur04, YT22]. reviewing [Añe11]. Reviews [Ano00c, Ano00e, Ano00d, Hay05, LB00, Lip07, Mar05, Sta02c, STB23]. Revises [Ano01j]. revision [MRS07, Sib17]. Revisited [AM04]. Revisiting [SGD05a, SGD05b, TZH22]. Revitalizing [MTD^{+}09]. Reviving [Hoh01]. RevKit [SFWD12]. Revolution [DOS99, GAS^{+}01, Moo01a, Sea99, AD04, Gom99, Moo01b, Moo03, Pri19, Ano99a]. Revolutionary [Ray99b, Ray99c, Ray01b]. rewards [Eds16, WG05]. rewriting [Sch91a]. RF [ASAAM^{+}19]. RF-based [ASAAM^{+}19]. RFC [Sta78a]. RFSFNS [VRS^{+}95, VRS^{+}99a, VRS^{+}99b]. Rhamdia [SBM^{+}10]. RHEED [HMYH22, Dan11, HMYH22]. Rhythm [Row02]. Richard [Cas02, Neu84, Fio03, Gay02, Wil02]. Richards [ORS^{+}14]. Richardson [Ano00e]. richtig [DF00]. Ride [Mau05, Sta04a]. Riding [SSC^{+}00]. riferimento [Mol01]. Riffraff [EKJ^{+}03]. Right [EKJ^{+}03, Sta97b, Sur01a]. Rights [Fal03, GB00, DPL^{+}91, Gom99, Mol01]. Ringelmann [SMS16]. Ripple [LO89]. RIPS [MWB89]. Risc [Ano00f, Ano00f, DXT^{+}18, GYW^{+}23, MMD^{+}22, PGW^{+}20]. RISC-V [DXT^{+}18, GYW^{+}23, MMD^{+}22, PGW^{+}20]. rise [Vai01]. Risiken [SG05]. Risikoanalyse [Feixx]. Risk [Ano18, GB00, LMPT22, MSC19, Sha10, Feixx, MD17]. Risks [Lev23, SG05, Neu99, SH11, Spe01, WG05]. Risky [Ano19]. rNMR [LSM09]. Robert [Kuc06, SD16]. robocode [Liu08]. Robot [SH19, Mau05]. Robotic [Bar00a, Coc03]. Robotics [Fly87b, IEE89]. Robots [Coc01a, Fly87a, Fly87b]. Robust [CX23, Neu99, Kir12]. Rogel [Bar16a]. Rogel-Salazar [Bar16a]. Rogue [Ano00i]. Role [Bar22, BNSW15, Men10, ACM93b, BCPS10, Far23, O’S03, SG06]. roles [BSP11, IC22]. ROM [Ano00a, Ano00b, Rod00]. ROMs [Ano01d]. room [Hol05]. Root [BMZ02]. roots [VRS^{+}95, VRS^{+}99a, VRS^{+}99b]. ROS [GTMR23]. ROSE [MWG08]. ROSS [McA85]. roster [Sur04]. rounding [FHL^{+}07]. route [SDL^{+}16]. router [CLL05]. routine [MN21]. Routines [GBDM77, SBD^{+}76]. Routing [LGW18, Sor06]. Row [Neu84]. Rowman [Aji17]. Rpath [LGA20]. RPC [Tie88, Tie93]. RS [SCH^{+}91b]. RS/6000 [SCH^{+}91b]. RSA [Bar00c]. RSOME [CX23]. RT [PBJ^{+}12, PH82]. RT-11 [PH82]. RTF [Kro00]. RTL [BGG^{+}15]. rule [AG22, Cla90, Mił10]. rule-based [AG22, Cla90, Mił10]. rules [Sta96b]. Run [Gag02, Fog06, Jon01]. Running [FSB^{+}01, McA85, SNF04]. Runtime [GB94, MB98, Mir03, MD22, Shi03, SPDQ22]. Rush [GB00]. Rydberg [SPAW17].

S [HMX21a, HMX21b, Neu84, Pen03]. S-FEM [HMX21a, HMX21b]. S/390 [Pen03]. SA [Rob20]. Safe [Maz15]. Safely [von88]. safety [EKUR10, Gar09, GEI^{+}11]. safety-critical [GEI^{+}11]. SafeWrite [Ano00j].

SAGE [Den13, Joy09a]. sagemath.org [Den13]. SAGU [Bro01]. SAINT [VSN22]. Sair [Mag04, DRP01, Jan01, Mag00, Mag01a, Mag01b, Mag01c, Mag04]. Salazar [Bar16a]. salesman [DC23]. Sam [Ano99a, Cas02]. SAMBA [Ano97c, MS02]. Samples [WM05]. Sampling [AL07]. Sams [P^{+}99]. samspil [MG05]. San [ACM92, IEE92c, IEE94b, Lei93a, USE94, USE02a, Kro99a]. sans [STS92]. Santa [Ano94b]. Saratoga [Kap92]. Sarge [Ano05b]. SASTBX [LHZ12]. SAT [McH92]. satellite [ODP15]. Sather [SO91]. satisfaction [GB06]. Satisfying [FvH03]. Savvy [Kro99b]. SaX [MSB09]. SC2000 [ACM00]. SCAI [AK95]. scalability [BZB17]. Scalable [XXCL19, AAA^{+}14, AAB^{+}05b, VBG^{+}10, Yad07]. scalar [SAHP15]. Scale [KGM^{+}16, KS11, PSP^{+}22, ZXB^{+}23, BOL14, BLG^{+}17, Bor09, BSC^{+}21, CJ19, CYL^{+}23, HXS20, Koc07, KT05, KL07, LSJ^{+}06, MNS19, ORS^{+}14, RB05a, RAMB18, RDZ20, SMS16, VBG^{+}10, VGP^{+}19, WFF18, XTY^{+}22]. Scaleable [PBH01]. Scaling [TZ22, KTTK17, QSX^{+}15]. SCAN [Cse99]. SCAN-98 [Cse99]. Scandinavian [AK95]. Scanner [Nic93, Pax95]. scanners [ACKT20]. Scattering [LHZ12, SDL^{+}16, TACA15, WFDK19]. SCC [SSC93]. Scenarios [WCHRM21, LLEL^{+}23]. scheduler [Mak03, UZ97]. Scheduling [Hag04, Kri90, LLEL^{+}23]. Scheme [DF00, MRN20, AFS81, AFS82, Bir93]. SCHEME-Manual [AFS81]. schemes [NMX19]. Schenker [Sha05]. Schlüsse [O’Rxx]. Schnittstelle [Per02]. scholarly [HK09]. Schools [PM00]. Schritt [Str94]. Schutz [NO03]. schweizerische [Sur01b]. Science [Adl00, ABC^{+}14, Ano01a, Bar01, CC03, MCGA22, DDA^{+}07, DD08, EKR91, FMFZ19, HETD09, Ing92, MAMC05]. Sciences [Hig93, MS91, MSNS91]. Scientific [ABC^{+}14, AAB^{+}04, Bad07, Bee01, Cse99, GDT^{+}02, GDT^{+}05, GvdHPR14, HW17b, Par03, Sto09, VOM12, Dan11, G^{+}01, Gal09, HR94, MM10, Pea16, Wol02, vdHGG^{+}13, Ano96e]. scientifiques [Rod00]. Scientists [Coc01a]. SCO [DF00, Mog03a, Mog03c, Mog03b, The04]. SCOPE [Gar09, NS01]. Scopira [DP09]. scoring [KRR23]. Scottish [Rit88]. Scotty [TGC^{+}21]. Screamer [Ano98]. screen [Gos83, Uni85a, Uni85e, Uni85c, Uni85d, Uni85f]. Scribe [Bla89b, Bla89a]. Scribus [Ano10, EKJ^{+}03, JP09b, PS^{+}09]. Scripted [CMTA19]. Scriptics [Ano00i]. Scripting [SL01, Bri09a, Bri09b, Fri97, Nob08]. scripts [BJM^{+}22, Kis90]. SCTE [MGFRG12]. Scuttles [Mog03b]. Scythe [PQM11]. SDK [Bar00c, Kro99b]. sDMD [ZJS^{+}20]. Search [CC04, HKA^{+}19, ACM05, XXCL19, AJLM18, Ban16, Ban17]. Searching [RM92]. Season [Eri99]. Sebastopol [Ano97a, Ano99a, Ano00a, Ano00b, Cas02]. Second [Ano91, Ano00c, BSK87, IEE92b, KD23]. second-order [KD23]. Secondary [LBF^{+}22, Ewe18]. Secrets [HW17b, Gil04]. Section [HW17a]. Sector [Fav23, Fre23, Hol23, LRBM23, VJ23, Hic04]. Sectors [LBF^{+}22]. Secure [Coh02, Lin02a, MCGA22, Sav23, ADF^{+}21, BM02, Bou05, For07, SS05a].

SecureBallot [ADF^{+}21]. Securing [PWA^{+}19]. Security [Ano01j, Ano04b, Ano18, Bra04, Coc01a, Cow03, CSD^{+}05, IEE92c, IEE05, IAS16, IKW23, Kim01b, Law02, Lin02a, MSW09, RSBP23, SJV^{+}05, SuS01, TB05, USE00b, WLC01, dCdCM14, AM18, Ano02a, Bau06b, Bau06a, FvH03, HKP02, Hic04, HJ07, KG20, Mac18, Mag01c, Mur20, NDDH^{+}21, Pay02, Sch11, Sie04, Sil13, Lin02a, McC05]. Security-critical [dCdCM14]. see [Dew07]. seeds [CAC09]. Seeing [GW10]. Seek [Coc01a]. Seeking [CRW^{+}04]. Seismic [FY18]. Select [Spi19, LMZT22]. Selected [Bar00c, BS14, BK14, Fio03, Gay02, YWA07]. Selecting [UMV15]. Selection [Sha04, AMOS19, BH11]. selectivity [BMS^{+}22]. Self [Ano03a, CRW^{+}04, Ell12, FKM^{+}11, PLO^{+}23, Sta79, Sta84, YXS^{+}19, ZFD21, jFFR16, HSX^{+}18, Sta80a, Sta81e, Sta81a, Sta81d, Sta81c, Sta81b]. Self-Admitted [PLO^{+}23, YXS^{+}19, ZFD21, HSX^{+}18]. Self-Contained [Ano03a]. Self-Documenting [Sta79, Sta84, Sta80a, Sta81e, Sta81a, Sta81d, Sta81c, Sta81b]. self-organized [jFFR16]. Self-Paced [Ell12]. Self-Service [CRW^{+}04]. Self-tuning [FKM^{+}11]. selfisher [BMS^{+}22]. SELINUX [McC05]. Selling [BBD^{+}96a]. Semantic [Coc01a]. semantics [GB20, Kli90]. semantics-supported [Kli90]. semaphores [Rog11]. semi [ABC18]. semi-structured [ABC18]. Semiconductor [GNGS17, AH19]. Sendmail [All02b]. Sense [Fal03, Mog03b, FFvdH01]. Sensing [Joh99, Eds16, ZPH^{+}15]. Sensitive [FQYS23, PKH07]. Sensor [BY14, DFCPSF15, GTMR23]. Sensornetzwerke [TNM17]. Sentences [MGM^{+}02]. Seoul [HY14]. Separation [WRDP17]. separations [DSK19, GDK21]. September [AT92, Ano94b, BSK87, CS99, FvdHJ10, Fur90, MG94, SM89b, USE94]. Sequence [AL07, Bar01, Haf01]. sequences [WN15]. sequential [ALGE12]. Series [BFC02, Fox08]. Serious [Per00]. Server [Ano96b, Ano96c, Ano97c, Ano00i, Ano00k, Ano01i, Bot03, Bra04, Kro99b, Reh01a, BMR^{+}23, WB07, AJ05, Kro99a]. Server-Lösungen [Bra04]. Serverless [DKK22, PCAJ^{+}23]. servers [KMG^{+}93, Sai01, TB05]. Service [BE06, CRW^{+}04, GB00, Nej12, ZZZ22, CRB^{+}18, Fow93, PPC^{+}15]. Service-Centric [BE06]. Service-Learning [Nej12]. Services [Ano00j, Ano04b, BE06, HOL^{+}07, Aji17, BM02, CLL05, Fri16, Ano96e]. Session [Wat85a, Joy09a, Wat87]. sessions [Dig82]. Set [Ano93b, Hae02, MCS12, MSZ^{+}01, Sch90b]. Set-Top [Hae02]. SETI [Bar01]. sets [Ano00f]. Setting [BHP^{+}01, EJS^{+}01, Hec99, Sta01b, Wil99, Amb15, GGH05]. setzt [Ano00f]. seventh [Cha13]. several [Mud97]. SGDL [Ano01j]. SGI [Bar00c, Kro99a]. Shackled [Sta04b]. Shadow [Sta04a]. shape [jFFR16]. Share [CWB^{+}04, CKB^{+}05, LMWM18, Ano15b]. Shared [BES^{+}01, BHP^{+}01, Fie99]. Shareware [Geh96, NO03]. Sharing [CH10, CSP^{+}03, Goo14, Har05, NMG11, Sta03c, FN21, MS02, Sin08, SSA08, Wii91b]. Shell [Ram94c, Ram94b, Fow93, GB20, Ram94a]. shells [ZLF^{+}22]. shifted [HKY^{+}21]. ship [LA10]. Ships [Ano01i, CTP^{+}22]. Shoah [Ano00j]. Shop [Hec99]. short [MDRN18, PSS^{+}07]. Shortage [CDsJ^{+}00, CMJ^{+}04]. Should [AS03, HR11, MFS15, VVM08, JDB09, SSAO04, Smi17, Spi21]. Showcase [USE00a, USE01a]. sichere [Lin02a]. Sicherheit [Bra04]. Side [BBD^{+}96a, GYW^{+}23]. side-channel [GYW^{+}23]. Sidebar [Wea03]. SIEM [VSdCCR23]. Sierra [BBE^{+}20]. SIG [Dig82]. SIGAPP [DGBH93]. SIGDOC [ACM89, ACM93a]. SIGGRAPH [ACM88]. sight [PSL21]. SIGMOD [FMA02]. Signage [CTP^{+}22]. signal [Gar09, KORP95, RA16]. SignalPlant [PNK^{+}23]. Signatures [BR95, LA10]. Significance [Rus09]. SIGOA [Abr81]. SIGPLAN [Abr81, FP95]. SIGPLAN/SIGOA [Abr81]. sim [HMYH22]. sim-trhepd-rheed [HMYH22]. Similarities [Cap13]. similarity [NRRS20]. Simphony [PGC21]. Simple [Gol06, MMP^{+}22, Wag03, CLS95, Kaw92, SP12]. Simplicity [Kim01a]. simulate [BFI^{+}21, JWC18, WGG16]. simulating [BM22, WFDK19, XAPK14]. Simulation [JPOB20, LL14, PGC21, SJW22, Ten93, WCHRM21, ASAB02, ASC^{+}21, BSW^{+}14, BMT^{+}20, CBRSH22, DPH16, DBP^{+}18, EHHH06, FTZ^{+}23, JD19, KDM17, KGT22, MBR21, SMRM^{+}17, SCR05, SCFR06, SWTC23, TL17, WHJ15, WGG^{+}19, ZJS^{+}20, ZC95]. simulations [Aki16, DGJH19, LPC^{+}15, MLZ^{+}23, MVF20, Nor23, ORI^{+}10, QSX^{+}15, TPK^{+}21b, VBG^{+}10]. Simulator [LFN^{+}11, ZKCS91, GGSRMPM20, HMYH22, KGW^{+}21, MGPB20]. simulators [AVA^{+}16]. Simultaneous [Joh18]. Sin [HR19]. Singapore [Ano06]. Single [Rie20, Coc03]. Single-Vendor [Rie20]. Sinkholes [Sor06]. Sinn [Fal03]. SIP [PM00]. sistema [RÓ01, VD01]. Sites [EJS^{+}01, PKGA22]. situational [KN93]. Six [Goo14, KKN^{+}21, MG12]. sixth [Ano94b]. Skepticism [RSAT19]. skill [JZ09]. skills [Bik96, HTU96]. Skin [Ano01j, Amb15, BSK^{+}15]. sky [Hol05]. SlicerRT [GKL^{+}14]. slicing [KFYI13]. slope [RAW^{+}16]. Slot [Sha03]. Slow [CPJ^{+}98, IKW23]. SM&A [Ano00j]. Small [GNR^{+}09, MH07, Ros02a, Fär05, Hay05, Ing92, LHZ12, RH21, Sin10a, LHZ12, NRG^{+}99]. small-angle [LHZ12]. small-world [Sin10a]. SmallEiffel [CZ99]. Smaller [Bar00c]. Smart [DDJ98a, DDJ99, DDJ98b]. smears [Ray01a]. Smells [DLT^{+}23, TPK21a, WFF18]. smelly [SK12]. SMEs [Lam09]. Smilei [DBP^{+}18]. SMMP [EHHH06]. SMMP-open-source [EHHH06]. smoothed [ZRZ^{+}21, CDR^{+}15, HMX21a, HMX21b]. SNA [Ano00j]. Sneaky [IKW23]. Snippet [Bak20]. Snort [WJM22]. Snow [DDJ98a, DDJ98b]. Snowbird [SC00]. Soapbox [Wil99]. Sobell [Men12]. SoC [AML^{+}10, GCE^{+}21]. Social [BAR16b, BCB^{+}17, MTM^{+}19, PMM17, PMM18, Waa09, CH11, KF17, Muw09, YLHW21]. Societal [Sta97b]. Society [Fio03, Gay02, SM89b, CPJ^{+}98]. Sociology [Gla00]. sociopolitical [Gla03b]. SoCs [PGW^{+}20]. SODECL [ALA20]. Softech [Ano01i]. Software [ACM88, Abe07, ASS^{+}23, AtHR11, ABC^{+}14, Aki16, APK14b, AMS03, Ano86, Ano95c, Ano96b, Ano96c, Ano96e, Ano97d, Ano98, Ano99d, Ano00i, Ano00j, Ano00k, Ano01a, Ano01e, Ano01i, Ano01j, Lin02b, Ano03b, Ano03c, Ano04b, Ano04c, Ano10, Ano15c, Ano16, ACHC11, BBD^{+}96a, BRH10, Bal19, Bar00b, Bar00a, Bar22, BS14, BKR^{+}20, BSFR22, Bea21, BPG94, BK14, BFC02, BAP00, BVLF14, BNSW15, Bla06, BSA22, BKHT21, Bon11, BJJ14, BSA14, Bra97, Bro19, BB02, Bro05, Bro96, BK02, BGL^{+}21, CK07, CFM08, CO12, CH10, Cha01b, CCSW10, CC03, CPJ^{+}98, Coc01a, Col09a, CK10, CW15a, CW15b, Cow03, CGK^{+}02, CMJ^{+}04, CRW^{+}04, CSD^{+}05, CGB^{+}05, Cur99, Cur02, DDJ98a, DDJ99, III01, De’15, Del01, Den13, DWP^{+}14, DKMB14, Doh01, DM15a, Ebe08, Ebe09, Edd96]. Software [EXA^{+}05, Ell12, Est06, Far91, FSB^{+}01, Feixx, FFvdH01, FFHL05, FFH^{+}05, FFHL07, FMFZ19, Fie88, Fie90a, Fie90b, Fio03, FK04, Fit11, Fox08, Fri97, FVD^{+}12, GP12, GKL^{+}14, GJMPAM^{+}14, GIA^{+}06, GF11, Gay02, Gil05, Gla03a, Gla08, GF17, Gol06, GBICMR13, GB21, Goo14, GAS^{+}01, Gra01, GM05, GNR^{+}09, Gra99, Gre18, GW09, GEMN07, HK03, HWZxx, HWZ01, mH00, Hal02, Ham99, Har99, HCH^{+}20, HLS^{+}13b, Hec99, HE17, HM19, HW17a, HW17b, Hol23, HY14, Hug95, HBGS19, IEE94a, IEE95b, Jak03, Jak04, Jam09, Jen01, Jen97, Jin18, JEB^{+}23, Joh02, Jor04, JS07, Kam11, Kam14a, Kam14b, Kar03, Keh94, KP76, KP81, KY16, KMF^{+}07, Koc07, Kos21, KS11, Kri03, Kro99b, Kro99a, Kro00, Kuk98, KSD^{+}12, Lam09, Lee99, LMM02, LPFD21, LZWH22, LL14, LRBM23, LRP11, Lit14]. Software [LO97, LBF^{+}22, LMZP19, MMP^{+}22, MTM^{+}19, Mar05, MH07, May06, McG01, McL05, Mee12, Men10, Mic04, MMB^{+}90, MN04, Mog01c, MS12, Mor08, MG94, MB16, NR03, Nej12, NO03, NK04, O’Rxx, OG07, Omb20, PMBM^{+}15, PM00, PMD13, Per00, PPRB07, PRRL12, Pom04, Pri19, PMG^{+}09, PKG^{+}10, PPG^{+}11, PBJ^{+}12, QSX^{+}15, Rac06, Raf23, RB05a, Rav00, RW87, Rie07, Rie15, RGCS14, Rob94c, RSAT19, Ros01a, Ros01b, Ros01c, RDZ20, RE04, SBDR22, San98, Sav23, SFF^{+}06, Sca06, SB08, SC02, ST10, SG05, SS06, SMRM^{+}17, Sie04, SCB04, Sim12, SSP17, SSP18, Spe01, SS04, DDJ98b, SG92, Sta96c, Sta98b, Sta02a, SVAGB20, Sta02c, SDD06, SV03, Stö04, TV99, TZ22, TBPS15, TRM16, Tay00, Thi22, Tot06, TWS^{+}22, Tro96d, Tro96a, Tro96c, Tro96b, Tro97, TDBEE11]. Software [The04, Ude97, UMV15, Veg06, VVM08, Waa09, Wal01, Wal99, WW01, WKS^{+}14, WM19, WCG22, WWSG21, Whe03, WP04, Wil02, WM05, WKB14, WL01, WKA^{+}08, XXAD21, XGF^{+}23, Yes12, YYL^{+}15, Zha16, ZRNA20, dCdCM14, dBLMT11, vdLLM09, vWHvW09, ACM93b, AKHG16, AMOS19, Aji17, AW07, ACB18, ATM22, APK14a, ABF^{+}14, AMWH19, AMC16, ALVV17, Amb15, Ano01g, Ano02a, Ano08b, Ant16, ASAB02, AKF21, ATHW92, Asu05, ASC^{+}21, BTL^{+}11, Bab02, BD03a, BVT06, BHMB03, Ban16, Ban17, BCPS10, BDAW15, BD03b, BJM^{+}22, BLG^{+}17, BMR^{+}23, BH11, BAR16b, BSK^{+}15, BR03, BMT^{+}20, BAE14, Bou05, Bro04, BMS^{+}22, BG12, BGL^{+}20, BGL^{+}22, CF07a, CK08, CGZ17, CFMRL11, CV13, CNSR23, CLM^{+}08, CJ17, CJ19, CYOS19, CG17, CKS16, CH11, Col09b, CH06a, CSEP14, CSP09, CHA06, CWHW12, CWZ06, CdSV07]. software [Cus04, DPH16, DSB^{+}16, DWJG02, DB05, DD17, DDA^{+}07, Don04, DM15b, Eds16, EKUR10, EHHH06, EMdL^{+}07, Emb06, EHP14, FHH11, Fog06, For12, jFFR16, Fow00, FvH03, FG16, Fri16, Fug03, FvdHJ10, GBG^{+}16, Gal60, Gal01, GL14, GEI^{+}11, GLT08, Gau03, GGT05, Ger03, GKP^{+}14, Goe07, Gom99, GM02, GSW08, GVOM09, GV16, GFZ16, GGH10, GW10, GFS05, HK09, HAC^{+}23, Har05, HOST05, HBR19, HR94, HLS^{+}13a, Hea09, HPM^{+}08, HNH03, HBB^{+}12, HETD09, HZS^{+}16, IHBS14, JCNS^{+}22, JK11, JK12, Joh92, JDB09, Jør01, Joy09a, JZ09, KOI94, KTF15, KKT17, KJ03, KNS18, KC22, KS02, KK17, KH05, Kop05, KS03, KFYI13, KG20, Kus05, eLAA^{+}23, LA10, LPC^{+}15, LG02, LSJ^{+}06, LSM09, LMZT22, LHZ12, LH03, LH14, Luc99b, LLS11, MG12, Mah03, MV05, MD17, MCS12]. software [McA08, MLA^{+}19, MTBS09, MMY^{+}19, MPE^{+}11, MFH02, Mol01, MAF22, MTD^{+}09, MRS07, MM10, MSR09, MSR10, Muw09, NRRS20, NYB10, NXC13, NN00, NMG11, Neu99, NGJ03, NDDH^{+}21, O’R99, O’S03, O’S04, OMA^{+}22, ORI^{+}10, Ous99, ODP15, PSSH16, PKGA22, PAB^{+}17, PSE04, Pay02, Ped05, PH16, PSDG18, PYM^{+}06, PSS^{+}07, PT91, Pya06, QLC^{+}12, Rad89, Raj13, RBM^{+}23, RCO20, RHW^{+}21, Rob05, RCGB^{+}22, RP08, Ros05, RC10, Rud10, RT05, Sal08, SSAO04, SBM^{+}10, SMO^{+}13, Sca05, SA15, Sch09, SHB^{+}20, SSR02, SIK^{+}13, Sil13, Sim05, Sin08, Sin10b, SK12, SCR05, SSS^{+}14, SM08, SSA08, Spi03, SG12, Spi21, CAC09, Sta09, SAOB02, SDD05, SG06, SDeaK^{+}09, Sur01b, TZH22, TLL^{+}14, TL17, TV13, THG23, VGP^{+}19, WLD^{+}17, WGS07, WHJ15, Wan21, WGG16, WGG^{+}19, WG06, WBB01, WBGM02, WZS^{+}18]. software [YL08, YLG05, YLXZ16, YLHW21, YZC22, Yap11, YA11, YT22, YA05, YKSH20, YMCF23, Yu06, ZKDP22, Zei03, ZDM10, ZW17, ZFY^{+}19, ZD05, Zic01, ZK21, dA15, vGPB10, vKSL03, vKvH03, Ano96e, Ano97b, Ano01i, Ano01j, Ano04b, BES^{+}01, Bol02, CK06g, CK06h, CF09, Jak04, Kro99b, Kro00, LMWM18, MP12, Mog99, Mog01a, Mog01b, Mol01, NO03, Rud10, SMS16, Séd02, SG05, Shi12, Sie99, Sie04, Spi03, St.04, Stö04, Sur01b, Ano97a, Cas02, Rus09]. Software-based [GF17]. software-defined [SSS^{+}14]. software-inspired [MSR09]. software-intensive [BGL^{+}22]. Softwarelizenzen [Jak03]. Softwaretools [Jor04]. SOGo [Mar11]. Solaris [DF00, Ano01j, Kro99a, Sec95, Woo01]. Solaris-compatible [Woo01]. Solaris-to-Linux [Ano01j]. Sold [RAH^{+}01]. Soldier [RAH^{+}01]. solid [BCP^{+}16, NGCI^{+}12]. solids [BFI^{+}21]. SOLIS [Bro04]. Solution [Ano96c, Ano00k, Hom00, LMW12, ST10, BBE^{+}20, EKUR10, GLMS18, LZ12, QSX^{+}15, VBG^{+}10, VSdCCR23, XTG^{+}11]. Solutions [AMS03, Ano00i, BdSI15, BGG^{+}94, Bra04, Cur99, Gan17, Kuk98, MB16, Nor02, PKP05, RDKT12, SVAGB20, Tro96d, Tro96a, Tro96c, Tro96b, Tro97, Hay05, Lam09, PWA^{+}19]. solvation [DSM^{+}19]. Solved [MZH22]. solver [BSW^{+}14, BZB17, BSC^{+}21, CDR^{+}15, DBLF16, FBY^{+}17, FTZ^{+}23, GDJG23, GB06, HXS20, HWM^{+}15, Hua17, KDM17, KD23, LMHL20, MVS15, ORS^{+}14, RJ21, SWTC23, TL17, XOTI22, YKK23, Zag14, ZCG17, DFU20]. Solvers [MCS12, DC23]. solves [Fri97]. Solving [Gro01, MTM^{+}19, Tie88, Tie93, FK99, jFFR16, HWL^{+}23, KGW^{+}21, KK17, SS23, WFV14, YSVM^{+}16, YSMA^{+}17]. Some [AS03, EKJ^{+}03, McL05, MI07, Wil71, Rob11, Bur04a]. SonicMQ [Ano00j]. Sony [MLWR18]. Sorting [CDsJ^{+}00]. Sorts [CDsJ^{+}00]. SOT [Ano00j]. Sound [Ano97c, Str94]. soup [Fis69]. Source [Abe07, ASS^{+}23, AtHR11, Adl00, ABC^{+}14, APCs22, Alf05, All02a, All02b, AM03, AM04, AMS03, Ang01, AWD^{+}18, ACC^{+}12, AHB^{+}09, Ano00d, Ano00f, Ano00k, Ano00l, Ano01e, Ano01i, Ano01j, Lin02b, Ano02b, Ano03a, Ano03b, Ano03c, Ano04b, Ano04c, Ano08c, Ano11, Ano15c, Ano16, Ano18, AML^{+}10, AS03, ALA20, ACHC11, BdP13, BGG^{+}15, BRH10, BMF^{+}16, Bal19, BC20a, Bar00c, Bar01, Bar22, BS14, BKR^{+}20, BSFR22, BdSI15, BY14, BlG12, BES^{+}01, BMZ14, Bax01, Bea21, BK14, BCB07, Bel22, BAP00, BM12, BVLF14, BNSW15, Bla06, BSA22, BKHT21, BNST99, Bol02, Bon11, BHP^{+}01, BGG^{+}00, BJJ14, BCB^{+}17, BSA14, Boy08, BKP05, Bro01, Bro19, BB02, BK02, BGL^{+}21, CC04, Cap12, CFM08, CCA^{+}19, CO12, Cha01a, CH10, Cha07, Cha98, CTP^{+}22, CSY^{+}04, CWM^{+}20, CZ22, CX23]. Source [CYL^{+}23, Chi01, CCSW10, CC03, Cho09, CBB06, CPJ^{+}98, Coc03, CPG^{+}04, Coh02, Coh03, CF09, Col09a, CK10, Cou20, Cow03, CdR99, CWB^{+}04, CMJ^{+}04, CRW^{+}04, CKB^{+}05, CH06b, CB12, Cur99, Dal02, DBBA10, DMJ05, DXT^{+}18, III01, De’15, DKK22, Del01, Den13, DWP^{+}14, DKMB14, DFP23, DKMT11, Doh01, DM15a, DFT21, DGC^{+}07, DMP^{+}02, Ebe07, Ebe08, Ebe09, Egy01, EE01, EJS^{+}01, EGK^{+}02, Ell12, Eng10, Ens04, Est06, FGBM14, Fal03, FL16, Far06, FSB^{+}01, Fav23, Feixx, FFvdH01, FFHL05, FFH^{+}05, FFHL07, Fer03, Fie89, Fit04, Fit11, Fox08, FT09, Fre23, FQYS23, FCTP21, GP12, Gag02, GKL^{+}14, Gal10, Gan17, GIA^{+}06, GF11, Gau07, GD12, GMBv20, GCE^{+}21, Gla99, Gla00, GGL21, GB21, Goo14, GAS^{+}01, Got05, Got07, Gra01, GGB17, GNR^{+}09, Gri16, GW09, Gro01]. Source [GEMN07, GNGS17, Gus20, Gut00, HK03, HKA^{+}19, Hae02, Haf01, Han00, HT21, HKP02, Har99, Har20, HCH^{+}20, HOL^{+}07, Hau01, Hec99, HR11, Her20, HE17, HM19, HW17a, HMKC12, Hoh01, Hol23, Hub04a, HBGS19, IAS16, IKW23, ILG10, IC23, JWC18, Jen01, JPOB20, Jin18, Joh99, Joh02, JJ00, Jon02, JS07, KC21, Kar03, KNS18, KGM^{+}16, Kenxx, KY16, KCAS23, KMF^{+}07, Kim01b, Knu99a, Knu99b, KKN^{+}21, KHA^{+}03, Kre03, KS11, KJRD16, Kri03, Kro99a, Kro00, Kuc06, KG01, KRB^{+}22, Lam09, LFN^{+}11, Law02, Law09, LMM02, Lev23, Lew99a, Lew99b, Li18, LPFD21, LMPT22, LZWH22, LLWM23, LL14, LMWM18, LRBM23, Lin08, LRP11, Lio96, Lit14, LC12b, Luc99a, LGW18, LBF^{+}22, MPG^{+}16, MSSvK08, MMD^{+}22, MMP^{+}22, MTM^{+}19, Mar01, MH07, May06, Maz15, McA19, McC99c, Mee12, MSW09]. Source [Men10, MFS15, MP12, MCGA22, MMD12, MN04, MS00, MK12, Mog01c, MOMM11, MS12, Mor08, MEB^{+}20, MB16, Nas04, Nej12, NO03, Noj01, O’D07, O’Rxx, OG07, Omb20, Oms03, Owe01, PSSH16, PMM17, PMM18, PLZ^{+}22, PMBM^{+}15, PM00, Pau04, PQM11, PMD13, Per00, PH16, PPRB07, PGW^{+}20, PZ20, PLO^{+}23, PK10, PRRL12, PGC21, PSP^{+}22, PFL^{+}12, PBH01, Pra03, Pri19, PMG^{+}09, PKG^{+}10, PPG^{+}11, PBJ^{+}12, QC18, Raf23, RSBP23, RB05a, RT12, Raj23, Ray98, Ray99b, Ray99a, Reh01b, RDKT12, Rie07, Rie10, Rie11, Rie15, Rie19, Rie20, Rie21, RCP^{+}12, RGCS14, RSAT19, Rob20, RLVdS21, Ros02a, Ros01a, RCB^{+}14, RLTD23, Rud10, SD16, SJV^{+}05, SBDR22, San98, San08, SS02, Sav23, SFF^{+}06, Sca06, Sca19, SB08, SC02, Sch19, SMS16, ST10, Sea99, Sea02, SG05]. Source [SS06, SK04, Sha10, SSM^{+}07, SSH22, Shi12, SAC^{+}15, Sie04, Sif00, SKSM19, SV19, SCB04, Sim12, SSP17, SSP18, SCDS15, SL01, SFWD12, Sor06, Sor01, Spe01, Spi03, Spi06, Spi11, Spi19, ACM05, Sta89a, Sta98b, SPS^{+}00, SPS^{+}02, SVAGB20, Sta06, Sta02c, SHK^{+}03, SJW22, STG19, SDD06, SF15, SV03, Sto99, Str02, SH19, Sur01b, TPK21a, TZ22, TGC^{+}20, TBPS15, TRM16, TNM17, Ter00, Thi22, TRB22, TTB09, Tot06, TGW^{+}22, TGS22, TWS^{+}22, TG15, TF21, The04, UMV15, UNF^{+}08, Van22, Veg06, VVM08, VJ23, VOM12, Waa09, WCHRM21, WW01, WKS^{+}14, WM19, WFW^{+}20, WCG22, Wat01, Wea03, Wen00, Wen02, WWSG21, WG00, WDK^{+}20, Whe03, WP04, WB02, WM05, WKB14, Wil99, WLC01, Wol03a, WBRH23, WL01, WKA^{+}08, Wut12, XWZ^{+}23, XXAD21, XGF^{+}23, YLL^{+}07, YMLT14, Zha16]. Source [ZRNA20, ZZZ22, ZXB^{+}23, dBLMT11, dlCKK15, vdLLM09, vWHvW09, ABC18, AKHG16, AMOS19, ADF^{+}21, AH19, Aji17, AW07, ACB18, Aki16, ASAAM^{+}19, ACB^{+}16, ALGE12, ATM22, APK14a, APK14b, ABF^{+}14, AJLM18, AMWH19, AMC16, AAB^{+}05a, ABNÅ05, AAA^{+}12, AAA^{+}14, ALVV17, AM18, ACKT20, Amb15, And01, And08, AVA^{+}16, Ano99a, Ano99c, Ano00g, Ano00h, Ano01f, Ano01g, Ano01h, Ano02a, Ano03e, Ano04a, Ano04d, Ano08b, Ano19, Ano21, Ant16, ASAB02, AKF21, AAB^{+}05b, AG22, AKMS23, Asu05, ASC^{+}21, APHV19, AZ17a, AZ17b, AFZ17, AFZ18, ATCZ19, AD04, BTL^{+}11, BOL14, BCR^{+}08, BHMB03, BMF^{+}19, Ban16, Ban17, BCPS10, BDAW15, BA15, BCP^{+}16, Bea04, BCvE^{+}05, BLG^{+}17, BYV08, BH11, BBE^{+}20, Ber22, BSW^{+}14, BCI^{+}09, BAR16b, BC20b, BSK^{+}15, BM22, BMT^{+}20, BFI^{+}21, Bon02, Bor09]. source [BCG^{+}14, BAE14, Bou05, BP14, BZB17, BMS^{+}22, BG12, BDP^{+}14, BSC^{+}21, BGL^{+}20, BGL^{+}22, But94, CF07a, CK08, CGZ17, CCA^{+}13, CMC^{+}15, CAWK22, CFGS05, Cap13, Car01, CZS^{+}21, CFL23, Cas19, CV13, CV22, CRB^{+}18, CF07b, CM06, CNSR23, CLL05, Cha11, CP04, CLM^{+}08, CJ17, CJ19, CYOS19, CFW17, CG17, CKS16, CH11, CKGW22, CBRSH22, Cio01, CFW01, Col05, Col09b, Com99, CH06a, CSEP14, CMTA19, CSP09, CSV^{+}07, CDR^{+}15, CHA06, CWHW12, CWZ06, CKB11, CdSV07, DPH16, DSK19, Dan11, DGJH19, DRM21, DB02, DSB^{+}16, DIK^{+}23, Dei10, DFCPSF15, DP09, DWJG02, DBP^{+}18, DD17, DSM^{+}19, DFU20, DTB05, DDA^{+}07, DD08, DM15b, DD10, DBLF16, DO16, Dwa04, ESM19, Eds16, EKUR10, EHHH06, EMdL^{+}07, Emb06, EHP14, ES23, Eub05, Ewe18, FLA^{+}16, FN21, Far23, FL15, Fei23]. source [FBY^{+}17, FTZ^{+}23, FHH11, Fog06, For12, For07, jFFR16, Fow00, FM10, Fra13, FRBRF19, Fra19, FG16, Fri06, Fri16, Fug03, FVD^{+}12, GBG^{+}16, Gal01, GLMS18, GGSRMPM20, GLCMC17, Gar09, GEI^{+}11, GLT08, Gau03, GMPS14, Gen99, GCK^{+}17, GGT05, GDK21, Ger03, GKP^{+}14, GDJG23, Gla03b, Gla04, Goe07, GF17, GRJS01, GSW08, GVOM09, Gv14, GV16, GFZ16, GPPT16, GTMR23, GYW^{+}23, GFD^{+}24, GGH05, GGH10, GW10, GFS05, HK09, HMO^{+}18, HMYH22, HAC^{+}23, HBC^{+}05, Har05, HBR19, HXS20, HFO^{+}12, Hay05, HLS^{+}13a, HLS^{+}13b, HWL^{+}23, HZ14, HMP^{+}15, HPM^{+}08, HPT17, Hic04, HL02, HBB^{+}12, HETD09, HJ07, Hol15, HSF^{+}15, HBZ09, HKvH16, HKY^{+}21, HC07, HYA20, HWM^{+}15, Hua17, HSX^{+}18, Hua23, HZS^{+}16, HMX21a, HMX21b, HM10, IHBS14, IDSM23, IC22, JP09a, JCNS^{+}22, JP09b, JH16, JEB^{+}23, JNN12, JK11]. source [JK12, Joh94b, Jon01, JDB09, Jør01, Joy08, Joy09a, Joy09b, JČMG11, JZ09, Kam21, KTF15, KKT17, KDM17, KTH^{+}22, KC22, KMG^{+}07, KHMA12, KPK^{+}17, Kle21, KS02, Koc07, Koc09, KGW^{+}21, KSH14, KKA^{+}19, KK17, KH05, Kor11, KT05, KL07, KGT22, KKA^{+}21, KSS^{+}23, KTTK17, KRR23, KFYI13, KSD^{+}12, KG20, KD23, Kus05, eLAA^{+}23, LA10, LLEL^{+}23, LPC^{+}15, LG02, LW03, Lei04, LSJ^{+}06, LFB^{+}21, LSM09, LR08, LGS^{+}17, LQ17, LMHL20, LMZT22, LRP21, Liu08, LHZ12, LRD^{+}19, Lla06, LQR17, LZ11a, LZ11b, LZ12, LH22, LH03, LH14, LGA20, MG12, Mac18, Mam01, MV05, MD17, MD18, MLZ^{+}23, MBTB21, Mar11, CCK21, Mar05, MSB09, MLMFN^{+}15, MFB23, MGPB20, Mas05, May17, MCS12, McA08, McC05, McC02b, MLA^{+}19, MDRN18, MWG08, MNS19, MZE13, MVF20, Mił10, MGYC18]. source [MSM10, MMY^{+}19, MPE^{+}11, MFH02, Mon03, MM04, Moo01b, MVAXP22, MAF22, MdL09, MTD^{+}09, MVS15, MRS07, MGFRG12, MSR09, MSR10, MCQF21, MLWR18, MOT^{+}18, Mur20, MRN20, Muw09, MQN19, NRRS20, NYB10, NXC13, NN00, NMG11, Neu99, NS05, NN20, NN21, NGJ03, NDDH^{+}21, NT06, NMS14, Nob08, Nor23, NZPWR22, NGCI^{+}12, NMX19, O’R99, O’S03, OMA^{+}22, ORS^{+}14, ODP15, PCAJ^{+}23, PM13, Pag07, PKGA22, PKB17, PAB^{+}17, PSE04, Pay02, Pea16, Ped05, PBOP07, PL05, PPC^{+}15, Pet06, Phi12, DARJ23, PM21, Pit16, PS^{+}09, PWA^{+}19, PNK^{+}23, PHT17, PPR19, PC13b, PYM^{+}06, Pot06, PSS^{+}07, Pow14, PSL21, PdSCJM22, Pya06, QB21, QLC^{+}12, QSX^{+}15, Qui00, Raj13, RZWW23, RBM^{+}23, RCO20, Ray01b, Ray01a, RHW^{+}21, RH21, RJ21, Rob05, RCGB^{+}22, RAMB18, RP08, RA16, Ros14, RNR17]. source [RDZ20, RC10, RT05, SS05a, SZ05, SBS20, SSAO04, Sam06, SBM^{+}10, SMO^{+}13, San01, Sca05, SA15, SPLD20, Sch09, Sch11, SGM^{+}08, Sch04, SHB^{+}20, Sea04, SRGCPB^{+}09, SMRM^{+}17, SSR02, SSS22, SIK^{+}13, SC16, SPAW17, Sil13, Sim05, Sin08, Sin10b, Sin10a, SK12, SHW^{+}21, SGNB08, SKB23, SCR05, SCFR06, SSS^{+}14, SM08, SAHP15, SSA08, SG12, Spi21, CAC09, Sta88a, Sta89b, SPG92, SP93, SP95, Sta96a, Sta98a, Sta09, SAOB02, SS23, Ste08, SDD05, SG06, SDL^{+}16, STB23, SDeaK^{+}09, SWTC23, TZH22, Tai13, TPSZ19, TLL^{+}14, TL17, Tay19, TPK^{+}21b, TTL06, TACA15, TG99, TV13, TGC^{+}21, THG20, THG23, TKSC20, UBR^{+}17, VGSN18, VBG^{+}10, VGP^{+}19, VSdCCR23, VOK^{+}22, Vir05, VSN22, VB19, WJM22, WCS20, WLD^{+}17, WACBL03, WFF18, WGS07, WHJ15, WFDK19, Wan21, WMLM22, WGG16]. source [WGG^{+}19, WBY^{+}08, Web04, Wes03, WG06, WFV14, WBB01, WBGM02, Wol02, WNS^{+}21, WSK^{+}22, Woo01, WG05, WPAV14, WMK^{+}17, WZS^{+}18, XTG^{+}11, XFS^{+}22, XMGM21, XMGM22, XAPK14, XTY^{+}22, XOTI22, Yad07, YLG05, YWA07, YLXZ16, YLHW21, YZC22, Yap11, Yes12, YA11, YT22, YA05, YKSH20, YMCF23, YKK23, You08, Yu06, YSC^{+}06, Zag14, ZVvDD11, ZFD21, ZKDP22, ZAC^{+}23, Zei03, ZK05, ZSW14, ZRZ^{+}21, ZE00, ZE03, ZDM10, ZW17, ZLF^{+}22, ZFY^{+}19, ZJS^{+}20, ZD05, ZWH21, ZWU22, ZCG17, Zic01, ZLL04, dA15, dCdCM14, dlVRB21, vGPB10, vKSL03, vKvH03, Lin02b, Bar00b, BES^{+}01, BW00, BR03, CFMRL11, CDsJ^{+}00, DOS99, DiB04, DFLS05, DH01, EW01, Eri00, Feixx, FK04, GA04a, GL14, Gom99, GF99, HK03, Hac98, Has05, HNH03, JV01, KGMI06, KT04, Lee99, Lin02a, LLS11, Lus04, MD04, Maj03, Mal02, MSZ^{+}01]. Source [Man03, Mau05, McG01, McL05, Moo01a, MM10, NR03, NRG^{+}99, NO03, NK04, O’Rxx, Oms03, Per05, Pud04, Ray99c, Ros01b, Ros01c, Ros00, RE04, SSC^{+}00, Sca04, SG05, SCSC04, Spe01, Spi03, SS04, St.04, Ste99, Sur01a, Sur01b, TH04, VSM06, Wal01, WP04, Gil06, Ens05]. Source* [GM05]. Source-Code [BHP^{+}01]. source-code-level [HC07]. Source-Driven [Fav23]. Source-Level [Sta89a, SPS^{+}00, SPS^{+}02, But94, Sta88a, Sta89b, SPG92, SP93, SP95, Sta96a, Sta98a]. Source/2 [Mau05]. Source/Open [Adl00]. Sourced [Coc01b]. SourceForge.net [Koc09, MG12]. Sourcerer [BOL14]. Sources [CKB^{+}05, DOS99, DCS05, Sea99, SAC^{+}15, von88, KRR23, Vie97]. sourcing [PSDG18]. sous [Hom00]. South [De’15, HY14]. Southeast [ACM95, CH06a]. Southeastcon [IEE92d]. Space [BES^{+}01, Bes03, WCHRM21, Eds16, Sie04, Wen02]. Space-Based [WCHRM21]. space-borne [Eds16]. spacegroups [AZ17a, AZ17b]. Spaces [FFH^{+}05]. SpagoBI [Fra13]. spam [Mau05]. Spanish [RÓ01, VD01]. Spare [CRW^{+}04]. SPARK [HMW15, CZ22]. SPARTACUS [FTZ^{+}23]. Spatial [AMS03, Ell12, MGR16]. spatio [MLMFN^{+}15]. spatio-dynamic [MLMFN^{+}15]. SPC [GB00]. SPDX [KKT17]. SPE [Gad88]. speaking [Sam06]. Speaks [RHS^{+}04]. Spec [Bar00c, ALGE12]. Special [Ano00e, Cod75, DBBA10, vKvH03, Joy09a]. specialisation [VSGM14]. Specialist [SM89b]. specialization [vKSL03]. Specialized [dlPRGB99, Tay19]. specific [AHG94, AZ17a, AZ17b, LPC^{+}15]. specific-purpose [LPC^{+}15]. Specification [Bar01, Coc03]. Specifications [AMS03, Kli90]. Spectra [YDZ19, LSM09, PSS^{+}07, SC16, WFDK19]. spectral [CMC^{+}15, CZS^{+}21, DBLF16]. spectral/ [CMC^{+}15]. Spectrally [HW17a]. Spectrum [PM00, Blo04, ZPH^{+}15]. Speech [Ano02b, Col09a, WRDP17]. Speech-Enabling [Ano02b]. speed [CKGW22, HYA20]. Speeding [DDJ99]. SpeedShop [SGM^{+}08]. Sperm [SBM^{+}10, SMO^{+}13]. SPH [CDR^{+}15]. SPHinXsys [ZRZ^{+}21]. Spielesammlung [CK06g, CK06h]. spin [WPAV14]. spin-adapted [WPAV14]. spinning [Ude97]. Spire [Ano96c]. spirit [Pet06]. SPL [MAF22]. sponsoring [Ros05]. Spontaneous [GF11]. Spoofing [YDZ19]. Sports [BY14]. spot [KHA^{+}03]. Spotlighting [Ano06]. Spotlights [Ano95c]. spots [BCI^{+}09]. spp [HW17a]. Spring [Dig82, Ano92]. Springs [Kap92]. Sprinting [Got07]. Sprite [Kup93]. Spyglass [Ano96b]. SQL [Tan11b]. SQLite [Ano04d]. Square [BMZ02, eLAA^{+}23]. SSA [Nov03, Nov04]. St [Waa09, Dol91, TSM88]. St. [ACM97]. Stabilität [Bra04]. Stability [BSA14, Bra04]. Stacey [Ano00l]. Stack [LZWH22, Sha03, Och12, ZFY^{+}19, PKGA22]. StackGuard [Wag03]. staff [NN20]. stage [SG06]. staging [Cou17]. staining [Amb15, BSK^{+}15]. Stake [GB00]. Stakeholder [Rie07]. Stallman [Cas02, Neu84, Fio03, Gay02, Wil02]. Standard [Neh04, PM00, Sto99, Yeo05, AHM^{+}07, EMD03, GHL^{+}04, Neh07, Rap94, Bar00c]. Standardization [Egy01]. Standardizing [Cou20]. Standards [Ano95e, BMZ14, GB00, LBF^{+}22, Opexx, PBH01, PKP02, PKP05, Rus14, Sto99, ATM22, CF07b, GKP^{+}14, MTBS09, Sim05]. Standards-Based [GB00]. STAR [Coo95a, Coo95b]. STAR/MPI [Coo95a, Coo95b]. Starch [Ano15a]. Stardock [Ano01j]. Staroffice [GGK99]. Started [McA19]. Starter [Ano97c]. Startup [Str02, Wal01]. STAT [Ano00i]. State [BlG12, BES^{+}01, LLdI00, MS00, DDHS03, Mak03]. States [DGBH93, BH07, DPL^{+}91, Dre94, JWC18]. Static [LMZP19, MRGP20, LFA92, NDDH^{+}21, RP08, THG23, ZLL04]. Statically [Sid04]. Station [Ano00j, FSB^{+}01]. Statistical [Ano00i, PQM11, RGCS14, BMS^{+}22, GVOM09, HFO^{+}12, WN15]. Statistics [Fri16, RHS^{+}04, Aji17]. StatSoft [Ano01j]. Status [PKG^{+}10, WW01, WBGM02]. staying [ACM93a]. Steele [Neu84]. Steering [Rie11]. Steindachneridion [SMO^{+}13]. step [Wat85b]. Steven [Gil06]. stiff [MCS12]. Still [CRW^{+}04, CSD^{+}05, LRBM23, MI07, Waa09, Wol04, Wol03b, Mud97, THG20]. Stochastic [ALA20, CX23, KF17, FTZ^{+}23, FHH11]. Stokes [HWL^{+}23, MVS15]. Stone [Ano99a]. Stop [Ray98]. storage [PPR19]. storia [Cor05]. stories [Cor05]. Story [Ano05a, Sto99, TV99, BH17, For12, Moo01b]. Straight [BHP^{+}01, EJS^{+}01, Sta01b]. strategic [Far23, SM08]. Strategies [BBM^{+}21, Egy01, LBF^{+}22, SMNF88, KMG^{+}07, Wes03]. Strategy [Coc03, CFGS05, Cus04, RSZ96]. Stream [Phi93, TGC^{+}21]. Streaming [Ano04b, WFV14]. Streaming-Media [Ano04b]. Streams [Ano00k]. strength [Mur20]. Stretched [Wut12]. Strictly [Ano00l]. Striking [Gal01]. String [Mor92]. strive [SSAO04]. Strongly [HOL^{+}07]. Structural [BNSW15, KSD^{+}12, ZRNA20, CFMRL11, SM08, eLAA^{+}23]. Structure [AFS81, AFS82, Coh82, Cra90, ZAC^{+}23, AFZ17, AFZ18, ATCZ19, Bow05, BSP11, CSV^{+}07, FLA^{+}16, GM84, Kam21, LZ11a, LZ11b, MSB09, MFB23]. structured [ABC18, San78b, Sta78b]. Structures [Gil05, JEB^{+}23, SSC^{+}00, AZ17a, AZ17b, EHP14, LZ12, RP08, ZLF^{+}22]. student [GSW08]. Students [Nej12, EMdL^{+}07, HETD09, MAMC05, MdL09, Spi21]. studied [SBM^{+}10]. Studienarbeit [Geh96]. Studies [Goo14, EKUR10, Emb06, Gal01, GYW^{+}23, MG12, MFH02]. Study [Ano04c, BNSW15, CFM08, CWM^{+}20, CASA22, Gau07, GMBv20, JEB^{+}23, KGM^{+}16, KCAS23, LMZP19, PMM17, PMM18, RSBP23, RÓ01, SC02, SCDS15, SDD06, WBB^{+}74, Zad02, ZXB^{+}23, AKHG16, AW07, ACB18, BLG^{+}17, BH11, BSW^{+}14, BGL^{+}20, CIC13, CV13, CNSR23, CLM^{+}08, CJ17, DRM21, DD17, DTB05, DDHS03, Gau03, Ger03, GPPT16, GGH05, GW10, HZS^{+}16, JK11, JK12, KFYI13, LLEL^{+}23, LGS^{+}17, LMZT22, CCK21, MLWR18, NXC13, NDDH^{+}21, OMA^{+}22, PKB17, PAB^{+}17, PSE04, RC10, SSS22, Shi12, SCR05, SSA08, THG20, TKSC20, VSGM14, WMLM22, YAS91, ZFY^{+}19, ZRGJ21, ZWU22, vKSL03, Geh96]. Studying [LZWH22, SIK^{+}13, ZVvDD11, ZWH21, ZWU22, CJ19, Eds16, KF17]. Stuff [BBD^{+}96a]. stunning [Sha05]. Stuttgart [BSK87]. SU-E-J-42 [GKL^{+}14]. SU-E-QI-18 [BJJ14]. SU-E-T-253 [DKMB14]. SU-E-T-465 [TBPS15]. SU-E-T-570 [PBJ^{+}12]. SU-E-T-816 [PPG^{+}11]. SU-FF-J-158 [GNR^{+}09]. SU-FF-T-118 [PMG^{+}09]. SU-FF-T-75 [KMF^{+}07]. SU-G-BRB-02 [KY16]. SU-GG-T-141 [PKG^{+}10]. SU-GG-T-393 [WKA^{+}08]. SU2 [ZAC^{+}23]. Subject [Ano04a]. Subroutines [Cod75]. Subscription [Ano95a]. Subset [Ano96e]. subsets [QC18]. subspace [HKY^{+}21]. Substring [Joh94a]. subsurface [JD19]. subsystem [Mit84, VSN22]. subtract [Sib17]. subtract-with-borrow [Sib17]. subtype [BR95]. Subworkshop [BAP00]. Succeed [Bro19, BR03, Gom99]. Success [Ens05, Gil06, Ray99a, Sto99, TV99, TGW^{+}22, CHA06, MP12, Sin10a, Web04]. Successful [Fre23, FCTP21, STG19, Fog06, SM08]. Successor [Ano95c]. Such [CPJ^{+}98, Mic04, Fie90b]. Suchen [Gün02]. Sudhanshu [TG15]. Suite [Ano96e, Ano02b, GvdHPR14, Kro99b, Kuc06, Fra13, GYW^{+}23, MM10, Vir05, Ano01i]. suites [ALVV17]. Suits [Sea02]. sul [Mol01]. Summary [BAP00, SZAB98]. Summer [Ano93c, USE90]. Summit [HDR03, HDR04, Ray98, BBE^{+}20]. Sun [Ano00j, Ano04a, Gal10, Kro99b, Sur04]. Sunk [Jon02]. SUPDUP [Sta78a]. Super [ZC95, Ano99b]. Supercomputer [Coc01a]. Supercomputers [Coc03, DDJ98a, DDJ98b, BBE^{+}20]. Superconducting [ZC95, BM22]. superoptimizer [GK92]. superscalar [UZ97]. SuperScreen [BM22]. supervised [SKB23]. supplementary [PKB17]. Supply [Har20]. Support [Ano00j, BOM97, Bee91a, Bee91b, Bra04, KMF^{+}07, Kro00, MSLH71, SZAB97, Bee17, Bla89a, Bla89b, BS05, Bro03, But95, Don04, Fra95, MWB89, RA16, Sin08, Sin10b, Wii91b, Yan90, Yan92, YWA07]. supported [Kli90]. Supporting [Han00, HOST05, PFL^{+}12, EKUR10, GKP^{+}14, KP93]. suppurativa [DSB^{+}16]. suppurativa/acne [DSB^{+}16]. surface [GBG^{+}16, WNS^{+}21]. Surgeons [Bar00a, Coc03]. Surgery [GIA^{+}06]. Suricata [WJM22]. Surveillance [BA15, Hol05]. Survey [BKR^{+}20, Con87, KKA^{+}19, TWS^{+}22, dBLMT11, BCG^{+}14, HNH03, IDSM23, Kri90, ZE00]. Survey-based [KKA^{+}19]. Surveyor [San78b, Sta78b]. Survivability [RT12]. survival [Coo91]. SUSE [Bau06b, Ano01j, RAH^{+}01]. Sustainability [GL14]. sustainable [dA15]. sustained [YLHW21]. SVGAlib [Lin00]. Swan [SGD00]. Swedish [Jön05c]. sweet [KHA^{+}03]. Swing [Hag04]. Swiss [Sur01b]. Switching [Ron05a]. SymbiFlow [MEB^{+}20]. Symbiodinium [HW17a]. Symbolic [Ano97d, Bra97, CCG^{+}02, Coo95b, Jef08, Lev95b, Lev95a, Mio90, Coo95a]. Symposium [ACM88, ACM93b, Ano93d, Ano94c, BBdD17, Cse99, Dig82, DGBH93, FP95, IEE90, IEE92b, IEE93, Jef08, Lev95b, Lev95a, Mio90, Ten93, USE00b, ACM94, Abr81, PT91]. SymPy [JČMG11]. synchronized [MSK05]. Synchrotron [SAC^{+}15]. Synergetic [Ano00k]. synergy [HPM^{+}08]. Syngress [Ano11, SD16]. Synopsis [Ano18]. Syntax [Kli90]. Syntax-directed [Kli90]. synthesis [CCA^{+}13]. Sys [Plo97]. Sysmon [SKB23]. System [ASWD18, Ano90a, Ano90c, Ano97c, Ano00e, Ano00i, Ano00j, Ano04b, ALA20, AHM^{+}07, Bro01, CTP^{+}22, Den13, DKMT11, G^{+}06, Gre80, GEMN07, Har94, Kro99b, Kro99a, McC02b, MB98, MS91, MSNS91, Mur94, PSSH16, PMG^{+}09, PKG^{+}10, PPG^{+}11, PBJ^{+}12, Shi03, TBPS15, TF21, VOK^{+}22, WLC01, WKA^{+}08, Zha16, ADF^{+}21, ABNÅ05, AAA^{+}12, And01, AAB^{+}05b, BGM99, Bea04, BCvE^{+}05, Big13, Bor88, Car89, CKH91, CK06b, CK06c, Cla90, Coc01a, Col05, Dig75a, Deo90, Don04, Eds16, GSW08, GPPT16, HLL^{+}95, JP09a, JK12, KGMI06, Kaw92, KN93, KW94, Kra05, KRR23, LR08, LQ17, LS04, Mac99, MBR21, MT94, MSM10, MWG^{+}91, MHP94, MQN19, Pag07, PL05, PH16, PH82, Phi12, QLC^{+}12, QSX^{+}15, RÓ01, RA16, SBS20, SP12, Sch04, SHN97, TMM^{+}13, TTL06, VD01]. system [WB07, Wii91a, Wil13, YM93, Ygg93, Ygg94, Ano01j, Pel89]. systematic [LC12a, TPSZ19, YT22]. Systeme [Gün02, Jor04, Cor00, Jor04]. Systems [Ano94b, Ano96c, Ano00f, Ano00i, Ano00j, Ano00k, BPG94, BSA14, CWM^{+}20, Coo95b, Cow03, CRW^{+}04, FK04, HWZ01, Hau01, IEE92a, IEE95a, ILG10, Jor04, LMZP19, MSC19, Maz15, MSW09, Mio90, Owe01, PG02, Pra03, Reh01b, SCSC04, Sha10, SVAGB20, TGC^{+}21, USE94, USE98b, ACM93a, Ahm08b, Ale92, AAB^{+}04, APK14a, APK14b, AMC16, Ano96d, AG22, BJWZ08, BM06, BSW95, BD03b, BLG^{+}17, BYV08, BH11, BMT^{+}20, BG12, BDP^{+}14, Bud10, CCA^{+}13, CJ19, CK06a, CFW01, Com99, Coo95a, CHA06, Dig82, DPH16, Don04, HAC^{+}23, HYA20, HZS^{+}16, Jae03, JCNS^{+}22, JWC18, JNN12, Joy08, Joy09b, JČMG11, KMG^{+}93, KTF15, KSS^{+}23, Lla06, LQR17, LLS11, MFB23, MSS95, MOT^{+}18, OMA^{+}22, RHW^{+}21, RCGB^{+}22, RAMB18, Sch91a, SRGCPB^{+}09, Sut02, TZH22, VGD^{+}97, VB19, WFF18, WGS07, YSC^{+}06, dA15]. systems [dlVRB21, Ano02b, Ano04a, Kro99b, Kro00, Kuk98, ZKDP22]. systems-on-chip [Don04]. Szeged [Cse99].

T [DKMB14, KMF^{+}07, PMG^{+}09, PKG^{+}10, PPG^{+}11, PBJ^{+}12, TBPS15, WKA^{+}08, Ano00l, HLS^{+}13a, HLS^{+}13b]. T.Rex [Ano00j]. T/TCP [Ano00l]. tables [Ano01h, Wil14]. tackle [Wol02]. Tactician [Ano96c]. Take [Ano93b]. Takes [XMGM21, XMGM22]. Taking [All02a, All02b, PM00]. Talarian [Kro99a]. tale [BH17]. TalentSoft [Ano97c]. Talk [Sta04a, TGS22]. Tangible [CGK^{+}02]. TapeWare [Ano00i]. Targeting [CDG97]. Task [SBDR22, DFU20]. task-based [DFU20]. Tasking [BOM97, MB98, Shi03, Ano01j]. Tasks [Plo97, SPDQ22]. Taxes [BES^{+}01]. Taxman [DDJ98a, DDJ98b]. taxonomy [May17]. TC [And03]. TCG [And03]. Tcl [Ass95, DF00, AG95, DDJ98a, Lor95, DDJ98b, ZK05]. Tcl/Tk [Ass95, AG95, Lor95, ZK05]. TCP [Ano00l, Str94, Ano00l]. TCP/IP [Str94]. TCPA [And03]. TDDFT [HAC^{+}23]. Teach [P^{+}99, BAE14]. Teaching [BM12, CV13, CSP09, DKMT11, GSW08, McA08, NN00, PPRB07, Col05, Rob11]. Team [Lus04, SMS16]. Teams [GS12]. tech [San01]. Technical [BHMB03, CRW^{+}04, MTM^{+}19, OMA^{+}22, PLO^{+}23, TKSC20, USE99, USE01b, USE02b, USE02c, YXS^{+}19, HSX^{+}18, MAMC05, Sin08, Sin10b, Tay19, USE98a, ZFD21]. Technique [Sta02c, FG92, YMCF23]. Techniques [LMZP19, PYM^{+}06, BPG94, GB06, GJS^{+}02, Ham07, Kra05, LF90, Qui00, TG15]. Technological [Men10, Waa09, Pya06]. Technologie [TNM17]. Technologies [Ano96c, Ano00i, Ano00j, Ano00k, CTP^{+}22, MS12, Wal01, BCHR12, Bor09, GP05, Kan12, Kro99a]. technologique [Cor00]. Technology [Ano88b, Ano96b, Ano97d, BMZ14, Bra97, Cho09, Cus04, DDJ98a, EXA^{+}05, Hoh01, Kro99b, Kro99a, Law09, DDJ98b, VGdlP01, Wat01, Cor05, GHM^{+}05, Kle21, LQ17, MTBS09, Sea04, Zic01]. Technology-Based [EXA^{+}05]. tecnologia [Cor05]. tecnologica [Zic01]. TECO [Bee86, Dig74, Dig75a, Dig80a, Dig80b, Dig82, Mur09, PH82, Uni77, Har77]. Tecplot [Ano96c]. Tekstredigering [Lan89]. Telecom [Ano00k, Kro00]. téléphonie [VSM06]. Telephony [PM00, Sug02a, Sug02b]. Tell [SV03]. Telnet [STS92]. Template [How98]. Templates [CWM^{+}20]. TENCON [Bao93]. tension [DB02]. Tensor [BSA22]. Tenth [ACM93b, IEE94b]. TenXpert [Ano96b]. TeraSpell [Laz98]. Termcap [Sta88d, Sta92c, Sta97c, Sta88d]. terminal [MYU89]. Test [Bar01, Kro99b, Lew99b, LMZP19, MCS12, WMLM22, WDK^{+}20, Ano01a, CMTA19, THG20, ZVvDD11]. Testament [PR96, WY94, W^{+}95]. Tester [Ano02b, Ano11, CB12]. Testing [AS97, Ano01j, BMZ14, DDJ99, EJS^{+}01, MZH22, ODP15, SL01, ZLL04, CMTA19, Kan12, LRP21, MGFRG12, She07, WN15]. Tests [MSSvK08]. tethering [BSP11]. Texas [IEE92c, IEE94b, IEE95b]. Texinfo [CS91, CS93, CS95, CS96, CS99, SC88b, Rob94d]. TEXlive’4 [Rod00]. TEXmacs [vdHGG^{+}13, vdH04, GvdHPR14]. Text [Abr81, CPJ^{+}98, Coh82, DF00, Fin91, Gre80, Kro00, Mud97, SBA92, Val93, BK91, Bk94, Dat85, Fin80a, Fin80b, FK90, GRJS01, HSX^{+}18, KB90, Raj13, Sch91a, Ude89, Dig75a, Dig75b]. Text-Oriented [Coh82, Sch91a]. Text-Processing [Gre80]. Textbook [Rad92]. textbooks [Ber22]. textual [Joh94b]. Textverarbeitung [Str94]. Textwerkzeuge [Gün02]. TGUI [DKMT11]. TH [DWP^{+}14, Zha16]. TH-A-19A-01 [DWP^{+}14]. TH-E-209-03 [Zha16]. Their [GB00, EKR91, JCNS^{+}22, Mah03, Neh07, WFF18, WGS07, ZWU22]. Them [Lev23]. theorem [LN92]. Theoretical [Boy13, MSLH71, Pya06]. Theory [Fin80a, Fin80b, RGCS14, CHA06, Fär05, FHH11, GCK^{+}17, HPT17, RAMB18, SHW^{+}21, YKSH20, SHW^{+}21]. theory/configuration [HPT17]. Therapy [Ano14, PMG^{+}09, PKG^{+}10, PPG^{+}11, PBJ^{+}12]. There [Bar00b, CPJ^{+}98, Mic04, Fie90b]. thermal [AH19]. thick [Sch09]. thin [GF17]. ThinAirApp [Ano01i]. Thing [CPJ^{+}98, Mic04, Fie90b]. Things [Gal10, Li18, MD18, TF21]. ThinkWrap [Sta96b]. Third [Ano87, Ano11, BPG94, FvdHJ10, IEE93]. those [Mud97]. Thought [Ano96c, CRW^{+}04, Mar01]. Thread [Ano00l]. Threads [Woo01, dlPRGB99]. Threads.h [Ano00i]. Threat [SSH22, Sta96b]. Threatens [EKJ^{+}03, Vai01]. threats [VSN22]. Three [ATM22, CKS16, Emb06, Kuk98, BTL^{+}11, EHP14, JJ91]. Three-dimensional [CKS16, EHP14]. Three-part [ATM22]. Throughput [LGW^{+}22]. Thursday [DMP^{+}02]. Tiered [DXT^{+}18]. tIGAr [Kam21]. TiGL [SKSM19]. Time [Ano01i, BFC02, FQYS23, McC99c, MFS15, PSR16, Reh01b, SSC^{+}00, YLL^{+}07, dlPRGB99, ACB18, CYOS19, Cur99, DVC^{+}07, GPPT16, GTMR23, HAC^{+}23, HZ14, HKvH16, HWM^{+}15, Hua17, Kam11, PSS^{+}07, Rui13, SBS20, Sta80a, Sta81d, Sta81c, Sta81b, TL17, VGD^{+}97, YSVM^{+}16, YSMA^{+}17]. Time-Based [MFS15]. time-dependent [YSVM^{+}16, YSMA^{+}17]. time-domain [HKvH16]. Time-Frequency [PSR16]. Time-Sensitive [FQYS23]. timely [QC18]. times [DRM21, Lin02b]. timescale [Mas05]. TIMESERIES [Ano97c]. Tinkering [Col09a]. Tiny [Bar00b, SG99]. Tips [Ste00a]. Tk [Ass95, AG95, Bea94, Lor95, ZK05]. TkPerl [Bea94]. TOC [ACM05]. today [WM01]. Together [OG07, ESM19]. toggles [PdSCJM22]. Tokyo [IEE94a]. tolerance [KTP95]. Tolerant [IEE90, IEE92a, IEE92b, IEE93, LQ17, Yad07]. Tom [SGD00]. tomography [CKS16, HFO^{+}12]. Too [RAH^{+}01]. Tool [Ano96e, Ano01i, Ano01j, Ano02b, Ano04b, Ber96, CCG^{+}02, GNR^{+}09, KMF^{+}07, Kim01b, Kro99b, Kro99a, Kro00, LOW91, Mam01, Sch03, Tan11a, Tan11b, UNF^{+}08, CCA^{+}13, DPH16, FL15, JCNS^{+}22, KOI94, LC12a, MSZ^{+}01, MGPB20, Mił10, DARJ23, PPR19, RJ21, SPLD20, SKB23, VSN22, WRSG92, You08]. ToolBox [LHZ12, GHH20, PSR16, Ano96d, AMR18, CV22, CKGW22, DSK19, GDK21, HSF^{+}15, JRA^{+}18, MBR21, PC13a, RZWW23, RAW^{+}16, RHR^{+}21, TACA15, FRAK15]. Toolkit [AG95, Ano96b, Ano96c, Ano00j, CB12, GIA^{+}06, HOL^{+}07, Kro00, Laz99, SFWD12, Wol03a, Wol03b, Bea94, HWM^{+}15, Hua17, Hua23, KP93, Qui00, Rac00, WFDK19, Kro99a, Ano11]. toolkits [FvH03, Kro99a]. Tools [Ano00j, Ano01i, Ano01j, Ano02b, Bar00b, Blo04, Cha01a, CSD^{+}05, DDJ99, DM15a, Ebe09, EGK^{+}02, GS00, Jor04, KP76, KP81, KC92, Kro99b, Kro99a, Kro00, Kuc06, Kuk98, Nas04, Omb20, PFL^{+}12, RMAM19, Rob94c, SCSC04, SSM^{+}07, SCDS15, SHN97, Sor06, SuS01, TRB22, Tot06, Wen00, Wol04, Wol03b, AVA^{+}16, Ano93a, BTL^{+}11, BM06, BGR89, CFW17, DM15b, EM93, Gv14, GGH05, HLL^{+}95, HBR19, HFO^{+}12, HL02, IDSM23, JP09b, JLL23, KTF15, Koc09, Kor11, MM10, MLWR18, Rac00, Sch90b, SSS^{+}14, THG23, Twi04, Vir05, WCS20, ZLL04]. Tools.h [Ano00i]. Toolset [Ano01i, Dol91, ZK05]. Top [Hae02, MG12, OSM94a, OSM94b]. Topic [JDB09]. topological [WZS^{+}18]. TOPS [Sta81c]. TOPS-20 [Sta81c]. Toronto [Ass95]. Torvalds [Flo94, Li94]. total [HMYH22]. total-reflection [HMYH22]. touch [ACM93a]. Toulouse [IEE93]. Tour [Mir03]. TowerEiffel [Ano96c]. Towers [ACM89]. Tozzi [Pri19]. TPC [Lla06]. TPC-C [Lla06]. TPCC [Lla06]. TPCC-UVa [Lla06]. TPU [Smi90]. Trace [MZG14, KSK09]. traceability [BG12, HZ14]. Traces [LZWH22, Bow05]. tracing [JLL23]. Track [USE98a, USE01b, USE02b, USE02c, Sta04a]. tracking [GTMR23, Joh94a, RDZ20]. Traction [Wea03]. trade [CFMRL11, Gil04]. trade-off [CFMRL11]. Trademark [Fal03]. Trademarks [III01, Gil04]. Traditional [LC12b, Wil13]. traffic [ACW04]. Training [EKJ^{+}03, GB00]. transaction [QB21]. Transactionalizing [RVLS14]. Transactions [Ano00l, Ano04c, Fra95]. TransactNet [Ano96b]. Transfer [BMR^{+}23, BG95, NGCI^{+}12]. Transform [Wut12]. transformation [GFZ16, QLC^{+}12]. Transforms [PSR16]. Transistors [Bar00b]. Transitioned [KKN^{+}21]. Translation [SS02]. translator [Lie92, Smy97]. Transparency [Hol23, PMBM^{+}15, Mur20]. Transparent [GGL21, IC23, PMBM^{+}15]. Transport [RLTD23, KPK^{+}17, KGW^{+}21, MBTB21]. Trap [May06, Sta04b, YSVM^{+}16]. traveling [DC23]. Treatment [BJJ14, DWP^{+}14, GNR^{+}09, PBJ^{+}12, TBPS15, Gen99]. Tree [AL07, Ano96b, Car89, CLS95, Ham90, Mer03, Nov03, Nov04]. tree-based [Car89]. Treemacs [Ham88, Ham90]. Trees [Sid04, Wen90]. Treibers [Bud10]. Trend [WM19]. Trends [Ahm08a, Edw98, WCS20]. trhepd [HMYH22]. TRI [ACM97]. TRI-Ada’97 [ACM97]. triage [JCNS^{+}22]. Tribute [Boy00]. Tricks [EJS^{+}01, Lus04, Ste00a, Rob11]. Tridia [Ano00i]. Trillian [Ano00h]. trio [CM06]. TRIPs [May06]. Triumph [Kim01a]. Triumphant [Mog99]. Trivial [CASA22]. Troff [Ano10]. Trondheim [AK95]. True [CAC09]. Truly [Mud97]. trunk [Jør01]. truss [RP08]. truss-type [RP08]. Trust [Bel22, Gal01, RNR17]. Trusted [And03, SZ05, SS05a]. Trustworthiness [dBLMT11]. Try [EKJ^{+}03]. Trying [CSP^{+}03]. Tucson [IEE05]. Tuesday [DMP^{+}02]. Tuning [CZ22, UZ97, FKM^{+}11]. turbines [RH21]. turbulent [CFCA13a, CFCA13b]. Turing [Bar00b]. Turkey [NRG^{+}99, YA11]. turnkey [SKB23]. turns [Ano01h]. TurnSafe [Ano00j]. Tutorial [Ell12, Kot90, SF15, Wel94b]. Tux [EKJ^{+}03]. Twelfth [USE98b]. TWENEX [Sta80a, Sta81d]. Twenty [IEE92b, IEE93, MS91, MSNS91, RSAT19]. Twenty-Fifth [MSNS91]. Twenty-Fourth [MS91]. Twenty-second [IEE92b]. Twenty-third [IEE93]. Twin [Abb12]. Twin-float [Abb12]. Twitter [WKS^{+}14]. Two [Ahm08a, BY14, Bec93, BE06, Gla08, MFH02, Pra03, SG99, Sta03a, BTL^{+}11, BM22, CGZ17, CAWK22, KT05, KL07, KSV16, MFB23, MSR09, Sta96b, Rus09]. two-dimensional [BM22, MFB23]. Two-Guys-in-a-Garage [Pra03]. two-phase [CAWK22]. two-way [KSV16]. twoWayGPBEFoam [LMHL20]. TX [ACM00]. Tyne [IEE90]. Type [MRGP20, BR95, RP08, TL17, WFV14, KK94]. Typed [HOL^{+}07, Sid04, Car89]. types [Kir12, Lal91]. Typesetting [Ano10].

Typography [Fur90].

U.S. [Bar00a, Bol02, CGB^{+}05, Hol23, MSZ02, PM00, Sca19]. Ubuntu [Cha13, Gre11a, Gre11b, ORC06, Teo13]. UCITA [Sta00a]. Uganda [Muw09]. UK [BBdD17, Ano14]. ultrasound [WHJ15]. Umfeld [EW01]. Umgang [DF00, Sie04, Ste00a]. Umgebung [CK06a, CK06c, CK06d, CK06e, CK06f, CK06g, CK06h]. UML [Coc01b]. Umple [LFB^{+}21]. Umsteiger [Ron05a]. unabhängige [Lin02b]. unauthorized [Ros00]. unbound [Hah94]. uncertainty [FL15]. Uncompromising [Bra04]. Uncovering [BKR^{+}20, VSN22]. Uncrackable [Ano03e]. undergraduate [DDA^{+}07]. Underpin [Ell12]. understand [AMWH19]. Understanding [ABC18, DRM21, GBICMR13, Jan08, KCAS23, Mor11, RMAM19, San08, SFF^{+}06, Sca06, SSA08, St.04, Win95, BG12, NXC13, ZFY^{+}19]. UNDI [BFI^{+}21]. undo [Yan90, Yan92]. Unearthing [SSS22]. UnForm [Ano00k]. UniCC [Mey18]. Unicode [Uni01, DM97, Mud97, Noj01]. Unified [Bro01, FTZ^{+}23]. unipolar [WSK^{+}22]. UniPress [Uni85e, Uni85c, Uni85d, Uni85f, Uni86]. Unique [Lev23]. unit [THG20]. United [AT92, BH07, DPL^{+}91, Dre94]. Units [BY14, Gre14]. Universal [HS15, Mey18, NRRS20, RA16]. universities [Fel93]. University [BSK87, MSZ02, Smy97, SM89b]. UNIX [Ano00i, DF00, Str94, AL92, Ano93b, BGR89, Bor88, Com84, CCA84, Coc01b, Coc03, Cor05, FY18, Far92, Gan95, Gos81, HTU96, Hah94, Hen92, JJ91, KP84, Lew88, Lio96, Mit84, Mor96, PDG^{+}87, Pal87, Pom04, RB92, Sal94, SBA92, SHN97, Uni85d, Ano03d, Fie90a, Fie90b, Fri97, Gil88, Har94, PDG^{+}88, Rob94a, Sor01, WM01]. Unix-Grundlagen [Str94]. Unleashed [NR03, Teo13]. Unreasonable [Ros01d]. unrolling [Cha92]. Unsinn [Fal03]. unsplit [LH22]. Untriviality [CASA22]. Unveiling [WFW^{+}20]. Unveils [Ano02b]. Update [Ano95d, Ano08c, Liu06, And08, XOTI22]. updated [Lea92, LMOS93, Sta92b]. Updates [Aki16, APK14b, Ano01i, Ano01j, Ano04b, GJMPAM^{+}14, HLS^{+}13b, QSX^{+}15, SMRM^{+}17, Yes12]. Upgrade [Ano96b, Dan11]. uphill [ES23]. upon [IEE90]. Urheber [Oms03]. Urheber- [Oms03]. Urheberrecht [Geh96, Sur01b]. urheberrechtliche [NO03]. urheberrechtlicher [Stö04]. URLLC [LLEL^{+}23]. USA [TG15, ACM00, FMA02, Kap92, USE90, USE94, USE99, USE00a, USE01b, USE02b, USE02c]. Usability [NT06, Raj23, WCG22, Fri06, HPM^{+}08, ZDM10]. Usage [Bla06, DLT^{+}23, GO99, BDP^{+}14, RHR^{+}21]. USB [Kro00]. Use [Bol02, Fri97, GA04b, GM05, HW17a, MGM^{+}02, PMM17, PMM18, PKP05, SA15, SG05, Tra95, Eig03, HBR19, KK94, MV05, Ped05, PL05, Ron05b, Sie99, Sil13, Sin08, VD01, YLG05, Yes12]. Used [CWM^{+}20, Ell12]. useful [Bec93]. Usenet [Coc01a]. Usenix [USE99, Ano88c, Ano90b, Ano90c, Ano90d, Ano92, Ano93c, Ano93d, Ano94c, Coc01a, Den99, The04]. User [ACM88, ABB^{+}92, ABB^{+}95, ABB^{+}99, Anoxx, CW15a, CW15b, DMBS79, DF00, EKJ^{+}03, UCLxx, Lea92, LO89, LO92, MMR95, Mey18, MC91, Raj23, Rob96, Ron05a, Sch90b, SMNF88, Dig74, Dig80a, Dig80b, Dat85, Fie90b, For12, FvH03, GIM07, Gos84, GW10, JZ09, KN93, KK99, Li91, Pal87, PDG^{+}88, PH82, SLC88, Uni85e, Uni85c, Uni86, ZDM10, BAR16b]. User-Agent [EKJ^{+}03]. user-based [ZDM10]. user-centered [For12]. User-Controlled [CW15a, CW15b]. user-developed [Fie90b]. User-friendly [Sch90b]. Users [Ano04b, BV87, McC02a, Ron05a, Sta80b, Veg06, Com84, CCA84, Lew88, Sta80a, Sta81d, Sta81c, Sta81b]. uses [Car89, Gom99]. USGS [PH82]. Using [Adk11, ALGE12, AG95, Ano00e, Ano00d, Bak20, BY14, Big13, BAE14, BSA14, Col05, DS00, DS02, DM15a, DM15b, FP94, GKL^{+}14, Gol06, Gui00, HETD09, Kos21, Liu08, LGW^{+}22, LBF^{+}22, MAMC05, MTM^{+}19, MSS95, Mit94, MOMM11, PG02, RE04, Sch03, SCDS15, Sor06, Spi11, Sta88e, SPG92, Sta92b, Sta99, Sta00b, Sta00c, Sta03b, Tan11b, TBPS15, TRB22, VMKB05, WCHRM21, Wal01, Wii91b, WM05, von88, Aji17, ACB18, ASAAM^{+}19, Amb15, AJ05, ASC^{+}21, BCHR12, BSK^{+}15, Bow05, BG12, Bud10, CKS16, CSP09, DSB^{+}16, Edd96, Eds16, FHH11, Fri16, GK92, GM94, GB06, GV16, HFO^{+}12, HC07, HSX^{+}18, JK12, Joy09a, Kam21, Koc09, KFYI13, KG20, eLAA^{+}23, LW03, LGA20, MWG^{+}90, MGR16, MLWR18, NN00, NMG11, Noj01, Och09, OK94, ODP15, PSSH16, PPR19, RÓ01]. using [RVLS14, SBM^{+}10, She07, SCR05, WHJ15, WMLM22, Wen90, WKB14, WFV14, XFS^{+}22, Yad07, ZAC^{+}23, ZLL04]. uso [VD01]. USPTO [GM05]. USRP [ZPH^{+}15]. usw [Ano01c]. Utah [SC00]. Utilities [Coc01b, JJ91]. Utility [Kro00, BCR^{+}08, Fri97, Rac06]. utilizando [RÓ01]. utilization [Amb15, KK17, SM08]. utilizing [BTL^{+}11, HK95]. UVa [Lla06]. üzemeltetése [Lás05].

v [CGK^{+}02, Sta02a, DXT^{+}18, GYW^{+}23, MMD^{+}22, PGW^{+}20, SM89b]. v.7.1 [Kuk98]. v0.7 [Hua17]. v1.0 [Ano00i, Laz99, PSP^{+}22]. v1.8 [Arc94]. v2107 [GDK21]. v3 [Car04]. v4.0 [Ano00k]. VA [ACM93b, ACM94, Kro99d]. valid [THG20]. Validated [Cse99]. Validation [Ano02b, AML^{+}10, BZB17, JK12, KSK09, DIK^{+}23, GFS05, eLAA^{+}23, LA10, LLEL^{+}23, DARJ23]. validity [Höp04]. Valley [Sta04a]. Valuable [PM00]. Value [MCS12, Ude97, Far23, Fra13, Sim05]. Valued [WCG22]. values [KT05]. VanillaSearch [Ano96c]. variability [HZS^{+}16]. Variable [FL16, Che95, MMY^{+}19]. Variable-Precision [FL16]. Variants [Mor96]. variation [VSGM14]. variational [MMY^{+}19, XOTI22]. various [DC23, Fri97, SHW^{+}21]. VARs [Tay00]. VArStation [Kro99d]. VAX [Dig80b]. VAX/VMS [Dig80b]. VDM [PT91]. vector [SAHP15]. VELAS [RZWW23]. Vendor [Rie20]. Vendors [BB02, Kim01b, MV05, RNR17]. Venture [Got05]. Ventures [PRRL12]. VERI [CYL^{+}23]. Verifiable [dCdCM14]. Verification [Ano01i, DIK^{+}23, JPOB20, ABF^{+}14, BCHR12, BP14, Kan12, LQR17]. verified [MRH23, ZK21]. verify [JH16]. Verilog [Kro99b, WB02]. Verlages [Oms03]. versatile [FG16]. Version [Ano96a, Ano98, Ano01b, AFZ17, ATCZ19, Bol02, DS99, DS00, EKJ^{+}03, Kot90, Kro99c, Kuk98, LLSt99, LLSt00, LSM^{+}99, LSM^{+}00, NRG^{+}99, Nic93, Per05, SSC^{+}00, Sim00, SSP17, SSP18, Sta97a, Sta99, SPS^{+}00, SM00a, SM00b, Sta00b, SM02, Ano00k, Ano01a, AFZ18, CS95, CS96, CS99, Che86, Che87b, Eat00, EHHH06, FLA^{+}16, G^{+}01, Gos83, Lea92, Lew88, LZ11b, LMOS93, LS04, Pax95, RCGB^{+}22, Sta92b, SMS04]. Versionen [DF00]. versions [DC23]. Verson [Ano97c]. versus [BES^{+}01, Bou05, DD17, GWT^{+}01, RM92, Shi12, Sin10b, dlPRGB99]. Vertragsrecht [Oms03]. Very [Ano94c]. vesicle [BSP11]. VHLL [Ano94c]. via [FvH03, JH16, MGYC18, QB21, SDL^{+}16]. viable [Ste08]. Vicious [NN16b, NN21]. Victoria [MG94]. video [FG16]. video/audio [FG16]. Videomodem [Ano00j]. Videophone [Man00]. videotape [SMNF88]. viele [Ste00a]. View [Ian02, Wei03]. Viewing [AM03]. Viewpoint [BB08, IC23, Raf23, San08, Spi21, Sta00a, Sta09]. Viewpoints [CK08, Stö04]. Views [Bar00b, Bar00c, Bar00a, Bar01, Coc01a, Coc01b, Coc03, DDJ98a, DDJ99, DDJ98b, Fei23]. vigilantism [Mau05]. vignetting [BTL^{+}11]. VIII [USE94]. Violations [MOMM11]. Violin [Fyk97]. Virkus [Kuc06]. Virtual [AAB^{+}05a, CDG97, DDJ99, GS12, LR11, LLWM23, MB98, Shi03, SCDS15, TB05, CK06a, CK06c, Gal01, WHJ15]. virtualization [G^{+}06]. virtualized [BJWZ08]. virtuellen [CK06a, CK06c, CK06d, CK06e, CK06f, CK06g, CK06h]. Virus [Ano00k]. Vis [Est06]. Vis-à-Vis [Est06]. Visible [CGK^{+}02]. Vision [BKP05, BSW95, McL92, MSS95]. Visual [Ano04d, Mon03, Ano01i, Ano01j]. visualisation [HBB^{+}12]. Visualization [Ano01j, Ano04b, SSM^{+}07, CSEP14, FHH11, Fri16, RZWW23, TL17, YA05, Aji17]. Visualizing [Joh94b]. visually [Phi12]. VLSI [IEE94c, QR92]. VM [G^{+}06, TB05]. VMEBus [Per02]. VMS [Dig80b, HBB^{+}12, Uni85f]. VMStools [HBB^{+}12]. Vocabulary [SBS20]. VOCAL [BK02]. Voice [KG01]. Voices [Ano99a, DOS99, Sea99]. VoIP [AML^{+}10, BK02]. Voltage [CCA^{+}19]. Volume [PBJ^{+}12, Bon93, LH22, Zag14]. Volunteer [KS11, BP14, MG12]. Volunteering [BKR^{+}20]. Volunteers [BSFR22]. vorbereiten [DF00]. Vorstellung [CK06b, CK06c, CK06e, CK06f, CK06g, CK06h]. Vote [CSD^{+}05]. Voter [CSD^{+}05]. Voting [CKB^{+}05, And08, PL05, ADF^{+}21]. VPR [MEB^{+}20]. VR [Wes00]. VRtuoso [Kuk98]. vs [Ano04c, CWB^{+}04, CMJ^{+}04, DPL^{+}91, Dwa04, For07]. VSP [FY18]. vulnerabilities [HM10, MD17, XTY^{+}22]. Vulnerability [CYL^{+}23, ACKT20, GYW^{+}23, KRR23]. vulnerable [PSL21]. VxWorks [PG02].

W [Ano04c]. Wacky [RAH^{+}01]. WADAS [ACM93b]. Wall [DDJ99]. WannierTools [WZS^{+}18]. WAP [CWB^{+}04]. War [Sta03a]. Wardialing [EKJ^{+}03]. Warfare [Cha07]. Warm [CK10]. warming [BB08]. warning [THG23]. warnings [MRS07]. Was [Kam14a, Kam14b, Mud97].

Washington [ACM93b, IEE89, IEE95a]. Wasted [KCAS23]. water [FHH11, ORS^{+}14, ODP15, DARJ23, ABC^{+}14]. Waterloo [ACM93a]. watershed [ORS^{+}14]. Watters [SD16]. Wave [Ano00i, PHT17, TL17, WGG16]. Wavelet [Kro00, PSR16]. waves [DBLF16]. Way [Bea21, CPJ^{+}98, DFT21, Gag02, Lus04, Ing92, KSV16]. Wayback [JCNS^{+}22]. Ways [BE06, JWC18]. WCL [Hen92]. WE-D-9A-06 [BVLF14]. weakened [NO03]. Weather [McC02b]. Web [Ano96b, Ano96c, Ano97c, Ano00d, MC91, TG15, Uni01, ACKT20, BMR^{+}23, Bor09, Chi93, EKUR10, HM10, NMX19, AMS03, Ano97d, Ano01i, Ano01j, Ano02b, Ano03e, Ano04b, Bra97, Coc01a, DDJ98a, DB02, EJS^{+}01, EKJ^{+}03, GSW08, GP05, Ham99, HBC^{+}05, Hau01, IAS16, Kro99b, KG01, LW03, MSW09, PM00, Per00, Sai01, DDJ98b, Ude97, Veg06, Wal99]. web-based [EKUR10, AMS03, Ano01j, GSW08]. Web-Enabled [KG01]. Web-mode [MC91, Chi93]. Web-Savvy [Kro99b]. web-scale [Bor09]. WebCompiler [Kro99b]. Weber [Gil06]. WebFountain [Ano03e]. WebSphere [AJ05]. Webtime [Ano98]. WebWork [WACBL03]. Weka [HBZ09]. WEP [Coc01b]. Werkzeuge [FG85]. wetland [MLMFN^{+}15]. where [Dew07]. Whether [Nag18, AMWH19]. Which [MSC19, WJM22, Car89, For07]. While [Bro19]. white [Ros00, Han00]. Who [ATM22, DWJG02, Lew97, Man92]. whom [JLH^{+}17]. WhyMP [MRH23]. WI [FMA02]. wichtigen [GGK99]. wichtigsten [CK06b, Gün02]. Wicked [Eub05]. Wide [DB02, BVT06, Bik96]. wide-area [BVT06]. widely [BM02]. Widgets [Tro96c]. WiFi [BCI^{+}09]. Wikipedia [Cap13]. Wild [DLT^{+}23]. Wiley [San01]. Will [CK08, Fly87b, HW17b, Ano00h]. Williams [Cas02]. wind [RH21]. Window [AG95, Ano90a, TGC^{+}21]. Windows [DF00, PKP02, PKP05, Rod00, Ano00h, Ano01j, CK06a, Gag02, HWZxx, HWZ01, PKP05, Rac00, STS92, Veg06]. Windows-Programmierung [PKP02, PKP05]. Windoze [CPJ^{+}98]. WINE [Gag02]. Wins [Bar00b, DiB04]. Winter [Ano90b]. WIP [MDRN18]. Wired [Coc01b]. Wireless [Ano01i, Ano01j, Ano02b, CWB^{+}04, Far06, Kuc06, SNF04, Vir05, DFCPSF15]. WISPER [Far06]. within [HMP^{+}15, HW17a, PPC^{+}15]. Without [EKJ^{+}03, Kos21, Mog03c, CH06a, Gre18]. withstand [Sta01a]. wizards [Ano01d, Neu84]. Wolfram [Ano00j]. Women [TGS22, TWS^{+}22]. Woods [Neu84]. Woodstock [Wol03a]. woody [Ano01c]. Woos [GAS^{+}01]. Word [Knu99a, Cra89]. Work [CGK^{+}02, HBGS19, Maj03, Car04, Geh96, Mah03, Wii91b]. Workbook [MMP^{+}22]. Worker [CDsJ^{+}00, CMJ^{+}04]. Workflow [AtHR11, Bak20, MPG^{+}16, TMM^{+}13]. workflows [KTTK17]. working [CFW17]. workload [VSGM14]. workloads [AVA^{+}16]. workplace [ACM93a]. Works [Hef97, Lus04, Väl04]. Workshop [AY93, Ano92, BPG94, FFvdH01, FFHL05, FFH^{+}05, IEE92a, QR92, ACM05, Ass95, Bik96, EKR91, HTU96, KJ03, DMP^{+}02]. workshops [BS14]. Workstation [Ano00i, Kro99d, Str94]. workstations [BGR89, Gad88].

World [AHB^{+}09, Bik96, Bon11, CC03, CPJ^{+}98, Fin91, GWT^{+}01, TG99, Cor05, Dvo04, Neu84, Phi12, Sal08, Sin10a, DB02]. Worlds [LR11, KHA^{+}03]. Worm [CWB^{+}04]. Worn [BY14]. Worries [SSC^{+}00]. Worth [HR11, THG23]. WOS [Ano01d]. WOS-Doku [Ano01d]. would [Dew07, Sta96b]. Woz [Bar00c]. Wristwatch [Man00]. Writers [DDJ99]. Writing [Chi97, Gli97, Hig93, Mam01, She01, Tro05, BGR89, KX86]. written [Lie92]. WSDL [Bar00c]. WSRP [YWA07].

X [Ano96c, Ano97c, Ano00i, Ste00a, AG95, Ano90a, Ber96, CG17, Fox08, Hoh01, MWG^{+}90, Tro96a, WFDK19, Ygg93, Ygg94]. X-Based [Ber96]. X-ray [CG17, WFDK19]. X-Window [AG95]. x86 [Ahm08b]. Xastir [MSM10]. XDesign [CG17]. XDoclet [WACBL03]. XEmacs [Aye01, Aye97]. Xen [Fox08, SJV^{+}05]. XFree [Ano01c]. XFree86 [DF00, SuS01, Hoh01]. XFree86/23.3.6 [DF00]. XII [USE98b]. XILINX [ZKCS91]. Xiran [Ano04b]. Xmgr [Vau96]. XML [Ano02b, Bar01, CBB06, GWT^{+}01, Kim01a, Kro00, Mam01, Noj01, Qui00, SSC^{+}00]. XML-based [Mam01]. XML-Oriented [Ano02b]. XMP [Kro99d]. XS [ZKCS91]. XS-XILINX [ZKCS91]. XtalComp [LZ12]. XtalOpt [AFZ17, AFZ18, ATCZ19, FLA^{+}16, LZ11a, LZ11b].

Y2K [SSC^{+}00]. YACC [DS88, DS99, DS00, DS02, Vol89, DS90]. YACC-compatible [DS88, DS99, DS00, DS02, DS90]. YADE [KGT22]. YAWL [AtHR11]. Year [Bes04, EKJ^{+}03, WB02, FRBRF19, GAS^{+}01]. Years [Ahm08a, Gri16, LRBM23, Par03, RSAT19, Pre16a, Pre16b, SHB^{+}20, Sta96b]. Yer [Far91]. Yggdrasil [AL07, Ygg94]. York [Smy97]. Yosemite [Ano00i]. Young [GAS^{+}01]. Yourself [P^{+}99, Sea04].

Z [Mor92, G^{+}06, TB05]. z/VM [G^{+}06, TB05]. z9 [AHM^{+}07]. Zapdos [DIK^{+}23]. Zeek [WJM22]. Zephyr [GJS^{+}02]. Zero [Ano04b]. zeros [SG99]. Zertifizierung [Mag04]. Zipf [MSSvK08]. ZKCM [Sai13]. Zondigo [Ano01j]. zSeries [BS05]. Zugriff [Per02]. zum [Mag04, Per02]. zur [FG85, Jor04, Oms03]. zusätzlich [Ano01c]. Zuse [BHP^{+}01].

References

Alsubaiee:2012:AOS

[AAA^{+}12] Sattam Alsubaiee, Yasser Altowim, Hotham Altwaijry, Alexander Behm, Vinayak Borkar, Yingyi Bu, Michael Carey, Raman Grover, Zachary Heilbron, Young-Seok Kim, Chen Li, Nicola Onose, Pouria Pirzadeh, Rares Vernica, and Jian Wen. ASTERIX: an open source system for “big Data’’ management and analysis (demo). Proceedings of the VLDB Endowment, 5(12):1898–1901, August 2012. CODEN ???? ISSN 2150-8097.

Alsubaiee:2014:ASO

[AAA^{+}14] Sattam Alsubaiee, Yasser Altowim, Hotham Altwaijry, Alexander Behm, Vinayak Borkar, Yingyi Bu, Michael Carey, Inci Cetindil, Madhusudan Cheelangi, Khurram Faraaz, Eugenia Gabrielova, Raman Grover, Zachary Heilbron, Young-Seok Kim, Chen Li, Guangqiang Li, Ji Mahn Ok, Nicola Onose, Pouria Pirzadeh, Vassilis Tsotras, Rares Vernica, Jian Wen, and Till Westmann. AsterixDB: a scalable, open source BDMS. Proceedings of the VLDB Endowment, 7(14):1905– 1916, October 2014. CODEN ???? ISSN 2150-8097.

Aliaga:2004:PGS

[AAB^{+}04] J. Aliaga, F. Almeida, J. M. Badia, S. Barrachina, V. Blanco, M. Castillo, U. Dorta, R. Mayo, E. S. Quintana, G. Quintana, C. Rodriguez, and F. de Sande. Parallelization of the GNU Scientific Library on heterogeneous systems. In Third International Workshop on Parallel and Distributed Computing. Third International Symposium on Algorithms, Models and Tools for Parallel Computing on Heterogeneous Networks. 5–7 July 2004, volume ??, pages 338–345. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. CODEN ???? ISSN ????

Alpern:2005:JRV

[AAB^{+}05a] B. Alpern, S. Augart, S. M. Blackburn, M. Butrico, A. Cocchi, P. Cheng, J. Dolby, S. Fink, D. Grove, M. Hind, K. S. McKinley, M. Mergen, J. E. B. Moss, T. Ngo, V. Sarkar, and M. Trapp. The Jikes Research Virtual Machine project: Building an open-source research community. IBM Systems Journal, 44(2):399–417, ???? 2005. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/
sj/442/alpern.pdf.

Appavoo:2005:EKO

[AAB^{+}05b] J. Appavoo, M. Auslander, M. Butrico, D. da Silva, O. Krieger, M. Mergen, M. Ostrowski, B. Rosenburg, R. W. Wisniewski, and J. Xenidis. Experiences with K42, an opensource, Linux-compatible, scalable operating-system kernel. IBM Systems Journal, 44(2):427–440, ???? 2005. CODEN

IBMSA7. ISSN 0018-8670. URL http://www.research.ibm.
com/journal/sj/442/appavoo.pdf.

Anderson:1992:LUG

[ABB^{+}92] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, S. Ostrouchov, and D. Sorensen. LAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-294-7. xv + 235 pp. LCCN QA76.73.F25 L36 1992.

Anderson:1995:LUG

[ABB^{+}95] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, S. Ostrouchov, and D. Sorensen. LAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 1995. ISBN 0-89871-345-5 (paperback). xix + 325 pp. LCCN QA76.73.F25 L36 1995.

Anderson:1999:LUG

[ABB^{+}99] E. Anderson, Z. Bai, C. Bischof, S. Blackford, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, and D. Sorensen. LAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, third edition, 1999. ISBN 0-89871-447-8. xxi + 407 pp. LCCN QA76.73.F25 L36 1999.

Abbott:2012:TFA

[Abb12] John Abbott. Twin-float arithmetic. Journal of Symbolic Computation, 47(5):536–551, May 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0747717111001970.

Ahalt:2014:WSS

[ABC^{+}14] Stan Ahalt, Larry Band, Laura Christopherson, Ray Idaszak, Chris Lenhardt, Barbara Minsker, Margaret Palmer, Mary Shelley, Michael Tiemann, and Ann Zimmerman. Water Science Software Institute: Agile and open source scientific software development. Computing in Science and Engineering, 16(3):18–26, May/June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Accioly:2018:USS

[ABC18] Paola Accioly, Paulo Borba, and Guilherme Cavalcanti. Understanding semi-structured merge conflict characteristics in open-source Java projects. Empirical Software Engineering, 23(4):2051–2085, August 2018. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://link.

springer.com/article/10.1007/s10664-017-9586-1.

Aberdour:2007:AQO

[Abe07] Mark Aberdour. Achieving quality in open source software. IEEE Software, 24(1):58–64, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Almeida:2014:COS

[ABF^{+}14] José Bacelar Almeida, Manuel Barbosa, Jean-Christophe Filliâtre, Jorge Sousa Pinto, and Bárbara Vieira. CAOVerif: an open-source deductive verification platform for cryptographic software implementations. Science of Computer Programming, 91 (part B)(??):216–233, October 1, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016764231200189X.

Alsberg:2005:GOS

[ABNÅ05] Bjørn K. Alsberg, Håvard Bjerke, Gunn M. Navestad, and Per-Olof Åstrand. GaussDal: An open source database management system for quantum chemical computations. Computer Physics Communications, 171(2):133–153, September 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002869.

Abrahams:1981:TMA

[Abr81] P. Abrahams, editor. Text manipulation: Proceedings of the ACM SIGPLAN/SIGOA symposium (Portland, OR, June 8– 10, 1981). ACM Press, New York, NY 10036, USA, 1981. ISBN 0-89791-043-5, 0-89791-050-8. LCCN QA76.7 .S54 v.16:6. Published as ACM SIGPLAN Notices, v. 16, no. 6, and ACM SIGOA newsletter, vol. 2, no. 1/2, spring/summer

Alaez:2016:TOS

[ACB^{+}16] Ricardo Marco Alaez, Jose M. Alcaraz Calero, Fatna Belqasmi, May El-Barachi, Mohamad Badra, and Omar Alfandi. Towards an open source architecture for multioperator LTE core networks. Journal of Network and Computer Applications, 75(??):101–109, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804516301916.

Akbarinasaji:2018:PBF

[ACB18] Shirin Akbarinasaji, Bora Caglayan, and Ayse Bener. Predicting bug-fixing time: a replication study using an open source software project. The Journal of Systems and Software, 136(??):173–186, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121217300365.

Anh:2012:CRR

[ACC^{+}12] Nguyen Duc Anh, Daniela S. Cruzes, Reidar Conradi, Martin Höst, and Xavier Franch. Collaborative resolution of requirements mismatches when adopting open source components. Lecture Notes in Computer Science, 7195:77–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-28714-5_7/.

Ayala:2011:FFA

[ACHC11] Claudia P. Ayala, Daniela Cruzes, Oyvind Hauge, and Reidar Conradi. Five facts on the adoption of open source software. IEEE Software, 28(2):95–99, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Amankwah:2020:ECC

[ACKT20] Richard Amankwah, Jinfu Chen, Patrick Kwaku Kudjo, and Dave Towey. An empirical comparison of commercial and open-source web vulnerability scanners. Software—Practice and Experience, 50(9):1842–1857, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

ACM:1988:PAS

[ACM88] ACM, editor. Proceedings of the ACM SIGGRAPH Symposium on User Interface Software, Banff, Alberta, Canada, October 17–19, 1988. ACM Press, New York, NY 10036, USA, 1988. ISBN 0-89791-283-7. LCCN QA76.9.U83A26 1988.

ACM:1989:CPS

[ACM89] ACM, editor. Conference proceedings / SIGDOC 89, November 8–10, 1989, Pittsburgh Hilton and Towers, Pittsburgh, Pennsylvania, SIGDOC — Conference Proceedings — 1989; 7th. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-337-X. LCCN QA 76.9 D6 S54 1989.

ACM:1992:PAC

[ACM92] ACM, editor. Proceedings of the 1992 ACM Conference on LISP and Functional Programming: papers presented at the conference, San Francisco, California, June 22–24, 1992. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-483-X. LCCN QA76.73.L23A26 1992. ACM order no. 552920.

ACM:1993:CPS

[ACM93a] ACM, editor. Conference proceedings: SIGDOC ’93, the 11th annual international conference, October 5-8, 1993, Waterloo, Ontario: getting in touch — staying in touch: moving online and multimedia systems into the workplace, SIGDOC 1993; 11th. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-630-1. LCCN QA76.9.D6I57 1993.

ACM:1993:WTA

[ACM93b] ACM, editor. WADAS ’93. Tenth Annual Washington Ada Symposium: Ada’s role in software engineering, June 28– July 1, 1993, McLean Hilton, McLean, VA: proceedings. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-609-3. LCCN QA76.73.A16W37 1993.

ACM:1994:AAA

[ACM94] ACM, editor. ADA in applications: 11th Annual ADA symposium — June 1994, McLean, VA. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-684-0. LCCN ????

ACM:1995:PAS

[ACM95] ACM, editor. Proceedings of the 33rd Annual Southeast Conference. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-747-2. LCCN ????

ACM:1997:PTA

[ACM97] ACM, editor. Proceedings of the TRI-Ada’97 Conference, November 9–13, 1997, St. Louis, MO. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-981-5. LCCN QA 76.73

A35 T75 1997. URL http://www.acm.org/pubs/contents/

proceedings/ada/269629/. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

ACM:2000:SHP

[ACM00] ACM, editor. SC2000: High Performance Networking and Computing. Dallas Convention Center, Dallas, TX, USA, November 4–10, 2000. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN

???? LCCN QA76.88. URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Staff:2005:FTL

[ACM05] ACM SIGSOFT Software Engineering Notes Staff. Frontmatter (TOC, letters, open source software (OSS) patent search engine, calendar of events, workshop and conference information). ACM SIGSOFT Software Engineering Notes, 30(2): 0, March 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Allaert:2004:EAT

[ACW04] Gaetan Allaert, Dirk Craeynest, and Philippe Waroquiers. European air traffic flow management: porting a large application to GNU/Linux. ACM SIGADA Ada Letters, 24(1):29– 37, March 2004. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Awazu:2004:BCO

[AD04] Yukika Awazu and Kevin C. Desouza. Brief communication: Open knowledge management: Lessons from the open source revolution. Journal of the American Society for Information Science and Technology: JASIST, 55(11):1016–1019, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agate:2021:SSO

[ADF^{+}21] Vincenzo Agate, Alessandra De Paola, Pierluca Ferraro, Giuseppe Lo Re, and Marco Morana. SecureBallot: a secure open source e-Voting system. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001776.

Adkins:2011:UGM

[Adk11] Lee C. Adkins. Using gretl for Monte Carlo experiments. Journal of Applied Econometrics, 26(5):880–885, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Adler:2000:OSO

[Adl00] Stephen Adler. Open source/open science 1999. Linux Journal, 70:??, February 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Abelson:1981:CNC

[AFS81] H. Abelson, R. M. Fano, and G. J. Sussman. Course notes CS 6.001: Structure and interpretation of computer programs, (chapter 1+2, guide to DEC-20, EMACS-intr., intr. to SCHEME, SCHEME-manual, chapter 3–7, problemsets 1– 9, quiz 1+2, final exam.). Report, Massachusetts Institute of Technology, Electrical Engineering and Computer Science Department, Cambridge, MA, USA, 1981.

Abelson:1982:CNC

[AFS82] H. Abelson, R. M. Fano, and G. J. Sussman. Course notes CS 6.001: Structure and interpretation of computer programs, (chapters 1-8, guide to DEC-20, intro. to EMACS, intro. to SCHEME). Report, Massachusetts Institute of Technology, Electrical Engineering and Computer Science Department, Cambridge, MA, USA, 1982.

Avery:2017:XNV

[AFZ17] Patrick Avery, Zackary Falls, and Eva Zurek. XtalOpt Version r10: an open-source evolutionary algorithm for crystal structure prediction. Computer Physics Communications, 217(??):210–211, August 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465517301005.

Avery:2018:XNV

[AFZ18] Patrick Avery, Zackary Falls, and Eva Zurek. XtalOpt version r11: an open-source evolutionary algorithm for crystal structure prediction. Computer Physics Communications, 222(??):418–419, January 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465517302989.

Alexander:1995:HCX

[AG95] P. Alexander and L. F. Gladden. How to create an X-window interface to Gnuplot and Fortran programs using the Tcl/Tk toolkit. Computers in physics, 9(1):57–??, 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Asad:2022:DAD

[AG22] Hafizul Asad and Ilir Gashi. Dynamical analysis of diversity in rule-based open source network intrusion detection systems. Empirical Software Engineering, 27(1):??, January 2022. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616

(electronic). URL https://link.springer.com/article/
10.1007/s10664-021-10046-w.

Ahmmed:2019:OSC

[AH19] Mohammad Shakil Ahmmed and Nazmul Huda. An opensource CFD model for computing thermal effect in the context of laser-induced semiconductor processing in photovoltaic applications. Journal of Computational Science, 34:55–65, May 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1877750318311190.

Anokwa:2009:OSD

[AHB^{+}09] Yaw Anokwa, Carl Hartung, Waylon Brunette, Gaetano Borriello, and Adam Lerer. Open source data collection in the developing world. Computer, 42(10):97–99, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alves:1994:CGA

[AHG94] J. Alves, M. Held, and M. Glesner. A code generator for an application specific pipelined processor. In Yuksel [Yuk94], pages 306–308 (vol. 1). ISBN 0-7803-1772-6. LCCN TK5101.A1M38 1994. Three volumes. IEEE Catalog No. 94CH3388-6.

Axnix:2007:OSD

[AHM^{+}07] C. Axnix, T. Hendel, M. Mueller, A. Nuñez Mencias, H. Penner, and S. Usenbinz. Open-standard development environment for IBM System z9 host firmware. IBM Journal of Research and Development, 51(1/2):195–??, January /March 2007. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-

8556 (electronic). URL http://www.research.ibm.com/
journal/rd/511/axnix.html.

Ahmad:2008:ATT

[Ahm08a] David Ahmad. Attack trends: Two years of broken crypto: Debian’s dress rehearsal for a global PKI compromise. IEEE Security & Privacy, 6(5):70–73, September/October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ahmad:2008:MCC

[Ahm08b] Feisal S. Ahmad. MINIX 3 C compiler performance: Comparing the Amsterdam Compiler Kit to the GNU Compiler Collection on x86 systems. Bachelor’s thesis, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, June 9, 2008. 27 +

26 (appendix) pp. URL http://www.minix3.org/theses/
ahmad-cc-performance.pdf;   http://www.minix3.org/
theses/ahmad-cc-performance_appendix.pdf.

Anand:2005:MPC

[AJ05] V. K. Anand and W. C. Jamison. A middleware performance characterization of Linux using IBM WebSphere Application Server. IBM Systems Journal, 44(2):353–367, ????

2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/sj/442/anand.pdf.

Ajiferuke:2017:BRS

[Aji17] Isola Ajiferuke. Book review: Statistics for library and information services: a primer for using open source R software for accessibility and visualization. Alon Friedman. Lanham, MD: Rowman & Littlefield Publishers, 2015. 376 pp. $95.00 (hardback). (ISBN 978-1-4422-4992-9). Journal of the Association for Information Science and Technology, 68(6):1594– 1595, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Almeida:2018:OSM

[AJLM18] Hayda Almeida, Ludovic Jean-Louis, and Marie-Jean Meurs. An open source and modular search engine for biomedical literature retrieval. Computational Intelligence, 34(1):200–218, February 2018. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic).

Aamodt:1995:SSC

[AK95] A. Aamodt and J. Komorowski, editors. SCAI ’95: 5th Scandinavian conference on artificial intelligence — May 1995, Trondheim, Norway, Scandinavian Conference on Artificial Intelligence 1995; 5th. IOS Press, Amsterdam, The Netherlands, 1995. ISBN 90-5199-221-1, 4-274-90046-0 (Ohmsha). ISSN 0922-6389 (print), 1879-8314 (electronic). LCCN Q334 .S3 1995.

Anzt:2021:CPR

[AKF21] Hartwig Anzt, Eileen Kuehn, and Goran Flegar. Crediting pull requests to open source research software as an academic contribution. Journal of Computational Science, 49: ??, February 2021. CODEN ???? ISSN 1877-7503 (print),

1877-7511 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1877750320305743.

Adams:2016:ESI

[AKHG16] Bram Adams, Ryan Kavanagh, Ahmed E. Hassan, and Daniel M. German. An empirical study of integration activities in distributions of open source software. Empirical Software Engineering, 21(3):960–1001, June 2016. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10664-015-9371-y.

Akimov:2016:SNU

[Aki16] Alexey V. Akimov. Software news and updates: Libra: an open-source “methodology discovery” library for quantum and classical dynamics simulations. Journal of Computational Chemistry, 37(17):1626–1649, June 30, 2016. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Assuncao:2023:HDM

[AKMS23] Wesley K. G. Assunção, Jacob Krüger, Sébastien Mosser, and Sofiane Selaoui. How do microservices evolve? An empirical analysis of changes in open-source microservice repositories. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121223001838.

Aho:1988:APL

[AKW88] Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger. The AWK Programming Language. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-07981-X. x + 210 pp. LCCN QA76.73.A95 A35 1988. See also [GAW87b, Mor87, Cor87].

Abrahams:1992:UI

[AL92] Paul W. Abrahams and Bruce R. Larson. UNIX for the Impatient. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-55703-7. xxvii + 559 pp. LCCN QA76.76.O63 A27 1992. Excellent, and thorough, coverage of UNIX, with chapters on the file system, utilities, shells, editors, Emacs, data manipulation, mail, network communications and resources, the X Window System, and a comparison of MS-DOS and UNIX.

Andersson:2007:MYS

[AL07] Samuel A. Andersson and Jens Lagergren. Motif Yggdrasil: Sampling sequence motifs from a tree mixture model. Journal of Computational Biology, 14(5):682–697, June 2007. CO-DEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.liebertpub.com/doi/abs/10.
1089/cmb.2007.R010;   https://www.liebertpub.com/doi/
pdf/10.1089/cmb.2007.R010.

Avramidis:2020:SOS

[ALA20] Eleftherios Avramidis, Marta Lalik, and Ozgur E. Akman. SODECL: an open-source library for calculating multiple orbits of a system of stochastic differential equations in parallel. ACM Transactions on Mathematical Software, 46(3):24:1– 24:21, September 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://dl.acm.org/

doi/10.1145/3385076.

Aleksander:1992:CCN

[Ale92] I. Aleksander. Capturing consciousness in neural systems. In Aleksander and Taylor [AT92], pages 17–22 (vol. 1). ISBN 0-444-89488-8. LCCN QA76.87.I56 1992. Two volumes.

Alfonsi:2005:OSC

[Alf05] Benjamin Alfonsi. Open source in the classroom. IEEE Distributed Systems Online, 6(6):??, June 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/csdl.computer.org/comp/mags/ds/2005/06/o6003.pdf.

Aldea:2012:USC

[ALGE12] Sergio Aldea, Diego R. Llanos, and Arturo González-Escribano. Using SPEC CPU2006 to evaluate the sequential and parallel code generated by commercial and open-source compilers. The Journal of Supercomputing, 59(1):486–498, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0920-8542&volume=
59&issue=1&spage=486.

Allman:2002:TOSa

[All02a] Eric Allman. Taking an open source project to market. In USENIX [USE02c], page ?? ISBN 1-880446-00-6. LCCN

QA76.8.U65 U84 2002. URL http://www.usenix.org/
publications/library/proceedings/usenix02/tech/techonefile.

html. Unpublished invited talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Allman:2002:TOSb

[All02b] Eric Allman. Taking an open source project to market: a parable of sendmail. In USENIX [USE02a], pages viii + 151. ISBN 1-880446-02-2. LCCN QA76.76.O63

B736 2002. URL http://www.usenix.org/publications/
library/proceedings/bsdcon02/tech.html. Unpublished

invited talk, BSDCON2002: Growing the BSD Community, February 11–14, 2002, Cathedral Hill Hotel, San Francisco, CA.

Altalhi:2017:ECO

[ALVV17] Abdulrahman H. Altalhi, J. M. Luna, M. A. Vallejo, and S. Ventura. Evaluation and comparison of open source software suites for data mining and knowledge discovery. Wiley Interdisciplinary Reviews: Data Mining and Knowledge Discovery, 7(3):??, 2017. ISSN 1942-4787 (print), 1942-4795 (electronic).

Allman:2003:EVO

[AM03] Eric Allman and Marshall Kirk McKusick. From the Editors: Viewing open source with an open mind. ACM Queue: Tomorrow’s Computing Today, 1(5):6–7, July/August 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2004:EOS

[AM04] Eric Allman and Marshall Kirk McKusick. From the Editors: Open source revisited. ACM Queue: Tomorrow’s Computing Today, 2(3):8–9, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Alves:2018:OIC

[AM18] Thiago Alves and Thomas Morris. OpenPLC: an IEC 61,131-3 compliant open source industrial controller for cyber security research. Computers & Security, 78(??):364–379, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818305388.

Amber:2015:CUL

[Amb15] Kyle T. Amber. Considerations for the utilization of ‘comparative analysis of colorimetric staining in skin using opensource software’ in an experimental setting. Experimental Dermatology, 24(9):717–718, 2015. ISSN 0906-6705 (print), 1600-0625 (electronic).

Alnaeli:2016:EEP

[AMC16] Saleh M. Alnaeli, Jonathan I. Maletic, and Michael L. Collard. An empirical examination of the prevalence of inhibitors to the parallelizability of open source software systems. Empirical Software Engineering, 21(3):1272–1301, June 2016. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10664-015-9385-5.

Apostolakos:2010:DIV

[AML^{+}10] Spyros Apostolakos, Apostolos Meliones, George Lykakis, Emmanuel Touloupis, and Vassilis Vlagoulis. Design, implementation and validation of an open source IP-PBX/ VoIP gateway multi-core SoC. International Journal of Parallel Programming, 38(3–4):288–302, June 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7458&volume=38&issue=3&spage=288.

Adewumi:2019:FFO

[AMOS19] Adewole Adewumi, Sanjay Misra, Nicholas Omoregbe, and Luis Fernandez Sanz. FOSSES: Framework for open-source software evaluation and selection. Software—Practice and Experience, 49(5):780–812, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Azzini:2018:DMP

[AMR18] Ivano Azzini, Ronal Muresano, and Marco Ratto. Dragonfly: a multi-platform parallel toolbox for MATLAB/Octave. Computer Languages, Systems and Structures, 52(??):21–42, June 2018. CODEN ???? ISSN 1477-8424 (print), 1873-

6866 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1477842417300660.

Anderson:2003:BWB

[AMS03] Geoffrey Anderson and Rafael Moreno-Sanchez. Building Web-based spatial information solutions around open specifications and open source software. Transactions in GIS, 7 (4):447–466, 2003. ISSN 1361-1682 (print), 1467-9671 (electronic).

Almeida:2019:IWH

[AMWH19] Daniel A. Almeida, Gail C. Murphy, Greg Wilson, and Michael Hoye. Investigating whether and how software developers understand open source software licensing. Empirical Software Engineering, 24(1):211–239, February 2019. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616

(electronic). URL http://link.springer.com/article/
10.1007/s10664-018-9614-9.

Anandakrishnan:1999:PEG

[Ana99] Sridhar Anandakrishnan. Penguins everywhere: GNU/ Linux in Antarctica. IEEE Software, 16(6):90–96, November/December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://computer.org/

software/so1999/s6090abs.htm;   http://dlib.computer.
org/so/books/so1999/pdf/s6090.pdf.

Anderson:2001:FOS

[And01] Annelise Anderson. FreeBSD: an open-source operating system for your personal computer. Bit Tree Press, Portola Valley, CA, USA, 2001. ISBN 0-9712045-1-9. xx + 423 pp. LCCN

QA76.76.O63 A49 2001. URL http://www.bittreepress.
com/FreeBSD/introbook/. Includes CD-ROM.

Anderson:2003:TCF

[And03] Ross Anderson. ‘trusted computing’ frequently asked questions — TC / TCG / LaGrande / NGSCB / Longhorn / Palladium / TCPA. World-Wide Web document., August 2003.

URL http://www.cl.cam.ac.uk/~rja14/tcpa-faq.html.

Anderson:2008:OSV

[And08] M. Anderson. Open-source voting — [update]. IEEE Spectrum, 45(10):13–14, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2011:MGD

[And11] Nathanael Anderson. Multiplatform GNU development. Linux Journal, 2011(209):1:1–1:??, September 2011. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anel:2011:IRC

[Añe11] Juan A. Añel. The importance of reviewing the code. Communications of the ACM, 54(5):40–41, May 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Angryk:2001:BRM

[Ang01] Rafal Angryk. Book review: MICO, An Open Source CORBA Implementation. PDCP: Parallel and Distributed Computing Practices, 4(3):??, September 2001. CODEN ???? ISSN 1097-2803.

Anonymous:1986:FIP

[Ano86] Anonymous. Free IBM-PC encryption software. Cryptologia, 10(4):224, October 1986. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). URL http://www.

informaworld.com/smpp/content~content=a741902944~db=
all~order=page.

Anonymous:1987:PAT

[Ano87] Anonymous, editor. Proceedings of AIDA-87. Third Annual Conference on Artificial Intelligence and Ada. George Mason Univ, Fairfax, VA, USA, 1987.

Anonymous:1988:MDD

[Ano88a] Anonymous. Macintosh download disk 16, 1988. 1 computer disk. Title supplied by cataloger. For fuller description of contents see list at Computing and Reserve Library Desk. 3dplt — Compact — DA Appl Font 1.03 — DA Artisto 1.41 — DA Blank Screen — DA Camera — DA ControlPanelPlus 0.85 — DA Coordinates — DA DeskZap 1.3 — DA Dvorak3.0+ — DA f(n)key — DA Fade to Black 2.0 — DA Font Grabber — DA Glass — DA Idle — DA MakePICTfile — DA Mouseometer — DA mouseprint — DA Multi-Scrap — DA New Idle1 — DA New Scrapbook2.0 — DA NumCaps 2.0 — DA ParmBlaster — DA ProMouse1.1 — DA Screendump2 — DA ScreenSave — DA Show Clip — DA Show PICTure — DA Stars1.3 — DA Studdclip — DA ZoomIdle 1.1 — EDIT 2.0d1 — Icon Exchanger 2.0 — Install Twelve-C Demo — Inventory — Keeper — MAC ID — MassCopier — microEMACS — Orion1.4 — PageSetupCustomizer — RamDisk+ — VideoWorks. Collection of freeware and shareware. System requirements: Macintosh. Macintosh (Computer).

Anonymous:1988:PFA

[Ano88b] Anonymous, editor. Proceedings of the Fourth Annual Artificial Intelligence and Advanced Computer Technology Conference. Tower Conference Management, Glen Ellyn, IL, USA, 1988.

Anonymous:1988:UPC

[Ano88c] Anonymous, editor. USENIX Proceedings. C++ Conference. USENIX Association, Berkeley, CA, USA, 1988.

Anonymous:1989:PAE

[Ano89] Anonymous, editor. Proceedings of the Autumn 1989 EUUG Conference. Eur. UNIX Syst. User Group, Buntingford, UK, 1989. ISBN 0-9513181-3-6. LCCN ????

Anonymous:1990:EXW

[Ano90a] Anonymous, editor. European X Window System Conference. CEP Consultants, Edinburgh, UK, 1990.

Anonymous:1990:PWU

[Ano90b] Anonymous, editor. Proceedings of the Winter 1990 USENIX Conference. USENIX Association, Berkeley, CA, USA, 1990.

Anonymous:1990:UAP

[Ano90c] Anonymous, editor. USENIX Association. Proceedings of the Fourth Large Installation System Administrator’s Conference. USENIX Association, Berkeley, CA, USA, 1990.

Anonymous:1990:UCC

[Ano90d] Anonymous, editor. USENIX C++ Conference. USENIX Association, Berkeley, CA, USA, 1990.

Anonymous:1991:SIM

[Ano91] Anonymous, editor. Second International Modula-2 Conference. Modula-2 and Beyond. Loughborough Univ. Technol, Loughborough, UK, 1991.

Anonymous:1992:PSE

[Ano92] Anonymous, editor. Proceedings of the Spring 1992 EurOpen/ USENIX Workshop. EurOpen, Buntingford, Herts, UK, 1992.

Anonymous:1993:GCT

[Ano93a] Anonymous. GNU C tools move to MS-DOS for H8/ 300 micronC. EDN, 38(5):66–??, March 1993. CODEN EDNSBH. ISSN 0012-7515, 0364-6637.

Anonymous:1993:GST

[Ano93b] Anonymous. GNU set to take on UNIX. Software Magazine, 13(6):33–??, April 1993. CODEN SMWMEQ. ISSN 0897-8085.

Anonymous:1993:PSU

[Ano93c] Anonymous, editor. Proceedings of the Summer 1993 USENIX Conference. USENIX Association, Berkeley, CA, USA, 1993.

Anonymous:1993:PUM

[Ano93d] Anonymous, editor. Proceedings of the USENIX Mach III Symposium. USENIX Association, Berkeley, CA, USA, 1993.

Anonymous:1994:FCP

[Ano94a] Anonymous, editor. 1993 FORML Conference Proceedings. Forth Modification Laboratory. 1993 FORML Conference. EuroFORTH ’93 Conference. EuroForth ’92 Conference. Forth Interest Group, Oakland, CA, USA, 1994.

Anonymous:1994:PSA

[Ano94b] Anonymous, editor. Proceedings of the sixth annual Embedded Systems Conference: Santa Clara, California, September 20– 23, 1994, number 2 in Proceedings of the Annual Embedded Systems Conference 1994. Miller Freeman Publications, San Francisco, CA, USA, 1994. ISBN 0-87930-355-7 (vol. 1), 0-87930-356-5 (vol. 2). LCCN ???? Two volumes.

Anonymous:1994:PUS

[Ano94c] Anonymous, editor. Proceedings of the USENIX Symposium on Very High Level Languages (VHLL). USENIX Association, Berkeley, CA, USA, 1994.

Anonymous:1995:NPR

[Ano95a] Anonymous. New products: Red Hat Linux Developers Package; CE Editor for Linux; Mathematica for Linux; Directories & References Corrections; Linux Resources; Advertisers Index; Consultants Directory; Subscription Information. Linux Journal, 19:??, November 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-

3827 (electronic). URL http://www.linuxjournal.com/
index.html;   http://www.linuxjournal.com/issue19/
ad19.html;   http://www.linuxjournal.com/lj/Extras/
consultants.dir.html;   http://www.linuxjournal.com/
lj/index.html;   http://www.linuxjournal.com/lj/ljstaff.
html;   http://www.linuxjournal.com/lj/ljsubsorder.
html.

Anonymous:1995:NPG

[Ano95b] Anonymous. A new project or a GNU project? Linux Journal, 13:??, May 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue13/gnu.html.

Anonymous:1995:SSF

[Ano95c] Anonymous. Software spotlights: F77 to F90 converters; GNU F77 (“g77”) – successor to “f2c”? ACM Fortran Forum, 14 (3):4–8, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:UGR

[Ano95d] Anonymous. Update GNAT report. Ada User Journal, 16(1): 27–??, March 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:WGG

[Ano95e] Anonymous. What’s GNU?: GNU coding standards. Linux Journal, 16:??, August 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue16/gnu16.html.

Anonymous:1995:WGPa

[Ano95f] Anonymous. What’s GNU?: Plan 9 (part 1 of 2). Linux Journal, 11:??, March 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue11/gnu11.html.

Anonymous:1995:WGPb

[Ano95g] Anonymous. What’s GNU?: Plan 9 (part 2 of 2). Linux Journal, 12:??, April 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue12/gnu12.html.

Anonymous:1996:CME

[Ano96a] Anonymous. Calc Manual, for Emacs Calc Version 2.02. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1996. ISBN 1-882114-18-3. ???? pp. LCCN ????

Anonymous:1996:NPA

[Ano96b] Anonymous. New products: Absoft Fortran 77 for Linux; Spyglass Client Web Technology Kit in Red Hat Linux; BLAST Communications Software for Linux; TenXpert CD Server Upgrade; TransactNet Web Interface Toolkit; 32 bit ODBC Driver for C-Tree Plus; Phonetics Data Remote. Linux Journal, 31:??, November 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1996:NPO

[Ano96c] Anonymous. New products: ObjectSpace Java Generic Library; ObjectSpace Web Toolkit; Innovative Software InvisibleWeb & Offline Proxy Server; Datacomm Internet and Intranet/Web Server with Cyrix 166MHz chip; ARDI Executor 2; TowerEiffel Release 2.0; Debian Linux 1.1; Open Systems Management COS/Print; Amtec Engineering Tecplot 7.0; Thought, Inc. VanillaSearch; Dimension X Liquid Reality Developer’s Kit; X Inside, Inc. Accelerated OpenGL Solution for Linux; Spire Technologies Tactician Plus. Linux Journal, 30:??, October 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://www.

linuxjournal.com/issue30/newprod.html.

Anonymous:1996:OCS

[Ano96d] Anonymous. Octave control systems toolbox: a MATLAB-like CACSD environment. Proceedings of the IEEE International Symposium on Computer-Aided Control System Design, pages 386–391, 1996. IEEE catalog number 96TH8136.

Anonymous:1996:SNA

[Ano96e] Anonymous. Software: New announcements: Free Fortran Subset ELF90; Fortran Plus from N. A. Software; ForeSys tool suite from Scientific Services; Fortran information resources; Fortran 90 information (August) from Michael Metcalf. ACM Fortran Forum, 15(2):31–35, August 1996. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:BRPh

[Ano97a] Anonymous. Book review: Programming with GNU software: By Mike Loukides and Andy Oram. O’Reilly, Sebastopol, CA. (1997). 244 pages. $39.95 (CD included). Computers and Mathematics with Applications, 33(9):148, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122197901577.

Anonymous:1997:BRPc

[Ano97b] Anonymous. Book review: Programming with GNU Software. Linux Journal, 38:??, June 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:NPS

[Ano97c] Anonymous. New products: SAMBA 1.9.17; Laptop Accelerated-X Display Server; MetaCard 2.1.1 and the MetaCard Starter Kit; pryzm; QuickStart; TIMESERIES by Empress; TalentSoft Web+ 3.0; Open Sound System Sound Drivers for Linux; Gnu Emacs Verson 20. Linux Journal, 44:??, December 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/2550.html.

Anonymous:1997:TNF

[Ano97d] Anonymous. Technology news: Fortran 90 news; free software; symbolic computing packages; Matlab 5; Web products. IEEE Computational Science & Engineering, 4(1):87–??, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print),

1558-190X (electronic). URL http://dlib.computer.org/
cs/books/cs1997/pdf/c1087.pdf.

Anonymous:1998:NPG

[Ano98] Anonymous. New products: GO-Global 1.5, GraphOn Corp.; InterBase 5 for Linux, InterBase Software Corp.; ObjectTeam for Linux, Cayenne Software, Inc.; NetWare for Linux 1.0 and KDE, Caldera, Inc.; 264DP Screamer Dual Alpha 21264, Microway, Inc.; NetBeans Developer 2.0, Beta 3, NetBeans, Inc.; Metro Link Motif Complete!, Metro Link Inc.; Debian GNU/ Linux 2.0 “Hamm”, Debian GNU/Linux; Journyx Webtime Version 2.0, Journyx LLC. Linux Journal, 55:91, 95, November 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1999:BROa

[Ano99a] Anonymous. Book review: OpenSources: Voices from the open source revolution: Edited by Chris DiBona, Sam Ockman and Mark Stone. O’Reilly, Sebastopol, CA. (1999). 272 pages. $24.95. Computers and Mathematics with Applications, 37(10):173, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0898122199903371.

Anonymous:1999:DLS

[Ano99b] Anonymous. Debian Linux ‘Super’ package buffer overflow. Network Security, 1999(3):3, March 1999. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1353485899900144.

Anonymous:1999:FOS

[Ano99c] Anonymous. Free open-source code to result in global private network. Network Security, 1999(11):4, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1353485899901319.

Anonymous:1999:LEF

[Ano99d] Anonymous. Letter to the Editor: On free software. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/1999-2/freesoftware.
html.

Anonymous:19xx:GGA

[Anoxx] Anonymous. GAWK: The GNU Awk User’s Guide. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, second edition, 19xx. ISBN 1-882114-27-2. ???? pp. LCCN ????

Anonymous:2000:BRLa

[Ano00a] Anonymous. Book review: Learning Debian GNU/Linux: By Bill McCarty. O’Reilly, Sebastopol, CA. (1999). 343 pages. $34.95 (CD-ROM included). Computers and Mathematics with Applications, 39(3–4):265, February 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122100900713.

Anonymous:2000:BRLb

[Ano00b] Anonymous. Book review: Learning Red Hat Linux: By Bill McCarty. O’Reilly, Sebastopol, CA. (1999). 378 pages. $34.95 (CD-ROM included). Computers and Mathematics with Applications, 39(3–4):265, February 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122100900695.

Anonymous:2000:BRL

[Ano00c] Anonymous. Book reviews: Linux Red Hat Certified Engineer Exam Cram by Andrew G. Feinberg; JavaScript Application Cookbook by Ralph Krause; Programming Pearls, Second Edition by Harvey Friedman. Linux Journal, 73:??, May 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.com/lj-issues/issue73/3831.
html;   http://noframes.linuxjournal.com/lj-issues/
issue73/3833.html;   http://noframes.linuxjournal.
com/lj-issues/issue73/3846.html.

Anonymous:2000:BROb

[Ano00d] Anonymous. Book reviews: Open Source Linux: Web Programming by Daniel Lazenby; Building Database Applications on the Web Using PHP3 by Gaelyne R. Gasson. Linux Journal, 76:??, August 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.com/lj-issues/issue76/3869.
html;   http://noframes.linuxjournal.com/lj-issues/
issue76/4060.html.

Anonymous:2000:BROa

[Ano00e] Anonymous. Book reviews: Oracle Database Administration by Charles Curley; Using Caldera Open Linux: Special Edition by Ben Crowder; Linux System Administration by Paul Almquist; Learning Debian GNU/Linux by Marjorie Richardson. Linux Journal, 71:??, March 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.com/lj-issues/issue71/3561.

html;   http://noframes.linuxjournal.com/lj-issues/
issue71/3670.html;   http://noframes.linuxjournal.
com/lj-issues/issue71/3815.html;   http://noframes.
linuxjournal.com/lj-issues/issue71/3821.html.

Anonymous:2000:HPS

[Ano00f] Anonymous. Hewlett–Packard setzt auf Linux — HP forciert die Portterung des Open-Source-Systems auf Intels Itanium und PA-Risc-CPUs. (German) [Hewlett–Packard sets up Linux — HP forces the porting of open-source systems to Intel’s Itanium and PA-RISC CPUs]. Computerwoche, 27(2):26, ???? 2000. ISSN 0170-5121.

Anonymous:2000:IAO

[Ano00g] Anonymous. IBM to adopt open source. Network Security, 2000(12):2, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1353485800120033.

Anonymous:2000:LBW

[Ano00h] Anonymous. Linux beats Windows — open-source Trillian Linux will be ready for Intel’s 64-bit Itanium processors. Computer Shopper, pages 254–255, June 2000. ISSN 0886-0556.

Anonymous:2000:NPAa

[Ano00i] Anonymous. New products: AVP for Linux/FreeBSD UNIX, Kaspersky Lab Ltd.; API PowerRAC Chassis 320, Alpha Processor Inc.; ODBC-ODBC Bridge, Easysoft Ltd.; LinkScan 6.1, Electronic Software Publishing Corporation; Metro-X Enhanced Server CD, Metro Link, Inc.; P-STAT Statistical Software, P-STAT, Inc.; System Manager in a Box v1.0, Pega-Soft Canada; PGI Workstation 3.1, PGI; Quick Restore 2.6, Workstation Solutions, Inc.; Threads.h++ and Tools.h++ Professional, Rogue Wave Software; Scriptics Connect 1.0, 1.1, Scriptics Corporation; TapeWare 6.2 Backup Software, Yosemite Technologies, Inc.; DoubleVision for Linux Systems, Tridia Corporation. Linux Journal, 71:??, March 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:NPF

[Ano00j] Anonymous. New products: FileZerver, Microtest, Inc.; AT75C310, Aplio Inc.; Eyelet GUI, MoJo Designs Inc.; J2SE 1.2.2 for Linux, Sun Microsystems, Inc.; GNUPro Tools for IA-64, Red Hat Software; Linux edition of “A Mother’s Shoah”, IL NewMedia Publishing; +One Station, Maxspeed Corporation; Parallel Computing Toolkit, Wolfram Research; Rave Systems RackMount-1UAXe, Rave Computer Association, Inc.; SafeWrite, TurnSafe Technologies, Inc.; Progress SonicMQ Adds Support for Linux, Progress Software Corporation; System Blocks, SM&A Corp.; T.Rex, Freemont Avenue Software, Inc.; Videomodem, COM One Services; SNA Gateway, Gcom, Inc.; Best Linux 2000, SOT Finnish Software Engineering Ltd. Linux Journal, 74:??, June 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:NPP

[Ano00k] Anonymous. New products: PerfectBACKUP+ 6.1, Merlin Software Technologies; Linux Driver for HIPPI 800, Essential Communication Corporation; Linux by Libranet, Libra Computer Systems Ltd.; Programming Development Kit, Macmillan Computer Publishing; Linux Anti-Virus Solution, DOLFIN.COM Inc.; OpenDesk.com version 1.0, HBE Software; UnForm v4.0, Synergetic Data Systems; Max for Linux, PlugSys International LLC; PizzaBox Linux Distribution, KYZO Ltd, Little Streams, The Abbotsbrook, Bourne End, Bucks; Appgen Linux Java Client and PowerWindows Applications, Appgen Business Software, Inc.; IVR Server, Open Source Telecom; Photogenics, Paul Nolan Ltd. Linux Journal, 70:??, February 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:SLT

[Ano00l] Anonymous. Strictly on-line: T/TCP: TCP for Transactions by Mark Stacey, Ivan Griffin and John Nelson; POSIX Thread Libraries by Felix Garcia and Javier Fernandez; Linux and Open-Source Applications by Peter Jones and M. B. Jorgenson; Laptops for Linux! by Jason Kroll. Linux Journal, 70:??, February 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.com/lj-issues/issue70/3075.
html;   http://noframes.linuxjournal.com/lj-issues/
issue70/3184.html;   http://noframes.linuxjournal.
com/lj-issues/issue70/3683.html;   http://noframes.
linuxjournal.com/lj-issues/issue70/3766.html.

Anonymous:2001:AFP

[Ano01a] Anonymous. Announcements: Free public beta test of Compaq Fortran 1.1; new release of coco; Lahey/Fujitsu Fortran 95 (LF95) Linux version 6.0; NAGWare f95 compiler release 4.1; N. A. Software FortranPlus compiler version 2.2; the Fortran Company; new journal: Journal of Computational Methods in Applied Science and Engineering. ACM Fortran Forum, 20(1):21–24, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:DGLa

[Ano01b] Anonymous. Debian GNU, Linux — i386, GNU, Hurd — i386, Version 2.2 — “Potato”, release 2 [Elektronische Ressource]. Lehmann, Köln, Germany, 2001. ISBN 3-931253-74-0. LCCN ???? Includes six CD-ROMs.

Anonymous:2001:DGLc

[Ano01c] Anonymous. Debian GNU, Linux 3.0 (final) [Elektronische Ressource]: woody; i386 (Intel, AMD, Cyrix usw.); zusätzlich auf CD: Openoffice 1.0, XFree 86 4.2, Modellbahnsteuerungssoftware, Debian GNU/Linux-Anwenderhandbuch, Linux-Anwenderhandbuch. Lehmann, Berlin, Germany, 2001. ISBN 3-931253-89-9. LCCN ???? Includes eight CD-ROMs.

Anonymous:2001:DGLb

[Ano01d] Anonymous. Debian GNU, Linux 3.0 (pre) [Elektronische Ressource]: wizards of OS; 6 CD-ROMs mit WOS-Doku I und II und neuem KNOPPIX. Lehmann, Berlin, Germany, 2001. ISBN 3-931253-81-3. LCCN ???? Includes six CD-ROMs.

Anonymous:2001:EOS

[Ano01e] Anonymous. Editorial: Open source and empirical software engineering. Empirical Software Engineering, 6(3):193–194, September 2001. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1017379030770.

Anonymous:2001:GBO

[Ano01f] Anonymous. Germany backs open source. Network Security, 2001(12):2, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1353485801012041.

Anonymous:2001:GEO

[Ano01g] Anonymous. Guest editorial: Open source software: investigating the software engineering, psychosocial and economic issues. Information Systems Journal, 11(4):273–276, 2001. ISSN 1350-1917 (print), 1365-2575 (electronic).

Anonymous:2001:OSM

[Ano01h] Anonymous. Open source movement turns tables on IP laws. Network Security, 2001(4):3–4, April 1, 2001. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1353485801004044.

Anonymous:2001:PFS

[Ano01i] Anonymous. Products: Free Software Foundation updates compiler toolset; IT Factory’s Lotus Tools Suite; Cardiff Software’s real-time document verification application; Great Bridge updates open source database; OpenPath Products’ wireless application authoring tool; Curl’s Web application development environment; ThinAirApp ships mobile Visual Basic IDE; Princeton Softech updates server archiving software; Codemesh releases Java communications environment. Computer, 34(9):112–114, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/books/co2001/pdf/r9112.pdf;
 http://www.computer.org/computer/co2001/r9112abs.
htm.

Anonymous:2001:PWB

[Ano01j] Anonymous. Products: Web-based remote administration tools; SGDL System’s 3D model development language kit; MigraTEC’s Solaris-to-Linux migration software; Visual Numerics updates C numerical library; Stardock’s Windows skin development software; InterNetwork’s new load capacity testing software; SuSE Linux for PowerPC; Raytheon updates network security tools; Tasking updates embedded development tools; ExoLab Group offers open-source databinding software; Omnicore Software’s Java development environment; Basis International releases Java-based business basic; Zondigo’s wireless software development kit; MDD introduces password administration software; StatSoft revises data visualization tool; Abaco updates mobile application development framework. Computer, 34(6):90–93, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814

(electronic). URL http://dlib.computer.org/co/books/
co2001/pdf/r6090.pdf.

Anonymous:2002:FIO

[Ano02a] Anonymous. A further investigation of open source software: community, co-ordination, code quality and security issues. Information Systems Journal, 12(1):3–5, 2002. ISSN 1350-1917 (print), 1365-2575 (electronic).

Anonymous:2002:PXO

[Ano02b] Anonymous. Products: XML-oriented FrameMaker from Adobe Systems; Breeze Factor’s XML data-binding tool for Java; Microsoft’s speech-enabling tools for Web applications; Agere Systems’s wireless LAN management tool; Macromedia debuts ColdFusion MX; OpenOffice.org unveils open source productivity suite; Integrated Measurement Systems’ validation tester. Computer, 35(6):82–83, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-

0814 (electronic). URL http://csdl.computer.org/dl/
mags/co/2002/06/r6082.htm;   http://csdl.computer.
org/dl/mags/co/2002/06/r6082.pdf.

Anonymous:2003:AOS

[Ano03a] Anonymous. Avanti: Open source compact, self-contained ILS in Java. Information Retrieval and Library Automation, 38(8):7–??, 2003. CODEN IRLAAQ. ISSN 0020-0220.

Anonymous:2003:CADa

[Ano03b] Anonymous. Call for articles: Developing with open source software. IEEE Software, 20(3):1, May/June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459

(electronic). URL http://dlib.computer.org/so/books/
so2003/pdf/s3001.pdf.

Anonymous:2003:CADb

[Ano03c] Anonymous. Call for articles: Developing with open source software. IEEE Software, 20(4):1, July/August 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://www.computer.org/software/so2003/s4001.
pdf.

Anonymous:2003:LUE

[Ano03d] Anonymous. Linux/Unix extend Red Hat 8’s functionality by adding Flash, extra fonts and Java. Personal computer world, 26(3):202–207, 2003. CODEN PCWODU. ISSN 0142-0232.

Anonymous:2003:NUP

[Ano03e] Anonymous. News 2.0: Uncrackable passwords; WebFountain drinks down the Web; embracing open source in India. ACM Queue: Tomorrow’s Computing Today, 1(5):8, July/August 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:CSI

[Ano04a] Anonymous. Closed subject: IBM and BEA Systems call for Sun to contribute Java as open-source code, but Sun’s CEO nixes the idea. Information Week, 996:24, 2004. CODEN INFWE4. ISSN 8750-6874.

Anonymous:2004:PIU

[Ano04b] Anonymous. Products: IMSI updates popular CAD tool; immersive data visualization for PC users; Zero G releases InstallAnywhere 6; CodeFutures updates database persistence tool; new streaming-media system from Xiran; Parasoft offers software for Web services security; ClearMail 2.0 now available; 10x Software’s open source IDE. Computer, 37(1):114, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/01/r1114.htm;   http:
//csdl.computer.org/dl/mags/co/2004/01/r1114.pdf.

Anonymous:2004:POV

[Ano04c] Anonymous. Progress: Open vs. closed [An Empirical Study of Open-Source and Closed-Source Software Products, by James W. Paulson et al., IEEE Transactions on Software Engineering, April 2004, pp.]. IEEE Spectrum, 41(8):50, August 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:VPP

[Ano04d] Anonymous. Visual programming: The pros and cons of open source, plus Java and SQLite. Personal computer world, 27 (5):212, 2004. CODEN PCWODU. ISSN 0142-0232.

Anonymous:2005:BSG

[Ano05a] Anonymous. The back story: Gnomes and reactors. IEEE Spectrum, 42(4):7, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:DGL

[Ano05b] Anonymous. Debian GNU/Linux 3.1 r0a “Sarge”. Lehmann, Berlin, Germany, 2005. ISBN 3-86541-064-2. LCCN ???? EUR9.95. Two DVD-ROMs.

Anonymous:2006:PGI

[Ano06] Anonymous, editor. Proceedings of Gelato ICE: Itanium Conference and Expo: Spotlighting Linux on Itanium-based Platforms, October 1-4, 2006, Biopolis, Singapore. ????,

????, 2006. ISBN ????  LCCN ????  URL http:/

/www.ice.gelato.org/;   http://www.ice.gelato.org/
about/oct06_presentations.php.

Anonymous:2008:BHH

[Ano08a] Anonymous. Black hats hit Red Hat. Network Security, 2008(9):2, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1353485808701028.

Anonymous:2008:OSS

[Ano08b] Anonymous. Open source software getting better. Network Security, 2008(6):1–2, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1353485808700709.

Anonymous:2008:UOS

[Ano08c] Anonymous. Update: Open-source baby. IEEE Spectrum, 45 (6):14, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2010:FTS

[Ano10] Anonymous. Free Typesetting Software: L^{A}TEX, Troff, Scribus, Figlet, Lout, Noweb, Freetype, Graphite, Groff. Books LLC, ????, 2010. ISBN 1-157-24708-3. 44 pp. LCCN ????

Anonymous:2011:BRPc

[Ano11] Anonymous. Book review: Penetration Tester’s Open Source Toolkit, by Jeremy Faircloth. Third edition. Syngress. ISBN 978-1-59749-627-8. Network Security, 2011(12):4, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-

9371 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135348581170126X.

Anonymous:2014:ICD

[Ano14] Anonymous. International Congress of Drug Therapy in HIV Infection 2–6 November 2014, Glasgow, UK. Journal of the International AIDS Society, 17:??, 2014. ISSN 1758-2652.

Anonymous:2015:BRGb

[Ano15a] Anonymous. Book review: The GNU Make Book, John Graham-Cumming. No Starch Press. ISBN 978-1-59327-649-2. Network Security, 2015(5):4, May 2015. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1353485815300362.

Anonymous:2015:BHS

[Ano15b] Anonymous. A brief history of SHARE. Web document., May

28, 2015. URL https://www.share.org/d/do/11532.

Anonymous:2015:OSS

[Ano15c] Anonymous. Open source software distribution held not exempt function. Bruce R. Hopkins’ Nonprofit Counsel, 32(4): 3–4, 2015. ISSN 1542-8419 (print), 1542-8427 (electronic).

Anonymous:2016:NOS

[Ano16] Anonymous. New open source software for high resolution microscopy. Optik & Photonik, 11(3):18, 2016. ISSN 1863-1460 (print), 2191-1975 (electronic).

Anonymous:2018:SOS

[Ano18] Anonymous. Synopsis: Open source security and risk analysis. Network Security, 2018(6):3, June 2018. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1353485818300515.

Anonymous:2019:ROS

[Ano19] Anonymous. Risky open source. Network Security, 2019(5): 19, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print),

1872-9371 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1353485819300625.

Anonymous:2021:PLP

[Ano21] Anonymous. Predicting the lifetime of pull requests in opensource projects. Journal of Software: Evolution and Process, 33(6):e2337:1–e2337:??, June 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anthes:2016:NOS

[Ant16] Gary Anthes. News: Open source software no longer optional. Communications of the ACM, 59(8):15–17, August 2016. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL http://cacm.acm.org/magazines/2016/
8/205050/fulltext.

Alami:2022:PRG

[APCs22] Adam Alami, Raúl Pardo, Marisa Leavitt Cohn, and Andrzej W sowski. Pull request governance in open source communities. IEEE Transactions on Software Engineering, 48 (12):4838–4856, December 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Avelino:2019:MAC

[APHV19] Guilherme Avelino, Leonardo Passos, Andre Hora, and Marco Tulio Valente. Measuring and analyzing code authorship in 1 + 118 open source projects. Science of Computer Programming, 176(??):14–32, May 1, 2019. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167642318300388.

Allen:2014:IOS

[APK14a] Bruce M. Allen, Paul K. Predecki, and Maciej Kumosa. Integrating open-source software applications to build molecular dynamics systems. Journal of Computational Chemistry, 35(9):756–764, 2014. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Allen:2014:SNU

[APK14b] Bruce M. Allen, Paul K. Predecki, and Maciej Kumosa. Software news and updates: Integrating open-source software applications to build molecular dynamics systems. Journal of Computational Chemistry, 35(9):756–764, April 5, 2014. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Arceneaux:1994:CGI

[Arc94] Joseph Arceneaux. CUG392 — GNU indent v1.8. C Users Journal, 12(1):88–??, January 1994. ISSN 0898-9788.

Akkas:1997:ITI

[AS97] A. Akkas and M. J. Schulte. Implementing and testing interval operations and intrinsics in the GNU Fortran compiler. Technical report, Lehigh University, Bethlehem, PA, USA,

December 1997. URL http://home.ku.edu.tr/~ahakkas/
publications/Impl_Test_Intv_Op.pdf.

Applewhite:2003:NSG

[AS03] Ashton Applewhite and Dale Strok. In the news: Should governments go open source?; Linux gets some respect; face to face with the experts. IEEE Software, 20(4):88–93, July/ August 2003. CODEN IESOEG. ISSN 0740-7459 (print),

0740-7459 (electronic). URL http://dlib.computer.org/
so/books/so2003/pdf/s4088.pdf.

Al-Sad:2019:RBD

[ASAAM^{+}19] Mohammad F. Al-Sa’d, Abdulla Al-Ali, Amr Mohamed, Tamer Khattab, and Aiman Erbad. RF-based drone detection and identification using deep learning approaches: an initiative towards a large open source drone database. Future Generation Computer Systems, 100(??):86–97, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X18330760.

Antoniades:2002:NSM

[ASAB02] I. P. Antoniades, I. Stamelos, L. Angelis, and G. L. Bleris. A novel simulation model for the development process of open source software projects. Software Process: Improvement and Practice, 7(3–4):173–188, September/December 2002. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ataei:2021:LOS

[ASC^{+}21] Mohammadmehdi Ataei, Vahid Shaayegan, Franco Costa, Sejin Han, Chul B. Park, and Markus Bussmann. LBfoam: an open-source software package for the simulation of foaming using the Lattice Boltzmann Method. Computer Physics Communications, 259(??):Article 107698, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030343X.

USENIX:1995:PTT

[Ass95] USENIX Association, editor. Proceedings of the Tcl/Tk Workshop, July 6–8, 1995, Toronto, Ontario, Canada, Proceedings of the Tcl Tk Workshop — USENIX Association 1995; 3rd. USENIX Association, Berkeley, CA, USA, 1995. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

Abrahao:2023:OSS

[ASS^{+}23] Silvia Abrahão, Miroslaw Staron, Alexander Serebrenik, Birgit Penzenstadler, and Rafael Capilla. Open source software: Communities and quality. IEEE Software, 40(4):96–99, 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Asundi:2005:NEE

[Asu05] Jai Asundi. The need for effort estimation models for open source software projects. ACM SIGSOFT Software Engineering Notes, 30(4):1–3, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alrabaee:2018:FRE

[ASWD18] Saed Alrabaee, Paria Shirani, Lingyu Wang, and Mourad Debbabi. FOSSIL: A resilient and efficient system for identifying FOSS functions in malware binaries. ACM Transactions on Privacy and Security (TOPS), 21(2):8:1–8:??, February 2018. ISSN 2471-2566 (print), 2471-2574 (electronic). URL https:

//dl.acm.org/citation.cfm?id=3175492.

Aleksander:1992:ANN

[AT92] Igor Aleksander and John Taylor, editors. Artificial neural networks, 2: proceedings of the 1992 International Conference on Artificial Neural Networks (ICANN-92), Brighton, United Kingdom, 4–7 September 1992. Elsevier, Amsterdam, The Netherlands, 1992. ISBN 0-444-89488-8. LCCN QA76.87.I56 1992. Two volumes.

Avery:2019:XNV

[ATCZ19] Patrick Avery, Cormac Toher, Stefano Curtarolo, and Eva Zurek. XtalOpt Version r12: an open-source evolutionary algorithm for crystal structure prediction. Computer Physics Communications, 237(??):274–275, April 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465518304089.

Adams:2011:OSS

[AtHR11] Michael Adams, Arthur H. M. ter Hofstede, and Marcello La Rosa. Open source software for workflow management: The case of YAWL. IEEE Software, 28(3):16–19, May/June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Arceneaux:1992:PGS

[ATHW92] J. Arceneaux, M. Tiemann, and D. V. Henkel-Wallace. The portability of GNU software. In Anonymous [Ano92], pages 89–103.

Aldrich:2022:TPH

[ATM22] Guenever Aldrich, Danny Tsang, and Jason McKenney. Three-part harmony for program managers who just don’t get it, yet: Open-source software, open standards, and agile software development. ACM Queue: Tomorrow’s Computing Today, 20(6):58–79, 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https://dl.acm.org/

doi/10.1145/3576027.

Arnold:2004:IPN

[AV04] T. W. Arnold and L. P. Van Doorn. The IBM PCIXCC: a new cryptographic coprocessor for the IBM eServer. IBM Journal of Research and Development, 48(3/4):475–487, ???? 2004. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-

8556 (electronic). URL http://www.research.ibm.com/
journal/rd/483/arnold.html;   http://www.research.
ibm.com/journal/rd/483/arnold.pdf.

Andujar:2016:OSF

[AVA^{+}16] Francisco J. Andújar, Juan A. Villar, Francisco J. Alfaro, José L. Sánchez, and Jesus Escudero-Sahuquillo. An opensource family of tools to reproduce MPI-based workloads in interconnection network simulators. The Journal of Supercomputing, 72(12):4601–4628, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Avetisyan:2006:IRA

[Ave06]  Arutyun I. Avetisyan.  The ISP RAS activities for

improving GCC for Itanium. In Anonymous [Ano06],

page ??  ISBN ????  LCCN ????  URL http://

www.ice.gelato.org/oct06/pres_pdf/gelato_ICE06oct_
gccimprov_avetisyan_ispras.pdf.

Aviram:1998:GON

[Avi98] Mariva H. Aviram. GNU offers a new kind of Java IDE. JavaWorld: IDG’s magazine for the Java community, 3(7), July 1998. CODEN ???? ISSN 1091-8906.

URL http://www.javaworld.com/javaworld/jw-07-1998/
jw-07-freebuilder.htm.

Ajila:2007:ESE

[AW07] Samuel A. Ajila and Di Wu. Empirical study of the effects of open source adoption on software development economics. The Journal of Systems and Software, 80(9):1517– 1529, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Angstadt:2018:MOS

[AWD^{+}18] Kevin Angstadt, Jack Wadden, Vinh Dang, Ted Xie, Dan Kramp, Westley Weimer, Mircea Stan, and Kevin Skadron. MNCaRT: An open-source, multi-architecture automataprocessing research and execution ecosystem. IEEE Computer Architecture Letters, 17(1):84–87, January/June 2018. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Abe:1993:PWD

[AY93] F. Abe and F. Yuasa, editors. Proceedings of Workshop on Distributed Computing and Network (KEK Proceedings 92-19). Nat. Lab. High Energy Phys, Ibaraki-ken, Japan, 1993.

Ayers:1997:CXG

[Aye97] Larry Ayers. A comparison of Xemacs and Gnu Emacs. Linux Journal, 34:??, February 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Ayers:2001:GEX

[Aye01] Larry Ayers. GNU Emacs and XEmacs. Prima Publishing, Roseville, CA, USA, 2001. ISBN 0-7615-2446-0. xxxv + 508 pp. LCCN QA76.76.T49 A94 2001. Includes CD-ROM.

Avery:2017:CRO

[AZ17a] Patrick Avery and Eva Zurek. Corrigendum to “RandSpg: an open-source program for generating atomistic crystal structures with specific spacegroups” [comput. phys. comm. 213 (2017) 208–216]. Computer Physics Communications, 220 (??):509, November 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465517301595.

See [AZ17b].

Avery:2017:ROS

[AZ17b] Patrick Avery and Eva Zurek. RandSpg: an open-source program for generating atomistic crystal structures with specific spacegroups. Computer Physics Communications, 213 (??):208–216, April 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465516303848.

Bayerl:2015:SFI

[BA15] Petra Saskia Bayerl and Babak Akhgar. Surveillance and falsification implications for open source intelligence investigations. Communications of the ACM, 58(8):62–69, August 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL http://cacm.acm.org/magazines/2015/
8/189852/fulltext.

Babu:2002:NHF

[Bab02] Satish Babu. New horizons of free software: An Indian perspective. TUGboat, 23(1):17–20, 2002. ISSN 0896-3207.

Bader:2007:FBG

[Bad07] Reinhold Bader. A Fortran binding for the GNU Scientific Library. ACM Fortran Forum, 26(2):4–11, August 2007. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Botana:2014:UFO

[BAE14] Francisco Botana, Miguel A. Abánades, and Jesús Escribano. Using a free open source software to teach mathematics. Computer Applications in Engineering Education, 22(4):728–735, 2014. CODEN CAPEED. ISSN 1061-3773 (print), 1099-0542 (electronic).

Baker:2020:CSU

[Bak20] Peter Baker. Code snippet: Using GNU Make to manage the workflow of data analysis projects. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v094c01;   https://www.jstatsoft.
org/index.php/jss/article/view/v094c01/v94c01.pdf.

Ballhausen:2019:FOS

[Bal19] M. Ballhausen. Free and open source software licenses explained. Computer, 52(6):82–86, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bansal:2016:EAS

[Ban16] Ankita Bansal. Empirical analysis of search based algorithms to identify change prone classes of open source software. Computer Languages, Systems and Structures, ??(??): 211–231, ???? 2016. CODEN ???? ISSN 1477-8424 (print),

1873-6866 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1477842416301397.

Bansal:2017:EAS

[Ban17] Ankita Bansal. Empirical analysis of search based algorithms to identify change prone classes of open source software. Computer Languages, Systems and Structures, 47 (Part 2)(??): 211–231, ???? 2017. CODEN ???? ISSN 1477-8424 (print),

1873-6866 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1477842416301397.

Baozong:1993:PTI

[Bao93] Yuan Baozong, editor. Proceedings / TENCON ’93, 1993 IEEE Region 10 Conference on Computer, Communication, Control, and Power Engineering, October 19–21, 1993, Beijing, Beijing International Convention Center, Beijing Continental Grand Hotel. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1233-3. LCCN QA75.5.I155 1993. IEEE Catalog No. 93CH3286-2.

Bendix:2000:SSC

[BAP00] Lars Bendix, Ulf Asklund, and Jonas Persson. Summary of the subworkshop on change management for open source software. Nordic Journal of Computing, 7(2):143–??, Summer 2000. CODEN NJCOFR. ISSN 1236-6064.

Baran:2000:NVf

[Bar00a] Nicholas Baran. News and views: Freenet: More anarchy for the Internet?; magnetic properties key to nanoengineering; nanoseconds not fast enough? here come femtoseconds; Caltech leads U.S. field in ACM programming contest; robotic surgeons may make fewer mistakes; free software for designing ICs. Dr. Dobb’s Journal of Software Tools, 25(6):18, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVM

[Bar00b] Nicholas Baran. News and views: More on tiny transistors; Open Source repository launched; design contest promotes new software tools; and then there’s a decryption contest; Fred Brooks wins ACM Turing Award. Dr. Dobb’s Journal of Software Tools, 25(3):18, March 2000. CODEN DDJOEB.

ISSN 1044-789X. URL http://sourceforge.net/.

Baran:2000:NVR

[Bar00c] Nicholas Baran. News and views: RSA algorithm in the public domain; Woz joins the Inventors Hall of Fame; entangled photons mean faster, smaller ICs; BEHEMOTH mothballed; Advanced Encryption Standard selected; SGI releases SDK as open source; WSDL spec released. Dr. Dobb’s Journal of Software Tools, 25(12):18, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2001:NVCb

[Bar01] Nicholas Baran. News and views: College Board to add Java test; astronomy project beats out algorithm for science prize; consortium to release XML business specification; open source developer agreement proposed; consortium hopes to eliminate computer failures; new linear algorithm for sequence analysis; distributing computing à la SETI catches on. Dr. Dobb’s Journal of Software Tools, 26(3):18, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://publicscience.

net/;   http://www.ddj.com/.

Barber:2016:BRE

[Bar16a] D. C. Barber. Book review: Essential MATLAB and Octave, by Jesus Rogel-Salazar. Contemporary Physics, 57(3):455–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bhatt:2016:SIO

[BAR16b] Punita Bhatt, Ali J. Ahmad, and Muhammad Azam Roomi. Social innovation with open source software: User engagement and development challenges in India. Technovation, 52–53(??):28–39, June/July 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0166497216000055.

Barba:2022:DRO

[Bar22] Lorena A. Barba. Defining the role of open source software in research reproducibility. Computer, 55(8):40–48, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bauer:2006:PPSb

[Bau06a] Mick Bauer. Paranoid penguin: security features in Debian 3.1. Linux Journal, 2006(145):??, May 2006. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bauer:2006:PPSa

[Bau06b] Mick Bauer. Paranoid penguin: security features in SUSE 10.0. Linux Journal, 2006(144):??, April 2006. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Baxter:2001:OSE

[Bax01] Michael Baxter. Open source in electronic design automation. Linux Journal, 82:162, 164–166, February 2001. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bowen:1991:HMC

[BB91] D. E. Bowen and A. C. Beers. A highly-portable Modula-2 compiler. In Anonymous [Ano91], pages 68–76.

Brown:2002:ROS

[BB02] Alan W. Brown and Grady Booch. Reusing open-source software and practices: The impact of open-source on commercial vendors. Lecture Notes in Computer Science, 2319:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2319/23190123.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2319/23190123.pdf.

Baraniuk:2008:VGW

[BB08] Richard G. Baraniuk and C. Sidney Burrus. Viewpoint: Global warming toward open educational resources. Communications of the ACM, 51(9):30–32, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baker:1996:SST

[BBD^{+}96a] Mary Baker, Bob Bruce, William H. Davidow, Michael Tiemann, and Linus Torvalds. Selling stuff that’s free: the commercial side of free software, 1996. URL http://www.usenix.

org/publications/library/proceedings/sd96/. Unpub-

lished invited talk at the USENIX 1996 Annual Technical Conference, January 22–26, 1996, San Diego, CA.

Beck:1996:LKI

[BBD^{+}96b] Michael Beck, Harold Bohme, Mirko Dzladzka, Ulrich Kunitz, Robert Magnus, and Dirk Verworner. Linux Kernel Internals. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-87741-4. xvii + 438 pp. LCCN QA76.76.O63 L55 1996. US$38.68. Forward by Linus Torvalds, creator of LINUX.

Burgess:2017:ISC

[BBdD17] Neil Burgess, Javier Bruguera, and Florent de Dinechin, editors. 2017 IEEE 24th Symposium on Computer Arithmetic (ARITH 24), London, UK, 24–26 July 2017. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2017. ISBN 1-5386-1966-0 (print), 1-5386-1965-2, 1-5386-1964-4. ISSN 1063-6889. LCCN QA76.9.C62

S95 2017. URL http://ieeexplore.ieee.org/servlet/
opac?punumber=8019911.

Bercea:2020:OSS

[BBE^{+}20] G. T. Bercea, A. Bataev, A. E. Eichenberger, C. Bertolli, and J. K. O’Brien. An open-source solution to performance portability for Summit and Sierra supercomputers. IBM Journal of Research and Development, 64(3/4):12:1–12:23, May/July 2020. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Bernardino:2021:IPG

[BBM^{+}21] Matheus Tavares Bernardino, Giuliano Belinassi, Paulo Meirelles, Eduardo Martins Guerra, and Alfredo Goldman. Improving parallelism in Git and GNU Compiler Collection: Strategies, difficulties, and lessons learned. IEEE Software, 38(5):92–100, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Buckland:1993:OPG

[BBNP93] M. K. Buckland, M. H. Butler, B. A. Norgard, and C. J. Plaunt. OASIS: prototyping graphical interfaces to networked information. In Bonzi [Bon93], pages 204–210. ISBN 0-938734-78-4. LCCN Z699.A1 A5 1993.

Bao:2020:AOS

[BC20a] Y. Bao and T. E. Carlson. Agile and open-source hardware. IEEE Micro, 40(4):6–9, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bhom:2020:HOS

[BC20b] Jihyun Bhom and Marcin Chrzaszcz. HEPLike: an open source framework for experimental likelihood evaluation. Computer Physics Communications, 254(??):Article 107235, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300692.

Belloni:2007:OSP

[BCB07] Mario Belloni, Wolfgang Christian, and Douglas Brown. Open source physics curricular material for quantum mechanics. Computing in Science and Engineering, 9(4):24–31, July/ August 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bosu:2017:PAS

[BCB^{+}17] A. Bosu, J. C. Carver, C. Bird, J. Orbeck, and C. Chockley. Process aspects and social dynamics of contemporary code review: Insights from open source development and industrial practice at Microsoft. IEEE Transactions on Software Engineering, 43(1):56–75, January 2017. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=7484733.

Bosu:2014:PIO

[BCG^{+}14] Amiangshu Bosu, Jeffrey Carver, Rosanna Guadagno, Blake Bassett, Debra McCallum, and Lorin Hochstein. Peer impressions in open source organizations: a survey. The Journal of Systems and Software, 94(??):4–15, August 2014. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121214000818.

Belt:2012:LEA

[BCHR12] Jason Belt, Patrice Chalin, John Hatcliff, and Robby. Leading-edge Ada verification technologies: highly automated Ada contract checking using Bakar Kiasan. ACM SIGADA Ada Letters, 32(3):3–4, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Bernaschi:2009:OOS

[BCI^{+}09] M. Bernaschi, F. Cacace, G. Iannello, M. Vellucci, and L. Vollero. OpenCAPWAP: an open source CAPWAP implementation for the management and configuration of WiFi hotspots. Computer Networks (Amsterdam, Netherlands: 1999), 53(2):217–230, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Beale:2016:OSC

[BCP^{+}16] Steven B. Beale, Hae-Won Choi, Jon G. Pharoah, Helmut K. Roth, Hrvoje Jasak, and Dong Hyup Jeon. Open-source computational model of a solid oxide fuel cell. Computer Physics Communications, 200(??):15–26, March 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465515003872.

Barbosa:2010:COS

[BCPS10] Luis S. Barbosa, Antonio Cerone, Alexander K. Petrenko, and Siraj A. Shaikh. Certification of open-source software: A role for formal methods? International Journal of Computer Systems Science and Engineering, 25(4):??, July 2010. CODEN CSSEEI. ISSN 0267-6192.

Baldassin:2008:OSB

[BCR^{+}08] Alexandro Baldassin, Paulo Centoducatte, Sandro Rigo, Daniel Casarotto, Luiz C. V. Santos, Max Schultz, and Olinto Furtado. An open-source binary utility generator. ACM Transactions on Design Automation of Electronic Systems., 13(2):27:1–27:??, April 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Becking:2005:MOS

[BCvE^{+}05] J. Becking, S. Course, G. van Enk, H. T. Hangyi, J. J. M. Lahaye, D. Ockeloen, R. Peters, H. Rosbergen, and R. van Wendel de Joode. MMBase: An open-source content management system. IBM Systems Journal, 44(2):381–??, ???? 2005. CODEN IBMSA7. ISSN 0018-8670.

Baiocchi:2003:GES

[BD03a] Giovanni Baiocchi and Walter Distaso. GRETL: Econometric software for the GNU generation. Journal of Applied Econometrics, 18(1):105–110, 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bayrak:2003:RBD

[BD03b] Coskun Bayrak and Chad Davis. The relationship between distributed systems and open software development. Communications of the ACM, 46(12):99–102, December 2003. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barry:2015:OSS

[BDAW15] David J. Barry, Charlotte H. Durkin, Jasmine V. Abella, and Michael Way. Open source software for quantification of cell migration, protrusions, and fluorescence intensities. Journal of Cell Biology, 209(1):163–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/209/1/163.

Bailey:2013:OSI

[BdP13] David H. Bailey and Marcos López de Prado. An open-source implementation of the critical-line algorithm for portfolio optimization. Algorithms (Basel), 6(1):169–196, March 2013. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https:/

/www.mdpi.com/1999-4893/6/1/169.

Browne:2014:COS

[BDP^{+}14] James C. Browne, Robert L. DeLeon, Abani K. Patra, William L. Barth, John Hammond, Matthew D. Jones, Thomas R. Furlani, Barry I. Schneider, Steven M. Gallo, Amin Ghadersohi, Ryan J. Gentner, Jeffrey T. Palmer, Nikolay Simakov, Martins Innus, Andrew E. Bruno, Joseph P. White, Cynthia D. Cornelius, Thomas Yearke, Kyle Marcus, Gregor von Laszewski, and Fugang Wang. Comprehensive, open-source resource usage measurement and analysis for HPC systems. Concurrency and Computation: Practice and Experience, 26(13):2191–2209, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barkat:2015:OSS

[BdSI15] Amine Barkat, Alysson Diniz dos Santos, and Sonia Ikken. Open source solutions for building IaaS clouds. Scalable Computing: Practice and Experience, 16(2):187–204, ???? 2015. CODEN ???? ISSN 1895-1767. URL https://www.scpe.

org/index.php/scpe/article/view/1089.

Brebner:2006:TWG

[BE06] Paul Brebner and Wolfgang Emmerich. Two ways to Grid: The contribution of Open Grid Services Architecture (OGSA) mechanisms to service-centric and resource-centric lifecycles.

Journal of Grid Computing, 4(1):115–131, March 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1570-7873&volume=4&issue=1&spage=115.

Beattie:1994:TPT

[Bea94] M. Beattie. TkPerl — a port of the Tk toolkit to Perl5. In Anonymous [Ano94c], pages 71–82.

Beard:2004:OSS

[Bea04] Jim Beard. An open-source system for electronic court filing. Linux Journal, 2004(122):3, June 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Beaton:2021:OSS

[Bea21] Wayne Beaton. Open source software engineering the Eclipse way. Computer, 54(6):59–63, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Becker:1993:TEG

[Bec93] Thomas Becker. Two extensions to GNU Emacs that are useful when editing TEX documents. TUGboat, 14(4):382– 386, December 1993. ISSN 0896-3207.

Beebe:1986:TIP

[Bee86] Nelson H. F. Beebe. TECO on the IBM PC. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 22, 1986. 62 pp.

Beebe:1991:ESa

[Bee91a] Nelson H. F. Beebe. L^{A}TEX editing support. Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 07, 1991. 28 pp. See also [Bee91b].

Beebe:1991:ESb

[Bee91b] Nelson H. F. Beebe. L^{A}TEX editing support. Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 07, 1991. 69 pp. This is an on-line Emacs INFO version of part of [Bee91a].

Beebe:2001:GSL

[Bee01] Nelson H. F. Beebe. GNU Scientific Library. World-Wide

Web document., 2001. URL https://www.math.utah.edu/
software/gsl.html.

Beebe:2017:DAS

[Bee17] Nelson H. F. Beebe. Decimal-arithmetic support in gcc compilers. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October

27, 2017. 13 pp. URL https://www.math.utah.edu/pub/
mathcw/dgcc/.

Bellomo:2000:DGL

[Bel00] Michael Bellomo. Debian GNU/Linux for dummies. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-0713-3. xxii + 324 pp. LCCN QA76.76.O63 B44978 2000.

Bellovin:2022:OST

[Bel22] Steven M. Bellovin. Open source and trust. IEEE Security & Privacy, 20(2):107–108, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

IEN034

[Ben78] C. J. Bennett. The GNOME controller, April 25, 1978. URL

http://www.cis.ohio-state.edu/htbin/ien/ien34.html.

Bernstein:1996:LCG

[Ber96] Daniel J. Bernstein. From the land of C++: Gnu Emacs: An X-based development tool. The X Journal: Computing Technology with the X Window System, 5(6):77–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Bergmann:2022:MOS

[Ber22] Seth Bergmann. Making open source textbooks, and diagrams with AlDraTex. TUGboat, 43(1):23–27, ???? 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/
tb43-1/tb133bergmann-diagrams.pdf.

Bauer:2001:LSY

[BES^{+}01] Barr Bauer, Jonathan Erickson, Richard Stallman, Mike Pentney, and Frank C. Earl. Letters: The state of your taxes; embedded space; more Open Source versus Free Software; shared source?; CORBA interoperability. Dr. Dobb’s Journal of Software Tools, 26(9):10, September 2001. CODEN DDJOEB.

ISSN 1044-789X. URL http://www.ddj.com/.

Beszedes:2003:OSM

[Bes03] Árpád Beszédes. Optimizing for space: Measurements and possibilities for improvement. In Hutton et al. [HDR03], pages 7–20. ISBN ???? LCCN ???? URL http://www.linux.org.

uk/~ajh/gcc/gccsummit-2003-proceedings.pdf.

Beszedes:2004:CBO

[Bes04] Árpád Beszédes. CSiBE Benchmark: One year perspective and plans. In Hutton et al. [HDR04], pages 7–15. ISBN ????

LCCN ???? URL http://people.redhat.com/lockhart/.
gcc04/MasterGCC-2side.pdf.

Belaire-Franch:2002:RPN

[BFC02] Jorge Belaire-Franch and Dulce Contreras. Recurrence plots in nonlinear time series analysis: Free software. Journal of Statistical Software, 7(9):1–18, ???? 2002. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v07/i09;   http://www.jstatsoft.org/
v07/i09/JSS_055.pdf;   http://www.jstatsoft.org/v07/
i09/updates.

Bonfa:2021:UOS

[BFI^{+}21] Pietro Bonfà, Jonathan Frassineti, Muhammad Maikudi Isah, Ifeanyi John Onuorah, and Samuele Sanna. UNDI: an opensource library to simulate muon-nuclear interactions in solids. Computer Physics Communications, 260(??):Article 107719, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303556.

Baker:1995:IAP

[BG95] T. P. Baker and E. W. Giering, III. Implementing Ada 9X protected objects and asynchronous transfer of control. International Journal of Mini and Microcomputers, 17(1):26–34, 1995. CODEN IJMMDE. ISSN 0702-0481.

Brorkens:2012:MUC

[BG12] Mark Brörkens and Ömer Gürsoy. Managing and understanding complex systems using traceability and open source software. INCOSE International Symposium, 22(1):2257–2265, 2012. ISSN 2334-5837.

Botsford:1994:PCI

[BGG^{+}94] J. Botsford, A. Gawman, M. Gentleman, E. Kidd, K. Lyons, and J. Slonim, editors. Proceedings. CASCON ’94. Integrated Solutions. Nat. Res. Council Canada, Ottawa, Ont., Canada, 1994.

Borchers:2000:POS

[BGG^{+}00] Robert Borchers, Susan Graham, Richard Gabriel, Todd Needham, and José Muñoz. Panel: Open source: IP in the Internet era. In ACM [ACM00], pages 79–80. ISBN

???? LCCN QA76.88. URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Balasubramanian:2015:EGL

[BGG^{+}15] Raghuraman Balasubramanian, Vinay Gangadhar, Ziliang Guo, Chen-Han Ho, Cherin Joseph, Jaikrishnan Menon, Mario Paulo Drumond, Robin Paul, Sharath Prasad, Pradip Valathol, and Karthikeyan Sankaralingam. Enabling GPGPU low-level hardware explorations with MIAOW: an open-source RTL implementation of a GPGPU. ACM Transactions on Architecture and Code Optimization, 12(2):21:1–21:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Butler:2020:MIO

[BGL^{+}20] Simon Butler, Jonas Gamalielsson, Björn Lundell, Christoffer Brax, Anders Mattsson, Tomas Gustavsson, Jonas Feist, and Erik Lönroth. Maintaining interoperability in open source software: a case study of the Apache PDFBox project. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121219302262.

Butler:2021:CCC

[BGL^{+}21] Simon Butler, Jonas Gamalielsson, Björn Lundell, Christoffer Brax, Johan Sjöberg, Anders Mattsson, Tomas Gustavsson, Jonas Feist, and Erik Lönroth. On company contributions to community open source software projects. IEEE Transactions on Software Engineering, 47(7):1381–1401, 2021. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Butler:2022:CCA

[BGL^{+}22] Simon Butler, Jonas Gamalielsson, Björn Lundell, Christoffer Brax, Anders Mattsson, Tomas Gustavsson, Jonas Feist, Bengt Kvarnström, and Erik Lönroth. Considerations and challenges for the adoption of open source components in software-intensive businesses. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121221002442.

B:1999:DIF

[BGM99] Praveen B., Deepak Gupta, and Rajat Moona. Design and implementation of a file system with on-the-fly data compression for GNU/Linux. Software—Practice and Experience, 29(10): 863–874, August 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bramer:2002:DGL

[BGO02] Michael Bramer, John Goerzen, and Ossama Othman. Debian GNU/Linux guide. Linuxland, München, Germany, 2002. ISBN 3-936759-00-6. 404 (est.) pp. LCCN ????

Beer:1989:DWT

[BGR89] M. D. Beer, S. M. George, and R. Rada. Developing writing tools for UNIX workstations. In Anonymous [Ano89], pages 31–36. ISBN 0-9513181-3-6. LCCN ????

Bolcer:2007:DCC

[BH07] John D. Bolcer and Robert B. Hermann. The development of computational chemistry in the United States. In Reviews in Computational Chemistry, volume 5, pages 1–63. Wiley, New York, NY, USA, January 2007. ISSN 1069-3599 (print), 1934-5372 (electronic). LCCN QD39.3.E46

R48. URL http://onlinelibrary.wiley.com/bookseries/
10.1002/SERIES6143.

Benlian:2011:CRI

[BH11] Alexander Benlian and Thomas Hess. Comparing the relative importance of evaluation criteria in proprietary and opensource enterprise application software selection — a conjoint study of ERP and office systems. Information Systems Journal, 21(6):503–525, 2011. ISSN 1350-1917 (print), 1365-2575 (electronic).

Berger:2017:AHA

[BH17] Lee R. Berger and John (John David) Hawks. Almost human: the astonishing tale of Homo naledi and the discovery that changed our human story. National Geographic, Washington, DC, USA, 2017. ISBN 1-4262-1811-7 (hardcover). 239 pp. LCCN GN284.5 .B47 2017.

Baldi:2003:TOO

[BHMB03] Stefan Baldi, Hauke Heier, and Anett Mehler-Bicher. Technical opinion: Open courseware and open source software. Communications of the ACM, 46(9):105–107, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bonzini:2001:LHG

[BHP^{+}01] Paolo Bonzini, Stuart Halloway, John Penry, Oluseyi Sonaiya, Bruce E. Hogman, Greg Bissell, Michael Hobbs, and Ben Laurie. Letters: Huge GCC executables; Java class loader; Department of Dumb Ideas; setting the record straight; the legacy of C#; DHTML source-code correction; shared libraries aren’t all bad; Zuse and Intel. Dr. Dobb’s Journal of Software Tools, 26(8):10, 12, August 2001. CODEN DDJOEB.

ISSN 1044-789X. URL http://www.ddj.com/.

Bigelow:2013:UGE

[Big13] D. Bigelow. Using the GNAT environment to maintain a large codebase inherited from another compilation system. Ada User Journal, 34(3):175–182, September 2013. CODEN AU-JOET. ISSN 1381-6551.

Bikerman:1996:WWC

[Bik96] Tania M. Bikerman. World wide computing: computing skills workshop, 1996–1997. Custom college series. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-064712-7. 376 pp. LCCN ????

Birkholz:1993:ELE

[Bir93] Matthew Birkholz. Emacs Lisp in Edwin Scheme. Thesis (m.s.), MIT Artificial Intelligence Laboratory, Cambridge, MA, USA, 1993. 81 (or 103??) pp. URL ftp://

publications.ai.mit.edu/ai-publications/1993/AITR-

1451.ps.Z. Also issued as Technical report AI-TR 1451.

Bosca:2014:SQQ

[BJJ14] R. Bosca, V. Johnson, and E. Jackson. SU-E-QI-18: QUAT-TRO: An open-source software package for quantitative imaging applications in assessing treatment response. Medical Physics, 41(6Part22):380–381, 2014. CODEN MPHYA6. ISSN 2473-4209.

Becker:2022:CPA

[BJM^{+}22] Benedikt Becker, Nicolas Jeannerod, Claude Marché, Yann Régis-Gianas, Mihaela Sighireanu, and Ralf Treinen. The CoLiS platform for the analysis of maintainer scripts in Debian software packages. International Journal on Software Tools for Technology Transfer (STTT), 24(5):717–733, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787

(electronic). URL https://link.springer.com/article/
10.1007/s10009-022-00671-1.

Balos:2008:OIA

[BJWZ08] Kazimierz Bałos, Marcin Jarzab, Damian Wieczorek, and Krzysztof Zieliński. Open interface for autonomic management of virtualized resources in complex systems — construction methodology. Future Generation Computer Systems, 24 (5):390–401, May 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bagley:1991:ATI

[BK91] S. C. Bagley and G. E. Kopec. Applications of text image editing. Proceedings of the SPIE — The International Society for Optical Engineering, 1460:71–79, 1991. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Bagley:1994:EIT

[Bk94] S. C. Bagley and G. E. kopec. Editing images of text. Communications of the ACM, 37(12):63–72, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bryan:2002:VOS

[BK02] David Bryan and David Kelly. VOCAL: Open source VoIP software for Linux. Embedded Linux Journal, 9:42–43, May/June 2002. CODEN ???? ISSN 1534-083X. URL

http://embedded.linuxjournal.com/magazine/issue09/
;   http://www.linuxdevices.com/articles/AT4384699491.
html.

Behr:2014:SPF

[BK14] Franz-Josef Behr and Barend Köbben. Selected papers from FOSS4G 2013: OSGeo’s global conference for open source geospatial software. Transactions in GIS, 18(4):477–479, 2014. ISSN 1361-1682 (print), 1467-9671 (electronic).

Bogart:2021:WHM

[BKHT21] Chris Bogart, Christian Kästner, James Herbsleb, and Ferdian Thung. When and how to make breaking changes: Policies and practices in 18 open source software ecosystems. ACM Transactions on Software Engineering and Methodology, 30 (4):42:1–42:56, July 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https://dl.acm.

org/doi/10.1145/3447245.

Bradski:2005:LBC

[BKP05] Gary Bradski, Adrian Kaehler, and Vadim Pisarevsky. Learning-based computer vision with Intel’s open source computer vision library. Intel Technology Journal, 9(2):119–130, May 2005. ISSN 1535-766X. URL ftp://download.intel.

com/technology/itj/2005/volume09issue02/art03_learning_
vision/vol09_art03.pdf;   http://developer.intel.com/
technology/itj/2005/volume09issue02/art03_learning_
vision/p01_abstract.htm.

Barcomb:2020:UPQ

[BKR^{+}20] A. Barcomb, A. Kaufmann, D. Riehle, K.-J. Stol, and B. Fitzgerald. Uncovering the periphery: A qualitative survey of episodic volunteering in free/libre and open source software communities. IEEE Transactions on Software Engineering, 46 (9):962–980, 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Black:1989:SSGa

[Bla89a] J. E. Black. SCRIBE support in GNU emacs. Technical information series 89CRD197, GE Research and Development Center, Schenectady, NY, USA, December 1989. v + 70 + 14 pp.

Black:1989:SSGb

[Bla89b] J. E. Black. Scribe support in GNU emacs. In ACM [ACM89], pages 125–135. ISBN 0-89791-337-X. LCCN QA 76.9 D6 S54 1989.

Blanchet:2006:CUI

[Bla06] Marc Blanchet. Configuration and Usage of IPv6-enabled Open Source Software, chapter 22, pages 381–384. Wiley, New York, NY, USA, 2006. ISBN 0-470-02874-2.

Baudis:2012:PSA

[BlG12] Petr Baudiš and Jean loup Gailly. PACHI: State of the art open source go program. Lecture Notes in Computer Science, 7168:24–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-31866-5_3/.

Behnamghader:2017:LSS

[BLG^{+}17] Pooyan Behnamghader, Duc Minh Le, Joshua Garcia, Daniel Link, Arman Shahbazian, and Nenad Medvidovic. A largescale study of architectural evolution in open-source software systems. Empirical Software Engineering, 22(3):1146–1193, June 2017. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
article/10.1007/s10664-016-9466-0.

Blossom:2004:GRT

[Blo04] Eric Blossom. GNU radio: Tools for exploring the radio frequency spectrum. Linux Journal, 2004(122):??, June 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bacon:2002:TOS

[BM02] Jean Bacon and Ken Moody. Toward open, secure, widely distributed services. Communications of the ACM, 45(6):59– 64, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barr:2006:PES

[BM06] Michael Barr and Anthony J. Massa. Programming embedded systems: with C and GNU development tools.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2006. ISBN 0-596-00983-6 (paperback). xxi + 301 pp. LCCN TK7895.E42 B37 2006eb; TK7895.E42. URL http://

www.loc.gov/catdir/enhancements/fy0715/2006287300-
d.html;   http://www.loc.gov/catdir/toc/fy0709/2006287300.
html;   http://www.oreilly.com/catalog/9780596009830.

Berger:2012:TOS

[BM12] Florian Berger and Wolfgang Müller. Towards an open source game engine for teaching and research. Lecture Notes in Computer Science, 7220:69–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31439-1_7/.

Bishop-VanHorn:2022:PSO

[BM22] Logan Bishop-Van Horn and Kathryn A. Moler. SuperScreen: an open-source package for simulating the magnetic response of two-dimensional superconducting devices. Computer Physics Communications, 280(??):Article 108464, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001837.

Braught:2018:MIP

[BMB^{+}18] Grant Braught, John Maccormick, James Bowring, Quinn Burke, Barbara Cutler, David Goldschmidt, Mukkai Krishnamoorthy, Wesley Turner, Steven Huss-Lederman, Bonnie Mackellar, and Allen Tucker. A multi-institutional perspective on H/FOSS projects in the computing curriculum. ACM Transactions on Computing Education, 18(2):7:1–7:??, July 2018. CODEN ???? ISSN 1946-6226.

Balkind:2016:OOS

[BMF^{+}16] Jonathan Balkind, Michael McKeown, Yaosheng Fu, Tri Nguyen, Yanqi Zhou, Alexey Lavrov, Mohammad Shahrad, Adi Fuchs, Samuel Payne, Xiaohua Liang, Matthew Matl, and David Wentzlaff. OpenPiton: an open source manycore research framework. Operating Systems Review, 50(2):217– 232, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Balkind:2019:OOS

[BMF^{+}19] Jonathan Balkind, Michael McKeown, Yaosheng Fu, Tri Nguyen, Yanqi Zhou, Alexey Lavrov, Mohammad Shahrad, Adi Fuchs, Samuel Payne, Xiaohua Liang, Matthew Matl, and David Wentzlaff. OpenPiton: an open source hardware platform for your research. Communications of the ACM, 62 (12):79–87, December 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://cacm.

acm.org/magazines/2019/12/241058/fulltext.

Benatto:2023:FCF

[BMR^{+}23] Leandro Benatto, Omar Mesquita, João L. B. Rosa, Lucimara S. Roman, Marlus Koehler, Rodrigo B. Capaz, and Graziâni Candiotto. FRET-Calc: a free software and web server for Förster Resonance Energy Transfer Calculation. Computer Physics Communications, 287(??):Article 108715, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000607.

Brooks:2022:ISO

[BMS^{+}22] Mollie E. Brooks, Valentina Melli, Esther Savina, Juan Santos, Russell Millar, Finbarr Gerard O’Neill, Tiago Veiga-Malta, Ludvig Ahm Krag, and Jordan Paul Feekings. Introducing selfisher: open source software for statistical analyses of fishing gear selectivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):1189–1197, August 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.
1139/cjfas-2021-0099.

Bonaccorso:2020:LPO

[BMT^{+}20] Fabio Bonaccorso, Andrea Montessori, Adriano Tiribocchi, Giorgio Amati, Massimo Bernaschi, Marco Lauricella, and Sauro Succi. LBsoft: a parallel open-source software for simulation of colloidal systems. Computer Physics Communications, 256(??):Article 107455, November 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302137.

Bertot:2002:PGS

[BMZ02] Yves Bertot, Nicolas Magaud, and Paul Zimmermann. A proof of GMP square root. Journal of Automated Reasoning, 29(3–4):225–252, September 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:/

/link.springer.com/article/10.1023/A:1021987403425.

Bauer:2014:ATM

[BMZ14] Rick Bauer, Ron Milford, and Li Zhen. Aligning technology and market drivers in an open source standards testing program. Computer, 47(11):30–36, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/csdl/mags/co/2014/11/
mco2014110030-abs.html.

Bollinger:1999:ROS

[BNST99] Terry Bollinger, Russell Nelson, Karsten Self, and Stephen Turnbull. Response: Open-source methods: Peering through the clutter. IEEE Software, 16(4):8–11, July/August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459

(electronic). URL http://dlib.computer.org/so/books/
so1999/pdf/s4006.pdf.

Bhowmik:2015:RSH

[BNSW15] Tanmay Bhowmik, Nan Niu, Prachi Singhania, and Wentao Wang. On the role of structural holes in requirements identification: an exploratory study on open-source software development. ACM Transactions on Management Information Systems (TMIS), 6(3):10:1–10:??, October 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Bollinger:2002:UFO

[Bol02] Terry Bollinger. Use of Free and Open-Source Software (FOSS) in the U.S. Department of Defense: Version: 1.2. Mitre report MP 02 W0000101, MITRE Corporation, Octo-

ber 28, 2002. 160 pp. URL http://www.egovos.org/pdf/
dodfoss.pdf.

Bajracharya:2014:SIL

[BOL14] Sushil Bajracharya, Joel Ossher, and Cristina Lopes. Sourcerer: an infrastructure for large-scale collection and analysis of open-source code. Science of Computer Programming, 79(??):241–259, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016764231200072X.

Baker:1997:LLA

[BOM97] T. P. Baker, Dong-Ik Oh, and Seung-Jin Moon. Low-level Ada tasking support for GNAT — performance and portability improvements. ACM SIGADA Ada Letters, 17(3):36–44, May/June 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bonzi:1993:APA

[Bon93] Susan Bonzi, editor. ASIS ’93: proceedings of the 56th ASIS annual meeting, 1993, volume 30, Columbus, Ohio, October 24–28, 1993. Learned Information, Medford, NJ, USA, 1993. ISBN 0-938734-78-4. LCCN Z699.A1 A5 1993.

Bonnet:2002:GPO

[Bon02] Philippe Bonnet. Going public: open-source databases and database research. In Franklin et al. [FMA02], page 633. ISBN ???? LCCN ???? ACM order number 475020.

Bonk:2011:FSW

[Bon11] Curtis J. Bonk. It’s a Free Software World After All: Opener #3: Availability of Open Source and Free Software, chapter 4, pages 139–161. Jossey-Bass, 2011. ISBN 1-118-26938-1.

Borenstein:1988:UER

[Bor88] N. S. Borenstein. UNIX emacs: a retrospective. lessons for flexible system design. In ACM [ACM88], pages 95–101. ISBN 0-89791-283-7. LCCN QA76.9.U83A26 1988.

Bortnikov:2009:OSG

[Bor09] Edward Bortnikov. Open-source grid technologies for webscale computing. ACM SIGACT News, 40(2):87–93, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bothner:2003:GCS

[Bot03] Per Bothner. GCC compile server. In Hutton et al. [HDR03], pages 21–34. ISBN ???? LCCN ???? URL http://www.

linux.org.uk/~ajh/gcc/gccsummit-2003-proceedings.pdf.

Boulanger:2005:OSV

[Bou05] A. Boulanger. Open-source versus proprietary software: Is one more secure and reliable than the other? IBM Systems Journal, 44(2):239–248, ???? 2005. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/
sj/442/boulanger.pdf.

Bowdidge:2005:RGU

[Bow05] Robert W. Bowdidge. Refactoring gcc using structure field access traces and concept analysis. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boyd:2000:PTH

[Boy00] Donald B. Boyd. Preface: A tribute to the halcyon days of QCPE. In Lipkowitz and Boyd [LB00], pages v–xvi. ISBN 0-471-36168-2 (hardcover), 0-470-12592-6 (e-book), 0-470-12619-1 (e-book). LCCN QD39.3.E46 R48 2000.

Boyd:2007:HCC

[Boy07] Donald B. Boyd. How computational chemistry became important in the pharmaceutical industry. In Lipkowitz [Lip07], chapter 7, pages 401–451. ISBN 0-470-08201-1 (hardcover), 0-470-11643-9 (e-book), 0-470-11644-7, 0-470-18907-X. LCCN QD39.3.E46 R54 v.23 2007. URL http://onlinelibrary.

wiley.com/doi/10.1002/9780470116449.ch7/summary.

Boykin:2008:IOS

[Boy08] Sam Boykin. Innovations: With open-source arms. Scientific American, 299(4):90–95, October 2008. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v299/n4/
full/scientificamerican1008-90.html;   http://www.
nature.com/scientificamerican/journal/v299/n4/pdf/
scientificamerican1008-90.pdf.

Boyd:2013:QCP

[Boy13] Donald B. Boyd. Quantum Chemistry Program Exchange, facilitator of theoretical and computational chemistry in pre-Internet history. In Strom and Wilson [SW13], pages 221–273. ISBN 0-8412-2716-0 (print), 0-8412-2717-9 (e-book). LCCN QD462.A1. Chapters based upon ten of the presentations at the Symposium “Pioneers of Quantum Chemistry” held March 28, 2011, at the 241st ACS National Meeting in Anaheim, CA, USA.

Breuer:2014:OSV

[BP14] Peter T. Breuer and Simon Pickin. Open source verification in an anonymous volunteer network. Science of Computer Programming, 91 (part B)(??):161–187, October 1, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167642313002037.

Becks:1994:NCT

[BPG94] K.-H. Becks and D. Perret-Gallix, editors. New computing techniques in physics research III: proceedings of the Third International Workshop on Software Engineering, Artificial Intelligence and Expert Systems for High Energy and Nuclear Physics: October 4–8, 1993, Oberammergau, Germany. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1994. ISBN 981-02-1699-8. LCCN QC793.47.E4I58 1993.

Baumgartner:1995:SLE

[BR95] G. Baumgartner and V. F. Russo. Signatures: a language extension for improving type abstraction and subtype polymorphism in C++. Software—Practice and Experience, 25 (8):863–889, August 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bonaccorsi:2003:WOS

[BR03] Andrea Bonaccorsi and Cristina Rossi. Why Open Source software can succeed. Research Policy, 32(7):1243–1258, 2003. CODEN ???? ISSN 0048-7333 (print), 1873-7625 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0048733303000519.

Branagan:1992:BRG

[Bra92] Linda Branagan. Book review: GNU Emacs. ;login: the USENIX Association newsletter, 17(2):34–??, March/April 1992. CODEN LOGNEM. ISSN 1044-6397.

Bramley:1997:TNF

[Bra97] Randall Bramley. Technology news: Fortran 90 news; free software; symbolic computing packages; Matlab 5; Web products. IEEE Computational Science & Engineering, 4(1): 87, 89–90, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

//dlib.computer.org/cs/books/cs1997/pdf/c1087.pdf;
 http://www.cs.berkeley.edu/~xiaoye;   http://www.
cs.cornell.edu/home/vavasis/qmg-home.html;   http:
//www.globus.org/sage; http://www.macsyma.com/;   http:
//www.mathworks.com/; http://www.netlib.org/benchmark/
linpackjava;   http://www.netlib.org/scalapack;   http:
//www.ucmp.berkeley.edu/subway/phylogen.html;   http:
//www.vni.com/products/wpd/jnl/jnl_1_0.html;   http:
//www.wolfram.com/look/cse.

Bramer:2004:DGL

[Bra04] Michael Bramer. Debian GNU LINUX 2.2: kompromisslos in Sicherheit & Stabilität: über 4000 Programmpakete: ausgereifte Desktop- & Server-Lösungen: mit Support & Handbuch. (German) [Debian GNU LINUX 2.2: Uncompromising in Security and Stability: Over 4000 program packages: Mature Desktop and Server Solutions: With Support and Handbook]. LinuxLand Internationale, Müenchen, Germany, 2004. ISBN 3-00-005785-4. ???? pp. LCCN ????

Balka:2010:HOO

[BRH10] Kerstin Balka, Christina Raasch, and Cornelius Herstatt. How open is open source? — software and beyond. Creativity and Innovation Management, 19(3):248–256, 2010. ISSN 0963-1690 (print), 1467-8691 (electronic).

Briot:2009:GSCa

[Bri09a] Emmanuel Briot. Gem #52: scripting capabilities in GNAT (part 1). ACM SIGADA Ada Letters, 29(2):37–39, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2009:GSCb

[Bri09b] Emmanuel Briot. Gem #54: scripting capabilities in GNAT (part 2). ACM SIGADA Ada Letters, 29(2):40–42, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Browning:1996:MFS

[Bro96] John Browning. Making free software pay. Scientific American, 274(1):35, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v274/n1/pdf/
scientificamerican0196-35a.pdf.

Brod:2001:SOS

[Bro01] Cesar Brod. SAGU: Open-source unified management system for educational institutions. In USENIX [USE01a], page ????

ISBN ???? LCCN ???? URL http://www.linuxshowcase.

org/brod.html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Brook:2003:FSG

[Bro03] Paul Brook. Fortran 95 support in GCC. In Hutton et al. [HDR03], pages 35–42. ISBN ???? LCCN ???? URL http://

www.linux.org.uk/~ajh/gcc/gccsummit-2003-proceedings.
pdf.

Brod:2004:ESB

[Bro04] Cesar Brod. EOF: SOLIS, a Brazilian free software cooperative. Linux Journal, 2004(120):12, April 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Brown:2005:EFS

[Bro05] Peter Brown. EOF: The Free Software Foundation at 20. Linux Journal, 2005(137):15, September 2005. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Brosgol:2019:HSS

[Bro19] B. M. Brosgol. How to succeed in the software business while giving away the source code: The AdaCore experience. IEEE Software, 36(6):17–22, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Burgess:1998:MFA

[BS98] Mark Burgess and Demosthenes Skipitaris. Managing filesystem ACLs with GNU/Cfengine. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN

1044-6397. URL http://www.usenix.org/publications/
login/1998-6/acls.html.

Borntrager:2005:PLS

[BS05] C. Bornträger and M. Schwidefsky. Providing Linux 2.6 support for the zSeries platform. IBM Systems Journal, 44 (2):331–340, ???? 2005. CODEN IBMSA7. ISSN 0018-

8670. URL http://www.research.ibm.com/journal/sj/
442/borntraeger.pdf.

Barbosa:2014:SCO

[BS14] Luís Soares Barbosa and Siraj Ahmed Shaikh. Selected contributions from the Open Source Software Certification (OpenCert) workshops. Science of Computer Programming, 91 (part B)(??):139–140, October 1, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167642314001713.

Bouktif:2014:PSO

[BSA14] Salah Bouktif, Houari Sahraoui, and Faheem Ahmed. Predicting stability of open-source software systems using combination of Bayesian classifiers. ACM Transactions on Management Information Systems (TMIS), 5(1):3:1–3:??, April 2014. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Bock:2022:MMG

[BSA22] Thomas Bock, Angelika Schmid, and Sven Apel. Measuring and modeling group dynamics in open-source software development: a tensor decomposition approach. ACM Transactions on Software Engineering and Methodology, 31(2):19:1–19:50, April 2022. CODEN ATSMER. ISSN 1049-331X (print),

1557-7392 (electronic). URL https://dl.acm.org/doi/10.
1145/3473139.

Bryngelson:2021:MOS

[BSC^{+}21] Spencer H. Bryngelson, Kevin Schmidmayer, Vedran Coralic, Jomela C. Meng, Kazuki Maeda, and Tim Colonius. MFC: an open-source high-order multi-component, multi-phase, and multi-scale compressible flow solver. Computer Physics Communications, 266(??):Article 107396, September 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301818.

Barcomb:2022:MEV

[BSFR22] Ann Barcomb, Klaas-Jan Stol, Brian Fitzgerald, and Dirk Riehle. Managing episodic volunteers in free/libre/open source software communities. IEEE Transactions on Software Engineering, 48(1):260–277, January 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Bullinger:1987:HII

[BSK87] H.-J. Bullinger, B. Shackel, and K. Kornwachs, editors. Human-computer Interaction — INTERACT ’87: proceedings of the Second IFIP Conference on Human-Computer Interaction, held at the University of Stuttgart, Federal Republic of Germany, 1–4 September, 1987. North-Holland Publishing Co., Amsterdam, The Netherlands, 1987. ISBN 0-444-70304-7. LCCN QA76.9.S88I325 1987.

Billings:2015:CAC

[BSK^{+}15] Paul C. Billings, Jenine K. Sanzari, Ann R. Kennedy, Keith A. Cengel, and John T. Seykora. Comparative analysis of colorimetric staining in skin using open-source software. Experimental Dermatology, 24(2):157–159, 2015. ISSN 0906-6705 (print), 1600-0625 (electronic).

Brown:2011:GPI

[BSP11] Frank C. Brown, Carmel H. Schindelhaim, and Suzanne R. Pfeffer. GCC185 plays independent roles in Golgi structure maintenance and AP-1-mediated vesicle tethering. Journal of Cell Biology, 194(5):779–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/194/5/779.

Barstow:1984:IPE

[BSS84] David R. Barstow, Howard E. Shrobe, and Erik Sandewall. Interactive Programming Environments. McGraw-Hill, New York, NY, USA, 1984. ISBN 0-07-003885-6. xii + 609 pp. LCCN QA76.6 .I5251 1984. US$34.95.

Batchelor:1995:MVA

[BSW95] Bruce G. Batchelor, Susan Snell Solomon, and Frederick M. Waltz, editors. Machine vision applications, architectures, and systems integration IV: Conference — October 1995, Philadelphia, PA, volume 2597 of Proceedings — SPIE The International Society for Optical Engineering 1995. SPIE Optical Engineering Press, Bellingham, WA, USA, 1995. ISBN 0-8194-1961-3. ISSN 0361-0748. LCCN TS510.S63 v.2597.

Bernabeu:2014:CCS

[BSW^{+}14] Miguel O. Bernabeu, James Southern, Nicholas Wilson, Peter Strazdins, Jonathan Cooper, and Joe Pitt-Francis. Chaste: a case study of parallelisation of an open source finite-element solver with applications to computational cardiac electrophysiology simulation. The International Journal of High Performance Computing Applications, 28(1):13–32, February 2014. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.sagepub.com/content/28/1/13.
full.pdf+html.

Babaloukas:2011:ETM

[BTL^{+}11] Georgios Babaloukas, Nicholas Tentolouris, Stavros Liatis, Alexandra Sklavounou, and Despoina Perrea. Evaluation of three methods for retrospective correction of vignetting on medical microscopy images utilizing two open source software tools. Journal of Microscopy, 244(3):320–324, 2011. ISSN 0022-2720 (print), 1365-2818 (electronic).

Buck:1982:EP

[Buc82a] Barbara Buck. EMACS primer. Prime Computer, Inc., Framingham, MA, USA, 1982. various pp.

Buck:1982:ERG

[Buc82b] Barbara Buck. EMACS reference guide. Prime Computer, Inc., Framingham, MA, USA, 1982. various pp.

Buddenhagen:2010:GRF

[Bud10] Olaf Buddenhagen. Gerätetreiber-Reengineering für Microkernel-Betriebssysteme am Beispiel eines Linux-KGI-Treibers für den Microkernel GNU Hurd (2010). (German) [Device-driver reengineering of microkernel operating systems using the example of a Linux driver for the KGI microkernel GNU Hurd (2010)]. Diplom Arbeit, Hochschule für Technik und Wirtschaft, Berlin, Germany, 2010. iv + 216 pp.

Bundy:1994:ADC

[Bun94] Alan Bundy, editor. Automated deduction, CADE-12: 12th International Conference on Automated Deduction, Nancy, France, June 26–July 1, 1994: proceedings, number 814 in Lecture Notes in Computer Science 1994. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58156-1, 0-387-58156-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D337I58 1994.

Bunks:2000:GG

[Bun00] Carey Bunks. Grokking the GIMP. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 0-7357-0924-6. xi + 342 pp. LCCN T385 .B852 2000. US$45.00.

Burgess:1995:GCE

[Bur95] Mark Burgess. The GNU configuration engine. Report 4, Høgskolen i Oslo, Avdeling for ingeniørutdanning, Oslo, Norway, 1995. ISBN 82-579-0104-0. 16 pp.

Burrell:2004:LEF

[Bur04a] Quentin L. Burrell. Letters to the Editor: Fitting Lotka’s Law: Some cautionary observations on a recent paper by Newby et al. (2003). Journal of the American Society for Information Science and Technology: JASIST, 55(13):1209–1210, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [NGJ03].

Burton:2004:BPI

[Bur04b] Andrew Burton. Building panoramic images in the GIMP. Linux Journal, 2004(119):8, March 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Busch:1999:GGG

[Bus99] David D. Busch. Great graphics with GIMP. Prima Publishing, Roseville, CA, USA, 1999. ISBN 0-7615-2407-X. xxx + 370 pp. LCCN T385 .B8664 2000. US$40.00. Edited by Kevin Harreld and Kim Spilker.

Butt:1994:RDS

[But94] Farooq Butt. Rapid development of a source-level debugger for PowerPC microprocessors. ACM SIGPLAN Notices, 29 (12):73–77, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Butt:1995:IFS

[But95] Farooq Butt. Implementing FORTRAN77 support in the GNU gdb debugger. ACM SIGPLAN Notices, 30(5):29–36, May 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ballance:1987:PII

[BV87] R. A. Ballance and M. L. Van De Vanter. Pan I: An introduction for users. Technical Report UCB/CSC 88/410, EECS-University of California, Berkeley, CA, September 1987. 60 pp.

Bevins:2014:WDO

[BVLF14] N. Bevins, M. Vanderhoek, S. Lang, and M. Flynn. WE-D-9A-06: Open source monitor calibration and quality control software for enterprise display management. Medical Physics, 41(6Part29):501, 2014. CODEN MPHYA6. ISSN 2473-4209.

Bakker:2006:WAD

[BVT06] Arno Bakker, Maarten Van Steen, and Andrew S. Tanenbaum. A wide-area Distribution Network for free software. ACM Transactions on Internet Technology (TOIT), 6(3):259–281, August 2006. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Beckman:2000:OSM

[BW00] Pete Beckman and Gregory V. Wilson. Open Source meets big iron. Dr. Dobb’s Journal of Software Tools, 25(6):30, 32, 34–35, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Boyer:1991:ACP

[BY91] Robert S. Boyer and Yuan Yu. AUTOMATED CORRECT-NESS PROOFS OF MACHINE CODE PROGRAMS FOR A COMMERCIAL MICROPROCESSOR. Technical Report TR-91-33, University of Texas, Austin, Austin, TX, USA, November 1991. 15 pp. prize (\ $1.50).

Boyer:1992:ACP

[BY92] R. S. Boyer and Yuan Yu. Automated correctness proofs of machine code programs for a commercial microprocessor. In Kapur [Kap92], pages 416–430. ISBN 3-540-55602-8. LCCN QA76.9.A96I57 1992.

Barshan:2014:RDS

[BY14] Billur Barshan and Murat Cihan Yüksek. Recognizing daily and sports activities in two open source machine learning environments using body-worn sensor units. The Computer Journal, 57(11):1649–1667, November 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/57/11/1649.

Ben-Yehuda:2008:OSF

[BYV08] Muli Ben-Yehuda and Eric Van Hensbergen. Open source as a foundation for systems research. Operating Systems Review, 42(1):2–4, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bricteux:2017:VSO

[BZB17] L. Bricteux, S. Zeoli, and N. Bourgeois. Validation and scalability of an open source parallel flow solver. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Staff:2009:LET

[CAC09] CACM Staff. Letters to the editor: True seeds of open source software. Communications of the ACM, 52(1):6, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Calcote:2010:APG

[Cal10] John Calcote. Autotools: a practitioner’s guide to GNU Autoconf, Automake, and Libtool. No Starch Press, San Francisco, CA, USA, 2010. ISBN 1-59327-206-5 (paperback). ???? pp. LCCN QA76.76.D47 C335 2010.

Cameron:1999:GEP

[Cam99] Debra Cameron. GNU Emacs Pocket Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-496-7. iii + 58 pp. LCCN QA76.76.T49 C348 1998. US$6.95. URL http://www.

oreilly.com/catalog/gnupr.

Cameron:2000:GEK

[Cam00] Debra Cameron. GNU Emacs — kurz & gut. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 3-89721-211-0. 60 pp. LCCN ???? German translation of [Cam99].

Cap:2012:BOS

[Cap12] Clemens H. Cap. Bitcoin das Open-Source-Geld. (German) [Bitcoin the open-source gold]. HMD Praxis der Wirtschaftsinformatik, 49(1):??, 2012. URL http://link.

springer.com/article/10.1007/BF03340666.

Capiluppi:2013:SCO

[Cap13] Andrea Capiluppi. Similarities, challenges and opportunities of Wikipedia content and open source projects. Journal of Software: Evolution and Process, 25(9):891–914, September 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Carr:1989:GTP

[Car89] D. J. Carr. Glue: a tree-based program development and maintenance system which uses explicit, typed, higher order cliches. Technical Report UIUCDCS-R-89-1495, University of Illinois at Urbana-Champaign, Urbana-Champaign, IL, USA, February 1989. 50 pp.

Carlson:2001:OSB

[Car01] R. Carlson. Open-source biology and its impact on industry. IEEE Spectrum, 38(5):15–17, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carlini:2004:PWL

[Car04] Paolo Carlini. Performance work in the libstdc++-v3 project. In Hutton et al. [HDR04], pages 17–24. ISBN ????

LCCN ???? URL http://people.redhat.com/lockhart/.
gcc04/MasterGCC-2side.pdf.

Cass:2002:FFR

[Cas02] Stephen Cass. Free as in freedom: Richard Stallman’s crusade for free software, by Sam Williams, O’Reilly, Sebastopol, CA, 2002, ISBN 0-596-00287-4, 240 pp., $22.96 [book review]. IEEE Spectrum, 39(6):56–57, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2019:FOS

[Cas19] Stephen Cass. The faces of open source: The humans behind the code — [resources geek life]. IEEE Spectrum, 56(3):18– 19, March 2019. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chowdhury:2022:UTP

[CASA22] Md Atique Reza Chowdhury, Rabe Abdalkareem, Emad Shihab, and Bram Adams. On the untriviality of trivial packages: an empirical study of npm JavaScript packages. IEEE Transactions on Software Engineering, 48(8):2695–2708, August 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Cao:2022:OSC

[CAWK22] Z. Cao, M. B. Agir, C. White, and K. Kontis. An open source code for two-phase rarefied flows: rarefiedMultiphaseFoam. Computer Physics Communications, 276(??):Article 108339, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000571.

Cuppens-Boulahia:2012:PTO

[CB12] Nora Cuppens-Boulahia. Penetration tester’s open source toolkit. Computers & Security, 31(4):630–632, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000600.

Christian:2006:OSX

[CBB06] Wolfgang Christian, Mario Belloni, and Douglas Brown. An open-source XML framework for authoring curricular material. Computing in Science and Engineering, 8(5):51–58, September/October 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cid:2022:OSF

[CBRSH22] Clara Cid, Aitor Baldomir, Miguel Rodríguez-Segade, and Santiago Hernández. An open-source framework for aircraft damage simulation in engine failure events. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S1877750322000904.

Chonacky:2003:SED

[CC03] Norman Chonacky and Dante Choi. Science and engineering databases in an open-source software world. Computing in Science and Engineering, 5(3):10–13, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.computer.org/comp/mags/cs/
2003/03/c3010abs.htm;   http://csdl.computer.org/dl/
mags/cs/2003/03/c3010.htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/c3010.pdf.

Cafarella:2004:BNO

[CC04] Mike Cafarella and Doug Cutting. Building Nutch: Open source search. ACM Queue: Tomorrow’s Computing Today, 2(2):54–61, April 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Colyer:2005:AOP

[CC05] A. Colyer and A. Clement. Aspect-oriented programming with AspectJ. IBM Systems Journal, 44(2):301–308, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://

www.research.ibm.com/journal/sj/442/colyer.pdf.

CCA:1984:CEMb

[CCA84] CCA Uniworks, Cambridge, MA, USA. CCA EMACS: manual for UNIX users, version 162.45z edition, October 1984. various pp.

Canis:2013:LOS

[CCA^{+}13] Andrew Canis, Jongsok Choi, Mark Aldham, Victor Zhang, Ahmed Kammoona, Tomasz Czajkowski, Stephen D. Brown, and Jason H. Anderson. LegUp: an open-source high-level synthesis tool for FPGA-based processor/accelerator systems. ACM Transactions on Embedded Computing Systems, 13(2): 24:1–24:??, September 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Celio:2019:BOS

[CCA^{+}19] C. Celio, P. Chiu, K. Asanovic, B. Nikolic, and D. Patterson. BROOM: An open-source out-of-order processor with resilient low-voltage operation in 28-nm CMOS. IEEE Micro, 39(2): 52–60, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cimatti:2002:NOT

[CCG^{+}02] Alessandro Cimatti, Edmund Clarke, Enrico Giunchiglia, Fausto Giunchiglia, Marco Pistore, Marco Roveri, Roberto Sebastiani, and Armando Tacchella. NuSMV 2: An OpenSource tool for symbolic model checking. Lecture Notes in Computer Science, 2404:359–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.de/link/service/
series/0558/bibs/2404/24040359.htm;   http://link.
springer.de/link/service/series/0558/papers/2404/24040359.
pdf.

Marimuthu:2021:HDO

[CCK21] Marimuthu C, Sridhar Chimalakonda, and Chandrasekaran K. How do open source app developers perceive API changes related to Android battery optimization? An empirical study. Software—Practice and Experience, 51(4):691–710, April 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Choi:2010:MFI

[CCSW10] Namjoo Choi, Indushobha Chengalur-Smith, and Andrew Whitmore. Managing first impressions of new open source software projects. IEEE Software, 27(6):73–77, November/ December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Carlson:1995:DDA

[CD95] William W. Carlson and Jesse M. Draper. Distributed data access in AC. ACM SIGPLAN Notices, 30(8):39–47, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Comar:1997:TGJ

[CDG97] C. Comar, G. Dismukes, and F. Gasperoni. Targeting GNAT to the Java Virtual Machine. In ACM [ACM97], pages 149–164. ISBN 0-89791-981-5. LCCN QA 76.73

A35 T75 1997. URL http://www.acm.org/pubs/contents/

proceedings/ada/269629/. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Cranor:1999:OSR

[CdR99] Charles D. Cranor and Theo de Raadt. Opening the source repository with anonymous CVS. In USENIX [USE99], page ?? ISBN 1-880446-33-2. LCCN ???? URL http://

www.openbsd.org/papers/anoncvs-paper.ps.

Crespo:2015:DOS

[CDR^{+}15] A. J. C. Crespo, J. M. Domínguez, B. D. Rogers, M. Gómez-Gesteira, S. Longshaw, R. Canelas, R. Vacondio, A. Barreiro, and O. García-Feal. DualSPHysics: Open-source parallel CFD solver based on Smoothed Particle Hydrodynamics (SPH). Computer Physics Communications, 187(??):204–216, February 2015. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465514003397.

Collins:2000:LMW

[CDsJ^{+}00] Terry Collins, Geoff Davis, Tarjei Tjxstheim [sic] Jensen, Raj Dash, DDJ, Rick Box, Dimitrios Souflis, James K. Yun, Guy Hammond, and DDJ. Letters: More worker shortage; music city; sorting through sorts; buy the book; E-address change; Open Source debate. Dr. Dobb’s Journal of Software Tools, 25(6):10, 12, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Currion:2007:OSS

[CdSV07] Paul Currion, Chamindra de Silva, and Bartel Van de Walle. Open source software for disaster management. Communications of the ACM, 50(3):61–65, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Caliari:2010:PFI

[CDSV10] Marco Caliari, Stefano De Marchi, Alvise Sommariva, and Marco Vianello. Padua2DM: fast interpolation and cubature at the Padua points in Matlab/Octave. Numerical Algorithms, 54(??):??, ???? 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=0&issue=0&spage=??.

Caliari:2011:PFI

[CDSV11] Marco Caliari, Stefano De Marchi, Alvise Sommariva, and Marco Vianello. Padua2DM: fast interpolation and cubature at the Padua points in Matlab/Octave. Numerical Algorithms, 56(1):45–60, January 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=1&spage=45.

Cameron:2005:LGE

[CEL^{+}05] Debra Cameron, James Elliot, Marc Loy, Eric Raymond, and Bill Rosenblatt. Learning GNU Emacs. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2005. ISBN 0-596-00648-9. xxii + 509 pp. LCCN

QA76.76.T49 C35 2005. URL ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/.

Clement-Fontaine:1998:LPG

[CF98] Mélanie Clément-Fontaine. La license publique générale GNU [logiciel libre]. Mémoire de d.e.a droit des créations immatérielles, Université de Montpellier, I Faculté de Droit, Montpellier, France, 1998. URL http://www.juriscom.net/

uni/mem/08/log01.pdf.

Camara:2007:IPO

[CF07a] Gilberto Câmara and Frederico Fonseca. Information policies and open source software in developing countries. Journal of the American Society for Information Science and Technology: JASIST, 58(1):121–132, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cerri:2007:OSO

[CF07b] Davide Cerri and Alfonso Fuggetta. Open standards, open formats, and open source. The Journal of Systems and Software, 80(11):1930–1937, November 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Colazo:2009:ILC

[CF09] Jorge Colazo and Yulin Fang. Impact of license choice on Open Source Software development activity. Journal of the American Society for Information Science and Technology: JASIST, 60(5):997–1011, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Carlberg:2013:CGM

[CFCA13a] Kevin Carlberg, Charbel Farhat, Julien Cortial, and David Amsallem. Corrigendum to “The GNAT method for nonlinear model reduction: Effective implementation and application to computational fluid dynamics and turbulent flows” [J. Comput. Phys. 242 (2013) 623–647]. Journal of Computational Physics, 250(??):713, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003628. See [CFCA13b].

Carlberg:2013:GMN

[CFCA13b] Kevin Carlberg, Charbel Farhat, Julien Cortial, and David Amsallem. The GNAT method for nonlinear model reduction: Effective implementation and application to computational fluid dynamics and turbulent flows. Journal of Computational Physics, 242(??):623–647, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001472. See corrigen-

dum [CFCA13a].

Capek:2005:HIO

[CFGS05] P. G. Capek, S. P. Frank, S. Gerdt, and D. Shields. A history of IBM’s open-source involvement and strategy. IBM Systems Journal, 44(2):249–257, ???? 2005. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/
sj/442/capek.pdf.

Casari:2023:BRB

[CFL23] Amanda Casari, Julia Ferraioli, and Juniper Lovato. Beyond the repository: Best practices for open source ecosystems researchers. ACM Queue: Tomorrow’s Computing Today, 21 (2):14–34, March 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https://dl.acm.org/

doi/10.1145/3595879.

Capra:2008:ESR

[CFM08] E. Capra, C. Francalanci, and F. Merlo. An empirical study on the relationship between software design quality, development effort and governance in open source projects. IEEE Transactions on Software Engineering, 34(6):765–782, November/ December 2008. CODEN IESEDJ. ISSN 0098-5589 (print),

1939-3520 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=4599582.

Capra:2011:FIO

[CFMRL11] Eugenio Capra, Chiara Francalanci, Francesco Merlo, and Cristina Rossi-Lamastra. Firms’ involvement in Open Source projects: a trade-off between software structural quality and popularity. The Journal of Systems and Software, 84(1): 144–161, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Claypool:2001:OSL

[CFW01] Mark Claypool, David Finkel, and Craig Wills. An open source laboratory for operating systems projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):145–148, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chilenski:2017:EME

[CFW17] M. A. Chilenski, I. C. Faust, and J. R. Walk. eqtools: Modular, extensible, open-source, cross-machine Python tools for working with magnetic equilibria. Computer Physics Communications, 210(??):155–162, January 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551630282X.

Ching:2017:XOS

[CG17] Daniel J. Ching and Dogă Gürsoy. XDesign: an open-source software package for designing X-ray imaging phantoms and experiments. Journal of Synchrotron Radiation, 24(2):537– 544, 2017. ISSN 0909-0495 (print), 1600-5775 (electronic).

Crawford:2005:FDU

[CGB^{+}05] Diane Crawford, Rajesh Gupta, Ashley Braganza, Raymond L. Robert, Hal Berghel, Richard Stallman, Michael Cusumano, Ephraim McLean, and Ralph Westfall. Forum: DARPA (and U.S.) opportunities lost; don’t ignore the CIO’s legal burdens; free is not open software; LEO lives; resolving the dataless dilemma in OO programming. Communications of the ACM, 48(7):11–13, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:FPW

[CGK^{+}02] Diane Crawford, Aram Gharib, Ken Kahn, B. Jagannathan, Richard Stallman, and Jim Densmore. Forum: In praise of work’s human dimension; make it visible, tangible, graphical; justifying anonymity; free software reality v. deception; already maxed out. Communications of the ACM, 45(7):11– 12, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Comar:1994:GPG

[CGS94] C. Comar, F. Gasperoni, and E. Schonberg. The GNAT project: a GNU-ADA94 compiler. Technique et Science Informatiques, 13(6):817–836, 1994. CODEN TTSIDJ. ISSN 0752-4072 (print), 2116-5920 (electronic).

Caneill:2017:DDT

[CGZ17] Matthieu Caneill, Daniel M. Germán, and Stefano Zacchiroli. The Debsources Dataset: two decades of free and open source software. Empirical Software Engineering, 22(3):1405– 1437, June 2017. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
article/10.1007/s10664-016-9461-5.

Conte:1991:BC

[CH91] T. M. Conte and W. W. Hwu. Benchmark characterization. In Milutinovic and Shriver [MS91], pages 365–372 (vol. 1). 4 vol.

Cook:2006:IEP

[CH06a] Ian Cook and Gavin Horobin. Implementing egovernment without promoting dependence: open source software in developing countries in Southeast Asia. Public Administration and Development, 26(4):279–289, 2006. ISSN 0271-2075 (print), 1099-162X (electronic).

Crowston:2006:AHO

[CH06b] Kevin Crowston and James Howison. Assessing the health of open source communities. Computer, 39(5):89–91, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chan:2010:DAK

[CH10] Johnny Chan and Kenneth Husted. Dual allegiance and knowledge sharing in open source software firms. Creativity and Innovation Management, 19(3):314–326, 2010. ISSN 0963-1690 (print), 1467-8691 (electronic).

Chou:2011:FAP

[CH11] Shih-Wei Chou and Mong-Young He. The factors that affect the performance of open source software development — the perspective of social capital and expertise integration. Information Systems Journal, 21(2):195–219, 2011. ISSN 1350-1917 (print), 1365-2575 (electronic).

Chamberlain:1991:LBB

[Cha91] Steve Chamberlain. LIB BFD, the Binary File Descriptor library. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1991. ISBN ???? ???? pp. LCCN ????

Charles:1992:LUP

[Cha92] H.-P. Charles. Loop unrolling for processors with instruction cache. In Quinton and Robert [QR92], pages 311–316. ISBN 0-444-89153-6. LCCN QA76.58.I57 1991.

Chassell:1997:PEL

[Cha97] Robert J. Chassell. Programming in Emacs Lisp: An Introduction. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, October 21, 1997. ISBN 1-882114-41-8. ???? pp. LCCN ????

Charles:1998:NOS

[Cha98] John Charles. In the news: Open source: Netscape pops the hood: NSF grant funds national network: Engineering center. IEEE Software, 15(4):79–82, July/August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459

(electronic). URL http://dlib.computer.org/so/books/
so1998/pdf/s4079.pdf.

Chassell:19xx:PEL

[Chaxx] Robert J. Chassell. Programming in Emacs Lisp: An Introduction. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 19xx. ISBN 1-882114-42-6. ???? pp. LCCN ????

Chan:2001:DID

[Cha01a] Jason Chan. Distributed intrusion detection with open source tools. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):20, 22, 24–25, August 2001. CODEN SYADE7.

ISSN 1061-2688. URL http://www.samag.com/.

Chassell:2001:SFC

[Cha01b] Robert Chassell. Software freedom: Creation and danger,

2001. URL http://www.linuxshowcase.org/tech.html.

Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Chassell:2001:IPE

[Cha01c] Robert J. Chassell. An Introduction to Programming in Emacs Lisp. GNU Press, Boston, MA, USA, 2001. ISBN 1-882114-43-4. 320 (est.) pp. LCCN ???? US$30. URL http://www.

gnupress.org/book4.html.

Chassell:2004:IPE

[Cha04] Robert J. Chassell. An Introduction to Programming in Emacs Lisp. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2004. ISBN 1-882114-56-6. ???? pp. LCCN ????

Crowston:2006:ISS

[CHA06] Kevin Crowston, James Howison, and Hala Annabi. Information systems success in free and open source software development: theory and measures. Software Process: Improvement and Practice, 11(2):123–148, March 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Charette:2007:OSW

[Cha07] R. N. Charette. Open-source warfare. IEEE Spectrum, 44(11): 26–32, November 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chao:2011:OSM

[Cha11] Lee Chao, editor. Open source mobile learning: mobile Linux applications. Information Science Reference, Hershey, PA, USA, 2011. ISBN 1-60960-613-2. ???? pp. LCCN LB1044.87 .O64 2011.

Chapelle:2013:BRO

[Cha13] Gregory Chapelle. Book review: The official Ubuntu book, seventh edition by Matthew Helmke and Amber Graner. ACM SIGSOFT Software Engineering Notes, 38(1):54–55, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:1986:GEB

[Che86] Pehong Chen. GNU Emacs BibTEX mode: version 1.5. Technical report UCB/CSD 87/317, University of California. Computer Science Division, Berkeley, CA, USA, 1986. ii + 34 pp.

Chen:1987:GEB

[Che87a] Pehong Chen. GNU Emacs BibTEX-mode. Technical Report UCB/CSD 87/317, University of California, Berkeley, Berkeley, CA, USA, 1987. ii + 57 pp. US$3.50.

Chen:1987:GET

[Che87b] Pehong Chen. GNU Emacs TEX mode: version 1.5. Report UCB/CSD 87/316, University of California, Computer Science Department, Berkeley, CA, USA, 1987. ii + 57 pp.

Cheng:1995:ECA

[Che95] H. H. Cheng. Extending C with arrays of variable length. Computer Standards and Interfaces, 17(4):375–406, September 1995. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Chen:2010:EIO

[CHE^{+}10] Yang Chen, Yuanjie Huang, Lieven Eeckhout, Grigori Fursin, Liang Peng, Olivier Temam, and Chengyong Wu. Evaluating iterative optimization across 1000 datasets. ACM SIGPLAN Notices, 45(6):448–459, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Childs:1993:GER

[Chi93] Bart Childs. GNU Emacs reference card (with web-

mode). ftp.cs.tamu.edu:/pub/tex-web/web/docs, Texas

A&M University, College Station, TX, USA, 1993.

Chin:1997:BRW

[Chi97] Eric Chin. Book review: Writing GNU Emacs Extensions. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/publications/login/1997-
12/emacs.html.

Chiariglione:2001:OSM

[Chi01] Leonardo Chiariglione. Open source in MPEG. Linux Journal, 83:126–126, 130, 132, March 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Chopra:2009:BRB

[Cho09] Samir Chopra. Book review: Open Source: Technology and Policy. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2382–2383, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ciccarelli:1978:IEE

[Cic78] Eugene Charles Ciccarelli. An introduction to the EMACS editor. AI memo 447, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, 1978. 23 pp.

Capiluppi:2013:EEF

[CIC13] Andrea Capiluppi and Daniel Izquierdo-Cortázar. Effort estimation of FLOSS projects: a study of the Linux kernel. Empirical Software Engineering, 18(1):60–88, February 2013. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10664-011-9191-7.

Cioffi:2001:LHO

[Cio01] D. F. Cioffi. Learning from hackers [open-source programming]. IEEE Spectrum, 38(6):77–79, June 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:2017:CLP

[CJ17] Boyuan Chen and Zhen Ming (Jack) Jiang. Characterizing logging practices in Java-based open source software projects — a replication study in Apache Software Foundation. Empirical Software Engineering, 22(1):330–374, February 2017. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10664-016-9429-5.

Chen:2019:ESL

[CJ19] Boyuan Chen and Zhen Ming (Jack) Jiang. Extracting and studying the logging–code-issue-introducing changes in Java-based large-scale open source software systems. Empirical Software Engineering, 24(4):2285–2322, August 2019. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616

(electronic). URL http://link.springer.com/article/
10.1007/s10664-019-09690-0.

Chryselius:2006:DQE

[CK06a] Toralf Chryselius and Andrea Kuntz. Debian unter Qemu Einführung in das Betriebssystem Debian Linux in der virtuellen Umgebung Qemu unter Windows. (German) [Debian under Qemu: Introduction in the Debian Linux operating systems in the Qemu virtual machine under Windows], volume 17 of Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-116-3 (book), 3-86768-716-1 (DVD). 159 pp. LCCN ????

Chryselius:2006:IDQ

[CK06b] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in Debian unter Qemu Einführung in das Betriebssystem Debian Linux in Qemu und Vorstellung der wichtigsten Internetprogramme. (German) [Internet Communication in Debian under Qemu: Introduction in the Debian Linux operating system in Qemu and creation of the most important Internet programs], volume 18 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-117-1 (book), 3-86768-717-X (DVD). 109 pp. LCCN ????

Chryselius:2006:IKQa

[CK06c] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in Kubuntu unter Qemu Einführung in das Betriebssystem Kubuntu und Vorstellung von Internetprogrammen in der virtuellen Umgebung Qemu. (German) [Internet Communication in Kubuntu under Qemu: Introduction to the Kubuntu operating system and creation of Internet programs in the Qemu virtual machine], volume 6 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-105-8 (Buch), 3-86768-705-6 (DVD). 107 pp. LCCN ????

Chryselius:2006:KQE

[CK06d] Toralf Chryselius and Andrea Kuntz. Kubuntu unter Qemu Einführung in das Betriebssystem Kubuntu Linux in der virtuellen Umgebung Qemu, volume 5 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-104-X (book), 3-86768-704-8 (DVD). 158 pp. LCCN ????

Chryselius:2006:LDQ

[CK06e] Toralf Chryselius and Andrea Kuntz. Lernprogramme in Debian unter Qemu Einführung in das Betriebssystem Debian und Vorstellung von Lernprogrammen in der virtuellen Umgebung Qemu, volume 19 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-118-X (book), 3-86768-718-8 (DVD). 141 pp. LCCN ????

Chryselius:2006:LKQc

[CK06f] Toralf Chryselius and Andrea Kuntz. Lernprogramme in Kubuntu unter Qemu Einführung in das Betriebssystem Kubuntu und Vorstellung von Lernprogrammen in der virtuellen Umgebung Qemu, volume 7 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-106-6 (book), 3-86768-706-4 (DVD). 152 pp. LCCN ????

Chryselius:2006:SKD

[CK06g] Toralf Chryselius and Andrea Kuntz. Software für Kinder in Debian unter Qemu Einführung in das Betriebssystem Debian und Vorstellung der Lern- und Spielesammlung Gcompris in der virtuellen Umgebung Qemu, volume 20 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-119-8 (book), 3-86768-719-6 (DVD). 113 pp. LCCN ????

Chryselius:2006:SKKa

[CK06h] Toralf Chryselius and Andrea Kuntz. Software für Kinder in Kubuntu unter Qemu Einführung in das Betriebssystem Kubuntu und Vorstellung der Lern- und Spielesammlung GCompris in der virtuellen Umgebung Qemu, volume 8 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-107-4 (book), 3-86768-707-2 (DVD). 108 pp. LCCN ????

Campbell-Kelly:2007:HHS

[CK07] Martin Campbell-Kelly. The history of the history of software. IEEE Annals of the History of Computing, 29(4):40–51, October/December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Campbell-Kelly:2008:VHR

[CK08] Martin Campbell-Kelly. Viewpoints: Historical reflections: Will the future of software be open source? Communications of the ACM, 51(10):21–23, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Conley:2010:PBW

[CK10] John P. Conley and Fan-Chin Kung. Private benefits, warm glow, and reputation in the free and open source software production model. Journal of Public Economic Theory, 12 (4):665–689, 2010. ISSN 1097-3923 (print), 1467-9779 (electronic).

Crawford:2005:FBL

[CKB^{+}05] Diane Crawford, Brad Kidwell, Jeffrey P. Buzen, David A. Patterson, Bradley Dilger, Brit J. Williams, Merle S. King, Chris Morris, David H. Jameson, Ralph Johnson, I. Samoladas, I. Stamelos, and L. Angelis. Forum: Bandwidth, latency, and physical reality; share disturbing information on e-voting; ensure quality assurance for bioinformatics applications; follow the money; share open source sources. Communications of the ACM, 48(1):11–13, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cruz:2011:FMM

[CKB11] F. A. Cruz, M. G. Knepley, and L. A. Barba. Fast Multipole Method for particle interactions: an open source parallel library component. In Tromeur-Dervout et al. [TDBEE11], pages 285–292. CODEN LNCSA6. ISBN 3-642-14437-3 (print), 3-642-14438-1 (e-book). ISSN 1439-7358 (print), 2197-7100 (electronic). LCCN ???? URL http://link.springer.

com/content/pdf/10.1007/978-3-642-14438-7_30. Pro-

ceedings of the twentieth meeting, Parallel CFD 2008, held May 19–22, 2008 in Lyon, France.

Christopher:2022:HOS

[CKGW22] Peter J. Christopher, Andrew Kadis, George S. D. Gordon, and Timothy D. Wilkinson. HoloGen: an open-source toolbox for high-speed hologram generation. Computer Physics Communications, 270(??):Article 108139, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002514.

Chatterjee:1991:EOD

[CKH91] A. Chatterjee, A. Khanna, and Y. Hung. ES-kit: an objectoriented distributed system. Concurrency, practice and experience, 3(6):525–539, December 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Choi:2016:TDP

[CKS16] Seung Woo Choi, Hee Jung Kwon, and Won Kyung Song. Three-dimensional printing using open source software and JPEG images from optical coherence tomography of an epiretinal membrane patient. Acta Ophthalmologica, page ??, 2016. ISSN 1755-375X (print), 1755-3768 (electronic).

Clarkson:1990:PRE

[Cla90] M. Clarkson. Praxis: a rule-based expert system for MAC-SYMA. In Miola [Mio90], pages 264–265. ISBN 3-540-52531-9. LCCN QA76.9.S88I576 1990.

Chan:2005:OOS

[CLL05] Ben C. B. Chan, John C. F. Lau, and John C. S. Lui. OPERA: an open-source extensible router architecture for adding new network services and protocols. The Journal of Systems and Software, 78(1):24–36, October 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2008:ESS

[CLM^{+}08] Weibing Chen, Jingyue Li, Jianqiang Ma, Reidar Conradi, Junzhong Ji, and Chunnian Liu. An empirical study on software development with open source components in the Chinese software industry. Software Process: Improvement and Practice, 13(1):89–100, January 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Close:1989:GM

[Clo89] Diane Barlow Close. The GAWK manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 0.12 beta edition, October 1989. viii + 152 pp.

Close:1991:GM

[Clo91] Diane Barlow Close. The GAWK manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 0.13, version 2.13 edition, December 1991. vii + 186 pp.

Chen:1995:STP

[CLS95] Tzer-Shyong Chen, Feipei Lai, and Rung-Ji Shang. A simple tree pattern matching algorithm for code generator. In IEEE [IEE95b], pages 162–167. ISBN 0-8186-7119-X. LCCN QA76.6 .I5 1995. IEEE Catalog No. 95CB35838.

Chae:2006:ATA

[CM06] Bongsug (Kevin) Chae and Roger McHaney. Asian trio’s adoption of Linux-based open source development. Communications of the ACM, 49(9):95–99, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cantwell:2015:NOS

[CMC^{+}15] C. D. Cantwell, D. Moxey, A. Comerford, A. Bolis, G. Rocco, G. Mengaldo, D. De Grazia, S. Yakovlev, J.-E. Lombard, D. Ekelschot, B. Jordi, H. Xu, Y. Mohamied, C. Eskilsson, B. Nelson, P. Vos, C. Biotto, R. M. Kirby, and S. J. Sherwin. Nektar++: an open-source spectral/hp element framework. Computer Physics Communications, 192(??):205–219, July 2015. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465515000533.

Crawford:2004:FOS

[CMJ^{+}04] Diane Crawford, David Monniaux, Ralph Johnson, Clark E. Hartsock, III, Ken Nickerson, Robert L. Glass, Marc Bellusci, and Alex Simonelis. Forum: Open source vs. capitalism and communism; what worker shortage?; design for pedagogy not just for software. Communications of the ACM, 47(4):11– 13, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coppola:2019:SGT

[CMTA19] Riccardo Coppola, Maurizio Morisio, Marco Torchiano, and Luca Ardito. Scripted GUI testing of Android open-source apps: evolution of test code and fragility causes. Empirical Software Engineering, 24(5):3205–3248, October 2019. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616

(electronic). URL http://link.springer.com/article/
10.1007/s10664-019-09722-9.

Chakroborti:2023:RCO

[CNSR23] Debasish Chakroborti, Sristy Sumana Nath, Kevin A. Schneider, and Chanchal K. Roy. Release conventions of open-source software: an exploratory study. Journal of Software: Evolution and Process, 35(1):e2499:1–e2499:??, January 2023. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cesar:2012:EOS

[CO12] Pablo Cesar and Wei Tsang Ooi. Embracing open source multimedia software. IEEE MultiMedia, 19(2):11–15, May/June 2012. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Cochran:2001:NVS

[Coc01a] Shannon Cochran. News and views: Scientists seek immersive reality; USENIX names lifetime achievement recipients [the GNU Project and the Kerberos network authentication system]; robots need programmers; evangelizing the Semantic Web; get your supercomputer software free; Usenet creator Jim Ellis dies; DARPA funds FreeBSD security initiative. Dr. Dobb’s Journal of Software Tools, 26(9):18, September 2001.

CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.

com/.

Cochran:2001:NVW

[Coc01b] Shannon Cochran. News and views: WEP [wired equivalent privacy]: Pining for the fjords?; bye bye Be; UNIX utilities open sourced; cosmology computer calculates creation; UML 2.0 infrastructure proposals. Dr. Dobb’s Journal of Software Tools, 26(11):18, November 2001. CODEN DDJOEB. ISSN

1044-789X. URL http://www.ddj.com/.

Cochran:2003:NVGb

[Coc03] Shannon Cochran. News and views: Government ponders open-source strategy; POSIX, Single UNIX specification merged; IBM plans new supercomputers; robotic surgeons have a heart. Dr. Dobb’s Journal of Software Tools, 28(2):14, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/documents/s=7790/ddj0302o/.

Cody:1975:FPS

[Cod75] W. J. Cody. The FUNPACK package of special function subroutines. ACM Transactions on Mathematical Software, 1(1): 13–25, March 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cohen:1982:TSC

[Coh82] E. Cohen. Text-oriented structure commands for structure editors. ACM SIGPLAN Notices, 17(11):45–49, November 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:2002:OSM

[Coh02] Fred Cohen. Is open source more or less secure? Network Security, 2002(7):17–19, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1353485802070113.

Cohen:2003:EOS

[Coh03] Stuart Cohen. EOF: The Open Source Development Lab. Linux Journal, 2003(113):14, September 2003. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Collison:2002:RGC

[Col02] Michael Collison. Retargeting the GNU C compiler. C/C++ Users Journal, 20(6):30, 32–36, June 2002. CODEN CCU-JEX. ISSN 1075-2838.

Cole:2005:UMT

[Col05] Jason R. Cole. Using Moodle: teaching with the popular open source course management system. O’Reilly Community Press, Sebastopol, CA, USA, 2005. ISBN 0-596-00863-5. xv + 219 pp. LCCN LB1028.68 .C65 2005. URL http://

www.oreilly.com/catalog/9780596008635.

Coleman:2009:CSL

[Col09a] Gabriella Coleman. Code is speech: Legal tinkering, expertise, and protest among free and open source software developers. Cultural Anthropology, 24(3):420–454, 2009. ISSN 0886-7356 (print), 1548-1360 (electronic).

Colford:2009:EFO

[Col09b] Scot Colford. Explaining free and open source software. Bulletin of the American Society for Information Science and Technology, 35(2):10–14, 2009. ISSN 2373-9223.

CCA:1984:CEMa

[Com84] Computer Corporation of America, Cambridge, MA, USA. CCA EMACS manual for UNIX users, version 162.43z edition, 1984. various pp.

Comerford:1999:POS

[Com99] Richard Comerford. The path to open-source systems. IEEE Spectrum, 36(5):25–31, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Conklin:1987:SH

[Con87] Jeff Conklin. A survey of hypertext. MCC Technical Report STP-356-86, Rev. 2, Microelectronics and Computer Technology Corporation, Austin, Texas, December 3, 1987. 69 pp. MCC Software Technology Program unrestricted nonproprietary report.

Cooke:1991:EAS

[Coo91] Chris Cooke. Emacs — a survival guide, [9] 1991.

Cooperman:1995:SBP

[Coo95a] G. Cooperman. STAR/MPI: binding a parallel library to interactive symbolic algebra systems. In Levelt [Lev95a], pages 126–132. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Cooperman:1995:SMB

[Coo95b] Gene Cooperman. STAR/MPI: Binding a parallel library to interactive symbolic algebra systems. In Levelt [Lev95b], pages 126–132. ISBN 0-89791-699-9. LCCN A 76.95 I59 1995.

Polytron:polyawk

[Cor87] Polytron Corporation. PolyAWK, 1987. 170 NW 167th Place, Beaverton, OR 97006. See also [AKW88].

Coris:2000:CTP

[Cor00] Marie Coris. Compétition technologique et propriété intellectuelle: quels enjeux pour les logiciels libres ? le cas du système d’exploitation GNU/Linux. Master’s thesis, Mém. D.E.A.: écon. de l’environnement innovation et aménagement, Bordeaux 4, France, 2000.

Cornelio:2005:MLG

[Cor05] Pietro Cornelio. Il mondo libero di GNU/Linux e UNIX BSD: storia, filosofia, tecnologia. (Italian) [The free world of GNU/Linux and BSD UNIX: stories, philosophies, technology]. Duke Italia, Milano, Italy, 2005. ISBN 88-86460-10-4. 104 pp. LCCN ???? Suppement to Linux Journal (Italian edition), number 59, October 2005.

Courtes:2013:FPM

[Cou13] Ludovic Courtès. Functional package management with Guix. arxiv.org, ??(??):1–11, May 20, 2013. URL https://arxiv.

org/abs/1305.4584.

Courtes:2017:CSG

[Cou17] Ludovic Courtès. Code staging in GNU Guix. ACM SIG-PLAN Notices, 52(12):41–48, December 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coughlan:2020:SOS

[Cou20] S. Coughlan. Standardizing open source license compliance with OpenChain. Computer, 53(11):70–74, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cowan:2003:SSO

[Cow03] Crispin Cowan. Software security for open-source systems. IEEE Security & Privacy, 1(1):38–45, January/February 2003. CODEN ???? ISSN 1540-7993 (print), 1558-

4046 (electronic). URL http://dlib.computer.org/sp/
books/sp2003/pdf/j1038.pdf;   http://www.computer.
org/security/j1038abs.htm.

Casarini:2001:GDE

[CP01] Paolo Casarini and Luca Padovani. The Gnome DOM engine. Markup Languages: Theory & Practice, 3(2):173–190, April 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626

(electronic). URL http://mitpress.mit.edu/catalog/
item/default.asp?sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=8010.

Chavan:2004:OSL

[CP04] Abhijeet Chavan and Shireen Pavri. Open-source learning management with moodle. Linux Journal, 2004(128):2, December 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Cody:2004:LBO

[CPG^{+}04] Robert Cody, Lawrence C. Paulson, Todd Greer, Rodney Bates, Daniel Myer, Gary Hewitt, Jay Michaelson, John C. Nash, and Josh Coates. Letters: Buffer overrun madness; open source: All grown up?; when is free really Free?; it’s not cool to be a bigot. ACM Queue: Tomorrow’s Computing Today, 2(5):10, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Clinton:1998:LBM

[CPJ^{+}98] Terry Clinton, Tom Parsons, Capers Jones, William Adams, Garth Klatt, Eric Haines, Ted Lewis, Philip Machanik, Stig Nilsson, Karl Reed, Howard R. Stearns, Neville Holmes, and John Brownie. Letters: The benefits of model-based integration; documentation is not green; picking on the overdog; buggy, slow windoze; there’s no such thing as free software; Linus’ law of open source development; bug-free development? no way; governmental IT planning and the Computer Society; text encoding questions; encoding the world’s languages. Computer, 31(11):4, 5–7, 11, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814

(electronic). URL http://dlib.computer.org/co/books/

co1998/pdf/ry004.pdf. Two letters discuss Unicode and Multicode [Mud97].

Cameron:1991:LGE

[CR91] Debra Cameron and Bill Rosenblatt. Learning GNU Emacs. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-84-6. xxvii + 411 pp. LCCN QA76.76.T49 C35 1991. URL http://www.oreilly.

com/catalog/9780937175842.

Cameron:1992:LGE

[CR92a] Debra Cameron and Bill Rosenblatt. Learning GNU Emacs. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-84-6.

Carlini:1992:GCC

[CR92b] Giuliano Carlini and Susan Rendina. Garbage collection for C programs. Dr. Dobb’s Journal of Software Tools, 17(11): 46–??, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Crabb:1989:PWP

[Cra89] D. Crabb. A perfect word processor at last? BYTE Magazine, 14(5):157–158, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cramer:1990:SMC

[Cra90] M. Cramer. Structure and mnemonics in computer and command languages. International Journal of Man-Machine Studies, 32(6):707–722, June 1990. CODEN IJMMBC. ISSN 0020-7373.

Cerrato:2018:CCO

[CRB^{+}18] Ivano Cerrato, Fulvio Risso, Roberto Bonafiglia, Kostas Pentikousis, Gergely Pongrácz, and Hagen Woesner. COM-POSER: A compact open-source service platform. Computer Networks (Amsterdam, Netherlands: 1999), 139(??):151–174, July 5, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128618301750.

Creel:2007:RFM

[Cre07] Michael Creel. I ran four million probits last night: HPC clustering with ParallelKnoppix. Journal of Applied Econometrics, 22(1):215–223, January–February 2007. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Crowder:2000:BRG

[Cro00] Ben Crowder. Book review: GNOME/GTK+ programming. Linux Journal, 79:54, November 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.com/lj-issues/issue79/4098.
html.

Cameron:1996:LGE

[CRR96] Debra Cameron, Bill Rosenblatt, and Eric S. Raymond. Learning GNU Emacs. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, September 1996. ISBN 1-56592-152-6. xxiv + 533 pp. LCCN QA76.76.T49 C35 1996. URL http://www.

oreilly.com/catalog/gnu2.

Crawford:2004:FPL

[CRW^{+}04] Diane Crawford, Peter G. Raeth, Mark Wallace, David Hawley, Donald Anselmo, Henry Ledgard, Dwayne Nelson, Yitzchak Gottlieb, Alex Simonelis, Coskun Bayrak, and Chad Davis. Forum: Principles of logical thought before technical details; why reinvent when we have open source?; still seeking software productivity; spare me the self-service; lineage of the Internet’s Open Systems Foundation. Communications of the ACM, 47(3):11–13, March 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chassell:1991:TGD

[CS91] Robert J. Chassell and Richard Stallman. Texinfo: the GNU documentation format: edition 2.04. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, second edition, April 1991. viii + 220 pp.

Chassell:1993:TGD

[CS93] Robert J. Chassell and Richard Stallman. Texinfo: the GNU documentation format. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2.19, for Texinfo version three edition, December 1993. ISBN 1-882114-46-9. x + 242 pp. LCCN ????

Chassell:1995:TGD

[CS95] Robert J. Chassell and Richard M. Stallman. Texinfo: The GNU Documentation Format (for Texinfo version 2.20, 28 February 1995). Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1995. ISBN 1-882114-63-9. LCCN ???? US$25.00.

Chassell:1996:TGD

[CS96] Robert J. Chassell and Richard M. Stallman. Texinfo: The GNU Documentation Format (for Texinfo version 2.23, 1 October 1996). Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, second edition, 1996. ISBN 1-882114-64-7. LCCN ???? US$25.00.

Chassell:1999:TGD

[CS99] Robert J. Chassell and Richard M. Stallman. Texinfo: The GNU Documentation Format (for Texinfo version 4.0, 28 September 1999). Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1999. ISBN 1-882114-67-1. x + 244 pp. LCCN ???? US$25.00.

Crawford:2005:FDP

[CSD^{+}05] Diane Crawford, Richard Stallman, Peter Denning, Jon Crowcroft, Herbert Kanner, Edwin D. Reilly, Jr., Len Cohen, Larry Brunelle, Philip Burgess, and Jonathan Grudin. Forum: To defend privacy, don’t collect personal data; look beyond abstraction to define computing; voter anonymity and vote security still impossible; hold the LEO and celebrate tools; must reading for software engineers; when is it OK to republish? Communications of the ACM, 48(5):11–13, May 2005.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Csendes:1999:DRC

[Cse99] Tibor Csendes, editor. Developments in Reliable Computing: Papers presented at the International Symposium on Scientific Computing, Computer Arithmetic, and Validated Numerics, SCAN-98, in Szeged, Hungary, volume 5(3) of Reliable Computing = Nadezhnye vychisleniia. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-6057-5. LCCN QA76.9.E94 D48 1999.

Cooper:2014:OSS

[CSEP14] Matthew L. Cooper, Clifford A. Shaffer, Stephen H. Edwards, and Sean P. Ponce. Open source software and the algorithm visualization community. Science of Computer Programming, 88(??):82–91, August 1, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167642314000021.

Crawford:2003:FLM

[CSP^{+}03] Diane Crawford, Richard M. Stallman, Richard Proudfoot, Sudip Bhattacharjee, Ram D. Gopal, G. Lawrence Sanders, Tom Pittman, Richard R. Brooks, and Dennis E. Hamilton. Forum: Legalize music sharing now; hold the games; give up trying to halt the flow of information; praise for Denning’s powerful chord of professionalism. Communications of the ACM, 46(9):13–15, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Costa-Soria:2009:TSA

[CSP09] Cristóbal Costa-Soria and Jennifer Pérez. Teaching software architectures and aspect-oriented software development using open-source projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):385, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Crawford:2007:POS

[CSV^{+}07] T. Daniel Crawford, C. David Sherrill, Edward F. Valeev, Justin T. Fermann, Rollin A. King, Matthew L. Leininger, Shawn T. Brown, Curtis L. Janssen, Edward T. Seidl, Joseph P. Kenny, and Wesley D. Allen. PSI3: an open-source Ab Initio electronic structure package. Journal of Computational Chemistry, 28(9):1610–1616, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2004:OSC

[CSY^{+}04] Kai Chen, Stephen R. Schach, Liguo Yu, Jeff Offutt, and Gillian Z. Heller. Open-source change logs. Empirical Software Engineering, 9(3):197–210, September 2004. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEMSE.0000027779.70556.d0.

Cheimaras:2022:LCO

[CTP^{+}22] Vasileios Cheimaras, Athanasios Trigkas, Panagiotis Papageorgas, Dimitrios Piromalis, and Emmanouil Sofianopoulos. A low-cost open-source architecture for a digital signage emergency evacuation system for cruise ships, based on IoT and LTE/4G technologies. Future Internet, 14(12):366, December 07, 2022. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/14/12/366.

Curley:1999:OSS

[Cur99] Charles Curley. Open source software for real-time solutions. Linux Journal, 66:??, October 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.com/lj-issues/issue66/3576.
html.

Curley:2002:EFS

[Cur02] Charles Curley. Emacs: the free software IDE. Linux Journal, 98:??, June 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://www.linuxjournal.

com/article.php?sid=5765.

Cusumano:2004:TSMc

[Cus04] Michael A. Cusumano. Technology strategy and management: Reflections on free and open software. Communications of the ACM, 47(10):25–27, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Castelluccia:2013:TEB

[CV13] Daniela Castelluccia and Giuseppe Visaggio. Teaching evidence-based software engineering: learning by a collaborative mapping study of open source software. ACM SIGSOFT Software Engineering Notes, 38(6):1–4, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cejnek:2022:PPO

[CV22] Matous Cejnek and Jan Vrba. Padasip: an open-source Python toolbox for adaptive filtering. Journal of Computational Science, 65:??, November 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S1877750322002460.

Courtes:2015:RUCa

[CW15a] Ludovic Courtès and Ricardo Wurmus. Reproducible and user-controlled software environments in HPC with Guix. arxiv.org, pages 1–13, 2015. URL https://arxiv.org/abs/

1506.02822.

Courtes:2015:RUCb

[CW15b] Ludovic Courtès and Ricardo Wurmus. Reproducible and user-controlled software environments in HPC with Guix. Lecture Notes in Computer Science, 9523:579–591, 2015. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crawford:2004:FDD

[CWB^{+}04] Diane Crawford, Steve Wildstrom, Hal Berghel, Charles Burnette, Jiming Liu, Peter J. Denning, Carl Binding, and Robert Swarr. Forum: Defense in depth gets the worm; put cognitive modes in CS and its curricula; don’t blame WAP [Wireless Application Protocol]; market share vs. peer approval in open source. Communications of the ACM, 47(2):11–13, February 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crowston:2012:FLO

[CWHW12] Kevin Crowston, Kangning Wei, James Howison, and Andrea Wiggins. Free/ libre open-source software development: What we know and what we do not know. ACM Computing Surveys, 44(2):7:1–7:35, February 2012. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Chen:2020:HCT

[CWM^{+}20] Lin Chen, Di Wu, Wanwangying Ma, Yuming Zhou, Baowen Xu, and Hareton Leung. How C++ templates are used for generic programming: an empirical study on 50 open source systems. ACM Transactions on Software Engineering and Methodology, 29(1):3:1–3:49, February 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3356579.

Cruz:2006:ECF

[CWZ06] David Cruz, Thomas Wieland, and Alexander Ziegler. Evaluation criteria for free/open source software products based on project analysis. Software Process: Improvement and Practice, 11(2):107–122, March 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Chen:2023:RPO

[CX23] Zhi Chen and Peng Xiong. RSOME in Python: An open-source package for robust stochastic optimization made easy. INFORMS Journal on Computing, 35(4):717–724, July/ August 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/
doi/full/10.1287/ijoc.2023.1291.

Cheng:2023:VLS

[CYL^{+}23] Yiran Cheng, Shouguo Yang, Zhe Lang, Zhiqiang Shi, and Limin Sun. VERI: a large-scale open-source components vulnerability detection in IoT firmware. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004606.

Chikano:2019:IOS

[CYOS19] Naoya Chikano, Kazuyoshi Yoshimi, Junya Otsuki, and Hiroshi Shinaoka. irbasis: Open-source database and software for intermediate-representation basis functions of imaginary-time Green’s function. Computer Physics Communications, 240(??):181–188, July 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046551930058X.

Colnet:1999:OEP

[CZ99] D. Colnet and O. Zendra. Optimizations of Eiffel programs: SmallEiffel, the GNU Eiffel Compiler. In Proceedings of Technology of Object-Oriented Languages and Systems, 7–10 June 1999, pages 341–350. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. CODEN ???? ISSN ????

Chen:2022:OSP

[CZ22] Donghua Chen and Runtong Zhang. An open source project for tuning and analyzing MapReduce performance in Hadoop and Spark. IEEE Software, 39(1):61–69, February 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carlson:2021:OSP

[CZS^{+}21] Max Carlson, Xiaoning Zheng, Hari Sundar, George Em Karniadakis, and Robert M. Kirby. An open-source parallel code for computing the spectral fractional Laplacian on 3D complex geometry domains. Computer Physics Communications, 261(??):Article 107695, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303416.

deAssisRangel2015

[dA15] João José de Assis Rangel and Anna Christine Azevedo Cordeiro. Free and open-source software for sustainable analysis in logistics systems design. Journal of Simulation, 9(1):27– 42, February 2015. ISSN 1747-7778 (print), 1747-7786 (electronic).

Daly:2002:AOS

[Dal02] Tim Daly. Axiom as open source. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):28, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Daniluk:2011:CCC

[Dan11] Andrzej Daniluk. Cooperative and competitive concurrency in scientific computing. A full open-source upgrade of the program for dynamical calculations of RHEED intensity oscillations. Computer Physics Communications, 182(6):1389– 1390, June 2011. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465511000804.

Pietri:2023:DVW

[DARJ23] Marco De Pietri, Javier Alguacil, Eduardo Rodríguez, and Rafael Juárez. Development and validation in water of FLUNED, an open-source tool for fluid activation calculations. Computer Physics Communications, 291(??):Article 108807, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001522.

DataGeneral:1985:DGE

[Dat85] Data General Corporation, Westboro, MA, USA. Data General EMACS text editor user’s manual, revision 00, February 1985 edition, 1985. various pp.

Davidson:1991:GCC

[Dav91] Andrew Davidson. Generic containers in C++. Dr. Dobb’s Journal of Software Tools, 16(8):50, 52, 124–125, August 1991. CODEN DDJSDM. ISSN 0884-5395.

Davis:2002:OSD

[DB02] Chad Davis and Coskun Bayrak. Open source development and the World Wide Web: a certain tension. ACM SIG-SOFT Software Engineering Notes, 27(5):93–97, September 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Devine:2005:BDF

[DB05] Dawn Devine and Michael Baxter. Belly dance and free software. Linux Journal, 2005(133):2, May 2005. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Damiani:2010:SIO

[DBBA10] Ernesto Damiani, Luis Barbosa, Peter T. Breuer, and Claudio A. Ardagna. Special issue: Open source certification. International Journal of Computer Systems Science and Engineering, 25(4):??, July 2010. CODEN CSSEEI. ISSN 0267-6192.

Ducrozet:2016:HOO

[DBLF16] Guillaume Ducrozet, Félicien Bonnefoy, David Le Touzé, and Pierre Ferrant. HOS-ocean: Open-source solver for nonlinear waves in open ocean based on High-Order Spectral method. Computer Physics Communications, 203(??):245–254, June 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465516300327.

delBianco:2011:SOS

[dBLMT11] Vieri del Bianco, Luigi Lavazza, Sandro Morasca, and Davide Taibi. A survey on open source software trustworthiness. IEEE Software, 28(5):67–75, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Derouillat:2018:SCO

[DBP^{+}18] J. Derouillat, A. Beck, F. Pérez, T. Vinci, M. Chiaramello, A. Grassi, M. Flé, G. Bouchard, I. Plotnikov, N. Aunai, J. Dargent, C. Riconda, and M. Grech. Smilei: a collaborative, open-source, multi-purpose particle-in-cell code for plasma simulation. Computer Physics Communications, 222(??):351–373, January 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465517303314.

Diaz:2000:GPB

[DC00] Daniel Diaz and Philippe Codognet. GNU Prolog: Beyond compiling Prolog to C. Lecture Notes in Computer Science, 1753:81–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/1753/17530081.htm;
 http://link.springer-ny.com/link/service/series/0558/
papers/1753/17530081.pdf.

Dal:2023:IID

[DC23] Deniz Dal and Esra Celik. Investigation of the impact of different versions of GCC on various metaheuristicbased solvers for traveling salesman problem. The Journal of Supercomputing, 79(11):12394–12440, July 2023. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-023-05152-z.

deCarnavalet:2014:CIV

[dCdCM14] Xavier de Carné de Carnavalet and Mohammad Mannan. Challenges and implications of verifiable builds for securitycritical open-source software. In Proceedings of the 30th Annual Computer Security Applications Conference, ACSAC ’14, pages 16–25. ACM Press, New York, NY 10036, USA, 2014. ISBN 1-4503-3005-3.

DiBona:2005:OSC

[DCS05] Chris DiBona, Danese Cooper, and Mark Stone, editors. Open Sources 2.0: the continuing evolution. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00802-3. xl + 445 pp. LCCN

QA76.76.S46 O643 2005. URL http://www.oreilly.com/
catalog/9780596008024.

Dionisio:2008:ICS

[DD08] John David N. Dionisio and Kam D. Dahlquist. Improving the computer science in bioinformatics through open source pedagogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):115–119, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL https://www.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2008.bib.

Draves:2010:FAO

[DD10] Scott Draves and Isabel Walcott Draves. The flame algorithm and its open source culture. Computer Graphics, 44(3):3:1– 3:??, August 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Dhir:2017:AOS

[DD17] Swati Dhir and Sanjay Dhir. Adoption of open-source software versus proprietary software: An exploratory study. Strategic Change, 26(4):363–371, 2017. CODEN STCHFT. ISSN 1086-1718 (print), 1099-1697 (electronic).

Dionisio:2007:OSS

[DDA^{+}07] John David N. Dionisio, Caskey L. Dickson, Stephanie E. August, Philip M. Dorin, and Ray Toal. An open source software culture in the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):70–74, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL https://www.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Dong:2003:FLG

[DDHS03] Yifei Dong, Xiaoqun Du, Gerard J. Holzmann, and Scott A. Smolka. Fighting livelock in the GNU i-protocol: a case study in explicit-state model checking. International Journal on Software Tools for Technology Transfer (STTT), 4(4): 505–528, August 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

DDJStaff:1998:NVK

[DDJ98a] DDJ Staff. News and views: Kudos for free software pioneers; PSCs: Personal supercomputers; smart dialing; let it snow…; math for the Web; the taxman changes; advances in nanoelectromechanical technology; Tcl goes it alone. Dr. Dobb’s Journal of Software Tools, 23(5):18, May 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Staff:1998:NVK

[DDJ98b] DDJ Staff. News and views: Kudos for free software pioneers; PSCs: Personal supercomputers; smart dialing; let it snow…; math for the Web; the taxman changes; advances in nanoelectromechanical technology; Tcl goes it alone. Dr. Dobb’s Journal of Software Tools, 23(5):18, May 1998. CO-DEN DDJOEB. ISSN 1044-789X.

DDJStaff:1999:NVSa

[DDJ99] DDJ Staff. News and views: Speeding up 3D modeling; Project Gutenberg; FSF honors Larry Wall; smart pens don’t make smart writers; power hogs; virtual fish: Java’s killer app?; evaluating testing tools; software patents con. Dr. Dobb’s Journal of Software Tools, 24(1):18, January 1999. CODEN DDJOEB. ISSN 1044-789X.

De:2015:OSS

[De’15] Rahul De’. Open Source Software in the Global South, page ?? Wiley, New York, NY, USA, 2015. ISBN 1-118-76777-2.

Deitcher:2010:COS

[Dei10] Avi Deitcher. The challenges of open source in the enterprise. Linux Journal, 2010(195):3:1–3:??, July 2010. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Delozier:2001:FOS

[Del01] Eric P. Delozier. Free and Open Source Software, page ?? Wiley, New York, NY, USA, 2001. ISBN 0-471-34608-X.

Deno:1999:NRB

[Den99] Cynthia Deno. New releases of *BSD and Debian Linux OSes given away at USENIX Annual Conference. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CO-DEN LOGNEM. ISSN 1044-6397. URL http://www.usenix.

org/publications/login/1999-6/newreleases.html.

Denny:2013:SOS

[Den13] J. K. Denny. SAGE: Open source mathematics software system (http://sagemath.org). College Mathematics Journal, 44(2):149–155, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/college.math.j.44.
2.149.

Deodhar:1990:GAI

[Deo90] S. Deodhar. GNU-Aid: Intelligent computer aided instruction system. In ACE ’90. Proceedings of [XVI Annual Convention and Exhibition of the IEEE In India], pages 14–16. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN ???? ISSN ????

Dewar:2007:BFW

[Dew07] Robert Dewar. Birds-of-a-feather: where would you like to see GNAT go? ACM SIGADA Ada Letters, 27(3):97–98, December 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dutt:2000:GBG

[DF00] Christoph Dutt and Joachim Freiburg. GiMP: Bilder gestalten, Fotos retuschieren; [Grundlagen der professionellen Bildbearbeitung, der Umgang mit Fotos, Grafiken und Text, Bilder für das Internet richtig vorbereiten; auf der CD: GIMP für Windows, SCO Unix, Debian GNU Linux, Solaris, OS/2 und BSD, Quelltext aller GIMP- und GTK-Versionen, über 300 Plug-ins in C, Perl, tcl, Python und Scheme, XFree86/23.3.6, GIMP User Manual als PDF-Dateien]. C und L, Böblingen, Germany, 2000. ISBN 3-932311-64-7. 522 + 98 pp. LCCN ???? Includes CD-ROM.

Delamo:2015:DOS

[DFCPSF15] Manuel Delamo, Santiago Felici-Castell, Juan J. Pérez-Solano, and Andrew Foster. Designing an open source maintenance-free Environmental Monitoring Application for Wireless Sensor Networks. The Journal of Systems and Software, 103(??):238–247, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121215000400.

Donnellan:2005:IOS

[DFLS05] Brian Donnellan, Brian Fitzgerald, Brian Lake, and John Sturdy. Implementing an Open Source knowledge base. IEEE Software, 22(6):92–95, November/December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Detti:2023:PBO

[DFP23] Andrea Detti, Ludovico Funari, and Luca Petrucci. Bench: An open-source factory of benchmark microservice applications. IEEE Transactions on Parallel and Distributed Systems, 34 (3):968–980, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Drost-Fromm:2021:OSC

[DFT21] I. Drost-Fromm and R. Tompkins. Open source community governance the Apache way. Computer, 54(4):70–75, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DiRenzo:2020:HSO

[DFU20] Mario Di Renzo, Lin Fu, and Javier Urzay. HTR solver: an open-source exascale-oriented task-based multi-GPU highorder code for hypersonic aerothermodynamics. Computer Physics Communications, 255(??):Article 107262, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300837.

Deaton:1993:ACS

[DGBH93] E. Deaton, K. M. George, H. Bergel, and G. Hedrick, editors. Applied Computing: States of the Art and Practice — 1993 Proceedings of the 1993 ACM/SIGAPP Symposium on Applied Computing. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-567-4. LCCN

QA76.76.A65 S95 1993. URL http://www.acm.org/pubs/
contents/proceedings/sac/162754/.

Duenas:2007:AEC

[DGC^{+}07] Juan C. Dueñas, Hugo A. Parada G., Félix Cuadrado, Manuel Santillán, and José L. Ruiz. Apache and Eclipse: Comparing open source project incubators. IEEE Software, 24(6):90–98, November/December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dasgupta:2019:FSO

[DGJH19] Adhiraj Dasgupta, Esteban Gonzalez-Juez, and Daniel C. Haworth. Flame simulations with an open-source code. Computer Physics Communications, 237(??):219–229, April 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551830403X.

Duncan:2001:OOS

[DH01] Andy Duncan and Sean Hull. Oracle and Open Source. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, April 2001. ISBN 0-596-00018-9. xvi + 406 pp. LCCN QA76.9.D3 D845 2001. US$39.95. URL http:

//www.oreilly.com/catalog/9780596000189;   http://
www.oreilly.com/catalog/oracleopen.

DiBona:2004:EWO

[DiB04] Chris DiBona. EOF: Why Open Source wins. Linux Journal, 2004(119):17, March 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

DEC:1974:TUG

[Dig74] Digital Equipment Corporation, Maynard, MA, USA. TECO user’s guide, 1974. v + 61 + 30 pp.

DEC:1975:DSI

[Dig75a] Digital Equipment Corporation, Maynard, MA, USA. DEC system 10: introduction to TECO (Text Editor and Corrector), 1975. 34 pp.

DEC:1975:TEC

[Dig75b] Digital Equipment Corporation, Maynard, MA, USA. Text Editor and Corrector program: programmer’s reference manual, revised edition, 1975. various pp.

DEC:1980:PTU

[Dig80a] Digital Equipment Corporation, Maynard, MA, USA. PDP-11 TECO user’s guide, 1980. iii + 150 pp.

DEC:1980:VVP

[Dig80b] Digital Equipment Corporation, Maynard, MA, USA. [VAX/ VMS Programmming card and PDP-11 TECO user’s guide], 1980. ???? pp.

DEC:1982:DMS

[Dig82] Digital Equipment Corporation, Maynard, MA, USA. Data management systems: SIG handout: 1982 Spring DECUS Symposium: May 10–14, 1982, Atlanta, Georgia: TECO sessions, 1982. 55 pp.

DeChant:2023:VVO

[DIK^{+}23] Corey DeChant, Casey Icenhour, Shane Keniley, Grayson Gall, Alexander Lindsay, Davide Curreli, and Steven Shannon. Verification and validation of the open-source plasma fluid code: Zapdos. Computer Physics Communications, 291(??):Article 108837, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523001820.

deIcaza:1998:GDE

[dILM98] Miguel de Icaza, Elliot Lee, and Federica Mena. The GNOME desktop environment. In USENIX [USE98a], page ??

ISBN ???? LCCN ???? URL http://www.usenix.org/
publications/library/proceedings/usenix98/freenix/
icaza.ps.

Decker:2022:PEO

[DKK22] Jonathan Decker, Piotr Kasprzak, and Julian Martin Kunkel. Performance evaluation of open-source serverless platforms for Kubernetes. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/7/234.

Detappe:2014:SOS

[DKMB14] A. Detappe, H. Korideck, G. Makrigiorgos, and R. Berbeco. SU-E-T-253: Open-source automatic software for quantifying biological assays of radiation effects. Medical Physics, 41 (6Part15):281–282, 2014. CODEN MPHYA6. ISSN 2473-4209.

Dinges:2011:OSA

[DKMT11] Gerlinde Dinges, Alexander Kowarik, Bernhard Meindl, and Matthias Templ. An open source approach for modern teaching methods: The interactive TGUI system. Journal of Statistical Software, 39(7):??, March 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v39/i07.

delaChevallerie:2015:FLH

[dlCKK15] David de la Chevallerie, Jens Korinth, and Andreas Koch. ffLink: a lightweight high-performance open-source PCI Express Gen3 interface for reconfigurable accelerators. ACM SIGARCH Computer Architecture News, 43(4):34–39, September 2015. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

delaPuente:1999:RTP

[dlPRGB99] Juan A. de la Puente, José F. Ruiz, and Jesús M. González-Barahona. Real-time programming with GNAT: Specialized kernels versus POSIX threads. ACM SIGADA Ada Letters, 19(2):73–77, June 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dong:2023:BWL

[DLT^{+}23] Yiwen Dong, Zheyang Li, Yongqiang Tian, Chengnian Sun, Michael W. Godfrey, and Meiyappan Nagappan. Bash in the wild: Language usage, code smells, and bugs. ACM Transactions on Software Engineering and Methodology, 32(1):8:1– 8:22, January 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https://dl.acm.org/

doi/10.1145/3517193.

delaVara:2021:ACC

[dlVRB21] Jose Luis de la Vara, Alejandra Ruiz, and Gaël Blondelle. Assurance and certification of cyber-physical systems: the AMASS open source ecosystem. The Journal of Systems and Software, 171(??):??, January 2021. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121220302120.

Do:1997:LEU

[DM97] James Do and Muhammed Mudawwar. Letters to the editor: Unicode misunderstood. Computer, 30(6):6, 9, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response and rebuttal to [Mud97].

Mauro:1999:IML

[dM99] Pancrazio de Mauro. Internationalizing messages in Linux programs. Linux Journal, 59:??, March 1999. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

ftp://ftp.ssc.com/pub/lj/listings/issue59/3023.tgz.

Dowling:2015:UFOa

[DM15a] Phelim Dowling and Kevin McGrath. Using free and open source tools to manage software quality. ACM Queue: Tomorrow’s Computing Today, 13(4):20, April 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/detail.cfm?id=2767182.

Dowling:2015:UFOb

[DM15b] Phelim Dowling and Kevin McGrath. Using free and open source tools to manage software quality. Communications of the ACM, 58(7):51–55, July 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://cacm.

acm.org/magazines/2015/7/188743/fulltext.

Dongarra:1979:LUG

[DMBS79] J. J. Dongarra, C. B. Moler, J. R. Bunch, and G. W. Stewart. LINPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1979. ISBN 0-89871-172-X (paperback). 320 pp. LCCN QA76.73 .L22 L5 1979, QA214 .L56 1979.

Darken:2005:DOS

[DMJ05] Rudy Darken, Perry McDowell, and Erik Johnson. The Delta3D open source game engine. IEEE Computer Graphics and Applications, 25(3):10–12, May/June 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Duparc:2002:WOS

[DMP^{+}02] Daniel Duparc, Bernard Mourrain, Bernard Parisse, Fabrice Rouillier, Marie-Françoise Roy, Nicolas Thiéry, and Paul Zimmermann. Workshop on Open Source Computer Algebra, Tuesday 21st – Thursday 23rd of May 2002, Lyon, France. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):24–27, March 2002. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Duy:2016:HDF

[DO16] Truong Vinh Truong Duy and Taisuke Ozaki. Hybrid and 4-D FFT implementations of an open-source parallel FFT package OpenFFT. The Journal of Supercomputing, 72(2):391–416, February 2016. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL http://link.springer.com/
article/10.1007/s11227-015-1568-8.

Doherty:2001:LOS

[Doh01] Sean Doherty. The law and open-source software. World-Wide Web document., October 29, 2001. URL http://www.

networkcomputing.com/1222/1222ws1.html.

Dollin:1991:HT

[Dol91] D. Dollin. The HP-ST toolset. In Prehn and Toetenel [PT91], pages 687–688. ISBN 0-387-54834-3 (New York) (vol. 1), 3-540-54834-3 (Berlin) (vol. 1), 0-540-54868-8 (New York) (vol. 2), 3-540-54868-8 (Berlin) (vol. 2). LCCN QA76.76.D47V36 1991.

Donato:2004:SPS

[Don04] Alberto Donato. A software platform to support dynamically reconfigurable systems-on-chip under the GNU/Linux operating system. Politecnico, Milano, Italy, 2004. ISBN ???? xvi + 95 pp. LCCN ????

DiBona:1999:OSV

[DOS99] Chris DiBona, Sam Ockman, and Mark Stone. Open Sources: Voices from the Open Source Revolution. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-582-3. viii + 272 pp. LCCN QA76.754

.O63 1999. US$24.95. URL http://www.oreilly.com/

catalog/opensources/. Includes a chapter The GNU Operating System and the Free Software Movement, by Richard Stallman, and a chapter Future of Cygnus Solutions: An Entrepreneur’s Account, by Michael Tiemann.

Demko:2009:SOS

[DP09] Aleksander B. Demko and Nick J. Pizzi. Scopira: an open source C++ framework for biomedical data analysis applications. Software—Practice and Experience, 39(6):641–660, April 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dagkakis:2016:MOS

[DPH16] Georgios Dagkakis, Ioannis Papagiannopoulos, and Cathal Heavey. ManPy: an open-source software tool for building discrete event simulation models of manufacturing systems. Software—Practice and Experience, 46(7):955–981, July 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Denning:1991:USV

[DPL^{+}91] Dorothy E. Denning, Donn B. Parker, Steven Levy, Eugene Spafford, Paula Hawthorn, Marc Rotenberg, J. J. Buck BloomBecker, and Richard Stallman. The United States vs. Craig Neidorf: a debate on electronic publishing, Constitutional rights and hacking. Communications of the ACM, 34 (3):22–43, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0001-0782/102869.html.

Drexl:1994:WPC

[Dre94] Josef Drexl. What is protected in a computer program?: copyright protection in the United States and Europe, volume 15 of IIC studies. VCH, Weinheim, Federal Republic of Germany, 1994. ISBN 3-527-28688-8. xvi + 124 pp. LCCN K1443.C6

D73 1994. DM98.00. URL http://www.loc.gov/catdir/
enhancements/fy0706/95158908-d.html;   http://www.
loc.gov/catdir/toc/wiley022/95158908.html.

Datta:2021:URB

[DRM21] Subhajit Datta, Reshma Roychoudhuri, and Subhashis Majumder. Understanding the relation between repeat developer interactions and bug resolution times in large open source ecosystems: a multisystem study. Journal of Software: Evolution and Process, 33(4):e2317:1–e2317:??, April 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Danesh:2001:SLG

[DRP01] Arman Danesh, James Russell, and Richard Petersen. SAIR Linux & GNU certified administrator: exam guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213205-1. xxv + 1008 pp. LCCN ????

Donnelly:1988:BYC

[DS88] Charles Donnelly and Richard M. Stallman. BISON— the YACC-compatible parser generator. Technical report, Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296,

1988. URL ftp://ftp.gnu.org/pub/gnu/bison/. Bison was

largely written by Robert Corbett, and made yacc-compati-ble by Richard Stallman. See also [Pax88].

Donnelly:1990:BYP

[DS90] Charles Donnelly and Richard Stallman. Bison: the Yacccompatible parser generator. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, Bison Version 1.12 edition, December 1990. iv + 96 pp.

Donnelly:1999:BMU

[DS99] Charles Donnelly and Richard Stallman. Bison: the YACCcompatible Parser Generator, Bison Version 1.28. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, January 14, 1999. ISBN 1-882114-45-0. ???? pp. LCCN ???? US$15.

Donnelly:2000:BMU

[DS00] Charles Donnelly and Richard Stallman. Bison Manual: Using the YACC-compatible Parser Generator, for Version 1.29. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2000. ISBN 1-882114-44-2. ???? pp. LCCN ????

Donnelly:2002:BMU

[DS02] Charles Donnelly and Richard M. Stallman. Bison Manual: Using the YACC-compatible Parser Generator. GNU Press, Boston, MA, USA, 2002. ISBN 1-882114-34-5. 130 (est.) pp.

LCCN ???? US$20.00. URL http://www.gnupress.org/
book10.html.

Daxhelet:2016:EER

[DSB^{+}16] M. Daxhelet, M. Suppa, F. Benhadou, V. Djamei, T. Tzellos, G. Ingvarsson, J. Boer, A. Martorell, J. R. Ingram, N. Desai, A. Nassif, J. Revuz, C. Hotz, V. Bettoli, I. E. Deckers, G. B. Jemec, E. Prens, C. C. Zouboulis, and V. del Marmol. Establishment of a European Registry for hidradenitis suppurativa/acne inversa by using an open source software. Journal of the European Academy of Dermatology and Venereology, 30 (8):1424–1426, 2016. ISSN 0926-9959 (print), 1468-3083 (electronic).

Damian:2019:OST

[DSK19] Santiago Márquez Damián, Federico Schaumburg, and Pablo A. Kler. Open-source toolbox for electromigrative separations. Computer Physics Communications, 237(??):244– 252, April 2019. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465518304077.

DiRemigio:2019:POS

[DSM^{+}19] Roberto Di Remigio, Arnfinn Hykkerud Steindal, Krzysztof Mozgawa, Ville Weijo, Hui Cao, and Luca Frediani. PCM-Solver: an open-source library for solvation modeling. International Journal of Quantum Chemistry, 119(1):e25685:1– e25685:??, January 5, 2019. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dinh-Trong:2005:FPR

[DTB05] T. T. Dinh-Trong and J. M. Bieman. The FreeBSD project: a replication case study of open source development. IEEE Transactions on Software Engineering, 31(6):481–494, June 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520

(electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1463231.

DuBoff:2002:LWN

[DuB02] Alan DuBoff. Letters: What? no more MontaVista Journeyman? what about the GNU cross-compiler? Embedded Linux Journal, 9:4, May/June 2002. CODEN ???? ISSN

1534-083X. URL http://embedded.linuxjournal.com/
magazine/issue09/;   http://www.linuxdevices.com/articles/
AT4384699491.html.

Duerinckx:1997:CRC

[Due97] G. Duerinckx. Cyclic redundancy checks in Ada95. ACM SIGADA Ada Letters, 17(1):41–53, January/February 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dumbill:2005:DGL

[Dum05] Edd Dumbill. Debain GNU/Linux: An Explorer‘s Notebook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00883-X. 256 (est.) pp. LCCN ???? EUR 24.00.

DeSutter:2007:LTC

[DVC^{+}07] Bjorn De Sutter, Ludo Van Put, Dominique Chanet, Bruno De Bus, and Koen De Bosschere. Link-time compaction and optimization of ARM executables. ACM Transactions on Embedded Computing Systems, 6(1):5:1–5:??, February 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dvorak:2003:NLO

[Dvo03] Zdeněk Dvořák. A new loop optimizer for GCC. In Hutton et al. [HDR03], pages 43–55. ISBN ???? LCCN ???? URL

http://www.linux.org.uk/~ajh/gcc/gccsummit-2003-proceedings.
pdf.

Dvorak:2004:DWI

[Dvo04] Zdeněk Dvořák. Declarative world inspiration. In Hutton et al. [HDR04], pages 25–36. ISBN ???? LCCN

???? URL http://people.redhat.com/lockhart/.gcc04/
MasterGCC-2side.pdf.

Dwan:2004:OSV

[Dwa04] Berni Dwan. Open source vs closed. Network Security, 2004(5):11–13, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1353485804000807.

Dempsey:2002:WOS

[DWJG02] Bert J. Dempsey, Debra Weiss, Paul Jones, and Jane Greenberg. Who is an open source software developer? Communications of the ACM, 45(2):67–72, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Desplanques:2014:TOS

[DWP^{+}14] M. Desplanques, K. Wang, J. Phillips, G. Gueorguiev, G. Baroni, and G. Sharp. TH-A-19A-01: An open source software for proton treatment planning in heterogeneous medium. Medical Physics, 41(6Part31):533, 2014. CODEN MPHYA6. ISSN 2473-4209.

Davidson:2018:COS

[DXT^{+}18] Scott Davidson, Shaolin Xie, Christopher Torng, Khalid Al-Hawai, Austin Rovinski, Tutu Ajayi, Luis Vega, Chun Zhao, Ritchie Zhao, Steve Dai, Aporva Amarnath, Bandhav Veluri, Paul Gao, Anuj Rao, Gai Liu, Rajesh K. Gupta, Zhiru Zhang, Ronald Dreslinski, Christopher Batten, and Michael Bedford Taylor. The Celerity open-source 511-core RISC-V tiered accelerator fabric: Fast architectures and design methodologies for fast chips. IEEE Micro, 38(2):30–41, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.computer.org/csdl/mags/mi/
2018/02/mmi2018020030-abs.html.

Dyer:2003:GDE

[Dye03] Russell Dyer. The GNOME 2 desktop environment. Linux Journal, 108:??, April 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Eaton:2014:GOR

[EaoGOBHW14] John W. (John Wesley) Eaton, David (Co author of GNU Octave) Bateman, Søren Hauberg, and Rik Wehbring. The GNU Octave 3.8 reference manual. Samurai Media Limited, London, USA, third edition, 2014. ISBN 988-13277-3-3 (paperback: vol. 1), 988-13277-4-1 (paperback: vol. 2). xii + 890 (two volumes) pp. LCCN QA 76.95 .E1 G68 2014.

Eaton:1997:GOH

[Eat97] John W. Eaton. GNU Octave: a high-level interactive language for numerical computations. Network Theory Ltd., Bristol, UK, 1997. ISBN 0-9541617-2-6. viii + 311 pp. LCCN ????

Eaton:2000:GOH

[Eat00] John W. (John Wesley) Eaton. GNU Octave: a high-level interactive language for numerical computations: edition 3 for Octave version 2.0.13. Network Theory Ltd., Bristol, UK, 2000. ISBN 0-9541617-2-6 (paperback). viii + 311 pp. LCCN ????

Eaton:2002:GOH

[Eat02] John W.. Eaton. GNU Octave: a high-level interactive language for numerical computations. Network Theory Ltd., Bristol, UK, third edition, 2002. ISBN 0-9541617-2-6. viii + 311 pp. LCCN ???? For Octave version 2.0.13, February 1997.

Eaton:2005:GOH

[Eat05] John W. Eaton. GNU Octave: a high-level interactive language for numerical computations. Network Theory Ltd., Bristol, UK, 2005. ISBN 0-9541617-2-6. viii + 311 pp.

LCCN ???? URL http://www.gbv.de/dms/ilmenau/toc/

520168526eaton.PDF. For Octave version 2.0.17 (stable).

Ebert:2007:OSD

[Ebe07] Christof Ebert. Open source drives innovation. IEEE Software, 24(3):105–109, May/June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2008:OSS

[Ebe08] Christof Ebert. Open source software in industry. IEEE Software, 25(3):52–53, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://csdl.

computer.org/comp/mags/so/2008/03/mso2008030052.pdf.

Ebert:2009:OSS

[Ebe09] Christof Ebert. Open source software: Guest Editor’s introduction: How open source tools can benefit industry. IEEE Software, 26(2):50–51, March/April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Eaton:2008:GOM

[EBH08] John W. Eaton, David Bateman, and Søren Hauberg. GNU Octave manual: a high-level interactive language for numerical computations. Network Theory Ltd., Bristol, UK, 2008. ISBN 0-9546120-6-X. x + 555 pp. LCCN ???? US$39.95. Version 3 for Octave version 3.0.2, August 2008.

Eddelbuttel:1996:SRO

[Edd96] Dirk Eddelbüttel. Software review: Object-oriented econometrics: Matrix programming in C++ using GCC and newmat. Journal of Applied Econometrics, 11(2):199–209, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eddelbuttel:2000:EO

[Edd00] Dirk Eddelbüttel. Econometrics with Octave. Journal of Applied Econometrics, 15(5):531–542, September–October 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Edelsohn:2004:HLL

[Ede04] David Edelsohn. High-level loop optimizations for GCC. In Hutton et al. [HDR04], pages 37–54. ISBN ???? LCCN

???? URL http://people.redhat.com/lockhart/.gcc04/
MasterGCC-2side.pdf.

Edson:2016:SOS

[Eds16] Curtis B. Edson. Out in space to open source: challenges and rewards of studying ecology using free software for space-borne remote sensing and geographic information system. Ecology, 97(10):2894–2896, 2016. CODEN ECOLAR. ISSN 0012-9658 (print), 1939-9170 (electronic).

Edwards:1998:ITC

[Edw98] John Edwards. Industry trends: The changing face of freeware. Computer, 31(10):11–13, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/co/books/co1998/pdf/
rx011.pdf.

El-Emam:2001:EOS

[EE01] Khaled El-Emam. Ethics and open source. Empirical Software Engineering, 6(4):291–292, December 2001. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1011962213685.

Edelsohn:2005:CGC

[EGH^{+}05] D. Edelsohn, W. Gellerich, M. Hagog, D. Naishlos, M. Namolaru, E. Pasch, H. Penner, U. Weigand, and A. Zaks. Contributions to the GNU Compiler Collection. IBM Systems Journal, 44(2):259–278, ???? 2005. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/
sj/442/edelsohn.pdf.

Ellson:2002:GOS

[EGK^{+}02] John Ellson, Emden Gansner, Lefteris Koutsofios, Stephen C. North, and Gordon Woodhull. Graphviz — open source graph drawing tools. Lecture Notes in Computer Science, 2265:483– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2265/22650483.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2265/22650483.pdf.

Egyedi:2001:SFC

[Egy01] T. Egyedi. Strategies for de facto compatibility: Standardization, proprietary and open source approaches to Java. Knowledge, Technology, and Policy, 14(2):113–128, July 1, 2001. CODEN ???? ISSN 0897-1986.

Eisenmenger:2006:EVS

[EHHH06] Frank Eisenmenger, Ulrich H. E. Hansmann, Shura Hayryan, and Chin-Kun Hu. An enhanced version of SMMP-open-source software package for simulation of proteins. Computer Physics Communications, 174(5):422–429, March 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505005850.

Echtle:1994:DCE

[EHP94] Klaus Echtle, Dieter Hammer, and David Powell, editors. Dependable computing — EDCC-1: first European Dependable Computing Conference, Berlin, Germany, October 4–6, 1994: proceedings, volume 852 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58426-9. LCCN QA76.9.F38 E33 1994.

Engelsen:2014:POS

[EHP14] Søren B. Engelsen, Peter I. Hansen, and Serge Pérez. POLYS 2.0: An open source software package for building threedimensional structures of polysaccharides. Biopolymers, 101 (7):733–743, 2014. ISSN 0006-3525 (print), 1097-0282 (electronic).

Eigler:2003:MPU

[Eig03] Frank Ch. Eigler. Mudflap: pointer use checking for C/C++. In Hutton et al. [HDR03], pages 57–69. ISBN

???? LCCN ???? URL http://www.linux.org.uk/~ajh/
gcc/gccsummit-2003-proceedings.pdf.

Ellis:2001:LOS

[EJS^{+}01] Steven Ellis, David Johnson, Mike Schmit, Jenny Jones, Simon Cooke, and Kurt Granroth. Letters: Open source Cobol; setting the Debian record straight; back to basics; load testing Web sites; open source hat tricks; KDE insider. Dr. Dobb’s Journal of Software Tools, 26(7):10, 12, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/.

Emery:2003:LHC

[EKJ^{+}03] Van Emery, Kathy, Jeremy, Nuno Vasconcellos, Craig, Mark Alford, Hiroshi Iwatani, Jesper Christensen, Robin Rowe, Mike Hjorleifsson, Ian, Tariq, LT, Charles, Chris Bruner, and John. Letters: Happy Chinese New Year, Tux; debit on the left, credit on the right; network desktop archive; MST helps Brazil’s poor; “wardialing” in 1979; GNOME 2 drops features of version 1; HTTP user-agent in Mozilla; ready to make movies; Linux training?; we are not riffraff; life without LJ is pain; don’t try to mimic another OS; Scribus progress; put maddog’s letter on the Web; freedom threatens some companies. Linux Journal, 110:6, 8, June 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/www.linuxjournal.com/article/6770.

Ehrig:1991:GGT

[EKR91] H. Ehrig, H.-J. Kreowski, and G. Rozenberg, editors. Graph grammars and their application to computer science: 4th international workshop, Bremen, Germany, March 5–9, 1990: proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-54478-X (New York), 3-540-54478-X (Berlin). LCCN QA75.5.G72 1991.

Egbring:2010:POS

[EKUR10] Marco Egbring, Gerd A. Kullak-Ublick, and Stefan Russmann. Phynx: an open source software solution supporting data management and web-based patient-level data review for drug safety studies in the general practice research database and other health care databases. Pharmacoepidemiology and Drug Safety, 19(1):38–44, 2010. ISSN 1053-8569 (print), 1099-1557 (electronic).

Laila:2023:MCO

[eLAA^{+}23] Umm e Laila, Najeed Ahmed, Asad Arfeen, Agha Yasir Ali, Mohammad Khurram, and Muzammil Ahmed Khan. Missioncritical open-source software adoption model validation using Partial Least Square — Structural Equation Modeling. Journal of Software: Evolution and Process, 35(2):e2514:1– e2514:??, February 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ellul:2012:CFO

[Ell12] Claire Ellul. Can free (and open source) software and data be used to underpin a self-paced tutorial on spatial databases? Transactions in GIS, 16(4):435–454, 2012. ISSN 1361-1682 (print), 1467-9671 (electronic).

Eigenmann:1993:PTO

[EM93] R. Eigenmann and P. McClaughry. Practical tools for optimizing parallel programs. In Tentner [Ten93], pages 160–165.

Embry:2006:TCS

[Emb06] Randall P. Embry. Three case studies in community-oriented, open-source software development. Linux Journal, 2006(142): 9, February 2006. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Easley:2003:MOS

[EMD03] Robert F. Easley, John G. Michel, and Sarv Devaraj. The MP3 open standard and the music industry’s response to Internet piracy. Communications of the ACM, 46(11):90–96, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ellis:2007:CHO

[EMdL^{+}07] Heidi J. C. Ellis, Ralph A. Morelli, Trishan R. de Lanerolle, Jonathan Damon, and Jonathan Raye. Can humanitarian open-source software development draw new students to CS? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):551–555, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engelfriet:2010:COS

[Eng10] Arnoud Engelfriet. Choosing an open source license. IEEE Software, 27(1):48–49, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ensmenger:2004:OSL

[Ens04] Nathan Ensmenger. Open source’s lessons for historians. IEEE Annals of the History of Computing, 26(4):104, 102– 103, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://csdl.

computer.org/dl/mags/an/2004/04/a4104.htm;   http:
//csdl.computer.org/dl/mags/an/2004/04/a4104.pdf.

Ensmenger:2005:SOS

[Ens05] Nathan Ensmenger. The Success of Open Source (review). Technology and Culture, 46(4):860–862, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
192291.

Epplin:2000:IDH

[Epp00] Jerry Epplin. Inside Debian Hurd. Dr. Dobb’s Journal of Software Tools, 25(12):21–22, 24, 26, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1999:EOS

[Eri99] Jonathan Erickson. Editorial: Open season. Dr. Dobb’s Journal of Software Tools, 24(2):8, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/

1999/9902/9902toc.htm;   http://www.ddj.com/ddj/1999/
1999_02/9902q/9902q.htm. Comments on the GNU open

software approach to software development, compared to the Microsoft model.

Erickson:2000:EOS

[Eri00] Jonathan Erickson. Editorial: Open Source, open projects. Dr. Dobb’s Journal of Software Tools, 25(3):10, March 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2001:EOR

[Eri01] Jonathan Erickson. Editorial: Open records, closed doors. Dr. Dobb’s Journal of Software Tools, 26(6):8, June 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/. Essay on open source code issues in government records.

Ertl:1994:PFE

[Ert94] M. A. Ertl. A portable Forth engine. In Anonymous [Ano94a], pages 253–257.

Eskandani:2023:UJF

[ES23] Nafise Eskandani and Guido Salvaneschi. The uphill journey of FaaS in the open-source community. The Journal of Systems and Software, 198(??):??, April 2023. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121222002655.

Eckert:2019:ATI

[ESM19] Remo Eckert, Matthias Stuermer, and Thomas Myrach. Alone or together? Inter-organizational affiliations of open source communities. The Journal of Systems and Software, 149(??):250–262, March 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121218302693.

Espedal:1996:RAB

[Esp96] Norvald Espedal. Realisering av brukergrensesnitt for AML i GNU Emacs. Hovedoppgave, ????, Stavanger, Norway, 1996. vii + 58 pp.

Esteve:2006:PPC

[Est06] Asunción Esteve. Patent protection of computer-implemented inventions vis-à-vis open source software. The Journal of World Intellectual Property, 9(3):276–300, 2006. ISSN 1422-2213 (print), 1747-1796 (electronic).

Eubanks:2005:WCJ

[Eub05] Brian D. Eubanks. Wicked cool Java: code bits, open-source libraries, and project ideas. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-061-5. xvii + 224 pp. LCCN QA76.73.J38 E92 2005; QA76.73.J38 E92 2005eb. URL http:

//www.oreilly.com/catalog/9781593270612.

Eckert:2001:HIU

[EW01] J. Eckert and T. Woigk. Haftung im Umfeld von Open Source. (German) [Liabilities in the field of Open Source]. World-Wide Web document., December 3, 2001.

Ewers:2018:OSR

[Ewe18] Helge Ewers. Open-source recombinant monoclonal secondary nanobodies. Journal of Cell Biology, 217(3):809–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/
3/809.

Eichler:2005:CJT

[EXA^{+}05] M. L. Eichler, P. R. Xavier, R. C. Araujo, R. C. Forte, and J. C. DelPino. Carbopolis: a Java technology-based free software for environmental education installing Carbopolis. Journal of Computers in Mathematics and Science Teaching, 24 (1):43–72, 2005. CODEN ???? ISSN 0731-9258 (print), 1943-5908 (electronic).

False:2003:MOS

[Fal03] Jana False. Markenrecht und Open Source — Sinn oder Unsinn? (German). [Trademark rights and open source—sense or nonsense?]. World-Wide Web document., 2003. URL http://

www.ifross.de/Fremdartikel/VortragLinuxTag2003.pdf.

Farris:1991:GYF

[Far91] Rick Farris. Get Yer Free Software Here! UNIX/world, 8(12): 95–??, December 1991. ISSN 0739-5922.

Farris:1992:MEU

[Far92] Rick Farris. Modern editors for UNIX. UNIX/world, 9(2): 73–80, February 1992. ISSN 0739-5922.

Farnqvist:2005:NTM

[Fär05] Tommy Färnqvist. Number theory meets cache locality: efficient implementation of a small prime FFT for the GNU multiple precision arithmetic library. Examensarbete, Institutionen för numerisk analys och datalogi, Kungliga Tekniska högskolan, Stockholm, Sweden, 2005. 37 pp.

Farrow:2006:WOS

[Far06] Rik Farrow. WISPER: Open source, long-distance wireless. ;login: the USENIX Association newsletter, 31(5):??, October 2006. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/publications/login/october-
2006-volume-31-number-5/wisper-open-source-long-distance-
wireless.

Faridian:2023:LOI

[Far23] Parisa Haim Faridian. Leading open innovation: the role of strategic entrepreneurial leadership in orchestration of value creation and capture in GitHub open source communities. Technovation, 119(??):??, January 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0166497222000931.

Favario:2023:TFO

[Fav23] Leonardo Favario. Toward a free and open source-driven public sector: an Italian journey. IEEE Software, 40(4):55–61, 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Felberg:2017:POS

[FBY^{+}17] Lisa E. Felberg, David H. Brookes, Eng-Hui Yap, Elizabeth Jurrus, Nathan A. Baker, and Teresa Head-Gordon. PB-AM: an open-source, fully analytical linear Poisson–Boltzmann solver. Journal of Computational Chemistry, 38(15):1275– 1282, June 5, 2017. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Furtado:2021:HSO

[FCTP21] Leonardo B. Furtado, Bruno Cartaxo, Christoph Treude, and Gustavo Pinto. How successful are open source contributions from countries with different levels of human development? IEEE Software, 38(2):58–63, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Foundation:1992:CGC

[FD92] Free Software Foundation and D. J. Delorie. CUG359 — GNU C/C++ for 386. C Users Journal, 10(3):117–??, March 1992. ISSN 0898-9788.

Feil:19xx:OSS

[Feixx] Thomas Feil. Open Source Software: eine rechtliche Risikoanalyse. (German) [Open Source software: a legal risk analysis]. World-Wide Web document., 19xx. URL http:

//www.recht-freundlich.de/download/OSS_Rechtliche_
Informationen.pdf.

Feitelson:2023:WDA

[Fei23] Dror G. Feitelson. “We do not appreciate being experimented on”: Developer and researcher views on the ethics of experiments on open-source projects. The Journal of Systems and Software, 204(??):??, October 2023. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121223001693.

Feldman:1993:AAU

[Fel93] M. B. Feldman. Ada83 and Ada9x in the universities: filling the pipeline. In ACM [ACM93b], page 106. ISBN 0-89791-609-3. LCCN QA76.73.A16W37 1993.

Ferris:2003:ACO

[Fer03] Paul Ferris. The age of corporate open source enlightenment. ACM Queue: Tomorrow’s Computing Today, 1(5):34– 44, July/August 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Feller:2005:OSA

[FFH^{+}05] Joseph Feller, Brian Fitzgerald, Scott Hissam, Karim Lakhani, and Walt Scacchi. Open Source Application Spaces: 5th Workshop on Open Source Software Engineering. ACM SIGSOFT Software Engineering Notes, 30(6):1–2, November 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feller:2005:CCC

[FFHL05] Joseph Feller, Brian Fitzgerald, Scott Hissam, and Karim Lakhani. Collaboration, conflict and control: report on the 4th Workshop on Open Source Software Engineering. ACM SIGSOFT Software Engineering Notes, 30(3):1–2, May 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feller:2007:PFO

[FFHL07] Joseph Feller, Brian Fitzgerald, Scott A. Hissam, and Karim R. Lakhani. Perspectives on Free and Open Source Software. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-56227-8. xxxi + 538 pp. LCCN ???? URL

http://mitpress.mit.edu/books/chapters/0262562278.

pdf. Foreword by Michael Cusumano. Epilogue by Clay Shirky.

Feller:2001:MSB

[FFvdH01] Joseph Feller, Brian Fitzgerald, and André van der Hoek. Making sense of the bazaar: 1st Workshop on Open Source Software Engineering. ACM SIGSOFT Software Engineering Notes, 26(6):51–52, November 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fischer:1985:MWG

[FG85] Gerhard Fischer and Rul Gunzenhäuser. Methoden und Werkzeuge zur Gestaltung benutzergerechter Computersysteme, volume 1 of Mensch Computer Kommunikation. de-Gruyter, Berlin, Germany; New York, NY, USA, 1985. ISBN 3-11-010070-3. xi + 282 pp. Also published in/as: summary rcvd Sep.1984.

Foster:1992:EIH

[FG92] C. E. Foster, III and H. C. Grossman. An empirical investigation of the Haifa register allocation technique in the GNU C compiler. In IEEE [IEE92d], pages 776–779 (vol. 2). ISBN 0-7803-0494-2. LCCN TK 7801 I117 1992. Two volumes. IEEE Catalog No. 92CH3094-0.

Friard:2016:BFV

[FG16] Olivier Friard and Marco Gamba. BORIS: a free, versatile open-source event-logging software for video/audio coding and live observations. Methods in Ecology and Evolution, 7(11): 1325–1330, 2016. ISSN 2041-210X.

Fagerholm:2014:OOS

[FGBM14] Fabian Fagerholm, Alejandro Sanchez Guinea, Jay Borenstein, and Jurgen Munch. Onboarding in open source projects. IEEE Software, 31(6):54–61, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (elec-

tronic). URL http://csdl.computer.org/csdl/mags/so/
2014/06/mso2014060054-abs.html.

Florence:2011:OSS

[FHH11] Robby Florence, Faisal Hossain, and David Huddleston. An open-source software for interactive visualization using C++ and OpenGL: Applications to stochastic theory education in water resources engineering. Computer Applications in Engineering Education, 19(1):48–55, 2011. CODEN CAPEED. ISSN 1061-3773 (print), 1099-0542 (electronic).

Fousse:2007:MMP

[FHL^{+}07] Laurent Fousse, Guillaume Hanrot, Vincent Lefèvre, Patrick Pélissier, and Paul Zimmermann. MPFR: a multiple-precision binary floating-point library with correct rounding. ACM Transactions on Mathematical Software, 33(2):1–15, June 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fiedler:1988:MBF

[Fie88] David Fiedler. More and better free software. C Users Journal, 6(11):91–??, November 1988. ISSN 0898-9788.

Fiedler:1989:HGS

[Fie89] David Fiedler. How to get source from the GNU project. C Users Journal, 7(1):109–??, January 1989. ISSN 0898-9788.

Fiedler:1990:FSH

[Fie90a] David Fiedler. The free software hit parade: a quick review of the most popular free Unix software. BYTE Magazine, 15(8): 85–86, 88, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:FSW

[Fie90b] David Fiedler. Free software!: When it comes to userdeveloped Unix programs, there is such a thing as a free lunch. BYTE Magazine, 15(6):97, 100, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fielding:1999:SLA

[Fie99] Roy T. Fielding. Shared leadership in the Apache project. Communications of the ACM, 42(4):42–43, April 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
journals/cacm/1999-42-4/p42-fielding/.

Finseth:1980:TPTa

[Fin80a] Craig A. Finseth. Theory and practice of text editors, or, a cookbook for an Emacs. Thesis (b.s.), M.I.T., Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1980. 103 pp. Supervised by David P. Reed.

Finseth:1980:TPTb

[Fin80b] Craig A. Finseth. Theory and practice of text editors or a cookbook for an emacs. Technical memoranda 165, Massachusetts Institute of Technology, Laboratory for Computer Science, Cambridge, MA, USA, 1980. 106 pp.

Finseth:1991:CTE

[Fin91] Craig A. Finseth. The Craft of Text Editing: Emacs for the Modern World. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97616-7 (New York), 3-540-97616-7 (Berlin). xii + 220 pp. LCCN QA76.76.T49 F56 1991. Contains extensive discussion of design issues for text editors, with examples from Emacs. Appendix B gives sources of numerous Emacs implementations. Appendix D summarizes the TECO command set.

Finston:2022:GEE

[Fin22a] Laurence Finston. A graphical ellipse envelope construction with GNU 3DLDF. TUGboat, 43(3):333–339, 2022. CODEN

???? ISSN 0896-3207. URL https://tug.org/TUGboat/
tb43-3/tb135finston-ellipse.pdf.

Finston:2022:IPG

[Fin22b] Laurence Finston. An introduction to GNU 3DLDF. TUGboat, 43(3):319–332, 2022. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb43-3/tb135finston-3dldf.
pdf.

Fioretti:2003:FSF

[Fio03] Marco Fioretti. Free software, free society: Selected essays of Richard M. Stallman. Linux Journal, 109:61–65, 67, May 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Fischer:1969:GS

[Fis69] Heinz Eduard Fischer. Gnu soup. Printed for Fred Cogswell, Fiddlehead Poetry Books, 1969. 48 pp. LCCN PR6056.I77 G6. Limited ed. of 500 copies. Poems.

Fitzgerald:2004:CLO

[Fit04] Brian Fitzgerald. A critical look at open source. Computer, 37(7):92–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/07/r7092.htm;   http:
//csdl.computer.org/dl/mags/co/2004/07/r7092.pdf.

Fitzgerald:2011:OSS

[Fit11] Brian Fitzgerald. Open source software: Lessons from and for software engineering. Computer, 44(10):25–30, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fraser:1990:LTE

[FK90] C. W. Fraser and B. Krishnamurthy. Live text (editing). Software—Practice and Experience, 20(8):851–858, August 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Feldman:1999:APS

[FK99] Michael B. Feldman and Elliot B. Koffman. Ada 95: problem solving and program design. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-36123-X. xv + 784 pp. LCCN QA76.73.A35 F43 1999. Includes CD-ROM.

Fitzgerald:2004:DIS

[FK04] Brian Fitzgerald and Tony Kenny. Developing an information systems infrastructure with Open Source software. IEEE Software, 21(1):50–55, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fursin:2011:MGM

[FKM^{+}11] Grigori Fursin, Yuriy Kashnikov, Abdul Wahid Memon, Zbigniew Chamski, Olivier Temam, et al. Milepost GCC: Machine learning enabled self-tuning compiler. International Journal of Parallel Programming, 39(3):296–327, June 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7458&volume=39&issue=3&spage=296.

Feinberg:2015:COS

[FL15] Jonathan Feinberg and Hans Petter Langtangen. Chaospy: an open source tool for designing methods of uncertainty quantification. Journal of Computational Science, 11:46–57, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-

7511 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S1877750315300119.

Fang:2016:OSV

[FL16] Xin Fang and Miriam Leeser. Open-source variable-precision floating-point library for major commercial FPGAs. ACM Transactions on Reconfigurable Technology and Systems, 9 (3):1–17, July 2016. ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2851507.

Falls:2016:XVR

[FLA^{+}16] Zackary Falls, David C. Lonie, Patrick Avery, Andrew Shamp, and Eva Zurek. XtalOpt version r9: an open-source evolutionary algorithm for crystal structure prediction. Computer Physics Communications, 199(??):178–179, February 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465515003690.

Floyd:1994:CLT

[Flo94] Michael A. Floyd. A conversation with Linus Torvalds. Dr. Dobb’s Journal of Software Tools, 19(5):27–??, May 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Flynn:1987:GR

[Fly87a] Anita M. Flynn. Gnat robots. AI Expert, 2(12):34–42, December 1987. ISSN 0888-3785.

Flynn:1987:GRH

[Fly87b] Anita M. Flynn. Gnat robots (and how they will change robotics). Technical Report Working Paper 295, Massachusetts Institute of Technology, Cambridge, MA, USA, June 1987.

Francalanci:2010:EAB

[FM10] Chiara Francalanci and Francesco Merlo. Empirical analysis of the bug fixing process in open source projects. International Journal of Computer Systems Science and Engineering, 25(4): ??, July 2010. CODEN CSSEEI. ISSN 0267-6192.

Franklin:2002:PAS

[FMA02] Michael Franklin, Bongki Moon, and Anastassia Ailamaki, editors. Proceedings of the ACM SIGMOD International Conference on Management of Data, June 3–6, 2002, Madison, WI, USA. ACM Press, New York, NY 10036, USA, 2002. ISBN ???? LCCN ???? ACM order number 475020.

Fernandez:2019:OSI

[FMFZ19] Daniel Méndez Fernández, Martin Monperrus, Robert Feldt, and Thomas Zimmermann. The open science initiative of the Empirical Software Engineering journal. Empirical Software Engineering, 24(3):1057–1060, June 2019. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
article/10.1007/s10664-019-09712-x;   http://link.
springer.com/content/pdf/10.1007/s10664-019-09712-
x.pdf.

Fursin:2008:MGM

[FMT^{+}08] Grigori Fursin, Cupertino Miranda, Olivier Temam, Mircea Namolaru, Elad Yom-Tov, Ayal Zaks, Bilha Mendelson, Edwin Bonilla, John Thomson, Hugh Leather, Chris Williams, and Michael O. Boyle. MILEPOST GCC: machine learning based research compiler. In ????, editor, GCC Summit ’08, pages 1–13. ????, 2008. ISBN ???? LCCN ???? URL http:

//gcc-ici.sourceforge.net/papers/fmtp2008.pdf.

Faridian:2021:AAA

[FN21] Parisa Haim Faridian and Donald O. Neubaum. Ambidexterity in the age of asset sharing: Development of dynamic capabilities in open source ecosystems. Technovation, ??(??): ??, ???? 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497218308186.

Fogel:2006:POS

[Fog06] Karl Fogel. Producing open source software: how to run a successful free software project. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-00759-0. xx + 279 pp. LCCN QA76.76.D47 F64 2006.

URL http://www.oreilly.com/catalog/9780596007591.

Ford:2007:OVC

[For07] Richard Ford. Open vs. closed: which source is more secure? ACM Queue: Tomorrow’s Computing Today, 5(1): 32–38, February 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Forbes:2012:CSO

[For12] Megan Forbes. CollectionSpace: a story of open-source software development and user-centered design. Bulletin of the American Society for Information Science and Technology, 38 (3):22–26, 2012. ISSN 2373-9223.

Fowler:1993:SS

[Fow93] G. Fowler. The shell as a service. In Anonymous [Ano93c], pages 267–277.

Fowler:2000:QOS

[Fow00] J. E. Fowler. QccPack: an open-source software library for quantization, compression, and coding. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http:/

/ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
838201;   http://www.ece.msstate.edu/~fowler/QccPack.

This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Foxwell:2008:RDG

[Fox08] Harry J. Foxwell. Review of ‘The Definitive Guide to the Xen Hypervisor’ (Prentice Hall Open Source Software Development Series), David Chisnall, Prentice Hall PTR, 2007, $49.99, ISBN 0-13-234971-X. ACM Queue: Tomorrow’s Computing Today, 6(3):69, May/June 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Fine:1994:UCP

[FP94] V. Fine and E. Potrebenikova. Using CERNLIB, PAW and GEANT packages on PCs under DJGPP. In Becks and Perret-Gallix [BPG94], pages 151–156. ISBN 981-02-1699-8. LCCN QC793.47.E4I58 1993.

Ferrante:1995:FAS

[FP95] Jeanne Ferrante and David Padua, editors. Fifth ACM SIG-PLAN Symposium on Principles and Practice of Parallel Programming, PPoPP, volume 30(8) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, August 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fu:2023:FOS

[FQYS23] Wenwen Fu, Wei Quan, Jinli Yan, and Zhigang Sun. Fenglin-I: An open-source time-sensitive networking chip enabling agile customization. IEEE Transactions on Computers, 72(1): 140–153, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Franky:1995:DPS

[Fra95] M. C. Franky. DGDBM: programming support for distributed transactions over replicated files. Operating Systems Review, 29(3):64–74, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Franceschini:2013:HMV

[Fra13] Monica Franceschini. How to maximize the value of big data with the open source SpagoBI suite through a comprehensive approach. Proceedings of the VLDB Endowment, 6(11):1170– 1171, August 2013. CODEN ???? ISSN 2150-8097.

Frazelle:2019:OSF

[Fra19] Jesse Frazelle. Open source firmware. Communications of the ACM, 62(10):34–38, October 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:

//cacm.acm.org/magazines/2019/10/239673/fulltext.

Frank:2015:FMG

[FRAK15] Florian Frank, Balthasar Reuter, Vadym Aizinger, and Peter Knabner. FESTUNG: A MATLAB/GNU Octave toolbox for the discontinuous Galerkin method, Part I: Diffusion operator. Computers and Mathematics with Applications, 70(1):11– 46, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0898122115001820.

Frantz:2019:ROS

[FRBRF19] Rafael Z. Frantz, Matheus H. Rehbein, Rodolfo Berlezi, and Fabricia Roos-Frantz. Ranking open source application integration frameworks based on maintainability metrics: a review of five-year evolution. Software—Practice and Experience, 49(10):1531–1549, October 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

FSF:1987:GBN

[Fre87] GNU’s bulletin: newsletter of the Free Software Foundation, page various, 1987. ISSN 1075-7813. The Foundation, Cambridge, MA, USA.

Frey:2023:HWL

[Fre23] Rasmus Frey. How we lead successful open-source collaborations in the Danish public sector. IEEE Software, 40(4):19–24, 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Friesenhahn:1997:AMP

[Fri97] Bob Friesenhahn. Autoconf makes for portable software — use of OS features and a freeware scripting utility solves application portability across various flavors of Unix. BYTE Magazine, 22(11):45–46, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friedman:2006:EBU

[Fri06] Nat Friedman. Eof: bringing usability to open source. Linux Journal, 2006(141):16, January 2006. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Friedman:2016:SLI

[Fri16] Alon Friedman. Statistics for library and information services: a primer for using open source R software for accessibility and visualization. Rowman and Littlefield, Lanham, UK, 2016. ISBN 1-4422-4992-7 (hardcover), 1-4422-4993-5 (e-book). xvi + 355 pp. LCCN Z669.8 .F75 2016.

Faulkner:2001:RNR

[FSB^{+}01] Andy Faulkner, Rich Smith, Brad Baylor, Jim Bailey, Paul Mack, Jim Lemaster, and Tom Hartel. Running a net radio station with open-source software. Linux Journal, 81:136, 138, 140, 142–143, January 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Foxwell:2009:PON

[FT09] Harry Foxwell and Christine Tran. Pro OpenSolaris: a New Open Source OS for Linux Developers and Administrators. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1891-6. xxi + 254 pp. LCCN QA76.76.O63 F59733 2009.

Feng:2023:SOS

[FTZ^{+}23] Kaikai Feng, Peng Tian, Jun Zhang, Fei Fei, and Dongsheng Wen. SPARTACUS: an open-source unified stochastic particle solver for the simulation of multiscale nonequilibrium gas flows. Computer Physics Communications, 284 (??):Article 108607, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003265.

Fuggetta:2003:OSS

[Fug03] Alfonso Fuggetta. Open source software — an evaluation. The Journal of Systems and Software, 66(1):77–90, April 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Furuta:1990:EPI

[Fur90] R. Furuta, editor. EP90: proceedings of the International Conference on Electronic Publishing, Document Manipulation and Typography, Gaithersburg, Maryland, September 1990. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-40246-8. LCCN Z286.E43I58 1990.

Fujioka:2012:AGM

[FVD^{+}12] Ei Fujioka, Edward Vanden Berghe, Ben Donnelly, Julio Castillo, Jesse Cleary, Chris Holmes, Sean McKnight, and Patrick Halpin. Advancing global marine biogeography research with open-source GIS software and cloud computing. Transactions in GIS, 16(2):143–160, 2012. ISSN 1361-1682 (print), 1467-9671 (electronic).

Fukuda:2010:MSI

[FvdHJ10] Komei Fukuda, Joris van der Hoeven, and Michael Joswig, editors. Mathematical software – ICMS 2010: third international congress on mathematical software, Kobe, Japan, September 13–17, 2010, proceedings, volume 6327 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-15581-2 (softcover). LCCN ????

Franke:2003:SHU

[FvH03] Nikolaus Franke and Eric von Hippel. Satisfying heterogeneous user needs via innovation toolkits: the case of Apache security software. Research Policy, 32(7):1199–1215, 2003. CODEN ???? ISSN 0048-7333 (print), 1873-7625 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0048733303000490.

Farfour:2018:SUG

[FY18] Mohammed Farfour and Wang Jung Yoon. Seismic UNIX and GNU Octave for VSP Data Processing and Interpretation, pages 73–92. Wiley, New York, NY, USA, 2018. ISBN 1-119-22751-8.

Fyk:1997:IPP

[Fyk97] J. Fyk. Intonational protention in the performance of melodic octaves on the violin. Lecture Notes in Computer Science, 1317:421–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganten:2000:DGL

[G^{+}00] Peter H. Ganten et al. Debian GNU, Linux: Grundlagen, Installation, Administration und Anwendung. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-65841-6, 3-540-66384-3. xvi + 792 pp. LCCN ????

Galassi:2001:GSL

[G^{+}01] Mark Galassi et al. GNU scientific library reference manual: edition 1.0 for GSL version 1.0. Network Theory Ltd., Bristol, UK, 2001. ISBN 0-9541617-0-X. xvi + 580 pp. LCCN ????

Ganten:2002:DGL

[G^{+}02] Peter H. Ganten et al. Debian GNU/Linux Power Pack: Grundlagen, Installation, Administration und Anwendung. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2002. ISBN 3-540-43250-7, 3-540-43267-1. 800 (est.) pp. LCCN ???? Includes four CD-ROMs.

Geiselhart:2006:IZV

[G^{+}06] Gregory Geiselhart et al., editors. IBM z/VM and Linux on IBM System z: virtualization cookbook for Red Hat Enterprise Linux 4. Number SG24-7272-00 in IBM redbooks. IBM Corporation, San Jose, CA, USA, 2006. ISBN 0-7384-9495-X (paperback). xiv + 218 pp. LCCN QA76.76.O63 I28 2006.

Gacek:2004:MMO

[GA04a] Cristina Gacek and Budi Arief. The many meanings of Open Source. IEEE Software, 21(1):34–40, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ganten:2004:DGLa

[GA04b] Peter H. Ganten and Wulf Alex. Debian GNU/Linux: Grundlagen, Installation, Administration und Anwendung. (German) [Debian GNU/Linux: Basis, Installation, Administration, and Use]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-43267-1. xxii + 946 pp. LCCN ????

Gadol:1988:SCL

[Gad88] S. Gadol. SPE — a Common Lisp environment on workstations. In Anonymous [Ano88b], pages 399–404.

Gagne:2002:WOS

[Gag02] Marcel Gagné. WINE: The open source way to run Windows applications. Sys Admin: The Journal for UNIX Systems Administrators, 11(1):8, 10, 12, 14, January 2002. CODEN SYADE7. ISSN 1061-2688.

Galler:1960:LEC

[Gal60] Bernard A. Galler. Letters to the Editor: Cost of software. Communications of the ACM, 3(4):A12, April 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gallivan:2001:SBB

[Gal01] Michael J. Gallivan. Striking a balance between trust and control in a virtual organization: a content analysis of open source software case studies. Information Systems Journal, 11 (4):277–304, 2001. ISSN 1350-1917 (print), 1365-2575 (electronic).

Galli:2004:GCP

[Gal04] Peter Galli. GPL could put heat on Microsoft. eWeek: The Enterprise Newsweekly, 21(48):1, 18, December 29, 2004.

Galassi:2009:GSL

[Gal09] Mark Galassi. GNU scientific library: reference manual. A GNU manual. Network Theory, Bristol, UK, third edition, 2009. ISBN 0-9546120-7-8. xvi + 573 pp. LCCN ????

Galvin:2010:PATb

[Gal10] Peter Baer Galvin. Pete’s all things Sun: Open source and free deduplication. ;login: the USENIX Association newsletter, 35(2):??, April 2010. CODEN LOGNEM. ISSN

1044-6397. URL https://www.usenix.org/publications/
login/april-2010-volume-35-number-2/petes-all-things-
sun-open-source-and-free.

Gancarz:1995:UP

[Gan95] Mike Gancarz. The UNIX philosophy. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1995. ISBN 1-55558-123-4. xix + 151 pp. LCCN QA76.76.O63G365 1995.

Ganten:2004:DGLb

[Gan04] Peter H. Ganten. Debian-GNU-Linux-Powerpack. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-66384-3. ???? pp. LCCN ????

Gangadharan:2017:OSS

[Gan17] G. R. Gangadharan. Open source solutions for cloud computing. Computer, 50(1):66–70, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/csdl/mags/co/2017/01/
mco2017010066-abs.html.

Garzik:2000:GCR

[Gar00] Jeff Garzik. Glibc: a comprehensive reference to GNU/Linux libC. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 1-57870-202-X. 500 (est.) pp. LCCN ???? Edited by Laurie Petrycki and others.

Gardinier:2009:OSD

[Gar09] Mark Gardinier. Open source development of a safety critical dual redundant (ada95/C++) signal control program environment (SCOPE). ACM SIGADA Ada Letters, 29(3):23–30, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goth:2001:NOM

[GAS^{+}01] Greg Goth, Scott Lorenz Andresen, Gil Alexander Shif, Paula Powers, and Scott Lorenz Andresen. In the news: The open market woos open source; awards & recognition: Anup Ghosh named Young Engineer of the Year; nextgeneration entertainment consoles: Not all fun and games; Asia Pacific region poised to follow India in software development revolution. IEEE Software, 18(2):104–111, March/April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459

(electronic). URL http://dlib.computer.org/so/books/
so2001/pdf/s2104.pdf.

Gaudeul:2003:APC

[Gau03] Alexandre Gaudeul. The (La)TEX project: a case study of open source software. TUGboat, 24(1):132–145, 2003. ISSN 0896-3207.

Gaudeul:2007:DOS

[Gau07] Alex Gaudeul. Do open source developers respond to competition? The L^{A}TEX case study. Review of Network Economics, 6(2):239–263, June 2007. CODEN ???? ISSN 2194-5993 (print), 1446-9022 (electronic). URL https:/

/econpapers.repec.org/article/bpjrneart/v_3a6_3ay_
3a2007_3ai_3a2_3an_3a9.htm;   https://www.degruyter.
com/view/j/rne.2007.6.issue-2/rne.2007.6.2.1119/rne.
2007.6.2.1119.xml.

FSF:1987:GM

[GAW87a] The GAWK manual. Free Software Foundation, Inc. 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1987. Also available via ANONYMOUS FTP to prep.ai.mit.edu. See also [AKW88].

FSF:gawk

[GAW87b] The GAWK manual. Free Software Foundation, Inc. 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1987. Also available via ANONYMOUS FTP to prep.ai.mit.edu. See also [AKW88].

Gay:2002:FSF

[Gay02] Joshua Gay, editor. Free Software, Free Society: Selected Essays of Richard M. Stallman. GNU Press, Boston, MA, USA, 2002. ISBN 1-882114-98-1. 220 (est.) pp. LCCN ???? US$24.95.

Geiring:1994:GAR

[GB94] E. W. Geiring and T. Baker. The GNU Ada Runtime Library (GNARL). In ACM [ACM94], pages 97–107. ISBN 0-89791-684-0. LCCN ????

Goth:2000:NNG

[GB00] Greg Goth and Cheryl Baltes. In the news: The next gold rush: Application service providers stake their claims in a redhot market: Are consumer rights at risk?: Standards-based training: Grand opening: SPC’s resource center. IEEE Software, 17(2):96–100, March/April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/books/so2000/pdf/s2096.pdf.

Granvilliers:2006:ARI

[GB06] Laurent Granvilliers and Frédéric Benhamou. Algorithm 852: RealPaver: an interval solver using constraint satisfaction techniques. ACM Transactions on Mathematical Software, 32(1):138–156, March 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://

www.sciences.univ-nantes.fr/info/perso/permanents/
granvil/papers/gbtoms05.pdf.

Greenberg:2020:EFS

[GB20] Michael Greenberg and Austin J. Blatt. Executable formal semantics for the POSIX shell. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):43:1–43:30, Jan-

uary 2020. URL https://dl.acm.org/doi/abs/10.1145/
3371111.

Gonzalez-Barahona:2021:BHF

[GB21] J. M. Gonzalez-Barahona. A brief history of free, open source software and its communities. Computer, 54(2):75–79, 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garbow:1977:MER

[GBDM77] B. S. Garbow, J. M. Boyle, J. J. Dongarra, and C. B. Moler. Matrix Eigensystem Routines—EISPACK Guide Extension, volume 51 of Lecture Notes in Computer Science, Editors: G.

Goos and J. Hartmanis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-08254-9, 3-540-08254-9. viii + 343 pp. LCCN QA193 .M381, QA267.A1,L43 no. 51.

Galland:2016:AOS

[GBG^{+}16] Olivier Galland, Håvard S. Bertelsen, Frank Guldstrand, Luc Girod, Rikke F. Johannessen, Fanny Bjugger, Steffi Burchardt, and Karen Mair. Application of open-source photogrammetric software MicMac for monitoring surface deformation in laboratory models. Journal of Geophysical Research: Solid Earth, 121(4):2852–2872, 2016. ISSN 2169-9313 (print), 2169-9356 (electronic).

Gonzalez-Barahona:2013:UHC

[GBICMR13] Jesus M. Gonzalez-Barahona, Daniel Izquierdo-Cortazar, Stefano Maffulli, and Gregorio Robles. Understanding how companies interact with free software communities. IEEE Software, 30(5):38–45, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Giri:2021:AIO

[GCE^{+}21] Davide Giri, Kuan-Lin Chiu, Guy Eichler, Paolo Mantovani, and Luca P. Carloni. Accelerator integration for open-source SoC design. IEEE Micro, 41(4):8–14, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Genova:2017:EOS

[GCK^{+}17] Alessandro Genova, Davide Ceresoli, Alisa Krishtal, Oliviero Andreussi, Robert A. DiStasio, Jr., and Michele Pavanello. eQE: an open-source density functional embedding theory code for the condensed phase. International Journal of Quantum Chemistry, 117(16), August 15, 2017. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

German:2012:MOS

[GD12] Daniel German and Massimiliano Di Penta. A method for open source license compliance of Java applications. IEEE Software, 29(3):58–63, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gibbons:2023:EOS

[GDJG23] Nicholas N. Gibbons, Kyle A. Damm, Peter A. Jacobs, and Rowan J. Gollan. Eilmer: an open-source multiphysics hypersonic flow solver. Computer Physics Communications, 282(??):Article 108551, January 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002703.

Gerlero:2021:EVO

[GDK21] Gabriel S. Gerlero, Santiago Márquez Damián, and Pablo A.

Kler. electroMicroTransport v2107: Open-source tool-

box for paper-based electromigrative separations. Computer Physics Communications, 269(??):Article 108143, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002551.

Galassi:2002:GSL

[GDT^{+}02] Mark Galassi, Jim Davies, James Theiler, Brian Gough, Gerard Jungman, Michael Booth, and Fabrice Rossi. GNU Scientific Library: Reference Manual. Network Theory Ltd., Bristol, UK, second edition, 2002. ISBN 0-9541617-3-4. xvi + 601 pp. LCCN QA76.73.C15.

Galassi:2005:GSL

[GDT^{+}05] Mark Galassi, Jim Davies, James Theiler, Brian Gough, Gerard Jungman, Michael Booth, and Fabrice Rossi. GNU Scientific Library: Reference Manual. Network Theory Ltd., Bristol, UK, second revised edition, 2005. ISBN 0-9541617-3-4. xvi + 601 pp. LCCN QA76.73.C15. URL http://www.network-

theory.co.uk/gsl/manual/.

Gehring:1996:SFS

[Geh96] Robert Gehring. Studienarbeit: Freeware, Shareware und Public Domain: Geschichte, Begrifflichkeit, Urheberrecht und Haftung. (German) [Study work: Freeware, shareware, and public domain: History, concepts, copyright, and liability]. World-Wide Web document., July 1996. URL http://ig.

cs.tu-berlin.de/oldstatic/sa/043.

Gary:2011:AMO

[GEI^{+}11] Kevin Gary, Andinet Enquobahrie, Luis Ibanez, Patrick Cheng, Ziv Yaniv, Kevin Cleary, Shylaja Kokoori, Benjamin Muffih, and John Heidenreich. Agile methods for open source safety-critical software. Software—Practice and Experience, 41(9):945–962, August 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gschwind:2007:OSE

[GEMN07] Michael Gschwind, David Erb, Sid Manning, and Mark Nutter. An open source environment for cell broadband engine system software. Computer, 40(6):37–47, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gengler:1999:NCG

[Gen99] Barbara Gengler. Now cryptography gets the ‘open source’ treatment. Network Security, 1999(6):6, June 1999. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1353485899900582.

German:2003:GPC

[Ger03] Daniel M. German. The GNOME project: a case study of open source, global software development. Software Process: Improvement and Practice, 8(4):201–215, October 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Grinzo:1999:PBC

[GF99] Lou Grinzo and Laryn Fernandes. Programmer’s bookshelf: Clarifying the Open Source movement. Dr. Dobb’s Journal of Software Tools, 24(9):119–120, September 1999. CODEN DDJOEB. ISSN 1044-789X.

Garzarelli:2011:OSS

[GF11] Giampaolo Garzarelli and Riccardo Fontanella. Open source software production, spontaneous input, and organizational learning. The American Journal of Economics and Sociology, 70(4):928–950, 2011. CODEN AJESA3. ISSN 0002-9246 (print), 1536-7150 (electronic).

Gokhin:2017:SBM

[GF17] David S. Gokhin and Velia M. Fowler. Software-based measurement of thin filament lengths: an open-source GUI for distributed deconvolution analysis of fluorescence images. Journal of Microscopy, 265(1):11–20, 2017. ISSN 0022-2720 (print), 1365-2818 (electronic).

Guo:2024:MOS

[GFD^{+}24] Yu-Chen Guo, Fan Feng, An Di, Shi-Qi Lu, and Ji-Chong Yang. MLAnalysis: an open-source program for high energy physics analyses. Computer Physics Communications, 294(??):Article 108957, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003028.

Gyimothy:2005:EVO

[GFS05] T. Gyimothy, R. Ferenc, and I. Siket. Empirical validation of object-oriented metrics on open source software for fault prediction. IEEE Transactions on Software Engineering, 31(10): 897–910, October 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=1542070.

Gu:2016:PFO

[GFZ16] X.-F. Gu, T. Furuhara, and W.-Z. Zhang. PTCLab: free and open-source software for calculating phase transformation crystallography. Journal of Applied Crystallography, 49(3): 1099–1106, 2016. CODEN JACGAR. ISSN 0021-8898 (print), 1600-5767 (electronic).

Grassi:2017:GAI

[GGB17] Paul Grassi, Mike Garcia, and Katie Boeckl. Government adopts an industry approach to open source collaboration. Computer, 50(11):78–81, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/csdl/mags/co/2017/11/
mco2017110078.html.

Gurbani:2005:CSO

[GGH05] Vijay K. Gurbani, Anita Garvert, and James D. Herbsleb. A case study of open source tools and practices in a commercial setting. ACM SIGSOFT Software Engineering Notes, 30(4): 1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gurbani:2010:MCO

[GGH10] Vijay K. Gurbani, Anita Garvert, and James D. Herbsleb. Managing a corporate open source software asset. Communications of the ACM, 53(2):155–159, February 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glinsky:1999:LIO

[GGK99] Peter Glinsky, Frank Gehrke, and Günter Klappheck. Das LINUX im Office Buch: [erfolgreiches Arbeiten mit wichtigen LINUX-Applikationen; die LINUX-Büropakete Staroffice und Applixware; das Bildbearbeitungsprogramm GIMP; die Datenbank PostgreSQL]. Sybex, Düsseldorf, Germany, 1999. ISBN 3-8155-0310-8. xx + 694 pp. LCCN ????

Goggins:2021:MOS

[GGL21] Sean P. Goggins, Matt Germonprez, and Kevin Lumbard. Making open source project health transparent. Computer, 54(8):104–111, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garcia-Garcia:2020:POS

[GGSRMPM20] Adrian Garcia-Garcia, Juan Carlos Saez, José Luis Risco-Martin, and Manuel Prieto-Matias. PBBCache: an opensource parallel simulator for rapid prototyping and evaluation of cache-partitioning and cache-clustering policies. Journal of Computational Science, 42:??, April 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S1877750319303229.

Georgas:2005:RIH

[GGT05] John C. Georgas, Michael M. Gorlick, and Richard N. Taylor. Raging incrementalism: harnessing change with open-source software. ACM SIGSOFT Software Engineering Notes, 30(4): 1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gorecki:2020:HAC

[GHH20] Jan Górecki, Marius Hofert, and Martin Holena. Hierarchical Archimedean copulas for MATLAB and Octave: The HACopula toolbox. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v093i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v093i10/v93i10.pdf.

Gellerich:2004:GBP

[GHL^{+}04] W. Gellerich, T. Hendel, R. Land, H. Lehmann, M. Mueller, P. H. Oden, and H. Penner. The GNU 64-bit PL8 compiler: Toward an open standard environment for firmware development. IBM Journal of Research and Development, 48(3/4): 543–556, ???? 2004. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.research.

ibm.com/journal/rd/483/gellerich.html;   http://www.
research.ibm.com/journal/rd/483/gellerich.pdf.

Gruber:2005:EPA

[GHM^{+}05] O. Gruber, B. J. Hargrave, J. McAffer, P. Rapicault, and T. Watson. The Eclipse 3.0 platform: Adopting OSGi technology. IBM Systems Journal, 44(2):289–??, ???? 2005. CO-DEN IBMSA7. ISSN 0018-8670.

Ghosh:2007:CCO

[Gho07] Rishab Aiyer Ghosh. CODE: Collaborative Ownership and the Digital Economy. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-57326-1. 384 (est.) pp. LCCN ????

Gary:2006:IOS

[GIA^{+}06] Kevin Gary, Luis Ibáñez, Stephen Aylward, David Gobbi, M. Brian Blake, and Keven Cleary. IGSTK: An open source software toolkit for image-guided surgery. Computer, 39(4): 46–53, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gilmore:1988:PBU

[Gil88] John Gilmore. Porting Berkeley Unix through the GNU C compiler. Submitted to USENIX 1988, but rejected as “not research”. Posted by its author to the TUHS mailing list on 21 May 2020., February 19, 1988. URL http://

mcvoy.com/lm/papers/porting-berkeley.pdf;   https:
//minnie.tuhs.org/pipermail/tuhs/2020-May/021204.html.

Gillespie:1993:CM

[Gil93] Dave Gillespie. Calc manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, GNU Emacs version 2.02 edition, 1993. ISBN 1-882114-18-3. x + 572 pp. LCCN ????

Gilbert:2004:EGP

[Gil04] Jill Gilbert. The entrepreneur’s guide to patents, copyrights, trademarks, trade secrets and licensing. Berkley Books, New York, NY, USA, 2004. ISBN 0-425-19409-4 (paperback). xii + 323 pp. LCCN KF2980 .G55 2004. See [Gil05], which serves as an appendix to this book.

Gilbert:2005:LSD

[Gil05] Jill Gilbert. Legal structures for data structures: The impact of patent and copyright law on software innovations. Master’s thesis, Marquette University, Milwaukee, WI, USA, 2005. See also [Gil04].

Gillespie:2006:BRS

[Gil06] Tarleton Gillespie. Book review: Steven Weber: The Success of Open Source. Isis, 97(3):592–593, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/510004.

GDT:2007:GUM

[GIM07] GIMP Documentation Team. GIMP user manual: GNU image manipulation program user manual. SoHoBooks, ????, USA, 2007. ISBN 1-4414-1932-2. 653 pp. LCCN T385 .G5426 2009.

Garcia:2011:KRR

[GJLT11] Saturnino Garcia, Donghwan Jeon, Christopher M. Louie, and Michael Bedford Taylor. Kremlin: rethinking and rebooting gprof for the multicore age. ACM SIGPLAN Notices, 46 (6):458–469, June 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garcia-Jacas:2014:SNU

[GJMPAM^{+}14] César R. García-Jacas, Yovani Marrero-Ponce, Liesner Acevedo-Martínez, Stephen J. Barigye, José R. Valdés-Martiní, and Ernesto Contreras-Torres. Software news and updates: QuBiLS–MIDAS: a parallel free-software for molecular descriptors computation based on multilinear algebraic maps. Journal of Computational Chemistry, 35(18):1395– 1409, July 5, 2014. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Guilan:2002:RCC

[GJS^{+}02] Dai Guilan, Tian Jinlan, Zhang Suqin, Jiang Weidu, and Dai Jun. Retargetable cross compilation techniques: comparison and analysis of GCC and Zephyr. ACM SIGPLAN Notices, 37(6):38–44, June 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Granlund:1992:EBU

[GK92] Torbjörn Granlund and Richard Kenner. Eliminating branches using a superoptimizer and the GNU C compiler. ACM SIGPLAN Notices, 27(7):341–352, July 1992. CO-DEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.

acm.org:80/pubs/citations/proceedings/pldi/143095/
p341-granlund/.

Gaitan:2014:SJC

[GKL^{+}14] J. Cifuentes Gaitan, N. Kirby, A. Lasso, L. Chin, C. Pinter, J. Pignol, G. Fichtinger, and J. Pouliot. SU-E-J-42: Customized deformable image registration using open-source software slicerrt. Medical Physics, 41(6Part7):164, 2014. CODEN MPHYA6. ISSN 2473-4209.

Graham:1982:GCG

[GKM82] Susan L. Graham, Peter B. Kessler, and Marshall K. Mckusick. Gprof: a call graph execution profiler. ACM SIGPLAN Notices, 17(6):120–126, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:2004:GCG

[GKM04] Susan L. Graham, Peter B. Kessler, and Marshall K. McKusick. gprof: a call graph execution profiler. ACM SIGPLAN Notices, 39(4):49–57, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghali:2014:POS

[GKP^{+}14] Fawaz Ghali, Ritesh Krishna, Simon Perkins, Andrew Collins, Dong Xia, Jonathan Wastling, and Andrew R. Jones. ProteoAnnotator — open source proteogenomics annotation software supporting PSI standards. PROTEOMICS, 14(23-24): 2731–2741, 2014. ISSN 1615-9853 (print), 1615-9861 (electronic).

Gamalielsson:2014:SOS

[GL14] Jonas Gamalielsson and Björn Lundell. Sustainability of Open Source software communities beyond a fork: How and why has the LibreOffice project evolved? The Journal of Systems and Software, 89(??):128–145, March 2014. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121213002744.

Glass:1999:NBL

[Gla99] Robert L. Glass. News briefs: Loyal opposition: Of open source, Linux . . . . . . and hype. IEEE Software, 16(1):128– ??, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://dlib.

computer.org/so/books/so1999/pdf/s1128.pdf.

Glass:2000:LOS

[Gla00] Robert L. Glass. Loyal opposition: The sociology of open source: Of cults and cultures. IEEE Software, 17(3):104–105, May/June 2000. CODEN IESOEG. ISSN 0740-7459 (print),

0740-7459 (electronic). URL http://dlib.computer.org/
so/books/so2000/pdf/s3104.pdf.

Glass:2003:EFS

[Gla03a] Robert L. Glass. Error-free software remains extremely elusive. IEEE Software, 20(1):104, 103, January/February 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459

(electronic). URL http://dlib.computer.org/so/books/
so2003/pdf/s1104.pdf.

Glass:2003:PPS

[Gla03b] Robert L. Glass. Practical programmer: a sociopolitical look at open source. Communications of the ACM, 46(11):21–23, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2004:PPL

[Gla04] Robert L. Glass. Practical programmer: a look at the economics of open source. Communications of the ACM, 47(2): 25–27, February 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2008:TME

[Gla08] Robert L. Glass. Two mistakes and error-free software: a confession. IEEE Software, 25(4):96, 95, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

tronic). URL http://bell.computer.org/dlcomments/.

Garcia-Lucas:2017:PPE

[GLCMC17] David García-Lucas, Gabriel Cebrián-Márquez, and Pedro Cuenca. Parallelization and performance evaluation of opensource HEVC codecs. The Journal of Supercomputing, 73(1): 495–513, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Glickstein:1997:WGE

[Gli97] Bob Glickstein. Writing GNU Emacs Extensions. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-261-1. xviii + 215 pp. LCCN QA76.76.T49G56 1997.

Ganino:2018:OPO

[GLMS18] Giulio Ganino, Domenico Lembo, Massimo Mecella, and Federico Scafoglieri. Ontology population for open-source intelligence: a GATE-based solution. Software—Practice and Experience, 48(12):2302–2330, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garzarelli:2008:OSS

[GLT08] Giampaolo Garzarelli, Yasmina Reem Limam, and Bjørn Thomassen. Open source software and economic growth: a classical division of labor perspective. Information Technology for Development, 14(2):116–135, 2008. ISSN 0268-1102 (print), 1554-0170 (electronic).

Garlan:1984:GIP

[GM84] David B. Garlan and Philip L. Miller. GNOME: an introductory programming environment based on a family of structure editors. ACM SIGPLAN Notices, 19(5):65–72, May 1984. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Granlund:1994:DII

[GM94] Torbjörn Granlund and Peter L. Montgomery. Division by invariant integers using multiplication. ACM SIG-PLAN Notices, 29(6):61–72, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL ftp:

//ftp.cwi.nl/pub/pmontgom/divcnst.psa4.gz;   ftp:/
/ftp.cwi.nl/pub/pmontgom/divcnst.psl.gz;   http://
www.acm.org:80/pubs/citations/proceedings/pldi/178243/
p61-granlund/.

Gordon:2002:LHQ

[GM02] O. E. Gordon and T. E. Malloy. On-line Homework/ quiz/ exam applet: Freely available Java software for evaluating performance on line. Behavior Research Methods, Instruments, and Computers, 34(2):241–244, May 1, 2002. CODEN BRM-CEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Graham:2005:UUC

[GM05] Stuart J. H. Graham and David C. Mowery. The use of USPTO “continuation” applications in the patenting of software: Implications for free and open source*. Law & Policy, 27(1):128–151, 2005. ISSN 0265-8240 (print), 1467-9930 (electronic).

Ghiotto:2020:NMC

[GMBv20] G. Ghiotto, L. Murta, M. Barros, and A. van der Hoek. On the nature of merge conflicts: A study of 2,731 open source Java projects hosted by GitHub. IEEE Transactions on Software Engineering, 46(8):892–915, 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Geerts:2014:TAF

[GMPS14] Floris Geerts, Giansalvatore Mecca, Paolo Papotti, and Donatello Santoro. That’s all folks!: Llunatic goes open source. Proceedings of the VLDB Endowment, 7(13):1565–1568, August 2014. CODEN ???? ISSN 2150-8097.

Gupta:2017:KSI

[GNGS17] Gagan Gupta, Tony Nowatzki, Vinay Gangadhar, and Karthikeyan Sankaralingam. Kickstarting semiconductor innovation with open source hardware. Computer, 50(6):50–59, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-

0814 (electronic). URL https://www.computer.org/csdl/
mags/co/2017/06/mco2017060050-abs.html.

Graves:2009:SFJ

[GNR^{+}09] E. Graves, G. Nelson, M. Rodriguez, H. Zhou, and P. Keall. SU-FF-J-158: An open source software tool for treatment planning for small animal conformal radiotherapy. Medical Physics, 36(6Part8):2513, 2009. CODEN MPHYA6. ISSN 2473-4209.

Goerzen:1999:DGL

[GO99] John Goerzen and Ossama Othman. Debian GNU/Linux: Guide to Installation and Usage. New Riders Publishing, Carmel, IN, USA, 1999. ISBN 0-7357-0914-9. 158 pp. LCCN QA76.76.O63 .G634 1999. US$24.99. URL http://www.

newriders.com/books/title.cfm?isbn=0735709149.

Goelker:2007:GPI

[Goe07] Klaus Goelker. GIMP 2 for photographers: image editing with open source software. Rockynook, Santa Barbara, CA, USA, 2007. ISBN 1-933952-03-2. ix + 186 pp. LCCN TR267.5.G56.

Golubitsky:2006:SSF

[Gol06] Chaos Golubitsky. Simple software flow analysis using GNU cflow. ;login: the USENIX Association newsletter, 31(2): 37–41, April 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/publications/login/2006-
04/pdfs/golubitsky.pdf.

Gomulkiewicz:1999:HCU

[Gom99] Robert W. Gomulkiewicz. How copyleft uses license rights to succeed in the Open Source software revolution and the implications for Article 2B. World-Wide Web docu-

ment., 1999. URL http://cyber.law.harvard.edu/is99/

Copyleft.htm. 36 Houston Law Review 179.

Goode:2014:EOI

[Goo14] Sigi Goode. Exploring organizational information sharing in adopters and non-adopters of open source software: Evidence from six case studies. Knowledge and Process Management, 21(1):78–89, 2014. ISSN 1092-4604 (print), 1099-1441 (electronic).

Gordon:1996:IG

[Gor96] Ian Gordon. Introduction to gawk. Linux Journal, 25:??, May 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Gosling:1981:UE

[Gos81] James Gosling. UNIX EMACS. Report, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, December 1981.

Gosling:1983:ESE

[Gos83] James Gosling. EMACS screen editor: version 264. UniPress Software, Inc., Edison, NJ, USA, 1983. various pp. Seven computer disks.

Gosling:1984:EUM

[Gos84] James Gosling. Emacs user’s manual. Pyramid Technology Corporation, Mountain View, CA, USA, revision a edition, 1984. various pp.

Goth:2005:OSM

[Got05] Greg Goth. Open source meets venture capital. IEEE Distributed Systems Online, 6(6):??, June 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/csdl.computer.org/comp/mags/ds/2005/06/o6002.pdf.

Goth:2007:STO

[Got07] Greg Goth. Sprinting toward open source development. IEEE Software, 24(1):88–91, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gough:2004:IGG

[Gou04] Brian Gough. An introduction to GCC for the GNU compilers gcc and g++. Network Theory Ltd., Bristol, UK, 2004. ISBN 0-9541617-9-3. iv + 116 pp. Foreword by Richard M. Stallman.

Gruber:2005:EIW

[GP05] O. Gruber and J. Ponzo. Eclipse: Integrating Web technologies. IBM Systems Journal, 44(2):279–??, ???? 2005. CODEN IBMSA7. ISSN 0018-8670.

Gaff:2012:OSS

[GP12] Brian M. Gaff and Gregory J. Ploussios. Open source software. Computer, 45(6):9–11, June 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Guan:2016:OSF

[GPPT16] Fei Guan, Long Peng, Luc Perneel, and Martin Timmerman. Open source FreeRTOS as a case study in realtime operating system evolution. The Journal of Systems and Software, 118(??):19–35, August 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121216300383.

Gray:1999:SMR

[Gra99] Bob Gray. Software mini-review: Red Hat Linux 6.0. ;login: the USENIX Association newsletter, 24(4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://

www.usenix.org/publications/login/1999-8/features/
redhat.html.

Graham:2001:RSO

[Gra01] Susan L. Graham. From research software to open source. Lecture Notes in Computer Science, 2000:195–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2000/20000195.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2000/20000195.pdf.

Greenberg:1980:PCM

[Gre80] B. S. Greenberg. Prose and CONS — Multics Emacs: a commercial text-processing system in LISP. In Conference Record of the 1980 LISP Conference, Stanford University. ACM Press, New York, NY 10036, USA, 1980.

Gregorio:2011:ITL

[Gre11a] Enrico Gregorio. Installing TEX Live 2010 on Ubuntu. TUGboat, 32(1):56–61, 2011. ISSN 0896-3207.

Gregorio:TB32-1-56

[Gre11b] Enrico Gregorio. Installing TEX Live 2010 on Ubuntu. TUGboat, 32(1):56–61, 2011. ISSN 0896-3207. URL https://tug.

org/TUGboat/tb32-1/tb100gregorio.pdf.

Grein:2014:PUG

[Gre14] C. K. W. Grein. Physical units with GNAT. Ada User Journal, 35(1):42–??, March 2014. CODEN AUJOET. ISSN 1381-6551.

Greenstein:2018:FSF

[Gre18] Shane Greenstein. Free software without a free lunch or free beer. IEEE Micro, 38(5):94–96, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.computer.org/csdl/mags/mi/
2018/05/mmi2018050094.html.

Griffith:2002:GCR

[Gri02] Arthur Griffith. GCC, the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 2002. ISBN 0-07-222405-3. xxiii + 647 pp. LCCN QA76.76.C65 G75 2002. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/bios/mh042/2003268319.html;
 http://www.loc.gov/catdir/description/mh031/2003268319.
html;   http://www.loc.gov/catdir/toc/mh031/2003268319.
html.

Grier:2016:FYO

[Gri16] David Alan Grier. Fifteen years to open source. Computer, 49(12):96, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/12/mco2016120096.html.

Goodrum:2001:OSA

[GRJS01] Abby A. Goodrum, Mark E. Rorvig, Ki-Tai Jeong, and Chitturi Suresh. An open source agenda for research linking text and image content features. Journal of the American Society for Information Science and Technology: JASIST, 52 (11):948–953, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gropp:2001:SCP

[Gro01] Bill Gropp. Solving CFD problems with open source parallel libraries. Lecture Notes in Computer Science, 1947:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1947/19470052.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1947/19470052.pdf.

Grinzo:2000:PBG

[GS00] Lou Grinzo and Jacques Surveyer. Programmer’s bookshelf: GNU tools and process patterns. Dr. Dobb’s Journal of Software Tools, 25(12):159–160, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Garcia:2002:ERI

[GS02] Rodrigo García García and Alfred Strohmeier. Experiences report on the implementation of EPTs for GNAT. ACM SIGADA Ada Letters, 22(4):22–27, December 2002. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Giaglis:2012:DEP

[GS12] G. M. Giaglis and D. Spinellis. Division of effort, productivity, quality, and relationships in FLOSS virtual teams: Evidence from the FreeBSD project. J.UCS: Journal of Universal Computer Science, 18(19):2625–??, ???? 2012. CODEN ????

ISSN 0948-6968. URL http://www.jucs.org/jucs_18_19/
division_of_effort_productivity.

Granlund:2004:GMG

[GSR^{+}04] Torbjörn Granlund, Gunnar Sjödin, Hans Riesel, Richard Stallman, Brian Beuning, Doug Lea, John Amanatides, Paul Zimmermann, Ken Weber, Per Bothner, Joachim Hollman, Bennet Yee, Andreas Schwab, Robert Harley, David Seal, Robert Harley, Torsten Ekedahl, Paul Zimmermann, Linus Nordberg, Kent Boortz, Kevin Ryde, Steve Root, Gerardo Ballabio, and Hans Thorsen. GNU MP: The GNU Multiple Precision Arithmetic Library. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, version 4.1.4 edition, September 21, 2004.

iv + 127 pp. URL ftp://ftp.gnu.org/gnu/gmp/gmp-4.1.
4.tar.gz;   http://www.swox.se/gmp/. GNU MP devel-

opment began in 1991. Earler versions are 1.0 (8-Aug-1991), 2.0 (24-Apr-1996), 3.0 (17-Apr-2000), and 4.0 (1-Dec-2001).

Gotel:2008:TSQ

[GSW08] Olly Gotel, Christelle Scharff, and Andrew Wildenberg. Teaching software quality assurance by encouraging student contributions to an open source Web-based system for the assessment of programming assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):214–218, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Guidolin:2023:HRO

[GTMR23] Mattia Guidolin, Luca Tagliapietra, Emanuele Menegatti, and Monica Reggiani. Hi-ROS: Open-source multi-camera sensor fusion for real-time people tracking. Computer Vision and Image Understanding: CVIU, 232(??):??, July 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1077314223000747.

Guinan:2000:CPD

[Gui00] Jamie Guinan. Cross-platform development using GCC. C/C++ Users Journal, 18(3):18–??, March 2000. CODEN CCUJEX. ISSN 1075-2838.

Gunther:2002:LGP

[Gün02] Karsten Günther. Linux ge-packt: [die wichtigsten Befehle für GNU/Linux-Systeme; Textwerkzeuge: Drucken, Konvertieren, Suchen und Anpassen; CDs erstellen unter Linux]. MITP Verlag, Bonn, Germany, 2002. ISBN 3-8266-0946-8. 560 (est.) pp. LCCN ???? EUR 14.95 (DE), EUR 15.40 (AT).

Gupta:2003:AAI

[Gup03] Sanjiv K. Gupta. Alias analysis for intermediate code. In Hutton et al. [HDR03], pages 71–78. ISBN ????

LCCN ???? URL http://www.linux.org.uk/~ajh/gcc/
gccsummit-2003-proceedings.pdf.

Gustavsson:2020:MOS

[Gus20] T. Gustavsson. Managing the open source dependency. Computer, 53(2):83–87, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gutmann:2000:OSC

[Gut00] Peter Gutmann. An open-source cryptographic coprocessor. In USENIX [USE00b], page ?? ISBN 1-880446-18-9.

LCCN ???? URL http://www.usenix.org/publications/
library/proceedings/sec2000/gutmann.html.

Guyot:2000:GBL

[Guy00] Paul Guyot. Are GPL and BSD License really compatible? World-Wide Web document., 2000. URL http://www.

kallisys.org/bsd-lite/bsd-gpl/?lg=en.

Green:2014:OST

[Gv14] Kevin R. Green and Lennaert van Veen. Open-source tools for dynamical analysis of Liley’s mean-field cortex model. Journal of Computational Science, 5(3):507–516, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S187775031300077X.

Greenwald:2016:PAA

[GV16] Chelsea Greenwald and Niranjan Venugopal. Poster-05: Automated analysis of MR distortion using a novel anthropomorphic phantom and open source software. Medical Physics, 43(8Part1):4936, 2016. CODEN MPHYA6. ISSN 2473-4209.

Gubinelli:2014:GTM

[GvdHPR14] Massimiliano Gubinelli, Joris van der Hoeven, François Poulain, and Denis Raux. GNU TEXMACS: towards a scientific office suite. In Hong and Yap [HY14], pages 562– 569. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (ebook). LCCN QA76.9.M35.

Granvik:2009:OOS

[GVOM09] Mikael Granvik, Jenni Virtanen, Dagmara Oszkiewicz, and Karri Muinonen. OpenOrb: Open-source asteroid orbit computation software including statistical ranging. Meteoritics & Planetary Science, 44(12):1853–1861, 2009. ISSN 1086-9379 (print), 1945-5100 (electronic).

Grodzinsky:2009:EIF

[GW09] Frances S. Grodzinsky and Marty J. Wolf. Ethical Interest in Free and Open Source Software, chapter 10, pages 245–271. Wiley, New York, NY, USA, 2009. ISBN 0-470-28181-2.

Gwebu:2010:SEE

[GW10] Kholekile L. Gwebu and Jing Wang. Seeing eye to eye? An exploratory study of free open source software users’ perceptions. The Journal of Systems and Software, 83(11):2287– 2296, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gohel:2001:LGK

[GWT^{+}01] Himanshu Gohel, Thomas Weidenfeller, Andrew D. Todd, Ze’ev Atlas, and Dave Paris. Letters: GNOME, KDE clarification; parsing XML; CPRM frenzy; deadlock detecting; Perl versus the world. Dr. Dobb’s Journal of Software Tools, 26(5):10, 12, May 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Guo:2023:TMS

[GYW^{+}23] Pengfei Guo, Yingjian Yan, Junjie Wang, Jingxin Zhong, Yanjiang Liu, and Jinsong Xu. Towards a metrics suite for evaluating cache side-channel vulnerability: Case studies on an opensource RISC-V processor. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003905.

Hackvan:1998:QOS

[Hac98] Stig Hackvän. Not quite Open Source, but close. Linux Today,

December 10, 1998. URL http://linuxtoday.com/news_
story.php3?ltsn=1998-12-10-001-05-NW-CY.

Hanasaki:2023:IRT

[HAC^{+}23] Kota Hanasaki, Zulfikhar A. Ali, Min Choi, Mauro Del Ben, and Bryan M. Wong. Implementation of real-time TDDFT for periodic systems in the open-source PySCF software package. Journal of Computational Chemistry, 44(9):980–987, April 5, 2023. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Haer:2002:LTS

[Hae02] Gregory Haer. The Linux4.TV set-top box open source project. Embedded Linux Journal, 7:12–15, January/ February 2002. CODEN ???? ISSN 1534-083X. URL

http://embedded.linuxjournal.com/magazine/issue07/
;   http://www.linuxdevices.com/articles/AT7194247162.
htm.

Hafner:2001:FOS

[Haf01] Ullrich Hafner. FIASCO — an open-source fractal image and sequence codec. Linux Journal, 81:152–155, January 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Hagog:2004:SMS

[Hag04] Mostafa Hagog. Swing modulo scheduling for GCC. In Hutton et al. [HDR04], pages 55–64. ISBN ???? LCCN

???? URL http://people.redhat.com/lockhart/.gcc04/
MasterGCC-2side.pdf.

Hahn:1994:UU

[Hah94] Harley Hahn. UNIX unbound. Osborne/McGraw-Hill, Berkeley, CA, USA, 1994. ISBN 0-07-882050-2 (paperback). xxxi + 792 pp. LCCN QA76.76.O63 H343 1994.

Hall:2002:FSB

[Hal02] Jon Hall. Free software in Brazil. Linux Journal, 101:??, September 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://www.linuxjournal.

com/article.php?sid=6125.

Hammerslag:1988:TM

[Ham88] D. Hammerslag. Treemacs manual. Technical Report UIUCDCS-R-88-1427, University of Illinois at Urbana-Champaign, Urbana-Champaign, IL, USA, May 1988. 25 pp.

Hammerslag:1990:TET

[Ham90] D. H. Hammerslag. Treemacs: an extensible tree editor. Technical Report UIUCDCS-R-90-1600, University of Illinois at Urbana-Champaign, Urbana-Champaign, IL, USA, May 1990. vi + 150 pp.

Hammel:1999:PWG

[Ham99] Michael J. Hammel. Programming Web graphics with perl and GNU software. Linux Journal, 64:??, August 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-

3827 (electronic). URL http://noframes.linuxjournal.
com/lj-issues/issue64/3464.html.

Hammel:2007:AGG

[Ham07] Michael J. Hammel. The artist’s guide to GIMP effects: creative techniques for photographers, artists, and designers. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-121-2 (paperback). xii + 348 pp. LCCN T385 .H329558 2007.

URL http://www.loc.gov/catdir/enhancements/fy0728/
2007001652-d.html;   http://www.loc.gov/catdir/toc/
ecip078/2007001652.html.

Hancock:2000:WHSa

[Han00] Bill Hancock. White House supporting open source code. Computers & Security, 19(7):577–578, November 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800070103.

Harvard:1977:HTM

[Har77] Harvard University and Harvard-Radcliffe Student Timesharing System, Cambridge, MA, USA. HRSTS Teco manual, second edition, 1977. 61 pp.

Harlander:1994:CSA

[Har94] Magnus Harlander. Central system administration in a heterogeneous Unix environment: GeNUAdmin. In USENIX Association [USE94], pages 1–8. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

Harrison:1999:MOS

[Har99] Mark Harrison. Managing open source software. ;login: the USENIX Association newsletter, 24(1s):??, January 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/1999-1/opensource.html.

Special issue on tools.

Harford:2000:GER

[Har00] Alex Harford. GIMP essential reference. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 0-7357-0911-4. xxi + 373 pp. LCCN T385 .H335 2000. US$24.95.

Hardaway:2005:SRC

[Har05] Donald E. Hardaway. Sharing research in the 21st Century: borrowing a page from open source software. Communications of the ACM, 48(8):125–128, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harutyunyan:2020:MYO

[Har20] N. Harutyunyan. Managing your open source supply chain — why and how? Computer, 53(6):77–81, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hassler:2005:OSL

[Has05] Vesna Hassler. Open Source libraries for information retrieval. IEEE Software, 22(5):78–82, September/October 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hausmann:2001:KEO

[Hau01] Simon Hausmann. Konqueror/Embedded: An open-source Web browser for embedded Linux systems. Embedded Linux Journal, 5:32–35, September/October 2001. CODEN ????

ISSN 1534-083X. URL http://embedded.linuxjournal.
com/magazine/issue05/;   http://www.linuxdevices.com/
articles/AT5068415887.html.

Haywood:2005:ROS

[Hay05] Stephen Haywood. Reviews: Open source solutions for small business problems. Linux Journal, 2005(134):15, June 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Hintzen:2012:VOS

[HBB^{+}12] Niels T. Hintzen, Francois Bastardie, Doug Beare, Gerjan J. Piet, Clara Ulrich, Nicolas Deporte, Josefine Egekvist, and Henrik Degel. VMStools: Open-source software for the processing, analysis and visualisation of fisheries logbook and VMS data. Fisheries Research, 115–116(??):31–43, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611003365.

Hanson:2005:IWA

[HBC^{+}05] V. L. Hanson, J. P. Brezin, S. Crayne, S. Keates, R. Kjeldsen, J. T. Richards, C. Swart, and S. Trewin. Improving Web accessibility through an enhanced open-source browser. IBM Systems Journal, 44(3):573–588, ???? 2005. CODEN

IBMSA7. ISSN 0018-8670. URL http://www.research.ibm.
com/journal/sj/443/hanson.html.

Hukal:2019:BCW

[HBGS19] P. Hukal, N. Berente, M. Germonprez, and A. Schecter. Bots coordinating work in open source software projects. Computer, 52(9):52–60, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harutyunyan:2019:IRF

[HBR19] Nikolay Harutyunyan, Andreas Bauer, and Dirk Riehle. Industry requirements for FLOSS governance tools to facilitate the use of open source software in commercial products. The Journal of Systems and Software, 158(??):Article 110390, December 2019. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121219301578.

Hornik:2009:OSM

[HBZ09] Kurt Hornik, Christian Buchta, and Achim Zeileis. Opensource machine learning: R meets Weka. Computational Statistics, 24(2):225–232, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://link.

springer.com/article/10.1007/s00180-008-0119-7.

Hou:2007:AIU

[HC07] Ting-Wei Hou and Fuh-Gwo Chen. An anomaly in an interpreter using GCC source-code-level register allocation. ACM SIGPLAN Notices, 42(4):9–13, April 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hasselbring:2020:OSR

[HCH^{+}20] W. Hasselbring, L. Carr, S. Hettrick, H. Packer, and T. Tiropanis. Open source research software. Computer, 53 (8):84–88, 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hutton:2003:PGD

[HDR03] Andrew J. Hutton, Stephanie Donovan, and C. Craig Ross, editors. Proceedings of the GCC Developers Summit May 25– 27, 2003, Ottawa, Ontario Canada. ????, ????, 2003. ISBN

???? LCCN ???? URL http://www.linux.org.uk/~ajh/
gcc/gccsummit-2003-proceedings.pdf.

Hutton:2004:PGD

[HDR04] Andrew J. Hutton, Stephanie Donovan, and C. Craig Ross, editors. Proceedings of the GCC Developers Summit, June 2nd–4th, 2004, Ottawa, Ontario, Canada. ????, ????, 2004.

ISBN ???? LCCN ???? URL http://people.redhat.com/
lockhart/.gcc04/MasterGCC-2side.pdf.

He:1995:FPI

[He95] Lei He. Floating point implementation for Motorola HC6811C: GNU C cross-compiler. Master’s thesis, Ryerson Polytechnic University, Toronto, Ontario, Canada, 1995.

Hislop:2017:HOS

[HE17] Gregory W. Hislop and Heidi J. C. Ellis. Humanitarian open source software in computing education. Computer, 50(10):98–101, October 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/csdl/mags/co/2017/10/
mco2017100098-abs.html.

Hearn:2009:RFS

[Hea09] Anthony C. Hearn. REDUCE is free software as of January 2009. ACM Communications in Computer Algebra, 43(1–2):

15–16, March/June 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hecker:1999:SSB

[Hec99] Frank Hecker. Setting up shop: The business of open-source software. IEEE Software, 16(1):45–51, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-

7459 (electronic). URL http://computer.org/software/
so1999/s1045abs.htm;   http://dlib.computer.org/so/
books/so1999/pdf/s1045.pdf.

Heffan:1997:CLC

[Hef97] Ira V. Heffan. Copyleft: Licensing collaborative works in the digital age. Stanford Law Review, 49(6):1487–1522, July 1997.

URL http://open-bar.org/docs/copyleft.pdf.

Heimlich:2016:GOI

[Hei16] O. Heimlich. GNU Octave interval package. Web document,

2016. URL http://octave.sourceforge.net/interval/.

Hennessey:1992:WDE

[Hen92] W. Hennessey. WCL: delivering efficient Common Lisp applications under UNIX. In ACM [ACM92], pages 260–269. ISBN 0-89791-483-X. LCCN QA76.73.L23A26 1992. ACM order no. 552920.

Hertzog:2004:DHT

[Her04] Raphaël Hertzog. Debian. Eyrolles, Paris, France, 2004. ISBN 2-212-11398-6. xi + 246 pp. LCCN ????

Herrera:2020:PCO

[Her20] A. Herrera. The promises and challenges of open source hardware. Computer, 53(10):101–104, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hislop:2009:UOS

[HETD09] Gregory W. Hislop, Heidi J. C. Ellis, Allen B. Tucker, and Scott Dexter. Using open source software to engage students in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1): 134–135, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Hawe:2012:KAD

[HFO^{+}12] David Hawe, Francisco R. Hernández Fernández, Liam O’Suilleabháin, Jian Huang, Eric Wolsztynski, and Finbarr O’Sullivan. Kinetic analysis of dynamic positron emission tomography data using open-source image processing and statistical inference tools. WIREs Computational Statistics, 4 (3):316–322, May/June 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Halme:1988:GED

[HH88] Heikki Halme and Juha Heinänen. GNU Emacs as a dynamically extensible programming environment. Software—Practice and Experience, 18(10):999–1009, October 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanwell:2021:OCJ

[HHG^{+}21] Marcus D. Hanwell, Chris Harris, Alessandro Genova, Mojtaba Haghighatlari, Muammar El Khatib, Patrick Avery, Johannes Hachmann, and Wibe Albert de Jong. Open Chemistry, JupyterLab, REST, and quantum chemistry. International Journal of Quantum Chemistry, 121(1):e26472:1– e26472:??, January 5, 2021. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hill:2005:DGL

[HHV05] Benjamin Mako Hill, David B. Harris, and Jaldhar Vyas. Debian GNU/Linux 3.1 bible. Wiley, New York, NY, USA, 2005. ISBN 0-7645-7644-5. xxvi + 640 pp. LCCN QA76.76.O63 H57135 2005. URL http://www.loc.gov/

catdir/toc/ecip054/2004027963.html.

Hichert:2004:OFS

[Hic04] Jan Hichert. Opinion: The financial sector’s move to open source security products. Network Security, 2004(3):19–20, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print),

1872-9371 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S135348580400056X.

Higham:1993:HWM

[Hig93] Nicholas J. Higham. Handbook of Writing for the Mathematical Sciences. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, June 1993. ISBN 0-89871-314-5. xii + 241 pp. LCCN QA42.H54 1993. US$21.50.

Hinckley:1987:OOE

[Hin87] K. Hinckley. An object oriented extension language for integrating disparate applications. In Bullinger et al. [BSK87], pages 529–533. ISBN 0-444-70304-7. LCCN QA76.9.S88I325 1987.

Hoepman:2007:IST

[HJ07] Jaap-Henk Hoepman and Bart Jacobs. Increased security through open source. Communications of the ACM, 50(1): 79–83, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hatakeyama:1995:IEJ

[HK95] T. Hatakeyama and H. Kakuda. The implementation and evaluation of a Japanese editor utilizing pronunciationinformation. Transactions of the Information Processing Society of Japan, 36(1):119–128, January 1995. CODEN JSGRD5. ISSN 0387-5806.

Haar:2003:JAO

[HK03] Tobias Haar and Gabriele Keck. Juristische Aspekte von Open-Source-Software: Offene Konkurrenz. (German) [Legal aspects of Open-Source software: Open competition]. World-Wide Web document., September 2003. URL http://www.

heise.de/ix/artikel/2003/09/079/.

Hafer:2009:AOS

[HK09] Lou Hafer and Arthur E. Kirkpatrick. Assessing open source software as a scholarly contribution. Communications of the ACM, 52(12):126–129, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hadjilambrou:2019:CCO

[HKA^{+}19] Zacharias Hadjilambrou, Marios Kleanthous, Georgia Antoniou, Antoni Portero, and Yiannakis Sazeides. Comprehensive characterization of an open source document search engine. ACM Transactions on Architecture and Code Optimization, 16(2):19:1–19:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hansen:2002:OSA

[HKP02] Marit Hansen, Kristian Köhntopp, and Andreas Pfitzmann. The open source approach — opportunities and limitations with respect to security and privacy. Computers & Security, 21(5):461–471, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802005163.

Hornikx:2016:OOS

[HKvH16] Maarten Hornikx, Thomas Krijnen, and Louis van Harten. openPSTD: the open source pseudospectral time-domain method for acoustic propagation. Computer Physics Communications, 203(??):298–308, June 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465516300443.

Hoshi:2021:OSL

[HKY^{+}21] Takeo Hoshi, Mitsuaki Kawamura, Kazuyoshi Yoshimi, Yuichi Motoyama, Takahiro Misawa, Youhei Yamaji, Synge Todo, Naoki Kawashima, and Tomohiro Sogabe. Kω — opensource library for the shifted Krylov subspace method of the form (zI − H)x = b. Computer Physics Communications, 258(??):Article 107536, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302551.

Hightower:2002:JTE

[HL02] Richard Hightower and Nicholas Lesiecki. Java tools for extreme programming: mastering open source tools including Ant, JUnit, and Cactus. Java open source library. Wiley, New York, NY, USA, 2002. ISBN 0-471-20708-X. xxvii + 516 pp. LCCN QA76.73.J38 H54 2002.

Harrold:1995:ASD

[HLL^{+}95] M. J. Harrold, L. Larsen, J. Lloyd, D. Nedved, M. Page, G. Rothermel, M. Singh, and M. Smith. Aristotle: a system for development of program analysis based tools. In ACM [ACM95], pages 110–119. ISBN 0-89791-747-2. LCCN ????

He:2013:PDO

[HLS^{+}13a] Yuye He, Chin Yee Liew, Nitin Sharma, Sze Kwang Woo, Yi Ting Chau, and Chun Wei Yap. PaDEL-DDPredictor: Open-source software for PD-PK-T prediction. Journal of Computational Chemistry, 34(7):604–610, 2013. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

He:2013:SNU

[HLS^{+}13b] Yuye He, Chin Yee Liew, Nitin Sharma, Sze Kwang Woo, Yi Ting Chau, and Chun Wei Yap. Software news and updates: PaDEL-DDPredictor: Open-source software for PD-PK-T prediction. Journal of Computational Chemistry, 34 (7):604–610, March 15, 2013. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Harrison:1989:IBP

[HM89] Michael A. Harrison and Ethan V. Munson. On integrated bibliography processing. Electronic Publishing—Origination, Dissemination, and Design, 2(4):193–209, December 1989. CODEN EPODEU. ISSN 0894-3982.

Huynh:2010:EIO

[HM10] Toan Huynh and James Miller. An empirical investigation into open source web applications’ implementation vulnerabilities. Empirical Software Engineering, 15(5):556–576, October 2010. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10664-010-9131-y.

Hjelsvold:2019:EEG

[HM19] Rune Hjelsvold and Deepti Mishra. Exploring and expanding GSE education with open source software development. ACM Transactions on Computing Education, 19(2): 12:1–12:??, February 2019. CODEN ???? ISSN 1946-6226.

URL https://dl.acm.org/ft_gateway.cfm?id=3230012.

Hoefling:2012:IOS

[HMKC12] Michael Hoefling, Michael Menth, Christian Kniep, and Marcus Camen. IBPM: An open-source-based framework for InfiniBand performance monitoring. Lecture Notes in Computer Science, 7201:236–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28540-0_17/.

Hallen:2018:OOS

[HMO^{+}18] Mark A. Hallen, Jeffrey W. Martin, Adegoke Ojewole, Jonathan D. Jou, Anna U. Lowegard, Marcel S. Frenkel, Pablo Gainza, Hunter M. Nisonoff, Aditya Mukund, Siyu Wang, Graham T. Holt, David Zhou, Elizabeth Dowd, and Bruce R. Donald. OSPREY 3.0: Open-source protein redesign for you, with powerful new features. Journal of Computational Chemistry, 39(30):2494–2507, November 15, 2018. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Helfer:2015:IOS

[HMP^{+}15] Thomas Helfer, Bruno Michel, Jean-Michel Proix, Maxime Salvo, Jérôme Sercombe, and Michel Casella. Introducing the open-source mfront code generator: Application to mechanical behaviours and material knowledge management within the PLEIADES fuel element modelling platform. Computers and Mathematics with Applications, 70(5):994–1023, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122115003132.

Harrold:1993:ECP

[HMR93] Mary Jean Harrold, Brian Malloy, and Gregg Rothermel. Efficient construction of program dependence graphs. ACM SIG-SOFT Software Engineering Notes, 18(3):160–170, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hoang:2015:SG

[HMW15] Duc Hoang, Yannick Moy, and Angela Wallenburg. SPARK 2014 and GNATprove. International Journal on Software Tools for Technology Transfer (STTT), 17(6):695–707, November 2015. CODEN ???? ISSN 1433-2779 (print),

1433-2787 (electronic). URL http://link.springer.com/
article/10.1007/s10009-014-0322-5.

Huo:2021:JJB

[HMX21a] Zenan Huo, Gang Mei, and Nengxiong Xu. juSFEM: a Juliabased open-source package of parallel Smoothed Finite Element Method (S-FEM) for elastic problems. Computers and Mathematics with Applications, 81(??):459–477, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122120300523.

Huo:2021:PJJ

[HMX21b] Zenan Huo, Gang Mei, and Nengxiong Xu. juSFEM: a Juliabased open-source package of parallel Smoothed Finite Element Method (S-FEM) for elastic problems. Computers and Mathematics with Applications, 81:459–477, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL https://www.sciencedirect.com/science/article/

pii/S0898122120300523. Development and Application of Open-source Software for Problems with Numerical PDEs.

Hanada:2022:STR

[HMYH22] Takashi Hanada, Yuichi Motoyama, Kazuyoshi Yoshimi, and Takeo Hoshi. sim-trhepd-rheed — open-source simulator of total-reflection high-energy positron diffraction (TRHEPD) and reflection high-energy electron diffraction (RHEED). Computer Physics Communications, 277(??):Article 108371, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552200090X.

Hertel:2003:MSD

[HNH03] Guido Hertel, Sven Niedner, and Stefanie Herrmann. Motivation of software developers in Open Source projects: an Internet-based survey of contributors to the Linux kernel. Research Policy, 32(7):1159–1177, 2003. CODEN ???? ISSN 0048-7333 (print), 1873-7625 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0048733303000477.

Handa:1993:MME

[HNT93] K. Handa, M. Nishikimi, and S. Tomura. Mule: A MULtilingual enhancement to GNU emacs. In Leiner [Lei93a], pages GAB 1–9.

Ho:1995:FPI

[Ho95] Lei Ho. Floating point implementation for Motorola HC6811C: GNU C cross-compiler. Master’s thesis, Department of Electrical Engineering, Ryerson Polytechnic University, Toronto, Ontario, Canada, 1995.

Hohndel:2001:RDT

[Hoh01] Dirk Hohndel. Reviving dead technology: How XFree86 and open source prevailed against the X Consortium, 2001. URL

http://www.linuxshowcase.org/tech.html. Unpublished

invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Hollenback:2005:GMY

[Hol05] Phil Hollenback. GNU motion: Your eye in the sky for computer room surveillance. Linux Journal, 2005(131):??, March 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Hastings:2007:IOS

[HOL^{+}07] Shannon Hastings, Scott Oster, Stephen Langella, David Ervin, Tahsin Kurc, and Joel Saltz. Introduce: An open source toolkit for rapid development of strongly typed Grid services. Journal of Grid Computing, 5(4):407–427, December 2007. CODEN ???? ISSN 1570-7873 (print), 1572-

9184 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1570-7873&volume=5&
issue=4&spage=407.

Hollander:2015:BNO

[Hol15] Barbara Gottfried Hollander. Bitcoins: navigating open source currency. Digital and information literacy. Rosen Publishing, New York, NY, USA, 2015. ISBN 1-4777-7930-2 (library bound), 1-4777-7931-0 (paperback), 1-4777-7932-9 (6pack). 48 pp. LCCN HG221.5 .H65 2015.

Holbrook:2023:CAT

[Hol23] Luke D. Holbrook. Clarity about transparency: Software transparency and open source software licenses in the U.S. public sector. IEEE Software, 40(4):71–76, 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hombourger:2000:SNC

[Hom00] Cédric Hombourger. Solution de network computing sous GNU/Linux. Mém. d.e.s.s.: Techniques nouvelles en informatique [t.n.i.], Université Montpellier II Sciences et Techniques du Languedoc, Montpellier 2, France, 2000. 59 pp.

Hoppner:2004:GPA

[Höp04] Julian P. Höppner. The GPL prevails: An analysis of the first-ever Court decision on the validity and effectivity of the GPL. SCRIPT, 1(4):??, 2004. URL http://www.law.ed.ac.

uk/ahrb/script-ed/issue4/GPL-case.asp.

Harrison:2005:SAO

[HOST05] W. Harrison, H. Ossher, S. Sutton, and P. Tarr. Supporting aspect-oriented software development with the Concern Manipulation Environment. IBM Systems Journal, 44 (2):309–318, ???? 2005. CODEN IBMSA7. ISSN 0018-

8670. URL http://www.research.ibm.com/journal/sj/
442/harrison.pdf.

Howes:1998:TPC

[How98] Brad Howes. Template processing classes for Python. Dr. Dobb’s Journal of Software Tools, 23(2):38, 40, 42, 44–46, 48, 100, February 1998. CODEN DDJOEB. ISSN 1044-789X.

Hepting:2008:CSB

[HPM^{+}08] Daryl H. Hepting, Lijuan Peng, Timothy J. Maciag, David Gerhard, and Brien Maguire. Creating synergy between usability courses and open source software projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):120–123, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

https://www.math.utah.edu/pub/mirrors/ftp.ira.uka.
de/bibliography/Misc/DBLP/2008.bib.

Hermann:2017:OSF

[HPT17] Gunter Hermann, Vincent Pohl, and Jean Christophe Tremblay. An open-source framework for analyzing N-electron dynamics. II. Hybrid density functional theory/configuration interaction methodology. Journal of Computational Chemistry, 38(28):2378–2387, October 30, 2017. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hatton:1994:HAS

[HR94] L. Hatton and A. Roberts. How accurate is scientific software? IEEE Transactions on Software Engineering, 20(10):785–797, October 1994. CODEN IESEDJ. ISSN 0098-5589 (print),

1939-3520 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=328993.

Held:2011:PCO

[HR11] Isaac Held and David Randall. Point/counterpoint: Open source climate model development is worth it /should climate models be open source? IEEE Software, 28(6):62–65, November/December 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hathhorn:2019:DCO

[HR19] C. Hathhorn and G. Rosu. Dealing with C’s original sin. IEEE Software, 36(5):24–28, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hofheinz:2015:GNU

[HS15] Dennis Hofheinz and Victor Shoup. GNUC: A new universal composability framework. Journal of Cryptology: the journal of the International Association for Cryptologic Research, 28(3):423–508, July 2015. CODEN JOCREQ. ISSN 0933-2790 (print), 1432-1378 (electronic). URL http://link.

springer.com/article/10.1007/s00145-013-9160-y.

Haney:1989:CBM

[HSC89] Daniel R. Haney, Richard Stallman, and Robert Corbett. CUG285 — BISON for MS-DOS. C Users Journal, 7(7):119– ??, July 1989. ISSN 0898-9788.

Horgue:2015:OST

[HSF^{+}15] P. Horgue, C. Soulaine, J. Franc, R. Guibert, and G. Debenest. An open-source toolbox for multiphase flow in porous media. Computer Physics Communications, 187(??):217–226, February 2015. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465514003403.

Huang:2018:ISA

[HSX^{+}18] Qiao Huang, Emad Shihab, Xin Xia, David Lo, and Shanping Li. Identifying self-admitted technical debt in open source projects using text mining. Empirical Software Engineering, 23(1):418–451, February 2018. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://link.

springer.com/article/10.1007/s10664-017-9522-4.

Hannig:2021:OSH

[HT21] Frank Hannig and Jurgen Teich. Open source hardware. Computer, 54(10):111–115, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hagan:1996:UCS

[HTU96] Kerry Hagan, Chris Alan Thyberg, and Carnegie Mellon University. UNIX: computing skills workshop, 1995–1996. College custom series. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-064681-3. 146 pp. LCCN ????

Huang:2017:IQV

[Hua17] Li Huang. iQIST v0.7: an open source continuous-time quantum Monte Carlo impurity solver toolkit. Computer Physics Communications, 221(??):423–424, December 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465517302898.

Huang:2023:AOS

[Hua23] Li Huang. ACFlow: an open source toolkit for analytic continuation of quantum Monte Carlo data. Computer Physics Communications, 292(??):Article 108863, November 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002084.

Hubicka:2003:PGA

[Hub03] Jan Hubička. Porting GCC to the AMD64 architecture. In Hutton et al. [HDR03], pages 79–105. ISBN ????

LCCN ???? URL http://www.linux.org.uk/~ajh/gcc/
gccsummit-2003-proceedings.pdf.

Hubbard:2004:OSC

[Hub04a] Jordan Hubbard. Open source to the core. ACM Queue: Tomorrow’s Computing Today, 2(3):24–31, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hubicka:2004:GCG

[Hub04b] Jan Hubička. The GCC call graph module: a framework for inter-procedural optimization. In Hutton et al. [HDR04], pages 65–77. ISBN ???? LCCN ???? URL http://people.

redhat.com/lockhart/.gcc04/MasterGCC-2side.pdf.

Hughes:1995:FFS

[Hug95] Phil Hughes. Freenets and free software. Dr. Dobb’s Developer Update, 2(2):2–3, February 1995. CODEN ???? ISSN 1079-8595.

Hunger:2001:DGL

[Hun01] Steven Hunger. Debian GNU/Linux bible. I D G Books Worldwide, Indianapolis, IN, USA, 2001. ISBN 0-7645-4710-0. ???? pp. LCCN QA76.76.O63 H847 2001.

Huppelshauser:2001:GGP

[Hüp01] Nicolas Hüppelshäuser. Die GNU General Public License und deutsches Recht. (German) [The GNU General Public License and German law]. World-Wide Web document., Febru-

ary 1, 2001. URL http://www.ifross.de/Fremdartikel/
seminararbeit.pdf.

Hoadley:2017:UOS

[HW17a] Kenneth D. Hoadley and Mark E. Warner. Use of open source hardware and software platforms to quantify spectrally dependent differences in photochemical efficiency and functional absorption cross section within the dinoflagellate Symbiodinium spp. Frontiers in Marine Science, 4, November 2017. ISSN 2296-7745.

Hocquet:2017:OIW

[HW17b] Alexandre Hocquet and Frédéric Wieber. “Only the initiates will have the secrets revealed”: Computational chemists and the openness of scientific software. IEEE Annals of the History of Computing, 39(4):40–58, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:/

/ieeexplore.ieee.org/document/8268025/.

He:2023:HHO

[HWL^{+}23] Xiaofeng He, Kun Wang, Tiegang Liu, Yiwei Feng, Bin Zhang, Weixiong Yuan, and Xiaojun Wang. HODG: highorder discontinuous Galerkin methods for solving compressible Euler and Navier–Stokes equations — an open-source component-based development framework. Computer Physics Communications, 286(??):Article 108660, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300005X.

Huang:2015:IOS

[HWM^{+}15] Li Huang, Yilin Wang, Zi Yang Meng, Liang Du, Philipp Werner, and Xi Dai. iQIST: an open source continuoustime quantum Monte Carlo impurity solver toolkit. Computer Physics Communications, 195(??):140–160, October 2015. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465515001629.

Hagerty:19xx:GSM

[HWZxx] Daniel Hagerty, Melissa Weisshaus, and Eli Zaretski. GNU Software for MS-Windows and MS-DOS. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 19xx. ISBN 1-882114-57-4. ???? pp. LCCN ???? Includes CD-ROM.

Hagerty:2001:GSM

[HWZ01] Daniel Hagerty, Melissa Weisshaus, and Eli Zaretski. GNU Software for MS-Windows and MS-DOS and Compatible Systems. GNU Press, Boston, MA, USA, second edition, 2001. ISBN 1-882114-58-2. 150 (est.) pp. LCCN ???? US$35.00.

URL http://www.gnupress.org/windows.html. Includes

CD-ROM.

Hasbestan:2020:POS

[HXS20] Jaber J. Hasbestan, Cheng-Nian Xiao, and Inanc Senocak. PittPack: an open-source Poisson’s equation solver for extreme-scale computing with accelerators. Computer Physics Communications, 254(??):Article 107272, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300898.

Hong:2014:MSI

[HY14] Hoon Hong and Chee Yap, editors. Mathematical Software — ICMS 2014: 4th International Conference, Seoul, South Korea, August 5–9, 2014, Proceedings, volume 8592. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (e-book). LCCN QA76.9.M35.

Hu:2020:API

[HYA20] Qinwen Hu, Se-Young Yu, and Muhammad Rizwan Asghar. Analysing performance issues of open-source intrusion detection systems in high-speed networks. Journal of Information Security and Applications (JISA), 51 (??):??, April 2020. CODEN ???? ISSN 2214-2126.

URL http://www.sciencedirect.com/science/article/
pii/S2214212619306003.

Heck:2014:HTJ

[HZ14] Petra Heck and Andy Zaidman. Horizontal traceability for just-in-time requirements: the case for open source feature requests. Journal of Software: Evolution and Process, 26(12): 1280–1296, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hunsen:2016:PBV

[HZS^{+}16] Claus Hunsen, Bo Zhang, Janet Siegmund, Christian Kästner, Olaf Leßenich, Martin Becker, and Sven Apel. Preprocessor-based variability in open-source and industrial software systems: An empirical study. Empirical Software Engineering, 21(2):449–482, April 2016. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10664-015-9360-1.

Ian:2002:IDA

[Ian02] Janis Ian. The Internet debacle — an alternative view. Performing Songwriter Magazine, 62:??, June 2002. URL http:/

/www.janisian.com/article-internet_debacle.html. See

also [Les03, Sta03c], and followup in the September/October issue of the magazine.

Imran:2016:WDA

[IAS16] A. Imran, S. Aljawarneh, and K. Sakib. Web data amalgamation for security engineering: Digital forensic investigation of open source cloud. J.UCS: Journal of Universal Computer Science, 22(4):494–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://www.jucs.

org/jucs_22_4/web_data_amalgamation_for.

Izquierdo:2022:ANC

[IC22] Javier Luis Cánovas Izquierdo and Jordi Cabot. On the analysis of non-coding roles in open source development. Empirical Software Engineering, 27(1):??, January 2022. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10664-021-10061-x.

Izquierdo:2023:VMT

[IC23] Javier Luis Cánovas Izquierdo and Jordi Cabot. Viewpoint: For a more transparent governance of open source. Communications of the ACM, 66(8):32–34, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.1145/3570635.

Ilg:2023:SCO

[IDSM23] Niclas Ilg, Paul Duplys, Dominik Sisejkovic, and Michael Menth. A survey of contemporary open-source honeypots, frameworks, and tools. Journal of Network and Computer Applications, 220(??):??, November 2023. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452300156X.

IEEE:1989:PII

[IEE89] IEEE, editor. Proceedings / 1989 IEEE International Conference on Robotics and Automation. Washington, DC. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1938-4. LCCN TJ 210.3 I44 1989. Three volumes. IEEE Catalog No. 89CH2750-8.

IEEE:1990:DPF

[IEE90] IEEE, editor. Digest of papers / Fault-Tolerant Computing, 20th International Symposium, June 26-28, 1990, Newcastle upon Tyne, England. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2051-X. LCCN QA 76.9 F38 I57 1990. IEEE Catalog No. 90CH2877-9.

IEEE:1992:DPI

[IEE92a] IEEE, editor. Digest of papers: the 1992 IEEE Workshop on Fault-Tolerant Parallel and Distributed Systems, July 6– 7, 1992, Amherst, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2870-7. LCCN QA76.9.F38I33 1992. IEEE catalog number 92TH0449-9.

IEEE:1992:FDP

[IEE92b] IEEE, editor. FTCS-22: digest of papers / the Twenty-second International Symposium on Fault-Tolerant Computing, July 8–10, 1992, Boston, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2875-8. LCCN QA 76.9 F38 I57 1992. IEEE Catalog No. 92CH3155-9.

IEEE:1992:PEA

[IEE92c] IEEE, editor. Proceedings / Eighth Annual Computer Security Applications Conference, San Antonio, Texas, November 30–December 4, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-3115-5 (paperback), 0-8186-3116-3 (microfiche), 0-8186-3117-1 (casebound). LCCN QA76.9.A25 C6375 1992. IEEE Computer Society Press order number 3115. IEEE catalog number 92TH04070-5.

IEEE:1992:PIS

[IEE92d] IEEE, editor. Proceedings / IEEE Southeastcon ’92, April 12–15, 1992, Birmingham, Alabama. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-7803-0494-2. LCCN TK 7801 I117 1992. Two volumes. IEEE Catalog No. 92CH3094-0.

IEEE:1993:FDP

[IEE93] IEEE, editor. FTCS-23: digest of papers: the Twenty-third International Symposium on Fault-Tolerant Computing, June 22-24, 1993, Toulouse, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, August 1993. ISBN 0-8186-3680-7. LCCN QA76.5.I58 1993.

IEEE:1994:FAS

[IEE94a] IEEE, editor. First Asia-Pacific Software Engineering Conference: proceedings, December 7–9, 1994, Tokyo, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6960-8 (paper), 0-8186-6961-6 (microfiche). LCCN QA76.758.A77 1994.

IEEE:1994:PTC

[IEE94b] IEEE, editor. Proceedings / the Tenth Conference on Artificial Intelligence for Applications, March 1–4, 1994, San Antonio, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5550-X. LCCN Q 334 C66 1994. IEEE Catalog No. 94CH3421-5.

IEEE:1994:PII

[IEE94c] IEEE, editor. Proceedings IEEE International Conference on Computer Design: VLSI in Computers and Processors (Cat. No.94CH35712). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6565-3. LCCN ????

IEEE:1995:PIC

[IEE95a] IEEE, editor. Proceedings of the International Conference on Multimedia Computing and Systems, May 15–18, 1995, Washington, DC. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7105-X. LCCN QA76.575.I623 1995. IEEE Catalog No. 95TH8066.

IEEE:1995:PNA

[IEE95b] IEEE, editor. Proceedings: the nineteenth annual International Computer Software and Applications Conference (COMPSAC ’95): August 9–11, 1995, Dallas, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7119-X. LCCN QA76.6 .I5 1995. IEEE Catalog No. 95CB35838.

IEEE:2005:PAC

[IEE05] IEEE, editor. Proceedings of the 21st Annual Computer Security Applications Conference 2005, December 05–09, 2005, Tucson, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2461-3. ISSN 1063-9527. LCCN L787.5. URL http:

//ieeexplore.ieee.org/servlet/opac?punumber=10467.

EEE Computer Society Order Number P2461.

Iaffaldano:2014:ROS

[IHBS14] G. Iaffaldano, R. Hawkins, T. Bodin, and M. Sambridge. REDBACK: Open-source software for efficient noise-reduction in plate kinematic reconstructions. Geochemistry, Geophysics, Geosystems, 15(4):1663–1670, 2014. ISSN 1525-2027.

Izquierdo:2019:OGD

[IHSR19] D. Izquierdo, N. Huesman, A. Serebrenik, and G. Robles. OpenStack gender diversity report. IEEE Software, 36(1):28– 33, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Davis:2001:PTS

[III01] Gervaise (Gerry) Davis III. Protecting trademarks on software in an open source environment: The Linux mark expe-

rience, 2001. URL http://www.linuxshowcase.org/tech.

html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Imtiaz:2023:OSF

[IKW23] Nasif Imtiaz, Aniqa Khanom, and Laurie Williams. Open or sneaky? Fast or slow? Light or heavy?: Investigating security releases of open source packages. IEEE Transactions on Software Engineering, 49(4):1540–1560, April 2023. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Iqbal:2010:POS

[ILG10] Syed Muhammad Zeeshan Iqbal, Yuchen Liang, and Hakan Grahn. ParMiBench — an open-source benchmark for embedded multiprocessor systems. IEEE Computer Architecture Letters, 9(2):45–48, July/December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ingber:1992:GAS

[Ing92] L. Ingber. GNU aids small science in a big way. GNU’s Bulletin, 1(12):9–10, 1992. CODEN ???? ISSN ????

Iwasaki:2002:DLB

[Iwa02] Hideya Iwasaki. Developing a Lisp-based preprocessor for TEX documents. Software—Practice and Experience, 32(14):1345– 1363, ???? 2002. CODEN SPEXBL. ISSN 0038-0644 (print),

1097-024X (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/98518913/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
98518913&PLACEBO=IE.pdf.

Jaeger:2005:GKE

[J^{+}05] Till Jaeger et al. Die GPL kommentiert und erklärt. (German) [The GPL discussed and explained]. O’Reilly, Köln, Germany, 2005. ISBN 3-89721-389-3. 182 pp. LCCN ????

Jaeger:2003:PBL

[Jae03] Andreas Jaeger. Porting to 64-bit Linux systems. In Hutton et al. [HDR03], pages 107–120. ISBN ????

LCCN ???? URL http://www.linux.org.uk/~ajh/gcc/
gccsummit-2003-proceedings.pdf.

Jakob:2003:FSR

[Jak03] Georg Jakob. Freie Softwarelizenzen — Rechtliche Aspekte der GPL, LGPL und BSDL. (German) [Free software licenses: Legal aspects of the GPL, LGPL, and BSDL]. World-Wide Web document (slide presentation)., 2003. URL http://

legacy.ffs.or.at/artikel/fs_lizenzen_slides.pdf.

Jakob:2004:FSG

[Jak04] Georg Jakob. Freiheit und Software. (German) [Freedom and software]. World-Wide Web document (slide presentation).,

2004. URL http://www.ffs.or.at/artikel/freiheit_sw.
pdf.

James:2009:FSC

[Jam09] Daniel James. Free Software for Creative People. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4302-1887-8. 450 pp. LCCN ????

Jang:2001:SLG

[Jan01] Michael H. Jang. Sair Linux/GNU installation and configuration. Coriolis Group Books, Scottsdale, AZ, USA, 2001. ISBN 1-57610-953-4. xxx + 351 pp. LCCN QA76.3 .J345 2001.

Janert:2008:GAU

[Jan08] Philipp Janert. Gnuplot in Action: Understanding Data with Graphs. Manning Publications, Greenwich, CT, USA, 2008. ISBN 1-933988-39-8. 275 (est.) pp. LCCN ???? US$35.00.

Joyner:2011:OSC

[JČMG11] David Joyner, Ondřej Čertík, Aaron Meurer, and Brian E. Granger. Open source computer algebra systems: SymPy.

ACM Communications in Computer Algebra, 45(3–4):225– 234, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jahanshahi:2022:WMT

[JCNS^{+}22] Hadi Jahanshahi, Mucahit Cevik, José Navas-Sú, Ayşe Başar, and Antonio González-Torres. Wayback Machine: a tool to capture the evolutionary behavior of the bug reports and their triage process in open-source software systems. The Journal of Systems and Software, 189(??):??, July 2022. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121222000565.

Jiang:2019:IDA

[JD19] Rui Jiang and Louis J. Durlofsky. Implementation and detailed assessment of a GNAT reduced-order model for subsurface flow simulation. Journal of Computational Physics, 379(??):192–213, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118307836.

Jones:2009:TIR

[JDB09] Paul Jones, Michael Day, and Alexander Ball. Topic 3: Institutional repositories should be built on open source software. Bulletin of the American Society for Information Science and Technology, 35(4):22–26, 2009. ISSN 2373-9223.

Joblin:2023:HHO

[JEB^{+}23] Mitchell Joblin, Barbara Eckl, Thomas Bock, Angelika Schmid, Janet Siegmund, and Sven Apel. Hierarchical and hybrid organizational structures in open-source software projects: a longitudinal study. ACM Transactions on Software Engineering and Methodology, 32(4):86:1–86:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3569949.

Jeffrey:2008:PAM

[Jef08] David Jeffrey, editor. Proceedings of the 21st annual meeting of the International Symposium on Symbolic Computation, IS-SAC 2008, July 20–23, 2008, Hagenberg, Austria. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-59593-904-0. LCCN ????

Jenson:1997:BRP

[Jen97] Rob Jenson. Book review: Programming with GNU software. ;login: the USENIX Association newsletter, 22(4):66–67, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Jenkins:2001:OSS

[Jen01] Steve Jenkins. Open-source software at the Aerodynamics Laboratory. Linux Journal, 90:70, 72–74, October 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-

3827 (electronic). URL http://noframes.linuxjournal.
com/lj-issues/issue90/4835.html.

Jeschke:2003:AFE

[Jes03a] Eric Jeschke. Antique film effects with the GIMP. Linux Journal, 2003(114):6, October 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Jeschke:2003:FPC

[Jes03b] Eric Jeschke. Fixing photo contrast with the GIMP. Linux Journal, 108:??, April 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Jeschke:2003:RRE

[Jes03c] Eric Jeschke. Removing red-eye with The GIMP. Linux Journal, 106:??, February 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Foss:2016:PFP

[jFFR16] Nicolai j. Foss, Lars Frederiksen, and Francesco Rullani. Problem-formulation and problem-solving in self-organized communities: How modes of communication shape project behaviors in the free open-source software community. Strategic Management Journal, 37(13):2589–2610, 2016. ISSN 0143-2095 (print), 1097-0266 (electronic).

Jeremias:2016:CEH

[JH16] Gunnar Jeremias and Mirko Himmel. Can everyone help verify the bioweapons convention? Perhaps, via open source monitoring. Bulletin of the Atomic Scientists, 72(6):412–417, 2016.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jin:2018:OMB

[Jin18] Zhi Jin. Open models: Beyond the open source software development. ACM SIGSOFT Software Engineering Notes, 43 (4):9–12, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jolitz:1991:PUT

[JJ91] W. F. Jolitz and L. G. Jolitz. Porting UNIX to the 386: three initial PC utilities. Dr. Dobb’s Journal of Software Tools, 16 (2):54, 56, 58–61, 111–112, 114–115, February 1991. CODEN DDJSDM. ISSN 0884-5395.

Jones:2000:LOS

[JJ00] Peter Jones and M. B. Jorgenson. Linux and open-source applications. Linux Journal, 70:??, February 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://noframes.linuxjournal.com/lj-
issues/issue/3683.html.

Johari:2011:ESE

[JK11] Kalpana Johari and Arvinder Kaur. Effect of software evolution on software metrics: an open source case study. ACM SIGSOFT Software Engineering Notes, 36(5):1–8, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johari:2012:VOO

[JK12] Kalpana Johari and Arvinder Kaur. Validation of object oriented metrics using open source software system: an empirical study. ACM SIGSOFT Software Engineering Notes, 37(1):1– 4, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jallad:2002:ICC

[JKS02] K. Jallad, J. Katz, and B. Schneier. Implementation of chosen-ciphertext attacks against PGP and GnuPG. In Lynn Batten and Jennifer Seberry, editors, Information security and privacy: 7th Australasian conference, ACISP 2002, Melbourne, Australia, July 3–5, 2002: proceedings, volume 2384, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-43861-0. LCCN QA76.9.A25 A278

2002. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t2384.htm;   http://www.counterpane.
com/pgp-attack.html.

Jiang:2017:WHD

[JLH^{+}17] Jing Jiang, David Lo, Jiahuan He, Xin Xia, Pavneet Singh Kochhar, and Li Zhang. Why and how developers fork what from whom in GitHub. Empirical Software Engineering, 22 (1):547–578, February 2017. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://link.

springer.com/article/10.1007/s10664-016-9436-6.

Janes:2023:OTT

[JLL23] Andrea Janes, Xiaozhou Li, and Valentina Lenarduzzi. Open tracing tools: Overview and critical comparison. The Journal of Systems and Software, 204(??):??, October 2023. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121223001887.

Johansson:2012:QOS

[JNN12] J. R. Johansson, P. D. Nation, and Franco Nori. QuTiP: an open-source Python framework for the dynamics of open quantum systems. Computer Physics Communications, 183 (8):1760–1772, August 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465512000835.

Johnson:1992:EPE

[Joh92] E. Johnson. Epsilon programmer’s editor (software review). Computers and the Humanities, 26(3):234–236, June 1992. CODEN COHUAD. ISSN 0010-4817 (print), 1572-8412 (electronic).

Johnson:1994:SMC

[Joh94a] J. H. Johnson. Substring matching for clone detection and change tracking. In Muller and Georges [MG94], pages 120–

126. ISBN 0-8186-6330-8. LCCN QA76.76.S64I58 1994.

Johnson:1994:VTR

[Joh94b] J. H. Johnson. Visualizing textual redundancy in legacy source. In Botsford et al. [BGG^{+}94], pages 9–18.

Johnson:1999:OSR

[Joh99] Shawana P. Johnson. Open source remote sensing effort. Linux Journal, 64:??, August 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.com/lj-issues/issue64/3418.
html.

Johnson:2002:OSS

[Joh02] Justin Pappas Johnson. Open source software: Private provision of a public good. Journal of Economics & Management Strategy, 11(4):637–662, 2002. ISSN 1058-6407 (print), 1530-9134 (electronic).

Johnson:2018:AAE

[Joh18] Robert W. Johnson. Algorithm 988: AMGKQ: An efficient implementation of adaptive multivariate Gauss–Kronrod quadrature for simultaneous integrands in Octave/MATLAB. ACM Transactions on Mathematical Software, 44(3):36:1– 36:19, April 2018. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://dl.acm.org/

citation.cfm?id=3157735.

Jones:2001:OSD

[Jon01] Paul Jones. Open (source)ing the doors for contributorrun digital libraries. Communications of the ACM, 44 (5):45–46, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/cacm/2001-44-5/p45-
jones/p45-jones.pdf;   http://www.acm.org/pubs/citations/
journals/cacm/2001-44-5/p45-jones/.

Jones:2002:HPC

[Jon02] Henry W. Jones, III. How a poor contract sunk an open-source deal. Linux Journal, 100:??, August 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/www.linuxjournal.com/article.php?sid=6025.

Jones:2005:GLA

[Jon05a] M. Tim Jones. GNU/Linux application programming. Charles River Media programming series. Charles River Media, Hingham, MA, USA, 2005. ISBN 1-58450-371-8. xxv + 486 pp. LCCN QA76.76.O63 J665 2005. URL http://www.loc.gov/

catdir/toc/ecip052/2004024882.html.

Jones:2005:OG

[Jon05b] M. Tim Jones. Optimization in GCC. Linux Journal, 2005 (131):11, March 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Jonsson:2005:MOB

[Jön05c] Per Jönsson. Modeller och beräkningar med GNU Octave. (Swedish) [Models and computations with GNU Octave]. Studentlitteratur, Lund, Sweden, 2005. ISBN 91-44-04360-0. 407 pp. LCCN ????

Jorgensen:2001:PIA

[Jør01] Niels Jørgensen. Putting it all in the trunk: incremental software development in the FreeBSD open source project. Information Systems Journal, 11(4):321–336, 2001. ISSN 1350-1917 (print), 1365-2575 (electronic).

Jordan:2004:ESL

[Jor04] Edmund Jordan. Embedded Systeme mit Linux programmieren: GNU-Softwaretools zur Programmierung ARMbasierender Systeme. (German) [Embedded Systems with Linux Programming: GNU Software Tools for Programming ARM-based Systems]. Franzis, Feldkirchen, Germany, 2004. ISBN 3-7723-5599-4. 384 pp. LCCN ????

Joyner:2008:OSC

[Joy08] David Joyner. Open source computer algebra systems: Axiom. ACM Communications in Computer Algebra, 42(1–2): 39–47, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Joyner:2009:ASS

[Joy09a] David Joyner. AMS special session on SAGE and mathematical research using open source software. ACM Communications in Computer Algebra, 43(2):49–54, June 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Joyner:2009:OSC

[Joy09b] David Joyner. Open source computer algebra systems: GAP. ACM Communications in Computer Algebra, 43(4):110–118, December 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jackin:2009:COS

[JP09a] B. J. Jackin and P. K. Palanisamy. A completely open source based computing system for computer generation of Fourier holograms. Computer Physics Communications, 180(10):1882–1887, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001246.

James:2009:CDM

[JP09b] Daniel James and Trevor Parsons. Crafting digital media: Audacity, Blender, Drupal, GIMP, Scribus, and other open source tools. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1888-6 (e-book), 1-4302-1887-8. xvi + 428 pp. LCCN QA76.76.I59 J36 2009eb.

Jiang:2020:PPF

[JPOB20] S. Jiang, P. Pan, Y. Ou, and C. Batten. PyMTL3: a Python framework for open-source hardware modeling, generation, simulation, and verification. IEEE Micro, 40(4):58–66, July/ August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jaust:2018:FAG

[JRA^{+}18] Alexander Jaust, Balthasar Reuter, Vadym Aizinger, Jochen Schütz, and Peter Knabner. FESTUNG: a MATLAB/GNU Octave toolbox for the discontinuous Galerkin method. Part III: Hybridized discontinuous Galerkin (HDG) formulation. Computers and Mathematics with Applications, 75(12):4505– 4533, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0898122118301895.

Joyner:2007:OSM

[JS07] David Joyner and William Stein. Open source mathematical software. Notices of the American Mathematical Society, 54 (10):1279, November 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.ams.

org/notices/200710/tx071001279p.pdf;   http://www.
sagemath.org/.

Jackson:2001:LOS

[JV01] Laura Elizabeth Jackson and Herbert Voß. LYX — an Open Source document processor. TUGboat, 22(1/2):32–41, March 2001. ISSN 0896-3207.

Jaschke:2018:OSM

[JWC18] Daniel Jaschke, Michael L. Wall, and Lincoln D. Carr. Open Source Matrix Product States: Opening ways to simulate entangled many-body quantum systems in one dimension. Computer Physics Communications, 225(??):59–91, April 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465517304204.

Jullien:2009:FCO

[JZ09] Nicolas Jullien and Jean-Benoît Zimmermann. Firms’ contribution to open-source software and the dominant user’s skill. European Management Review, 6(2):130–139, 2009. ISSN 1740-4754 (print), 1740-4762 (electronic).

Kamp:2011:SIP

[Kam11] Poul-Henning Kamp. The software industry IS the problem: The time has come for software liability laws. ACM Queue: Tomorrow’s Computing Today, 9(9):10, September 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamp:2014:BQS

[Kam14a] Poul-Henning Kamp. The bikeshed: Quality software costs money — Heartbleed was free. ACM Queue: Tomorrow’s Computing Today, 12(6):10, June 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamp:2014:QSC

[Kam14b] Poul-Henning Kamp. Quality software costs money — Heartbleed was free. ACM Queue: Tomorrow’s Computing Today, 12(6):10, June 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamensky:2021:OSI

[Kam21] David Kamensky. Open-source immersogeometric analysis of fluid–structure interaction using FEniCS and tIGAr. Computers and Mathematics with Applications, 81(??):634–648, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0898122120300481.

Kanig:2012:LEA

[Kan12] Johannes Kanig. Leading-edge Ada verification technologies: combining testing and verification with GNATTest and GNATProve — the Hi-Lite Project. ACM SIGADA Ada Letters, 32(3):5–6, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Kapur:1992:ADC

[Kap92] D. Kapur, editor. Automated deduction, CADE-11: 11th International Conference on Automated Deduction, Saratoga Springs, NY, USA, June 15–18, 1992: proceedings, volume 607 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55602-8. LCCN QA76.9.A96I57 1992.

Karels:2003:COS

[Kar03] Michael J. Karels. Commercializing open source software. ACM Queue: Tomorrow’s Computing Today, 1(5):46–55, July/August 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kawaguti:1992:SDC

[Kaw92] M. Kawaguti. A simple distributed computing system. In Verkerk and Wojcik [VW92], pages 877–880.

Kopec:1990:EIT

[KB90] G. E. Kopec and S. C. Bagley. Editing images of text. In Furuta [Fur90], pages 207–220. ISBN 0-521-40246-8. LCCN Z286.E43I58 1990.

Kortright:1992:CCT

[KC92] E. Kortright and D. Cordes. Cnest and cscope: Tools for the literate programming environment. In IEEE [IEE92d], pages 604–609 (vol. 2). ISBN 0-7803-0494-2. LCCN TK 7801 I117 1992. Two volumes. IEEE Catalog No. 92CH3094-0.

Kapitsaki:2021:MRO

[KC21] Georgia M. Kapitsaki and Georgia Charalambous. Modeling and recommending open source licenses with findOSSLicense. IEEE Transactions on Software Engineering, 47(5):919–935, May 2021. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Kaur:2022:EFA

[KC22] Rajdeep Kaur and Kuljit Kaur Chahal. Exploring factors affecting developer abandonment of open source software projects. Journal of Software: Evolution and Process, 34(9): e2484:1–e2484:??, September 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khatoonabadi:2023:WCU

[KCAS23] Sayedhassan Khatoonabadi, Diego Elias Costa, Rabe Abdalkareem, and Emad Shihab. On wasted contributions: Understanding the dynamics of contributor-abandoned pull requests — a mixed-methods study of 10 large open-source projects. ACM Transactions on Software Engineering and Methodology, 32(1):15:1–15:??, January 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/doi/10.1145/3530785.

Kurnia:2023:NOS

[KD23] Ruddy Kurnia and Guillaume Ducrozet. NEMOH: Open-source boundary element solver for computation of first- and secondorder hydrodynamic loads in the frequency domain. Computer Physics Communications, 292(??):Article 108885, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002308.

Karimi:2017:PNO

[KDM17] M. Karimi, H. Droghetti, and D. L. Marchisio. PU-Foam: a novel open-source CFD solver for the simulation of polyurethane foams. Computer Physics Communications, 217(??):138–148, August 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465517300978.

Kehoe:1994:PFS

[Keh94] Brendan Kehoe. Portability and free software. Dr. Dobb’s Developer Update, 1(11):2–??, October 1994. CODEN ???? ISSN 1079-8595.

Kent:2002:GLD

[Ken02] Jon Kent. GNU/Linux DVD player review. Linux Journal, 99:??, July 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://www.linuxjournal.

com/article/5644.

Kennedy:20xx:POS

[Kenxx] Dennis M. Kennedy. A primer on open source licensing legal issues: Copyright, copyleft and copyfuture. World-Wide

Web document., 20xx. URL http://www.denniskennedy.
com/opensourcedmk.pdf.

Kavaler:2017:SAO

[KF17] David Kavaler and Vladimir Filkov. Stochastic actor-oriented modeling for studying homophily and social influence in OSS projects. Empirical Software Engineering, 22(1):407–435, February 2017. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
article/10.1007/s10664-016-9431-y.

Kula:2013:MPA

[KFYI13] Raula Gaikovina Kula, Kyohei Fushida, Norihiro Yoshida, and Hajimu Iida. Micro process analysis of maintenance effort: an open source software case study using metrics based on program slicing. Journal of Software: Evolution and Process, 25 (9):935–955, 2013. ISSN 2047-7473 (print), 2047-7481 (electronic).

Kuhlins:2001:WEV

[KG01] S. Kuhlins and D. Gutacker. Web-enabled voice over IP call center. an open source based implementation. Lecture Notes in Computer Science, 2094:590–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2094/20940590.htm;   http:

//link.springer-ny.com/link/service/series/0558/papers/
2094/20940590.pdf.

Kumar:2020:MCS

[KG20] Rakesh Kumar and Rinkaj Goyal. Modeling continuous security: a conceptual model for automated DevSecOps using open-source software over cloud (ADOC). Computers & Security, 97(??):Article 101967, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302406.

Kazman:2016:EEA

[KGM^{+}16] R. Kazman, D. Goldenson, I. Monarch, W. Nichols, and G. Valetto. Evaluating the effects of architectural documentation: A case study of a large scale open source project. IEEE Transactions on Software Engineering, 42(3):220–260, March 2016. CODEN IESEDJ. ISSN 0098-5589 (print),

1939-3520 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=7230299.

Kawaguchi:2006:MAC

[KGMI06] Shinji Kawaguchi, Pankaj K. Garg, Makoto Matsushita, and Katsuro Inoue. MUDABlue: an automatic categorization system for Open Source repositories. The Journal of Systems and Software, 79(7):939–953, July 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kozicki:2022:IHP

[KGT22] Janek Kozicki, Anton Gladky, and Klaus Thoeni. Implementation of high-precision computation capabilities into the open-source dynamic simulation framework YADE. Computer Physics Communications, 270(??):Article 108167, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002794.

Koch:2021:DOS

[KGW^{+}21] Timo Koch, Dennis Gläser, Kilian Weishaupt, Sina Ackermann, Martin Beck, Beatrix Becker, Samuel Burbulla, Holger Class, Edward Coltman, Simon Emmert, Thomas Fetzer, Christoph Grüninger, Katharina Heck, Johannes Hommel, Theresa Kurz, Melanie Lipp, Farid Mohammadi, Samuel Scherrer, Martin Schneider, Gabriele Seitz, Leopold Stadler, Martin Utz, Felix Weinhardt, and Bernd Flemisch. DuMu^{x}3 — an open-source simulator for solving flow and transport problems in porous media with a focus on model coupling. Computers and Mathematics with Applications, 81(??):423–443, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0898122120300791.

Koponen:2005:OSS

[KH05] Timo Koponen and Virpi Hotti. Open source software maintenance process framework. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Koru:2003:LOS

[KHA^{+}03] A. Güneṡ Koru, Warren Harrison, Omar Alonso, Martin Fowler, and Holger Eichelberger. Letters: Open source research ethics; balancing both worlds; the agile sweet spot; more recent references. IEEE Software, 20(6):8–11, November/December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://csdl.

computer.org/dl/mags/so/2003/06/s6008.pdf.

Kilamo:2012:POS

[KHMA12] Terhi Kilamo, Imed Hammouda, Tommi Mikkonen, and Timo Aaltonen. From proprietary to open source — growing an open source ecosystem. The Journal of Systems and Software, 85(7):1467–1478, July 2012. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121211001683.

Kim:2001:TSJ

[Kim01a] Eugene Eric Kim. A triumph of simplicity: James Clark on markup languages and XML. Dr. Dobb’s Journal of Software Tools, 26(7):56, 58–60, July 2001. CODEN DDJOEB. ISSN

1044-789X. URL http://www.ddj.com/.

Kim:2001:OSI

[Kim01b] Gene Kim. Open source issues and opportunities for security tool vendors. In USENIX [USE01a], page ???? ISBN ????

LCCN ???? URL http://www.linuxshowcase.org/tech.

html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Kirtchev:2012:NRE

[Kir12] Hristian Hristov Kirtchev. A new robust and efficient implementation of controlled types in the GNAT compiler. ACM SIGADA Ada Letters, 32(3):43–50, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Kislitzin:1990:NMS

[Kis90] K. Kislitzin. Network monitoring by scripts. In Anonymous [Ano90c], pages 101–105.

Kitani:1994:MID

[Kit94] T. Kitani. Merging information by discourse processing for information extraction. In IEEE [IEE94b], pages 412–418. ISBN 0-8186-5550-X. LCCN Q 334 C66 1994. IEEE Catalog No. 94CH3421-5.

Katre:2003:RPW

[KJ03] S. A. Katre and Manjusha Joshi. Report on the Pune workshop on L^{A}TEX and free mathematical software. TUGboat, 24 (2):288–290, 2003. ISSN 0896-3207.

Krishnamurthy:2016:PDP

[KJRD16] Rajiv Krishnamurthy, Varghese Jacob, Suresh Radhakrishnan, and Kutsal Dogan. Peripheral developer participation in open source projects: an empirical analysis. ACM Transactions on Management Information Systems (TMIS), 6(4): 14:1–14:??, January 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Kawaguti:1994:CUI

[KK94] Minato Kawaguti and Norio Kitajima. Concurrent use of interactive TEX previewer with an Emacs-type editor. TUGboat, 15(3):293–300, September 1994. ISSN 0896-3207.

Kylander:1999:GUM

[KK99] Olof Kylander and Karin Kylander. The GIMP user’s manual. Coriolis Group Books, Scottsdale, AZ, USA, 1999. ISBN 1-57610-520-2. xxiv + 895 pp. LCCN T385 .K866 1999. US$49.99, CAN$73.99, UK£34.99.

Koo:2017:CUP

[KK17] Hyung-Min Koo and In-Young Ko. Construction and utilization of problem-solving knowledge in open source software environments. The Journal of Systems and Software, 131(??):402–418, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121216300899.

Koloniaris:2019:SBI

[KKA^{+}19] Stavros Koloniaris, George Kousiouris, Dimosthenis Anagnostopoulos, Mara Nikolaidou, and Konstantinos Tserpes. Survey-based investigation, feature extraction and classification of Greek municipalities maturity for open source adoption and migration prospects. The Journal of Systems and Software, 158(??):Article 110431, December 2019. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121219302055.

Krause:2021:OOS

[KKA^{+}21] Mathias J. Krause, Adrian Kummerländer, Samuel J. Avis, Halim Kusumaatmaja, Davide Dapelo, Fabian Klemens, Maximilian Gaedtke, Nicolas Hafen, Albert Mink, Robin Trunk, Jan E. Marquardt, Marie-Luise Maier, Marc Haussmann, and Stephan Simonis. OpenLB — open source lattice Boltzmann code. Computers and Mathematics with Applications, 81(??):258–288, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0898122120301875.

Kochhar:2021:MCO

[KKN^{+}21] Pavneet Singh Kochhar, Eirini Kalliamvakou, Nachiappan Nagappan, Thomas Zimmermann, and Christian Bird. Moving from closed to open source: Observations from six transitioned projects to GitHub. IEEE Transactions on Software Engineering, 47(9):1838–1856, September 2021.

Kapitsaki:2017:ALC

[KKT17] Georgia M. Kapitsaki, Frederik Kramer, and Nikolaos D. Tselikas. Automating the license compatibility process in open source software with SPDX. The Journal of Systems and Software, 131(??):386–401, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121216300905.

Koru:2007:ICC

[KL07] A. Güneş Koru and Hongfang Liu. Identifying and characterizing change-prone classes in two large-scale open-source products. The Journal of Systems and Software, 80(1):63–73, January 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kleinke:2021:CLL

[Kle21] Brian Kleinke. Challenges and lessons learned introducing an evolving open source technology into an established legacy Ada and C++ program. ACM SIGADA Ada Letters, 40(2): 70–72, April 2021. CODEN AALEE5. ISSN 1094-3641 (print),

1557-9476 (electronic). URL https://dl.acm.org/doi/10.
1145/3463478.3463485.

Klingler:1990:SSE

[Kli90] Carol Diane Klingler. Syntax-directed semantics-supported editing of algebraic specifications. Project report (m.s.), Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1990. xi + 131 pp.

Khullar:2007:SFO

[KMF^{+}07] D. Khullar, A. Molineu, D. Followill, W. Bosch, V. Willcut, T. Simpson, T. Ju, and J. Deasy. SU-FF-T-75: An opensource software tool to support film-based IMRT quality assurance. Medical Physics, 34(6Part8):2418, 2007. CODEN MPHYA6. ISSN 2473-4209.

Kahle:1993:IDS

[KMG^{+}93] B. Kahle, H. Morris, J. Goldman, T. Erickson, and J. Curran. Interfaces for distributed systems of information servers.

Journal of the American Society for Information Science, 44 (8):453–467, September 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kell:2007:FOS

[KMG^{+}07] L. T. Kell, I. Mosqueira, P. Grosjean, J-M. Fromentin, D. Garcia, R. Hillary, E. Jardim, S. Mardle, M. A. Pastoors, J. J. Poos, F. Scott, and R. D. Scott. FLR: an open-source framework for the evaluation and development of management strategies. ICES Journal of Marine Science, 64(4): 640–646, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/4/640/640024.

Kobayashi:1993:MGM

[KN93] Y. Kobayashi and M. Nagata. A method for generating messages on the on-line help system based on a user model and a situational model. In Smith and Salvendy [SS93], pages 621– 626 (vol. 2). ISBN 0-444-89540-X. LCCN QA76.9.I58 I593 1993. Two volumes.

Katz:2018:PYS

[KNS18] Daniel S. Katz, Kyle E. Niemeyer, and Arfon M. Smith. Publish your software: Introducing the Journal of Open Source Software (JOSS). Computing in Science and Engineering, 20 (3):84–88, ???? 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Knudsen:1999:AAO

[Knu99a] Craig Knudsen. AbiWord: AbiSource’s open source word processor. Linux Journal, 64:??, August 1999. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://noframes.linuxjournal.com/lj-
issues/issue64/3545.html.

Knudsen:1999:OSA

[Knu99b] Craig Knudsen. Open source with Applix. Linux Journal, 65:??, September 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue65/3574.html.

Koch:2007:SEO

[Koc07] Stefan Koch. Software evolution in open source projects — a large-scale investigation. Journal of Software Maintenance and Evolution: Research and Practice, 19(6):361–382, November 2007. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Koch:2009:EES

[Koc09] Stefan Koch. Exploring the effects of SourceForge.net coordination and communication tools on the efficiency of open source projects using data envelopment analysis. Empirical Software Engineering, 14(4):??, August 2009. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10664-008-9086-4.

Kakeshita:1994:FCS

[KOI94] T. Kakeshita, M. Oda, and Y. Imamura. Fall-in C: a software tool for pitfall detection in C programs. In IEEE [IEE94a], pages 256–265. ISBN 0-8186-6960-8 (paper), 0-8186-6961-6 (microfiche). LCCN QA76.758.A77 1994.

Kopper:2005:LEC

[Kop05] Karl Kopper. The Linux Enterprise Cluster: build a highly available cluster with commodity hardware and free software. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-036-4. xxiv + 430 pp. LCCN QA76.58 .K67 2005; QA76.58 K67 2005eb. URL http://www.oreilly.com/

catalog/9781593270360.

Koprowski:2020:CMP

[Kop20] Przemysław Koprowski. CQF Magma package. ACM Communications in Computer Algebra, 54(2):53–56, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3427218.
3427224.

Koranne:2011:HOS

[Kor11] Sandeep Koranne. Handbook of open source tools. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-7719-8 (e-book). xxxvi + 484 pp. LCCN QA76.76.S46 K67 2011.

Kneip:1995:ACE

[KORP95] J. Kneip, M. Ohmacht, K. Ronner, and P. Pirsch. Architecture and C++-programming environment of a highly parallel image signal processor. Microprocessing and Microprogramming, 41(5-6):391–408, October 1995. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

Kosior:2021:HFG

[Kos21] Wojciech Kosior. How i fought to graduate without using nonfree software. GNU Web essay., April 2021.

URL https://www.gnu.org/education/how-i-fought-to-
graduate-without-using-non-free-software.html.

Kotz:1990:GTV

[Kot90] David Kotz. GNUPLOT L^{A}TEX Tutorial Version 2.0. Computer Science Department, Duke University, Durham, NC, USA, February 1990. See also [WKC^{+}90].

Kotz:1991:CG

[Kot91] David Kotz. CUG334 GNUPLOT. C Users Journal, 9(4): 110–??, April 1991. ISSN 0898-9788.

Kernighan:1976:ST

[KP76] Brian W. Kernighan and P. J. Plauger. Software Tools. Addison-Wesley, Reading, MA, USA, 1976. ISBN 0-201-03669-X. 338 pp. LCCN QA76.6 .K42 1976.

Kernighan:1981:STP

[KP81] Brian W. Kernighan and P. J. Plauger. Software Tools in Pascal. Addison-Wesley, Reading, MA, USA, 1981. ISBN 0-201-10342-7. ix + 366 pp. LCCN QA76.6 .K493. US$13.95.

Kernighan:1984:UPE

[KP84] Brian W. Kernighan and Rob Pike. The UNIX Programming Environment. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1984. ISBN 0-13-937699-2 (hardcover), 0-13-937681-X (paperback). x + 357 pp. LCCN QA76.76.O63 K48 1984.

Knister:1993:IDT

[KP93] M. Knister and A. Prakash. Issues in the design of a toolkit for supporting multiple group editors. Computing Systems, 6(2): 135–166, Spring 1993. CODEN CMSYE2. ISSN 0895-6340.

Kernighan:1999:PP

[KP99] Brian W. Kernighan and Rob Pike. The Practice of Programming. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-61586-X. xii + 267 pp. LCCN QA76.6 .K48 1999. US$24.95,

CAN$37.50. URL http://cm.bell-labs.com/cm/cs/tpop/
code.html;   http://cseng.aw.com/bookdetail.qry?ISBN=
0-201-61586-X&ptype=0;   http://tpop.awl.com.

Kirsanskas:2017:QOS

[KPK^{+}17] Gediminas Kirsanskas, Jonas Nyvold Pedersen, Olov Karlström, Martin Leijnse, and Andreas Wacker. QmeQ 1.0: an open-source Python package for calculations of transport through quantum dot devices. Computer Physics Communications, 221(??):317–342, December 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465517302515.

Krafft:2005:DSC

[Kra05] Martin F. Krafft. The Debian system: concepts and techniques. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-069-0. 605 pp. LCCN QA76.76.O63 K68 2005; QA76.76.O63 K68 2005eb. URL http://www.loc.gov/

catdir/toc/ecip0515/2005019963.html.

Kurth:2022:OSP

[KRB^{+}22] Andreas Kurth, Wolfgang Rönninger, Thomas Benz, Matheus Cavalcante, Fabian Schuiki, Florian Zaruba, and Luca Benini. An open-source platform for high-performance non-coherent on-chip communication. IEEE Transactions on Computers, 71 (8):1794–1809, August 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Krebs:2000:GQL

[Kre00] W. G. Krebs. GNU Queue: Linux clustering made easy. Linux Journal, 79:144–??, November 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.com/lj-issues/issue79/4208.
html.

Kretchmar:2003:OSN

[Kre03] James Kretchmar. Open Source Network Administration. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-046210-1. 256 (est.) pp. LCCN TK5105.5 .K74 2003. US$44.99.

Krishnamurthy:1990:BSP

[Kri90] S. M. Krishnamurthy. A brief survey of papers on scheduling for pipelined processors. ACM SIGPLAN Notices, 25(7):97– 106, July 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kristan:2003:OSS

[Kri03] William B. Kristan. Open-source software for ecologists: a quick introduction to open-source software. The Bulletin of the Ecological Society of America, 84(1):30–33, 2003. ISSN 0012-9623 (print), 2327-6096 (electronic).

Kroeker:1999:NTNb

[Kro99a] Kirk L. Kroeker. New tools: Net development: Entrust’s Open Toolkits; IBM’s Parallel Enterprise Server; Talarian’s message-queueing product; SGI’s open source SAN management software. software development: ParaSoft’s automatic error-detection tool; Aonix’s GUI management system for Solaris 7. component technology: IONA Technologies’ EJB Toolkit; ProtoVIew’s diagramming tool. Computer, 32(6):107–110, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/r6107.pdf.

Kroeker:1999:NTNa

[Kro99b] Kirk L. Kroeker. New tools: Net development: Sun’s Java embedded server; MetaCreation’s Web-savvy graphics tool; WebCompiler’s HTML packaging tool. component technology: ProtoVIew Development’s ActiveX tools; Avilon Software’s load balancing component system. software development: Verilog’s test checker; Red Hat and Metrowerks’ development tools for Linux; the Object Factory’s optimization tool; Acumen Systems’s SDK for imaging; Aonix’s process-oriented lifecycle environment; Baan’s embedded software development suite. Computer, 32(5):103–107, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814

(electronic). URL http://dlib.computer.org/co/books/
co1999/pdf/r5103.pdf.

Kroll:1999:CRL

[Kro99c] Jason Kroll. CodeWarrior for Red Hat, Linux, GNU edition, version 4. Linux Journal, 66:??, October 1999. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://noframes.linuxjournal.com/lj-
issues/issue66/3654.html.

Kroll:1999:VLW

[Kro99d] Jason Kroll. VA Linux workstation VArStation XMP. Linux Journal, 67:??, November 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue67/3653.html.

Kroeker:2000:PIM

[Kro00] Kirk L. Kroeker. Products: iCompression’s MPEG-2 encoder reference design kit for USB; free project management support tools; Force Computers’ protocol software for telecom integration; Data Junction’s free XML tool; Fookes Software’s text editor for programmers; Research Systems’ IDL wavelet toolkit; Be opens source code to BeOS 5; JafSoft’s utility for converting text files into RTF. Computer, 33(5):98–100, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814

(electronic). URL http://dlib.computer.org/co/books/
co2000/pdf/r5098.pdf.

Kuhn:2023:CVS

[KRR23] Philipp Kühn, David N. Relke, and Christian Reuter. Common vulnerability scoring system prediction based on open source intelligence information sources. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823001967.

Koch:2002:ECO

[KS02] Stefan Koch and Georg Schneider. Effort, co-operation and co-ordination in an open source software project: GNOME. Information Systems Journal, 12(1):27–42, 2002. ISSN 1350-1917 (print), 1365-2575 (electronic).

Kuhn:2003:FGG

[KS03] Bradley Kuhn and Chris Sontag. Face-off: Is the GPL good for the software industry? Network World, 20(40): 45, October 6, 2003. ISSN 0887-7661 (print), 1944-7655

(electronic). URL http://www.nwfusion.com/columnists/

2003/1006faceoffno.html;   http://www.nwfusion.com/
columnists/2003/1006faceoffyes.html.

Krishna:2011:APV

[KS11] Raj P. M. Krishna and K. G. Srinivasa. Analysis of projects and volunteer participation in large scale free and open source software ecosystem. ACM SIGSOFT Software Engineering Notes, 36(2):1–5, March 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kulp:2012:SIM

[KSD^{+}12] Daniel W. Kulp, Sabareesh Subramaniam, Jason E. Donald, Brett T. Hannigan, Benjamin K. Mueller, Gevorg Grigoryan, and Alessandro Senes. Structural informatics, modeling, and design with an open-source Molecular Software Library (MSL). Journal of Computational Chemistry, 33(20): 1645–1661, 2012. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Koenka:2014:IOS

[KSH14] Israel Joel Koenka, Jorge Sáiz, and Peter C. Hauser. Instrumentino: an open-source modular Python framework for controlling Arduino based experimental instruments. Computer Physics Communications, 185(10):2724–2729, October 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465514002112.

Kanade:2009:VGO

[KSK09] Aditya Kanade, Amitabha Sanyal, and Uday P. Khedker. Validation of GCC optimizers through trace generation. Software—Practice and Experience, 39(6):611–639, April 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kress:2023:COS

[KSS^{+}23] Fabian Kreß, Vladimir Sidorenko, Patrick Schmidt, Julian Hoefer, Tim Hotfilter, Iris Walter, Tanja Harbaum, and Jürgen Becker. CNNParted: an open source framework for efficient Convolutional Neural Network inference partitioning in embedded systems. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128623002049.

Kramarev:2016:ITW

[KSV16] D. Kramarev, A. Sakzad, and E. Viterbo. Implementation of a two-way relay network with compute-and-forward in GNU Radio. Transactions on Emerging Telecommunications Technologies, 27(4):484–493, 2016. ISSN 2161-3915.

Koru:2004:DHM

[KT04] A. Gunes Koru and Jeff Tian. Defect handling in medium and large Open Source projects. IEEE Software, 21(4):54– 61, July/August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Koru:2005:CHC

[KT05] A. G. Koru and J. Tian. Comparing high-change modules and modules with the highest measurement values in two large-scale open-source products. IEEE Transactions on Software Engineering, 31(8):625–642, August 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1498769.

Kapitsaki:2015:ILT

[KTF15] Georgia M. Kapitsaki, Nikolaos D. Tselikas, and Ioannis E. Foukarakis. An insight into license tools for open source software systems. The Journal of Systems and Software, 102(??):72–87, April 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121214002945.

Katzner:2022:OSI

[KTH^{+}22] Todd Katzner, Eve Thomason, Karrin Huhmann, Tara Conkling, Camille Concepcion, Vince Slabe, and Sharon Poessel. Open-source intelligence for conservation biology. Conservation Biology, 36(6):e13988:1–e13988:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kermarrec:1995:IEF

[KTP95] Yvon Kermarrec, Laurent Nana Tchamnda, and Laurent Pautet. Implementing an efficient fault tolerance mechanism in Ada 9X: an early experiment with GNAT. Ada User Journal, 16(4):224–228, December 1995. CODEN AUJOET. ISSN 0268-652X.

Krieger:2017:BOS

[KTTK17] Michael T. Krieger, Oscar Torreno, Oswaldo Trelles, and Dieter Kranzlmüller. Building an open source cloud environment with auto-scaling resources for executing bioinformatics and biomedical workflows. Future Generation Computer Systems, 67(??):329–340, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X16300218.

Kuc:2006:ROS

[Kuc06] Bernard Kuc. Review of “Open Source Wireless Java Tools Suite by Robert Virkus,” Apress, 2005, $39.99, ISBN 1-59059-503-3. ACM Queue: Tomorrow’s Computing Today, 4(4):48, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Vir05].

Kukuk:1998:NPR

[Kuk98] Amy Kukuk. New products: Raritan MasterConsole MX4, Raritan Computer, Inc.; VRtuoso, Bittco Solutions; Debian GNU/Linux ARM, Debian GNU/Linux; LinuxCAD, Software Forge, Inc.; Perspective for Java, Three D Graphics; MetaCard 2.2, MetaCard Corp.; EtherPage Version 3.0, Personal Productivity Tools, Inc.; D 3 Linux v.7.1, Pick Systems. Linux Journal, 56:80, 95, December 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Kupfer:1993:SM

[Kup93] M. D. Kupfer. Sprite on mach. In Anonymous [Ano93d], pages 307–322.

Kushner:2005:OSM

[Kus05] D. Kushner. Open-source media [mobile multimedia software]. IEEE Spectrum, 42(2):50–51, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kozuch:1994:CES

[KW94] M. Kozuch and A. Wolfe. Compression of embedded system programs. In IEEE [IEE94c], pages 270–277. ISBN 0-8186-6565-3. LCCN ????

Kingsbury:1986:EEW

[KX86] Barry M. Kingsbury and John Xenakis. EMACS extension writing guide. Prime Computer, Natick, MA, USA, second edition, 1986.

Kerns:2016:SGB

[KY16] J. Kerns and D. Yaldo. SU-G-BRB-02: An open-source software analysis library for linear accelerator quality assurance. Medical Physics, 43(6Part24):3631, 2016. CODEN MPHYA6. ISSN 2473-4209.

Lapierre:2010:PEV

[LA10] Fabian D. Lapierre and Marc Acheroy. Performance enhancement and validation of the open-source software for modeling of ship infrared signatures (OSMOSIS). Journal of Computational and Applied Mathematics, 234(7):2342–2349, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037704270900572X.

Lallemand:1991:DAD

[Lal91] E. Lallemand. Development of an aid to definition of data types in LOTOS. Bulletin Scientifique de l’Association des Ingénieurs Electriciens sortis de l’Institut Electrotechnique Montefiore, 104(1):3–19, 1991. CODEN BURMA2. ISSN 0302-2676.

Lamastra:2009:SIC

[Lam09] Cristina Rossi Lamastra. Software innovativeness. a comparison between proprietary and free/open source solutions offered by Italian SMEs. R&D Management, 39(2):153–169, 2009. ISSN 0033-6807 (print), 1467-9310 (electronic).

Langmyhr:1989:TMG

[Lan89] Dag F. Langmyhr. Tekstredigering med GNU-Emacs. Kompendium 33, Institutt for informatikk, Oslo, Norway, 1989. 49 pp.

Pere:2005:GLR

[Lás05] Pere László. GNU/LINUX rendszerek üzemeltetése. (Hungarian) []. Kiskapu, Pécs, Hungary, 2005. ISBN ???? ???? pp. LCCN ????

Lattner:2003:ANG

[Lat03] Chris A. Lattner. Architecture for a next generation GCC. In Hutton et al. [HDR03], pages 121–133. ISBN ????

LCCN ???? URL http://www.linux.org.uk/~ajh/gcc/
gccsummit-2003-proceedings.pdf.

Lawton:2002:OSS

[Law02] George Lawton. Open source security: Opportunity or oxymoron? Computer, 35(3):18–21, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/co/books/co2002/pdf/
r3018.pdf;   http://www.computer.org/computer/co2002/
r3018abs.htm.

Lawton:2009:TNC

[Law09] George Lawton. Technology news: The changing face of open source. Computer, 42(5):14–17, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lazenby:1998:TE

[Laz98] Daniel Lazenby. TeraSpell 97 for Emacs. Linux Journal, 48:??, April 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-

3827 (electronic). URL http://noframes.linuxjournal.
com/lj-issues/2663.html.

Lazenby:1999:CGT

[Laz99] Daniel Lazenby. Cygnus GNUPro toolkit for Linux, v1.0. Linux Journal, 66:??, October 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.com/lj-issues/issue66/3455.
html.

Lipkowitz:2000:RCC

[LB00] Kenny B. Lipkowitz and Donald B. Boyd, editors. Reviews in Computational Chemistry, volume 15. Wiley, New York, NY, USA, January 2000. ISBN 0-471-36168-2 (hardcover), 0-470-12592-6 (e-book), 0-470-12619-1 (e-book). xxxv + 323 pp. LCCN QD39.3.E46 R48 2000.

Lundell:2022:ESU

[LBF^{+}22] Björn Lundell, Simon Butler, Thomas Fischer, Jonas Gamalielsson, Christoffer Brax, Jonas Feist, Tomas Gustavsson, Andrew Katz, Bengt Kvarnström, Erik Lönroth, and Anders Mattsson. Effective strategies for using open source software and open standards in organizational contexts: Experiences from the primary and secondary software sectors. IEEE Software, 39(1):84–92, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lin:2012:ATE

[LC12a] Kai-Wei Lin and Peng-Sheng Chen. An assistance tool employing a systematic methodology for GCC retargeting. Software—Practice and Experience, 42(1):19–36, January 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Llanos:2012:DBT

[LC12b] John Wilmar Castro Llanos and Silvia Teresita Acuña Castillo. Differences between traditional and open source development activities. Lecture Notes in Computer Science, 7343:131–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31063-8_11/.

Lecarme:2013:BGC

[LD13] Olivier Lecarme and Karine Delvare. The book of GIMP: a complete guide to nearly everything. No Starch Press, San Francisco, CA, USA, 2013. ISBN 1-59327-383-5. xvii + 656 pp. LCCN TR267.5.G56 L4313 2012. URL http://

www.loc.gov/catdir/enhancements/fy1214/2012020781-
b.html;   http://www.loc.gov/catdir/enhancements/fy1214/
2012020781-d.html.

Lea:1988:LGC

[Lea88] Douglas Lea. libg++, the GNU C++ library. In USENIX Association [USE88], pages 243–256.

Lea:1992:UGG

[Lea92] Doug Lea. User’s guide to the GNU C++ library: last updated 19 February, 1991 for version 1.39.0. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1992. iv + 134 pp.

Lea:1993:GCL

[Lea93] D. Lea. The GNU C++ library. C++ Report, 5(5):24–27, June 1993. CODEN CRPTE7. ISSN 1040-6042.

Leary:1994:NCD

[Lea94] K. Leary. Numerical C and DSP. Dr. Dobb’s Journal of Software Tools, 19(8):18–24, 90, August 1994. CODEN DDJSDM. ISSN 1044-789X.

Lee:1999:OSS

[Lee99] Steve H. Lee. Open Source software licensing. World-Wide

Web document., 1999. URL http://cyber.law.harvard.
edu/openlaw/gpl.pdf.

Leiner:1993:IIN

[Lei93a] B. Leiner, editor. INET ’93: International networking conference — August 1993, San Francisco, CA, Proceedings — INET 1993. Internet Society, Reston, VA, USA, 1993.

Leitz:1993:LEI

[Lei93b] Martin Leitz. LMAKE — Entwurf und Implementierung eines parallelen Make-Programmes. Master’s thesis, Technische Universität Braunschweig, Braunschweig, Germany, 1993.

Leibovitch:2004:EOS

[Lei04] Evan Leibovitch. EOF: Open source is for pigs. Linux Journal, 2004(124):13, August 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Leon:1998:GIM

[Leó98] Gonzalo León. The GNU’s Image Manipulation Program. Computer Graphics, 32(3):25–27, August 1998. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). Discusses the current status of GIMP, available from http://www.gimp.org, a developing competitor to Adobe Photoshop and other image manipulation packages.

Lessig:1999:COL

[Les99] Lawrence Lessig. Code and other laws of cyberspace. Basic Books, New York, 1999. ISBN 0-465-03912-X. xii + 297 pp. LCCN K564.C6 L47 1999. URL http://code-is-law.org.

Lessig:2001:COL

[Les01] Lawrence Lessig. El código y otras leyes del ciberespacio. Taurusesdigital, Madrid, Spain, 2001. ISBN 84-306-0428-6. 540 pp. LCCN K564.C6 L4718 2001.

Lesk:2003:CLR

[Les03] Michael Lesk. Chicken Little and the recorded music crisis. IEEE Security & Privacy, 1(5):73–75, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (elec-

tronic). URL http://csdl.computer.org/comp/mags/sp/
2003/05/j5073abs.htm;   http://csdl.computer.org/dl/
mags/sp/2003/05/j5073.pdf. See also [Ian02, Sta03c].

Levelt:1995:IPI

[Lev95a] A. H. M. Levelt, editor. ISSAC ’95: Proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation: July 10–12, 1995, Montreal, Canada, ISSAC — Proceedings — 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Levelt:1995:IIS

[Lev95b] A. H. M. (Antonius Henricus Maria) Levelt, editor. ISSAC’95: proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation, July 10–12, 1995, Montréal, Canada, ISSAC — Proceedings. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN A 76.95 I59 1995.

Levy:2023:CRU

[Lev23] Matthew L. Levy. Cybersecurity risks unique to open source and what communities are doing to reduce them. Computer, 56(6):78–83, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1988:GEL

[Lew88] Bil Lewis. GNU Emacs Lisp manual: Emacs version 18 for UNIX users. ????, ????, March 1988. various pp.

Lewin:1997:MGW

[Lew97] David I. Lewin. Mr. GNU’s who. Computers in physics, 11(5):406–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822578.

Lewis:1999:BCA

[Lew99a] Ted Lewis. Binary critic: Asbestos pajamas: An open source dialogue. Computer, 32(4):112, 108–111, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-

0814 (electronic). URL http://dlib.computer.org/co/
books/co1999/pdf/r4112.pdf. Continues the debate about

open source software development, notably in GNU/Linux [Lew99b].

Lewis:1999:BCO

[Lew99b] Ted Lewis. Binary critic: The open source acid test. Computer, 32(2):128, 125–127, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

dlib.computer.org/co/books/co1999/pdf/r2128.pdf. See

continuation in [Lew99a].

Li:1990:CTC

[LF90] C.-C. J. Li and W. K. Fuchs. CATCH — compiler-assisted techniques for checkpointing. In IEEE [IEE90], pages 74–81. ISBN 0-8186-2051-X. LCCN QA 76.9 F38 I57 1990. IEEE Catalog No. 90CH2877-9.

Long:1992:CSC

[LFA92] Junsheng Long, W. K. Fuchs, and J. A. Abraham. Compilerassisted static checkpoint insertion. In IEEE [IEE92a], pages 58–65. ISBN 0-8186-2870-7. LCCN QA76.9.F38I33 1992. IEEE catalog number 92TH0449-9.

Lethbridge:2021:UMD

[LFB^{+}21] Timothy C. Lethbridge, Andrew Forward, Omar Badreddin, Dusan Brestovansky, Miguel Garzon, Hamoud Aljamaan, Sultan Eid, Ahmed Husseini Orabi, Mahmoud Husseini Orabi, Vahdat Abdelzad, Opeyemi Adesina, Aliaa Alghamdi, Abdulaziz Algablan, and Amid Zakariapour. Umple: Modeldriven development for open source and education. Science of Computer Programming, 208(??):??, August 1, 2021. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167642321000587.

Lambert:2011:PBO

[LFN^{+}11] Emmanuel Lambert, Martin Fiers, Shavkat Nizamov, Martijn Tassaert, Steven G. Johnson, Peter Bienstman, and Wim Bogaerts. Python bindings for the Open Source Electromagnetic Simulator Meep. Computing in Science and Engineering, 13 (3):53–65, May/June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lawrie:2002:IDO

[LG02] Tony Lawrie and Cristina Gacek. Issues of dependability in open source software development. ACM SIGSOFT Software Engineering Notes, 27(3):34–37, May 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lucey:2020:CRE

[LGA20] Sean M. Lucey, Sarah K. Gaichas, and Kerim Y. Aydin. Conducting reproducible ecosystem modeling using the open source mass balance model Rpath. Ecological Modelling, 427(??):Article 109057, July 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301290.

Li:2017:EOS

[LGS^{+}17] Deguang Li, Bing Guo, Yan Shen, Junke Li, and Yanhui Huang. The evolution of open-source mobile applications: An empirical study. Journal of Software: Evolution and Process, 29(7):??, July 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lum:2018:OSD

[LGW18] Oliver Lum, Bruce Golden, and Edward Wasil. An opensource desktop application for generating arc-routing benchmark instances. INFORMS Journal on Computing, 30(2): 361–370, Spring 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.
org/doi/abs/10.1287/ijoc.2017.0785.

Lou:2022:OHT

[LGW^{+}22] Wenqi Lou, Lei Gong, Chao Wang, Zidong Du, and Xuehai Zhou. OctCNN: a high throughput FPGA accelerator for CNNs using Octave convolution algorithm. IEEE Transactions on Computers, 71(8):1847–1859, August 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lougee-Heimer:2003:COI

[LH03] R. Lougee-Heimer. The Common Optimization INterface for Operations Research: Promoting open-source software in the operations research community. IBM Journal of Research and Development, 47(1):57–66, ???? 2003. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.coin-or.org/;   http://www.
research.ibm.com/journal/rd/471/lougee.pdf.

Luan:2014:IPD

[LH14] Shao-Pu Luan and Chin-Yu Huang. An improved Pareto distribution for modelling the fault data of open source software. Software Testing, Verification and Reliability, 24(6):416–437, 2014. ISSN 0960-0833 (print), 1099-1689 (electronic).

Lopez:2022:GOS

[LH22] Joaquín López and Julio Hernández. gVOF: an opensource package for unsplit geometric volume of fluid methods on arbitrary grids. Computer Physics Communications, 277(??):Article 108400, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522001199.

Lai:1993:AAD

[LHCH93] Feipei Lai, Shu-Lin Hwang, Tzer-Shyong Chen, and Chia-Rung Hsieh. Arden — architecture development environment. In Baozong [Bao93], pages 5–9 (vol.1). ISBN 0-7803-1233-3. LCCN QA75.5.I155 1993. IEEE Catalog No. 93CH3286-2.

Liu:2012:SAS

[LHZ12] Haiguang Liu, Alexander Hexemer, and Peter H. Zwart. The Small Angle Scattering ToolBox (SASTBX): an open-source software for biomolecular small-angle scattering. Journal of Applied Crystallography, 45(3):587–593, 2012. CODEN JAC-GAR. ISSN 0021-8898 (print), 1600-5767 (electronic).

Li:1991:CEC

[Li91] Ping-Hsuan Li. Classifying Emacs commands for a better user interface. Thesis (m.s.), University of Florida, Gainesville, FL, USA, 1991. vii + 52 pp.

Li:1994:ILT

[Li94] Sing Li. An interview with Linus Torvalds. Dr. Dobb’s Journal of Software Tools, 19(5):26–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Li:2018:IPI

[Li18] YangQun Li. An integrated platform for the Internet of Things based on an open source ecosystem. Future Internet, 10(11): 105, October 31, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/10/11/105.

Lieberman:1992:GGE

[Lie92] J. E. Lieberman. GRef2End: a GeoRef to EndNote bibliography translator written in awk. Computers and Geosciences, 18(9):1271–1275, October 1992. CODEN CGEODT. ISSN 0098-3004 (print), 1873-7803 (electronic).

Link:2000:LGP

[Lin00] Jay Link. Linux graphics programming with SVGAlib. Coriolis Group Books, Scottsdale, AZ, USA, 2000. ISBN 1-57610-524-5. xxii + 513 pp. LCCN T385 .L564 2000.

Lingmann:2002:DSK

[Lin02a] Thomas Lingmann. Datenverschlüsselung: sichere Kommunikation mit Linux und BSD: Security mit Open Source. (German) [Data encoding: Secure communication with Linux and BSD: Security with Open Source]. C & L, Böblingen, Germany, 2002. ISBN 3-932311-87-8 (??invalid checksum??). 476 (est.) pp. LCCN ????

Anonymous:2002:LUM

[Lin02b] Anonymous, editor. Die Linuxzeitung: das unabhängige Monatsblatt für GNU/Linux, freie Software und Open Source. (German) [The Linux Times: the independent monthly for GNU/Linux, Free Software, and Open Source], 2002. URL

http://www.die.linuxzeitung.de/. ????, Berlin, Germany.

On-line journal.

Lindberg:2008:IPO

[Lin08] Van Lindberg. Intellectual Property and Open Source. O’Reilly Media, Sebastopol, CA, USA, 2008. ISBN 0-596-51796-3 (paperback). xv + 371 pp. LCCN K1519.C6 L56 2008.

Lions:1996:LCU

[Lio96] John Lions. Lions’ Commentary on UNIX 6th Edition, with Source Code. Computer classics revisited. Peer-to-Peer Communications, San Jose, CA 95164-0218, USA, 1996. ISBN 1-57398-013-7. 254 pp. LCCN QA76.76.O63 L562 1996.

US$29.96; CAN$41.95. URL http://www.peer-to-peer.
com/catalog/opsrc/lions.html. With forewords by Den-

nis M. Ritchie and Ken Thompson. Prefatory notes by Peter H. Salus and Michael Tilson; a Historical Note by Peter H. Salus; and Appreciations by Greg Rose, Mike O’Dell, Berny Goodheart, Peter Collinson, and Peter Reintjes. Originally circulated as two restricted-release volumes: “UNIX Operating System Source Code Level Six”, and “A Commentary on the UNIX Operating System”.

Lipkowitz:2007:RCC

[Lip07] Kenny B. Lipkowitz, editor. Reviews in Computational Chemistry, volume 23. Wiley, New York, NY, USA, 2007. ISBN 0-470-08201-1 (hardcover), 0-470-11643-9 (e-book), 0-470-11644-7, 0-470-18907-X. LCCN QD39.3.E46 R54 v.23 2007.

Litts:2014:WOS

[Lit14] Thom Litts. Why open source GIS software is not for me. Fisheries, 39(5):217, 2014. ISSN 0363-2415 (print), 1548-8446 (electronic).

Liu:2006:UOP

[Liu06] Shin-Ming Liu. Update on the Osprey Project, the alternative GCC backend for Itanium. In Anonymous [Ano06], page ??

ISBN ???? LCCN ???? URL http://www.ice.gelato.org/
oct06/pres_pdf/gelato_ICE06oct_osprey_liu_hp.pdf.

Liu:2008:UOS

[Liu08] Peter L. Liu. Using open-source robocode as a Java programming assignment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):63–67, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lin:2014:RBQ

[LL14] C. T. Lin and Y. F. Li. Rate-based queueing simulation model of open source software debugging activities. IEEE Transactions on Software Engineering, 40(11):1075–1099, November 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520

(electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6891380.

Llanos:2006:TUO

[Lla06] Diego R. Llanos. TPCC-UVa: an open-source TPC-C implementation for global performance measurement of computer systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):6–15, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lebl:2000:GSF

[LLdI00] George Lebl, Elliot Lee, and Miguel de Icaza. GNOME, its state and future. Linux Journal, 70:??, February 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://noframes.linuxjournal.com/lj-
issues/issue70/3754.html.

Larranaga:2023:OSI

[LLEL^{+}23] Ana Larrañaga, M. Carmen Lucas-Estañ, Sandra Lagén, Zoraze Ali, Imanol Martinez, and Javier Gozalvez. An open-source implementation and validation of 5G NR configured grant for URLLC in ns-3 5G LENA: a scheduling case study in industry 4.0 scenarios. Journal of Network and Computer Applications, 215(??):??, June 2023. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523000577.

Lewis:1990:GEL

[LLG90] Bil Lewis, Daniel LaLiberte, and GNU Manual Group. GNU Emacs Lisp reference manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1.03 edition, December 1990. ISBN 1-882114-10-8. xiv + 570 pp. LCCN ????

Lewis:1993:GEL

[LLG93] Bil Lewis, Daniel LaLiberte, and GNU Manual Group. GNU Emacs Lisp reference manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2.1 edition, September 1993. ISBN 1-882114-40-X. xiv + 570 pp. LCCN ????

Lewis:1994:GEL

[LLG94] Bil Lewis, Daniel LaLiberte, and GNU Manual Group. The GNU Emacs Lisp reference manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2.3 edition, June 1994. ISBN 1-882114-40-X. LCCN ????

Lundell:2011:PPO

[LLS11] Björn Lundell, Brian Lings, and Anna Syberfeldt. Practitioner perceptions of Open Source software in the embedded systems area. The Journal of Systems and Software, 84(9):1540–1549, September 2011. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121211000719.

Lewis:1999:GEL

[LLSt99] Bil Lewis, Dan LaLiberte, Richard Stallman, and the GNU Manual Group. GNU Emacs Lisp Reference Manual, for Emacs Version 20.4. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1999. ISBN 1-882114-72-8. ???? pp. LCCN ????

Lewis:2000:GEL

[LLSt00] Bil Lewis, Dan LaLiberte, Richard Stallman, and the GNU Manual Group. GNU Emacs Lisp Reference Manual, for Emacs Version 21. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2000. ISBN 1-882114-73-6. 974 pp.

LCCN ???? URL http://www.gnupress.org/book3.html.

Two volumes.

Li:2023:IDP

[LLWM23] Na Li, XiaoLing Liu, Yu Wang, and Musa Mojarad. Improving dynamic placement of virtual machines in cloud data centers based on open-source development model algorithm. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10723-023-09651-4.

Li:2020:TOS

[LMHL20] Dongyue Li, Daniele Marchisio, Christian Hasse, and Dirk Lucas. twoWayGPBEFoam: an open-source Eulerian QBMM solver for monokinetic bubbly flows. Computer Physics Communications, 250(??):Article 107036, May 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303728.

Lever:2002:OSS

[LMM02] John A. Lever, Chris Maresca, and Robert Munro. 9.2.4 open source software assessment for the Naval Oceanographic Office. INCOSE International Symposium, 12(1):423–429, 2002. ISSN 2334-5837.

Loosemore:1993:GCL

[LMOS93] Sandra Loosemore, Roland MacGrath, Andrew Oram, and Richard M. Stallmann. The GNU C Library reference manual: last updated 20 June 1993 for version 1.07 beta. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1993. xvi + 654 pp. LCCN ????

Li:2022:OAR

[LMPT22] Xiaozhou Li, Sergio Moreschini, Fabiano Pecorelli, and Davide Taibi. OSSARA: Abandonment risk assessment for embedded open source components. IEEE Software, 39(4):48–53, July/ August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Logg:2012:ASD

[LMW12] Anders Logg, Kent-Andre Mardal, and Garth Wells, editors. Automated Solution of Differential Equations by the Finite Element Method: The FEniCS Book, volume 84 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSA6.

ISBN 3-642-23098-9 (print), 3-642-23099-7 (e-book). ISSN 1439-7358 (print), 2197-7100 (electronic). xiii + 723 pp.

LCCN ???? URL http://fenicsproject.org;   http:
//fenicsproject.org/book/;   http://link.springer.
com/book/10.1007/978-3-642-23099-8;   http://www.springerlink.
com/content/978-3-642-23099-8;   https://launchpad.

net/fenics-book. The software developed by the FEniCS Project is free for all to use and modify (licensed under the GNU (L)GPL), and so is this book.

Linaaker:2018:MCO

[LMWM18] J. Linåker, H. Munir, K. Wnuk, and C. E. Mols. Motivating the contributions: an Open Innovation perspective on what to share as Open Source Software. The Journal of Systems and Software, 135(??):17–36, January 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121217302169.

Luo:2019:HDS

[LMZP19] Q. Luo, K. Moran, L. Zhang, and D. Poshyvanyk. How do static and dynamic test case prioritization techniques perform on modern software systems? An extensive study on GitHub projects. IEEE Transactions on Software Engineering, 45(11): 1054–1080, November 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Li:2022:EFM

[LMZT22] Xiaozhou Li, Sergio Moreschini, Zheying Zhang, and Davide Taibi. Exploring factors and metrics to select open source software components for integration: an empirical study. The Journal of Systems and Software, 188(??):??, June 2022. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121222000267.

Leach:1992:MTP

[LN92] J. Leach and S. Nieva. MIZ-PR: a theorem prover for polymorphic and recursive functions. In Voronkov [Vor92], pages 481–483. ISBN 0-387-55727-X (New York), 3-540-55727-X (Berlin). LCCN QA76.63.I55 1992.

Liu:1989:ELR

[LO89] Sying-Syang Liu and Roger Ogando. An Emacs-based logical ripple effect analyzer prototype user’s manual. Technical Report SERC-TR-32-F, Software Engineering Research Centre, ????, September 1989.

Lius:1992:OFU

[LO92] S. Lius and R. Ogando. The object finder user’s manual. Technical Report SERC-TR-56-F, Software Engineering Research Centre, ????, January 1992.

Loukides:1997:PGS

[LO97] Mike Loukides and Andy Oram. Programming with GNU Software. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-112-7. xiv + 244 pp. LCCN QA76.76.O63L65 1997. US$39.95. URL http://

www.ora.com/catalog/prognu/;   http://www.oreilly.
com/catalog/prognu. Includes CD-ROM.

Logan:1999:AGG

[Log99] Syd Logan. The artists’ guide to the Gimp. Linux Journal, 64:??, August 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue64/3352.html.

Loki:2004:CFG

[Lók04] Gábor Lóki. Code factoring in GCC. In Hutton et al. [HDR04], pages 79–84. ISBN ???? LCCN ???? URL http://

people.redhat.com/lockhart/.gcc04/MasterGCC-2side.
pdf.

Loosemore:2015:GCL

[Loo15] Sandra Loosemore. GNU C Library 2.22 reference manual. Samurai Media Limited, Hong Kong, 2015. ISBN 988-8381-07-5 (vol. 1), 988-8381-08-3 (vol. 2). xx + 1075 (2 volumes) pp. LCCN ????

Lord:1995:AGI

[Lor95] Thomas Lord. An anatomy of guile: The interface to Tcl/Tk. In Association [Ass95], pages 95–114. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

Loui:1996:WGA

[Lou96] Ronald P. Loui. Why GAWK for AI? ACM SIGPLAN Notices, 31(8):8–9, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Love:2006:RGD

[Lov06] Robert Love. Rapid GNOME development with Mono. Linux Journal, 2006(143):8, March 2006. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Liu:1991:OFA

[LOW91] Sying-Syang Liu, Roger Ogando, and Norman Wilde. The object finder: a design recovery tool. Technical Report SERC-TR-46-F, Software Engineering Research Centre, ????, January 1991.

Lauricella:2015:JSP

[LPC^{+}15] Marco Lauricella, Giuseppe Pontrelli, Ivan Coluzza, Dario Pisignano, and Sauro Succi. JETSPIN: a specific-purpose open-source software for simulations of nanofiber electrospinning. Computer Physics Communications, 197(??):227–238, December 2015. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465515002994.

Li:2021:CCC

[LPFD21] Renee Li, Pavitthra Pandurangan, Hana Frluckaj, and Laura Dabbish. Code of conduct conversations in open source software projects on Github. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):19:1– 19:31, April 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.org/doi/10.1145/3449093.

Li:2017:HDT

[LQ17] Xigao Li and Lin Qian. A hybrid disaster-tolerant model with DDF technology for MooseFS open-source distributed file system. The Journal of Supercomputing, 73(5):2052–2068, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lomuscio:2017:MOS

[LQR17] Alessio Lomuscio, Hongyang Qu, and Franco Raimondi. MCMAS: an open-source model checker for the verification of multi-agent systems. International Journal on Software Tools for Technology Transfer (STTT), 19(1):9–30, February 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787

(electronic). URL http://link.springer.com/article/
10.1007/s10009-015-0378-x;   http://link.springer.
com/content/pdf/10.1007/s10009-015-0378-x.pdf.

Li:2008:MOS

[LR08] Jinhu Li and Jeffrey S. Racine. Maxima: An open source computer algebra system. Journal of Applied Econometrics, 23(4):515–523, June 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lam:2011:EMI

[LR11] Shyong (Tony) K. Lam and John Riedl. Expressing my inner gnome: Appearance and behavior in virtual worlds. Computer, 44(7):103–105, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Linaaker:2023:OSS

[LRBM23] Johan Linåker, Gregorio Robles, Deborah Bryant, and Sachiko Muto. Open source software in the public sector: 25 years and still in its infancy. IEEE Software, 40(4):39–44, 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Liu:2019:WFA

[LRD^{+}19] Bohan Liu, Guoping Rong, Liming Dong, He Zhang, Danni Chen, Tiange Chen, Yuyan Chen, and Tiantian Zhang. What are the factors affecting the handover process in open source development? The Journal of Systems and Software, 153(??):238–254, July 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121219300718.

Lindman:2011:MOS

[LRP11] Juho Lindman, Matti Rossi, and Anna Puustell. Matching open source software licenses with corresponding business models. IEEE Software, 28(4):31–35, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lima:2021:AEH

[LRP21] Luan P. Lima, Lincoln S. Rocha, and Matheus Paixao. Assessing exception handling testing practices in open-source libraries. Empirical Software Engineering, 26(5):??, September 2021. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616

(electronic). URL https://link.springer.com/article/
10.1007/s10664-021-09983-3.

Loosemore:2004:GCL

[LS04] Sandra Loosemore and Richard Stallman. The GNU C library: system and network applications for GNU C Libraries version 2.3.x. A GNU manual. GNU Press, Boston, MA, USA, version 2.3.x. edition, 2004. ISBN 1-882114-24-8. xiv + 606 pp. LCCN QA76.73.C15 L66 2004.

Li:1994:CFC

[LSF94] Chung-Chi Jim Li, E. M. Stewart, and W. K. Fuchs. Compiler-assisted full checkpointing. Software—Practice and Experience, 24(10):871–886, October 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leptos:2006:MOS

[LSJ^{+}06] Kyriacos C. Leptos, David A. Sarracino, Jacob D. Jaffe, Bryan Krastins, and George M. Church. MapQuant: Opensource software for large-scale protein quantification. PRO-TEOMICS, 6(6):1770–1782, 2006. ISSN 1615-9853 (print), 1615-9861 (electronic).

Loosemore:1999:GCL

[LSM^{+}99] Sandra Loosemore, Richard M. Stallman, Roland McGrath, Andrew Oram, and Ulrich Drepper. The GNU C Library Reference Manual (for Version 2.1 Beta). Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 0.08 draft edition, 1999. ISBN 1-882114-53-1. ???? pp. LCCN ????

Loosemore:2000:GCL

[LSM^{+}00] Sandra Loosemore, Richard M. Stallman, Roland McGrath, Andrew Oram, and Ulrich Drepper. GNU C Library Reference Manual, for Version 2.x. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2000. ISBN 1-882114-54-X. ???? pp. LCCN ???? Two volumes.

Loosemore:2001:GCL

[LSM^{+}01] Sandra Loosemore, Richard M. Stallman, Roland McGrath, Andrew Oram, and Ulrich Drepper. The GNU C Library Reference Manual. GNU Press, Boston, MA, USA, 2001. ISBN 1-882114-55-8. xxii + 625–1178 pp. LCCN ???? US$60.00.

URL http://www.gnupress.org/book6.html. Two volumes.

For glibc version 2.2.x.

Lewis:2009:ROS

[LSM09] Ian A. Lewis, Seth C. Schommer, and John L. Markley. rNMR: open source software for identifying and quantifying metabolites in NMR spectra. Magnetic Resonance in Chemistry, 47(S1):S123–S126, 2009. ISSN 0749-1581 (print), 1097-458X (electronic).

Loosemore:1996:GCL

[LSMO96] Sandra Loosemore, Richard M. Stallman, Roland McGrath, and Andrew Oram. The GNU C Library reference manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1996. ISBN 1-882114-53-1. xvi + 674 pp. LCCN ???? 0.07 draft, last updated 21 May 1996.

Lucas:1999:MPL

[Luc99a] Michael Lucas. Maintaining patch levels with open source BSDs. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):35–36, 38, 40, 42, September 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.samag.com/.

Lucky:1999:FSR

[Luc99b] R. W. Lucky. Free software [reflections]. IEEE Spectrum, 36 (5):24, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lussier:2004:NTH

[Lus04] Stephane Lussier. New tricks: How Open Source changed the way my team works. IEEE Software, 21(1):68–72, January/ February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lee:2003:OSW

[LW03] James Lee and Brent Ware. Open source Web development with LAMP: using Linux, Apache, MySQL, Perl, and PHP. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-77061-X (paperback). xxxiv + 460 pp. LCCN QA76.76.D47 L435 2003.

Lonie:2011:XOS

[LZ11a] David C. Lonie and Eva Zurek. XtalOpt: An open-source evolutionary algorithm for crystal structure prediction. Computer Physics Communications, 182(2):372–387, February 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465510003140.

Lonie:2011:XVR

[LZ11b] David C. Lonie and Eva Zurek. XtalOpt version r7: An opensource evolutionary algorithm for crystal structure prediction. Computer Physics Communications, 182(10):2305–2306, October 2011. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046551100213X.

Lonie:2012:IDC

[LZ12] David C. Lonie and Eva Zurek. Identifying duplicate crystal structures: XtalComp, an open-source solution. Computer Physics Communications, 183(3):690–697, March 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465511003699.

Lefevre:2016:OBB

[LZ16] Vincent Lefèvre and Paul Zimmermann. Optimized Binary64 and Binary128 arithmetic with GNU MPFR. Report hal-01502326, Inria Grenoble — Rhône-Alpes, LIP — Laboratoire de l’Informatique du Parallélisme and Inria Nancy — Grand Est, LORIA — ALGO — Department of Algorithms, Computation, Image and Geometry, April 2016. 10 pp. URL

https://hal.inria.fr/hal-01502326. To appear in IEEE

ARITH’2016 proceedings London, UK (24–26 July, 2016).

Lefevre:2017:OBB

[LZ17] V. Lefèvre and P. Zimmermann. Optimized Binary64 and Binary128 arithmetic with GNU MPFR. In Burgess et al. [BBdD17], pages 18–26. ISBN 1-5386-1966-0 (print), 1-5386-1965-2, 1-5386-1964-4. ISSN 1063-6889. LCCN QA76.9.C62

S95 2017. URL http://ieeexplore.ieee.org/servlet/
opac?punumber=8019911.

Li:2022:SPL

[LZWH22] Heng Li, Haoxiang Zhang, Shaowei Wang, and Ahmed E. Hassan. Studying the practices of logging exception stack traces in open-source software projects. IEEE Transactions on Software Engineering, 48(12):4907–4924, December 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

MacKinnon:1999:LOS

[Mac99] James G. MacKinnon. The Linux operating system: Debian GNU/Linux. Journal of Applied Econometrics, 14(4):443– 452, July–August 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

MacCarty:2002:LDG

[Mac02] Bill MacCarty. Learning Debian GNU Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 1-56592-705-2. ???? pp. LCCN QA76.76.O63 M372 1999. US$34.95. Includes CD-ROM.

Mackey:2018:BOS

[Mac18] Tim Mackey. Building open source security into agile application builds. Network Security, 2018(4):5–8, April 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1353485818300321.

Moreira:2022:OSS

[MAF22] Rodrigo André Ferreira Moreira, Wesley K. G. Assunção, and Eduardo Figueiredo. Open-source software product line extraction processes: the ArgoUML-SPL and Phaser cases. Empirical Software Engineering, 27(4):??, July 2022. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10664-021-10104-3.

Maginnis:2000:SLG

[Mag00] Tobin Maginnis. Sair Linux and GNU certification: Level I: Installation and configuration. Wiley, New York, NY, USA, 2000. ISBN 0-471-36978-0. xix + 284 pp. LCCN QA76.3 .M3235 2000.

Maginnis:2001:SLGa

[Mag01a] Tobin Maginnis. Sair Linux and GNU certification level I: installation and configuration. Wiley, New York, NY, USA, 2001. ISBN 0-471-41797-1. xix + 508 pp. LCCN QA76.3 .M32337 2001.

Maginnis:2001:SLGb

[Mag01b] Tobin Maginnis. Sair Linux and GNU certification level I: networking. Wiley, New York, NY, USA, 2001. ISBN 0-471-36977-2. xx + 476 pp. LCCN QA76.3 .M32337 2000.

Maginnis:2001:SLGc

[Mag01c] Tobin Maginnis. Sair Linux and GNU certification level I: security, ethics, and privacy. Wiley, New York, NY, USA, 2001. ISBN 0-471-36975-6. xix + 395 pp. LCCN QA76.3 .M32339 2001.

Maginnis:2004:GLZ

[Mag04] Tobin Maginnis. GNU-Linux-Zertifizierung: Prüfungsvorbereitung zum Sair Linux and GNU Certified Professional/Administrator (LCP/LCA). (German) [GNU/Linux Certification: Examination Preparation for Sair Linux and GNU Certified Professional/Administrator (LCP/LCA)]. dpunkt-Verlag, Heidelberg, Germany, 2004. ISBN ???? ???? pp. LCCN ????

Mahony:2003:GCH

[Mah03] Siobhán O. Mahony. Guarding the commons: how community managed software projects protect their work. Research Policy, 32(7):1179–1198, 2003. CODEN ???? ISSN 0048-7333 (print), 1873-7625 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0048733303000489.

Mahout:2013:GA

[Mah13] Vincent Mahout. The GNU Assembler, pages 227–238. Wiley, New York, NY, USA, 2013. ISBN 1-118-56212-7.

Majerus:2003:CEM

[Maj03] Laura A. Majerus. Court evaluates meaning of “derivative work” in an Open Source license. World-Wide Web doc-

ument., 2003. URL http://library.findlaw.com/2003/
Jun/16/132811.html.

Makarov:2003:FSA

[Mak03] Vladimir N. Makarov. The finite state automaton based pipeline hazard recognizer and instruction scheduler in GCC. In Hutton et al. [HDR03], pages 135–149. ISBN ????

LCCN ???? URL http://www.linux.org.uk/~ajh/gcc/
gccsummit-2003-proceedings.pdf.

Makarov:2004:FRP

[Mak04] Vladimir N. Makarov. Fighting register pressure in GCC. In Hutton et al. [HDR04], pages 85–103. ISBN ???? LCCN

???? URL http://people.redhat.com/lockhart/.gcc04/
MasterGCC-2side.pdf.

Malcolm:2002:POS

[Mal02] Jeremy Malcolm. Problems in Open Source licensing. World-Wide Web document., 2002. URL http://www.ilaw.com.

au/public/licencearticle.html.

Mamlin:2001:OSX

[Mam01] B. Mamlin. An open-source XML-based Java implementation of the medical gopher order and note writing tool. Journal of Biomedical Informatics, 35(SUPP):826–??, 2001. CODEN JBIOBL. ISSN 1532-0464.

Marques:2005:UOI

[MAMC05] Nuno C. Marques, Francisco Azevedo, Carmen Morgado, and Jorge F. Custódio. Using Octave to introduce programming to technical science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3): 198–202, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maner:1992:APE

[Man92] W. Maner. Adapting a programming editor (EMACS) for persons who are blind: issues, initiatives and problems. Journal of Microcomputer Applications, 15(1):13–19 (or 13–20??), January 1992. CODEN JMIADO. ISSN 0745-7138 (print), 1096-374X (electronic).

Mann:2000:GLW

[Man00] Steve Mann. A GNU/Linux wristwatch videophone. Linux Journal, 75:??, July 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue75/3993.html.

Mantz:2003:ROS

[Man03] Reto Mantz. Rechtsfragen von Open Source. (German) [Legal questions about Open Source]. World-Wide Web document.,

2003. URL http://www.retosphere.de/opensource/Rechtsfragen_
von_Open_Source.pdf.

Marti:2001:TCD

[Mar01] Don Marti. Thought crimes, databases, kernel hacking and other news from the O’Reilly Open Source Conference. Linux Journal, 90:??, October 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://www.

linuxjournal.com/articles/tradeshow/0039.html. Web

only.

Marcotte:2003:PUG

[Mar03] Ludovic Marcotte. Programming under GNUstep — an introduction. Linux Journal, 108:??, April 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Marti:2005:ROS

[Mar05] Don Marti. Reviews: Open source licensing: Software freedom and intellectual property law. Linux Journal, 2005(129):19, January 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Marcotte:2011:SSG

[Mar11] Ludovic Marcotte. SOGo — open-source groupware. Linux Journal, 2011(201):1:1–1:??, January 2011. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Marzolla:2022:QNM

[Mar22] Moreno Marzolla. Queueing networks and Markov chains analysis with the Octave queueing package. ACM SIGMET-RICS Performance Evaluation Review, 49(4):47–52, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3543146.
3543158.

Massey:2005:LAL

[Mas05] Bart Massey. Longitudinal analysis of long-timescale open source repository data. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Matz:2003:DIG

[Mat03] Michael Matz. Design and implementation of the graph coloring register allocator for GCC. In Hutton et al. [HDR03], pages 151–169. ISBN ???? LCCN ???? URL http://www.

linux.org.uk/~ajh/gcc/gccsummit-2003-proceedings.pdf.

Maurer:2005:NOS

[Mau05] Jim Maurer. News 2.0: Open Source/2; anti-spam activism … or vigilantism?; ride, robot, ride. ACM Queue: Tomorrow’s Computing Today, 3(7):8, September 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maxwell:2001:LCK

[Max01] Scott Andrew Maxwell. Linux Core Kernel commentary. Coriolis Group Books, Scottsdale, AZ, USA, 2001. ISBN 1-58880-149-7. xviii + 717 pp. LCCN QA76.76.O63 M373337 2001.

May:2006:ETT

[May06] Christopher May. Escaping the TRIPs’ trap: The political economy of free and open source software in Africa. Political Studies, 54(1):123–146, 2006. ISSN 0032-3217 (print), 1467-9248 (electronic).

Mayer:2017:TCL

[May17] Philip Mayer. A taxonomy of cross-language linking mechanisms in open source frameworks. Computing, 99(7):701–724, July 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mazzini:2015:CPO

[Maz15] Silvia Mazzini. The CONCERTO Project: an open source methodology for designing, deploying, and operating reliable and safe CPS systems. Ada User Journal, 36(4):264–267, December 2015. CODEN AUJOET. ISSN 1381-6551.

Millet:1998:PGT

[MB98] Laurent Millet and Ted Baker. Porting the GNAT tasking runtime system to the Java Virtual Machine. Lecture Notes in Computer Science, 1411:19–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1411/14110019.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1411/14110019.pdf.

Murugesan:2016:OSC

[MB16] San Murugesan and Irena Bojanova. Open-source cloud software solutions. In G. R. Gangadharan, Deepnarayan Tiwari, Lalit Sanagavarapu, Shakti Mishra, Abraham Williams, and Srimanyu Timmaraju, editors, Encyclopedia of Cloud Computing, chapter 12, pages 139–149. Wiley, New York, NY, USA, 2016. ISBN 1-118-82193-9.

Maclean:2021:TEF

[MBR21] John Maclean, J. E. Bunder, and A. J. Roberts. A toolbox of equation-free functions in Matlab /Octave for efficient system level simulation. Numerical Algorithms, 87(4):1729–1748, August 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01027-z.

Marcolongo:2021:QOS

[MBTB21] Aris Marcolongo, Riccardo Bertossa, Davide Tisi, and Stefano Baroni. QEHeat: an open-source energy flux calculator for the computation of heat-transport coefficients from first principles. Computer Physics Communications, 269(??):Article 108090, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521002022.

Motl:1991:UMG

[MC91] Mark B. Motl and Bart Childs. A user’s manual for GNU

Emacs’ web-mode. ftp.cs.tamu.edu:/pub/tex-web/web/

docs, Texas A&M University, College Station, TX, USA,

McArthur:1985:RRE

[McA85] David McArthur. Running ROSS in an Emacs environment. The Rand paper series P-7088, Rand Corporation, Santa Monica, CA, USA, April 1985. 10 pp.

McAndrew:2008:TCO

[McA08] Alasdair McAndrew. Teaching cryptography with open-source software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):325–329, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

McAffer:2019:GSO

[McA19] J. McAffer. Getting started with open source governance. Computer, 52(10):92–96, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McCarty:1999:LDG

[McC99a] Bill McCarty. Learning Debian GNU/Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-705-2. xiii + 343 pp. LCCN QA76.76.O63

M372 1999. US$34.95. URL http://www.oreilly.com/
catalog/debian.

McCarty:1999:LRL

[McC99b] Bill McCarty. Learning Red Hat Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-627-7. xiii + 378 pp. LCCN QA76.76.O63 M379

1999. US$34.95. URL http://www.oreilly.com/catalog/
9781565926271;   http://www.oreilly.com/catalog/redhat.

McConnell:1999:EOS

[McC99c] Steve McConnell. From the editor: Open-source methodology: Ready for prime time? IEEE Software, 16(4):6–8, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print),

0740-7459 (electronic). URL http://dlib.computer.org/
so/books/so1999/pdf/s4006.pdf.

McCarty:2002:LRL

[McC02a] Bill McCarty. Learning Red Hat Linux: a Guide to Red Hat Linux for New Users. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002.

ISBN 0-596-00071-5. xx + 346 pp. LCCN QA76.76.O63

M376 2002. US$34.95. URL http://safari.oreilly.com/
0596000715;   http://www.oreilly.com/catalog/9780596000714;
 http://www.oreilly.com/catalog/redhat2.

McClellen:2002:WDD

[McC02b] Chris McClellen. Weather data distribution and system monitoring: An open-source ORB for a rainy day. Dr. Dobb’s Journal of Software Tools, 27(1):69–70, 72–74, January 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.

com/ftp/2002/2002_01/weather.txt.

McCarty:2003:LRL

[McC03] Bill McCarty. Learning Red Hat Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2003. ISBN 0-596-00469-9. xvi + 319 pp. LCCN

QA76.76.O63 M376 2003. URL http://www.oreilly.com/
catalog/9780596004699.

McCarty:2004:LRE

[McC04] Bill McCarty. Learning Red Hat Enterprise Linux and Fedora. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2004. ISBN 0-596-00589-X. xvii + 326 pp. LCCN QA76.76.O63 M376 2004. URL http://

www.oreilly.com/catalog/9780596005894.

McCarty:2005:SNO

[McC05] Bill McCarty. SELINUX: NSA’s open source Security Enhanced Linux. O’Reilly Media, Sebastopol, CA, USA, 2005. ISBN 0-596-00716-7. xiii + 238 pp. LCCN

TK5105.59 .M37 2005. URL http://www.oreilly.com/
catalog/9780596007164.

McGowan:2001:LIO

[McG01] David McGowan. Legal implications of Open-Source software. University of Illinois Law Review, 2001(1):241–300,

2001. URL http://www.law.umn.edu/uploads/images/
254/McGowanD-OpenSourceFinal.pdf.

Milewicz:2022:SFO

[MCGA22] Reed Milewicz, Jeffrey Carver, Samuel Grayson, and Travis Atkison. A secure future for open-source computational science and engineering. Computing in Science and Engineering, 24(4):65–69, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

McHugh:1992:EBD

[McH92] J. McHugh. An EMACS based downgrader for the SAT. In IEEE [IEE92c], pages 228–237. ISBN 0-8186-3115-5 (paperback), 0-8186-3116-3 (microfiche), 0-8186-3117-1 (casebound). LCCN QA76.9.A25 C6375 1992. URL http://www.cl.cam.

ac.uk/~fapp2/steganography/bibliography/1013.html.

Reprinted in ‘Computer and Network Security’.

McLauchlan:1992:HLV

[McL92] P. F. McLauchlan. HORATIO: libraries for vision applications. Technical Report OUEL 1967/92, Oxford University, Oxford, UK, December 1992. 72 pp.

McLaughlin:2005:ISP

[McL05] Laurianne McLaughlin. Inside the software patents debate: Some good news for Open Source developers. IEEE Software, 22(3):102–105, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://csdl.

computer.org/comp/mags/so/2005/03/s3102.pdf.

Muller:2021:PPO

[MCQF21] Heiko Müller, Sonia Castelo, Munaf Qazi, and Juliana Freire. From papers to practice: the openclean open-source data cleaning library. Proceedings of the VLDB Endowment, 14 (12):2763–2766, July 2021. CODEN ???? ISSN 2150-

8097. URL https://dl.acm.org/doi/10.14778/3476311.
3476339.

Mazzia:2012:TSS

[MCS12] Francesca Mazzia, Jeff R. Cash, and Karline Soetaert. A test set for stiff initial value problem solvers in the open source software R: Package deTestSet. Journal of Computational and Applied Mathematics, 236(16):4119–4131, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-

1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042712001331.

Madanmohan:2004:OSR

[MD04] T. R. Madanmohan and Rahul De’. Open Source reuse in commercial firms. IEEE Software, 21(6):62–69, November/ December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mansfield-Devine:2017:OSS

[MD17] Steve Mansfield-Devine. Open source software: determining the real risk posed by vulnerabilities. Network Security, 2017(1):7–12, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1353485817300053.

Mansfield-Devine:2018:OSI

[MD18] Steve Mansfield-Devine. Open source and the Internet of Things. Network Security, 2018(2):14–19, February 2018. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1353485818300163.

Moy:2022:PCG

[MD22] Yannick Moy and Claire Dross. Proving the correctness of GNAT light runtime library. ACM SIGADA Ada Letters, 42 (1):65–67, June 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://dl.acm.org/

doi/10.1145/3577949.3577959.

Morelli:2009:FEI

[MdL09] Ralph Morelli and Trishan de Lanerolle. Foss 101: engaging introductory students in the open source movement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):311–315, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Mehrotra:2018:OSR

[MDRN18] Pavan Mehrotra, Sabar Dasgupta, Samantha Robertson, and Paul Nuyujukian. An open-source realtime computational platform (short WIP paper). ACM SIGPLAN Notices, 53 (6):109–112, June 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Murray:2020:SVO

[MEB^{+}20] K. E. Murray, M. A. Elgammal, V. Betz, T. Ansell, K. Rothman, and A. Comodi. SymbiFlow and VPR: An open-source design flow for commercial and novel FPGAs. IEEE Micro, 40 (4):49–57, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mecklenburg:2005:MPG

[Mec05] Robert Mecklenburg. Managing Projects with GNU Make. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2005. ISBN 0-596-00610-1. xviii + 280 pp. LCCN QA76.76.U84 O73 2005. US$29.95.

Meeker:2012:FSO

[Mee12] Heather J. Meeker. Free Software And Open Source, chapter 2, pages 11–26. Wiley, New York, NY, USA, 2012. ISBN 1-119-19770-8.

Meijers:1992:OEB

[Mei92] F. Meijers. The OPAL event builder; practical experience with C++ in data acquisition. In Verkerk and Wojcik [VW92], pages 180–183.

Mengesha:2010:RTF

[Men10] Nigussie Tadesse Mengesha. The role of technological frames of key groups in open source software implementation in a developing country context. The Electronic Journal of Information Systems in Developing Countries, 43(1):1–19, 2010. ISSN 1681-4835.

Mendell:2012:BRP

[Men12] Matt Mendell. Book review: A practical guide to Fedora and Red Hat enterprise Linux by Mark G. Sobell. ACM SIGSOFT Software Engineering Notes, 37(1):36, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Merrill:2003:GGN

[Mer03] Jason Merrill. GENERIC and GIMPLE: a new tree representation for entire functions. In Hutton et al. [HDR03], pages 171–193. ISBN ???? LCCN ???? URL http://www.linux.

org.uk/~ajh/gcc/gccsummit-2003-proceedings.pdf.

Meyer:2018:UUL

[Mey18] Jan Max Meyer. UniCC: A universal LALR(1) Parser Generator User Manual. Phorward Software Technologies, Castroper Str. 65, 44357 Dortmund, Germany, 2018. URL https:

//phorward.info/products/unicc/unicc.pdf.

Martinez:2023:BOS

[MFB23] Emanuel A. Martínez, Juan I. Beltrán Fínez, and Flavio Y. Bruno. BinPo: an open-source code to compute the band structure of two-dimensional electron systems. Computer Physics Communications, 284(??):Article 108595, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522003149.

Mockus:2002:TCS

[MFH02] Audris Mockus, Roy T. Fielding, and James D. Herbsleb. Two case studies of open source software development: Apache and Mozilla. ACM Transactions on Software Engineering and Methodology, 11(3):309–346, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Michlmayr:2015:WHS

[MFS15] Martin Michlmayr, Brian Fitzgerald, and Klaas-Jan Stol. Why and how should open source projects adopt time-based releases? IEEE Software, 32(2):55–63, March/April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (elec-

tronic). URL http://www.computer.org/csdl/mags/so/
2015/02/mso2015020055-abs.html.

Muller:1994:ICS

[MG94] Hausi A. Muller and Mari Georges, editors. International Conference on Software Maintenance: proceedings, September 19–23, 1994, Victoria, British Columbia, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6330-8. LCCN QA76.76.S64I58 1994.

Matzon:2005:GGP

[MG05] Mads Matzon and Kristian Gøtrik. GNU General Public Licenses (GPL) samspil med dansk ophavs- og aftaleret. (Danish) [GNU General Public Licenses (GPL) cooperation with Danish author law and contract law]. Videnskabsbutikken, Københavns Universitet, Copenhagen, Denmark, 2005. ISBN 87-91337-69-0. 61 pp. LCCN ????

M:2012:ESG

[MG12] Krishna Raj P. M. and Srinivasa K. G. Empirical studies of global volunteer collaboration in the development of free and open source software: analysis of six top ranked projects in sourceforge.net. ACM SIGSOFT Software Engineering Notes, 37(2):1–11, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mostaco-Guidolin:2012:SOS

[MGFRG12] Luiz C. Mostaço-Guidolin, Rafael B. Frigori, Leonid Ruchko, and Ricardo M. O. Galvão. SCTE: an open-source Perl framework for testing equipment control and data acquisition. Computer Physics Communications, 183(7):1511–1518, July 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465512000665.

Miranda:2002:HUG

[MGM^{+}02] Javier Miranda, Francisco Guerra, Ernestina Martel, José Martín, and Alexis González. How to use GNAT to efficiently preprocess new Ada sentences. Lecture Notes in Computer Science, 2361:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2361/
23610179.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2361/23610179.pdf.

Martiradonna:2020:ASO

[MGPB20] Sergio Martiradonna, Alessandro Grassi, Giuseppe Piro, and Gennaro Boggia. 5G-air-simulator: an open-source tool modeling the 5G air interface. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):Article 107151, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128619317359.

Municchi:2016:HES

[MGR16] Federico Municchi, Christoph Goniva, and Stefan Radl. Highly efficient spatial data filtering in parallel using the opensource library CPPPO. Computer Physics Communications, 207(??):400–414, October 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465516301606.

Miller:2018:IPR

[MGYC18] Daniel Bradford Miller, William Bradley Glisson, Mark Yampolskiy, and Kim-Kwang Raymond Choo. Identifying 3D printer residual data via open-source documentation. Computers & Security, 75(??):10–23, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404818300324.

Hall:2000:MLF

[mH00] Jon “maddog” Hall. My life and free software. Linux Journal, 74:??, June 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue74/4047.html.

Martin:2007:OSA

[MH07] Ken Martin and Bill Hoffman. An open source approach to developing software in a small organization. IEEE Software, 24 (1):46–53, January/February 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Muller:1994:OSR

[MHP94] G. Muller, M. Hue, and N. Peyrouze. Operating system: results of the FTM experiment. In Echtle et al. [EHP94], pages 491–508. ISBN 3-540-58426-9. LCCN QA76.9.F38 E33 1994.

Mirtchovski:2007:WSD

[MI07] Andrey Mirtchovski and Latchesar Ionkov. Why some dead OSes still matter. ;login: the USENIX Association newsletter, 32(5):5–12, October 2007. CODEN LOGNEM. ISSN

1044-6397. URL https://www.usenix.org/publications/
login/october-2007-volume-32-number-5/why-some-dead-
oses-still-matter.

Michaelson:2004:TNS

[Mic04] Jay Michaelson. There’s no such thing as a free (software) lunch. ACM Queue: Tomorrow’s Computing Today, 2(3):40– 47, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Milkowski:2010:DOS

[Mił10] Marcin Miłkowski. Developing an open-source, rule-based proofreading tool. Software—Practice and Experience, 40(7):

543–566, June 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Miola:1990:DIS

[Mio90] A. Miola, editor. Design and Implementation of Symbolic Computation Systems. International Symposium DISCO ’90. Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-540-52531-9. LCCN QA76.9.S88I576 1990.

Miranda:2003:GTG

[Mir03] Javier Miranda. A guided tour to the GNAT runtime. Ada User Journal, 24(1):55–??, March 2003. CODEN AUJOET. ISSN 1381-6551.

Miranda:2007:TCO

[Mir07] J. Miranda. Towards certification of object-oriented code with the GNAT compiler. Ada User Journal, 28(3):178–??, September 2007. CODEN AUJOET. ISSN 1381-6551.

Mitchell:1984:ISU

[Mit84] William H. Mitchell. An Icon subsystem for UNIX emacs. Technical report TR 84-8, Dept. of Computer Science, The University of Arizona, Tucson, AZ, USA, 1984. 17 pp.

Mittag:1994:UGC

[Mit94] L. Mittag. Using GNU C for cross-development. In Anonymous [Ano94b], pages 45–50. ISBN 0-87930-355-7 (vol. 1), 0-87930-356-5 (vol. 2). LCCN ???? Two volumes.

Mittag:1995:CDG

[Mit95] L. Mittag. Cross debugging with GNU GDB. Embedded Systems Programming, 8(8):28–30, 32, August 1995. CODEN EYPRE4. ISSN 1040-3272.

Mobius:2012:OOS

[MK12] Jan Möbius and Leif Kobbelt. OpenFlipper: an open source geometry processing and rendering framework. Lecture Notes in Computer Science, 6920:488–500, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/978-3-642-27413-8_31.

Medeiros:2019:IMC

[MLA^{+}19] Flávio Medeiros, Gabriel Lima, Guilherme Amaral, Sven Apel, Christian Kästner, Márcio Ribeiro, and Rohit Gheyi. An investigation of misunderstanding code patterns in C open-source software projects. Empirical Software Engineering, 24(4):1693–1726, August 2019. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://link.

springer.com/article/10.1007/s10664-018-9666-x.

Martinez-Lopez:2015:OSS

[MLMFN^{+}15] Javier Martínez-López, Julia Martínez-Fernández, Babak Naimi, María F. Carreño, and Miguel A. Esteve. An open-source spatio-dynamic wetland model of plant community responses to hydrological pressures. Ecological Modelling, 306(??):326–333, June 24, 2015. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380014005973.

Munir:2018:OIU

[MLWR18] Hussan Munir, Johan Linåker, Krzysztof Wnuk, and Per Runeson. Open innovation using open source tools: a case study at Sony Mobile. Empirical Software Engineering, 23 (1):186–223, February 2018. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://

link.springer.com/article/10.1007/s10664-017-9511-
7;   http://link.springer.com/content/pdf/10.1007/
s10664-017-9511-7.pdf.

Mao:2023:DDL

[MLZ^{+}23] Runze Mao, Minqi Lin, Yan Zhang, Tianhan Zhang, Zhi-Qin John Xu, and Zhi X. Chen. DeepFlame: a deep learning empowered open-source platform for reacting flow simulations. Computer Physics Communications, 291(??):Article 108842, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552300187X.

Moodley:2004:CMP

[MM04] K. Moodley and H. Murrell. A colour-map plugin for the open source, Java based, image processing package, ImageJ.

Computers and Geosciences, 30(6):609–618, 2004. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Mucci:2010:OSP

[MM10] Philip J. Mucci and Tushar Mohan. An Open Source performance tools software suite for scientific computing. Concurrency and Computation: Practice and Experience, 22(2): 206–216, February 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Minsky:1990:SP

[MMB^{+}90] M. Minsky, J. McCarthy, R. S. Boyer, R. Stallman, and S. Garfinkel. Software patents. Dr. Dobb’s Journal of Software Tools, 15(11):56, 58, 62, 65–67, 70–73, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Milovanovic:2012:IFM

[MMD12] M. Milovanović, M. Minović, and S. Dusan. Interoperability framework for multimodal biometry: Open source in action. J.UCS: Journal of Universal Computer Science, 18(11):1558–??, ???? 2012. CODEN ????

ISSN 0948-6968. URL http://www.jucs.org/jucs_18_11/
interoperability_framework_for_multimodal.

Mallasen:2022:POSa

[MMD^{+}22] David Mallasén, Raul Murillo, Alberto A. Del Barrio, Guillermo Botella, Luis Piñuel, and Manuel Prieto-Matias. PERCIVAL: Open-source posit RISC-V core with quire capability. IEEE Transactions on Emerging Topics in Computing, 10(3):1241–1252, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mandanici:2022:SPP

[MMP^{+}22] Andrea Mandanici, Giuseppe Mandaglio, Giovanni Pirrotta, Valeria Conti Nibali, and Giacomo Fiumara. Simple physics with Python: a workbook on introductory physics with opensource software. Computing in Science and Engineering, 24 (2):74–78, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mashayekhi:1995:URA

[MMR95] V. Mashayekhi, M. Maley, and J. Riedl. User recovery of audio operations. In IEEE [IEE95a], pages 329–334. ISBN 0-8186-7105-X. LCCN QA76.575.I623 1995. IEEE Catalog No. 95TH8066.

Misawa:2019:MOS

[MMY^{+}19] Takahiro Misawa, Satoshi Morita, Kazuyoshi Yoshimi, Mitsuaki Kawamura, Yuichi Motoyama, Kota Ido, Takahiro Ohgoe, Masatoshi Imada, and Takeo Kato. mVMC-open-source software for many-variable variational Monte Carlo method. Computer Physics Communications, 235(??):447–462, February 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465518303102.

Mitasova:2004:GOS

[MN04] Helena Mitasova and Markus Neteler. GRASS as open source free software GIS: Accomplishments and perspectives. Transactions in GIS, 8(2):145–154, 2004. ISSN 1361-1682 (print), 1467-9671 (electronic).

Mohankumar:2021:IOR

[MN21] N. Mohankumar and A. Natarajan. InvFD, an OCTAVE routine for the numerical inversion of the Fermi–Dirac integral. Computer Physics Communications, 267(??):Article 108062, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001740.

Mengerink:2019:EOR

[MNS19] Josh G. M. Mengerink, Jeroen Noten, and Alexander Serebrenik. Empowering OCL research: a large-scale corpus of open-source data from GitHub. Empirical Software Engineering, 24(3):1574–1609, June 2019. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://

link.springer.com/article/10.1007/s10664-018-9641-
6;   http://link.springer.com/content/pdf/10.1007/
s10664-018-9641-6.pdf.

Moffitt:2002:GAE

[Mof02] Nick Moffitt. GAR: Automating entire OS builds. Embedded Linux Journal, 9:38–41, May/June 2002. CODEN ???? ISSN

1534-083X. URL http://embedded.linuxjournal.com/
magazine/issue09/;   http://gar.lnx-bbc.org/wiki/GarChitecture?
action=print.

Moglen:1999:ATF

[Mog99] Eben Moglen. Anarchism triumphant: Free Software and the death of copyright. World-Wide Web document.,

1999. URL http://emoglen.law.columbia.edu/my_pubs/
anarchism.html.

Moglen:2001:FSMa

[Mog01a] Eben Moglen. Free Software matters: Enforcing the GPL, I. World-Wide Web document., August 12, 2001. URL http:/

/emoglen.law.columbia.edu/publications/lu-12.html.

Moglen:2001:FSMb

[Mog01b] Eben Moglen. Free Software matters: Enforcing the GPL, II. World-Wide Web document., September 10, 2001. URL http:

//emoglen.law.columbia.edu/publications/lu-13.html.

Moglen:2001:FSMc

[Mog01c] Eben Moglen. Free software matters: Free software or open source? World-Wide Web document., February 15, 2001.

URL http://emoglen.law.columbia.edu/publications/
lu-07.html.

Moglen:2003:QSH

[Mog03a] Eben Moglen. Questioning SCO: a hard look at nebulous claims. World-Wide Web document., August 1, 2003.

URL http://www.gnu.org/philosophy/sco/questioning-
sco.html.

Moglen:2003:SSS

[Mog03b] Eben Moglen. SCO scuttles sense, claiming GPL invalidity. World-Wide Web document., August 18, 2003. URL http:/

/www.fsf.org/licensing/sco/sco-preemption.html.

Moglen:2003:SFR

[Mog03c] Eben Moglen. SCO: Without fear and without research. World-Wide Web document., November 24, 2003. URL http:

//www.gnu.org/philosophy/sco/sco-without-fear.html.

Molino:2001:SLR

[Mol01] Donato Molino. Il software libero in riferimento alle recenti disposizioni legislative sul diritto d’autore. (Italian) [Free Software in the light of recent legislative dispositions on the rights of authors]. World-Wide Web document., July

2001. URL http://www.prosa.it/philosophy/opinioni/
donato.shtml.

Monden:2011:GGU

[MOMM11] Akito Monden, Satoshi Okahara, Yuki Manabe, and Kenichi Matsumoto. Guilty or not guilty: Using clone metrics to determine open source licensing violations. IEEE Software, 28 (2):42–47, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Moncur:2003:MYV

[Mon03] Michael G. Moncur. MySQL: your visual blueprint to open source database management. Visual read less, learn more. Wiley, New York, NY, USA, 2003. ISBN 0-7645-1692-

2. xv + 290 pp. LCCN QA76.3.S67 M66 2003. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/bios/wiley045/2002110260.
html;   http://www.loc.gov/catdir/description/wiley039/
2002110260.html;   http://www.loc.gov/catdir/toc/wiley031/
2002110260.html.

Moody:2001:RCL

[Moo01a] Glyn Moody. Rebel code: Linux and the Open Source Revolution. Allen Lane, London, UK and New York, NY, USA, 2001. ISBN 0-7139-9520-3. viii + 334 pp. LCCN QA76.76.O63 M645 2001.

Moody:2001:RCI

[Moo01b] Glyn Moody. The rebel code: the inside story of Linux and the open source revolution. Perseus Publishers, Cambridge, MA, USA, 2001. ISBN 0-7382-0333-5. viii + 334 pp. LCCN QA76.76.O63 M663 2001. US$27.50.

Moore:2003:R

[Moo03] J. T. S. Moore. Revolution OS. Documentary film (85 minutes) from Wonderview Productions., September 30,

2003. URL http://revolution-os.com/;   http://www.
amazon.com/Revolution-Os-Linus-Torvalds/dp/B0000A9GLO.

MKS:awk

[Mor87] Mortice Kern Systems, Inc. MKSAWK, 1987. 35 King Street North, Waterloo, Ontario, Canada, Tel: (519) 884-2251. See also [AKW88].

Morris:1991:CPC

[Mor91] W. G. Morris. CCG: a prototype coagulating code generator. ACM SIGPLAN Notices, 26(6):45–58, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moreland:1992:CMS

[Mor92] Carl Moreland. CUG372 — Mouse++, String++, and Z++ classes. C Users Journal, 10(12):122–??, December 1992. ISSN 0898-9788.

Morin:1996:MUV

[Mor96] Richard Morin. Mach-based UNIX variants. SunExpert Magazine, 7(9):30, 32, 33, September 1996. ISSN 1053-9239. Discusses the University of Utah’s Mach 4 project (see http://www.cs.utah.edu/projects/flux/mach4/html/) the GNU Hurd project (see http://www.cs.pdx.edu/ trent/gnu/hurd/) and the commercial Mach Ten implementation that runs Mach on top of MacOS (see http://www.tenon.com/).

Morozov:2008:OSS

[Mor08] Igor Morozov. Open-source software framework integrates data analysis. Eos, Transactions American Geophysical Union, 89(29):261–262, 2008. ISSN 0096-3941 (print), 2324-9250 (electronic).

Morgan:2011:UCC

[Mor11] Cliff Morgan. Understanding the Creative Commons licence. Learned Publishing, 24(1):51–53, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Moses:2012:MPH

[Mos12] Joel Moses. Macsyma: a personal history. Journal of Symbolic Computation, 47(2):123–130, February 2012. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0747717110001483.

Murgia:2018:EQQ

[MOT^{+}18] Alessandro Murgia, Marco Ortu, Parastou Tourani, Bram Adams, and Serge Demeyer. An exploratory qualitative and quantitative analysis of emotions in issue report comments of open source systems. Empirical Software Engineering, 23(1): 521–564, February 2018. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://link.

springer.com/article/10.1007/s10664-017-9526-0.

Montenbruck:2000:APC

[MP00] Oliver Montenbruck and Thomas Pfleger. Astronomy on the personal computer. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67221-4. xv + 310 pp. LCCN QB51.3.E43 M6613 2000. Translated by Storm Dunlop, with a foreword by Richard M. West.

Midha:2012:FAS

[MP12] Vishal Midha and Prashant Palvia. Factors affecting the success of Open Source Software. The Journal of Systems and Software, 85(4):895–905, April 2012. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016412121100286X.

Mittag:2011:CAO

[MPE^{+}11] Anja Mittag, Fernanda E. Pinto, Denise C. Endringer, Attila Tarnok, and Dominik Lenz. Cellular analysis by opensource software for affordable cytometry. Scanning, 33(1): 33–40, 2011. ISSN 0161-0457 (print), 1932-8745 (electronic).

Madec:2016:GOS

[MPG^{+}16] Morgan Madec, François Pecheux, Yves Gendrault, Elise Rosati, Christophe Lallement, and Jacques Haiech. GeNeDA: An open-source workflow for design automation of gene regulatory networks inspired from microelectronics. Journal of Computational Biology, 23(10):841–855, October 2016. CO-DEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.liebertpub.com/doi/abs/10.
1089/cmb.2015.0229;   https://www.liebertpub.com/doi/
pdf/10.1089/cmb.2015.0229.

Muzammal:2019:RBD

[MQN19] Muhammad Muzammal, Qiang Qu, and Bulat Nasrulin. Renovating blockchain with distributed databases: an open source system. Future Generation Computer Systems, 90 (??):105–117, January 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167739X18308732.

Maruyama:1994:COL

[MR94] K. Maruyama and N. Raguideau. Concurrent object-oriented language ’COOL’. ACM SIGPLAN Notices, 29(9):105–114, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Melo:2020:TIC

[MRGP20] Leandro T. C. Melo, Rodrigo G. Ribeiro, Breno C. F. Guimarães, and Fernando Magno Quintão Pereira. Type inference for C: Applications to the static analysis of incomplete programs. ACM Transactions on Programming Languages and Systems, 42(3):15:1–15:71, December 2020. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3421472.

Melquiond:2023:WFV

[MRH23] Guillaume Melquiond and Raphaël Rieu-Helft. WhyMP, a formally verified arbitrary-precision integer library. Journal of Symbolic Computation, 115(??):74–95, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-

855X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0747717122000657.

Muta:2020:EOS

[MRN20] Abhinav Muta, Prabhu Ramachandran, and Pawan Negi. An efficient, open source, iterative ISPH scheme. Computer Physics Communications, 255(??):Article 107283, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300953.

Moser:2007:EAC

[MRS07] Raimund Moser, Barbara Russo, and Giancarlo Succi. Empirical analysis on the correlation between GCC compiler warnings and revision numbers of source files in five industrial software projects. Empirical Software Engineering, 12(3):295– 310, June 2007. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10664-006-9029-x.

Milutinovic:1991:PTA

[MS91] V. Milutinovic and B. D. Shriver, editors. Proceedings of the Twenty-Fourth Annual Hawaii International Conference on System Sciences (Cat. No.91TH0350-9). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. 4 vol.

Mitchell:2000:GSS

[MS00] Mark Mitchell and Alexander Samuel. GCC 3.0: The state of the source. In USENIX [USE00a], page ?? ISBN 1-880446-17-0. LCCN ???? URL http://www.usenix.org/

publications/library/proceedings/als2000/mitchell.
html.

Maginnis:2002:SLG

[MS02] Tobin Maginnis and Sair Development Team. Sair Linux and GNU certification: level II, Samba and resource sharing. Wiley, New York, NY, USA, 2002. ISBN 0-471-40535-3. ???? pp. LCCN QA76.3 .M3237 2002.

Matloff:2008:ADG

[MS08] Norman S. Matloff and Peter Jay Salzman. The art of debugging with GDB, DDD, and Eclipse. No Starch Press, San Francisco, CA, USA, 2008. ISBN 1-59327-174-3 (paperback), 1-59327-002-X. xiv + 264 pp. LCCN QA76.9.D43 M35 2008.

Moreno-Sanchez:2012:FOS

[MS12] Rafael Moreno-Sanchez. Free and Open Source Software for Geospatial Applications (FOSS4G): a mature alternative in the geospatial technologies arena. Transactions in GIS, 16(2): 81–88, 2012. ISSN 1361-1682 (print), 1467-9671 (electronic).

Monnier:2020:EEL

[MS20] Stefan Monnier and Michael Sperber. Evolution of Emacs Lisp. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):74:1–74:55, June 2020. URL https:/

/dl.acm.org/doi/abs/10.1145/3386324.

Martin-Samos:2009:SOS

[MSB09] Layla Martin-Samos and Giovanni Bussi. SaX: An open source package for electronic-structure and optical-properties calculations in the GW approximation. Computer Physics Communications, 180(8):1416–1425, August 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000496.

Mackita:2019:ERM

[MSC19] Masky Mackita, Soo-Young Shin, and Tae-Young Choe. ER-MOCTAVE: a risk management framework for IT systems which adopt cloud computing. Future Internet, 11(9):195, September 10, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/11/9/195.

Miranda:2005:IAS

[MSK05] Javier Miranda, Edmond Schonberg, and Hristian Kirtchev. The implementation of Ada 2005 synchronized interfaces in the GNAT compiler. ACM SIGADA Ada Letters, 25(4):41–48, December 2005. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Matsen:1971:CST

[MSLH71] F. A. Matsen, Harrison Shull, Peter Lykos, and Frank Harris. Computational support for theoretical chemistry: report of a conference held at the National Institutes of Health Bethesda, Maryland, May 8–9, 1970. Report, National Academy of Sciences, Washington, DC, USA, 1971. vii + 51 pp. URL https:

//books.google.com/books?id=ZUArAAAAYAAJ&pg=PR3. See

also the sequel report [WBB^{+}74].

Miranda:2003:DCP

[MSM^{+}03] Javier Miranda, Edmond Schonberg, Miguel Masmano, Jorge Real, and Alfons Crespo. Dynamic ceiling priorities in GNAT implementation report. ACM SIGADA Ada Letters, 23(4): 24–27, December 2003. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mills:2010:XOS

[MSM10] Curt E. Mills, Steve Stroh, and Laura Shaffer Mills. Xastir: open-source client for the automatic packet reporting system.

Linux Journal, 2010(189):2:1–2:??, January 2010. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Milutinovic:1991:PTH

[MSNS91] V. Milutinovic, B. D. Shriver, J. F. Nunamaker, Jr., and R. H. Sprague, Jr., editors. Proceedings of the Twenty-Fifth Hawaii International Conference on System Sciences (Cat. No.91TH0394-7). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2420-5. LCCN ???? 4 vol.

Muller-Seitz:2009:OSS

[MSR09] Gordon Müller-Seitz and Guido Reger. Is open source software living up to its promises? Insights for open innovation management from two open source software-inspired projects. R&D Management, 39(4):372–381, 2009. ISSN 0033-6807 (print), 1467-9310 (electronic).

Muller-Seitz:2010:NBS

[MSR10] Gordon Müller-Seitz and Guido Reger. Networking beyond the software code? An explorative examination of the development of an open source car project. Technovation, 30(11–12):627–634, November/December 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721000088X.

Miller:1995:UGC

[MSS95] J. W. V. Miller, M. Shridhar, and B. N. Shabestari. Using GNU C to develop PC-based vision systems [2597-31]. In Batchelor et al. [BSW95], pages 253–258. ISBN 0-8194-1961-3. ISSN 0361-0748. LCCN TS510.S63 v.2597.

Maillart:2008:ETZ

[MSSvK08] T. Maillart, D. Sornette, S. Spaeth, and G. von Krogh. Empirical tests of Zipf’s Law mechanism in open source Linux distribution. Physical Review Letters, 101(21):218701, November 19, 2008. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.

org/doi/10.1103/PhysRevLett.101.218701.

Meike:2009:ISS

[MSW09] Michael Meike, Johannes Sametinger, and Andreas Wiesauer. Internet security: Security in open source Web content management systems. IEEE Security & Privacy, 7(4):44–51, July/ August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mangalam:2001:GOS

[MSZ^{+}01] H. Mangalam, J. Stewart, J. Zhou, K. Schlauch, M. Waugh, G. Chen, A. D. Farmer, G. Colello, and J. W. Weller. GeneX: An Open Source gene expression database and integrated tool set. IBM Systems Journal, 40(2):552–569, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://

www.research.ibm.com/journal/sj/402/mangalam.html;
 http://www.research.ibm.com/journal/sj/402/mangalam.
pdf;   http://www.research.ibm.com/journal/sj/402/
mangalam.txt.

Mowery:2002:LPI

[MSZ02] David C. Mowery, Bhaven N. Sampat, and Arvids A. Ziedonis. Learning to patent: Institutional experience, learning, and the characteristics of U.S. university patents after the Bayh–Dole Act, 1981–1992. Management Science, 48(1):73– 89, January 2002. CODEN MSCIAM. ISSN 0025-1909 (print),

1526-5501 (electronic). URL http://pubsonline.informs.
org/doi/pdf/10.1287/mnsc.48.1.73.14278.

Maskit:1994:MPS

[MT94] D. Maskit and S. Taylor. A message-driven programming system for fine-grain multicomputers. Software—Practice and Experience, 24(10):953–980, October 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Megias:2009:FTA

[MTBS09] David Megias, Wouter Tebbens, Lex Bijlsma, and Francesc Santanach. Free technology academy: a European initiative for distance education about free software and open standards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):70–74, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Morelli:2009:RCE

[MTD^{+}09] Ralph Morelli, Allen Tucker, Norman Danner, Trishan R. De Lanerolle, Heidi J. C. Ellis, Ozgur Izmirli, Danny Krizanc, and Gary Parker. Revitalizing computing education through free and open source software for humanity. Communications of the ACM, 52(8):67–75, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marsan:2019:TSS

[MTM^{+}19] J. Marsan, M. Templier, P. Marois, B. Adams, K. Carillo, and G. L. Mopenza. Toward solving social and technical problems in open source software ecosystems: Using causeand-effect analysis to disentangle the causes of complex problems. IEEE Software, 36(1):34–41, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mudawwar:1997:MTM

[Mud97] Muhammed F. Mudawwar. Multicode: a truly multilingual approach to text encoding: Unicode was designed to extend ASCII for encoding text in different languages, but it still have several important drawbacks. multicode overcomes those drawbacks. Computer, 30(4):37–43, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See also response and rebuttal [DM97] and letter [CPJ^{+}98].

Murdock:1994:ODG

[Mur94] Ian Murdock. Overview of the Debian GNU/Linux system. Linux Journal, 6:??, October 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

noframes.linuxjournal.com/lj-issues/issue6/debian.
html.

Murphy:2009:BT

[Mur09] Dan Murphy. The beginnings of TECO. IEEE Annals of the History of Computing, 31(4):110–115, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Murray:2020:OSS

[Mur20] Dale Murray. Open source and security: why transparency now equals strength. Network Security, 2020(7):17–19, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-

9371 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1353485820300829.

Muwanguzi:2009:AOS

[Muw09] Samuel Muwanguzi. Adoption of open source software (OSS) for Uganda: a social construction perspective. Proceedings of the American Society for Information Science and Technology, 46(1):1–4, 2009. ISSN 2373-9231.

Mannaert:2005:UOS

[MV05] Herwig Mannaert and Kris Ven. The use of open source software platforms by independent software vendors: issues and opportunities. ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Moratilla-Vega:2022:OSC

[MVAXP22] M. A. Moratilla-Vega, M. Angelino, H. Xia, and G. J. Page. An open-source coupled method for aeroacoustics modelling. Computer Physics Communications, 278(??):Article 108420, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001394.

Milasinovic:2020:DOS

[MVF20] Danko Z. Milasinovic, Arso M. Vukicevic, and Nenad D. Filipovic. dfemtoolz: an open-source C++ framework for efficient imposition of material and boundary conditions in finite element biomedical simulations. Computer Physics Communications, 249(??):Article 106996, April 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303376.

Mortensen:2015:OHL

[MVS15] Mikael Mortensen and Kristian Valen-Sendstad. Oasis: a high-level/high-performance open source Navier–Stokes solver. Computer Physics Communications, 188(??):177–188, March 2015. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465514003786.

Mangaser:1989:CPS

[MWB89] A. A. Mangaser, Y. Wang, and S. E. Butner. Concurrent programming support for a multimanipulator experiment on RIPS. In IEEE [IEE89], pages 853–859 (vol. 2). ISBN 0-8186-1938-4. LCCN TJ 210.3 I44 1989. Three volumes.

Muller:1990:PAU

[MWG^{+}90] H. Muller, J. Winckler, S. Grzybek, M. Otte, B. Stoll, F. Equoy, and N. Higelin. PASTIS-program animation using X. In Anonymous [Ano90a], pages 104–111.

Muller:1991:PAS

[MWG^{+}91] H. Muller, J. Winckler, S. Grzybek, M. Otte, B. Stoll, F. Equoy, and N. Higelin. The program animation system PASTIS. Journal of Visualization and Computer Animation, 2(1):26–33, January–March 1991. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Meneely:2008:RRE

[MWG08] Andrew Meneely, Laurie Williams, and Edward F. Gehringer. ROSE: a repository of education-friendly open-source projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):7–11, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Miura:1989:AKD

[MYU89] K. Miura, K. Yonezawa, and Y. Uesugi. Advanced Kanji display terminal. Anritsu Technical Bulletin, September 1989. CODEN ANTKAE. ISSN 0003-5211.

Mikram:2013:PCP

[MZE13] J. Mikram, F. Zinoun, and A. El Abdllaoui. POINCARÉ CODE: a package of open-source implements for normalization and computer algebra reduction near equilibria of coupled ordinary differential equations. Computer Physics Communications, 184(9):2204–2213, September 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465513001380.

Mihajlovic:2014:DIQ

[MZG14] Bojan Mihajlović, Zeljko Zilić, and Warren J. Gross. Dynamically instrumenting the QEMU emulator for Linux process trace generation with the GDB debugger. ACM Transactions on Embedded Computing Systems, 13(5s):167:1–167:??, November 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Miller:2022:RCF

[MZH22] Barton P. Miller, Mengxiao Zhang, and Elisa R. Heymann. The relevance of classic fuzz testing: Have we solved this one? IEEE Transactions on Software Engineering, 48 (6):2028–2039, June 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL https://arxiv.

org/abs/2008.06537;   https://ieeexplore.ieee.org/
document/9309406.

Nagappan:2018:RWG

[Nag18] M. Nagappan. Reconsidering whether GOTO is harmful. IEEE Software, 35(3):93–95, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Naishlos:2004:AG

[Nai04] Dorit Naishlos. Autovectorization in GCC. In Hutton et al. [HDR04], pages 105–117. ISBN ???? LCCN

???? URL http://people.redhat.com/lockhart/.gcc04/
MasterGCC-2side.pdf.

Nash:2004:EGJ

[Nas04] Michael P. Nash. Explorer’s Guide to Java Open Source Tools. Manning Publications, Greenwich, CT, USA, 2004. ISBN 1-932394-19-2. 300 (est.) pp. US$39.95. URL http://www.

manning.com/catalog/view.php?book=nash.

Nguyen-Duc:2021:ASA

[NDDH^{+}21] Anh Nguyen-Duc, Manh Viet Do, Quan Luong Hong, Kiem Nguyen Khac, and Anh Nguyen Quang. On the adoption of static analysis for software security assessment — a case study of an open-source e-government project. Computers & Security, 111(??):Article 102470, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002947.

Negus:2015:GGP

[Neg15] Christopher Negus. GNU General Public License, pages 860– 872. Wiley, New York, NY, USA, 2015. ISBN 1-119-20953-6.

Neher:2004:CCI

[Neh04] Markus Neher. CoStLy: Complex interval Standard functions Library. Preprint 04/18, Fakultät für Mathematik, Universität Karlsruhe, Karlsruhe, Germany, 2004.

URL http://iamlasun8.mathematik.uni-karlsruhe.de/~
ae16/CoStLy.html.

Neher:2007:CSF

[Neh07] Markus Neher. Complex standard functions and their implementation in the CoStLy library. ACM Transactions on Mathematical Software, 33(1):2:1–2:27, March 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Nejmeh:2012:HFO

[Nej12] Brian A. Nejmeh. The humanitarian free and open-source software project: Engaging students in service-learning through building software. In Ralph Morelli, Trishan de Lanerolle, and Allen Tucker, editors, Service-Learning in the Computer and Information Sciences, chapter 5, pages 117– 136. Wiley, New York, NY, USA, 2012. ISBN 1-118-31913-3.

Neumann:1984:RHD

[Neu84] Peter G. Neumann. Review of The hacker’s dictionary: a guide to the world of computer wizards by Guy L. Steele, Donald R. Woods, Raphael A. Finkel, Mark R. Crispin, Richard M. Stallman, and Geoffrey S. Goodfellow. Harper and Row Publishers 1983. ACM SIGSOFT Software Engineering Notes, 9(1):12–15, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1999:IRRa

[Neu99] Peter G. Neumann. Inside risks: Robust open-source software. Communications of the ACM, 42(2):128, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
journals/cacm/1999-42-2/p128-neumann/.

Neumann:2000:GPR

[Neu00] Sven Neumann. Gimp Pocket Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-731-1. iii + 97 pp. LCCN T385 .N47 2000. US$9.95. Translated by Sven Riedel.

Novaresio:2012:OSL

[NGCI^{+}12] Valerio Novaresio, María García-Camprubí, Salvador Izquierdo, Pietro Asinari, and Norberto Fueyo. An open-source library for the numerical modeling of mass-transfer in solid oxide fuel cells. Computer Physics Communications, 183(1):125–146, January 2012. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046551100275X.

Newby:2003:OSS

[NGJ03] Gregory B. Newby, Jane Greenberg, and Paul Jones. Open source software development and Lotka’s Law: Bibliometric patterns in programming. Journal of the American Society for Information Science and Technology: JASIST, 54(2):169– 178, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See comments [Bur04a].

Nicol:1993:FLS

[Nic93] G. T. Nicol. Flex: The Lexical Scanner Generator, for Flex Version 2.3.7. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1993. ISBN 1-882114-21-3. 128 (est.) pp. LCCN

???? URL http://www.gnupress.org/book11.html.

Norris:2004:MCD

[NK04] Jeffrey S. Norris and Poul-Henning Kamp. Mission-critical development with Open Source software: Lessons learned. IEEE Software, 21(1):42–49, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nelson:2011:SEU

[NMG11] Adam Nelson, Tim Menzies, and Gregory Gay. Sharing experiments using open-source software. Software—Practice and Experience, 41(3):283–305, March 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nishiura:2014:PCP

[NMS14] Daisuke Nishiura, Miki Y. Matsuo, and Hide Sakaguchi. ppohDEM: Computational performance for open source code of the discrete element method. Computer Physics Communications, 185(5):1486–1495, May 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465514000472.

Ntantogian:2019:EPH

[NMX19] Christoforos Ntantogian, Stefanos Malliaros, and Christos Xenakis. Evaluation of password hashing schemes in open source web platforms. Computers & Security, 84(??):206–224, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818308332.

Nelson:2000:TCN

[NN00] Daniel Nelson and Yau Man Ng. Teaching computer networking using open source software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3): 13–16, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Neville-Neil:2016:GL

[NN16a] George Neville-Neil. GNL is not Linux. ACM Queue: Tomorrow’s Computing Today, 14(1):40, January 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/detail.cfm?id=2909572.

Neville-Neil:2016:KVG

[NN16b] George V. Neville-Neil. Kode vicious: GNL is not Linux. Communications of the ACM, 59(4):31–32, April 2016. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://cacm.acm.org/magazines/2016/4/
200171/fulltext.

Neville-Neil:2020:NPG

[NN20] George V. Neville-Neil. The non-psychopath’s guide to managing an open-source project: Respect your staff, learn from others, and know when to let go. QUEUE, 18(6):30–36,

December 2020. URL https://dl.acm.org/doi/10.1145/
3442632.3447645.

Neville-Neil:2021:KVN

[NN21] George V. Neville-Neil. Kode Vicious: The non-psychopath’s guide to managing an open source project. Communications of the ACM, 64(4):25–27, April 2021. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:

//dl.acm.org/doi/10.1145/3451229.

Niebuhr-Offermann:2003:AUS

[NO03] Tanja Niebuhr-Offermann. Der abgeschwächte urheberrechtliche Schutz von Software: Shareware, Public-Domain-Software, Open-Source Software. (German) [The weakened copyright protection of software: shareware, publicdomain-software, Open-Source software]. World-Wide Web document., June 2003. URL http://www.ifross.de/

Fremdartikel/SeminararbeitNiebuhr-Offermann.pdf.

Noble:2008:GMY

[Nob08] Michael S. Noble. Getting more from your multicore: exploiting OpenMP from an open-source numerical scripting language. Concurrency and Computation: Practice and Experience, 20(16):1877–1891, November 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Noji:2001:OSP

[Noj01] Kentaro Noji. Open source project for Unicode locales for Linux using Unicode databases, collation keys and a XML based locale data. In Unicode Consortium [Uni01], page ??

ISBN ???? LCCN ???? URL http://www.unicode.org/
iuc/iuc18/papers/a2-paper.pdf.

Noronha:2002:ILS

[Nor02] Frederick Noronha. Indian language solutions for GNU/Linux. Linux Journal, 103:??, November 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/www.linuxjournal.com/article.php?sid=6282.

Norouzi:2023:PPM

[Nor23] H. R. Norouzi. PhasicFlow: a parallel, multi-architecture open-source code for DEM simulations. Computer Physics Communications, 291(??):Article 108821, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001662.

Novillo:2003:TSN

[Nov03] Diego Novillo. Tree SSA: a new optimization infrastructure for GCC. In Hutton et al. [HDR03], pages 181–193. ISBN

???? LCCN ???? URL http://www.linux.org.uk/~ajh/
gcc/gccsummit-2003-proceedings.pdf.

Novillo:2004:DIT

[Nov04] Diego Novillo. Design and implementation of Tree SSA. In Hutton et al. [HDR04], pages 119–130. ISBN ???? LCCN

???? URL http://people.redhat.com/lockhart/.gcc04/
MasterGCC-2side.pdf.

Narduzzo:2003:MAG

[NR03] A. Narduzzo and A. Rossi. Modularity in action: GNU/ Linux and Free/Open Source software development model unleashed. World-Wide Web document., May 12, 2003. URL

http://opensource.mit.edu/papers/narduzzorossi.pdf.

Neal:1999:LSL

[NRG^{+}99] John Neal, Thiadmer Riemersma, Jeff Genender, Torpum Jannak, Richard A. Clarke, William C. Brown, Jeffrey Simmers, and Kemal Gencay. Letters: The Small language; Java 2 graphics; cross-platform independence; version control; median filters; Open Source in Turkey. Dr. Dobb’s Journal of Software Tools, 24(12):10, 12, December 1999. CODEN DDJOEB. ISSN 1044-789X.

Nafi:2020:UCL

[NRRS20] Kawser Wazed Nafi, Banani Roy, Chanchal K. Roy, and Kevin A. Schneider. A universal cross language software similarity detector for open source software categorization. The Journal of Systems and Software, 162(??):??, April 2020. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121219302651.

Neubauer:2001:ELD

[NS01] Matthias Neubauer and Michael Sperber. Down with Emacs Lisp: Dynamic scope analysis. ACM SIGPLAN Notices, 36 (10):38–49, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://cristal.inria.fr/ICFP2001/Abstracts/50.html.

Neus:2005:OMC

[NS05] A. Neus and P. Scherf. Opening minds: Cultural change with the introduction of open-source collaboration methods.

IBM Systems Journal, 44(2):215–225, ???? 2005. CODEN

IBMSA7. ISSN 0018-8670. URL http://www.research.ibm.
com/journal/sj/442/neus.pdf.

Nichols:2006:UPO

[NT06] David M. Nichols and Michael B. Twidale. Usability processes in open source projects. Software Process: Improvement and Practice, 11(2):149–162, March 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Neamtiu:2013:TBU

[NXC13] Iulian Neamtiu, Guowu Xie, and Jianbo Chen. Towards a better understanding of software evolution: an empirical study on open-source software. Journal of Software: Evolution and Process, 25(3):193–218, 2013. ISSN 2047-7473 (print), 2047-7481 (electronic).

Nagy:2010:OAO

[NYB10] Del Nagy, Areej M. Yassin, and Anol Bhattacherjee. Organizational adoption of open source software: barriers and remedies. Communications of the ACM, 53(3):148–151, March 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nourisa:2022:COS

[NZPWR22] Jalil Nourisa, Berit Zeller-Plumhoff, and Regine Willumeit-Römer. CppyABM: an open-source agent-based modeling library to integrate C++ and Python. Software—Practice and Experience, 52(6):1337–1351, June 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Olsson:1990:DAG

[OCH90a] Ronald A. Olsson, Richard H. Crawford, and W. Wilson Ho. Dalek: A GNU, improved programmable debugger. In Proceedings of the 1990 Usenix Summer Conference, Anaheim, CA, pages 221–232. USENIX Association, Berkeley, CA, USA, June 1990.

Olsson:1990:DGI

[OCH90b] Ronald A. Olsson, Richard H. Crawford, and W. Wilson Ho. Dalek: A GNU, improved programmable debugger. In USENIX Association [USE90], pages 221–232. LCCN QA76.8.U65 U81 1990.

Ochem:2009:GCA

[Och09] Quentin Ochem. Gem #56: creating Ada to Java calls using GNAT-AJIS. ACM SIGADA Ada Letters, 29(2):46–49, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2012:GDS

[Och12] Quentin Ochem. Gem #95: dynamic stack analysis in GNAT. ACM SIGADA Ada Letters, 32(1):46–48, April 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

ODonnell:2007:MOS

[O’D07] Casey O’Donnell. Making an open source case for offshoring commentary. IEEE Transactions on Professional Communication, 50(2):85–87, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Owen:2015:TWD

[ODP15] Harry Jon Foord Owen, Clare Duncan, and Nathalie Pettorelli. Testing the water: detecting artificial water points using freely available satellite data and open source software. Remote Sensing in Ecology and Conservation, 1(1):61–72, 2015. ISSN 2056-3485.

Obrenovic:2007:OSS

[OG07] Željko Obrenovic and Dragan Gaševic. Open source software: All you do is put it together. IEEE Software, 24(5):86–95, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ohl:1992:CGM

[Ohl92] Thorsten Ohl. CUG368 — GNUlib for MS-DOS. C Users Journal, 10(11):74–??, November 1992. ISSN 0898-9788.

Oda:1994:PDC

[OK94] M. Oda and T. Kakeshita. Pitfall detection of C programs using pattern matching. Transactions of the Information Processing Society of Japan, 35(11):2427–2436, November 1994. CODEN JSGRD5. ISSN 0387-5806.

Oman:1989:OLA

[Oma89] P. Oman. An objective look at C++ environments. IEEE Software, 6(2):100–103, 105, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Openja:2022:TDF

[OMA^{+}22] Moses Openja, Mohammad Mehdi Morovati, Le An, Foutse Khomh, and Mouna Abidi. Technical debts and faults in open-source quantum software systems: an empirical study. The Journal of Systems and Software, 193(??):??, November 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-

1228 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0164121222001480.

Ombredanne:2020:FOS

[Omb20] P. Ombredanne. Free and open source software license compliance: Tools for software composition analysis. Computer, 53 (10):105–109, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Omsels:2003:OSD

[Oms03] Hermann-Josef Omsels. Open Source und das deutsche Urheber- und Vertragsrecht (Beitrag zur Festschrift für Paul Hertin, mit freundlicher Genehmigung des Verlages C. H. Beck). (German) [Open Source and the German copyright and contract law (contribution to the Festschrift for Paul Hertin, with friendly permission of the publisher C. H. Beck)]. World-Wide Web document., 2003. URL http://www.ifross.de/

Fremdartikel/Festschriftbeitrag.PDF.

OSI:20xx:LSA

[Opexx] Open Source Initiative. Licenses & standards: About open source licenses. Web site, 20xx. URL http://opensource.

org/licenses.

OReilly:1999:LOS

[O’R99] Tim O’Reilly. Lessons from open-source software development. Communications of the ACM, 42(4):32–37, April 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
journals/cacm/1999-42-4/p32-o_reilly/.

OReilly:20xx:SOS

[O’Rxx] Tim O’Reilly. Schlüsse aus der Open-Source-Software-Entwicklung. (German) [key to Open Source software development]. World-Wide Web document., 20xx. URL http://

www.heise.de/tp/deutsch/special/wos/6433/1.html.

Oxer:2006:UH

[ORC06] Jonathan Oxer, Kyle Rankin, and Bill Childers. Ubuntu Hacks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-52720-9. xix + 426 pp. LCCN QA76.76.O63 2006. EUR 29.00. URL http://

www.oreilly.com/catalog/9780596527204.

Oskooi:2010:MFF

[ORI^{+}10] Ardavan F. Oskooi, David Roundy, Mihai Ibanescu, Peter Bermel, J. D. Joannopoulos, and Steven G. Johnson. Meep: a flexible free-software package for electromagnetic simulations by the FDTD method. Computer Physics Communications, 181(3):687–702, March 2010. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550900383X.

Orgogozo:2014:OSM

[ORS^{+}14] L. Orgogozo, N. Renon, C. Soulaine, F. Hénon, S. K. Tomer, D. Labat, O. S. Pokrovsky, M. Sekhar, R. Ababou, and M. Quintard. An open source massively parallel solver for Richards equation: Mechanistic modelling of water fluxes at the watershed scale. Computer Physics Communications, 185(12):3358–3371, December 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465514002719.

OSullivan:2002:MCA

[O’S02] Maureen O’Sullivan. Making copyright ambidextrous: An expose of copyleft. Journal of Information, Law & Technology (JILT), 2002(3), 2002. ISSN 1361-4169.

URL http://elj.warwick.ac.uk/jilt/02-3/osullivan.
html;   http://www2.warwick.ac.uk/fac/soc/law/elj/
jilt/2002_3/osullivan/.

OSullivan:2003:PEQ

[O’S03] Maureen O’Sullivan. The pluralistic, evolutionary, quasi-legal role of the GNU General Public Licence in free/libre/open source software (“FLOSS”). European intellectual property review, 25(8):340–348, August 2003. ISSN 0142-0461.

OSullivan:2004:EFS

[O’S04] Maureen O’Sullivan. EOF: Free software licenses. Linux Journal, 2004(122):11, June 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Ohtani:1994:EIT

[OSM94a] T. Ohtani, H. Sawamura, and T. Minami. EUODHILOS-II on top of GNU epoch. Lecture Notes in Computer Science, 814: 816–820, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohtani:1994:ETG

[OSM94b] T. Ohtani, H. Sawamura, and T. Minami. EUODHILOS-II on top of GNU Epoch. In Bundy [Bun94], pages 816–820. ISBN 3-540-58156-1, 0-387-58156-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D337I58 1994.

Oram:1991:MPM

[OT91] Andrew Oram and Steve Talbott. Managing Projects with Make. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1991. ISBN 0-937175-90-0. xiv + 136 pp. LCCN QA76.76.O63 T35 199.

URL http://www.oreilly.com/catalog/9780937175903;
 http://www.oreilly.com/catalog/make2.

Ousterhout:1999:FSN

[Ous99] John Ousterhout. Free software needs profit. Communications of the ACM, 42(4):44–45, April 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/citations/journals/
cacm/1999-42-4/p44-ousterhout/.

Owen:2001:OSP

[Owe01] J. Michael Owen. An open-source project for modeling hydrodynamics in astrophysical systems. Computing in Science and Engineering, 3(6):54–59, November/December 2001.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X

(electronic). URL http://computer.org/cise/cs2001/
c6054abs.htm;   http://dlib.computer.org/cs/books/
cs2001/pdf/c6054.pdf.

Pedersen:1999:STY

[P^{+}99] Jesper Pedersen et al. Sams Teach Yourself Emacs in 24 Hours. Sams.net Pub., Indianapolis, IN, USA, 1999. ISBN 0-672-31594-7. xvi + 445 pp. LCCN QA76.76.T49 S34 1999. US$29.99.

Paschali:2017:ROS

[PAB^{+}17] Maria-Eleni Paschali, Apostolos Ampatzoglou, Stamatia Bibi, Alexander Chatzigeorgiou, and Ioannis Stamelos. Reusability of open source software across domains: a case study. The Journal of Systems and Software, 134(??):211–227, December 2017. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121217301991.

Page:2007:AOS

[Pag07] William S. Page. Axiom: open source computer algebra system. ACM Communications in Computer Algebra, 41(3):114, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Palmer:1987:DEOb

[Pal87] James E. Palmer. The design and evaluation of online help for UNIX Emacs: Capturing the user in menu design. CDC technical report series 39, Communications Design Center, Carnegie Mellon University, Pittsburgh, PA, USA, November 1987. 17 pp.

Parker:1991:MNE

[Par91] Tim Parker. Moving up to new editors. UNIX review, 9(9): 69–75, September 1991. CODEN UNRED5. ISSN 0742-3136.

Park:2003:HQC

[Par03] Buhm Soon Park. The ‘hyperbola of quantum chemistry’: the changing practice and identity of a scientific discipline in the early years of electronic digital computers, 1945–65. Annals of Science, 60(3):219–247, 2003. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.ingentaconnect.com/content/
tandf/tasc/2003/00000060/00000003/art00001.

Paulson:2004:OSD

[Pau04] Linda Dailey Paulson. Open source databases move into the marketplace. Computer, 37(7):13–??, July 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/dl/mags/co/2004/07/r7013.
htm;   http://csdl.computer.org/dl/mags/co/2004/07/
r7013.pdf.

Paxson:1988:FFL

[Pax88] Vern Paxson. flex — fast lexical analyzer generator. GNU software package., 1988. URL http://flex.sourceforge. net/. See also [DS88] for the bison parser generator.

Paxson:1995:FVF

[Pax95] Vern Paxson. Flex, version 2.5: a fast scanner generator. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2.5 edition, March 1995. ISBN ???? ???? pp. LCCN ????

Payne:2002:SOS

[Pay02] Christian Payne. On the security of open source software. Information Systems Journal, 12(1):61–78, 2002. ISSN 1350-1917 (print), 1365-2575 (electronic).

Poslad:2001:OSS

[PBH01] Stefan Poslad, Phil Buckle, and Rob Hadingham. Open source, standards and scaleable agencies. Lecture Notes in Computer Science, 1887:296–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
1887/18870296.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/1887/18870296.pdf.

Pyakuryal:2012:SIH

[PBJ^{+}12] A. Pyakuryal, I. Bacchus, S. Jang, G. Narayanasamy, M. Gopalakrishnan, D. Pokhrel, J. Luo, V. Sathiaseelan, and B. Mittal. SU-E-T-570: Improvement to the Histogram Analysis in Radiation Therapy (HART): An open source software system for the multi-dimensional dose-volume histogram analysis in Digital Image Communication in Medicine — Radiation Therapy (DICOM-RT) treatment plans. Medical Physics, 39(6Part19):3836–3837, 2012. CODEN MPHYA6. ISSN 2473-4209.

Pedroni:2007:OSP

[PBOP07] Michela Pedroni, Till Bay, Manuel Oriol, and Andreas Pedroni. Open source projects in programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):454–458, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Poppe:2013:CMO

[PC13a] Koen Poppe and Ronald Cools. CHEBINT: a MATLAB/ Octave toolbox for fast multivariate integration and interpolation based on Chebyshev approximations over hypercubes. ACM Transactions on Mathematical Software, 40(1):2:1–2:13, September 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Porenta:2013:CCI

[PC13b] Jernej Porenta and Mojca Ciglaric. Comparing commercial IP reputation databases to open-source IP reputation algorithms. International Journal of Computer Systems Science and Engineering, 28(1):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Padulano:2023:LOS

[PCAJ^{+}23] Vincenzo Eduardo Padulano, Pablo Oliver Cortés, Pedro Alonso-Jordá, Enric Tejedor Saavedra, Sebastián Risco, and Germán Moltó. Leveraging an open source serverless framework for high energy physics computing. The Journal of Supercomputing, 79(8):8940–8965, May 2023. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-022-05016-y.

Palmer:1987:DEOa

[PDG^{+}87] J. E. Palmer, T. M. Duffy, K. Gomoll, T. Gomoll, J. Palmquist-Richards, and J. A. Trumble. The design and evaluation of online help for UNIX EMACS: access mechanisms. In Bullinger et al. [BSK87], pages 461–466. ISBN 0-444-70304-7. LCCN QA76.9.S88I325 1987.

Palmer:1988:DEO

[PDG^{+}88] J. Palmer, T. Duffy, K. Gomoll, T. Gomoll, J. Richards-Palmquist, and J. A. Trumble. The design and evaluation of online help for Unix EMACS: capturing the user in menu design. IEEE Transactions on Professional Communication, 31(1):44–51, March 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Prutchi:2022:HAF

[PdSCJM22] Eduardo S. Prutchi, Heleno de S. Campos Junior, and Leonardo G. P. Murta. How the adoption of feature toggles correlates with branch merges and defects in open-source projects? Software—Practice and Experience, 52(2):506–536, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pearce:2016:RIO

[Pea16] J. M. Pearce. Return on investment for open source scientific hardware development. Science and Public Policy, 43(2):192– 195, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Peck:2008:BGN

[Pec08] Akkana Peck. Beginning GIMP: from novice to professional. The expert’s voice in open source. Apress, Berkeley, CA, USA, second edition, 2008. ISBN 1-4302-1070-2 (paperback). xxvi + 557 pp. LCCN T385 .P395 2008.

Pedersen:2005:EOS

[Ped05] Palle Pedersen. EOF: Open-source use accelerates software development. Linux Journal, 2005(140):17, December 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Peltonen:1989:GEI

[Pel89] Kyle G. Peltonen. A GNU Emacs interface to the Community Information System project. Thesis (b.s.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1989. 3 + 26 pp. Supervised by David K. Gifford.

Penner:2003:PGI

[Pen03] Hartmut E. Penner. Porting GCC to the IBM S/390 platform. In Hutton et al. [HDR03], pages 195–212. ISBN ????

LCCN ???? URL http://www.linux.org.uk/~ajh/gcc/
gccsummit-2003-proceedings.pdf.

Perdue:2000:PPB

[Per00] Tim Perdue. PHP4 and PostgreSQL: Building serious Web applications with open-source software. Linux Journal, 80:134, 136, 138, 140, 142, 144, December 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://noframes.linuxjournal.com/lj-
issues/issue80/4343.html.

Perkonigg:2002:RSG

[Per02] Christian Perkonigg. Realisierung einer Schnittstelle für den GNU Debugger zum Zugriff auf ein Background Debug Modul über den VMEBus. Diplomarbeit am fb-informatik, Fachhochschule Landshut, Landshut, Germany, 2002. 21 pp. Includes CD-ROM.

Perens:2005:OSD

[Per05] Bruce Perens. The Open Source definition: Version 1.9. World-Wide Web document., 2005. URL http://www.

opensource.org/docs/definition.php.

Pesch:1993:GID

[Pes93] R. H. Pesch. GNU info: a decade of hypertext experience. In ACM [ACM93a], pages 233–240. ISBN 0-89791-630-1. LCCN QA76.9.D6I57 1993.

Petullo:2005:DGA

[Pet05] Mike Petullo. Developing GNOME applications with Java. Linux Journal, 2005(135):5, July 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Petreley:2006:ERSc

[Pet06] Nicholas Petreley. Etc/rant: the spirit of open source. Linux Journal, 2006(144):??, April 2006. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Pollino:2012:COS

[PFL^{+}12] Maurizio Pollino, Grazia Fattoruso, Luigi La Porta, Antonio Bruno Della Rocca, and Valentina James. Collaborative open source geospatial tools and maps supporting the response planning to disastrous earthquake events. Future Internet, 4 (2):451–468, May 07, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.com/1999-5903/4/2/451.

Parkinson:2002:HIS

[PG02] Paul Parkinson and Franco Gasperoni. High-integrity systems development for integrated modular avionics using VxWorks and GNAT. Lecture Notes in Computer Science, 2361:163– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2361/23610163.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2361/23610163.pdf.

Ploeg:2021:SOS

[PGC21] S. Ploeg, H. Gunther, and R. M. Camacho. Simphony: An open-source photonic integrated circuit simulation framework. Computing in Science and Engineering, 23(1):65–74, 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Petrisko:2020:BAO

[PGW^{+}20] D. Petrisko, F. Gilani, M. Wyse, D. C. Jung, S. Davidson, P. Gao, C. Zhao, Z. Azad, S. Canakci, B. Veluri, T. Guarino, A. Joshi, M. Oskin, and M. B. Taylor. BlackParrot: An agile open-source RISC-V multicore for accelerator SoCs. IEEE Micro, 40(4):93–102, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Phillips:1982:UML

[PH82] Richard E. Phillips and Phoebe L. Hauff. USGS mineralogy laboratory user’s guide to the TECO editing program for the DEC RT-11 operating system: (part C of the USGS mineralogy laboratory user’s guide to the DEC RT-11 operating system). Open-file report 82-177, U.S. Geological Survey, Denver, CO, USA, 1982. ii + 39 pp.

Petersen:2016:OOH

[PH16] O. Petersen and L. Holst. OP31.05: Obstetric homemonitoring including CTG in complicated pregnancies: development and results of system, based on free available open source software. Ultrasound in Obstetrics & Gynecology, 48: 154–155, 2016. ISSN 0960-7692 (print), 1469-0705 (electronic).

Phillips:1993:SDQ

[Phi93] N. C. K. Phillips. Stream driven query processing in a database. In Deaton et al. [DGBH93], pages 593–597. ISBN 0-89791-567-4. LCCN QA76.76.A65 S95 1993. URL http://

www.acm.org/pubs/contents/proceedings/sac/162754/.

Phillips:2012:GCRa

[Phi12] Lee Phillips. Gnuplot cookbook: over 80 recipes to visually explore the full range of features of the world’s preeminent open source graphing system: [quick answers to common problems]. Packt Pub., Olton Birmingham, UK, 2012. ISBN 1-84951-724-X (paperback), 1-84951-725-8 (ebook). iv + 205 pp. LCCN

T385 .P45 2012. URL https://www.packtpub.com/big-
data-and-business-intelligence/gnuplot-cookbook.

Pohl:2017:OSF

[PHT17] Vincent Pohl, Gunter Hermann, and Jean Christophe Tremblay. An open-source framework for analyzing N-electron dynamics. I. Multideterminantal wave functions. Journal of Computational Chemistry, 38(17):1515–1527, June 30, 2017. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pilz:2009:IGG

[Pil09] Jürgen Pilz, editor. Interfacing Geostatistics and GIS. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-33235-9 (hardcover), 1-281-92749-X, 3-540-33236-7 (e-book). xv + 282 pp. LCCN

QE33.2.S82 I58 2009. URL http://www.springerlink.com/
openurl.asp?genre=book%26isbn=978-3-540-33236-7.

Pinto:2002:PGC

[Pin02] Paulo Pinto. Programming in GNOME with C++. C/C++ Users Journal, 20(5):28–??, May 2002. CODEN CCUJEX. ISSN 1075-2838.

Pittenger:2016:KYO

[Pit16] Mike Pittenger. Know your open source code. Network Security, 2016(5):11–15, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1353485816300484.

Pino:2010:IOS

[PK10] Alexandros Pino and Georgios Kouroupetroglou. ITHACA: An open source framework for building component-based augmentative and alternative communication applications. ACM Transactions on Accessible Computing, 2(4):14:1–14:??, June 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Park:2017:ESS

[PKB17] Jihun Park, Miryung Kim, and Doo-Hwan Bae. An empirical study of supplementary patches in open source projects. Empirical Software Engineering, 22(1):436–473, February 2017. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10664-016-9432-x.

Pyakuryal:2010:SGC

[PKG^{+}10] A. Pyakuryal, A. Kepka, M. Gopalakrishnan, S. Jang, V. Sathiaseelan, and B. B. Mittal. SU-GG-T-141: Current status of the Histogram Analysis in Radiation Therapy (HART): An open-source software system. Medical Physics, 37(6Part17):3217, 2010. CODEN MPHYA6. ISSN 2473-4209.

Papoutsoglou:2022:AOS

[PKGA22] Maria Papoutsoglou, Georgia M. Kapitsaki, Daniel German, and Lefteris Angelis. An analysis of open source software licensing questions in Stack Exchange sites. The Journal of Systems and Software, 183(??):??, January 2022. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121221002107.

Pearce:2007:EFS

[PKH07] David J. Pearce, Paul H. J. Kelly, and Chris Hankin. Efficient field-sensitive pointer analysis of C. ACM Transactions on Programming Languages and Systems, 30(1):4:1–4:42, November 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Prinz:2002:CPB

[PKP02] Peter Prinz and Ulla Kirch-Prinz. C für PCs: [auf Basis des neuen Standards ANSI C 99; Anwendungen: Windows-Programmierung, Grafik und hardwarenahe Programmierung; auf CD: Microsoft C/C++-Compiler (book edition), GNU C/C++-Compiler, Programmbeispiele, Musterlösungen]. MITP Verlag, Bonn, Germany, 2002. ISBN 3-8266-0784-8. 830 (est.) pp. LCCN ???? Includes CD-ROM.

Prinz:2005:CEP

[PKP05] Peter Prinz and Ulla Kirch-Prinz. C — Einführung und professionelle Anwendung: [auf Basis des Standards ANSI C 99; Anwendungen: Windows-Programmierung, Grafik und hardwarenahe Programmierung; auf CD: Microsoft C/C++- Compiler Book edition, GNU C/C++-Compiler, Beispiele, Musterlöungen]. (German) [C — Introductory and Professional Use [based on ANSI C 99]; Use: Windows Programming, Graphics and Hardware Programming; with CD: Microsoft C/C++-Compiler Book edition, GNU C/C++- Compiler, Examples, and Solutions]. mitp-Verlag, Bonn, Germany, 2005. ISBN 3-8266-1580-8. 800 pp. LCCN ???? Includes one CD-ROM.

Penha-Lopes:2005:WUO

[PL05] Joana Matos Penha-Lopes. Why use an open source e-voting system? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):412, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Plonka:1997:MSA

[Plo97] Dave Plonka. Managing sys admin tasks with GNATS. Sys Admin: The Journal for UNIX Systems Administrators, 6(2): 51–??, February 1997. CODEN SYADE7. ISSN 1061-2688.

Pinna:2023:ISA

[PLO^{+}23] Andrea Pinna, Maria Ilaria Lunesu, Stefano Orrù, , and Roberto Tonelli. Investigation on self-admitted technical debt in open-source blockchain projects. Future Internet, 15(7):

232–??, July 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/15/7/232.

Parker:1991:CDD

[PLS^{+}91] Donn D. Parker, Stephen Levy, Eugene Spafford, Paula Hawthorn, Marc Rotenberg, J. J. Buck BloomBecker, and Richard Stallman. Colleagues debate Denning’s comments. Communications of the ACM, 34(3):33–41, March 1991. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pantiuchina:2022:WDD

[PLZ^{+}22] Jevgenija Pantiuchina, Bin Lin, Fiorella Zampetti, Massimiliano Di Penta, Michele Lanza, and Gabriele Bavota. Why do developers reject refactorings in open-source projects? ACM Transactions on Software Engineering and Methodology, 31 (2):23:1–23:23, April 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https://dl.acm.

org/doi/10.1145/3487062.

Paulson:2000:NBU

[PM00] Linda Dailey Paulson and Orren Merton. News briefs: U.S. picks new encryption standard; better software with open source; taking a SIP of Internet telephony; schools may hold valuable spectrum; getting a feel for the Web. Computer, 33(12):20–23, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/rz020.pdf.

Pagano:2013:HDO

[PM13] Dennis Pagano and Walid Maalej. How do open source communities blog? Empirical Software Engineering, 18(6): 1090–1124, December 2013. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10664-012-
9211-2.

Piris:2021:DOS

[PM21] Mario Piris and Ion Mitxelena. DoNOF: an open-source implementation of natural-orbital-functional-based methods for quantum chemistry. Computer Physics Communications, 259(??):Article 107651, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520303167.

Pauliuk:2015:LIE

[PMBM^{+}15] Stefan Pauliuk, Guillaume Majeau-Bettez, Christopher L. Mutel, Bernhard Steubing, and Konstantin Stadler. Lifting industrial ecology modeling to a new level of quality and transparency: a call for more transparent publications and a collaborative open source software framework. Journal of Industrial Ecology, 19(6):937–949, 2015. ISSN 1088-1980 (print), 1530-9290 (electronic).

Peng:2013:LOS

[PMD13] Gang Peng, Jifeng Mu, and C. Anthony Di Benedetto. Learning and open source software license choice. Decision Sciences, 44(4):619–643, 2013. ISSN 0011-7315 (print), 1540-5915 (electronic).

Pyakuryal:2009:SFI

[PMG^{+}09] A. Pyakuryal, K. Myint, M. Gopalakrishnan, S. Jang, V. Sathiaseelan, J. Logemann, and B. Mittal. SU-FF-T-118: Improvements to the Histogram Analysis in Radiation Therapy (HART) open-source software system. Medical Physics, 36(6Part10):2547, 2009. CODEN MPHYA6. ISSN 2473-4209.

Palyart:2017:SSI

[PMM17] M. Palyart, G. C. Murphy, and V. Masrani. A study of social interactions in open source component use. IEEE Transactions on Software Engineering, PP(99):1, ???? 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=8049385.

Palyart:2018:SSI

[PMM18] Marc Palyart, Gail C. Murphy, and Vaden Masrani. A study of social interactions in open source component use. IEEE Transactions on Software Engineering, 44(12):1132–1145, December 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL https://ieeexplore.ieee.

org/document/8049385/.

Plesinger:2023:DOS

[PNK^{+}23] Filip Plesinger, Petr Nejedly, Zuzana Koscova, Maurice Rohr, Ivo Viscor, Radovan Smisek, Adam Ivora, Pavel Leinveber, Karol Curila, and Christoph Hoog Antink. DeepPlayer: an open-source SignalPlant plugin for deep learning inference. Software—Practice and Experience, 53(2):455–464, February 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pomerantz:2004:LHU

[Pom04] Gregory M. Pomerantz. The legal history of UNIX and free software. Web document., June 19, 2004.

URL http://wiki.tuhs.org/lib/exe/fetch.php?media=
publications:theses:gmp_thesis.pdf.

Potter:2006:OSF

[Pot06] Bruce Potter. Open source firewall alternatives. Network Security, 2006(1):16–17, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1353485806703257.

Powell:2000:JDG

[Pow00] Dennis Powell. Judgement day for the GPL?: Determining the legality of the GPL. World-Wide Web document., June 26,

2000. URL http://www.linuxplanet.com/linuxplanet/
reports/2000/1.

Powers:2014:OSCa

[Pow14] Shawn Powers. The open-source classroom: encrypting your cat photos. Linux Journal, 2014(237):8:1–8:??, January 2014. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Petcu:2015:CRO

[PPC^{+}15] Dana Petcu, Silviu Panica, Ciprian Crăciun, Marian Neagul, and Calin Şandru. Cloud resource orchestration within an open-source component-based platform as a service. Concurrency and Computation: Practice and Experience, 27(9): 2443–2469, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pyakuryal:2011:SAH

[PPG^{+}11] A. Pyakuryal, D. Pokhrel, M. Gopalakrishnan, S. Jang, J. Luo, Evgin Gocer, V. Sathiaseelan, and B. Mittal. SU-E-T-816: Application of the Histogram Analysis in Radiation Therapy (HART): An open source software system. Medical Physics, 38(6Part23):3678–3679, 2011. CODEN MPHYA6. ISSN 2473-4209.

Ponnuswamy:2019:FRS

[PPR19] P. Priya Ponnuswamy, R. Vidhya Priya, and C. P. Shabari Ram. File retrieval and storage in the open source cloud tool using digital bipartite and digit compact prefix indexing method. Concurrency and Computation: Practice and Experience, 31(14):e5307:1–e5307:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Petrenko:2007:TSE

[PPRB07] Maksym Petrenko, Denys Poshyvanyk, Václav Rajlich, and Joseph Buchta. Teaching software evolution in open source. Computer, 40(11):25–31, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pemstein:2011:SSL

[PQM11] Daniel Pemstein, Kevin M. Quinn, and Andrew D. Martin. The Scythe Statistical Library: An open source C++ library for statistical computation. Journal of Statistical Software, 42 (12):??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v42/i12.

Purcell:1996:LBG

[PR96] John Purcell and Amanda Robinson, editors. The Linux Bible: The GNU Testament. Yggdrasil Computing, Inc., Berkeley, CA, USA, fourth edition, 1996. ISBN 1-883601-10-X (??), 1-883601-20-7. v + 1886 pp. LCCN QA76.76.O63

L5458 1996. US$39.95. URL http://www.yggdrasil.com/
Products/bible4.html. Includes CD ROM.

Prasad:2003:OSJ

[Pra03] G. Prasad. Open Source Java: Fortune 500 systems at twoguys-in-a-garage prices. Cutter IT Journal, 16(5):10–15, 2003. CODEN ???? ISSN 1522-7383.

Preining:TB26-3-241

[Pre05] Norbert Preining. TEX Live for Debian. TUGboat, 26(3): 241–242, 2005. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb26-3/tb84preining.pdf.

Preining:TB29-1-136

[Pre08] Norbert Preining. TEX (Live) on Debian. TUGboat, 29(1): 136–139, 2008. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb29-1/tb91preining.pdf.

Preining:2016:YTL

[Pre16a] Norbert Preining. 10 years of TEX Live in Debian. TUGboat, 37(1):45–47, 2016. ISSN 0896-3207.

Preining:TB37-1-45

[Pre16b] Norbert Preining. 10 years of TEX Live in Debian. TUGboat, 37(1):45–47, 2016. ISSN 0896-3207.

Priestley:2019:FPH

[Pri19] Mark Priestley. For Fun and Profit: A History of the Free and Open Source Software Revolution by Christopher J. Tozzi (review). Technology and Culture, 60(2):655–657, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
726960.

Piva:2012:OSS

[PRRL12] Evila Piva, Francesco Rentocchini, and Cristina Rossi-Lamastra. Is open source software about innovation? Collaborations with the open source community and innovation performance of software entrepreneurial ventures. Journal of Small Business Management, 50(2):340–364, 2012. ISSN 0047-2778 (print), 1540-627X (electronic).

Pittman:2009:SOS

[PS^{+}09] Gregory Pittman, Christoph Schäfer, et al., editors. Scribus: open source desktop publishing: the official manual. FLES Books, Lostwithiel, Cornwall, UK, 2009. ISBN 0-9560780-0-1 (paperback). 439 pp. LCCN Z253.532 S33 S434 2009.

Pinto:2018:COS

[PSDG18] Gustavo Pinto, Igor Steinmacher, Luiz Felipe Dias, and Marco Gerosa. On the challenges of open-sourcing proprietary software projects. Empirical Software Engineering, 23(6):3221– 3247, December 2018. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://link.

springer.com/article/10.1007/s10664-018-9609-6.

Paulson:2004:ESO

[PSE04] J. W. Paulson, G. Succi, and A. Eberlein. An empirical study of open-source and closed-source software products. IEEE Transactions on Software Engineering, 30(4):246– 256, April 2004. CODEN IESEDJ. ISSN 0098-5589 (print),

1939-3520 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1274044.

Prana:2021:SMH

[PSL21] Gede Artha Azriadi Prana, Abhishek Sharma, and David Lo. Out of sight, out of mind? How vulnerable dependencies affect open-source projects. Empirical Software Engineering, 26(4): ??, July 2021. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL https://link.springer.com/
article/10.1007/s10664-021-09959-3.

Politikos:2022:DVO

[PSP^{+}22] Dimitris V. Politikos, Nikolaos Sykiniotis, Georgios Petasis, Pavlos Dedousis, Alba Ordoñez, Rune Vabø, Aikaterini Anastasopoulou, Endre Moen, Chryssi Mytilineou, Arnt-Børre Salberg, Archontia Chatzispyrou, and Ketil Malde. DeepOtolith v1.0: an open-source AI platform for automating fish age reading from otolith or scale images. Fishes, 7(3):121, May 29, 2022. CODEN ???? ISSN 2410-3888. URL https://www.

mdpi.com/2410-3888/7/3/121.

Prusa:2016:DWT

[PSR16] Zdeněk Průsa, Peter L. Søndergaard, and Pavel Rajmic. Discrete wavelet transforms in the large time-frequency analysis toolbox for MATLAB/GNU Octave. ACM Transactions on Mathematical Software, 42(4):32:1–32:23, July 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://dl.acm.org/citation.cfm?id=2839298.

Poullet:2007:AQS

[PSS^{+}07] Jean-Baptiste Poullet, Diana M. Sima, Arjan W. Simonetti, Bart De Neuter, Leentje Vanhamme, Philippe Lemmerling, and Sabine Van Huffel. An automated quantitation of short echo time MRS spectra in an open source software environment: AQSES. NMR in Biomedicine, 20(5):493–504, 2007. ISSN 0952-3480 (print), 1099-1492 (electronic).

Palmer:2016:PES

[PSSH16] P. J. Palmer, M. A. Sinclair, C. E. Siemieniuch, and M. J. De Henshaw. A practical example of a software factory: building a custom application for analysing EU Cyber Physical System (CPS) projects using open source software components. IN-COSE International Symposium, 26(1):336–351, 2016. ISSN 2334-5837.

Prehn:1991:VFS

[PT91] S. Prehn and W. J. Toetenel, editors. VDM ’91: formal software development methods: 4th international symposium of VDM Europe, Noordwijkerhout, The Netherlands, October 21–25, 1991: proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-54834-3 (New York) (vol. 1), 3-540-54834-3 (Berlin) (vol. 1), 0-540-54868-8 (New York) (vol. 2), 3-540-54868-8 (Berlin) (vol. 2). LCCN QA76.76.D47V36 1991.

Puder:2004:MOS

[Pud04] Arno Puder. MICO: An Open Source CORBA implementation. IEEE Software, 21(4):17–19, July/August 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Punjani:2004:RRG

[Pun04] Mukta Punjani. Register rematerialization in GCC. In Hutton et al. [HDR04], pages 131–139. ISBN ???? LCCN

???? URL http://people.redhat.com/lockhart/.gcc04/
MasterGCC-2side.pdf.

Plaga:2019:SFD

[PWA^{+}19] Sven Plaga, Norbert Wiedermann, Simon Duque Anton, Stefan Tatschner, Hans Schotten, and Thomas Newe. Securing future decentralised industrial IoT infrastructures: Challenges and free open source solutions. Future Generation Computer Systems, 93(??):596–608, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167739X18314043.

Pyati:2006:CTM

[Pya06] Ajit Pyati. A critical theoretical model for library-led technological development: a case of open source software and libraries. Proceedings of the American Society for Information Science and Technology, 43(1):1–4, 2006. ISSN 2373-9231.

Porter:2006:TPI

[PYM^{+}06] Adam Porter, Cemal Yilmaz, Atif M. Memon, Arvind S. Krishna, Douglas C. Schmidt, and Aniruddha Gokhale. Techniques and processes for improving the quality and performance of open-source software. Software Process: Improvement and Practice, 11(2):163–176, March 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

PyramidTechnology:1984:ERD

[Pyr84] Pyramid Technology Corporation, Mountain View, CA, USA. Emacs release description, revision a edition, 1984. various pp.

Phipps:2020:COS

[PZ20] S. Phipps and S. Zacchiroli. Continuous open source license compliance. Computer, 53(12):115–119, December 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Qi:2021:EEO

[QB21] Kan Qi and Barry Boehm. Effort estimation of open source Android projects via transaction analysis. Journal of Software: Evolution and Process, 33(1):e2253:1–e2253:??, January 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Quick:2018:DFI

[QC18] Darren Quick and Kim-Kwang Raymond Choo. Digital forensic intelligence: Data subsets and open source intelligence (DFINT + OSINT): a timely and cohesive mix. Future Generation Computer Systems, 78 (part 2)(??):558–567, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X16308639.

Qian:2012:FOS

[QLC^{+}12] Yu Qian, Yue Liu, John Campbell, Elizabeth Thomson, Y. Megan Kong, and Richard H. Scheuermann. FCSTrans: An open source software system for FCS file conversion and data transformation. Cytometry Part A, 81A(5):353–356, 2012. ISSN 1552-4922 (print), 1552-4930 (electronic).

Quinton:1992:APV

[QR92] Patrice Quinton and Yves Robert, editors. Algorithms and parallel VLSI architectures II: proceedings of the International Workshop, Algorithms and Parallel VLSI Architectures II, Château de Bonas, Gers, France, June 3–6, 1991. Elsevier, Amsterdam, The Netherlands, 1992. ISBN 0-444-89153-6. LCCN QA76.58.I57 1991.

Qin:2015:SNU

[QSX^{+}15] Xinming Qin, Honghui Shang, Hongjun Xiang, Zhenyu Li, and Jinlong Yang. Software news & updates: HONPAS: a linear scaling open-source solution for large system simulations. International Journal of Quantum Chemistry, 115(10): 647–655, May 15, 2015. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Quin:2000:OSX

[Qui00] Liam Quin. Open source XML database toolkit: resources and techniques for improved development. Wiley, New York, NY, USA, 2000. ISBN 0-471-37522-5 (paperback). xiv + 434 pp. LCCN QA76.76.H94 Q56 2000.

Ronglong:2016:SOS

[RA16] Suthat Ronglong and Chonlameth Arpnikanondt. Signal: an open-source cross-platform universal messaging system with feedback support. The Journal of Systems and Software, 117(??):30–54, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121216000510.

Racine:2000:CTG

[Rac00] J. Racine. The Cygwin tools: a GNU toolkit for Windows. Journal of Applied Econometrics, 15(3):331–341, 2000. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Racine:2006:SRG

[Rac06] Jeff Racine. Software review: gnuplot 4.0: a portable interactive plotting utility. Journal of Applied Econometrics, 21(1): 133–141, January–February 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Radel:1989:FTS

[Rad89] Jon Radel. “Free” TEX software for IBM PCs. TUGboat, 10 (2):202, July 1989. ISSN 0896-3207.

Rada:1992:CTH

[Rad92] Roy Rada. Converting a textbook to hypertext. ACM Transactions on Information Systems, 10(3):294–315, July 1992. CODEN ATISET. ISSN 1046-8188. URL http://www.acm.

org:80.

Raffel:2023:VBM

[Raf23] Colin Raffel. Viewpoint: Building machine learning models like open source software. Communications of the ACM, 66 (2):38–40, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3545111.

Richardson:2001:LEO

[RAH^{+}01] Bruce Richardson, Anonymous, Nathan Hokanson, Ken O. Burtch, Jim V., Jerel Crosland, Paul Taylor, Sheldon Dubrowin, Paul Dale Roberts, Dean Provins, Kathy Lynn, and Andre Lessa. Letters to the editor: Offended; A real bastard; common misconception; Ada boy!; wacky names; penultimate Linux box?; SuSe too loosa; LJ interactive; sold on Soldier; groff is great; what’s up with Ogg?; changes to the Python Developer’s Handook. Linux Journal, 83:6, 141–142, March 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Raja:2013:ACC

[Raj13] Uzma Raja. All complaints are not created equal: text analysis of open source software defect reports. Empirical Software Engineering, 18(1):117–138, February 2013. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10664-012-9197-9.

Rajanen:2023:OSU

[Raj23] Mikko Rajanen. Open source usability and user experience. Computer, 56(2):106–110, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramey:1994:BGS

[Ram94a] Chet Ramey. Bash — the GNU shell. ;login: the USENIX Association newsletter, 19(6):25–33, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Ramey:1994:WGBb

[Ram94b] Chet Ramey. What’s GNU — bash — the GNU shell (part 2 of 2). Linux Journal, 4:??, August 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Ramey:1994:WGBa

[Ram94c] Chet Ramey. What’s GNU: bash — the GNU shell, part 1 of 2. Linux Journal, 3:??, July 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Roch:2018:PAO

[RAMB18] Loïc M. Roch, Tyanko Aleksiev, Riccardo Murri, and Kim K. Baldridge. Performance analysis of open-source distributed file systems for practical large-scale molecular ab initio, density functional theory, and GW + BSE calculations. International Journal of Quantum Chemistry, 118(1):??, January 5, 2018. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Raportirenko:1994:GPS

[Rap94] A. M. Raportirenko. GSL: a portable standard Lisp interpreter. In Becks and Perret-Gallix [BPG94], pages 645–650. ISBN 981-02-1699-8. LCCN QC793.47.E4I58 1993.

Rau:2004:PBC

[Rau04] Lars Rau. Phänomenologie und Bekämpfung von ‘Cyberpiraterie’: Eine kriminologische und kriminalpolitische Analyse. (German) [Phenomenology and the Fight against ‘Cyberpiracy’: A criminological and criminal-political analysis]. Doctoral thesis, Justus-Liebig-Universität, Gießen, Germany,

2004. 342 pp. URL http://geb.uni-giessen.de/geb/
volltexte/2004/1819/.

Ravicher:2000:FCS

[Rav00] Daniel Ravicher. Facilitating collaborative software development: The enforceability of mass-market public software licenses. Virginia Journal of Law and Technology, 5(11):1522–

1687, 2000. URL http://www.vjolt.net/vol5/issue3/
v5i3a11-Ravicher.html.

Reuter:2016:FMG

[RAW^{+}16] Balthasar Reuter, Vadym Aizinger, Manuel Wieland, Florian Frank, and Peter Knabner. FESTUNG: a MATLAB/ GNU Octave toolbox for the discontinuous Galerkin method, Part II: Advection operator and slope limiting. Computers and Mathematics with Applications, 72(7):1896–1925, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122116304606.

Raymond:1991:NHD

[Ray91] Eric Raymond. The New Hacker’s Dictionary. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-68069-6. xx + 433 pp. LCCN PN6231.E4 H3 1991. This book corresponds to version 2.9.6 of the on-line jargon file. The latest (at the time of writing) is version 2.9.12 (jargon2912.txt.z) which is avail-

able by anonymous ftp from prep.ai.mit.edu (in /pub/gnu)
or wuarchive.wustl.edu (in mirrors/gnu). Changes since

the publication of this book can be found in the file jargonupd.z. (*.z are files compressed by GNU zip (gzip)).

Raymond:1998:SPO

[Ray98] Eric Raymond. Stop the presses: Open source summit. Linux Journal, 50:??, 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://www.ssc.com/ lj/issue50/2918.html. Discusses Perl, Python, and Tcl.

Raymond:1999:ILO

[Ray99a] Eric Raymond. Interview: Linux and open-source success. IEEE Software, 16(1):85–89, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459

(electronic). URL http://dlib.computer.org/so/books/
so1999/pdf/s1085.pdf.

Raymond:1999:CBM

[Ray99b] Eric S. Raymond. The Cathedral & the Bazaar: Musings on Linux and Open Source by an Accidental Revolutionary. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-724-9. xi + 268 pp. LCCN QA76.76.O63 R396 1999. US$19.95. URL http://

www.oreilly.com/catalog/cb.

Raymond:1999:CBP

[Ray99c] Eric S. Raymond. The Cathedral and the Bazaar: Musings on Linux and Open Source by an Accidental Revolutionary. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, February 1999. ISBN 0-596-00108-8. 208 pp. LCCN QA76.76.O63 R396 1999. US$16.95. URL http:

//www.oreilly.com/catalog/9780596001087;   http://
www.oreilly.com/catalog/cathbazpaper.

Raymond:2001:WMS

[Ray01a] E. S. Raymond. Why Microsoft smears — and fears — open source. IEEE Spectrum, 38(8):14–15, August 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Raymond:2001:CBM

[Ray01b] Eric S. Raymond. The cathedral and the bazaar: musings on Linux and open source by an accidental revolutionary. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, revised edition, 2001. ISBN 0-596-00131-2. xiv + 241 pp. LCCN QA76.76.O63 R397 2001. URL http:/

/www.oreilly.com/catalog/9780596001315.

Roberts:1992:UDG

[RB92] Ralph Roberts and Mark Boyd. UNIX Desktop Guide to Emacs. Programming series. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1992. ISBN 0-672-30171-7. xxiii + 504 pp. LCCN QA76.76.T49 R62 1992. US$27.95, CDN$34.95.

Raja:2005:IQL

[RB05a] Uzma Raja and Evelyn Barry. Investigating quality in largescale open source software. ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ritsko:2005:Pb

[RB05b] John J. Ritsko and Alex Birman. Preface. IBM Systems Journal, 44(2):213–??, ???? 2005. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/
sj/442/preface.pdf.

Rappoport:2023:LMO

[RBM^{+}23] Dmitrij Rappoport, Samuel Bekoe, Luke Nambi Mohanam, Scott Le, Naje’ George, Ziyue Shen, and Filipp Furche. Libkrylov: a modular open-source software library for extremely large on-the-fly matrix computations. Journal of Computational Chemistry, 44(11):1105–1118, April 30, 2023. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Roy:2010:NMF

[RC10] C. K. Roy and J. R. Cordy. Near-miss function clones in open source software: an empirical study. Journal of Software Maintenance and Evolution: Research and Practice, 22(3): 165–189, April 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Rothenberg:2014:WOS

[RCB^{+}14] Christian Esteve Rothenberg, Roy Chua, Josh Bailey, Martin Winter, Carlos N. A. Correa, Sidney C. de Lucena, Marcos Rogerio Salvador, and Thomas D. Nadeau. When open source meets network control planes. Computer, 47(11):46–54, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print),

1558-0814 (electronic). URL http://csdl.computer.org/
csdl/mags/co/2014/11/mco2014110046-abs.html.

Robles:2022:DEE

[RCGB^{+}22] Gregorio Robles, Andrea Capiluppi, Jesus M. Gonzalez-Barahona, Björn Lundell, and Jonas Gamalielsson. Development effort estimation in free/open source software from activity in version control systems. Empirical Software Engineering, 27(6):??, November 2022. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL https://link.

springer.com/article/10.1007/s10664-022-10166-x.

Rashid:2020:MEP

[RCO20] Mehvish Rashid, Paul M. Clarke, and Rory V. O’Connor. A mechanism to explore proactive knowledge retention in open source software communities. Journal of Software: Evolution and Process, 32(3):e2198:1–e2198:??, March 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rigby:2012:CPR

[RCP^{+}12] Peter Rigby, Brendan Cleary, Frederic Painchaud, Margaret-Anne Storey, and Daniel German. Contemporary peer review in action: Lessons from open source development. IEEE Software, 29(6):56–61, November/December 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Remondino:2012:LCO

[RDKT12] Fabio Remondino, Silvio Del Pizzo, Thomas P. Kersten, and Salvatore Troisi. Low-cost and open-source solutions for automated image orientation — a critical overview. Lecture Notes in Computer Science, 7616:40–54, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34234-9_5/.

Rousseau:2020:SPT

[RDZ20] Guillaume Rousseau, Roberto Di Cosmo, and Stefano Zacchiroli. Software provenance tracking at the scale of public source code. Empirical Software Engineering, 25(4):2930– 2959, July 2020. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
article/10.1007/s10664-020-09828-5.

Ruffin:2004:UOS

[RE04] Michel Ruffin and Christof Ebert. Using Open Source software in product development: A primer. IEEE Software, 21(1):82– 86, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rehman:2001:CRH

[Reh01a] Rafeeq Ur Rehman. Configuring Red Hat’s high-availability server. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):43–44, 46, 48, April 2001. CODEN SYADE7.

ISSN 1061-2688. URL http://www.samag.com/.

Rehman:2001:OSR

[Reh01b] Rafeeq Ur Rehman. Open source real-time operating systems. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):91–95, January 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.com/.

Reid:1993:IE

[Rei93] Jonathan Makoto Reid. Internationalizing Emacs. Thesis (m.s.), University of Illinois at Urbana-Champaign, Urbana-Champaign, IL 61801, USA, 1993. v + 79 pp.

Rigby:2014:PRO

[RGCS14] Peter C. Rigby, Daniel M. German, Laura Cowen, and Margaret-Anne Storey. Peer review on open-source software projects: Parameters, statistical models, and theory. ACM Transactions on Software Engineering and Methodology, 23 (4):35:1–35:??, August 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Reinauer:2021:DAO

[RH21] Tobias Reinauer and Ulrich Elmer Hansen. Determinants of adoption in open-source hardware: a review of small wind turbines. Technovation, 106(??):??, August 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000705.

Reuter:2021:FOU

[RHR^{+}21] Balthasar Reuter, Hennes Hajduk, Andreas Rupp, Florian Frank, Vadym Aizinger, and Peter Knabner. FESTUNG 1.0: Overview, usage, and example applications of the MATLAB/ GNU Octave toolbox for discontinuous Galerkin methods. Computers and Mathematics with Applications, 81(??):3–41, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0898122120303254.

Rossini:2004:ESS

[RHS^{+}04] A. J. Rossini, Richard M. Heiberger, Rodney A. Sparapani, Martin Mächler, and Kurt Hornik. Emacs speaks statistics. Journal of Computational and Graphical Statistics, 13(1):247– 261, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Raza:2021:NCO

[RHW^{+}21] Akber Raza, Chengkuan Hong, Xian Wang, Anshuman Kumar, Christian R. Shelton, and Bryan M. Wong. NIC-CAGE: an open-source software package for predicting optimal control fields in photo-excited chemical systems. Computer Physics Communications, 258(??):Article 107541, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302587.

Riehle:2007:EMO

[Rie07] Dirk Riehle. The economic motivation of open source software: Stakeholder perspectives. Computer, 40(4):25–32, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2010:ECO

[Rie10] Dirk Riehle. The economic case for open source foundations. Computer, 43(1):86–90, January 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2011:CSO

[Rie11] Dirk Riehle. Controlling and steering open source projects. Computer, 44(7):93–96, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2015:HOS

[Rie15] Dirk Riehle. How open source is changing the software developer’s career. Computer, 48(5):51–57, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/csdl/mags/co/2015/05/
mco2015050051-abs.html.

Riehle:2019:IOS

[Rie19] D. Riehle. The innovations of open source. Computer, 52 (4):59–63, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2020:SVO

[Rie20] D. Riehle. Single-vendor open source firms. Computer, 53 (4):68–72, 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2021:OSD

[Rie21] Dirk Riehle. The open source distributor business model. Computer, 54(12):99–103, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ritchie:1988:SHC

[Rit88] Russell A. Ritchie. Scottish HCI centre extensions for GNU emacs. Report AMU8829/01S (also Strathclyde University computer science report; no. CS/HCI/8829/01), Scottish HCI Centre, Edinburgh, Scotland, March 9, 1988. 22 pp.

Riesch:2021:MOS

[RJ21] Michael Riesch and Christian Jirauschek. mbsolve: an opensource solver tool for the Maxwell–Bloch equations. Computer Physics Communications, 268(??):Article 108097, November 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002095.

Rudmark:2023:BPH

[RLTD23] Daniel Rudmark, Juho Lindman, Andreas Tryti, and Brede Dammen. Beyond procurement: How Entur navigated the open source journey to advance public transport. IEEE Software, 40(4):62–70, 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Roque:2021:ICH

[RLVdS21] João V. Roque, João D. Lopes, Mário P. Véstias, and José T. de Sousa. IOb-cache: a high-performance configurable opensource cache. Algorithms (Basel), 14(8), August 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https:/

/www.mdpi.com/1999-4893/14/8/218.

Rada:1992:SVB

[RM92] R. Rada and C. Murphy. Searching versus browsing in hypertext. Hypermedia, 4(1):1–30, 1992. CODEN HYPEEW. ISSN 0955-8543.

Rhodes:1999:PPD

[RM99] Neil Rhodes and Julie McKeehan. Palm Programming: The Developer’s Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-525-4. xxi + 457 pp. LCCN QA76.5 .R56 1999. US$32.95. URL

http://www.oreilly.com/catalog/palmprog. Includes CD-

ROM.

Rigger:2019:UGB

[RMAM19] Manuel Rigger, Stefan Marr, Bram Adams, and Hanspeter Mössenböck. Understanding GCC builtins to develop better tools. arXiv.org, ??(??):??, July 1, 2019. URL https://

arxiv.org/abs/1907.00863.

Roumani:2017:ATE

[RNR17] Yaman Roumani, Joseph K. Nwankpa, and Yazan F. Roumani. Adopters’ trust in enterprise open source vendors: an empirical examination. The Journal of Systems and Software, 125(??):256–270, March 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL //www.sciencedirect.com/science/article/pii/
S0164121216302503.

Rodriguez:2001:EIL

[RÓ01] Ameneiros Rodríguez and Ibán Óscar. Estudio e implementación de una LAN para PYMES utilizando GNU/LINUX como sistema operativo. (Spanish) [Study and implementation of a LAN with PYMES using GNU/Linux as operating system]. E.U. Politécnica, Ferrol, Spain, 2001. ISBN ???? LCCN ???? Includes one CD-ROM.

Robbins:1994:WGGa

[Rob94a] Arnold Robbins. What’s GNU?: GNU’s not Unix! Linux Journal, 1:??, March 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue1/whatsgnu.html.

Robbins:1994:WGGb

[Rob94b] Arnold Robbins. What’s GNU: groff. Linux Journal, 7: ??, November 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue7/whatsgnu.7.html.

Robbins:1994:WGS

[Rob94c] Arnold Robbins. What’s GNU?: Software tools. Linux Journal, 2:??, April/May 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue2/gnu2.html.

Robbins:1994:WGT

[Rob94d] Arnold Robbins. What’s GNU: Texinfo. Linux Journal, 6:??, October 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue6/texinfo.html.

Robbins:1995:WG

[Rob95a] Arnold Robbins. What’s GNU? Linux Journal, 10:??, February 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Robbins:1995:WGPa

[Rob95b] Arnold Robbins. What’s GNU? [Plan 9 Part I]. Linux Journal, 10:??, March 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://interactive.

linuxjournal.com/article/1012.

Robbins:1995:WGPb

[Rob95c] Arnold Robbins. What’s GNU? Plan 9 Part II. Linux Journal, 10:??, April 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

interactive.linuxjournal.com/article/0062;   http:
//portal.acm.org/citation.cfm?id=324810.

Robbins:1996:EAP

[Rob96] Arnold Robbins. Effective AWK Programming: a User’s Guide for GNU AWK. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1996. ISBN 0-916151-88-3. viii + 322 pp. LCCN QA76.73.A95 R63 1996. US$27.00. URL

http://www.ssc.com/ssc/eap/.

Robbins:1997:EAP

[Rob97] Arnold Robbins. Effective AWK Programming. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, second edition, 1997. ISBN 1-57831-000-8. x + 322 pp. LCCN QA76.73.A95 R63 1997. US$27.00. URL http://www.ssc.

com/ssc/eap/.

Roberts:2005:GMO

[Rob05] Keith A. Roberts. Generic methodology for open source software development. ACM SIGSOFT Software Engineering Notes, 30(2):1–5, March 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robbins:2011:GAT

[Rob11] Arnold Robbins. GNU awk 4.0: teaching an old bird some new tricks. Linux Journal, 2011(209):5:1–5:??, September 2011. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Robson:2020:ISO

[Rob20] R. Robson. IEEE SA Open: Engaging industry, academia, and researchers in open source development. Computer, 53 (6):53–56, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rodriguez:2000:LLG

[Rod00] Dominique Rodriguez. L’essentiel de L^{A}TEX et GNU-Emacs: manuel de réalisation de documents scientifiques, CD-ROM TEXlive’4 GNU-Emacs 20.5 pour Windows, exercices corrigés. Informatiques. Série Réseaux et télécoms. Dunod, Paris, France, 2000. ISBN 2-10-004814-7. ISSN 1622-5694. xv + 352 pp. LCCN ???? Includes CD-ROM.

Rogers:2009:GBBa

[Rog09a] Pat Rogers. Gem #35: bounded buffer package in GNAT hierarchy (part 1). ACM SIGADA Ada Letters, 29(1):54–56, April 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2009:GBBb

[Rog09b] Pat Rogers. Gem #37: bounded buffer package in GNAT hierarchy (part 2). ACM SIGADA Ada Letters, 29(1):58–60, April 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2011:GGS

[Rog11] Pat Rogers. Gem #81: GNAT semaphores. ACM SIGADA Ada Letters, 31(2):33–35, August 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ronneburg:2001:DGLa

[Ron01a] Frank Ronneburg. Debian GNU, Linux-Anwenderhandbuch. Lehmann, Berlin, Germany, 2001. ISBN 3-931253-73-2. 600 (est.) pp. LCCN ???? Includes CD-ROM.

Ronneburg:2001:DGLb

[Ron01b] Frank Ronneburg. Debian GNU, Linux-Anwenderhandbuch. Addison-Wesley, Reading, MA, USA, 2001. ISBN 3-8273-1748-7. 600 (est.) pp. LCCN ???? Includes CD-ROM.

Ronneburg:2005:DGLb

[Ron05a] Frank Ronneburg. Debian GNU/Linux Anwenderhandbuch: für Einsteiger, Umsteiger und Fortgeschritten. (German) [Debian GNU/Linux User Handbook: for Beginning, Switching, and Advanced Users]. Addison Wesley in Pearson Education Deutschland, München, Germany, 2005. ISBN 3-8273-2148-4. 744 pp. LCCN ???? EUR49.95. Includes one DVD.

Ronneburg:2005:DGLa

[Ron05b] Frank Ronneburg. Debian GNU/Linux: installation, administration, exploitation. (French) [Debian GNU/Linux: installation, administration, use]. Campus Press, Paris, France, 2005. ISBN 2-7440-1941-0. 624 pp. LCCN ???? Includes one CD-ROM.

Rondeau:2015:GRE

[Ron15] Thomas W. Rondeau. On the GNU Radio ecosystem. In Oliver Holland, Hanna Bogucka, and Arturas Medeisis, editors, Opportunistic Spectrum Sharing and White Space Access, pages 25–48. Wiley, New York, NY, USA, 2015. ISBN 1-119-05724-8.

Rosenberg:2000:OSU

[Ros00] Donald K. Rosenberg. Open Source: the unauthorized white papers. IDG Books Worldwide, Foster City, CA, 2000. ISBN 0-7645-4660-0 (paperback). xxvi + 374 pp. LCCN QA76.6.

R679 2000. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http://www.loc.gov/catdir/toc/
wiley021/00044952.html;   http://www.stromian.com/
Book/.

Rosen:2001:GLN

[Ros01a] Lawrence Rosen. Geek law: Naming open-source software. Linux Journal, 90:88, October 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Rosen:2001:NYOa

[Ros01b] Lawrence Rosen. Naming your Open Source software, Part I. World-Wide Web document., 2001. URL http://www.

rosenlaw.com/html/GL6.pdf.

Rosen:2001:NYOb

[Ros01c] Lawrence Rosen. Naming your Open Source software, Part II. World-Wide Web document., 2001. URL http://www.

rosenlaw.com/html/GL7.pdf.

Rosen:2001:UFI

[Ros01d] Lawrence Rosen. The unreasonable fear of infection. World-Wide Web document., 2001. URL http://www.rosenlaw.

com/html/GPL.PDF.

Rose:2002:ASO

[Ros02a] Julian Rose. Anatomy of a small open-source kernel for DSPs. Embedded Linux Journal Online, July 27, 2002. URL http:/

/www.linuxdevices.com/articles/AT6489405871.html.

Rosen:2002:BL

[Ros02b] Lawrence Rosen. Bad law. World-Wide Web document., 2002.

URL http://www.rosenlaw.com/html/GL11.pdf.

Rosen:2002:LD

[Ros02c] Lawrence Rosen. License defamation. World-Wide Web doc-

ument., 2002. URL http://www.rosenlaw.com/html/GL12.
pdf.

Rosen:2005:BIS

[Ros05] Jean-Pierre Rosen. On the benefits for industrials of sponsoring free software developments. Ada User Journal, 26(4): 308–??, December 2005. CODEN AUJOET. ISSN 1381-6551.

Ross:2014:OSD

[Ros14] P. E. Ross. Open-source drones for fun and profit. IEEE Spectrum, 51(3):54–59, March 2014. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rowe:2002:GFG

[Row02] Robin Rowe. GFX film: GIMP at rhythm & hues. Linux Journal, 95:34–37, March 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Romeo:2008:JCO

[RP08] Francesco Romeo and Gianluca Padoan. JTruss: a CADoriented educational open-source software for static analysis of truss-type structures. Computer Applications in Engineering Education, 16(4):280–288, 2008. CODEN CAPEED. ISSN 1061-3773 (print), 1099-0542 (electronic).

Robles:2019:TYO

[RSAT19] G. Robles, I. Steinmacher, P. Adams, and C. Treude. Twenty years of open source software: From skepticism to mainstream. IEEE Software, 36(6):12–15, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rahman:2023:SMO

[RSBP23] Akond Rahman, Shazibul Islam Shamim, Dibyendu Brinto Bose, and Rahul Pandita. Security misconfigurations in open source Kubernetes manifests: an empirical study. ACM Transactions on Software Engineering and Methodology, 32 (4):99:1–99:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https://dl.acm.

org/doi/10.1145/3579639.

Rybin:1996:AGP

[RSKF96] S. Rybin, A. Strohmeier, A. Kuchumov, and V. Fofanov. ASIS for GNAT: From the prototype to the full implementation. Lecture Notes in Computer Science, 1088:298–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rybin:1996:AGG

[RSZ96] S. Rybin, A. Strohmeier, and E. Zueff. ASIS for GNAT: goals, problems and implementation strategy. ACM SIGADA Ada Letters, 16(2):39–49, March/April 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ruffolo:2005:LCE

[RT05] Joe Ruffolo and Ron Terry. Linux in the classroom: An experience with Linux and open-source software in an educational environment. Linux Journal, 2005(133):6, May 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Raja:2012:DEM

[RT12] U. Raja and M. J. Tretter. Defining and evaluating a measure of open source project survivability. IEEE Transactions on Software Engineering, 38(1):163–174, January/ February 2012. CODEN IESEDJ. ISSN 0098-5589 (print),

1939-3520 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6127835.

Rivas:2015:MAP

[RTH15] Mario Aldea Rivas, Héctor Pérez Tijero, and Michael González Harbour. Multiprocessor Ada platform based on MaRTE OS and GNAT. ACM SIGADA Ada Letters, 35(1):74–79, April 2015. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rudolphi:2010:ELO

[Rud10] Felix Rudolphi. Ein elektronisches Laborjournal als Open-Source-Software. (German) [An electronic laboratory journal as open source software]. Nachrichten aus der Chemie, 58(5): 548–550, 2010. CODEN NACHFB. ISSN 1439-9598 (print), 1868-0054 (electronic).

Ruiz:2002:GMG

[Rui02] José Antonio Ortega Ruiz. GNU MDK: GNU MIX Development Kit. GNU Press, Boston, MA, USA, 2002. ISBN 1-882114-62-0. 110 pp. LCCN ???? US$16.95. URL http:/

/www.gnupress.org/book16.html.

Ruiz:2013:GRT

[Rui13] José F. Ruiz. Going real-time with Ada 2012 and GNAT. ACM SIGADA Ada Letters, 33(1):45–52, April 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Russel:1988:PCE

[Rus88] Elizabeth Russel. PBE customizer: an expert aide for customizing PBE emacs. Master’s project, Boston University, College of Liberal Arts, Computer Science Dept., Boston, MA, USA, 1988. 32 pp. Also BUCS technical report 88-005.

Russell:2009:TBC

[Rus09] Andrew L. Russell. Two Bits: The Cultural Significance of Free Software (review). Technology and Culture, 50(4):964– 966, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/363449.

Russell:2014:OSD

[Rus14] Andrew L. Russell. Open Standards and the Digital Age: History, Ideology, and Networks. Cambridge studies in the emergence of global enterprise. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-03919-3 (hardcover), 1-107-61204-7 (paperback). xvii + 306 pp. LCCN T59.2.U6 R87 2014.

Ruan:2014:TLC

[RVLS14] Wenjia Ruan, Trilok Vyas, Yujie Liu, and Michael Spear. Transactionalizing legacy code: an experience report using GCC and Memcached. ACM SIGARCH Computer Architecture News, 42(1):399–412, March 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Rich:1987:FRS

[RW87] Charles Rich and Richard C. Waters. Formalizing reusable software components in the programmer’s apprentice. Technical Report AI Memo 954, Massachusetts Institute of Technology, Cambridge, MA, USA, February 1987. 28 pp.

Ritchie:1989:MEG

[RW89] R. A. Ritchie and G. R. S. Weir. Menu-based extensions to GNU emacs. In Sutcliffe and Macaulay [SM89b], pages 245– 257. ISBN 0-521-38430-3. LCCN QA76.9.H85 B75 1989.

Ran:2023:VOS

[RZWW23] Zheng Ran, Chunming Zou, Zunjie Wei, and Hongwei Wang. VELAS: an open-source toolbox for visualization and analysis of elastic anisotropy. Computer Physics Communications, 283(??):Article 108540, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002594.

Scalise:2015:UOS

[SA15] Kathleen Scalise and Diane D. Allen. Use of open-source software for adaptive measurement: Concerto as an R-based computer adaptive development and delivery platform. British Journal of Mathematical and Statistical Psychology, 68(3): 478–496, 2015. ISSN 0007-1102 (print), 2044-8317 (electronic).

Shkarin:2015:OSG

[SAC^{+}15] Andrei Shkarin, Evelina Ametova, Suren Chilingaryan, Timo Dritschler, Andreas Kopmann, Matthias Vogelgesang, Roman Shkarin, and Sergey Tsapko. An open source GPU accelerated framework for flexible algebraic reconstruction at synchrotron light sources. Fundamenta Informaticae, 141(2–3):259–274, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Solano-Altamirano:2015:DCO

[SAHP15] J. M. Solano-Altamirano and Julio M. Hernández-Pérez. DensToolKit: a comprehensive open-source package for analyzing the electron density and its derivative scalar and vector fields. Computer Physics Communications, 196(??):362–371, November 2015. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046551500274X.

SDT:2001:SLG

[Sai01] Sair Development Team. Sair Linux and GNU certification level II: Apache and Web servers. Wiley, New York, NY, USA, 2001. ISBN 0-471-40537-X. xix + 300 pp. LCCN A76.3 .M3236 2001.

SDT:2002:SLG

[Sai02] Sair Development Team. Sair Linux and GNU certification: level II core concepts and practices. Wiley, New York, NY, USA, 2002. ISBN 0-471-40538-8. xvi + 399 pp. LCCN ????

SaiToh:2013:ZCL

[Sai13] Akira SaiToh. ZKCM: a C++ library for multiprecision matrix computation with applications in quantum information. Computer Physics Communications, 184(8):2005–2020, August 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465513001306.

Salzman:1988:OLA

[Sal88] I. J. Salzman. An objective look at C++ compilers. UNIX review, 6(11):81, 83, 85–86, 88, November 1988. CODEN UNRED5. ISSN 0742-3136.

Salus:1994:QCU

[Sal94] Peter H. Salus. A quarter century of UNIX. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54777-5. xii + 256 pp. LCCN QA76.76.O63 S342 1994.

Salus:2008:DGP

[Sal08] Peter H. Salus. The daemon, the gnu, and the penguin: how free and open software is changing the world. Reed Media Services, Keller, TX, USA, 2008. ISBN 0-9790342-3-X. 204 (est.) pp. LCCN ????

Saltzer:2020:OL

[Sal20] J. H. Saltzer. The origin of the MIT License. IEEE Annals of the History of Computing, 42(4):94–98, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Samuelson:2006:LSI

[Sam06] Pamela Samuelson. Legally speaking: IBM’s pragmatic embrace of open source. Communications of the ACM, 49(10): 21–25, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sandewall:1978:PIE

[San78a] Erik Sandewall. Programming in an interactive environment: the “Lisp” experience. ACM Computing Surveys, 10(1):35– 71, March 1978. CODEN CMSVAN. ISSN 0010-4892. See also [Sta78b, San78b].

Sandewall:1978:SFS

[San78b] Erik Sandewall. Surveyor’s forum: Structured editing with a Lisp. ACM Computing Surveys, 10(4):507–508, December 1978. CODEN CMSVAN. ISSN 0010-4892. See [San78a, Sta78b].

Sanders:1998:ELO

[San98] James Sanders. Edgewatch: Linux, open source, and software’s future. IEEE Software, 15(5):88–91, September/ October 1998. CODEN IESOEG. ISSN 0740-7459 (print),

0740-7459 (electronic). URL http://dlib.computer.org/
so/books/so1998/pdf/s5088.pdf.

Sandred:2001:MOS

[San01] Jan Sandred. Managing open source projects: a Wiley tech brief. Wiley, New York, NY, USA, 2001. ISBN 0-471-18917-0 (e-book), 0-471-40396-2. xvii + 189 pp. LCCN QA76.76.D47 S27 2001eb.

Santini:2003:BCI

[San03] Simone Santini. The profession: Bringing copyright into the information age. Computer, 36(8):104, 102–103, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-

0814 (electronic). URL http://csdl.computer.org/dl/
mags/co/2003/08/r8104.htm;   http://csdl.computer.
org/dl/mags/co/2003/08/r8104.pdf.

Santos:2008:VUP

[San08] Carlos Santos, Jr. Viewpoint: Understanding partnerships between corporations and the open source community: a research gap. IEEE Software, 25(6):96–97, November/ December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stamelos:2002:CQA

[SAOB02] Ioannis Stamelos, Lefteris Angelis, Apostolos Oikonomou, and Georgios L. Bleris. Code quality analysis in open source software development. Information Systems Journal, 12(1):43–60, 2002. ISSN 1350-1917 (print), 1365-2575 (electronic).

Savage:2023:NLS

[Sav23] Neil Savage. News: Locking down secure open source software. Communications of the ACM, 66(5):13–14, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3586584.

Schlogl:2008:BFO

[SB08] Alois Schlögl and Clemens Brunner. BioSig: a free and open source software library for BCI research. Computer, 41(10):

44–50, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schoonover:1992:GEU

[SBA92] Michael A. Schoonover, John S. Bowie, and William R. Arnold. GNU Emacs: UNIX Text Editing and Programming. Hewlett-Packard Press series. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56345-2. xxvii + 609 pp. LCCN QA76.76.T49S36.

Smith:1976:MER

[SBD^{+}76] B. T. Smith, J. M. Boyle, J. J. Dongarra, B. S. Garbow, Y. Ikebe, V. C. Klema, and C. B. Moler. Matrix Eigensystem Routines—EISPACK Guide, volume 6 of Lecture Notes in Computer Science, Editors: G. Goos and J. Hartmanis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 0-387-06710-8. vii + 387 pp. LCCN QA193 .M37, QA267.A1 L43 no.6.

Samuel:2022:DNC

[SBDR22] Binny M. Samuel, Hillol Bala, Sherae L. Daniel, and V. Ramesh. Deconstructing the nature of collaboration in organizations open source software development: The impact of developer and task characteristics. IEEE Transactions on Software Engineering, 48(10):3969–3987, October 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Sanches:2010:SMR

[SBM^{+}10] Eduardo Antônio Sanches, Robie Allan Bombardelli, Ronan Maciel Marcos, Giovano Neumann, César Pereira Rebechi De Toledo, and Elizabeth Romagosa. Sperm motility of Rhamdia quelen studied using computer-assisted analysis by open-source software. Aquaculture Research, 42(1):153–156, 2010. ISSN 1355-557X (print), 1365-2109 (electronic).

Sajedi-Badashian:2020:VTB

[SBS20] Ali Sajedi-Badashian and Eleni Stroulia. Vocabulary and time based bug-assignment: a recommender system for opensource projects. Software—Practice and Experience, 50(8): 1539–1564, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1988:ILL

[SC88a] W. W. Smith and R. H. Campbell. Introduction to leif language descriptions. Technical Report UIUCDCS-R-88-1444, University of Illinois at Urbana-Champaign, Urbana-Champaign, IL, USA, July 1988. 78 pp.

Stallman:1988:TGD

[SC88b] Richard Stallman and Robert J. Chassell. Texinfo: the GNU documentation format. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1.1 edition, May 1988. vi + 102 pp.

Storer:2000:DPD

[SC00] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2000: proceedings, Data Compression Conference, March 28–30, 2000, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838139. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Schmitz:2002:POS

[SC02] Patrice-Emmanuel Schmitz and Sébastien Castiaux. Pooling open source software: An IDA feasibility study: Interchange of data between administrations. Technical report, European Commission, DG Enterprise, June 2002. 147 pp. URL http:

//europa.eu.int/ISPO/ida/export/files/en/1115.pdf.

Schreppers:2008:ACC

[SC08] Walter Schreppers and Annie Cuyt. Algorithm 871: a C/ C++ precompiler for autogeneration of multiprecision programs. ACM Transactions on Mathematical Software, 34(1): 5:1–5:20, January 2008. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sibaev:2016:PFO

[SC16] Marat Sibaev and Deborah L. Crittenden. PyVCI: a flexible open-source code for calculating accurate molecular infrared spectra. Computer Physics Communications, 203(??):290– 297, June 2016. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465516300418.

Scacchi:2004:FOS

[Sca04] Walt Scacchi. Free and Open Source development practices in the game community. IEEE Software, 21(1):59–66, January/ February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Scacchi:2005:OBE

[Sca05] Walt Scacchi. OpenEC/B: electronic commerce and free/open source software development. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scacchi:2006:UOS

[Sca06] Walt Scacchi. Understanding Open Source Software Evolution, chapter 9, pages 181–205. Wiley, New York, NY, USA, 2006. ISBN 0-470-87180-6, 0-470-87181-4, 0-470-87182-2 (online). LCCN QA76.76.D47 S66135 2006.

Scanlon:2019:CFO

[Sca19] T. P. Scanlon. Critical factors for open source advancement in the U.S. Department of Defense. IEEE Software, 36(6): 29–33, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sim:2004:LRO

[SCB04] Deborah Sim, Monifa Clayton, and Brian Boyer. Legal research on open source software. World-Wide Web document.,

2004. URL http://www.daviddfriedman.com/Academic/
Course_Pages/21st_century_issues/21st_century_law/
open_source_legal_01.htm.

Siotto:2015:APS

[SCDS15] Eliana Siotto, Marco Callieri, Matteo Dellepiane, and Roberto Scopigno. Ancient polychromy: Study and virtual reconstruction using open source tools. Journal on Computing and Cultural Heritage (JOCCH), 8(3):16:1–16:??, May 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Smith:2006:ABS

[SCFR06] Neil Smith, Andrea Capiluppi, and Juan Fernández-Ramil. Agent-based simulation of open source evolution. Software Process: Improvement and Practice, 11(4):423–434, July 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Schmidt:1990:GPH

[Sch90a] D. C. Schmidt. gperf: a perfect hash function generator. In Anonymous [Ano90d], pages 87–101.

Schumacker:1990:UIS

[Sch90b] F. Schumacker. User-friendly interface with a set of LOTOS tools. Bulletin Scientifique de l’Association des Ingénieurs Electriciens sortis de l’Institut Electrotechnique Montefiore, 103(1):3–20, 1990. CODEN BURMA2. ISSN 0302-2676.

Schurr:1991:PTH

[Sch91a] A. Schurr. PROGRESS-editor: a text-oriented hybrid editor for programmed graph rewriting systems. In Ehrig et al. [EKR91], page 67. ISBN 0-387-54478-X (New York), 3-540-54478-X (Berlin). LCCN QA75.5.G72 1991.

Stephens:1991:ILP

[SCH^{+}91b] C. Stephens, B. Cogswell, J. Heinlein, G. Palmer, and J. P. Shen. Instruction level profiling and evaluation of the IBM RS/6000. ACM SIGARCH Computer Architecture News, 19 (3):180–189, May 1991. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Schiele:2003:BUC

[Sch03] Robert Schiele. Building and using a cross development tool chain. In Hutton et al. [HDR03], pages 213–222. ISBN ????

LCCN ???? URL http://www.linux.org.uk/~ajh/gcc/
gccsummit-2003-proceedings.pdf.

Schwarz:2004:AOS

[Sch04] Brett Schwarz. Asterisk open-source PBX system. Linux Journal, 2004(118):6, February 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Schneider:2009:TFO

[Sch09] K. G. Schneider. The thick of the fray: Open source software in libraries in the first decade of this century. Bulletin of the American Society for Information Science and Technology, 35 (2):15–19, 2009. ISSN 2373-9223.

Schonberg:2010:OHG

[Sch10] E. Schonberg. Origins and history of GNAT. Ada User Journal, 31(1):42–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Schryen:2011:OSS

[Sch11] Guido Schryen. Is open source security a myth? Communications of the ACM, 54(5):130–140, May 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schoettle:2019:OSL

[Sch19] H. Schoettle. Open source license compliance — why and how? Computer, 52(8):63–67, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:2005:SOS

[SCR05] Neil Smith, Andrea Capiluppi, and Juan F. Ramil. A study of open source software evolution data using qualitative simulation. Software Process: Improvement and Practice, 10(3):287– 300, July 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Serrano:2004:POS

[SCSC04] Nicolas Serrano, Sonia Calzada, Jose Mari Sarriegui, and Ismael Ciordia. From proprietary to Open Source tools in information systems development. IEEE Software, 21(1):56–58, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

SM-D:2016:BRA

[SD16] SM-D. Book review: Automating Open Source Intelligence [edited by Robert Layton and Paul Watters, Syngress. ISBN 978-0-12-802916-9]. Network Security, 2016(3):4, March 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1353485816300241.

Stewart:2005:OPD

[SDD05] Katherine J. Stewart, David P. Darcy, and Sherae L. Daniel. Observations on patterns of development in open source software projects. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stewart:2006:OCA

[SDD06] Katherine J. Stewart, David P. Darcy, and Sherae L. Daniel. Opportunities and challenges applying functional data analysis to the study of open source software evolution. Statistical Science, 21(2):167–178, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1154979819.

Sun:2009:OSD

[SDeaK^{+}09] Y. Sun, P. Davis, e. a. Kosmacek, F. Ianzini, and M. A. Mackey. An open-source deconvolution software package for 3-D quantitative fluorescence microscopy imaging. Journal of Microscopy, 236(3):180–193, 2009. ISSN 0022-2720 (print), 1365-2818 (electronic).

Sturmberg:2016:EOS

[SDL^{+}16] Björn C. P. Sturmberg, Kokou B. Dossou, Felix J. Lawrence, Christopher G. Poulton, Ross C. McPhedran, C. Martijn de Sterke, and Lindsay C. Botten. EMUstack: an open source route to insightful electromagnetic computation via the Bloch mode scattering matrix method. Computer Physics Communications, 202(??):276–286, May 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465516000035.

Searls:1999:OSV

[Sea99] Doc Searls. Open sources: Voices from the open source revolution. Linux Journal, 63:??, July 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.com/lj-issues/3415.html.

Searls:2002:LSO

[Sea02] Doc Searls. Linux for suits: Open-source radio. Linux Journal, 93:54–56, 58–59, January 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/issue93/article.php?sid=
5571.

Searls:2004:DIH

[Sea04] Doc Searls. DIY-IT: How Linux and open source are bringing do-it-yourself to information technology. Linux Journal, 2004 (118):4, February 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Sechrest:1995:OGS

[Sec95] John Sechrest. Opinion: Gnu and Solaris. ;login: the USENIX Association newsletter, 20(4):5–??, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Sedallian:2002:GRL

[Séd02] Valérie Sédallian. Garanties et responsabilités dans les logiciels libres. (French) [Guarantees and responsibilities in Free Software]. World-Wide Web document., September 1, 2002.

URL http://www.juriscom.net/pro/2/da20020901.pdf.

Sabbah:2005:I

[SF05] Danny Sabbah and Daniel Frye. Introduction. IBM Systems Journal, 44(2):??, ???? 2005. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/
sj/442/intro.pdf.

Stol:2015:ISA

[SF15] Klaas-Jan Stol and Brian Fitzgerald. Inner source — adopting open source development practices in organizations: A tutorial. IEEE Software, 32(4):60–67, July/August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (elec-

tronic). URL http://www.computer.org/csdl/mags/so/
2015/04/mso2015040060-abs.html.

Scacchi:2006:GEU

[SFF^{+}06] Walt Scacchi, Joseph Feller, Brian Fitzgerald, Scott Hissam, and Karim Lakhani. Guest editorials: Understanding free/ open source software development processes. Software Process: Improvement and Practice, 11(2):95–105, March 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Soeken:2012:ROS

[SFWD12] Mathias Soeken, Stefan Frehse, Robert Wille, and Rolf Drechsler. RevKit: An open source toolkit for the design of reversible circuits. Lecture Notes in Computer Science, 7165: 64–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-29517-1_6/.

Stallman:1992:ASP

[SG92] R. Stallman and S. Garfinkel. Against software patents. Communications of the ACM, 35(1):17, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Segura:1999:EGT

[SG99] J. Segura and A. Gil. ELF and GNOME: Two tiny codes to evaluate the real zeros of the Bessel functions of the first kind for real orders. Computer Physics Communications, 117(3):250–262, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001933.

Sedlmaier:2005:RBR

[SG05] Roman Sedlmaier and Jan Gigerich. Rechtliche Bedingungen und Risiken der Landeshauptstadt München für den Einsatz von Open-Source Software. (German) [Legal conditions and risks for the City of Munich from the use of Open-Source software]. World-Wide Web document., 2005. URL http://

www.jurpc.de/aufsatz/20050010.htm.

Stewart:2006:MRD

[SG06] Katherine J. Stewart and Sanjay Gosain. The moderating role of development stage in free/open source software project performance. Software Process: Improvement and Practice, 11 (2):177–191, March 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Spinellis:2012:OAO

[SG12] Diomidis Spinellis and Vaggelis Giannikas. Organizational adoption of open source software. The Journal of Systems and Software, 85(3):666–682, March 2012. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121211002512.

Swan:2000:TSG

[SGD00] Tom Swan, Brian Gill, and Erik Dafforn. Tom Swan’s GNU C++ for Linux. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2153-8. xii + 831 pp. LCCN QA76.73.C153 S93 2000. Includes CD-ROM with Mandrake-Linux 6.0, X Class library for C, and the entire source code from the book.

Skibilski:2005:RDB

[SGD05a] Przemysław Skibiński, Szymon Grabowski, and Sebastian Deorowicz. Revisiting dictionary-based compression. Software— Practice and Experience, 35(15):1455–1476, December 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Skibinski:2005:RDB

[SGD05b] Przemysław Skibiński, Szymon Grabowski, and Sebastian Deorowicz. Revisiting dictionary-based compression. Software— Practice and Experience, 35(15):1455–1476, December 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schulz:2008:OSO

[SGM^{+}08] Martin Schulz, Jim Galarowicz, Don Maghrak, William Hachfeld, David Montoya, and Scott Cranford. Open — SpeedShop: an open source infrastructure for parallel performance analysis. Scientific Programming, 16(2–3):105–121, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Slogsnat:2008:OSH

[SGNB08] David Slogsnat, Alexander Giese, Mondrian Nüssle, and Ulrich Brüning. An open-source HyperTransport core. ACM Transactions on Reconfigurable Technology and Systems, 1(3): 14:1–14:??, September 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Sojer:2011:LRA

[SH11] Manuel Sojer and Joachim Henkel. License risks from ad hoc reuse of code from the Internet. Communications of the ACM, 54(12):74–81, December 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Suguitan:2019:BHO

[SH19] Michael Suguitan and Guy Hoffman. Blossom: a handcrafted open-source robot. ACM Transactions on Human–Robot Interaction (THRI), 8(1):2:1–2:27, March 2019. CODEN ???? ISSN 2573-9522 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3310356.

Shapiro:1995:PAP

[Sha95] Jim Shapiro. Prototyping algorithms in perl. Linux Journal, 16:??, August 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL ftp://ftp.ssc.

com/pub/lj/issue16/perls.pl.gz;   http://noframes.
linuxjournal.com/lj-issues/issue16/;   http://noframes.
linuxjournal.com/lj-issues/issue16/perls.pl.txt.

Sharma:2003:OSS

[Sha03] Naveen N. S. Sharma. Optimal stack slot assignment in GCC. In Hutton et al. [HDR03], pages 223–228. ISBN ????

LCCN ???? URL http://www.linux.org.uk/~ajh/gcc/
gccsummit-2003-proceedings.pdf.

Sharma:2004:AMS

[Sha04] Naveen Sharma. Addressing mode selection in GCC. In Hutton et al. [HDR04], pages 141–148. ISBN ???? LCCN

???? URL http://people.redhat.com/lockhart/.gcc04/
MasterGCC-2side.pdf.

Shaffer:2005:MSS

[Sha05] Kris Shaffer. Make stunning Schenker graphs with GNU Lilypond. Linux Journal, 2005(140):??, December 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Shatnawi:2010:QIA

[Sha10] R. Shatnawi. A quantitative investigation of the acceptable risk levels of object-oriented metrics in open-source systems. IEEE Transactions on Software Engineering, 36(2):216–225, March/April 2010. CODEN IESEDJ. ISSN 0098-5589 (print),

1939-3520 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=5383377.

Scott:2020:CAC

[SHB^{+}20] N. S. Scott, A. Hibbert, J. Ballantyne, S. Fritzsche, A. L. Hazel, D. P. Landau, D. W. Walker, and Z. Was. CPC’s 50th anniversary: Celebrating 50 years of open-source software in computational physics. Computer Physics Communications, 252(??):Article 107269, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520300886.

Sheppo:1987:ERG

[She87] Marion Sheppo. EMACS reference guide. Prime Computer, Inc., Natick, MA, USA, 2nd for release 21.0 edition, 1987. various pp.

Sheets:2001:WGA

[She01] John R. Sheets. Writing Gnome Applications. Addison-Wes-ley, Reading, MA, USA, 2001. ISBN 0-201-65791-0. xxv + 449 pp. LCCN QA76.76.D47 S4885 2000.

Sheridan:2007:PTC

[She07] Flash Sheridan. Practical testing of a C99 compiler using output comparison. Software—Practice and Experience, 37 (14):1475–1488, November 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shippy:2003:PGT

[Shi03] P. J. Shippy. Porting the Gnat tasking runtime system to the Java Virtual Machine. Literary Review, 47(2):119–120, 2003. CODEN ???? ISSN 0024-4589.

Shirazi:2012:FOS

[Shi12] Farid Shirazi. Free and Open Source Software versus Internet content filtering and censorship: a case study. The Journal of Systems and Software, 85(4):920–931, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121211002883.

Steinbeck:2003:CDK

[SHK^{+}03] C. Steinbeck, Y. Han, S. Kuhn, O. Horlacher, E. Luttmann, and E. Willighagen. The Chemistry Development Kit (CDK):

An open-source Java library for chemo- and bioinformatics. Journal of Chemical Information and Computer Sciences, 43 (2):493–500, 2003. CODEN ???? ISSN 0095-2338 (print), 1520-5142 (electronic).

Snyder:1997:TUS

[SHN97] Garth Snyder, Trent R. Hein, and Evi Nemeth. Tools for UNIX system administrators. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-665431-2. LCCN QA76.76.O63 T45 1997. Companion CD-ROM to UNIX system administration handbook, 2nd ed..

Steele:1993:FCE

[SHS^{+}93] Guy L. Steele, Jr., Danny Hillis, Richard Stallman, Gerald J. Sussman, Marvin Minsky, John McCarthy, John Backus, Fernando Corbató, and Ronald E. Anderson. Forum: Code of ethics reconsidered. Communications of the ACM, 36(7):17– 20, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Singh:2021:DOS

[SHW^{+}21] Vijay Singh, Uthpala Herath, Benny Wah, Xingyu Liao, Aldo H. Romero, and Hyowon Park. DMFTwDFT: an opensource code combining Dynamical Mean Field Theory with various density functional theory packages. Computer Physics Communications, 261(??):Article 107778, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030388X.

Sibidanov:2017:RSB

[Sib17] Alexei Sibidanov. A revision of the subtract-with-borrow random number generators. Computer Physics Communications, 221(??):299–303, December 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465517302916.

Sidwell:2003:HGB

[Sid03] Nathan Sidwell. How to get the best from g++. In Hutton et al. [HDR03], pages 229–242. ISBN ????

LCCN ???? URL http://www.linux.org.uk/~ajh/gcc/
gccsummit-2003-proceedings.pdf.

Sidwell:2004:STT

[Sid04] Nathan Sidwell. Statically typed trees in GCC. In Hutton et al. [HDR04], pages 149–168. ISBN ???? LCCN

???? URL http://people.redhat.com/lockhart/.gcc04/
MasterGCC-2side.pdf.

Siepmann:1999:LHF

[Sie99] Jürgen Siepmann. Lizenz- und haftungsrechtliche Fragen bei der kommerziellen Nutzung Freier Software. (German) [License and legal claim questions from commercial use of Free Software]. World-Wide Web document., 1999. URL http://

www.jurpc.de/aufsatz/19990163.htm.

Siepmann:2004:FSR

[Sie04] Jürgen Siepmann. Freie Software — Rechtsfreier Raum? Rechtssicherheit im Umgang mit Open Source Software. (German) [Free Software — no legal space? legal security in association with Open Software software]. World-Wide

Web document., 2004. URL http://www.ostc.de/FreieSW-
RechtsfreierRaum.pdf.

Sifry:2000:OSG

[Sif00] David Sifry. Open-source group calendaring: GCTP and OpenFlock. In USENIX [USE00a], page ?? ISBN 1-880446-17-0. LCCN ???? URL http://www.usenix.org/

publications/library/proceedings/als2000/sifry.html.

Shihab:2013:SRO

[SIK^{+}13] Emad Shihab, Akinori Ihara, Yasutaka Kamei, Walid M. Ibrahim, Masao Ohira, Bram Adams, Ahmed E. Hassan, and Ken ichi Matsumoto. Studying re-opened bugs in open source software. Empirical Software Engineering, 18(5):1005–1042, October 2013. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10664-012-9228-6.

Silic:2013:DUO

[Sil13] Mario Silic. Dual-use open source security software in organizations — dilemma: Help or hinder? Computers & Security, 39 (part B)(??):386–395, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404813001326.

Simpson:2000:CVC

[Sim00] Michael Simpson. CVS version control and branch management. Dr. Dobb’s Journal of Software Tools, 25(10):108, 110– 114, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Simon:2005:VOS

[Sim05] K. D. Simon. The value of open standards and open-source software in government environments. IBM Systems Journal, 44(2):227–238, ???? 2005. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/
sj/442/simon.pdf.

Simon:2012:IOS

[Sim12] Phil Simon. Implementing open source software. In Steve King, editor, The Next Wave of Technologies, chapter 16, pages 285–303. Wiley, New York, NY, USA, 2012. ISBN 1-119-19994-8.

Singh:2008:MTS

[Sin08] Vandana Singh. Modes of technical support in open source software — peer to peer digital knowledge creation, sharing and re-use. Proceedings of the American Society for Information Science and Technology, 45(1):1–3, 2008. ISSN 2373-9231.

Singh:2010:SWE

[Sin10a] Param Vir Singh. The small-world effect: The influence of macro-level properties of developer collaboration networks on open-source project success. ACM Transactions on Software Engineering and Methodology, 20(2):6:1–6:??, August 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Singh:2010:CTS

[Sin10b] Vandana Singh. Comparison of technical support for open source software versus proprietary software. Proceedings of the American Society for Information Science and Technology, 47 (1):1–2, 2010. ISSN 2373-9231.

Sailer:2005:BMB

[SJV^{+}05] Reiner Sailer, Trent Jaeger, Enriquillo Valdez, Ramon Caceres, Ronald Perez, Stefan Berger, John Linwood Griffin, and Leendert van Doorn. Building a MAC-based security architecture for the Xen open-source hypervisor. In IEEE [IEE05], pages 276–285. ISBN 0-7695-2461-3. ISSN 1063-

9527. LCCN L787.5. URL http://ieeexplore.ieee.org/
servlet/opac?punumber=10467. EEE Computer Society Or-

der Number P2461.

Steiner:2022:DOS

[SJW22] Lukas Steiner, Matthias Jung, and Norbert Wehn. DRAM-Sys4.0: an open-source simulation framework for in-depth DRAM analyses. International Journal of Parallel Programming, 50(2):217–242, April 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL https://

link.springer.com/article/10.1007/s10766-022-00727-
4.

Shankar:2004:COS

[SK04] K. S. Shankar and Helmut Kurth. Certifying open source— the Linux experience. IEEE Security & Privacy, 2(6):28– 33, November/December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://csdl.

computer.org/dl/mags/sp/2004/06/j6028.htm;   http:
//csdl.computer.org/dl/mags/sp/2004/06/j6028.pdf.

Singh:2012:ERM

[SK12] Satwinder Singh and K. S. Kahlon. Effectiveness of refactoring metrics model to identify smelly and error prone classes in open source software. ACM SIGSOFT Software Engineering Notes, 37(2):1–11, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smiliotopoulos:2023:DLM

[SKB23] Christos Smiliotopoulos, Georgios Kambourakis, and Konstantia Barbatsalou. On the detection of lateral movement through supervised machine learning and an open-source tool to create turnkey datasets from Sysmon logs. International Journal of Information Security, 22(6):1893–1919, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270

(electronic). URL https://link.springer.com/article/
10.1007/s10207-023-00725-8.

Siggel:2019:TOS

[SKSM19] Martin Siggel, Jan Kleinert, Tobias Stollenwerk, and Reinhold Maierl. TiGL: An open source computational geometry library for parametric aircraft design. Mathematics in Computer Science, 13(3):367–389, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Smith:2001:AOS

[SL01] Larry Smith and Cameron Laird. Android: Open-source scripting for testing and automation. Dr. Dobb’s Journal of Software Tools, 26(7):99–102, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/

ftp/2001/2001_07/android.txt.

Smith:1988:LLO

[SLC88] William W. (William Wayne) Smith, Daniel LaLiberte, and Roy Harold Campbell. The leif language oriented editor: user manual, language description manual, and installation guide. Report UIUCDCS-R-88-1444, UILU-ENG-88-1749, Dept. of Computer Science, University of Illinois at Urbana-Champaign, 1304 W. Springfield, Urbana, IL 61801, USA, July 1988. 41 + 30 + 5 pp.

Stallman:1989:GMP

[SM89a] Richard Stallman and Roland McGrath. GNU make: a program for directing recompilation. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 0.25 Beta for make, Version 3.57 Beta edition, October 24, 1989. vi + 118 pp.

Sutcliffe:1989:PCV

[SM89b] Alistair Sutcliffe and Linda Macaulay, editors. People and computers V: proceedings of the fifth conference of the British Computer Society Human-Computer Interaction Specialist Group, University of Nottingham, 5–8 September 1989. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-38430-3. LCCN QA76.9.H85 B75 1989.

Stallman:2000:GMPa

[SM00a] Richard M. Stallman and Roland McGrath. GNU Make: a Program for Directing Recompilation, for Version 3.79. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, April 04, 2000. ISBN 1-882114-80-9. ???? pp. LCCN ????

Stallman:2000:GMPb

[SM00b] Richard M. Stallman and Roland McGrath. GNU Make: a Program for Directing Recompilation, for Version 3.79. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2000. ISBN 1-882114-81-7. ???? pp. LCCN ????

Stallman:2002:GMP

[SM02] Richard M. Stallman and Roland McGrath. GNU Make: a Program for Directing Recompilation, for GNU Make Version 3.79.1. GNU Press, Boston, MA, USA, 2002. ISBN 1-882114-82-5. 196 (est.) pp. LCCN ???? US$20.00. URL http://

www.gnupress.org/book8.html.

Sohn:2008:SAS

[SM08] So Young Sohn and Min Seok Mok. A strategic analysis for successful open source software utilization based on a structural equation model. The Journal of Systems and Software, 81(6):1014–1024, June 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Smith:1990:EET

[Smi90] Norman E. Smith. Emulating EMACS with TPU. The VAX professional, 12(4):22–26, 1990. CODEN VAXPEN. ISSN 8750-9628.

Smith:2017:WWS

[Smi17] Arfon Smith. Why we should give credit to code creators. Physics World, 30(3):38–41, 2017. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

stacks.iop.org/2058-7058/30/i=3/a=37.

Shneiderman:1988:UIS

[SMNF88] Ben Shneiderman, Thomas Malone, Donald Norman, and James Foley. User interface strategies ’88 (videotape), 1988. US$1,800.00. From Computing Reviews: “User interface strategies ’88 was a two-day satellite TV course, taught October 5 and 12, 1988, and organized by Ben Shneiderman. The course features four outstanding researchers in humancomputer interaction: Ben Shneiderman, Thomas W. Malone, Donald A. Norman, and James D. Foley. All four speakers are not only leading researchers in their respective areas, but also excellent communicators. This package consists of 10 hours of videotape (eight hours of lectures and two hours of discussion) and four books of supplementary materials. These materials consist of more than 400 pages and contain all the transparencies used in the presentations, annotated bibliographies and relevant papers (except for Malone’s area), and a transcript of Norman’s lectures. … The programming environment features the NeWS window system with pie menus, the EMACS-editor with tab windows, and a ‘pseudo-scientific visualizer’ for PostScript dictionaries.”.

Sanches:2013:SMP

[SMO^{+}13] E. A. Sanches, R. M. Marcos, R. Y. Okawara, D. Caneppele, R. A. Bombardelli, and E. Romagosa. Sperm motility parameters for steindachneridion parahybae based on open-source software. Journal of Applied Ichthyology, 29(5):1114–1122, 2013. ISSN 0175-8659 (print), 1439-0426 (electronic).

Shah:2017:SNU

[SMRM^{+}17] Jindal K. Shah, Eliseo Marin-Rimoldi, Ryan Gotchy Mullen, Brian P. Keene, Sandip Khan, Andrew S. Paluch, Neeraj Rai, Lucienne L. Romanielo, Thomas W. Rosch, Brian Yoo, and Edward J. Maginn. Software news and updates: Cassandra: an open source Monte Carlo package for molecular simulation. Journal of Computational Chemistry, 38(19):1727–1739, July 15, 2017. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Stallman:2004:GMP

[SMS04] Richard Stallman, Roland McGrath, and Paul D. Smith. GNU make: a program for directed recompilation: GNU make version 3.81. A GNU manual. GNU Press, Boston, MA, USA, 2004. ISBN 1-882114-83-3. vi + 184 pp. LCCN QA76.76.T49 S735 2004.

Scholtes:2016:ARL

[SMS16] Ingo Scholtes, Pavlin Mavrodiev, and Frank Schweitzer. From Aristotle to Ringelmann: a large-scale analysis of team productivity and coordination in Open Source Software projects. Empirical Software Engineering, 21(2):642– 683, April 2016. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10664-015-9406-4.

Smyth:1997:GGN

[Smy97] Gavin Smyth. GNAT: The GNU New York University Ada Translator: a high-quality, low-cost Ada 95 compiler. Dr. Dobb’s Journal of Software Tools, 22(12):86, 88, 89–90, 105– 107, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Schneider:2006:GCC

[SNC^{+}06] Mary D. Schneider, Nima Najand, Sana Chaker, Justin M. Pare, Julie Haskins, Sarah C. Hughes, Tom C. Hobman, John Locke, and Andrew J. Simmonds. Gawky is a component of cytoplasmic mRNA processing bodies required for early Drosophila development. Journal of Cell Biology, 174 (3):349–??, July 2006. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/174/3/349.

Stankovic:2004:MWG

[SNF04] Jan Stankovic, Markus Neteler, and Roberto Flor. Mobile wireless GRASS GIS for handheld computers running GNU/ Linux. Transactions in GIS, 8(2):225–233, 2004. ISSN 1361-1682 (print), 1467-9671 (electronic).

Schmidt:1991:CES

[SO91] Heinz W. Schmidt and Stephen M. Omohundro. CLOS, Eiffel, and Sather: a comparison. Technical Report TR-91-047, International Computer Science Institute, Berkeley, CA, September 1991.

Sorfa:2001:OSI

[Sor01] Petr Sorfa. Open source IDEs for Linux/Unix. C/C++ Users Journal, 19(3):8–??, March 2001. CODEN CCUJEX. ISSN 1075-2838.

Sorenson:2006:ERS

[Sor06] Eric Sorenson. Enterprise routing sinkholes using Linux and open source tools. ;login: the USENIX Association newsletter, 31(4):??, August 2006. CODEN LOGNEM. ISSN

1044-6397. URL https://www.usenix.org/publications/
login/august-2006-volume-31-number-4/enterprise-routing-
sinkholes-using-linux-and-open.

Stallman:1993:DGG

[SP93] Richard Stallman and Roland H. Pesch. Debugging with GDB: the GNU source-level debugger. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 4.09 for GDB version 4.9 edition, 1993. vi + 174 pp. Previous edition published under title: The GDB manual. August 1993.

Stallman:1995:DGG

[SP95] Richard Stallman and Roland H. Pesch. Debugging with GDB: the GNU source-level debugger. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 4.12, for GDB version 4.14 edition, January 1995. ISBN 1-882114-08-6. vi + 184 pp. LCCN QA76.6 .S693 1995.

Schonberg:2012:ISD

[SP12] Edmond Schonberg and Vincent Pucci. Implementation of a simple dimensionality checking system in Ada 2012. ACM SIGADA Ada Letters, 32(3):35–42, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Sibalic:2017:AOS

[SPAW17] N. Sibalić, J. D. Pritchard, C. S. Adams, and K. J. Weatherill. ARC: an open-source library for calculating properties of alkali Rydberg atoms. Computer Physics Communications, 220(??):319–331, November 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465517301972.

Silvestri:2022:ERM

[SPDQ22] Emiliano Silvestri, Alessandro Pellegrini, Pierangelo Di Sanzo, and Francesco Quaglia. Effective runtime management of tasks and priorities in GNU OpenMP applications. IEEE Transactions on Computers, 71(10):2632–2645, October 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Speichert:2001:HOS

[Spe01] Horst Speichert. Haftungsrisiko Open Source Software?. (German) [Liability risks of Open Source software?]. World-Wide

Web document., 2001. URL http://www.medienkultur-
stuttgart.de/source/frameset.htm?../thema02/2archiv/
news6/mks6OSS.htm.

Stallman:1992:UGG

[SPG92] Richard Stallman, Roland H. Pesch, and John Gilmore. Using GDB: a guide to the GNU source-level debugger. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 4.01, for GDB version 4.4 edition, 1992. vi + 172 pp.

Spindler:2003:ROS

[Spi03] Gerald Spindler. Rechtsfragen der Open Source Software. (German) [Legal questions of Open-Source software]. World-Wide Web document., 2003. URL http://www.vsi.de/

inhalte/aktuell/studie_final_safe.pdf.

Spinellis:2006:OSP

[Spi06] Diomidis Spinellis. Open source and professional advancement. IEEE Software, 23(5):70–71, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2011:CUO

[Spi11] Diomidis Spinellis. Choosing and using open source components. IEEE Software, 28(3):96, 95, May/June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2019:HSO

[Spi19] D. Spinellis. How to select open source components. Computer, 52(12):103–106, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spinellis:2021:VWC

[Spi21] Diomidis Spinellis. Viewpoint: Why computing students should contribute to open source software projects. Communications of the ACM, 64(7):36–38, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.1145/3437254.

Schmidmayer:2020:EOS

[SPLD20] Kevin Schmidmayer, Fabien Petitpas, Sébastien Le Martelot, and Éric Daniel. ECOGEN: an open-source tool for multiphase, compressible, multiphysics flows. Computer Physics Communications, 251(??):Article 107093, June 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303959.

Stallman:2000:DGG

[SPS^{+}00] Richard Stallman, Roland Pesch, Stan Shebs, et al. Debugging with GDB: The GNU Source-Level Debugger, for Version 5. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2000. ISBN 1-882114-77-9. ???? pp. LCCN ????

Stallman:2002:DGG

[SPS^{+}02] Richard M. Stallman, Roland Pesch, Stan Shebs, et al. Debugging with GDB: The GNU Source-Level Debugger. GNU Press, Boston, MA, USA, 2002. ISBN 1-882114-88-4. viii + 344 pp. LCCN QA76.9.D43 D422 2003. US$25.00. URL

http://www.gnupress.org/book7.html.

Sendin-Rana:2009:IPF

[SRGCPB^{+}09] Pablo Sendín-Raña, Francisco J. González-Castaño, Enrique Pérez-Barros, Pedro S. Rodríguez-Hernández, Felipe Gil-Castiñeira, and José M. Pousada-Carballo. Improving the performance and functionality of Mondrian open-source OLAP systems. Software—Practice and Experience, 39(3):279–298, March 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1993:HIP

[SS93] Michael J. Smith and Gavriel Salvendy, editors. Humancomputer interaction: proceedings of the Fifth International Conference on Human-Computer Interaction, Orlando, Florida, (HCI International ’93), August 8–13, 1993. Elsevier, Amsterdam, The Netherlands, 1993. ISBN 0-444-89540-X. LCCN QA76.9.I58 I593 1993. Two volumes.

Sato:2002:EOS

[SS02] Hiroki Sato and Keitaro Sekine. Experiences on an open source translation effort in Japan. In USENIX [USE02a], pages 19–26. ISBN 1-880446-02-2. LCCN QA76.76.O63

B736 2002. URL http://www.usenix.org/publications/
library/proceedings/bsdcon02/sato.html.

Spinellis:2004:GEI

[SS04] Diomidis Spinellis and Clemens Szyperski. Guest Editors’ introduction: How is Open Source affecting software development? IEEE Software, 21(1):28–33, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

tronic). URL http://csdl.computer.org/comp/mags/so/
2004/01/s1028.pdf.

Sadeghi:2005:TMS

[SS05a] Ahmad-Reza Sadeghi and Christian Stüble. Towards multilaterally secure computing platforms — with open source and trusted computing. Information Security Technical Report, 10(2):83–95, ???? 2005. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1363412705000221.

Schoblick:2005:DGL

[SS05b] Gabriele Schoblick and Robert Schoblick. Debian GNU/ Linux. bhv, Bonn, Germany, 2005. ISBN 3-8266-8151-7. 767 pp. LCCN ???? EUR22.95. Includes two CD-ROMs.

Serrano:2006:OSS

[SS06] Nicolás Serrano and José María Sarriegi. Open source software ERPs: a new alternative for an old need. IEEE Software, 23 (3):94–97, May/June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stark:2023:OSP

[SS23] Justina Stark and Ivo F. Sbalzarini. An open-source pipeline for solving continuous reaction–diffusion models in image-based geometries of porous media. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S1877750323001783.

Sowe:2008:UKS

[SSA08] Sulayman K. Sowe, Ioannis Stamelos, and Lefteris Angelis. Understanding knowledge sharing activities in free/open source software projects: an empirical study. The Journal of Systems and Software, 81(3):431–446, March 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Samoladas:2004:OSS

[SSAO04] Ioannis Samoladas, Ioannis Stamelos, Lefteris Angelis, and Apostolos Oikonomou. Open source software development should strive for even greater code maintainability. Communications of the ACM, 47(10):83–87, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

SSC:1984:SRC

[SSC93] SSC staff. SCC Reference Cards. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1984–1993. These are some good, inexpensive reference/tutorial cards on UNIX commands, Bourne shell, Korn shell, emacs, vi, C, C++, etc …. e.g. the new “UNIX System Command Summary for SVR4.2/Solaris 2.1” (ISBN: 0-916151-61-1) …. Contact Belinda Frazier (bel@ssc.com) or sales@ssc.com for more info.

Samwel:2000:LDS

[SSC^{+}00] Bart Samwel, Jiri Soukup, Glenn Crist, Evan Easton, Ron Ruble, David A. Rogers, Al Stevens, Bruce MacDonald, and Scott Venckus. Letters: Data structures as objects; real (Netscape) time; riding the XML bandwagon; porting to CE; nothing new about Open Source; Y2K worries?; version control. Dr. Dobb’s Journal of Software Tools, 25(2):12, 14, February 2000. CODEN DDJOEB. ISSN 1044-789X.

Shim:2022:DLO

[SSH22] WooChul Shim, Hyejin Shin, and Yong Ho Hwang. On data licenses for open source threat intelligence. IEEE Security & Privacy, 20(4):8–22, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sherman:2007:OST

[SSM^{+}07] William R. Sherman, Simon Su, Philip A. McDonald, Yi Mu, and Frederick Harris, Jr. Open-source tools for immersive environmental visualization. IEEE Computer Graphics and Applications, 27(2):88–91, March/April 2007. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2017:EBS

[SSP17] V. B. Singh, M. Sharma, and H. Pham. Entropy based software reliability analysis of multi-version open source software. IEEE Transactions on Software Engineering, PP(99):

1, ???? 2017. CODEN IESEDJ. ISSN 0098-5589 (print),

1939-3520 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=8081836.

Singh:2018:EBS

[SSP18] V. B. Singh, Meera Sharma, and Hoang Pham. Entropy based software reliability analysis of multi-version open source software. IEEE Transactions on Software Engineering, 44 (12):1207–1223, December 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL https://

ieeexplore.ieee.org/document/8081836/.

Sharma:2002:FCH

[SSR02] Srinarayan Sharma, Vijayan Sugumaran, and Balaji Rajagopalan. A framework for creating hybrid-open source software communities. Information Systems Journal, 12(1):7–25, 2002. ISSN 1350-1917 (print), 1365-2575 (electronic).

Snyder:2014:OSS

[SSS^{+}14] Jason Snyder, Deepan Seeralan, Shereef Sayed, Jeffery Wilson, Carl B. Dietrich, Stephen H. Edwards, and Jeffrey H. Reed. Open source software-defined radio tools for education, research, and rapid prototyping. International Journal on Software Tools for Technology Transfer (STTT), 16(1):67– 80, February 2014. CODEN ???? ISSN 1433-2779 (print),

1433-2787 (electronic). URL http://link.springer.com/
article/10.1007/s10009-012-0241-2.

Sharma:2022:UOS

[SSS22] Pankajeshwara Nand Sharma, Bastin Tony Roy Savarimuthu, and Nigel Stanger. Unearthing open source decision-making processes: a case study of Python enhancement proposals. Software—Practice and Experience, 52(10):2312–2346, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

StLaurent:2004:UOS

[St.04] Andrew M. St. Laurent. Understanding Open Source and Free Software Licensing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00581-4. 208 (est.) pp. LCCN ???? US$24.95. URL http://www.

oreilly.com/catalog/osfreesoft/book/.

Schwarz:2010:HCP

[ST10] Michael Schwarz and Yuri Takhteyev. Half a century of public software institutions: Open source as a solution to hold-up problem. Journal of Public Economic Theory, 12(4):609–639, 2010. ISSN 1097-3923 (print), 1467-9779 (electronic).

RFC0746

[Sta78a] R. Stallman. RFC 746: SUPDUP graphics extension, March

17, 1978. URL ftp://ftp.internic.net/rfc/rfc746.txt;
 https://www.math.utah.edu/pub/rfc/rfc746.txt. Sta-

tus: UNKNOWN. Not online.

Stallman:1978:SFS

[Sta78b] Richard M. Stallman. Surveyor’s forum: Structured editing with a Lisp. ACM Computing Surveys, 10(4):505–507, December 1978. CODEN CMSVAN. ISSN 0010-4892. See [San78a, San78b].

Stallman:1979:EEC

[Sta79] Richard M. Stallman. EMACS: The extensible, customizable, self-documenting display editor. Report AI Memo 519, 519A, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, June 1979. 29 pp.

Stallman:1980:EMT

[Sta80a] Richard Stallman. EMACS manual for TWENEX users: a reference manual for the extensible, customizable, selfdocumenting real-time display editor. AI memo 555, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, 1980. iv + 203 pp.

Stallman:1980:EMI

[Sta80b] Richard M. Stallman. EMACS manual for ITS users. Technical Report AI Memo 554, Massachusetts Institute of Technology, Cambridge, MA, USA, June 1980. 218 pp.

Stallman:1981:EECb

[Sta81a] R. M. Stallman. EMACS: the extensible customizable selfdocumenting display editor. In Abrahams [Abr81], pages 147– 156. ISBN 0-89791-043-5, 0-89791-050-8. LCCN QA76.7 .S54 v.16:6. Published as ACM SIGPLAN Notices, v. 16, no. 6, and ACM SIGOA newsletter, vol. 2, no. 1/2, spring/summer

Stallman:1981:EMU

[Sta81b] Richard Stallman. EMACS manual for ITS users: a reference manual for the extensible, customizable, self-documenting real-time display editor. AI memo 554 554, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, 1981. iv + 224 pp. This manual corresponds to EMACS version 162.

Stallman:1981:EMTb

[Sta81c] Richard Stallman. EMACS manual for TOPS-20 users: a reference manual for the extensible, customizable, selfdocumenting real-time display editor. Ai memo 556, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, 1981. various pp.

Stallman:1981:EMTa

[Sta81d] Richard Stallman. EMACS manual for TWENEX users: a reference manual for the extensible, customizable, selfdocumenting real-time display editor. AI memo 555, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, October 1981. iv + 236 pp. This manual corresponds to EMACS version 162.

Stallman:1981:EECa

[Sta81e] Richard M. Stallman. EMACS: the extensible, customizable self-documenting display editor. AI memo 519a, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, March 1981. 28 pp.

Stallman:1984:EEC

[Sta84] Richard M. Stallman. EMACS: The extensible, customizable, self-documenting display editor. In Interactive Programming Environments [BSS84], pages 300–325. ISBN 0-07-003885-6. LCCN QA76.6 .I5251 1984. US$34.95.

Stallman:1985:GM

[Sta85] Richard Stallman. The GNU manifesto. Dr. Dobb’s Journal of Software Tools, 10(3):30–??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Stallman:1986:GEMa

[Sta86a] Richard Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 4th, Emacs version 17 edition, February 1986. viii + 255 pp.

Stallman:1986:GEMb

[Sta86b] Richard Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 5th, Emacs version 18 for UNIX users edition, October 1986. x + 284 pp.

Stallman:1986:GEMc

[Sta86c] Richard Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 6th, Emacs version 18 for UNIX users edition, March 1986. x + 284 pp.

Stallman:1987:GEM

[Sta87] Richard Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 6th, version 18 edition, March 1987. viii + 284 pp.

Stallman:1988:GMG

[Sta88a] Richard Stallman. GDB manual: the GNU source-level debugger. Free Software Foundation, Inc., Cambridge? Mass., 2nd, GDB version 2.5 edition, February 1988. ii + 63 pp.

Stallman:1988:GEMa

[Sta88b] Richard Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 6th, Emacs version 18 edition, February 1988. x + 290 pp.

Stallman:1988:GEMb

[Sta88c] Richard Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 6th, Emacs version 18 for UNIX users edition, 1988. x + 294 pp.

Stallman:1988:TTL

[Sta88d] Richard Stallman. Termcap: the termcap library and data base. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, second edition, November 1988. ii + 66 pp.

Stallman:1988:UPG

[Sta88e] Richard M. Stallman. Using and porting GNU CC. Technical report, Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1988. Electronic mail: rms@prep.ai.mit.edu. Software also available via ANONYMOUS FTP to prep.ai.mit.edu.

Stallman:1989:GMGa

[Sta89a] R. M. Stallman. GDB manual (the GNU source-level debugger). Technical report, Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, January 1989. Third Edition, GDB version 3.1.

Stallman:1989:GMGb

[Sta89b] Richard Stallman. GDB manual: the GNU source-level debugger. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, third, GDB version 3.4 edition, October 1989. iv + 76 pp.

Stallman:1992:GEM

[Sta92a] Richard Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 7th, Emacs version 18 edition, September 1992. ISBN 1-882114-01-9. xii + 340 pp. LCCN ????

Stallman:1992:UPG

[Sta92b] Richard Stallman. Using and porting GNU CC: last updated 15 February 1992 for version 2.0. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1992. vi + 330 pp. Preliminary draft, which will change.

Stallman:1992:TMT

[Sta92c] Richard M. Stallman. The Termcap Manual: The Termcap Library and Data Base. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, second edition, December 1992. ISBN ???? ???? pp. LCCN ????

Stallman:1993:GEMa

[Sta93a] Richard Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 8th, Emacs version 19 edition, June 1993. ISBN 1-882114-02-7. xiv + 392 pp. LCCN ????

Stallman:1993:GEMb

[Sta93b] Richard M. Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 9th, Emacs version 19.19 edition, August 1993. ISBN 1-882114-03-5. xiv + 404 pp. LCCN ???? This is the official manual for GNU Emacs. It is available both as a typeset document, and online in the Emacs info system.

Stallman:1994:GEM

[Sta94] Richard Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 10th, Emacs version 19.26 edition, July 1994. ISBN 1-882114-04-3. xiv + 442 pp. LCCN ???? Includes GNU Emacs reference card.

Stallman:1995:GEM

[Sta95] Richard Stallman. GNU Emacs manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 11th, Emacs version 19.29 edition, June 1995. ISBN 1-882114-52-3. xiv + 470 pp. LCCN ???? Includes GNU Emacs reference card.

Stallman:1996:DGG

[Sta96a] Richard Stallman. Debugging with GDB: the GNU sourcelevel debugger. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, August 1996. ISBN 1-882114-09-4. vi + 188 pp. LCCN QA76.9.D43 S73 1996. For GDB version 4.16.

Stallman:1996:TAN

[Sta96b] Richard Stallman. ThinkWrap — Americans now face the threat of two years in prison for indecent network postings; it would be helpful if they could access precise rules for avoiding imprisonment. Datamation, 42(5):98–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Stallman:1996:FSD

[Sta96c] Richard M. Stallman. The free software definition. World-Wide Web document., 1996. URL http://www.gnu.org/

philosophy/free-sw.html.

Stallman:1997:GEM

[Sta97a] Richard Stallman. GNU Emacs Manual, for Version 20.1. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, thirteenth edition, 1997. ISBN 1-882114-06-X. ???? pp. LCCN ????

Stallman:1997:SDR

[Sta97b] Richard Stallman. Societal dimensions: The right to read. Communications of the ACM, 40(2):85–87, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL http://www.acm.org/pubs/citations/
journals/cacm/1997-40-2/p85-stallman/.

Stallman:1997:TMT

[Sta97c] Richard M. Stallman. The Termcap Manual: The Termcap Library and Data Base. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, third edition, 1997. ISBN 1-882114-87-6.

68 (est.) pp. LCCN ???? URL http://www.gnupress.org/
book12.html.

Stallman:1998:DGG

[Sta98a] Richard Stallman. Debugging with GDB: the GNU sourcelevel debugger. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, fifth edition, April 1998. ISBN 1-882114-75-2. vi + 194 pp. LCCN QA76.9.D43 S73 1998. For GDB version 4.17.

Stallman:1998:WFS

[Sta98b] Richard M. Stallman. Why “free software” is better than “open source”. World-Wide Web document., 1998.

URL http://www.gnu.org/philosophy/free-software-
for-freedom.html.

Stallman:1999:UPG

[Sta99] Richard M. Stallman. Using and Porting the GNU Compiler Collection, For GCC Version 2.95. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1999. ISBN 1-882114-37-X, 1-882114-38-8. x + 574 pp. LCCN ???? US$50.00.

Stallman:2000:VWW

[Sta00a] Richard Stallman. Viewpoint: Why we must fight UCITA. Communications of the ACM, 43(6):27–28, June 2000. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://en.wikipedia.org/wiki/UCITA;
http://www.acm.org/pubs/citations/journals/cacm/2000-
43-6/p27-stallman/.

Stallman:2000:UPGa

[Sta00b] Richard M. Stallman. Using and Porting GNU CC, for Version 2.95. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2000. ISBN 1-882114-38-8. ???? pp. LCCN ????

Stallman:2000:UPGb

[Sta00c] Richard M. Stallman. Using and Porting the Gnu Compiler Collection (GCC). iUniverse, Inc., 2021 Pine Lake Road, Suite 100, Lincoln, NE 68512, USA, 2000. ISBN 0-595-10035-X. 556 (est.) pp. LCCN ????

Stallman:2001:CFW

[Sta01a] Richard Stallman. Can freedom withstand e-books? Communications of the ACM, 44(3):111, March 2001. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL http://www.acm.org/pubs/citations/
journals/cacm/2001-44-3/p111-stallman/.

Stallman:2001:LSD

[Sta01b] Richard Stallman. Letters: Setting the Debian record straight. Dr. Dobb’s Journal of Software Tools, 26(4):10, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www. ddj.com/. Comments on the GNU HURD kernel, the GNU Project, and the confusion between the Open Source Movement and the Free Software Foundation, and between GNU and Linux.

Stallman:2002:FFS

[Sta02a] Richard Stallman. Forum: Free software reality v. deception. Communications of the ACM, 45(7):11–12, July 2002.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stallman:2002:GEM

[Sta02b] Richard M. Stallman. GNU Emacs Manual. GNU Press, Boston, MA, USA, fifteenth edition, 2002. ISBN 1-882114-85-X. 644 (est.) pp. LCCN QA76.76.T49.S73 2002. US$45.00.

URL http://www.gnupress.org/emacs15.html.

Stark:2002:PRQ

[Sta02c] Jacqueline Stark. Peer reviews as a quality management technique in open-source software development projects. Lecture Notes in Computer Science, 2349:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2349/23490340.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2349/23490340.pdf.

Stallman:2003:RWT

[Sta03a] Richard Stallman. A response to ‘A War on Two Fronts’. Iterations: an interdisciplinary journal of software history, 1 (1):1–4, January 13, 2003. ISSN 1541-843X. URL http:/

/www.cbi.umn.edu/iterations/stallman.html;   http:
//www.cbi.umn.edu/iterations/stallman.pdf.

Stallman:2003:UGG

[Sta03b] Richard Stallman. Using GCC: the GNU compiler collection reference manual. GNU Press, Boston, MA, USA, 2003. ISBN 1-882114-39-6. 432 pp. LCCN ????

Stallman:2003:FLM

[Sta03c] Richard M. Stallman. Forum: Legalize music sharing now. Communications of the ACM, 46(9):13, September 2003. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ian02, Les03].

Staff:2004:NTE

[Sta04a] CACM Staff. News track: Experts dispute passport plan; Parliament endorses open access; A shadow over the valley; paint by numbers; talk of the future; emotional ride. Communications of the ACM, 47(10):9–10, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stallman:2004:FSJ

[Sta04b] Richard Stallman. Free but shackled — the Java trap. World-Wide Web document, April 12, 2004. URL http://www.gnu.

org/philosophy/java-trap.html.

Stanik:2006:NOS

[Sta06] John Stanik. News 2.0: Open source gets Mac attack; oh no, not another consortium!; get that chip out of my drink. ACM Queue: Tomorrow’s Computing Today, 4(7):10, September 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stallman:2009:VWO

[Sta09] Richard Stallman. Viewpoint: Why ‘open source’ misses the point of free software. Communications of the ACM, 52(6):31– 33, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stallman:2012:PEP

[Sta12] Richard Stallman. The problems of the (earlier) Plan 9 license. Web essay., June 10, 2012. URL http://www.gnu.

org/philosophy/plan-nine.html. See [MI07] for comments

on this essay, and how its objections were resolved in newer releases of Plan 9.

Sultana:2023:CRO

[STB23] Sayma Sultana, Asif Kamal Turzo, and Amiangshu Bosu. Code reviews in open source projects: how do gender biases affect participation and outcomes? Empirical Software Engineering, 28(4):??, July 2023. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL https://link.

springer.com/article/10.1007/s10664-023-10324-9.

Stevens:1993:GQP

[Ste93] Roger T. Stevens. Get quick, professional plots with GNU-PLOT. C Users Journal, 11(6):99–??, June 1993. ISSN 0898-9788.

Stevens:1995:CP

[Ste95] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 20(8):121–??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1999:CPO

[Ste99] Al Stevens. C programming: Open Source, open mike. Dr. Dobb’s Journal of Software Tools, 24(9):94, 96–98, September 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_09/storch.zip.

Steil:2000:GOA

[Ste00a] Michael Steil. GNOME: [objektorientiert arbeiten mit X; das GNU Network Object Model Environment optimal einrichten; der gekonnte Umgang mit der Oberfläche; viele Tips und Tricks; Programmieren mit dem GTK+]. Computer-&- Literatur-Verlag, Böblingen, Germany, 2000. ISBN 3-932311-60-4. 606 pp. LCCN ???? Includes CD-ROM.

Stevens:2000:CPB

[Ste00b] Al Stevens. C programming: Building Quincy 2000. Dr. Dobb’s Journal of Software Tools, 25(11):133–134, 136, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPP

[Ste01] Al Stevens. C programming: Play Mysty for me, I’m on the QT with GCC. Dr. Dobb’s Journal of Software Tools, 26(9): 107–109, 111, September 2001. CODEN DDJOEB. ISSN

1044-789X. URL http://www.ddj.com/.

Stein:2008:CWC

[Ste08] William A. Stein. Can we create a viable free open source alternative to Magma, Maple, Mathematica and Matlab? In Jeffrey [Jef08], pages 5–6. ISBN 1-59593-904-0. LCCN ????

Steinmacher:2019:LMG

[STG19] I. Steinmacher, C. Treude, and M. A. Gerosa. Let me in: Guidelines for the successful onboarding of newcomers to open source projects. IEEE Software, 36(4):41–49, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Stoughton:1999:SRO

[Sto99] Nicholas M. Stoughton. Standard report: Open source — a standards success story? ;login: the USENIX Association newsletter, 24(2):??, April 1999. CODEN LOGNEM. ISSN

1044-6397. URL http://www.usenix.org/publications/
login/standards/25.opensource.html.

Stormann:2004:FSH

[Stö04]  Carsten Störmann.  Freie Software — haftungs- und

gewährleistungsrechtliche Fragen unter Beachtung urheberrechtlicher Gesichtspunkte. (German) [Free Software — Claims and guarantee questions with regard to copyright viewpoints]. PhD thesis, Fakultät für Wirtschaftswissensschaften, Universität Paderborn, Paderborn, Germany,

2004. 135 pp. URL http://www.ifross.de/Fremdartikel/
DiplomarbeitStoermann.pdf.

Stodden:2009:LFR

[Sto09] Victoria Stodden. The legal framework for reproducible scientific research: Licensing and copyright. Computing in Science and Engineering, 11(1):35–40, January/February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Strobel:1994:LUJ

[Str94] Stefan Strobel. LINUX: Unix für jedermann; der PC als Workstation; Installation Schritt für Schritt, Unix-Grundlagen und TCP/IP, Editieren mit dem Emacs, DOS-Programme unter LINUX, Textverarbeitung mit LaTeX 2e, Sound, Grafik, Netzwerkspiele. Vogel-Verlag, Postfach 67 40, D-8700 Würzburg, Germany, 1994. ISBN 3-8259-1330-9. 100 pp. LCCN ???? 49.00 DM; 358.00 ÖS; 49.00 Sfr.

Strauss:2002:LOS

[Str02] Daryll Strauss. Life in an open source startup, 2002. Unpublished invited talk, 2002 USENIX Annual Technical Conference, June 10–15, 2002, Monterey, California, USA.

Stading:1992:CBB

[STS92] James Stading, Tom Thompson, and Ben Smith. Creating Bit-Mapped buttons: A Windows file manager, Telnet for Macs, and dired sans emacs. BYTE Magazine, 17(10):275–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sugar:2002:GBTa

[Sug02a] David Sugar. GNU Bayonne is for telephony. Linux Journal, 100:??, August 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://www.

linuxjournal.com/article.php?sid=6077.

Sugar:2002:GBTb

[Sug02b] David Sugar. GNU Bayonne is for telephony. Embedded Linux Journal Online, August 1, 2002. URL http://www.

linuxdevices.com/articles/AT5162780509.html.

Sury:2001:OSO

[Sur01a] Ursula Sury. Open-Source ist nicht Open-Right. (German) [Open Source is not Open Right]. World-Wide Web doc-

ument., May 2001. URL http://www.svifsi.ch/revue/
pages/issues/n015/in015Mosaic.html.

Sury:2001:OSS

[Sur01b] Ursula Sury. Open-Source-Software und das schweizerische Urheberrecht. (German) [Open-Source software and Swiss copyright]. World-Wide Web document., 2001. URL http:

//www.ch-open.ch/html/events/2001/OSS-Event.pdf;
http://www.ch-open.ch/html/events/2001/oss_recht.html.

Surveyer:2004:SAO

[Sur04] J. Surveyer. Sun adds to opensource Java IDE roster: a review of NetBeans Java IDE. Application Development Trends, 11 (9):48, 2004. CODEN ???? ISSN 1073-9564.

SuSE:2001:SLP

[SuS01] SuSE. SuSE Linux 7.1 Professional: über 2.000 Programme auf CD & DVD; Kernel 2.4, GNU parted, Security, Networking, Development, XFree86 4.0.2, KDE 2.0.1 + Multimedia Tools & Games; [professionelles Linux-Betriebssystem für Inel- & AMD-PCs]. SuSE GmbH, Nürnberg, Germany, 2001. ISBN 3-934678-58-0. ???? pp. LCCN ????

Sutter:2002:ESF

[Sut02] Ed Sutter. Embedded systems firmware demystified. CMP Books, 6600 Silacci Way, Gilroy, CA 95020, USA, 2002. ISBN 1-57820-099-7. xii + 364 pp. LCCN TK7895.E42 S875 2002. Includes CD-ROM.

Stone:2003:NSF

[SV03] Adam Stone and Danna Voth. In the news: Software flaws: To tell or not to tell?; open source in the US government. IEEE Software, 20(1):70–73, January/February 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459

(electronic). URL http://dlib.computer.org/so/books/
so2003/pdf/s1070.pdf.

SilvaBorges:2019:HDD

[SV19] H. Silva Borges and M. T. Valente. How do developers promote open source projects? Computer, 52(8):27–33, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stamelos:2020:OSS

[SVAGB20] Ioannis Stamelos, Iraklis Varlamis, Dimosthenis Anagnostopoulos, and Jesus M. Gonzalez-Barahona. Open source systems: Enterprise software and solutions. The Journal of Systems and Software, 163(??):??, May 2020. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121220300236.

Schuwer:2015:IBO

[SvGH15] Robert Schuwer, Michiel van Genuchten, and Les Hatton. On the impact of being open. IEEE Software, 32(5):81–83, September/October 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http://www.computer.

org/csdl/mags/so/2015/05/mso2015050081.html.

Strom:2013:PQC

[SW13] E. Thomas Strom and Angela K. Wilson, editors. Pioneers of Quantum Chemistry, volume 1122 of ACS Symposium Series. American Chemical Society, Washington, DC, USA, January 2013. ISBN 0-8412-2716-0 (print), 0-8412-2717-9 (e-book). xi + 330 + 2 pp. LCCN QD462.A1. Chapters based upon ten of the presentations at the Symposium “Pioneers of Quantum Chemistry” held March 28, 2011, at the 241st ACS National Meeting in Anaheim, CA, USA.

Stallman:2015:GPG

[SW15] Richard Stallman and Zachary Weinberg. GCC 5.2. GNU CPP reference manual. Samurai Media Limited, Hong Kong, 2015. ISBN 988-8381-70-9. ii + 80 pp. LCCN ????

Sun:2023:DOS

[SWTC23] Jie Sun, Yiqing Wang, Baolin Tian, and Zheng Chen. detonationFoam: an open-source solver for simulation of gaseous detonation based on OpenFOAM. Computer Physics Communications, 292(??):Article 108859, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002047.

Safford:2005:TCO

[SZ05] David Safford and Mimi Zohar. Trusted computing and open source. Information Security Technical Report, 10(2):74–82, ???? 2005. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-

605X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1363412705000191.

Schulte:1997:AIS

[SZAB97] M. J. Schulte, V. Zelov, A. Akkas, and J. C. Burley. Adding interval support to the GNU Fortran compiler. Technical report, Lehigh University, Bethlehem, PA, USA,

September 1997. URL http://home.ku.edu.tr/~ahakkas/
publications/Add_Intv_Sup_Comp.pdf.

Schulte:1998:SAP

[SZAB98] M. J. Schulte, V. Zelov, A. Akkas, and J. C. Burley. Summary of accomplishments and plans for future research on the interval enhanced GNU Fortran compiler. Technical report, Lehigh University, Bethlehem, PA, USA, June 1998. URL http://

home.ku.edu.tr/~ahakkas/publications/Summary.pdf.

Schulte:1999:IEG

[SZAB99] Michael J. Schulte, Vitaly Zelov, Ahmet Akkas, and James Craig Burley. The interval-enhanced GNU Fortran compiler. In Csendes [Cse99], pages 311–321. ISBN 0-7923-6057-5. LCCN QA76.9.E94 D48 1999.

Thierry:2015:IDO

[TACA15] Bertrand Thierry, Xavier Antoine, Chokri Chniti, and Hasan Alzubaidi. μ-diff: an open-source Matlab toolbox for computing multiple scattering problems by disks. Computer Physics Communications, 192(??):348–362, July 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465515001125.

Taibi:2013:ROS

[Tai13] Fathi Taibi. Reusability of open-source program code: a conceptual model and empirical investigation. ACM SIGSOFT Software Engineering Notes, 38(4):1–5, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tange:2011:GPC

[Tan11a] Ole Tange. GNU Parallel: The command-line power tool. ;login: the USENIX Association newsletter, 36(1):??, February 2011. CODEN LOGNEM. ISSN 1044-6397. URL https:

//www.usenix.org/publications/login/february-2011-
volume-36-number-1/gnu-parallel-command-line-power-
tool.

Tange:2011:GSC

[Tan11b] Ole Tange. GNU SQL: a command line tool for accessing different databases using DBURLs. ;login: the USENIX Association newsletter, 36(2):??, April 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.usenix.

org/publications/login/april-2011-volume-36-number-
2/gnu-sql-command-line-tool-accessing-different.

Taylor:1999:BI

[Tay99] Ian Lance Taylor. BFD Internals. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1999. ISBN ???? ???? pp. LCCN ????

Taylor:2000:VIM

[Tay00] Dean Taylor. VARs: Increasing margins through free software. Linux Journal, 76:??, August 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Taylor:2019:BFO

[Tay19] Michael B. Taylor. Bootstrapping a future of open source, specialized hardware: technical perspective. Communications of the ACM, 62(12):78, December 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:

//cacm.acm.org/magazines/2019/12/241057/fulltext.

Turk:2005:VLS

[TB05] D. Turk and J. Bausch. Virtual Linux servers under z/VM: Security, performance, and administration issues. IBM Systems Journal, 44(2):341–??, ???? 2005. CODEN IBMSA7. ISSN 0018-8670.

Tanny:2015:SIA

[TBPS15] S. Tanny, J. Bogue, E. Parsai, and N. Sperling. SU-E-T-465: Implementation of an automated collision detection program using open source software for the pinnacle treatment planning system. Medical Physics, 42(6Part19):3441, 2015. CO-DEN MPHYA6. ISSN 2473-4209.

Tromeur-Dervout:2011:PCF

[TDBEE11] Damien Tromeur-Dervout, Gunther Brenner, David R. Emerson, and Jocelyne Erhel, editors. Parallel Computational Fluid Dynamics 2008: Parallel Numerical Methods, Software Development and Applications, volume 74 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSA6. ISBN 3-642-14437-3 (print), 3-642-14438-1 (e-book). ISSN 1439-7358 (print), 2197-7100 (electronic).

LCCN ???? URL http://link.springer.com/book/10.
1007/978-3-642-14438-7;   http://www.springerlink.
com/content/978-3-642-14438-7. Proceedings of the twen-

tieth meeting, Parallel CFD 2008, held May 19–22, 2008 in Lyon, France.

Tentner:1993:HPC

[Ten93] A. M. Tentner, editor. High Performance Computing Symposium 1993. Grand Challenges in Computer Simulation. Proceedings of the 1993 Simulation Multiconference on the High Performance Computing Symposium. SCS, San Diego, CA, USA, 1993.

Teodorovici:2013:BRU

[Teo13] Vasile G. Teodorovici. Book review: Ubuntu unleashed, 2013 edition by Matthew Helmke with Andrew Hudson and Paul Hudson. ACM SIGSOFT Software Engineering Notes, 38(4): 41, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Terpstra:2000:OSC

[Ter00] John H. Terpstra. Open source community. In ACM [ACM00], page 96. ISBN ???? LCCN QA76.88. URL http://www.

sc2000.org/proceedings/info/fp.pdf.

Turchet:2021:EAO

[TF21] Luca Turchet and Carlo Fischione. Elk audio OS: an open source operating system for the Internet of Musical Things. ACM Transactions on Internet of Things (TIOT), 2(2):12:1– 12:18, May 2021. CODEN ???? ISSN 2691-1914 (print),

2577-6207 (electronic). URL https://dl.acm.org/doi/10.
1145/3446393.

Tiemann:1999:BGW

[TG99] Michael Tiemann and Byron Gillespie. The brave GNU world: Building a fast open-source compiler for IA-32 meant knowing what to change and how. Performance Computing, 17(11):25– 29, October 1999. CODEN UNRED5. ISSN 0742-3136. Discusses compiler optimization issues for the superscalar Intel Pentium II processors.

Tsalis:2015:BRH

[TG15] Nikolaos Tsalis and Dimitris Gritzalis. Book review: Hacking Web Intelligence: Open Source Intelligence and Web Reconnaissance Concepts and Techniques, Sudhanshu Chauhan, Nutan Kumar Panda, Elsevier Publications, USA (2015). Computers & Security, 55(??):113, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404815001236.

Tang:2020:OOS

[TGC^{+}20] X. Tang, E. Giacomin, B. Chauviere, A. Alacchi, and P. Gaillardon. OpenFPGA: An open-source framework for agile prototyping customizable FPGAs. IEEE Micro, 40(4):41–48, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Traub:2021:SGE

[TGC^{+}21] Jonas Traub, Philipp Marian Grulich, Alejandro Rodríguez Cuéllar, Sebastian Breß, Asterios Katsifodimos, Tilmann Rabl, and Volker Markl. Scotty: General and efficient open-source window aggregation for stream processing systems. ACM Transactions on Database Systems, 46(1):1:1– 1:46, April 2021. CODEN ATDSD3. ISSN 0362-5915 (print),

1557-4644 (electronic). URL https://dl.acm.org/doi/10.
1145/3433675.

Trinkenreich:2022:WOS

[TGS22] Bianca Trinkenreich, Marco Aurelio Gerosa, and Igor Steinmacher. Women in open source: We need to talk about it. Computer, 55(12):145–149, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trinkenreich:2022:PGE

[TGW^{+}22] Bianca Trinkenreich, Mariam Guizani, Igor Wiese, Tayana Conte, Marco Gerosa, Anita Sarma, and Igor Steinmacher. Pots of gold at the end of the rainbow: What is success for open source contributors? IEEE Transactions on Software Engineering, 48(10):3940–3953, October 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Thomas:2004:OSE

[TH04] Dave Thomas and Andy Hunt. Open Source ecosystems. IEEE Software, 21(4):89–91, July/August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

ITUG:2003:LRM

[The03] The Indian TEX Users Group. The L^{A}TEX Reference Manual. GNU Press, Boston, MA, USA, 2003. ISBN 1-882114-70-1. 250 (est.) pp. LCCN ???? US$40.00 (est.). URL http://

www.gnupress.org/book15.html.

USENIX-Board:2004:OLU

[The04] The USENIX Board of Directors. An open letter from the USENIX Association rebutting SCO’s position on open source software. ;login: the USENIX Association newsletter, 29 (2):??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/publications/login/2004-
04/openpdfs/Usenix.pdf.

Trautsch:2020:UIT

[THG20] Fabian Trautsch, Steffen Herbold, and Jens Grabowski. Are unit and integration test definitions still valid for modern Java projects? An empirical study on open-source projects. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-

1228 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0164121219301955.

Trautsch:2023:ASA

[THG23] Alexander Trautsch, Steffen Herbold, and Jens Grabowski. Are automated static analysis tools worth it? An investigation into relative warning density and external software quality on the example of Apache open source projects. Empirical Software Engineering, 28(3):??, May 2023. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10664-023-10301-2.

Thiem:1999:KAD

[Thi99] Uwe Thiem. KDE application development. New Riders Publishing, Carmel, IN, USA, 1999. ISBN 1-57870-201-1. xiii + 189 pp. LCCN QA76.76.A65 T4713 1999. URL http://www.

newriders.com/books/title.cfm?isbn=1578702011.

Thimbleby:2022:NNO

[Thi22] Harold Thimbleby. NHS number open source software: Implications for digital health regulation and development. ACM Transactions on Computing for Healthcare (HEALTH), 3(4): 42:1–42:??, October 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https://dl.acm.org/

doi/10.1145/3538382.

Thompson:1990:KLEa

[Tho90a] T. Thompson. Keynote — a language and extensible graphic editor for music. In Anonymous [Ano90b], pages 89–100.

Thompson:1990:KLEb

[Tho90b] T. Thompson. Keynote — a language and extensible graphic editor for music. Computing Systems, 3(2):331–357, Spring 1990. CODEN CMSYE2. ISSN 0895-6340.

Thorup:1992:GEF

[Tho92] Kresten Krab Thorup. GNU emacs as a front end to L^{A}TEX. TUGboat, 13(3):304–308, October 1992. ISSN 0896-3207.

Tiemann:1988:SRP

[Tie88] M. D. Tiemann. Solving the RPC problem in GNU C++. In Anonymous [Ano88c], pages 343–361.

Tiemann:1990:EHI

[Tie90] M. D. Tiemann. An exception handling implementation for C++. In Anonymous [Ano90d], pages 215–232.

Tiemann:1993:SRP

[Tie93] Michael D. Tiemann. Solving the RPC problem in GNU C++. In Waldo [Wal93], pages 217–233. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Tsoukalas:2020:TDF

[TKSC20] Dimitrios Tsoukalas, Dionysios Kehagias, Miltiadis Siavvas, and Alexander Chatzigeorgiou. Technical debt forecasting: an empirical study on open-source repositories. The Journal of Systems and Software, 170(??):??, December 2020. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121220301904.

Tavakkol:2017:CGA

[TL17] Sasan Tavakkol and Patrick Lynett. Celeris: a GPUaccelerated open source software with a Boussinesq-type wave solver for real-time interactive simulation and visualization. Computer Physics Communications, 217(??):117–127, August 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465517300784.

Tan:2014:BCO

[TLL^{+}14] Lin Tan, Chen Liu, Zhenmin Li, Xuanhui Wang, Yuanyuan Zhou, and Chengxiang Zhai. Bug characteristics in open source software. Empirical Software Engineering, 19(6):1665– 1705, December 2014. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10664-013-
9258-8.

Taura:2013:DIG

[TMM^{+}13] Kenjiro Taura, Takuya Matsuzaki, Makoto Miwa, Yoshikazu Kamoshida, Daisaku Yokoyama, Nan Dun, Takeshi Shibata, Choi Sung Jun, and Jun’ichi Tsujii. Design and implementation of GXP make — a workflow system based on make.

Future Generation Computer Systems, 29(2):662–672, February 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X11001087.

Tech:2017:BTO

[TNM17] Robin P. G. Tech, Konstanze E. K. Neumann, and Wendelin Michel. Blockchain-technologie und open-source- sensornetzwerke. In Interdisziplinäre Perspektiven zur Zukunft der Wertschöpfung, pages 93–108. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., December

2017. URL http://link.springer.com/chapter/10.1007/
978-3-658-20265-1_8.

Torvalds:1999:LE

[Tor99] Linus Torvalds. The Linux edge. Communications of the ACM, 42(4):38–39, April 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org:80/pubs/citations/journals/cacm/1999-42-4/
p38-torvalds/.

Toth:2006:EOS

[Tot06] Kal Toth. Experiences with open source software engineering tools. IEEE Software, 23(6):44–52, November/December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tamburri:2021:ECS

[TPK21a] D. A. Tamburri, F. Palomba, and R. Kazman. Exploring community smells in open-source: An automated approach. IEEE Transactions on Software Engineering, 47(3):630–652, March 2021. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Tekgul:2021:DOS

[TPK^{+}21b] Bulut Tekgül, Petteri Peltonen, Heikki Kahila, Ossi Kaario, and Ville Vuorinen. DLBFoam: an open-source dynamic load balancing model for fast reacting flow simulations in OpenFOAM. Computer Physics Communications, 267(??): Article 108073, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001855.

Tamburri:2019:DCP

[TPSZ19] Damian A. Tamburri, Fabio Palomba, Alexander Serebrenik, and Andy Zaidman. Discovering community patterns in open-source: a systematic approach and its evaluation. Empirical Software Engineering, 24(3):1369–1417, June 2019. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616

(electronic). URL http://link.springer.com/article/
10.1007/s10664-018-9659-9;   http://link.springer.
com/content/pdf/10.1007/s10664-018-9659-9.pdf.

Trans:1995:UPI

[Tra95] K. Trans. The use of planning and interpretation in developing an on- line EMACS help facility. In Aamodt and Komorowski [AK95], pages 463–467. ISBN 90-5199-221-1, 4-274-90046-0 (Ohmsha). ISSN 0922-6389 (print), 1879-8314 (electronic). LCCN Q334 .S3 1995.

Tomas:2022:DAE

[TRB22] Joana Tomás, Deolinda Rasteiro, and Jorge Bernardino. Data anonymization: an experimental evaluation using open-source tools. Future Internet, 14(6):167, May 30, 2022. CODEN ????

ISSN 1999-5903. URL https://www.mdpi.com/1999-5903/
14/6/167.

Tattar:2016:AOS

[TRM16] Prabhanjan Narayanachar Tattar, Suresh Ramaiah, and B. G. Manjunath. Appendix A: Open Source Software — An Epilogue, pages 627–630. Wiley, New York, NY, USA, 2016. ISBN 1-119-15274-7.

Troan:1996:FSSb

[Tro96a] Erik Troan. Free software solutions: Basic X programming in Python. The X Journal: Computing Technology with the X Window System, 5(6):84–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Troan:1996:FSSd

[Tro96b] Erik Troan. Free software solutions: Linux 2.0. The X Journal: Computing Technology with the X Window System, 6(2):72, 70, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http://www.sigs.com/

publications/docs/txjr/9611/txjr9611.toc.html.

Troan:1996:FSSc

[Tro96c] Erik Troan. Free software solutions: Meta widgets in Python. The X Journal: Computing Technology with the X Window System, 5(7):??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

Troan:1996:FSSa

[Tro96d] Erik Troan. Free software solutions: The Python language. The X Journal: Computing Technology with the X Window System, 5(5):96–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Troan:1997:FSS

[Tro97] Eric Troan. Free software solutions: From Python to Java. UNIX Developer, 1(1):77–78, January/February 1997. ISSN 1090-2279.

Tromey:2004:GNA

[Tro04] Tom Tromey. Gcj: the new ABI and its implications. In Hutton et al. [HDR04], pages 169–174. ISBN ???? LCCN

???? URL http://people.redhat.com/lockhart/.gcc04/
MasterGCC-2side.pdf.

Tromey:2005:WGF

[Tro05] Tom Tromey. Writing a GCC front end. Linux Journal, 2005 (133):5, May 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Tanenbaum:1988:MAS

[TSM88] Andrew S. Tanenbaum, Johan W. Stevenson, and Jost Muller. MINIX for the ATARI ST and MINIX manual for the ATARI ST. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, version 1.1 edition, 1988. ISBN 0-13-584392-8 (disks), 0-13-584434-7 (manual). LCCN QA76.76.O63. 9 computer disks.

Toral:2009:MML

[TTB09] S. L. Toral, R. Martínez Torres, and F. Barrero. Modelling mailing list behaviour in open source projects: the case of ARM Embedded Linux. J.UCS: Journal of Universal Computer Science, 15(3):648–??, ???? 2009. CODEN

???? ISSN 0948-6968. URL http://www.jucs.org/jucs_
15_3/modelling_mailing_list_behaviour.

Thain:2006:HML

[TTL06] Douglas Thain, Todd Tannenbaum, and Miron Livny. How to measure a large open-source distributed system. Concurrency and Computation: Practice and Experience, 18(15): 1989–2019, December 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tackett:1999:PCE

[TV99] Buford D. Tackett and Buddy Van Doren. Process control for error-free software: a software success story. IEEE Software, 16(3):24–29, May/June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/so1999/s3024abs.htm;   http:
//dlib.computer.org/so/books/so1999/pdf/s3024.pdf.

Toby:2013:GIG

[TV13] Brian H. Toby and Robert B. Von Dreele. GSAS-II: the genesis of a modern open-source all purpose crystallography software package. Journal of Applied Crystallography, 46(2):544– 549, 2013. CODEN JACGAR. ISSN 0021-8898 (print), 1600-5767 (electronic).

Twitty:2004:GTE

[Twi04] William B. Twitty. GNU tools for embedded development. Osborne/McGraw-Hill, Berkeley, CA, USA, 2004. ISBN 1-57820-130-6. 352 pp. LCCN ???? UK£34.00.

Trinkenreich:2022:WPO

[TWS^{+}22] Bianca Trinkenreich, Igor Wiese, Anita Sarma, Marco Gerosa, and Igor Steinmacher. Women’s participation in open source software: a survey of the literature. ACM Transactions on Software Engineering and Methodology, 31(4):81:1–81:??, October 2022. CODEN ATSMER. ISSN 1049-331X (print),

1557-7392 (electronic). URL https://dl.acm.org/doi/10.
1145/3510460.

Tan:2022:SOS

[TZ22] Xin Tan and Minghui Zhou. Scaling open source software communities: Challenges and practices of decentralization. IEEE Software, 39(1):70–75, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tagra:2022:RRB

[TZH22] Ankur Tagra, Haoxiang Zhang, and Ahmed E. Hassan. Revisiting reopened bugs in open source software systems. Empirical Software Engineering, 27(4):??, July 2022. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10664-022-10133-6.

Urbancic:2017:FOS

[UBR^{+}17] Vasja Urbančič, Richard Butler, Benjamin Richier, Manuel Peter, Julia Mason, Frederick J. Livesey, Christine E. Holt, and Jennifer L. Gallop. Filopodyan: an open-source pipeline for the analysis of filopodia. Journal of Cell Biology, 216(10): 3405–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/10/3405.

IEN013

[UCLxx] UCL. GNOME user’s guide, ???? ??, 19xx. URL http://

www.cis.ohio-state.edu/htbin/ien/ien13.html.

Udell:1989:ETE

[Ude89] J. Udell. Extensible text editors for programmers. BYTE Magazine, 14(3):197–198, 200–202, 204, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:VFS

[Ude97] Jon Udell. The value of free software — freeware gems like Apache, Linux, and Perl keep the Web spinning. BYTE Magazine, 22(12):109–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ullah:2015:SBR

[UMV15] Najeeb Ullah, Maurizio Morisio, and Antonio Vetro. Selecting the best reliability model to predict residual defects in open source software. Computer, 48(6):50–58, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/csdl/mags/co/2015/06/
mco2015060050-abs.html.

Utke:2008:OFM

[UNF^{+}08] Jean Utke, Uwe Naumann, Mike Fagan, Nathan Tallent, Michelle Strout, Patrick Heimbach, Chris Hill, and Carl Wunsch. OpenAD/F: a modular open-source tool for automatic differentiation of Fortran codes. ACM Transactions on Mathematical Software, 34(4):18:1–18:36, July 2008. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

UTACC:1977:TPR

[Uni77] University of Texas at Austin. Computation Center, Austin, TX, USA. TECO pocket reference list: DECsystem-10, 1977. 8 pp.

UniPress:1985:ESE

[Uni85a] UniPress Software, Inc. Emacs screen editor for IBM-PC/MS-DOS, 1985. 1 program file (IBM–PC) on 2 computer disks manual.

UniPress:1985:UEN

[Uni85b] UniPress Emacs newsletter, page various, 1985. UniPress Software, Edison, NJ, USA.

UniPress:1985:UESb

[Uni85c] UniPress Software, Inc., Edison, NJ, USA. UniPress EMACS screen editor: MLISP user’s guide, version 2.0 edition, 1985. various pp.

UniPress:1985:UESc

[Uni85d] UniPress Software, Inc., Edison, NJ, USA. UniPress EMACS screen editor: UNIX EMACS reference manual, version 2.0 edition, 1985. various pp.

UniPress:1985:UESa

[Uni85e] UniPress Software, Inc., Edison, NJ, USA. UniPress EMACS screen editor: user’s guide, version 2.0 edition, 1985. various pp.

UniPress:1985:UESd

[Uni85f] UniPress Software, Inc., Edison, NJ, USA. UniPress EMACS screen editor: VMS EMACS reference manual, version 2.0 edition, 1985. various pp.

UniPress:1986:UEU

[Uni86] UniPress Software, Inc., Edison, NJ, USA. UniPress Emacs user’s guide, version 2.10 and later edition, 1986. various pp.

UC:2001:EIU

[Uni01] Unicode Consortium, editor. Eighteenth International Unicode Conference (IUC18) Unicode and the Web: the Global Connection, April 24–27, 2001, Hong Kong. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.org, 2001. ISBN ???? LCCN
???? URL http://www.unicode.org/iuc/iuc18.

USENIX:1988:UPC

[USE88] USENIX Association, editor. USENIX proceedings: C++ Conference, Denver, CO, October 17–21, 1988. USENIX Association, Berkeley, CA, USA, 1988.

USENIX:1990:PSU

[USE90] USENIX Association, editor. Proceedings of the Summer 1990 USENIX Conference: June 11–15, 1990, Anaheim, California, USA. USENIX Association, Berkeley, CA, USA, 1990. LCCN QA76.8.U65 U81 1990.

USENIX:1994:PES

[USE94] USENIX Association, editor. Proceedings of the Eighth Systems Administration Conference (LISA VIII): September 19– 23, 1994, San Diego, CA, USA. USENIX Association, Berkeley, CA, USA, 1994. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

USENIX:1998:PFT

[USE98a] USENIX, editor. Proceedings of the FreeNIX Track: USENIX 1998 annual technical conference: June 15–19, 1998, New Orleans, LA. USENIX Association, Berkeley, CA, USA, 1998.

ISBN ???? LCCN ???? URL http://www.usenix.org/
publications/library/proceedings/usenix98/freenix/.

USENIX:1998:PTS

[USE98b] USENIX, editor. Proceedings of the Twelfth Systems Administration Conference (LISA XII). USENIX Association, Berkeley, CA, USA, 1998. ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL http://db.usenix.org/

publications/library/proceedings/lisa98.

USENIX:1999:UAT

[USE99] USENIX, editor. Usenix Annual Technical Conference. June 6–11, 1999. Monterey, California, USA. USENIX Association, Berkeley, CA, USA, 1999. ISBN 1-880446-33-2.

LCCN ???? URL http://db.usenix.org/publications/
library/proceedings/usenix99.

USENIX:2000:PAL

[USE00a] USENIX, editor. Proceedings of the 4th Annual Linux Showcase and Conference, Atlanta, October 10–14, 2000, Atlanta, Georgia, USA. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-17-0. LCCN ???? URL http:

//www.usenix.org/publications/library/proceedings/
als2000/.

USENIX:2000:PNU

[USE00b] USENIX, editor. Proceedings of the Ninth USENIX Security Symposium, August 14–17, 2000, Denver, Colorado. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-18-9.

LCCN ???? URL http://www.usenix.org/publications/
library/proceedings/sec2000.

USENIX:2001:PAL

[USE01a] USENIX, editor. Proceedings of the 5th Annual Linux Showcase and Conference, November 5–10, 2001, Oakland, CA. USENIX Association, Berkeley, CA, USA, 2001. ISBN ????

LCCN ???? URL http://www.linuxshowcase.org/tech.
html.

USENIX:2001:PFT

[USE01b] USENIX, editor. Proceedings of the FREENIX Track: 2001 USENIX Annual Technical Conference: June 25–30, 2001, Marriott Copley Place Hotel, Boston, Massachusetts, USA. USENIX Association, Berkeley, CA, USA, 2001. ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL http:

//www.usenix.org/publications/library/proceedings/
usenix01/freenix01/technical.html.

USENIX:2002:PBF

[USE02a] USENIX, editor. Proceedings of BSDCon 2002: February 11–14, 2002, Cathedral Hill Hotel, San Francisco, CA.

USENIX Association, Berkeley, CA, USA, 2002. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736 2002. URL http:

//www.usenix.org/publications/library/proceedings/
bsdcon02/tech.html.

USENIX:2002:PFT

[USE02b] USENIX, editor. Proceedings of the FREENIX Track: 2002 USENIX Annual Technical Conference, June 10–15, 2002, Monterey, California, USA. USENIX Association, Berkeley, CA, USA, 2002. ISBN 1-880446-01-4. LCCN QA76.8.U65

P765 2002. URL http://www.usenix.org/publications/
library/proceedings/usenix02/.

USENIX:2002:PGT

[USE02c] USENIX, editor. Proceedings of the General Track: 2002 USENIX Annual Technical Conference, June 10–15, 2002, Monterey, California, USA. USENIX Association, Berkeley, CA, USA, 2002. ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.usenix.org/publications/
library/proceedings/usenix02/.

Unger:1997:TGI

[UZ97] A. Unger and E. Zehendner. Tuning the GNU instruction scheduler to superscalar microprocessors. In Proceedings of the 23rd EUROMICRO Conference EUROMICRO 97. ‘New Frontiers of Information Technology’, pages 275–282. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN ???? ISSN ????

Vaidhyanathan:2001:CCR

[Vai01] Siva Vaidhyanathan. Copyrights and copywrongs: the rise of intellectual property and how it threatens creativity. New York University Press, New York, NY, USA, 2001. ISBN 0-8147-8806-8. xi + 243 pp. LCCN Z642 .V35 2001.

Valdes:1991:LLB

[Val91] R. Valdes. Little languages, big questions. Dr. Dobb’s Journal of Software Tools, 16(9):16–18, 20, 22, 25, September 1991. CODEN DDJSDM. ISSN 0884-5395.

Valdes:1993:TEA

[Val93] Ray Valdés. Text editors: Algorithms and architectures. Dr. Dobb’s Journal of Software Tools, 18(4):38–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Valimaki:2004:GGP

[Väl04] Mikko Välimäki. GNU General Public License and the distribution of derivative works. World-Wide Web document.,

2004. URL http://www.soberit.hut.fi/~msvalima/gpl_
derivative.pdf.

Vancsa:2022:FOO

[Van22] Ildikó Vancsa. The four opens: Open source beyond the code. Computer, 55(6):81–84, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaught:1996:GGX

[Vau96] Andy Vaught. Graphing with Gnuplot and Xmgr. Linux Journal, 28:??, August 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Vazquez:2016:NDI

[Váz16] R. Vázquez. A new design for the implementation of isogeometric analysis in Octave and Matlab: GeoPDEs 3.0. Computers and Mathematics with Applications, 72(3):523–554, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0898122116302681.

Voyiatzis:2019:OSI

[VB19] Evangelos Voyiatzis and Michael C. Böhm. An open-source implementation of a quasi harmonic approach to compute the intramolecular entropy of particle systems. Computer Physics Communications, 234(??):286–293, January 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465518302558.

Valiev:2010:NCS

[VBG^{+}10] M. Valiev, E. J. Bylaska, N. Govind, K. Kowalski, T. P. Straatsma, H. J. J. Van Dam, D. Wang, J. Nieplocha, E. Apra, T. L. Windus, and W. A. de Jong. NWChem: a comprehensive and scalable open-source solution for large scale molecular simulations. Computer Physics Communications, 181(9):1477–1489, September 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465510001438.

Vaamonde:2001:ISG

[VD01] Fernández Vaamonde and Manuel David. Implantación de un sistema de gestión centralizada de paquetes deb para su uso en configuraciones Debian GNU/Linux. (Spanish) [Implantation of a system of centralized management of deb packages for its use in configurations of Debian GNU/Linux ]. Technical report, Trabajos fin de carrera de la Facultade de Informática de A Coruña. Departamento de Electrónica e Sistemas, Facultade de Informática Universidade da Coruña, Coruña, Spain, 2001. 111 pp. Final engineering project. Includes one CD-ROM.

vanderHoeven:2004:GT

[vdH04] Joris van der Hoeven. GNU TeXmacs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):24–25, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vanderHoeven:2013:GTS

[vdHGG^{+}13] Joris van der Hoeven, Andrey Grozin, Massimiliano Gubinelli, Grégoire Lecerf, François Poulain, and Denis Raux. GNU TEXmacs: a scientific editing platform. ACM Communications in Computer Algebra, 47(1/2):59–61, March 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

vanderLinden:2009:OSD

[vdLLM09] Frank van der Linden, Björn Lundell, and Pentti Marttiin. Open source development: Commodification of industrial software: a case for open source. IEEE Software, 26(4):77– 83, July/August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Veglis:2006:OSW

[Veg06] Andreas Veglis. Open source Web software for Windows users. IEEE Distributed Systems Online, 7(7):??, July 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.org/comp/mags/ds/2006/07/
o7005.pdf.

Vaughan:2000:GAA

[VETT00] Gary V. Vaughan, Ben Elliston, Tom Tromey, and Ian Lance Taylor. GNU Autoconf, Automake and Libtool. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 1-57870-190-2. xx + 390 pp. LCCN QA76.76.O63 G598 2000. US$40.00.

URL http://sources.redhat.com/autobook/;   http://
sources.redhat.com/autobook/autobook/autobook_toc.
html;   http://sources.redhat.com/autobook/download.
html;   http://www.newriders.com/books/title.cfm?isbn=
1578701902.

Vestal:1997:RMD

[VGD^{+}97] S. Vestal, L. Guerby, R. Dewar, D. McConnell, and B. Lewis. Reimplementing a multiprocess distributed paradigm for realtime systems in Ada 95. ACM SIGADA Ada Letters, 17(5): 93–99, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vardanega:2001:ACR

[VGdlP01] Tullio Vardanega, Rodrigo García, and Juan Antonio de la Puente. An application case for Ravenscar technology: Porting OBOSS to GNAT/ORK. Lecture Notes in Computer Science, 2043:392–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2043/
20430392.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2043/20430392.pdf.

Vassallo:2019:LSE

[VGP^{+}19] Carmine Vassallo, Giovanni Grano, Fabio Palomba, Harald C. Gall, and Alberto Bacchelli. A large-scale empirical exploration on refactoring activities in open source software projects. Science of Computer Programming, 180(??):1–15, July 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print),

1872-7964 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167642318302557.

vanGurp:2010:CPR

[vGPB10] Jilles van Gurp, Christian Prehofer, and Jan Bosch. Comparing practices for reuse in integration-oriented software product lines and large open source software projects. Software—Practice and Experience, 40(4):285–312, April 10, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanGumster:2010:GB

[vGS10] Jason van Gumster and Robert Shimonski. GIMP bible. Wiley, New York, NY, USA, 2010. ISBN 0-470-52397-2 (paperback), 0-470-63642-4 (e-book). xxxv + 722 pp. LCCN T385 .V36 2010.

Valdivia-Garcia:2018:CPB

[VGSN18] Harold Valdivia-Garcia, Emad Shihab, and Meiyappan Nagappan. Characterizing and predicting blocking bugs in open source projects. The Journal of Systems and Software, 143(??):44–58, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121218300530.

Valimaki:2004:CCC

[VH04] Mikko Välimäki and Herkko Hietanen. The challenges of creative commons licensing. World-Wide Web document., 2004.

URL http://www.hiit.fi/u/hietanen/docs/cri06-04_A_
Vaelimaeki_Hietanen.pdf.

vonHagen:2003:DGG

[vHW03] William von Hagen and Kurt Wall. The Definitive Guide to GCC. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-109-7. xxviii + 519 pp. LCCN QA76.76.C65 W36 2004.

vonHagen:2006:DGG

[vHW06] William von Hagen and Kurt Wall. The Definitive Guide to GCC. Apress, Berkeley, CA, USA, second edition, 2006. ISBN 1-59059-585-8. 584 (est.) pp. LCCN ????

Vieth:1997:GEE

[Vie97]  Ulrik Vieth. A GNU Emacs editing mode for  and

sources. TUGboat, 18(1):12–16, March 1997. ISSN 0896-3207.

Virkus:2005:PJP

[Vir05] Robert Virkus. Pro J2ME Polish: open source wireless Java tools suite. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-503-3 (hardcover). ???? pp. LCCN QA76.73.J38 V57 2005.

URL http://www.loc.gov/catdir/enhancements/fy0663/
2005016571-d.html;   http://www.loc.gov/catdir/toc/
ecip0514/2005016571.html.

Viseur:2023:COO

[VJ23] Robert Viseur and Nicolas Jullien. CommunesPlone: an original open source model of resource pooling in the public sector. IEEE Software, 40(4):46–54, 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

vonKrogh:2003:CJS

[vKSL03] Georg von Krogh, Sebastian Spaeth, and Karim R. Lakhani. Community, joining, and specialization in open source software innovation: a case study. Research Policy, 32(7):1217– 1241, 2003. CODEN ???? ISSN 0048-7333 (print), 1873-

7625 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0048733303000507.

vonKrogh:2003:SIO

[vKvH03] Georg von Krogh and Eric von Hippel. Special issue on open source software development. Research Policy, 32(7):1149– 1157, July 2003. CODEN ???? ISSN 0048-7333 (print), 1873-

7625 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0048733303000544.

Venton:2005:ULH

[VMKB05] T. Venton, M. Miller, R. Kalla, and A. Blanchard. Using Linux for hardware bring up, development, and manufacturing. IBM Systems Journal, 44(2):319–329, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/sj/442/venton.pdf.

Vijaykumar:2022:MPO

[VOK^{+}22] Nandita Vijaykumar, Ataberk Olgun, Konstantinos Kanellopoulos, F. Nisa Bostanci, Hasan Hassan, Mehrshad Lotfi, Phillip B. Gibbons, and Onur Mutlu. MetaSys: a practical open-source metadata management system to implement and evaluate cross-layer optimizations. ACM Transactions on Architecture and Code Optimization, 19(2):26:1–26:29, June 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3505250.

Volkman:1989:BGB

[Vol89] Victor R. Volkman. Bison: A GNU breed of YACC. C Users Journal, 7(8):117–??, August 1989. ISSN 0898-9788.

Volkman:1996:MG

[Vol96] Victor R. Volkman. Micro-C and GNU. C/C++ Users Journal, 14(1):88–??, January 1996. CODEN CCUJEX. ISSN 1075-2838.

Voras:2012:ECC

[VOM12] Ivan Voras, Marin Orlić, and Branko Mihaljević. An early comparison of commercial and open-source cloud platforms for scientific environments. Lecture Notes in Computer Science, 7327:164–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30947-2_20/.

vonBechtolsheim:1988:UEE

[von88] Stephan von Bechtolsheim. Using the Emacs editor to safely edit TEX sources. TEXniques, Publications for the TEX community, 7:195–202, 1988.

Voronkov:1992:LPA

[Vor92] A. Voronkov, editor. Logic Programming and Automated Reasoning. International Conference LPAR ’92 Proceedings, volume 624 of Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-55727-X (New York), 3-540-55727-X (Berlin). LCCN QA76.63.I55 1992.

Vrahatis:1995:RPP

[VRS^{+}95] M. N. Vrahatis, O. Ragos, T. Skiniotis, F. A. Zafiropoulos, and T. N. Grapsa. RFSFNS: a portable package for the numerical determination of the number and the calculation of roots of Bessel functions. Computer Physics Communications, 92(2–3):252–266, December 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001159. See erratum [VRS^{+}99b].

Vrahatis:1999:EBP

[VRS^{+}99a] M. N. Vrahatis, O. Ragos, T. Skiniotis, F. A. Zafiropoulos, and T. N. Grapsa. Erratum to: RFSFNS: a portable package for the numerical determination of the number and the calculation of roots of Bessel functions [Comput. Phys. Commun. 92 (1995) 252–266]. Computer Physics Communications, 117(3):290, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046559800109X. See [VRS^{+}95].

Vrahatis:1999:ESP

[VRS^{+}99b] M. N. Vrahatis, O. Ragos, T. Skiniotis, F. A. Zafiropoulos, and T. N. Grapsa. Erratum to: RFSFNS: a portable package for the numerical determination of the number and the calculation of roots of Bessel functions [Comput. Phys. Commun. 92 (1995) 252–266]. Computer Physics Communications, 117(3):290, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046559800109X. See [VRS^{+}95].

Vazao:2023:IEG

[VSdCCR23] Ana Paula Vazão, Leonel Santos, Rogério Luís de C. Costa, and Carlos Rabadão. Implementing and evaluating a GDPRcompliant open-source SIEM solution. Journal of Information Security and Applications (JISA), 75(??):??, June 2023. CODEN ???? ISSN 2214-2126. URL http://www.

sciencedirect.com/science/article/pii/S2214212623000935.

Vasilescu:2014:VSW

[VSGM14] Bogdan Vasilescu, Alexander Serebrenik, Mathieu Goeminne, and Tom Mens. On the variation and specialisation of workload — a case study of the Gnome ecosystem community. Empirical Software Engineering, 19(4):955–1008, August 2014. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10664-013-9244-1.

VanMeggelen:2006:ATO

[VSM06] Jim Van Meggelen, Jared Smith, and Leif Madsen. Asterisk: la téléphonie Open Source. O’Reilly, Paris, France, édition française edition, 2006. ISBN 2-84177-394-9. xvi + 414 pp. LCCN ????

Vlachos:2022:SOS

[VSN22] Vasileios Vlachos, Yannis C. Stamatiou, and Sotiris Nikoletseas. The SAINT observatory subsystem: an open-source intelligence tool for uncovering cybersecurity threats. International Journal of Information Security, 21(5):1091–1106, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270

(electronic). URL https://link.springer.com/article/
10.1007/s10207-022-00599-2.

Ven:2008:SYA

[VVM08] Kris Ven, Jan Verelst, and Herwig Mannaert. Should you adopt open source software? IEEE Software, 25(3):54–59, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Verkerk:1992:PIC

[VW92] C. Verkerk and W. Wojcik, editors. Proceedings of the International Conference on Computing in High Energy Physics ’92. Proceedings (CERN 92-07). CERN, Geneve, Switzerland, 1992.

vonWangenheim:2009:EOS

[vWHvW09] Christiane Gresse von Wangenheim, Jean Carlo Rossa Hauck, and Aldo von Wangenheim. Enhancing open source software in alignment with CMMI-DEV. IEEE Software, 26(2):59– 67, March/April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Viega:1998:MGM

[VWM98] John Viega, Barry Warsaw, and Ken Manheimer. Mailman: The GNU Mailing List Manager. In USENIX [USE98b], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63

S97 1998. URL http://www.usenix.org/publications/
library/proceedings/lisa98/viega.html.

Welsh:1995:LBG

[W^{+}95] Matt Welsh et al. The Linux bible: the GNU testament. Yggdrasil Computing, Inc., Berkeley, CA, USA, third edition, 1995. ISBN 1-883601-12-6. 1596 pp. LCCN QA 76.76 O63 L56 1995.

Waaijer:2009:HRO

[Waa09] Marlen K. Waaijer. Handbook of Research on Open Source Software: Technological, Economic, and Social Perspectives (St. Amant, K. and Still, B., Eds.; 2007) [Book Review]. IEEE Transactions on Professional Communication, 52(1):109–112, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walnes:2003:JOS

[WACBL03] Joe Walnes, Ara Abrahamian, Mike Cannon-Brookes, and Patrick A. Lightbody. Java Open source programming: with XDoclet, JUnit, WebWork, Hibernate (Java Open Source Library). Wiley, New York, NY, USA, 2003. ISBN 0-471-46362-0. xx + 459 pp. LCCN QA76.73.J38 J3785 2003.

Wagle:2003:SSB

[Wag03] Perry M. Wagle. StackGuard: Simple buffer overflow protection for GCC. In Hutton et al. [HDR03], pages 243–255.

ISBN ???? LCCN ???? URL http://www.linux.org.uk/~
ajh/gcc/gccsummit-2003-proceedings.pdf.

Waldo:1993:ECL

[Wal93] Jim Waldo, editor. The Evolution of C++: Language Design in the Marketplace of Ideas. USENIX Association and MIT Press, Berkeley, CA, USA and Cambridge, MA, USA, 1993. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993.

Wallace:1999:PWG

[Wal99] Shawn P. Wallace. Programming Web Graphics with Perl and GNU Software. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-478-9. xiv + 454 pp. LCCN QA76.73.P22 W34 1999. US$29.95.

URL http://www.oreilly.com/catalog/prowg/.

Wall:2001:ST

[Wal01] David A. E. Wall. Software technologies: Using Open Source for a profitable startup. Computer, 34(12):158–160, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-

0814 (electronic). URL http://dlib.computer.org/co/
books/co2001/pdf/rz158.pdf;   http://www.computer.
org/computer/co2001/rz158abs.htm.

Wang:2021:OSS

[Wan21] Jinyong Wang. Open source software reliability model with nonlinear fault detection and fault introduction. Journal of Software: Evolution and Process, 33(12):e2385:1–e2385:??, December 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Warkus:2004:OGD

[War04] Matthias Warkus. The official GNOME 2 developer’s guide. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-030-5. 497 pp. LCCN QA76.9.U83 W375 2004; QA76.9.U83 W375 2004eb. URL http://www.oreilly.com/

catalog/9781593270308.

Waters:1985:PAS

[Wat85a] R. C. Waters. The programmer’s apprentice: A session with KBEmacs. IEEE Transactions on Software Engineering, SE-11(11):1296–1320, November 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1701948.

Waters:1985:KST

[Wat85b] Richard C. Waters. KBEmacs: a step toward the programmer’s apprentice. Technical report AI-TR-753, Massachusetts Institute of Technology, Artificial Intelligence Laboratory., Cambridge, MA, USA, 1985. 236 pp.

Waters:1987:PAS

[Wat87] R. C. Waters. The programmer’s apprentice: a session with KBEmacs. In Anonymous [Ano87], page 1.

Waters:1994:CPE

[Wat94] R. C. Waters. Cliche-based program editors. ACM Transactions on Programming Languages and Systems, 16(1):102–150, January 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Watanabe:2001:MOS

[Wat01] Takayuki Watanabe. Merits of open-source resolution to resolve a digital divide in information technology. Lecture Notes in Computer Science, 2105:92–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2105/21050092.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2105/21050092.pdf.

Williams:2002:IVO

[WB02] Stephen Williams and Michael Baxter. Icarus Verilog: Opensource Verilog more than a year later. Linux Journal, 99: ??, July 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://www.linuxjournal.

com/article/6001.

Walfield:2007:CGH

[WB07] Neal H. Walfield and Marcus Brinkmann. A critique of the GNU Hurd multi-server operating system. Operating Systems Review, 41(4):30–39, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wiberg:1974:SNC

[WBB^{+}74] Kenneth B. Wiberg, Lawrence S. Bartell, Jacob Bigeleisen, Robert B. K. Dewar, Frank E. Harris, F. A. Matsen, Harrison Shull, and Lawrence C. Snyder. A study of a national center for computation in chemistry. Report, National Academy of Sciences, Washington, DC, USA, 1974. ix + 79 pp. URL

https://books.google.com/books?id=qTYrAAAAYAAJ. This

is a sequel to an earlier report [MSLH71].

Witten:2001:GOD

[WBB01] Ian H. Witten, David Bainbridge, and Stefan Boddie. Greenstone: Open-source DL software. Communications of the ACM, 44(5):47, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/pubs/articles/journals/cacm/2001-44-
5/p47-witten/p47-witten.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/2001-44-5/p47-witten/.

Wolf:2002:OSS

[WBGM02] Marty J. Wolf, Kevin Bowyer, Don Gotterbarn, and Keith Miller. Open source software: intellectual challenges to the status quo. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):317–318, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Wolter:2023:OSL

[WBRH23] Thomas Wolter, Ann Barcomb, Dirk Riehle, and Nikolay Harutyunyan. Open source license inconsistencies on GitHub.

ACM Transactions on Software Engineering and Methodology, 32(5):110:1–110:??, September 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/3571852.

Watson:2008:BOS

[WBY^{+}08] Richard T. Watson, Marie-Claude Boudreau, Paul T. York, Martina E. Greiner, and Donald Wynn, Jr. The business of open source. Communications of the ACM, 51(4):41–46, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wang:2014:OOG

[WCA^{+}14] Dongming Wang, Xiaoyu Chen, Wenya An, Lei Jiang, and Dan Song. OpenGeo: An open geometric knowledge base. In Hong and Yap [HY14], pages 240–245. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (e-book). LCCN QA76.9.M35.

Wang:2022:HDO

[WCG22] Wenting Wang, Jinghui Cheng, and Jin L. C. Guo. How do open source software contributors perceive and address usability?: Valued factors, practices, and challenges. IEEE Software, 39(1):76–83, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wainer:2021:MSS

[WCHRM21] Gabriel Wainer, Román Cárdenas, Kevin Henares, and Cristina Ruiz-Martín. Modeling and simulation of space-based pandemic scenarios using an open-source platform. Computing in Science and Engineering, 23(4):80–84, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Walker:2020:OST

[WCS20] Andrew Walker, Tomas Cerny, and Eungee Song. Open-source tools and benchmarks for code-clone detection: past, present, and future trends. ACM SIGAPP Applied Computing Review, 19(4):28–39, January 2020. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3381307.3381310.

Whatmough:2020:CAR

[WDK^{+}20] P. N. Whatmough, M. Donato, G. G. Ko, S. K. Lee, D. Brooks, and G. Wei. CHIPKIT: An agile, reusable opensource framework for rapid test chip development. IEEE Micro, 40(4):32–40, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Weathersby:2003:SFD

[Wea03] John M. Weathersby, Jr. Sidebar: a foot in the door: Can open source find traction in government? ACM Queue: Tomorrow’s Computing Today, 1(5):52–53, July/August 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Weber:2004:SOS

[Web04] Steve Weber. The success of open source. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01292-5. viii + 312 pp. LCCN QA76.76.S46 W43 2004. URL http://www.

loc.gov/catdir/toc/fy045/2003056916.html.

Wehr:2003:ROO

[Weh03] Kenneth Wehr. Review: openoffice.org 1.0 resource kit. Linux Journal, 2003(115):14, November 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Weinberg:2003:MPV

[Wei03] Zachary Weinberg. A maintenance programmer’s view of GCC. In Hutton et al. [HDR03], pages 257–268. ISBN ????

LCCN ???? URL http://www.linux.org.uk/~ajh/gcc/
gccsummit-2003-proceedings.pdf.

Welsh:1994:EFF

[Wel94a] Matt Welsh. Emacs: Friend or foe. Linux Journal, 5:??, September 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Welsh:1994:TEP

[Wel94b] Matt Welsh. Tutorial: Emacs for programmers. Linux Journal, 6:??, October 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Welsh:1995:ILK

[Wel95] Matt Welsh. Implementing loadable kernel modules for Linux. Dr. Dobb’s Journal of Software Tools, 20(5):18–20, 22, 24, 96, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Wendt:1990:FCG

[Wen90] Alan L. Wendt. Fast code generation using automaticallygenerated decision trees. ACM SIGPLAN Notices, 25(6): 9–15, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p9-wendt/.

Wen:2000:OSI

[Wen00] Bobby S. Wen. Open-source intrusion detection tools for Linux. Linux Journal, 78:??, October 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Wen:2002:CSO

[Wen02] Howard Wen. Crystal Space: an open-source 3-D graphics engine. Linux Journal, 97:92–94, May 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

West:2000:ADG

[Wes00] Adrian West. Advanced 3-D graphics: GNU Maverik — A VR micro-kernel. Linux Journal, 77:??, September 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

West:2003:HOO

[Wes03] Joel West. How open is open enough?: Melding proprietary and open source platform strategies. Research Policy, 32(7): 1259–1285, 2003. CODEN ???? ISSN 0048-7333 (print),

1873-7625 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0048733303000520.

Wang:2019:EOS

[WFDK19] Y. L. Wang, G. Fabbris, M. P. M. Dean, and G. Kotliar. EDRIXS: an open source toolkit for simulating spectra of resonant inelastic X-ray scattering. Computer Physics Communications, 243(??):151–165, October 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519301353.

Walter:2018:CST

[WFF18] Bartosz Walter, Francesca Arcelli Fontana, and Vincenzo Ferme. Code smells and their collocations: a large-scale experiment on open-source systems. The Journal of Systems and Software, 144(??):1–21, October 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121218301109.

Witherden:2014:POS

[WFV14] F. D. Witherden, A. M. Farrington, and P. E. Vincent. PyFR: an open source framework for solving advectiondiffusion type problems on streaming architectures using the flux reconstruction approach. Computer Physics Communications, 185(11):3028–3040, November 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465514002549.

Wang:2020:UED

[WFW^{+}20] Zhendong Wang, Yang Feng, Yi Wang, James A. Jones, and David Redmiles. Unveiling elite developers’ activities in open source projects. ACM Transactions on Software Engineering and Methodology, 29(3):16:1–16:35, July 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3387111.

Wettstein:2000:GML

[WG00] Greg Wettstein and Johannes Grosen. Gaining the middleground: a Linux-based open-source middleware initiative. In USENIX [USE00a], page ?? ISBN 1-880446-17-0.

LCCN ???? URL http://www.usenix.org/publications/
library/proceedings/als2000/wettstein.html.

Woods:2005:OSE

[WG05] Dan Woods and Gautam Guliani. Open source for the enterprise: managing risks, reaping rewards. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-10119-8. xiii + 217 pp. LCCN T58.5 .W66

West:2006:COI

[WG06] Joel West and Scott Gallagher. Challenges of open innovation: the paradox of firm investment in open-source software. R&D Management, 36(3):319–331, 2006. ISSN 0033-6807 (print), 1467-9310 (electronic).

Warren:2016:GOS

[WGG16] Craig Warren, Antonios Giannopoulos, and Iraklis Giannakis. gprMax: Open source software to simulate electromagnetic wave propagation for Ground Penetrating Radar. Computer Physics Communications, 209(??):163–170, December 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465516302533.

Warren:2019:CBG

[WGG^{+}19] Craig Warren, Antonios Giannopoulos, Alan Gray, Iraklis Giannakis, Alan Patterson, Laura Wetter, and Andre Hamrah. A CUDA-based GPU engine for gprMax: Open source FDTD electromagnetic simulation software. Computer Physics Communications, 237(??):208–218, April 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465518303990.

Wang:2007:MEO

[WGS07] Yi Wang, Defeng Guo, and Huihui Shi. Measuring the evolution of open source software systems with their communities. ACM SIGSOFT Software Engineering Notes, 32(6):7:1–7:??, November 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wheeler:2003:WOS

[Whe03] David A. Wheeler. Why open source software / free software (OSS/FS)? look at the numbers! World-Wide Web document., December 31, 2003. URL http://www.dwheeler. com/oss_fs_why.html. Also available in other formats, and in translations to other languages.

Wang:2015:BVS

[WHJ15] Yu Wang, Emily Helminen, and Jingfeng Jiang. Building a virtual simulation platform for quasistatic breast ultrasound elastography using open source software: a preliminary investigation. Medical Physics, 42(9):5453–5466, 2015. CODEN MPHYA6. ISSN 2473-4209.

Wiil:1991:IDE

[Wii91a] U. K. Wiil. Issues in the design of EHTS: a multiuser hypertext system for collaboration. In Milutinovic et al. [MSNS91], pages 629–639 (vol. 2). ISBN 0-8186-2420-5. LCCN ???? 4 vol.

Wiil:1991:UES

[Wii91b] U. K. Wiil. Using events as support for data sharing in collaborative work. Informatik, Informationen Reporte, 1991. ISSN 0233-2582.

Wilkinson:1971:SCN

[Wil71] J. H. Wilkinson. Some comments from a numerical analyst. Journal of the ACM, 18(2):137–147, April 1971. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilson:1999:SOS

[Wil99] Greg Wilson. Soapbox: Is the open-source community setting a bad example? IEEE Software, 16(1):23–25, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print),

0740-7459 (electronic). URL http://dlib.computer.org/
so/books/so1999/pdf/s1023.pdf.

Williams:2002:FFR

[Wil02] Sam Williams. Free as in Freedom: Richard Stallman’s Crusade for Free Software. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00287-4. xii + 225 pp. LCCN QA76.76.C73 W55

2002. US$22.95. URL http://www.oreilly.com/catalog/
freedom.

Wilson:TB34-2-132

[Wil13] Peter Wilson. Glisterings: a font of fleurons; Fonts, /Linux, and ; Mixing traditional and system fonts. TUGboat, 34(2): 132–135, 2013. ISSN 0896-3207. URL https://tug.org/

TUGboat/tb34-2/tb107glister.pdf.

WilsonGarcia:2014:GLD

[Wil14] Edscott Wilson García. GNU libdbh (disk-based hash tables). SourceForge software archive., April 9, 2014.

URL http://sourceforge.net/projects/dbh;   http://
www.gnu.org/software/libdbh/.

Winter:1995:UAL

[Win95] R. Winter. Understanding assembly language listings from C compilers. Embedded Systems Programming, 8(11):94–96, 98, 101, 103–113, November 1995. CODEN EYPRE4. ISSN 1040-3272.

Withers:1990:CG

[Wit90] Robert Withers. CUG333 gAWK. C Users Journal, 8(12): 126–??, December 1990. ISSN 0898-9788.

Waleed:2022:WOS

[WJM22] Abdul Waleed, Abdul Fareed Jamali, and Ammar Masood. Which open-source IDS? Snort, Suricata or Zeek. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128622002420.

Wilken:1993:EMA

[WK93] K. D. Wilken and T. Kong. Efficient memory access checking. In IEEE [IEE93], pages 566–575. ISBN 0-8186-3680-7. LCCN QA76.5.I58 1993.

Wu:2008:SGI

[WKA^{+}08] Y. Wu, D. Khullar, A. Apte, J. Alaly, J. Matthews, W. Bosch, K. Ulin, and J. Deasy. SU-GG-T-393: Improvements to the computational environment for radiotherapy research opensource software system. Medical Physics, 35(6Part15):2815, 2008. CODEN MPHYA6. ISSN 2473-4209.

Willmes:2014:BRD

[WKB14] Christian Willmes, Daniel Kürner, and Georg Bareth. Building research data management infrastructure using open source software. Transactions in GIS, 18(4):496–509, 2014. ISSN 1361-1682 (print), 1467-9671 (electronic).

Williams:1990:GIP

[WKC^{+}90] Thomas Williams, Colin Kelley, John Campbell, David Kotz, and Russell Lang. GNUPLOT—An Interactive Plotting Program. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, August 31, 1990. Available in several Internet archives, including the Free Software Foundation collection on prep. ai.mit.edu. GNUPLOT can produce output for many different devices, including L^{A}TEX picture mode, PostScript, and the X Window System. See also [Kot90].

Wang:2014:MOS

[WKS^{+}14] Xiaofeng Wang, Ilona Kuzmickaja, Klaas-Jan Stol, Pekka Abrahamsson, and Brian Fitzgerald. Microblogging in open source software development: The case of Drupal and Twitter. IEEE Software, 31(4):72–80, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wu:2001:OSS

[WL01] Ming-Wei Wu and Ying-Dar Lin. Open source software development: An overview. Computer, 34(6):33–38, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-

0814 (electronic). URL http://dlib.computer.org/co/
books/co2001/pdf/r6033.pdf;   http://www.computer.
org/computer/co2001/r6033abs.htm.

Witten:2001:DOS

[WLC01] Brian Witten, Carl Landwehr, and Michael Caloyannides. Does open source improve system security? IEEE Software, 18(5):57–61, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/so2001/s5057abs.htm;   http:
//dlib.computer.org/so/books/so2001/pdf/s5057.pdf.

Wallace:2017:OIO

[WLD^{+}17] Byron C. Wallace, Marc J. Lajeunesse, George Dietz, Issa J. Dahabreh, Thomas A. Trikalinos, Christopher H. Schmid, and Jessica Gurevitch. OpenMEE: Intuitive, open-source software for meta-analysis in ecology and evolutionary biology. Methods in Ecology and Evolution, 8(8):941–947, 2017. ISSN 2041-210X.

Wrightson:2001:MUY

[WM01] Katherine Wrightson and Joe Merlino. Mastering Unix: Your comprehensive guide to today’s most powerful OS. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2817-3. xlii + 897 pp. LCCN QA76.76.O63 W75 2001. Featured on the CD: desktop environments, including Gnome and KDE; servers, including Samba-ache, and Sendmail; utilities, including GNU Emacs, Perl, and NEdit.

Williams:2005:EPP

[WM05] Christopher J. Williams and Christine M. Moffitt. Estimation of pathogen prevalence in pooled samples using maximum likelihood methods and open-source software. Journal of Aquatic Animal Health, 17(4):386–391, 2005. ISSN 0899-7659 (print), 1548-8667 (electronic).

Wang:2019:OSS

[WM19] Jinyong Wang and Xiaoping Mi. Open source software reliability model with the decreasing trend of fault detection rate. The Computer Journal, 62(9):1301–1312, September 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://academic.oup.com/comjnl/
article/62/9/1301/5146169.

Wu:2017:ALI

[WMK^{+}17] Yuhao Wu, Yuki Manabe, Tetsuya Kanda, Daniel M. German, and Katsuro Inoue. Analysis of license inconsistency in large collections of open source projects. Empirical Software Engineering, 22(3):1194–1222, June 2017. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616

(electronic). URL http://link.springer.com/article/
10.1007/s10664-016-9487-8.

Wang:2022:TAM

[WMLM22] Yuqing Wang, Mika V. Mäntylä, Zihao Liu, and Jouni Markkula. Test automation maturity improves product quality — quantitative study of open source projects using continuous integration. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0164121222000280.

Wang:2015:SDB

[WN15] Yongge Wang and Tony Nicol. On statistical distance based testing of pseudo random sequences and experiments with PHP and Debian OpenSSL. Computers & Security, 53(??):44–64, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000693.

Woo:2021:OSM

[WNS^{+}21] Mino Woo, Robert T. Nishida, Mario A. Schriefl, Marc E. J. Stettler, and Adam M. Boies. Open-source modelling of aerosol dynamics and computational fluid dynamics: Nodal method for nucleation, coagulation, and surface growth. Computer Physics Communications, 261(??):Article 107765, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520303829.

Woehr:1994:GKG

[Woe94a] J. Woehr. Getting to know GNU. Embedded Systems Programming, 7(2):34–35, 37, 39–40, February 1994. CODEN EYPRE4. ISSN 1040-3272.

Woehr:1994:WG

[Woe94b] J. Woehr. What’s GNU? Embedded Systems Programming, 7 (1):70–72, 74, January 1994. CODEN EYPRE4. ISSN 1040-3272.

Wolfe:1998:BP

[Wol98] Rosalee Wolfe. Beyond paint. Computer Graphics, 32(3): 29–31, August 1998. CODEN CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). Discusses and compares GIMP (GNU Image Manipulation Program), Graphics Workshop, NIH Image/Scion Image, Paint Shop Pro, and Webfx.

Wolfe:2002:OSP

[Wol02] A. Wolfe. An open-source platform to tackle scientific computing. IEEE Spectrum, 39(6):22–24, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolfe:2003:TEP

[Wol03a] Alexander Wolfe. Toolkit: Eclipse: a platform becomes an open-source Woodstock. ACM Queue: Tomorrow’s Computing Today, 1(8):14–16, November 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2003:TGT

[Wol03b] Alexander Wolfe. Toolkit: GNU tools, still relevant? ACM Queue: Tomorrow’s Computing Today, 1(9):14, 16–17, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2003:GTS

[Wol04] Alexander Wolfe. GNU tools, still relevant? ACM Queue: Tomorrow’s Computing Today, 1(9):14, 16–17, December/ January 2003/2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wood:2001:BOS

[Woo01] John Wood. Building an open-source Solaris-compatible threads library. In USENIX [USE01b], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL http:

//www.usenix.org/publications/library/proceedings/
usenix01/freenix01/wood.html.

Wiebe:2004:OSS

[WP04] Andreas Wiebe and Felix Prändl. Open Source Software — Rechtliche Rahmenbedingungen nach Österreichischem Recht. (German) [Open-Source software — legal boundary conditions for Austrian law]. Österreichische Juristen-Zeitung, 17:628–637, September 2004. URL http://

recht.wu-wien.ac.at/INSTITUT/PR/informationsrecht/
Rechtsinformationen/Abteilung/OpenSource/ArtikelWiebeOEJZHeft1704.
pdf.

Wouters:2014:CFO

[WPAV14] Sebastian Wouters, Ward Poelmans, Paul W. Ayers, and Dimitri Van Neck. CheMPS2: a free open-source spinadapted implementation of the density matrix renormalization group for ab initio quantum chemistry. Computer Physics Communications, 185(6):1501–1514, June 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465514000496.

Wilkinson:1971:LA

[WR71] James H. Wilkinson and Christian Reinsch, editors. Linear Algebra, volume II of Handbook for Automatic Computation, Editors: F. L. Bauer, A. S. Householder, F. W. J. Olver, H. Rutishauser, K. Samelson and E. Stiefel. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05414-6, 3-540-05414-6. viii + 439 pp. LCCN QA251 .W67.

Wood:2017:BSS

[WRDP17] Sean U. N. Wood, Jean Rouat, Stéphane Dupont, and Gueorgui Pironkov. Blind speech separation and enhancement with GCC-NMF. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(4):745–755, ???? 2017. CODEN

???? ISSN 2329-9290. URL http://ieeexplore.ieee.org/
document/7828077/.

Wright:2000:BGG

[Wri00] Peter Wright. Beginning GTK+/GNOME Programming. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-381-1. xi + 613 pp. LCCN QA76.9.U83 W754 2000. US$39.99.

Wang:1992:EAT

[WRSG92] Weigang Wang, R. Rada, K. Strickland, and C. Ghaoui. An expertext authoring tool. Information and Decision Technologies, 18(2):101–114, 1992. CODEN IDTEEI. ISSN 0923-0408.

Woo:2022:OSM

[WSK^{+}22] Mino Woo, Mario A. Schriefl, Markus Knoll, Adam M. Boies, Marc E. J. Stettler, Simone Hochgreb, and Robert T. Nishida. Open-source modelling of aerosol dynamics and computational fluid dynamics: Bipolar and unipolar diffusion charging and photoelectric charging. Computer Physics Communications, 278(??):Article 108399, September 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001187.

Wuttke:2012:LFT

[Wut12] Joachim Wuttke. Laplace–Fourier transform of the stretched exponential function: Analytic error bounds, double exponential transform, and open-source implementation “libkww”. Algorithms (Basel), 5(4):604–628, December 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/5/4/604.

Wall:2004:DGG

[WvH04] Kurt Wall and William von Hagen. The Definitive Guide to GCC. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-109-7. xxviii + 519 pp. LCCN QA76.76.C65W36 2004.

Wang:2001:OSS

[WW01] Huaiqing Wang and Chen Wang. Open source software adoption: a status report. IEEE Software, 18 (2):90–95, March/April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/so2001/s2090abs.htm;   http:
//dlib.computer.org/so/books/so2001/pdf/s2090.pdf.

Wessel:2021:DDM

[WWSG21] Mairieli Wessel, Igor Wiese, Igor Steinmacher, and Marco Aurelio Gerosa. Don’t disturb me: Challenges of interacting with software bots on open source software projects. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):301:1–301:21, October 2021. CODEN ???? ISSN

2573-0142 (electronic). URL https://dl.acm.org/doi/10.
1145/3476042.

Welsh:1994:LBG

[WY94] Matt Welsh and Yggdrasil Computing Inc. The Linux bible: the GNU testament. Yggdrasil Computing, San Jose, CA, USA, second expanded edition, 1994. ISBN 1-883601-10-X (paperback). 1176 pp. LCCN ????

Wu:2018:WOS

[WZS^{+}18] QuanSheng Wu, ShengNan Zhang, Hai-Feng Song, Matthias Troyer, and Alexey A. Soluyanov. WannierTools: an opensource software package for novel topological materials. Computer Physics Communications, 224(??):405–416, March 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0010465517303442.

Xiong:2014:BOS

[XAPK14] Qingang Xiong, Soroush Aramideh, Alberto Passalacqua, and Song-Charng Kong. BIOTC: an open-source CFD code for simulating biomass fast pyrolysis. Computer Physics Communications, 185(6):1739–1746, June 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465514000459.

Xia:2022:PHI

[XFS^{+}22] Tianpei Xia, Wei Fu, Rui Shu, Rishabh Agrawal, and Tim Menzies. Predicting health indicators for open source projects (using hyperparameter optimization). Empirical Software Engineering, 27(6):??, November 2022. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10664-022-10171-0.

Xu:2023:LLI

[XGF^{+}23] Sihan Xu, Ya Gao, Lingling Fan, Zheli Liu, Yang Liu, and Hua Ji. LiDetector: License incompatibility detection for open source software. ACM Transactions on Software Engineering and Methodology, 32(1):22:1–22:??, January 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/doi/10.1145/3518994.

Xiao:2008:GEI

[Xia08] Nong Xiao. Guest Editor’s introduction: Best papers from the GCC 2006 Conference. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(2):105– 106, ???? 2008. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http://www.informaworld.

com/smpp/content~content=a791170372.

Xin:2021:ITC

[XMGM21] Reynold Xin, Wes McKinney, Alan Gates, and Chris Mc-Cubbin. It takes a community: The open-source challenge. ACM Queue: Tomorrow’s Computing Today, 19(5):115–136, September 2021. CODEN AQCUAE. ISSN 1542-7730 (print),

1542-7749 (electronic). URL https://dl.acm.org/doi/10.
1145/3494834.3498568.

Xin:2022:ITC

[XMGM22] Reynold Xin, Wes McKinney, Alan Gates, and Chris McCubbin. It takes a community — : the open source challenge. Communications of the ACM, 65(5):48–55, May 2022. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3498568.

Xu:2022:OIC

[XOTI22] RuQing G. Xu, Tsuyoshi Okubo, Synge Todo, and Masatoshi Imada. Optimized implementation for calculation and fastupdate of Pfaffians installed to the open-source fermionic variational solver mVMC. Computer Physics Communications, 277(??):Article 108375, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522000947.

Xia:2011:MMD

[XTG^{+}11] Bing Xia, Zheng-Fu Tai, Yu-Cheng Gu, Bang-Jing Li, Li-Sheng Ding, and Yan Zhou. MyMolDB: a micromolecular database solution with open source and free components. Journal of Computational Chemistry, 32(13):2942–2948, October 2011. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xu:2022:DCV

[XTY^{+}22] Rongze Xu, Zhanyong Tang, Guixin Ye, Huanting Wang, Xin Ke, Dingyi Fang, and Zheng Wang. Detecting code vulnerabilities by learning from large-scale open source repositories. Journal of Information Security and Applications (JISA), 69(??):??, September 2022. CODEN ???? ISSN

2214-2126. URL http://www.sciencedirect.com/science/
article/pii/S221421262200148X.

Xia:2023:LLA

[XWZ^{+}23] Xiaoya Xia, Wei Wang, Shengyu Zhao, Sikang Bian, and Rong Wang. Lessons learned from the Ant Group Open Source Program Office. Computer, 56(4):92–97, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xu:2021:DAO

[XXAD21] Fei Xu, Qingang Xiong, Vadym Aizinger, and Guillaume Ducrozet. Development and application of open-source software for problems with numerical PDEs. Computers and Mathematics with Applications, 81(??):1–2, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122120304557.

Xue:2019:ASC

[XXCL19] Y. Xue, Z. Xu, M. Chandramohan, and Y. Liu. Accurate and scalable cross-architecture cross-OS binary code search with emulation. IEEE Transactions on Software Engineering, 45(11):1125–1149, November 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Yoo:2005:OSS

[YA05] Terry S. Yoo and Michael J. Ackerman. Open source software for medical image processing and visualization. Communications of the ACM, 48(2):55–59, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yildirim:2011:FFF

[YA11] Nihan Yildirim and Hacer Ansal. Foresighting FLOSS (free/ libre/open source software) from a developing country perspective: the case of Turkey. Technovation, 31(12):666–678, December 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001052.

Yacko:1988:GEB

[Yac88] Nancy Jane Yacko. A GNU Emacs browser for C++ code development. Thesis (m.s.), University of Illinois at Urbana-Champaign, Urbana-Champaign, IL 61801, USA, 1988. v + 33 pp.

Yadava:2007:BDB

[Yad07] Himanshu Yadava. The Berkeley DB book: the definite guide to creating scalable and fault-tolerant, embedded database applications using the open source Berkeley DB library. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-672-2. xx + 442 pp. LCCN QA76.9.D3 Y33 2007.

Yang:1990:ERP

[Yan90] Y. Yang. Experimental rapid prototype of undo support. Information and Software Technology, 32(9):625–635, November 1990. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Yang:1992:ADU

[Yan92] Y. Yang. Anatomy of the design of an undo support facility. International Journal of Man-Machine Studies, 36(1):81–95, January 1992. CODEN IJMMBC. ISSN 0020-7373.

Yap:2011:PDO

[Yap11] Chun Wei Yap. PaDEL-descriptor: An open source software to calculate molecular descriptors and fingerprints. Journal of Computational Chemistry, 32(7):1466–1474, 2011. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yasumura:1991:CSC

[YAS91] M. Yasumura, M. Arisawa, and N. Saito. A case study of computer literacy education. Joho-Shori (J. Information Processing Soc. Japan), 32(12):1310–1317, 1991. CODEN JOSHA4. ISSN 0447-8053.

Yang:2019:EOS

[YDZ19] J. Yang, R. K. Das, and N. Zhou. Extraction of octave spectra information for spoofing attack detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12): 2373–2384, December 2019. ISSN 2329-9304.

Yeoh:2005:BAL

[Yeo05] C. Yeoh. Building applications for the Linux Standard Base. IBM Systems Journal, 44(2):369–??, ???? 2005. CODEN IBMSA7. ISSN 0018-8670.

Yesylevskyy:2012:SNU

[Yes12] Semen O. Yesylevskyy. Software news and updates: Pteros: Fast and easy to use open-source C++ library for molecular analysis. Journal of Computational Chemistry, 33(19):1632– 1636, July 15, 2012. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yggdrasil:1993:YLG

[Ygg93] Yggdrasil Computing, Inc. Yggdrasil Linux/GNU/X operating system, 1993. ISSN 1069-3955. 1 computer laser optical disk.

Yggdrasil:1994:YLG

[Ygg94] Yggdrasil Linux/GNU/X operating system, 1994. ISSN 1069-3955. Yggdrasil Computing, Inc., Berkeley, CA, USA. CD-ROM. Updated quarterly.

You:2023:DOS

[YKK23] Hojun You, Juhyun Kim, and Chongam Kim. Deneb: an open-source high-performance multi-physical flow solver based on high-order DRM-DG method. Computer Physics Communications, 286(??):Article 108672, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000176.

Yoon:2020:JOS

[YKSH20] Hongkee Yoon, Taek Jung Kim, Jae-Hoon Sim, and Myung Joon Han. Jx: an open-source software for calculating magnetic interactions based on magnetic force theory. Computer Physics Communications, 247(??):Article 106927, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519302991.

Yalta:2008:GLP

[YL08] A. Talha Yalta and Riccardo Lucchetti. The GNU/Linux platform and freedom respecting software for economists. Journal of Applied Econometrics, 23(2):279–286, 2008. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yan:2005:UOS

[YLG05] N. Yan, D. Leip, and K. Gupta. The use of open-source software in the IBM corporate portal. IBM Systems Journal, 44 (2):419–??, ???? 2005. CODEN IBMSA7. ISSN 0018-8670.

Yang:2021:DIS

[YLHW21] Xuan Yang, Xiao Li, Daning Hu, and Harry Jiannan Wang. Differential impacts of social influence on initial and sustained participation in open source software projects. Journal of the Association for Information Science and Technology, 72(9): 1133–1147, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yang:2007:ERM

[YLL^{+}07] Byung-Sun Yang, Junpyo Lee, SeungIl Lee, Seongbae Park, Yoo C. Chung, Suhyun Kim, Kemal Ebcioglu, Erik Altman, and Soo-Mook Moon. Efficient register mapping and allocation in LaTTe, an open-source Java just-in-time compiler. IEEE Transactions on Parallel and Distributed Systems, 18 (1):57–69, January 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2016:MAR

[YLXZ16] Jianfeng Yang, Yu Liu, Min Xie, and Ming Zhao. Modeling and analysis of reliability of multi-release open source software incorporating both fault detection and correction processes. The Journal of Systems and Software, 115(??):102– 110, May 2016. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121216000200.

Yamamura:1993:AFM

[YM93] K. Yamamura and T. Matsuura. Adapting the Fujitsu’s mainframe operating system MSP to the open computing environment. In Abe and Yuasa [AY93], pages 40–45.

Yoon:2023:PFO

[YMCF23] Tae Jun Yoon, Katie A. Maerzke, Robert P. Currier, and Alp T. Findikoglu. PyOECP: a flexible open-source software library for estimating and modeling the complex permittivity based on the open-ended coaxial probe (OECP) technique. Computer Physics Communications, 282(??):Article 108517, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002363.

Yangui:2014:COS

[YMLT14] Sami Yangui, Iain-James Marshall, Jean-Pierre Laisne, and Samir Tata. CompatibleOne: The open source cloud broker. Journal of Grid Computing, 12(1):93–109, March 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://link.springer.com/article/
10.1007/s10723-013-9285-0.

Younker:2008:FAP

[You08] Jeff Younker. Foundations of agile Python development: [Python, agile project methods, and a comprehensive open source tool chain!]. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2008. ISBN 1-59059-981-0. xxi + 393 pp. LCCN QA76.73.P98.

Yu:2006:MKO

[YSC^{+}06] Liguo Yu, Stephen R. Schach, Kai Chen, Gillian Z. Heller, and Jeff Offutt. Maintainability of the kernels of open-source operating systems: a comparison of Linux with FreeBSD, NetBSD, and OpenBSD. The Journal of Systems and Software, 79 (6):807–815, June 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Young-S:2017:OGI

[YSMA^{+}17] Luis E. Young-S., Paulsamy Muruganandam, Sadhan K. Adhikari, Vladimir Loncar, Dusan Vudragović, and Antun Balaz. OpenMP GNU and Intel Fortran programs for solving the time-dependent Gross–Pitaevskii equation. Computer Physics Communications, 220(??):503–506, November 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465517302321.

Young-S:2016:OFP

[YSVM^{+}16] Luis E. Young-S., Dusan Vudragović, Paulsamy Muruganandam, Sadhan K. Adhikari, and Antun Balaz. OpenMP Fortran and C programs for solving the time-dependent Gross–Pitaevskii equation in an anisotropic trap. Computer Physics Communications, 204(??):209–213, July 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551630073X.

Yilmaz:2022:QEM

[YT22] Nebi Yilmaz and Ayça Kolukisa Tarhan. Quality evaluation models or frameworks for open source software: a systematic literature review. Journal of Software: Evolution and Process, 34(6):e2458:1–e2458:??, June 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Yu:2006:IPM

[Yu06] Liguo Yu. Indirectly predicting the maintenance effort of open-source software. Journal of Software Maintenance and Evolution: Research and Practice, 18(5):311–332, September 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Yuksel:1994:MEC

[Yuk94] O. Yuksel, editor. 7th Mediterranean Electrotechnical Conference: proceedings, April 12–14, 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1772-6. LCCN TK5101.A1M38 1994. Three volumes. IEEE Catalog No. 94CH3388-6.

Yang:2007:IWS

[YWA07] X. Yang, X. D. Wang, and R. Allan. Investigation of WSRP support in selected open-source portal frameworks. Concurrency and Computation: Practice and Experience, 19(12): 1729–1738, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2019:ACL

[YXS^{+}19] M. Yan, X. Xia, E. Shihab, D. Lo, J. Yin, and X. Yang. Automating change-level self-admitted technical debt determination. IEEE Transactions on Software Engineering, 45(12): 1211–1229, December 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Yi:2015:ESF

[YYL^{+}15] Qiuping Yi, Zijiang Yang, Jian Liu, Chen Zhao, and Chao Wang. Explaining software failures by cascade fault localization. ACM Transactions on Design Automation of Electronic Systems., 20(3):41:1–41:??, June 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2022:EAE

[YZC22] Jianfeng Yang, Ming Zhao, and Jing Chen. ELS algorithm for estimating open source software reliability with masked data considering both fault detection and correction processes. Communications in Statistics: Theory and Methods, 51(19): 6792–6817, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1866610.

Zack:2001:DUG

[Zac01] Randy Zack. Debugging under GNU/Linux. C/C++ Users Journal, 19(2):??, February 2001. CODEN CCUJEX. ISSN 1075-2838.

Zanella:2023:TOS

[ZAC^{+}23] Alice Zanella, Luca Abergo, Francesco Caccia, Myles Morelli, and Alberto Guardone. Towards an open-source framework for fluid-structure interaction using SU2, MB-Dyn and preCICE. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0377042723001553.

Zadok:2002:OAC

[Zad02] Erez Zadok. Overhauling Amd for the ’00s: a case study of GNU Autotools. In USENIX [USE02b], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002. URL http:

//www.usenix.org/publications/library/proceedings/
usenix02/tech/freenix/zadok.html.

Zaghi:2014:OSF

[Zag14] S. Zaghi. OFF, Open source Finite volume Fluid dynamics code: a free, high-order solver based on parallel, modular, object-oriented Fortran API. Computer Physics Communications, 185(7):2151–2194, July 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465514001283.

Zhou:1995:OSS

[ZC95] J. J. Zhou and Moon-Jung Chung. Object-oriented simulation for the superconducting super collider. Transactions of the Society for Computer Simulation, 12(1):1–25, March 1995. CODEN TSCSEV. ISSN 0740-6797.

Zendra:2001:CAG

[ZC01] Olivier Zendra and Dominique Colnet. Coping with aliasing in the GNU Eiffel Compiler implementation. Software—Practice and Experience, 31(6):601–613, May 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhu:2017:DOS

[ZCG17] Lianhua Zhu, Songze Chen, and Zhaoli Guo. dugksFoam: an open source OpenFOAM solver for the Boltzmann model equation. Computer Physics Communications, 213(??):155– 164, April 2017. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465516303642.

Zhou:2005:OSS

[ZD05] Ying Zhou and Joseph Davis. Open source software reliability model: an empirical approach. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhao:2010:EIU

[ZDM10] Luyin Zhao, Fadi P. Deek, and James A. McHugh. Exploratory inspection — a user-based learning method for improving open source software usability. Journal of Software Maintenance and Evolution: Research and Practice, 22(8): 653–675, December 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zhao:2000:SQR

[ZE00] Luyin Zhao and Sebastian Elbaum. A survey on quality related activities in open source. ACM SIGSOFT Software Engineering Notes, 25(3):54–57, May 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhao:2003:QAU

[ZE03] Luyin Zhao and Sebastian Elbaum. Quality assurance under the open source development model. The Journal of Systems and Software, 66(1):65–75, April 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zeitlyn:2003:GED

[Zei03] David Zeitlyn. Gift economies in the development of open source software: anthropological reflections. Research Policy, 32(7):1287–1291, 2003. CODEN ???? ISSN 0048-7333 (print),

1873-7625 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0048733303000532.

Zampetti:2021:SAT

[ZFD21] Fiorella Zampetti, Gianmarco Fucci, and Massimiliano Di Penta. Self-admitted technical debt practices: a comparison between industry and open-source. Empirical Software Engineering, 26(6):??, November 2021. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10664-021-10031-3.

Zheng:2019:TUB

[ZFY^{+}19] Wei Zheng, Chen Feng, Tingting Yu, Xibing Yang, and Xiaoxue Wu. Towards understanding bugs in an open source cloud management stack: an empirical study of OpenStack software bugs. The Journal of Systems and Software, 151 (??):210–223, May 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0164121219300354.

Zhang:2016:TDH

[Zha16] D. Zhang. TH-E-209-03: Development of an in-house CT dose monitoring and management system based on open-source software resources — pearls and pitfalls. Medical Physics, 43(6Part47):3902, 2016. CODEN MPHYA6. ISSN 2473-4209.

Ziccardi:2001:DAN

[Zic01] Giovanni Ziccardi. Il diritto d’autore nell’era digitale: evoluzione tecnologica e copyright: Internet, mp3, DivX;-), open source, Gnu/Linux, free software, mezzi di protezione. (Italian) [Copyright in the digital age: evolution, technology, and copyright: Internet, mp3, DivX;-), open source, Gnu/ Linux, free software, means of protection]. Il sole 24 ore, Milan, Italy, 2001. ISBN 88-324-4459-3. xxi + 343 pp. LCCN KE1809 .Z53 2001.

Zimmermann:2010:RCG

[Zim10] Paul Zimmermann. Reliable computing with GNU MPFR. Lecture Notes in Computer Science, 6327:42–45, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/978-3-642-15582-6_8.pdf.

Zheng:2020:SOS

[ZJS^{+}20] Size Zheng, Leili Javidpour, Muhammad Sahimi, Katherine S. Shing, and Aiichiro Nakano. sDMD: an open source program for discontinuous molecular dynamics simulation of protein folding and aggregation. Computer Physics Communications, 247(??):Article 106873, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302607.

Zeller:2005:EOS

[ZK05] Andreas Zeller and Jens Krinke. Essential open source toolset: programming with Eclipse, JUnit, CVS, Bugzilla, Ant, Tcl/ Tk and more. Wiley, New York, NY, USA, 2005. ISBN 0-470-84445-0 (paperback). xii + 392 pp. LCCN QA76.76.D47

Z45 2005. URL http://www.loc.gov/catdir/toc/ecip053/
2004026271.html.

Zimmerman:2021:MFE

[ZK21] Alexander G. Zimmerman and Julia Kowalski. Mixed finite elements for convection-coupled phase-change in enthalpy form: Open software verified and applied to 2D benchmarks. Computers and Mathematics with Applications, 84(??):77–96, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0898122120304338.

Zien:1991:XXF

[ZKCS91] Jason Zien, Jackson Kong, Pak K. Chan, and Martine Schlag. XS-XILINX 2000/3000 FPGA simulator. ucsc-crl-91-42, University of California, Santa Cruz, Santa Cruz, CA, USA, Oc-

tober 1991. 39 pp. URL ftp://ftp.cse.ucsc.edu/pub/tr/
ucsc-crl-91-42.ps.Z;   mailto::rnalib@ftp.cs.ucsc.

edu. prize (\ $5.00).

Zampetti:2022:ECS

[ZKDP22] Fiorella Zampetti, Ritu Kapur, Massimiliano Di Penta, and Sebastiano Panichella. An empirical characterization of software bugs in open-source Cyber–Physical Systems. The Journal of Systems and Software, 192(??):??, October 2022. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121222001315.

Zhao:2022:OSF

[ZLF^{+}22] Han Zhao, Xiangbei Liu, Andrew H. Fletcher, Ru Xiang, John T. Hwang, and David Kamensky. An open-source framework for coupling non-matching isogeometric shells with application to aerospace structures. Computers and Mathematics with Applications, 111(??):109–123, April 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122122000657.

Zitser:2004:TSA

[ZLL04] Misha Zitser, Richard Lippmann, and Tim Leek. Testing static analysis tools using exploitable buffer overflows from open source code. ACM SIGSOFT Software Engineering Notes, 29(6):97–106, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhao:2015:JEB

[ZPH^{+}15] Yanxiao Zhao, Jems Pradhan, Jun Huang, Yu Luo, and Lina Pu. Joint energy-and-bandwidth spectrum sensing with GNU radio and USRP. ACM SIGAPP Applied Computing Review, 14(4):40–49, January 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/2724928.2724932.

Zhou:2021:ESO

[ZRGJ21] Zhide Zhou, Zhilei Ren, Guojun Gao, and He Jiang. An empirical study of optimization bugs in GCC and LLVM. The Journal of Systems and Software, 174(??):??, April 2021. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121220302740.

Zhang:2020:BNB

[ZRNA20] Lixia Zhang, Leonardo O. Rodrigues, Niven R. Narain, and Viatcheslav R. Akmaev. bAIcis: a novel Bayesian network structural learning algorithm and its comprehensive performance evaluation against open-source software. Journal of Computational Biology, 27(5):698–708, May 2020. CO-DEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.liebertpub.com/doi/abs/10.
1089/cmb.2019.0210;   https://www.liebertpub.com/doi/
pdf/10.1089/cmb.2019.0210.

Zhang:2021:SOS

[ZRZ^{+}21] Chi Zhang, Massoud Rezavand, Yujie Zhu, Yongchuan Yu, Dong Wu, Wenbin Zhang, Jianhang Wang, and Xiangyu Hu. SPHinXsys: an open-source multi-physics and multiresolution library based on smoothed particle hydrodynamics. Computer Physics Communications, 267(??):Article 108066, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001788.

Zhang:2014:COS

[ZSW14] Tao Zhang, Wei Shu, and Min-You Wu. CUIRRE: an opensource library for load balancing and characterizing irregular applications on GPUs. Journal of Parallel and Distributed Computing, 74(10):2951–2966, October 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001269.

Zaidman:2011:SCE

[ZVvDD11] Andy Zaidman, Bart Van Rompaey, Arie van Deursen, and Serge Demeyer. Studying the co-evolution of production and test code in open source and industrial developer test processes through repository mining. Empirical Software Engineering, 16(3):325–364, June 2011. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10664-010-9143-7.pdf.

Zhao:2017:IEO

[ZW17] Rongying Zhao and Mingkun Wei. Impact evaluation of open source software: an Altmetrics perspective. Scientometrics, 110(2):1017–1033, February 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2204-y.

Zhou:2021:SBH

[ZWH21] Jiayuan Zhou, Shaowei Wang, and Ahmed E. Hassan. Studying backers and hunters in bounty issue addressing process of open source projects. Empirical Software Engineering, 26(4): ??, July 2021. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL https://link.springer.com/
article/10.1007/s10664-021-09979-z.

Zhou:2022:SDT

[ZWU22] Jiayuan Zhou, Shaowei Wang, and Naoyasu Ubayashi. Studying donations and their expenses in open source projects: a case study of GitHub projects collecting donations through open collectives. Empirical Software Engineering, 27(1):??, January 2022. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL https://link.springer.com/
article/10.1007/s10664-021-10060-y.

Zhao:2023:LSE

[ZXB^{+}23] Yutong Zhao, Lu Xiao, Andre B. Bondi, Bihuan Chen, and Yang Liu. A large-scale empirical study of real-life performance issues in open source projects. IEEE Transactions on Software Engineering, 49(2):924–946, February 2023. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Zhao:2022:QAI

[ZZZ22] Defu Zhao, Qunying Zou, and Milad Boshkani Zadeh. A QoS-aware IoT service placement mechanism in fog computing based on open-source development model. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https://link.

springer.com/article/10.1007/s10723-022-09604-3.